A Complete Bibliography of Publications in the Journal of Cell Biology: 1990–1994

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
             WWW URL: http://www.math.utah.edu/~beebe/

04 June 2019 Version 1.00

Title word cross-reference

1 [CPU^{+}90, CKD90, DG90, GDKP92, ILH^{+}90, KYLV90, LKD^{+}94, MHR94, NVR92, SSCD90, YM90b, ZDO^{+}91]. 100 [WB92a]. 15 [MGD^{+}92]. 17 [MMY^{+}92, WLH92]. 175 [OKH92]. 18 [GBOB90]. 180 [SM93a]. 2 [ATD^{+}90, LM91, LLK91, LKD^{+}94, PT92, WGP^{+}93]. 3 [LSRC93, LGSB93, MH91, ROBN^{+}92]. 4 [CKD90, KWW^{+}94]. 43 [JFA^{+}91]. 440 [CKB93]. 5 [BCG^{+}92, KVE92, KYBC94, RYKA90]. 6 [EG91, GKY^{+}94, GK93a, GK92a, GI93, HWvF90, JK92, KWK92, KHV^{+}93, MH93b, RR92a, SGBN^{+}93]. 67 [HR94b]. 85 [WP93]. 92 [LWF^{+}91]. ^{(2+)} [GeKdV^{+}91]. ^{2+} [BAP^{+}90, BGB94, BZB^{+}93, CK91, CF94, DVT92, EP94, FMR^{+}92, GPF94, GZC^{+}91, GDKP92, HFN92, HBP^{+}93, JTS^{+}91, JFA^{+}91, JSA^{+}93, KISY91, KATS90, KSFG91, KC90, MC91b, MOKF92, PNM^{+}94, RBB^{+}94, TWMS90, TSBS92, VPC^{+}91, VPP^{+}93, YC93, vKWHF94]. a [KDT93]. → [BS94c, RPPB94].

-active [HSUS90]. -alpha [KYLV90]. -amino [MGD^{+}92]. -anchored [KGA^{+}93]. -ATPase [KKFD90, MNME^{+}93, VKS^{+}90]. -ATPase-defective [MR94]. -aza-2’-deoxycytidine [BCG^{+}92]. -beta-glucan [RPPB94]. -binding [LS90, WHK^{+}93]. -biphosphate [DVT92]. -bp [MMY^{+}92]. -Ca [FMR^{+}92]. -class [NVR92]. -day [GBOB90]. -deficient [MCEG90]. -Deoxyglucose [PT92]. -dependent [LGSB93, OHP^{+}94, BGK^{+}94, HDS90, SBL^{+}91, vKWH^{+}91]. -directed [KYBC94]. -encoded [KWK92, RR92a]. -expressing [GKY^{+}94]. -factor [KDT93]. -growth-inhibited [SSCD90]. -HETE [TCDH93]. -hydroxy- [ROBN^{+}92]. -independent [CPU^{+}90, MSQ^{+}90, LBF92, GeKdV^{+}91]. -induced [YM90b]. -integrins [LKD^{+}94]. -kD [CKB93, HR94b, LWF^{+}91, SM93a, WB92a, ST94]. -kilodalton [OKH92, WP93]. -linked [CKD90]. -lipoxygenase [KVE92]. -macroglobulin [ATD^{+}90]. -mediated [GDKP92, JFA^{+}91, KWW^{+}94, LSRC93, MH91, PMHS92]. -methylglutaryl [ROBN^{+}92]. -methyltetrahydrofolate [RYKA90]. -microglobulin [LLK91]. -mobilizing [TTR^{+}90]. -mRNA-induced [PSN^{+}91]. -negative [MWLC^{+}94]. -nm [LM91, WGP^{+}93]. -phosphatase [EG91]. -phosphate [GK92a, GI93, JK92, KHV^{+}93, MH93b, SGBN^{+}93, ZDO^{+}91]. -phosphate-independent [GK93a]. -phosphate-specific [HWvF90]. -phosphate/insulin-like [JK92]. -regulated [FF93, SLH92]. -related [DG90, MHR94]. -release [JSA^{+}93]. -residue [WLH92]. -stimulated [TLWA94]. -subunit [ILH^{+}90]. -trisphosphate [SRV^{+}90]. -type [TN94].

/calmodulin [WLWL94, BAP^{+}90, KISY91]. /calmodulin- [WLWL94]. /calmodulin-dependent [BAP^{+}90, KISY91]. /Cl [PA91]. /Cl-cotransport [PA91]. /K [PA91].

0 [DJA^{+}90]. 0-2A [DJA^{+}90]. 000 [FF93, ZFF92]. 000-D [FF93, ZFF92]. 04 [MJM91].

1 [AMBN91, ASL^{+}94, BBZ94, BBK90, BRWB91, BF90, BLSR92, BCB93, BMB^{+}92, CdPA^{+}93, CSCSA94, CDS^{+}91, CYCA90, CPM^{+}91, DCY^{+}94, DMM^{+}91, DSdF^{+}90, DGAB^{+}93, DS93, DSR^{+}93, DCCH92, ES92, FIH^{+}94, GAHB94, HNP94, IRH^{+}91, INY^{+}93, JG94, JS90, KSST91, LBH93, LMH^{+}94, LFG^{+}90, LGI^{+}93, LRSB94, LCLG94, MH93a, MGD^{+}92, MGS90, MSD92b, NNHB^{+}92, NS93, NHA^{+}92, PLB^{+}92, PKG^{+}94, RM94, RMFA92, RPS^{+}92, SAF90, SNG^{+}91, TSH^{+}93, TOK93, VTSS94, WKD^{+}91, YOWM91, ZJA^{+}94, ZSM^{+}93, ZB94, vKWH^{+}91]. 1- [LSRC93]. 1/beta [DvHB^{+}91, HRC^{+}90]. 1/BMP [GKY^{+}94]. 1/ICAM [CdPA^{+}93, CPU^{+}90]. 10 [BIPG91]. 1080 [CML90]. 110 [DLSB^{+}90, DLSK91]. 110-kD [DLSB^{+}90, DLSK91]. 110-kD-calmodulin [CSM90]. 115 [MK93, WCR92]. 115-kD [WCR92]. 12 [COLL^{+}90, TCDH93]. 120 [BGI^{+}90, CCW^{+}92, PC91a]. 120-kD [BGI^{+}90, PC91a]. 13 [OD93]. 13-A [OD93]. 14 [NHK^{+}94]. 14-kD [NHK^{+}94]. 140 [LPM^{+}91, MVM91, PZP^{+}91, ZMS^{+}91]. 15th [KWFM91].

160/180 [FWBS93]. 175K [SOH^{+}90]. 180 [CG91, FWBS93]. 180-kD [CG91]. 19.5S [BN91, BMGN92]. 193 [ISN^{+}94]. 19K [GK90]. 1A [NFL90]. 1B [BRA^{+}94]. 1C [AGGS92]. 1M [AUT^{+}91].

2 [BCG^{+}92, BCBK92, CH93, GLQRB91, HH90, JG94, KA93, LQNRB90, NDdC92, PDA93, PH92, Ree94, SJMK93, SBG93, ZSM^{+}93]. 2-3 [OKH92]. 20 [JKS93, KCB^{+}92, LCMP90, LML^{+}94, MK91a, MJM92]. 200 [NMP^{+}92]. 200-kD [NMP^{+}92]. 205K [IFLG90]. 21 [KDP^{+}92]. 21-kD [KDP^{+}92]. 22 [DJ93b]. 22-kD [DJ93b]. 22/SR13 [SSWS92]. 220 [IYN^{+}91, INY^{+}93]. 220-kD [IYN^{+}91, INY^{+}93]. 228 [BBZ94]. 228-kD [BBZ94]. 22S [BHS94]. 23 [WC90a]. 25 [MVV^{+}91, WHLW90]. 25-dihydroxyvitamin [BPDM90]. 25-kD [MVV^{+}91, WHLW90]. 252a [UYA^{+}91]. 26 [RTHB90]. 27 [FF93, MIU^{+}92, SNFN91, YMW^{+}90]. 27-kD [MIU^{+}92, SNFN91, YMW^{+}90]. 275 [RBR90]. 275-kD [RBR90]. 280 [GYE^{+}90]. 29 [BHS94, LBL^{+}91]. 29-kD [BHS94]. 2A [DJA^{+}90, FBML90]. 2Aadult [WN92, WWN92]. 2Aperinatal [WN92]. 2B [PWY^{+}92]. 2C [UOKH93]. 2u [DYS^{+}90]. 2u-globulin [DYS^{+}90]. 2X [DAM^{+}93]. 2X-myosin [DAM^{+}93].

3 [KLST93, ACSM^{+}94, CdPA^{+}93, CSMdPSM94, CDS^{+}91, CLD^{+}91, DBC90, EUH91, LMH^{+}94, LCT^{+}93, PH92, YGM^{+}94]. 3’-untranslated [KLST93]. 30 [ZFF92]. 34 [FF93, KBG91, LM91]. 34-kD [KBG91, LM91]. 37 [CJO92]. 37-kD [CJO92]. 39 [HP92, RMB90]. 39-kD [HP92, RMB90]. 3B [LKXS93]. 3T3 [AB91, BS94c, CJC93, DCM^{+}93, FA91, FGSBZ93, GTP^{+}92, GHR91, HHR^{+}90, JS90, KATS90, LSAH90, MMD91, PT92, RPH^{+}92, WTRD92, ZBPV92]. 3T3-1 [JS90]. 3T3-L1 [CJC93, MMD91, RPH^{+}92].

4 [CPU^{+}90, FP91, JP94, MGB90, MH93a, MHYK92, OVB^{+}94, PTK^{+}93, VTSS94]. 4-dihydropyridine [YAJ91]. 4-dihydroxybenzoate [NCBS90]. 4-galactosyltransferase [ELS93, HS92, YHM^{+}94]. 4/bullous [CKG^{+}90]. 40 [HKS^{+}92, JAB^{+}91]. 41 [JCSH93]. 42 [EOL^{+}90]. 42-kD [EOL^{+}90]. 420 [ATD^{+}90]. 420-kD [ATD^{+}90]. 42Sp48 [CPM^{+}91]. 42Sp50 [DMM^{+}91]. 43 [AC93, CSH^{+}92, GB90, MBS91, PBCK93, PMM91, PNR^{+}93, SCM^{+}93, WC93, YH93, dCV90]. 43-kD [AC93, PMM91, PNR^{+}93, SCM^{+}93, YH93]. 440 [KOB91]. 440-kD [KOB91]. 466 [WBL^{+}91]. 466-kD [WBL^{+}91]. 467 [KT92]. 467-kD [KT92]. 47 [MGG93]. 47-kD [MGG93]. 4TM [YLWS94].

5 [ZZR94, MHYK92]. 5’-noncoding [ZZR94]. 5-trisphosphate [FNM^{+}92, GSV93, MM93b, YC93]. 5-trisphosphate-dependent [SJIC92]. 5-trisphosphate-induced [HN90]. 50 [BHHG90]. 50-kD [BHHG90]. 54 [BEGN91, HD91b, RWL^{+}90, ZBW93]. 54-kD [BEGN91, HD91b, RWL^{+}90, ZBW93]. 55 [WLPB92]. 55-kD [WLPB92]. 58 [BRO^{+}91a]. 58-kD [BRO^{+}91a]. 58K [CSP90]. 5A [RHB^{+}93, PH92]. 5b [PH92]. 5’nucleotidase [SMSH92].

6 [BZF^{+}92, BLB^{+}91, JS90, KIS^{+}94, WBWH93]. 6-kD [KIS^{+}94]. 6-sialyltransferase [BZF^{+}92, WLH92]. 6/beta [HRC^{+}90, SCJ^{+}91]. 60 [BS94b, BFTR93, WST^{+}90]. 60-70-kD [WST^{+}90]. 60-kD [BS94b, BFTR93]. 62 [JS92, MW92]. 62-kD [JS92, MW92]. 67 [MWH^{+}91, SDW^{+}93]. 67-kD [MWH^{+}91]. 68 [dSK90]. 68- [LPA^{+}93]. 68-kD [dSK90]. 69 [HFL94]. 69-kD [HFL94]. 6A/E [SWFB91].

7/alpha [RPS^{+}92]. 70 [IMK^{+}92, WWS^{+}94]. 70-kD [IMK^{+}92, WWS^{+}94]. 70K [KKBL92]. 70S [FPGS91]. 72 [CL92, LPA^{+}93, RM94]. 72-kD [CL92, LPA^{+}93, RM94]. 74 [PMPV92]. 74-kD [PMPV92]. 7H6 [ZSM^{+}93]. 7S [LPA^{+}93, YMO92].

8 [KHHR92]. 80/ [WP93]. 82 [TKHM91]. 82-kD [TKHM91]. 83 [KW92a]. 83-kD [KW92a]. 87 [GW94b]. 88 [DW91]. 88-kD [DW91].

9 [SPKV94]. 9-O-acetylation [SPKV94]. 92 [RSM^{+}94]. 92-kD [RSM^{+}94]. 9E3 [MGB90]. 9E3/CEF [MGB90]. 9E3/CEF-4 [MGB90].

A- [KMF^{+}91, VDS^{+}92]. A-associated [BDK92]. A-chain [JRCW^{+}90]. A-domain [RUH^{+}94]. A-sensitive [PRAK94]. A-treated [SPHvD91, IDT^{+}92]. a/b [CL91b]. a1 [TRV^{+}90]. A23187 [KVE92]. A23187-induced [KVE92]. A4 [NRS^{+}93a]. A431 [DPCP91, MCM93, SSJ^{+}94, SKK^{+}91]. AA [OTWH92]. AA4 [OSK^{+}92]. Aberrant [BEO90, BF91]. ability [CL94, FMLD92, LMB^{+}94b, SRHK94, SLWF91, TKP^{+}91, TG94]. Abl [VJB^{+}94]. able [BLSR92, CHBGL91, KBCM^{+}90]. Abnormal [DOW92, KS94, AWH^{+}94, MSSD93, STS91]. abnormalities [BF91]. abolished [KC90, LKS94]. ABP [CCW^{+}92, GYE^{+}90]. ABP-120 [CCW^{+}92]. ABP-280 [GYE^{+}90]. abrogates [LD92, PRB^{+}91]. absence [BPM93, CL93, CWBB93, DLMS91, GHZD90, KAF^{+}93, LSR^{+}90a, MBC90, MOT90, SNE90, SBB91b, SS90b, TG94]. absent [DMM^{+}91, MGB90]. abundance [LB92a, SDK^{+}93]. abundant [SGWE94]. Acanthamoeba [BK90, BBK92, DP92, KLBBK91, MAA^{+}94, MZP91, RKB^{+}90, SRP90, VAL^{+}93]. accelerate [DMWJ90]. acceptable [BSD^{+}92]. acceptor [SSRM90]. acceptors [MMC91]. access [BF90, LCPAM94]. accessible [HH90, IDT^{+}92, SPD^{+}90]. accompanied [OHF^{+}94]. accompanies [BTR92]. accompany [FMJ91]. accompanying [HFAGM94]. account [SWD94, SY90b, WDT^{+}92]. accounting [ATV^{+}93]. accumulate [TSP90]. accumulates [Ang91, BHC^{+}94, SABF^{+}90]. Accumulation [CLDV90, DI91, KSB^{+}93, OB93, WRDK92, WF90, BCYC94, CLG^{+}92, DE90, FM94a, FDT^{+}91, GHR91, JS90, KGK90, NPST94, ON94, PMKH91, TWC93, YM93, YC93]. accurate [WKMG92]. acetate [CSS^{+}94]. acetate-mediated [CSS^{+}94]. acetylated [BSW91]. acetylation [SBT93, SPKV94]. Acetylcholine [SF90b, BP93, BRO^{+}91a, BSM90, BM90, CH92, FGH92, Fro91, GFV^{+}91, GSM93, HMC91, HCS94, LC91, PRGC91, PBD93, PNR^{+}93, RGHC91, RB90, SCL91, SCM^{+}93, SBK^{+}91, TU90, TGR^{+}92, VBA90, Wal94, WHH^{+}93, YH93]. Acetylcholinesterase [Ang91, AHM94, DMB92b, MVTJ94, RR92b, Rot90]. acetylgalactosaminylphosphotransferase [BTSL91]. acetylglucosaminidase [MGDS93]. AChR [NR90, PC90, PMM91, WBS91b]. acid [ASR91, AGGS92, BG91a, BH94, BG91b, BR91, BMGC91, CSR^{+}92, DRJ^{+}91, EG91, FPR90, FRL93, GK90, GHR91, GPS92, HE93, JKLG94, KHS^{+}90, KGA^{+}93, LFC94, LFB94, Low92, ML91, MGD^{+}92, NRS^{+}93a, OVB^{+}94, PHSA93, PLB^{+}91, RHH92, RELS^{+}91, SRG^{+}94, SGW91a, SVB94, TSLR90, VTSS94, WBE91, WGTCY94, YYR92, YSK^{+}94, ZCP^{+}94, ZZR94]. acid-mediated [LFB94]. Acidic [FSY^{+}91, BT93, BSR^{+}90, CL94, FM94a, HPM^{+}94, SFM^{+}90, WSL91]. acidification [HSvD93a, MPVL92, RHR90, ST94, ZBPV92]. acids [MYT94, SPKV94, SDR^{+}94, SY91, VWS90, VFC^{+}91, WRP^{+}91]. acinar [JGK^{+}94, PMHS92]. Acquisition [CT91, BMPW90]. acrosomal [OD93, SAJ^{+}94]. across [CPD^{+}93, GFW92, HMB^{+}93, KHHR92, MF90, PCD^{+}92, RG92, WMB^{+}94]. act [HMS^{+}90, LCPAM94, NS90, SSK^{+}90]. ACT3 [CM94b]. Actin [AP92, DCTG90, GIAS93, HSB^{+}94, NPR^{+}94, SFC^{+}94, SS90b, TCP90, AGKC92, Apg91, AHW90, BTSL91, BMS92, BPTS94, Bea91, BB93, BS94c, BMS^{+}91, BGI^{+}90, CLZ91, CCFZ92, CKMC94, CW90a, CW90b, CBRW92, CFW93, CSCSA94, CPSS92, CMZ90, CCR93, CSSL93, CSG92, CWHH92, CGF^{+}93, CM93, CPB94, DPCP91, DGGG93, DP92, DCL^{+}92, DEH^{+}91, ENS91, EC93, FMR^{+}93, FB93, FF93, FZ93, FKLS92, FBML90, FW93, Fow90, FKH^{+}93, FKT93, FH90, GMNH93, GT94, GCMDP91, GYE^{+}90, GeKdV^{+}91, HE93, HD91a, Har92, HKC92, HG93b, HSG94, HHL94, HWD94, HNH91, HLPL90, HCYS90, HMT^{+}92, JBJH^{+}92, KLST93, KZS94, KTG90, LKSR94, LB92b, LSG92, LG93b, LJHC94, LNSSA93, LDS^{+}92, MWSP91, MZP91, Mat94, MWD94, MVCC93, MKB^{+}94, MA90, MVV^{+}91, MLM^{+}93, MPM^{+}94, ND91, OB93, OTT92, OD93, PSHK92, PB91]. actin [PMTB94, PMW92, SW90, SYO^{+}91, SGC^{+}94, SLG92, SLHG93, SAJ^{+}94, Sch94, SKT^{+}90, SHTL93, SL90a, SY90b, SM91d, TM92, TDT92, TDWT92, TSS91, TOS^{+}94, TWV94, VDTT91, WLWL94, WDWT90, WPW92a, WPW92b, WPBF94, WMG^{+}93, WOC90, WP93, YNST93, YKY92, ZFF92, ZDP92, dAWL^{+}90, dHvBeHVB92, vDS94, CSNZ92, VJB^{+}94]. actin- [WLWL94]. actin-based [TOS^{+}94, YNST93]. actin-binding [ENS91, GYE^{+}90, HKC92, HMT^{+}92, LJHC94, TWV94, WMG^{+}93, WP93, dHvBeHVB92]. actin-bundling [ZFF92]. actin-containing [BTSL91, CPSS92]. Actin-facilitated [AP92]. actin-gelling [SHTL93]. actin-modulating [SYO^{+}91]. actin-nucleating [SL90a]. Actin-related [SFC^{+}94, MVCC93, PMTB94, SGC^{+}94, Sch94]. actin-scruin [OD93]. actin/gelsolin [DPCP91]. acting [CSR^{+}92, DHS93]. actinin [CPSS92, CMB92, FKB^{+}90b, HKC92, MWSP91, Mat94, MWD94, MA90, OPB90, PB91, RFF92, WPW92b, WHK^{+}93]. actinin/alpha [LDS^{+}92]. Actinocoryne [FCF92]. actins [MAA^{+}94]. action [BMGC91, CL92, DLSB^{+}90, FMR90, GS94, HR90c, JSG^{+}92, LDNM92, LKD^{+}94, OTT92, PS94, SBR91, TTR^{+}90, WMS^{+}93]. actions [KDT^{+}91]. Activated [BPM93, HTHC93, LSC^{+}93, ATD^{+}90, AJPB90, BB93, CXS^{+}94, GSR^{+}93, KHvdL^{+}92, MC91b, NPR^{+}94, SLP91, SBG93, TFT94, VSVP93b, LKD^{+}94]. activates [ASE90, HG94]. activating [BLSR92, KFC^{+}94, KHHR92, LDT^{+}92, TSBN94, WW94a, ZMMP90]. Activation [HE93, LB92a, NHA^{+}92, PRB^{+}91, RSW^{+}91, SSG93, SADJGP94, SZB94a, TBL^{+}92, WK92, vKWH^{+}91, ABM^{+}93, ATSE90, BFLS94, BSP^{+}94, DVT92, DFC^{+}92, FAMR92, FAS^{+}93, FTT^{+}93, GZC^{+}91, HRT^{+}91, Har92, IKL^{+}94, JM92, JKLG94, KNR93, LBH93, LWK91, LHS^{+}92, LPM^{+}91, LGPM90, MSC92, OHF^{+}94, OSK^{+}92, RLWK93, RHB^{+}93, SWD92, STL^{+}90, SOA^{+}93, SCMU93, SMM90, SNG^{+}91, SSJ90, SWFB91, SR92e, TSBS92, WNTT94, WHN^{+}94]. Activation-dependent [NHA^{+}92, WK92]. activator [CML90, EGP^{+}91, EMC^{+}90, FAMR92, GSS^{+}91, LSP^{+}93, QvMWV^{+}91, SR90a, SPD^{+}90, SLP91, SWL90, YM90a]. activator-mediated [SR90a]. activators [SS91b, SFY^{+}94]. Active [KATS90, BGRW91, ET91, HSUS90, KM92a, KMSW91, LCC^{+}93, NRS^{+}93a, ODPL93, OPL94, SOY93, dAWS^{+}90]. actively [MMH^{+}92, SAVK94]. activin [BG94, SK91]. activities [BSR^{+}90, HWS90, HSW90, MKCE92, MSQ^{+}90, PvBeHB^{+}91, SK91, SFO^{+}91, TRD^{+}92]. activity [BG91a, BTRB93, BDK92, BMB^{+}92, CXS^{+}94, CSM90, CG91, CQYD94, DPW91, EOL^{+}90, FF93, FAMR92, FRL93, GMW^{+}94, GFWM92, GC90, HWLT94, HXM^{+}94, ITSR90, JPJF91, JRK^{+}94, JS90, KHW93, LML^{+}94, LEM93, Mah91, MSC92, PRC^{+}90, RTHB90, RGT^{+}93, RMFA92, RB90, SR90a, SST92a, SM92a, SL90a, SLS^{+}90, SDCD93, SVK^{+}92, SOKS91, SSM92b, TSBN94, TVG^{+}93, TWC94, UHFG92, VMS^{+}91, VJB^{+}94, WIF^{+}91, WRHW92, XMW^{+}93, YM90a, YKM^{+}92, YGM^{+}94, ZZP94]. activity-dependence [FRL93]. actomyosin [AP92]. actophorin [MWSP91, MZP91]. acts [CNSC90, CSSL93, EPNN91, RLB^{+}91]. acute [Mac92]. acyl [DKG92]. acylation [PS94, PGAR90]. adapted [LBL^{+}91]. adaptin [Rob90, Rob93]. adaptor [CJC93]. adaptors [HSvD93a, SBR93, SCS92]. addition [Hen90]. additional [AJL^{+}90, BBJ^{+}91]. addressin [BRWB91, MSB^{+}91]. adducin [JGO^{+}91]. adenomatous [HHH^{+}93]. adenovirus [GK90, WFCN94]. adenoviruses [KHHC93]. adenylate [BMV90, JPJF91]. adenylyl [IKL^{+}94, SSG93, ZS94]. Adherence [SDR^{+}90, JTS^{+}91, KCYH92, KHvdL^{+}92, RPS^{+}92]. Adherence-dependent [SDR^{+}90]. adherens [CJO92, MLRJ92, TOA^{+}91]. adherent [GAHB94, LNSSA93, MM90]. Adhesion [FJN93, vdWvKvKdB^{+}92, MIS93, NVR92, AAM92, AORB90, AMBN91, AUT^{+}91, ACSM^{+}94, AHS^{+}92, BTSL91, BRA^{+}94, BSSL^{+}91, BMTTG94, BKB^{+}90, BGK^{+}94, BLR^{+}92, BJ90, Blo92, BF93, BCB93, BDB90, BGM^{+}91, BTR92, CL91a, CPU^{+}90, CdPA^{+}93, CDS^{+}91, CPSS92, CWBK90, CLD^{+}91, CMS^{+}90a, CLK^{+}91, CMB91, CdJC93, DMB93, DCY^{+}94, DI91, DG90, DGAB^{+}93, DS93, DAS^{+}92, DKM^{+}92, DGGR92, EFD^{+}93, EHB^{+}90, EWP^{+}92, EWP^{+}93a, FT93, FvBuHWS92, FMK^{+}94, FBS^{+}91, GTP^{+}92, GAS^{+}90a, GMB^{+}91, GFM93, HJI^{+}93, HWLT94, HGDA93, HMS^{+}90, HMY92, HDS90, HSP93, HNP94, HWS90, ISF^{+}92, IRH^{+}91, IYN^{+}91, JZAVP93, KKAS90a, KKAS90b, KES^{+}92, KT92, KSL^{+}94, KRdlR^{+}92, KWW^{+}94, KA90, KYBC94, LCMP90, LWV92, LVB^{+}94, LSC^{+}93, LFB94, LPM^{+}91, LEM93, MBW^{+}94, MNME^{+}93, MH93a, MHT^{+}92, MMJ^{+}90, ML94, MKCE92, MSQ^{+}90, MRSN93, MFS^{+}94, MMI^{+}91, MULA^{+}91, MJ92]. adhesion [NIT94, NYA^{+}91, NGP^{+}91, NS93, NSWH90, NHH^{+}91, OVB^{+}94, PW93, PTMB94, PFSS92, RLB^{+}91, RWG^{+}92, RCJ^{+}94, RUH^{+}94, RM94, RLKB91, RA94, SWD92, SSG93, SEC^{+}93, SMM90, SNG^{+}91, SGN^{+}93, SPKV94, SLWF91, SCJ^{+}91, SEMH92, STo^{+}92, STC93, SM92b, TYM^{+}92, TSK^{+}94, TRLG90, TSS^{+}92, VR92, WNTT94, WS91a, WS91b, WHH^{+}93, YED^{+}93, ZS94, ZBH94, ZFA^{+}93, ZMMP90, ZDR^{+}92, vKWH^{+}91, vKWHF94]. adhesion-defective [SEC^{+}93]. Adhesion-dependent [FJN93, ZS94]. adhesions [CKG^{+}90, CML90, HSCR94, HSP93, TGB90, YCO^{+}93]. Adhesive [HLC^{+}93, KFW^{+}91, ALE^{+}92, BSEF^{+}93, BGLB94, CL91a, DR91, FTT^{+}93, OK90, PMG^{+}92, PAFC92, RH92b, SCS90, SSM92b, TRP^{+}93, WIF^{+}90, WGM^{+}93, WF94]. adipocytes [LVTB93, MDKB94, MMD91, RPH^{+}92]. adipocytic [CJC93]. adipogenesis [SLB^{+}94]. adipose [SGG^{+}91]. adjacent [HAG^{+}93, KBG91, KGA^{+}93, TAWJ91]. adjoining [NKW^{+}92]. ADP [EOO^{+}94, JvCH^{+}94, NHK^{+}94, SBR91, SSK^{+}90, TSH^{+}93]. ADP-ribosylation [SBR91, EOO^{+}94]. ADP/ATP [SSK^{+}90]. adrenal [SMJ^{+}94]. adrenergic [KCDR91]. adsorbed [MJ92]. Adult [EERM^{+}91, BML^{+}90, BOH^{+}92, BBB91, CG90, CKL^{+}91, DIYSS92, DMB93, KYLV90, LWB91, LC91, ME91, SBK^{+}91, SST92b, WMvdK94, YRBBP90]. aequorins [KRCT93]. AF [RLB^{+}91]. AF/R1 [RLB^{+}91]. affect [HWL^{+}90, KZS94, PMW92]. affected [HG93a, YYR92]. affecting [GW94a, MHGC90, SLK91, WG92b, WNM^{+}94]. affects [SSFD93, WYM^{+}92, YMO92, YHM^{+}94, vKWHF94]. afferent [MSB^{+}91]. Affinity [FKHG93, BTRB93, BMK^{+}90, CLP^{+}94, CE94, GT94, HHL94, MAA^{+}94, MWH^{+}91, SRB92, TBL^{+}92, VKR^{+}91, VPC^{+}91]. affinity-high [VPC^{+}91]. afflicted [DGF^{+}94]. aFGF [OR92]. African [HPM^{+}94, MG94a]. after [BO91, CCFZ92, ET94, FGS^{+}92, FFB^{+}93, FKT93, KHTD92, KBP91, KT93, LCK^{+}92, PB91, RBD92, SST92b, TLWA94, THL^{+}90, ZZP94, vGvMS^{+}91]. Ag [PMM90]. against [CYC91, ESL^{+}90, FL92, GKKS90, IMK^{+}92, JD91, WKY^{+}93]. agarose [MAA^{+}94]. agents [ZCH90]. agglutinins [HKS90]. aggregate [LC94]. aggregation [BCBK92, BKFT93, CH91, FWF90, IBD92, KBA92, MH92, NR90, SR92e, THBW94, TSH^{+}93, Wal94, YDS92]. aggressive [KDS^{+}91]. aging [PZV^{+}91]. agonist [GBS90, JWHPM93, KCDR91]. agonist-induced [GBS90]. AGp110 [SGM^{+}90a]. Agrin [HCS94, NR90, LC91, Wal94]. Agrin-induced [HCS94, NR90, Wal94]. AHNAK [SB93b]. AIDS [BBB^{+}94]. airway [DOW92]. alanine [NFYI93, PTD90]. alanine/glyoxylate [PTD90]. Albumin [NXHMJ93, GB92]. alcohol [WGR^{+}92]. aldehyde [MYT94]. aldolase [PT92]. aldosterone [DFC^{+}94]. aldosterone-responsive [DFC^{+}94]. algal [WGP^{+}93]. alkaline [Low92, YMO92]. all-or-none [HBP^{+}93]. allele [SBG93, ZRW^{+}94]. alleles [BO91, FKB^{+}90b]. allogeneic [PRB94]. allows [LSR^{+}90a]. along [BRL^{+}93, CSR^{+}92, EHB^{+}90, RA90, SAF90, SFA90]. Alpha [GHS^{+}93, LBF92, ACC^{+}94, ATD^{+}90, AJPB90, BPDM90, BZF^{+}92, BVS^{+}93, BGLB94, BMM^{+}94, CPSS92, CKG^{+}90, CWBK90, CH93, CNSC90, CH92, CTQ91, CPM^{+}91, CMB92, DGF^{+}94, DBM94, DG90, DvHB^{+}91, DMM^{+}91, DGGG93, DBC90, DFC^{+}94, DYS^{+}90, EMGS93, ECAG93, FKHG93, FPP91, FAYM90, FSSA92, FKB^{+}90b, GCZ^{+}92, GSFS92, GAvdM91, GSP^{+}94, GAHB94, GDKP92, GSM93, HRC^{+}90, HC94, HWW^{+}93, HKC92, HAN^{+}93, HLC^{+}93, ILH^{+}90, ISF^{+}92, IBD92, IITG94, JO92, JGO^{+}91, JJS92, KWW^{+}94, KLST93, KDS^{+}91, KAF^{+}93, KYBC94, KWR94, KYLV90, KQJ91, LLSM92, LCD^{+}93, LHT^{+}91, LB92a, LGSB93, LGI^{+}93, LRSB94, LDS^{+}92, MWSP91, MBC^{+}91, MH91, MH93a, Mat94, MWD94, MGD^{+}92, MA90, MJM92, MS94, MR91, MSFW92, MMY^{+}92, ND91, NHH^{+}91, NDdC92, OHS92, OUS^{+}93, OKH92, OVB^{+}94, OPB90, PRF^{+}93, PB91, PLB^{+}92, RG92, RCJ^{+}94, RFF92]. alpha [RPS^{+}92, SMK91, SADJGP94, SMM90, SM93b, SJG93, SJMK93, SR90b, SWF^{+}92, SLM^{+}90, SCJ^{+}91, SSA^{+}90, SdAN^{+}91, SBL^{+}91, STC93, TMP^{+}91, TYM^{+}92, TK94, TRS^{+}90, TTVV^{+}90, TGCT90, VKR^{+}91, VHM^{+}93, VLH^{+}93, WRHW92, WPW92b, WOC91, WGM^{+}93, WWWB90, WKD^{+}91, WBL^{+}91, WFCN94, WHK^{+}93, WLH92, YCO^{+}93, YKM^{+}92, YGM^{+}94, ZVF91, ZMV^{+}93, ZB94, dCTG92, dCQTR91, BPM93, NIT94, SRK^{+}94]. alpha-actin [LDS^{+}92, ND91]. alpha-actinin [CPSS92, CMB92, FKB^{+}90b, HKC92, MWSP91, Mat94, MA90, OPB90, PB91, RFF92, WPW92b, WHK^{+}93]. alpha-actinin-binding [MWD94]. alpha-amidating [MJM92]. alpha-beta [TK94]. alpha-cardiac [KLST93]. alpha-catenin [OUS^{+}93]. alpha-factor [SR90b]. alpha-fetoprotein [TGCT90]. alpha-galactosidase [IBD92]. alpha-helical [WBL^{+}91, dCTG92]. alpha-latrotoxin-induced [TTVV^{+}90]. alpha-lytic [FSSA92]. alpha-mannosidase [MR91, VHM^{+}93]. alpha-protein [GCZ^{+}92]. alpha-satellite [MMY^{+}92, OHS92]. alpha-smooth [DGGG93]. alpha-spectrin [LCD^{+}93]. alpha-spectrin-deficient [DGF^{+}94]. alpha-subunit [GSM93]. also [DJA^{+}90, KZS94]. alter [LW92, PMLM94, PPD92]. Alteration [MMD91, CLG^{+}92]. Alterations [DS90b, CCAF90, IITG94, MHM91, SBWV90]. Altered [ALE^{+}92, EZF^{+}94, HDC^{+}91, JLL^{+}94, AWH^{+}94, CCW^{+}92, EB92, KIV93, PMKH91, SLHG93, SH91, WB92b]. altering [CLZ91, PMA91]. Alternative [SES94, CPU^{+}90, DBC90, HE91a, PAC^{+}92, SJMK93]. Alternatively [SMRG91, CE94, GTH90, KEBD91, MULA^{+}91, OKMB91, PJC^{+}90, SSB^{+}91, WS92]. alters [BR91, LB92a]. alveolar [ML94, WHK90]. alveoli [SMCA91]. Alzheimer [WDB^{+}92, Kos94]. Alzheimer-like [WDB^{+}92]. amastigotes [LEM93]. amebae [FWF90]. amidating [MJM92]. amiloride [SVK^{+}92]. amiloride-sensitive [SVK^{+}92]. Amino [JJS92, NRS^{+}93a, SVB94, WBE91, BR91, BSP^{+}94, DMJ^{+}94, FPR90, GK90, GJJM94, KGA^{+}93, Low92, MYT94, MGD^{+}92, OVB^{+}94, RPE^{+}90, RHH92, SMRG91, SDR^{+}94, SW93b, SMM^{+}91, VWS90, VTSS94, VFC^{+}91, WTH^{+}94]. Amino- [JJS92]. amino-propeptide [SMRG91]. amino-terminal [DMJ^{+}94, SW93b, SMM^{+}91, WTH^{+}94]. amino-terminally [RPE^{+}90]. Aminopeptidase [KCY92, WHF^{+}90, NRS93b]. aminotransferase [NFYI93, PTD90]. amoeba [KPAY91]. amoebae [HZC93]. amoeboid [JT93]. AMOG [GAS^{+}90a]. among [BSG^{+}92, CHVF91, Gau90, KMBR90, LCMP90, TND^{+}93, WBW^{+}94]. AMP [BT93, DES^{+}91, HG94, SAH^{+}92, ZZP94]. AMP-dependent [ZZP94]. amphibian [JG92, PG92, RMG90, SBT93, WMCG91]. amphiglycan [DvdSM^{+}92]. amphiregulin [JSG^{+}92]. amplification [HS90b, LBL^{+}91]. amplification-associated [LBL^{+}91]. amplitudes [KK92]. amyloid [DUVV93, NRS^{+}93a]. analogs [GT94]. analogue [KA91b, MP94, PMKH91]. analogues [MPEG93, TRD^{+}92]. analyses [DGL^{+}90, SLS^{+}90]. Analysis [BOH^{+}92, FMCR93, HKC92, HWL^{+}90, IFLG90, PRGC91, SPvV94, SR92d, AR91, AFHDD90, AKK^{+}93, BTBO94, BWLK93, BBK^{+}93, BM90, CLZ91, CLK^{+}91, CBM^{+}94, DJ93a, DMM93, DWY^{+}93, FGK^{+}90, FE90, FFKL92, FKH^{+}93, FKT93, HLA^{+}90, HGDA93, HDES94, KBP91, Koz91, LCC^{+}93, LSP^{+}94, LTPB92, MECE93, MWD94, MG91b, NIT94, PDTG92, PDSB92, QCCD92, ROE^{+}90, SK94a, SM90, SKSC94, SGC^{+}94, SYG91, SEH^{+}93, SCS90, SBC^{+}93, TASJ92, TSBN94, VBV^{+}90, WHK^{+}93, WWN92, ZSH^{+}93, dCQTR91]. analyzed [LFF91]. Analyzing [Lam94]. Anaphase [RSCS94, BGH93, HSSC92, MECE93, MCG92, SY94a, SM92a, SW92, SMTW94, SH92, TCS^{+}94, WKMG92, ZCH90, MS92]. anchor [BMTTG94, Car91, HGW^{+}91, HFD91, HAG^{+}93, HD90, MMLS92, MC91c, MC91d, MC92, MB93, NWvH94, RMB90, SWFB91, VM93, VM94, XMW^{+}93]. Anchorage [MOT90, AJPB90, FGSBZ93, GA91, HGDA93, RZS^{+}94, TISH94, TTE^{+}94, UK92]. anchorage-dependent [FGSBZ93, RZS^{+}94]. anchorage-independent [GA91, HGDA93, TISH94, UK92]. anchored [BPM93, BGRW91, CS90, DLD^{+}90, DKM^{+}92, ES91, KHW93, LD91, LD92, LFG^{+}90, Low92, MC91c, RYKA90, SHH94, VR92, XMW^{+}93, YOWM91, ZLC^{+}93, vZDF^{+}93, HLRBE93, KGA^{+}93, MNV93]. anchoring [CKG^{+}90, CAS^{+}92, LSK^{+}94, RSB^{+}90, RLKB91, SMM90, SVB94]. anchors [MMC91]. Androgen [DYS^{+}90]. anemia [DGF^{+}94]. angiogenesis [BRIA^{+}94, LIAJS94, NT94, TRLG90, YM90b]. Angiogenic [SRR90, BDZ^{+}92, PBM^{+}90, TVG^{+}93]. angiotensin [SJIC92]. animal [BG94, CW90a, CW90b, PMKH91, WTSP91]. anion [RLWK93]. anisotropy [GT93]. Ankyrin [DMB93, KWFM91, LFB94, GRFB94, GORB91, KDTB90, KB91b, KOB91, NSWH90, PBB^{+}91]. Ankyrin-binding [DMB93, LFB94]. ankyrin-deficient [KB91b]. ankyrin-fodrin [GORB91]. ankyrinB [CKB93]. ankyrins [OKMB91]. Annexin [EGW^{+}93, FFS^{+}93, CE94, HG93a, SPH^{+}91, SSJ^{+}94, WG92b, WMB^{+}94]. annexins [SH90a, SBH92]. annulate [DLMS91]. anomalies [TFC94]. antagonist [JWHPM93, PSE^{+}90]. antagonizes [SR92e]. Antennapedia [BGLJ^{+}93]. anterior [DSJB^{+}94, PNB91]. anterograde [KSYN^{+}94, SSVE91]. anterogradely [DP94, HSYYK90, HSYK^{+}91]. Anti [GAHH94, CL91a, CH93, FAYM90, GMM^{+}93, TVG^{+}93]. anti-adhesive [CL91a]. anti-angiogenic [TVG^{+}93]. anti-beta [CH93]. anti-integrin [FAYM90, GMM^{+}93]. Anti-tumor [GAHH94]. antibiotic [SMHO92]. Antibodies [BTSL91, GKKS90, HTBF94, IMK^{+}92, BBRE90, BI90, BLFQ93, CYC91, DB90, FAYM90, GMM^{+}93, MBG93, MBS91, MLRJ92, NEL^{+}91, SGBN^{+}93, SOH^{+}90, SBBS90, YSK90]. antibody [BMK^{+}90, BFL90, CH93, DSR^{+}93, FFSW91, GAHH94, GG92, Has93, HIJ90, KCYH92, NYA^{+}91, NSKM90, OSK^{+}92, PTMB94, RKB^{+}90, RGT^{+}93, RMG90, ST94, SDW^{+}93, SOKS91, WTS93, ZSM^{+}93]. antibody-coated [PTMB94]. antibody-triggered [DSR^{+}93]. anticentromere [BBRE90]. anticoagulantly [dAWS^{+}90]. anticytoskeletal [PSN^{+}91]. Antigen [PIRB92, Apg91, BHHG90, CKG^{+}90, DJA^{+}90, EFD^{+}93, HKS^{+}92, HEF92, JAB^{+}91, KES^{+}92, KAP^{+}93, LVB^{+}94, MWBD90, MIU^{+}92, NIK^{+}93, NKW^{+}92, OKH92, PMM90, RAGC94, SDW^{+}93, SOH^{+}90, SL90b, TSS^{+}92, WFD92, WCC^{+}90, WKY^{+}93, YOWM91, ZFA^{+}93, aUFQ^{+}92, vKWHF94]. antigen-1 [vKWHF94]. antigen-induced [Apg91]. antigen-processing [RAGC94]. Antigen-receptor [PIRB92]. antigenic [SL90b]. antigenically [CLCC90]. antigens [BZF^{+}92, BNI^{+}94, KYA^{+}94, KHvdL^{+}92, MBC90, REW90, KKBL92]. antiparallel [WDB^{+}92]. antipeptide [BI90]. Antiribophorin [YSK90]. Antisense [SBWV90, FGSBZ93, FNV^{+}92, IITG94, SLS^{+}90, SVD90, TSK^{+}94, WSL91, WTK^{+}92]. antiserum [FL92]. antithrombin [dAWS^{+}90]. aorta [BBK90, dAWS^{+}90]. aortic [MCF93]. AP [BCBK92, PDA93, RMFA92, TOK93]. AP-1 [BMB^{+}92, RMFA92, TOK93]. AP-2 [BCBK92, PDA93]. APC [HBB94]. Apical [CvdE91, DHP93, GORB91, GPRB93, SWS^{+}92, AKM94, BS94a, BMM90, CSS^{+}94, GC93, GIAS93, GAS90b, HLRBE93, HC94, HH90, LSMRB90, LSP^{+}91, LTS93, LTWH92, RBR90, SMSH92, SZKM91, SVK^{+}92, WNBAS90, WHF^{+}90]. Aplysia [KS90, KKBK92]. apoAI [DCH^{+}92]. apoB [DCH^{+}92]. apolipoprotein [DAW^{+}94]. apolipoproteins [DHP93]. apoprotein [TMI^{+}91]. Apoptosis [BBvH^{+}94, KHaA92, BRD94, BMM^{+}94, DJ93a, ET94, EZF^{+}94, FM94b, FF94, LCC^{+}93, LWL^{+}92, OHF^{+}94, RZS^{+}94, SOWDK94, ZZL^{+}91]. APP [DUVV93, DUVV93]. apparatus [BN91, BKR90, DAW^{+}94, DHPG94, DLSB^{+}90, DLSK91, DAIS90, FRC^{+}91, GK92a, HH94a, HSS93, JS92, KA91b, KRB91, MP94, MJG^{+}91, MCJ^{+}93, MSLS91, NPH^{+}93, NRS93b, PMKH91, RDH^{+}92, Rob90, SBW90, SYCA94, TASJ92, WMG^{+}93, WNM^{+}94, WHW^{+}91, WGTCY94]. apparent [HKS^{+}92]. apparently [HTD90]. appear [MM90]. appearance [HM92a, dlRKR^{+}90, dlR91]. appears [BMK^{+}90, NKW^{+}92]. appendage [PDA93]. approach [SBR91]. aqueous [FV91, ZBS^{+}94]. arachidonic [GHR91]. Architecture [AR93, HWM^{+}93, HASW94, Ral93]. area [HGS^{+}94]. ARF [HPR93, TOA^{+}93, WW94a]. ARF1 [ZRW^{+}94]. ARFp13 [WW94a]. Arg [BSSL^{+}91, BDB90, DG90, EUH91, YYY^{+}91, YED^{+}93]. arising [TQL^{+}90]. arm [KKM91, KIO^{+}94, MSFW92, MGW92, PMLM94, SMK91, STK^{+}93, SY90a, SOY93]. armadillo [OUS^{+}93, PMG^{+}92]. arms [GOP^{+}94, GAvdM91, MOM^{+}92, PRSS90, PR91, PMS92, PPD92, SS92a, TK94]. aromatic [NRS93b]. Arp1 [PMTB94, Sch94]. arranged [ABR^{+}92]. Arrangement [MOM^{+}92, OVB^{+}94]. array [BK91]. arrest [BBS92, BS94c, FKB^{+}90a, GN92, Jas93, LKD^{+}94, WBS^{+}91a, YMO92]. arrest-specific [BBS92, BS94c]. arrested [FA91, HWF92]. arrival [SPW93]. arrive [HLRBE93]. arterial [MRBP^{+}90]. arthritic [SFM^{+}90]. articular [LCLG94]. articulins [MB92]. artificial [DUVV93]. ascidian [SSJ90, STH^{+}93]. ASGP [FGS91]. asialoglycoprotein [HKSL90, WL91, YL93]. Asp [DG90, BSSL^{+}91, BDB90, EUH91, GEKL^{+}94, YYY^{+}91, YED^{+}93]. asparagine [HGB^{+}91]. aspects [GZC^{+}91, IK94, SLI^{+}91, YHM^{+}94]. Aspergillus [LOOM93, MM93a, OMRM93, OOM90]. assay [GXHS94]. assays [SS91a]. assemblages [ECO93]. assemble [DBM94, ECO93, GG90]. assembled [BSB92, HW90, HM93, SM91b]. assembles [HEF92, MMY^{+}92]. assemblies [DEH^{+}91]. assemblin [LM91, WGP^{+}93]. Assembly [ABR^{+}92, BGMM^{+}93, BSM90, CFFL93, CL93, DAR93, GWMC90, GFV^{+}91, HWvF90, HWW94, JG92, KB91a, LPA^{+}93, RPE^{+}90, Rob93, SCL91, SDE^{+}93, VFC^{+}91, AP92, BRG^{+}90, BCBK92, BDS92, BSD^{+}92, BM90, BDW92, BSHB90, BTR92, CSSL93, CCAF90, CF90, DLMS91, DGL^{+}90, DDW94, DCTG90, DES^{+}91, DCL^{+}92, FAWM94, FCFL94, FAYM90, GT94, GW94b, GF94, GeKdV^{+}91, HKKO^{+}94, HODF94, HW92, HWM^{+}93, HPH^{+}91, HNPN94, HM91, HCKJ92, HYD93, HHLS92, JJB^{+}93, JR92, KL90, KS94, Knu93, KHKO^{+}93, KQJ91, LTYY94, Lam94, LLK91, LMW^{+}94, MCF91, MKSF93, MGG93, MLRJ92, MKG93, MHG^{+}90, MSLS91, MPTW90, MR92, MHDT94, MG91b, NYA^{+}91, NWD90, ND92, ORS90a, PKN91, PRGC91, PDA93, PPFM90, RMK91, RPPB94, RMR92, RGHC91, Rot90, SCM90, SM91a, SY90a, Sch91, SPvV94, Smi94, TH92a, TDWT92, TM90, TTE^{+}94]. assembly [TYF90, UHFG92, VVB^{+}92, VGC94, VBS^{+}93, VM93, WRA93, WOC90, WKY^{+}93, WCF92, WC90b, ZMV^{+}93, ZCP^{+}94, SRP90]. assembly-defective [WC90b]. assessed [HKSL90, LSP^{+}94]. assessment [DKG92, JKLG94]. assignment [KSE^{+}92, VNS^{+}94]. associate [MCG92]. associated [AU90, AC93, ABD^{+}94, ALE^{+}92, AVP92, AS94, BGGW^{+}94, BBMN94, BSW91, BF93, BN91, BMGN92, BHHG90, BM91b, BWCG93, CYCA90, CYC91, CPB94, DJ93a, DH93, DP94, DMWJ90, DWY^{+}93, EOL^{+}90, FBLL90, FRSL90, FAU^{+}93, FSM^{+}93, FCD^{+}93, GW94a, HRT^{+}91, HSCR94, HKS^{+}92, HHH^{+}93, HAS92, INY^{+}93, JS92, KC90, KKFD90, LA92, LM91, LWL^{+}92, LDT^{+}92, LBL^{+}91, LSR90b, LGSB93, LMW^{+}94, LK93a, MK93, MBB^{+}93, MBHG91, MH93b, MEM94, MC91b, MIU^{+}92, MMLG94, MPVB91, NHK^{+}94, NSKM90, NFL90, OC91, OSJS94, OBBS90, PC91a, PvBeHB^{+}91, POGS90, PDTG92, PMHS92, PBB^{+}91, PHB^{+}91, PBCP91, PE93, PFSS92, PTD90, RAGC94, RBR90, RTC^{+}94, Rob90, SCMB92, SZB94a, SPD^{+}90, SDW^{+}93, SDR^{+}90, SJM^{+}94, TASJ92, TFT94, TSS^{+}92, TLSW93, TAVC91, UOKH93, VGM^{+}93, WBAP91, WGP^{+}93, WFD^{+}91, WC93, WKS^{+}93, WDB^{+}92, WB92a]. associated [YFS92, YMW^{+}90, ZABM94, ZSM^{+}93, ZMMP90, vKWHF94, BDK92, CB92, MSPB90]. associates [BM91a, CSCSA94, HFL94, HSYK^{+}91, JCSH93, MKG93, OBC^{+}92, TMHKO94]. Association [BNI^{+}94, CPSS92, KSVM92, MTC^{+}91, MMD93a, SRK^{+}94, TG94, WLPB92, ZC92, BZT93, BB93, CLCC90, CLG^{+}92, DLSK91, FKS93, FIH^{+}94, HNC^{+}92, HSM^{+}93, KWGS92, KN91, LVBK92, LFF91, MECE93, MBC90, MBZL91, MS94, MPM^{+}94, NI93, NHA^{+}92, OUS^{+}93, PIRB92, RHH92, SLB^{+}94, SBL^{+}91, WNTT94, WGD^{+}93, WHW^{+}91, ZDZSBT94, DG90]. Associations [WKY^{+}93]. aster [FAWM94, WCR93]. asters [VBAK91]. Astral [SH92, HBR90, PSHK92, RA90]. astrochondrin [SNRS93]. Astrocyte [BMS92, SCO^{+}91, LSR90b, LCK^{+}92, SNRS93, HTHC93]. Astrocyte-derived [SCO^{+}91]. astrocyte/T [HTHC93]. Astrocytes [ASG^{+}93, DDFC94, HTHC93, SCO^{+}91]. astrocytic [WSL91]. astrocytoma [CL94]. asymmetric [AHM94, SHLS93, YMW^{+}90]. asymmetrically [LMSH94]. asymmetry [KIO^{+}94]. asynaptic [LKS94]. Atlantic [BSW91]. ATP [BMH^{+}93, FKT93, NFYI93, PMLM94, RPH^{+}92, SC90a, SSK^{+}90, UNY90, WRB^{+}94, YM90a, ZZL^{+}91]. ATP-dependent [BMH^{+}93, NFYI93]. ATP-insensitive [PMLM94]. ATPase [MR94, CS91, DBM94, GMW^{+}94, GAS^{+}90a, GC93, GORB91, JHGR93, KKFD90, KCM^{+}93, LVA^{+}90, MBW^{+}94, MNME^{+}93, MBB^{+}93, NT90, SWS^{+}92, TWC94, VKS^{+}90, VPC^{+}91, WRB^{+}94, ZABM94]. ATPase-like [KCM^{+}93]. ATPases [CF94, SGWE94]. atrophy [RFF92]. AtT [JKS93, KCB^{+}92, MK91a, MJM92]. AtT-20 [JKS93, KCB^{+}92, MK91a, MJM92]. AtT20 [TH91, TH92b]. attach [KCC90]. attached [GA92]. Attachment [MHM91, AR91, ARW^{+}94, BMM^{+}94, BCG^{+}92, Car91, CNSC90, CSRS^{+}90, DSQ^{+}93, FW91, GAvdM91, HRC^{+}90, HD91a, HBR90, HWW94, KGA^{+}93, KKES91, LHH^{+}91, LCPAM94, MCDA^{+}93, MC91c, MC91d, MC92, MC94, PPJF91, RA90, RSCS94, WRHW92, WOC91]. attenuates [SHB90, WGRW91]. atypical [HLL94, KCDR91]. Aurintricarboxylic [BG91a]. auto [REW90]. auto- [REW90]. autoantibodies [PBCP91]. autoantigen [BBZ94, CAS^{+}92]. Autocrine [IBR94, JSG^{+}92, LD92, LK93b, MVM90, OMBL90, RDBHG90, WMS^{+}93, WTK^{+}92]. autocrine-negative [WTK^{+}92]. autoimmune [STS91]. autoinhibition [SLS^{+}90]. autonomous [WE92]. autophagic [BTBO94, Dun90a, Dun90b, HGB^{+}91, LDT^{+}92, NF92, THL^{+}90]. autophagic-lysosomal [HGB^{+}91]. autophagosomes [BTBO94]. Autophagy [TBT^{+}92, Dun90a, Dun90b, Hol93, KKG^{+}90, RSC^{+}93]. autophosphorylation [BSHW91, CGW94, LD91, LG93a]. autoregulatory [BWV92]. Auxilin [AU90]. auxotrophy [DRJ^{+}91]. availability [HFN92]. averaging [SGW^{+}91b]. avian [AGB^{+}91, FBLL90, HR90a, KB91a, LWB91, MGB90, RBR90, VKRD94, WAS^{+}91]. avidity [LB92a]. avoidance [FRSL90]. AVP [PRB^{+}91]. axes [CSR^{+}92]. axial [CHG94]. axis [MBG93, RMMH93, RK94]. axon [BB90, BPK91, BA92, JB93, KSYN^{+}94, MMD^{+}93b, NPST94, OH93, SRB92, TK91]. axon-like [BPK91]. Axonal [EMGS93, HLWL92, PSD92, BA93, BRL^{+}93, KHTD92, LPK92, LHT^{+}91, NFL90, PLS92, PSG94, RMK91, SR92b, TKHM91, VHW^{+}92]. axonally [SKD^{+}91]. axonemal [HHG^{+}94]. axoneme [DAR93, PRSS90]. axonemes [KK92, PMLM94, SS92a]. axonin [KSC^{+}91, SKD^{+}91]. axonin-1 [KSC^{+}91, SKD^{+}91]. axonogenesis [MBS91]. axons [BSB92, CKB93, HLWL92, Hol93, KDTB90, LPK92, MHGC90, OH92, OBBS90, PLS92, VPVC91, WFD92]. axotomy [FFB^{+}93]. aza [BCG^{+}92].

B [INM^{+}93, SDR^{+}90, CL91b, BP91, BZT93, Blo91, BZB^{+}93, BMGC91, CLCC90, CPA^{+}93, CBE90, DIDB^{+}91, ECAG93, GK93a, GI93, HAN^{+}93, HLR92, HEF92, IITG94, KSB^{+}93, vdWvKvKdB^{+}92, KAF^{+}93, KMF^{+}91, LD91, LMW^{+}94, MCG92, MMLG94, MMY^{+}92, NVWR91, OHF^{+}94, PLB^{+}91, PSGE92, RSM^{+}94, SRG^{+}94, SPKV94, SB93c, SW93b, SGGC^{+}90, SH92, VDS^{+}92, VGM^{+}93, WBH^{+}90, WEB90, YKM^{+}92, ZVF91]. B- [SGGC^{+}90]. B-cadherin [NVWR91]. B-cdc2 [PLB^{+}91]. B-crystallin [IITG94]. B-dependent [VDS^{+}92]. B-fragment [SRG^{+}94]. B-type [KMF^{+}91]. B1 [PH91, SF90a]. B2 [GN92, HPH^{+}91, HPWN93, KSE^{+}92, KN91]. B27 [LLK91]. B3 [FKS93]. B4 [DDW94]. b5 [PAC^{+}92]. BA [AGGS92]. Bactenecins [ZLG^{+}90]. bacteria [MB94, PH92, YFS92]. bacterial [CD91, LSA91, TCP90]. BALB [FA91, GTP^{+}92]. BALB/c [FA91]. Balbiani [OOBJ92]. band [FNS^{+}91, Lut91, MNS90]. banded [KLM^{+}91]. bands [DAIS90]. barbed [SYO^{+}91]. bare [Fuc90]. barrier [DDFC94, Gum93, HKS90, NCM^{+}93, REW90, RHP^{+}91, SH90b]. barrier-specific [SH90b]. basal [Ang91, AHM94, HSS93, JD91, MK91a, MS94, MPVB91, SEBH90, TASJ92]. based [AV91, ISF^{+}91, IYN^{+}91, KSYN^{+}94, PTK^{+}93, SS91b, SON^{+}94, TOS^{+}94, WBL^{+}91, YNST93]. basement [DB91, GBOB90, HMY92, KT92, MBC^{+}91, RLKB91, SY90a, SCJ^{+}91, SF91, SBB91b, STA^{+}94, WGM^{+}93, YCC92]. Basic [FAMR92, AAM^{+}93, ASDC91, ATSE90, BDH^{+}94, BGRW91, CL91b, FMR90, GBOB90, GFL92, HHF93, JBJH^{+}92, MMR91, PBM^{+}90, SR90a, SSC90, TSR90, VLH^{+}93]. basis [BMS^{+}91, HWL^{+}90, SLS^{+}90]. baskets [GA92]. Basolateral [BS94a, PIS94, AKK^{+}93, BMM90, CvdE91, CPP^{+}94, GIAS93, GAS90b, HPP^{+}93b, HH90, NI93, THL^{+}90, WNBAS90, YPH^{+}92, ZLC^{+}93]. basolaterally [NLFRB91]. basonuclin [TG94]. basophilic [Apg91, OSK^{+}92]. BB [BRIA^{+}94, OTWH92]. BC3H1 [BEO90]. bcl [BMM^{+}94, Ree94]. bcl-2 [BMM^{+}94, Ree94]. BDNF [KG94, MMW^{+}92]. be [AL91, CH93, CPM^{+}91, FM94a, vdWvKvKdB^{+}92, MM90, RNS92, SPKV94, TDT92, VGMS91]. bearing [YFS92, MCJ^{+}93]. beat [TT94, WYM^{+}92]. before [FWF90, HWvF90, vtHvM90, HLC^{+}93, KT93, LSK^{+}94, NLFRB91, SPW93]. BEHAB [JKH94]. behavior [AR91, AW92, CO91, FKT93, Has93, HLB90, HIJ90, OH92, QvMWV^{+}91, SOE^{+}91, SW93a, TK91, TH92a]. belong [YLWS94]. belonging [SR92b]. belongs [MGS90]. Bem1p [PZB^{+}94]. Bem2 [KFC^{+}94]. Bem2p [PZB^{+}94]. beneath [CBE90, MHYK92]. benzodiazepine [ESL^{+}90]. Beta [BFLS94, FJN94, HG93b, HNH91, HAK94, LKSR94, LNSSA93, MTJM93, ODK^{+}93, PPZ^{+}93, AW91, ACC^{+}94, ALE^{+}92, ASMC^{+}92, BBK^{+}93, BRA^{+}94, BML^{+}90, BZF^{+}92, BRIA^{+}94, BBMN94, BHS91, BGLB94, BSHB90, BLT^{+}90, BFTR93, CPB^{+}93, CKG^{+}90, CWBK90, CSNZ92, CH93, CNSC90, CAS^{+}92, CTQ91, CKD90, DGL^{+}90, DBM94, DG90, DvHB^{+}91, DGGG93, DBSK90, DFC^{+}94, EMGS93, Eri93, ELS93, FKHG93, FAMR92, FAYM90, GHS^{+}93, GCZ^{+}92, GSFS92, GSS^{+}92, GAS^{+}90a, GAvdM91, GLAB93, GAHB94, GRFB94, GDKP92, HRC^{+}90, HGDA93, HS92, HHR^{+}90, HMY92, HMLL94, HSG94, HNP94, HLWL92, HR90b, HLC^{+}93, HBB94, ISF^{+}92, JGL^{+}93, JJB^{+}93, JRCW^{+}90, JLL^{+}94, JGO^{+}91, JRSB90, vdWvKvKdB^{+}92, KWW^{+}94, KWFM91, KFW^{+}91, KLST93, KZS94, KCDR91, KDS^{+}91, KYBC94, KW92a, KWR94, KHR91, KNR93, KBT92, KCYH92, KCC90, KA94, KQJ91, LBF92, LPA^{+}90, LFWC92, LSRC93]. beta [LLSM92, LHT^{+}91, LLK91, LGSB93, LCK^{+}92, LSG92, LCLG94, LKD^{+}94, LGPM90, MBW^{+}94, MMG90, MHR94, MBC^{+}91, MH91, MBG93, MELD94, NS90, NVR92, OSKN94, OVB^{+}94, OPB90, PRF^{+}93, PPHS^{+}92, PH94a, PH94b, PMG^{+}92, PSJ^{+}91, PBM^{+}90, PDWB92, PKG^{+}94, RH92b, RCJ^{+}94, RHH92, RUH^{+}94, RRD93, RPPB94, RDBHG90, RPS^{+}92, STK^{+}93, SHS^{+}93, SADJGP94, SKSC94, SLG92, SH91, SAJ^{+}94, SSW94, SCMU93, SBWV90, SYG91, SMM90, SM93b, SJG93, SEH^{+}93, SSCD90, SPKV94, SLWF91, SEAB91, SLM^{+}90, SCJ^{+}91, SWHCW91, SR92c, SSFD93, STC93, TMHKO94, TMP^{+}91, TYM^{+}92, TK94, TRS^{+}90, TGD93, VBvG^{+}94, VKR^{+}91, VLH^{+}93, WOC91, WGM^{+}93, WKD^{+}91, WFCN94, WLH92, WTK^{+}92, YM90b, YCO^{+}93, YHM^{+}94, ZMV^{+}93, ZB94, ZZR94, FAS^{+}93, LCPAM94, MPIDD93, MGW92, NRS^{+}93a, SCO^{+}91, STL^{+}90, SOA^{+}93, SSK^{+}93, TLXM90, TMI^{+}91, TMHKO94]. beta- [LSG92, SLG92]. Beta-actin [LKSR94, HSG94, KZS94]. beta-adrenergic [KCDR91]. Beta-catenin [HAK94, BVF^{+}93, HNP94, HBB94, KW92a, MBG93, PMG^{+}92]. beta-cells [CAS^{+}92, GCZ^{+}92, KA94]. Beta-COP [ODK^{+}93, PPZ^{+}93]. beta-cytoplasmic [KLST93]. beta-dynein [PKG^{+}94]. beta-galactoside [BZF^{+}92, WLH92]. beta-glucan [BSHB90]. beta-glucosylasparagine [SSW94]. beta-nerve [PPHS^{+}92]. beta-receptor [JRCW^{+}90, SWHCW91]. beta-receptors [BRIA^{+}94]. beta-spectrin [BBMN94, DBSK90, KWFM91, PDWB92]. beta-subunit [CPB^{+}93, SKSC94]. beta-tubulin [HLWL92, PH94b, SH91]. beta-turn [BSW^{+}92]. Beta-very [MTJM93]. beta/A4 [NRS^{+}93a]. beta/IC1 [MGW92]. Betaglycan [LCPAM94]. between [AFA^{+}94, ASL^{+}94, BKWD94, BR94b, BT93, BPS90, BMGC91, CL91a, CLD^{+}91, CCR93, CMB92, DPCP91, DDFC94, ES91, EC93, FTPM^{+}94, GOP^{+}94, GRC^{+}91, HH94a, HP92, HHL94, HPP^{+}93b, JKS93, KKAS90a, KRMK93, KGM92, KHF94, KP90, MG94a, ML91, NDM^{+}94, OCMB91, OPB90, POGS90, PYA90, PZB^{+}94, PCP^{+}91, RHB^{+}90, RPPB94, SJIC92, SPHvD91, SL90b, SDE^{+}93, SR92b, SMWB93, TOS^{+}94, VVR^{+}90, WHLW90, WBW^{+}94, WKY^{+}93, WN92, ZFA^{+}93, eSP91]. BeWo [CvdE91, CSvdE93]. BFA [BS94a]. bFGF [OR92, SMJ^{+}94]. BH2 [PMM90]. BH2-Ag [PMM90]. BHK [RTHB90]. Bidirectional [CSvdE93]. bidirectionally [BPS90]. bifunctional [GPH^{+}90]. BIK1 [BSF90]. Bilateral [aUFQ^{+}92]. bilayer [BVR^{+}93, TSS^{+}92, VM93]. bilayers [CDS^{+}91]. bile [CHBGL91, WTSP91, WMB^{+}94]. BIMA [MM93a]. bimC4 [OMRM93]. bind [AS94, FB93, GG90, MSD92b, PHMH92, PZP^{+}91]. Binding [HNR90, HK91, HWM90, KGR94, OSK^{+}92, TNS^{+}94, TSS91, AA94, ATV^{+}93, BKSB94, BT92, BKB^{+}90, Bea91, BHS91, BTRB93, BMK^{+}90, BZT94, BG91b, BKFT93, BMWT93, BM91b, BSR^{+}90, BML^{+}94, BFTR93, CYCA90, CL91b, CJB^{+}93, CSR^{+}92, CG91, CHG94, DMB93, DCY^{+}94, DAR93, DP92, DKM^{+}92, DWJP93, ENS91, EUH91, EEC^{+}92, FLUS92, FMCR93, FAS^{+}93, FAYM90, Fow90, FTT^{+}93, GSFS92, GWO^{+}93, GFH90, GF94, GYE^{+}90, GT93, HTHC93, HJI^{+}93, HMS^{+}90, HKC92, HAS92, Hen90, HAK93, HWvF90, HNP94, HR94b, HWS90, HSW90, HSB^{+}94, HLC^{+}93, HMT^{+}92, IFLG90, ISDM92, JJ92, JKH94, JS90, JLL^{+}94, JGB^{+}93, KM92a, KSB^{+}93, KHW93, KHF94, KLBBK91, LMH^{+}94, LIAJS94, LWV92, LGSB93, LJHC94, LFB94, LCPAM94, LEM93, Mac92, MKCE92, MJG^{+}91, MSQ^{+}90, MWD94, MWH^{+}91, MB90, MKB^{+}94, MBHG91, MGPPH93, MSSD93]. binding [MMI^{+}91, MHYK92, MSFW92, MKF91, NYA^{+}91, NS93, NIK^{+}93, NF90, ND92, NT90, OWF^{+}93, OKH92, OVB^{+}94, PSZ^{+}93, PWY^{+}92, PHSA93, PTMB94, PMLM94, PBCP91, PE93, RR92a, RWG^{+}92, RG92, RK90a, RDH^{+}92, RKB^{+}90, RWL^{+}90, SLKS94, SPRvD91, SFM^{+}90, SM93a, SOH^{+}90, SGN^{+}93, SLM^{+}90, SSH94, SSA^{+}90, SCF^{+}93, SOKS91, SOY93, SGM90b, TMHKO94, TYM^{+}92, TKHM91, TBL^{+}92, TBKF^{+}92, TSH^{+}93, TYSS92, TTE^{+}94, TRV^{+}90, TAVC91, TGB90, TWV94, VJB^{+}94, VKR^{+}91, VPC^{+}91, VLH^{+}93, VTSS94, WRDK92, WLWL94, WIF^{+}91, WPW92a, WPW92b, WMG^{+}93, WB92a, WP93, YSK^{+}94, YKM^{+}92, ZBW93, aUFQ^{+}92, dWLS91, dAWS^{+}90, dHvBeHVB92, ATSE90, ASE90, CSKZ94, CRWS93, GAS90b, JvCH^{+}94, JGK^{+}94, LS90, MSS^{+}94, NPR^{+}94, ONN91, RDDR93, SPBB92, WDP^{+}94, WHK^{+}93, YTN^{+}94, dWLS91]. binding-like [WIF^{+}91]. binds [BS94b, BFL90, BK91, CNSC90, CKD90, DMJ^{+}94, Fow90, FMK^{+}94, GAHH94, JJ92, JBJH^{+}92, KWFM91, LXM91, MMS93, MVM91, SAJ^{+}94, SNH^{+}94, SOKS91, SS92b, TB90, ZMS^{+}91, ZDR^{+}92]. Biochemical [CPB94, EEC^{+}92, HAN^{+}93, MG91b, SM90, TOK93, CCT94, LSP^{+}94, RMK^{+}94, RS91, SMC92, WHK^{+}93, XL91]. biochemically [CHW94, Wat90]. Biogenesis [COLL^{+}90, DHPG94, LHFV93, SSC^{+}94, YAJ90, YAJ91, CSJD91, DW91, HNvdK^{+}94, HMS94, HVK91, LTS93, RSS^{+}94, SS93, TFTH91, WTH^{+}94, WBH^{+}90, ZHL93]. Biogenetic [LQNRB90, GLQRB91]. Biological [TRD^{+}92, YGM^{+}94, BTRB93, JRK^{+}94, XMW^{+}93]. biologically [BGRW91, ODPL93, OPL94]. Biophysical [DES^{+}91]. biopolymer [JETS91]. Biosynthesis [DCH^{+}92, KGA^{+}93, PDC91, VPVC91, WBWH93, NHTP90, OTWH92]. biosynthetic [RPPB94]. biotin [SW92]. biotin-tubulin [SW92]. BIP [BM91a, BHFS93, BS93b, FGH92, KBA92, KKBK92, SBL^{+}91, VMR90, VPC^{+}91]. BiP/GRP78 [VMR90]. biphasic [BLB^{+}91, PCD^{+}92]. biphosphate [DVT92]. bipotential [WHN^{+}94]. birefringent [UNY90]. bladder [BT93, RGJ91, YLWS94]. blastomeres [HFN92, SN93]. blastula [LSK^{+}94]. blastulae [GHZD90]. bloc [HKKO^{+}94]. block [FM94b, MGDS93, SGBN^{+}93, TGD93, vdBRD^{+}93]. blockers [DMC^{+}91]. blocking [NXHMJ93, WB93a]. blocks [DBWS94, EFD^{+}93, GAHH94, KL90, LTP^{+}92, LCM^{+}94, NSKM90, PMA91, TYM^{+}92, VMR90]. blood [DDFC94, GRC^{+}91, HD91a, REW90, RHP^{+}91, SH90b]. blood-brain [DDFC94, REW90, RHP^{+}91, SH90b]. BMP- [GKY^{+}94]. bodies [BDS92, CFPCL92, CFFL93, CT91, FCFL94, KB91b, MSC92, TH92c, VGMS91, WRP^{+}91]. body [BR94b, BBB91, CFFL93, FFS^{+}93, HWW94, HKS90, JD91, KDKF93, MBG93, MM93a, MPVB91, OSJS94, RK90b, SCF^{+}93, SHOA92, TASJ92, WGBB91, WHC^{+}93]. body-associated [MPVB91, OSJS94]. bodywall [GW94b]. bond [MGW92, SBSG92, YDS92]. bonded [OPL94]. bonds [ZFA^{+}93]. Bone [KYK^{+}94, AWH^{+}94, BGRW91, EFTJ^{+}93, GDB^{+}94, GKY^{+}94, LHH^{+}91, MMI^{+}91, MHG^{+}90, PWY^{+}92, SSG^{+}90, SAVK94, VBvG^{+}94, WBWH93, YKI^{+}91]. bop2 [KIO^{+}94]. bop2-1 [KIO^{+}94]. border [CSM90, EK92, HWM90, MMJ^{+}90, WHFM93]. borders [VKS^{+}90]. BOS1 [SNFN91]. both [ASR91, ASNH93, BB90, BLR^{+}94, BO91, BPS90, DDW90, FT93, HSYYK90, HH90, JS90, KIN^{+}94, MBS91, MS92, MHGC90, OH93, SRG^{+}94, SM93a, TRH^{+}92, VGRC^{+}92, WTH^{+}94, ZMS^{+}91]. bound [CRS94, CG91, HPR93, LVK^{+}91, NT94, SRB92, TOA^{+}93, TSS91, WC90a, NDM^{+}94]. boundaries [BGH90]. Bovine [GMQ^{+}93, SLH92, AU90, HBS92, KYA^{+}94, KSFG91, KNR93, KVF^{+}90, LYC^{+}94, MOT90, STM91, XMW^{+}93, ZVF91, ZLG^{+}90]. box [MMY^{+}92]. bp [MMY^{+}92]. BR96 [GAHH94]. Bradyrhizobium [HWS90, HSW90]. Brain [HSYYK90, AU90, BML^{+}90, BSW91, DMB93, DDFC94, EEC^{+}92, ECM^{+}92, FMR^{+}93, GJK^{+}91, GFM93, JKH94, JCO^{+}94, KYA^{+}94, KWK92, KOB91, LCK^{+}92, MMW^{+}92, NEL^{+}91, OKMB91, REW90, RHP^{+}91, SM90, SH90b, STZG91, SLNN^{+}91, YOWM91, aUFQ^{+}92]. brain-derived [FMR^{+}93, MMW^{+}92]. branching [BGLJ^{+}93, STA^{+}94]. Bravo [KRdlR^{+}92, dlRKR^{+}90, dlR91]. Bravo/ [KRdlR^{+}92]. Breaching [NCM^{+}93]. breakdown [BAP^{+}90, BMS92, FCF92, HAS90, MIS93, TSBS92]. Breaking [Gum93]. breast [JZAVP93, LWL^{+}92]. brefeldin [CWBB93, DLSB^{+}90, DLSK91, IDT^{+}92, KRB91, LSGD^{+}91, LTWH92, MCM92b, NMP^{+}92, PRAK94, PHSvD92, RTC^{+}94, SPHvD91, SFSG91, TGD93, TH92b, WB92b, YCBM94, KDLS92, MMD^{+}93b, RB92, CP90]. brefeldin-sensitive [NMP^{+}92]. bridge [SCF^{+}93, WDWT90]. bristle [CKMC94]. broken [PMS^{+}91]. brown [KCDR91, LVTB93, SGG^{+}91]. Brownian [HLPL90]. brucei [FTPM^{+}94, HAS92, MWLC^{+}94, RG94]. brush [CSM90, EK92, HWM90, WHFM93]. BSA [CVH91]. BSC [ES92]. BSC-1 [ES92]. bud [FMS93, PZB^{+}94, YTN^{+}94]. budding [AJL^{+}90, BDH^{+}94, EOO^{+}94, GHK94, JLC93, JCSH93, KHP91a, KLERG94, KDO^{+}94, LMSA91, PSHK92, PNPB94, RS91, SS91a, ZPS^{+}92]. buffer [SN93]. bulk [CVK^{+}94, MPM93, PMD^{+}93]. bulk-phase [CVK^{+}94]. bullfrog [GH91a]. bullosa [CHVF91]. bullous [CKG^{+}90]. bumetanide [PA91]. bumetanide-sensitive [PA91]. bundle [CKMC94, SAJ^{+}94, WOC90]. bundled [HLPL90, UOKH93]. bundles [GH91a, LWK91, MWSP91, MH94, SLPB93]. Bundling [MA90, FF93, FFW91, VJB^{+}94, ZFF92]. burst [BLSR92, ZB94]. butyric [SRG^{+}94]. Bypassing [SY94a].

C [TCS^{+}94, CMS^{+}90b, DPW91, DPND90, Eri93, EZF^{+}94, GMLD92, GSWW93, JCO^{+}94, LSAH90, MC91a, MC93, RMFA92, RNG94, SDR^{+}90, SMWB93, Apg91, BG94, BGRW91, CXS^{+}94, CETK^{+}94, CE94, DDW90, DVT92, DJA^{+}90, DY93, DCL^{+}92, GCZ^{+}92, GTP^{+}92, GPF94, GG90, GRFB94, HRT^{+}91, HWW^{+}93, HPWN93, JO92, JG92, JK93, KHHC93, Mac92, MSI^{+}94, MBS91, MWLC^{+}94, OWF^{+}93, PIRB92, PvBeHB^{+}91, PBCK93, SPH^{+}91, SZB94a, SFY^{+}94, SR92e, TCDH93, UKH^{+}91, VWT^{+}91, WRP^{+}91, WKM^{+}92, WHN^{+}94, FA91, SBG93, VJB^{+}94]. C- [DDW90]. C-cadherin [BG94]. C-dependent [PIRB92, PBCK93, TCDH93]. c-erbA [MWS^{+}93]. c-erbB-2 [SBG93]. c-fos [CMS^{+}90b, GMLD92, GSWW93, RMFA92, SDR^{+}90]. c-Ha-rasVal [LSAH90]. c-jun [EZF^{+}94, SDR^{+}90]. c-met [JCO^{+}94, RNG94, SMWB93]. c-mos-transformed [DPW91]. C-reactive [Mac92]. c-src [DPND90, Eri93]. C2 [FGH92, OWF^{+}93, VGRC^{+}92]. C2C12 [CCT94, KYK^{+}94]. C3bi [PMM90]. Ca [FMR^{+}92, GPF94, BAP^{+}90, BGK^{+}94, BGB94, BZB^{+}93, CK91, CF94, DVT92, EP94, FF93, GeKdV^{+}91, GZC^{+}91, GDKP92, HFN92, HBP^{+}93, JTS^{+}91, JFA^{+}91, JSA^{+}93, KISY91, KATS90, KSFG91, KC90, LBF92, LS90, MC91b, MOKF92, PMHS92, PNM^{+}94, RBB^{+}94, SLH92, SBL^{+}91, TLWA94, TWMS90, TSBS92, VPC^{+}91, VPP^{+}93, WLWL94, WHK^{+}93, YC93, vKWHF94]. Ca2 [HDS90, TTR^{+}90, vKWH^{+}91]. CaaX [FKS93, KN91]. CaaX-dependent [FKS93]. Caco [vtHvM90, GLQRB91, HH90, LQNRB90]. Caco-2 [vtHvM90, GLQRB91, HH90, LQNRB90]. cactus [WS93b]. Cadherin [MHT^{+}92, BTSL91, BGK^{+}94, CSMG90, CLK^{+}91, FBS^{+}91, HNPN94, HNP94, HAK94, HBB94, IYN^{+}91, KAP^{+}93, KWW^{+}94, MNME^{+}93, MRSN93, NIT94, NVWR91, NHS^{+}94, NSWH90, WNTT94, WHH^{+}93, BTSL91, BZ90, BGB94, BCB93, BG94, CO91, DAS^{+}92, FRL93, JFA^{+}91, KW92a, NVWR91, NGP^{+}91, VR92, WJ92]. cadherin-based [IYN^{+}91]. cadherin-catenin [HAK94]. cadherin-like [CSMG90, CLK^{+}91]. Cadherin-mediated [MHT^{+}92, MRSN93, NIT94]. cadherin/beta [BVF^{+}93]. cadherin/catenin [HNPN94]. cadherins [ASL^{+}94, BCF^{+}93, HW94, INY^{+}93, KSFG91, DSJB^{+}94]. Caenorhabditis [DE93, ECO93, FW91, GW94a, GW94b, HDC^{+}91, HW93a, HWW94, LA92, LH92, SJMK93, WW94b]. Caenorhaditis [CMM94]. Caffeine [DMWJ90, JK93]. caged [FKT93]. CAK [TSBN94]. CAL1 [VMS^{+}91]. Calcineurin [CF94, TH90]. Calcineurin-dependent [CF94]. Calcium [Fuj93, HCYK91, PMA91, SN93, WDT^{+}92, WHFM93, AKC92, BKSB94, BU94, BS91, BL94, CM94a, CNZ^{+}92, CCG^{+}94, CPA^{+}93, Cit92, DSR^{+}93, DCTG90, DP93, DB90, FMCR93, FAF^{+}93, FMJ91, GSS^{+}93, GC91, HN90, HM92b, HWW^{+}93, HMB^{+}93, JLC^{+}90, KII^{+}91, KRCT93, KBP91, KM94, KJ90, LHS^{+}93, LWK91, LOOM93, MM90, MHM91, MMM^{+}92, MLS94, MM91, MHG^{+}90, NS93, NIK^{+}93, NT90, PSZ^{+}93, PBCP91, QH94, RRG^{+}94, SAH^{+}92, SPH^{+}91, Sch93, SCF^{+}93, SSJ90, SMCA91, TNS^{+}94, TT94, TS91, TIA^{+}92, TFDS91, VCFM94, VR92, WCW91, ZCH90, BPZ^{+}92, NT90]. calcium-binding [FMCR93, NIK^{+}93, NT90, SCF^{+}93, TNS^{+}94]. Calcium-calmodulin [WHFM93]. calcium-dependent [HM92b, NS93, TIA^{+}92, VR92]. calcium-evoked [SPH^{+}91]. calcium-mediated [CCG^{+}94, QH94]. calcium-specific [KJ90]. calcium-stimulated [WCW91]. Calcium/ [BPZ^{+}92]. caldesmon [HHY^{+}93, VCL93, WLWL94, YYM90]. caldesmon-containing [VCL93]. caliber [MHGC90, NPST94]. californica [KS90]. Calmodulin [BD92, CSM90, BKSB94, BPZ^{+}92, BSD94, Dav92, EEC^{+}92, ECM^{+}92, GT93, HWL^{+}90, LOOM93, PJS^{+}94b, SLS^{+}90, SHOA92, WHFM93, ZC92, dAWL^{+}90]. calmodulin- [WLWL94]. calmodulin-binding [BKSB94, EEC^{+}92]. calmodulin-dependent [BAP^{+}90, KISY91]. calpactin [NSH90, SPH^{+}91]. calpain [FTT^{+}93]. calreticulin [AKC92]. calsequestrin [VPC^{+}91]. calvaria [ASNH93]. CAM [CMK^{+}90, FvBuHWS92, KA90, BGM^{+}91, FMK^{+}94, GMB^{+}91, GPRB93, KKAS90a, KKAS90b, KRdlR^{+}92, LVB^{+}94, MKCE92, MLRJ92, MJM91, WWD94, SES94]. CAM-stimulated [WWD94]. CAM/ [FMK^{+}94]. camera [KII^{+}91]. cAMP [DDW90, GFWM92, HC94, KM94, LSGD^{+}91, MJM91, NS90, PMHS92, RBD92, RGHC91, SRG^{+}94, SGW91a, SVD90, YR93]. cAMP-dependent [GFWM92, SGW91a, RBD92]. cAMP-independent [DDW90, LSGD^{+}91, RBD92]. can [ASR91, AL91, BO91, BKFT93, Car91, CH93, DBM94, DS90c, EHB^{+}90, vdWvKvKdB^{+}92, LCPAM94, MPK^{+}94, QH94, SWD94, SPKV94, SMWB93, VGMS91, WDT^{+}92, WWD94]. canalicular [GHS^{+}93]. canaliculi [CHBGL91, WTSP91]. cancer [LWL^{+}92, PCG^{+}94, PBCP91, VKRD94]. cancer-associated [LWL^{+}92, PBCP91]. candidates [DMB93]. canine [BBGK90, NSWH90, NT90]. cap [ISDM92, WPS91, WC90a]. CAP-23 [WC90a]. capable [MF90, MPAF91]. capacity [VPC^{+}91]. capillary [DDFC94, GMM^{+}93, IPF^{+}90, SOPA94]. capper [TDWT92]. capping [AGKC92, AC92, DGF^{+}94, HIJ90, MH94, MWBD90, SYO^{+}91, SMC92, SKSC94, WPW92b]. caps [BG94, MMM91, WPBF94]. capsid [LMW^{+}94]. capsulatus [TM90]. capture [HBR90, NW94]. CapZ [Bea91, HMT^{+}92]. Carbohydrate [HWS90, HSW90, NS93, BZF^{+}92, EWP^{+}92, EWP^{+}93a, FWD^{+}92, HSM^{+}93, ISF^{+}91, KKAS90a, MGS90, WS91a, WS91b, YOWM91]. carbohydrate-based [ISF^{+}91]. carbohydrate-derivatized [WS91a, WS91b]. Carboxy [TIA^{+}92, AJPB90, GK90, GSFS92, HW92, HMS^{+}90, JJS92, MSQ^{+}90, NMPN90, SLWF91, VHB93, VFC^{+}91, WTH^{+}94]. carboxy- [WTH^{+}94]. carboxy-terminal [AJPB90, HW92, HMS^{+}90, JJS92, MSQ^{+}90, SLWF91, VFC^{+}91]. carboxyl [KN91, MYT94, NPST94, SG92]. carboxyl-terminal [MYT94]. carboxypeptidase [VWS90]. carcinoembryonic [EFD^{+}93, ZFA^{+}93]. carcinogenesis [NGP^{+}91]. carcinoma [ACC^{+}94, BT93, FBS^{+}91, JSG^{+}92, JZAVP93, KWW^{+}94, KMSW91, LFWC92, RWG90, SEC^{+}93, SBN92, SK91, SSFD93, TTR^{+}90, WNTT94, WHH^{+}93, WTK^{+}92]. carcinomas [HHH^{+}93, SSB93]. cardiac [BKW91, FSL90, FMR^{+}92, FKH^{+}93, JSA^{+}93, KLST93, LDS^{+}92, LSS^{+}90, MBZL91, MLS94, MM93b, SLL^{+}90, SDCD93, TFT94]. cardiomyocytes [DIYSS92, EERM^{+}91]. carrier [PNM^{+}94, SSK^{+}90]. carriers [SP94]. carrot [PJS^{+}92]. carrying [NHK^{+}94]. carteri [KRMK93]. cartilage [AWH^{+}94, BR94a, CLG^{+}92, CMS^{+}90b]. case [NFYI93]. Casein [KMN92, RRD93, YSBM91]. castellanii [BK90, KLBBK91, MZP91]. CAT [DAMS91]. catabolic [WWHA91]. catalase [KLR^{+}93]. catalytic [FCW91, JO92]. catalyzed [SBR91]. catenin [BVF^{+}93, HNPN94, HNP94, HAK94, HBB94, KW92a, MBG93, NIT94, NHS^{+}94, OUS^{+}93, OK92, PMG^{+}92, WNTT94]. catenins [NIT94]. cation [DCCH92, GK92a, KCM^{+}93, MH93a, MH93b]. cation-independent [GK92a, MH93b]. cations [EUH91]. cause [DB90, HKS^{+}92, RFF92, SHOA92, vDS94]. caused [BG91a, BF91, CBM^{+}94, DRJ^{+}91, OGM^{+}93, RLD^{+}91]. causes [CSCSA94, GN92, JM92, ON94, SCMU93, SBG93, TFC94, WBS^{+}91a]. caveolae [CYR^{+}94, FNM^{+}92, Fuj93, LTS93, PJS94a, SOPA94, SSDK^{+}94, SYCA94, SFY^{+}94]. caveolae-mediated [SOPA94]. Caveolin [LTS93, SYCA94, LSV^{+}94, SSTL93, SLB^{+}94]. caveolin-rich [LSV^{+}94, SSTL93, SLB^{+}94]. CBF2 [JLC93]. cCRP [SCMB92]. CD11 [NS90]. CD11/CD18 [NS90]. CD11a [PIRB92]. CD11a/CD18 [PIRB92]. CD11a/CD18-cytoskeleton [PIRB92]. CD11b [DSdF^{+}90, DGAB^{+}93, DS93, MSD92b, RUH^{+}94, GAHB94, ZB94]. CD11b/CD18 [DSdF^{+}90, DGAB^{+}93, DS93, MSD92b, RUH^{+}94, GAHB94, ZB94]. CD18 [DSdF^{+}90, DGAB^{+}93, DS93, MSD92b, RUH^{+}94, GAHB94, NS90, ZB94]. CD18-cytoskeleton [PIRB92]. CD2 [CLD^{+}91]. CD20 [BZB^{+}93]. CD22 [SPKV94]. CD29 [KCYH92]. CD3 [WCC^{+}90]. CD31 [AMBN91, DCY^{+}94]. CD4 [PMBL^{+}92]. CD43 [NXHMJ93, YNST93]. CD44 [CNU92, BBJ^{+}91, CKP91, CMK^{+}90, FSK^{+}92, HTHC93, HLH^{+}92, HHH^{+}93, JJ92, LWV92, LFB94, MTS^{+}94, NI93, PHSA93, SGVS^{+}93, TOS^{+}94]. CD44E [BBJ^{+}91]. CD44H [TBVS92]. CD45 [ACSM^{+}94]. CD54 [DSdF^{+}90]. CD62 [MSD^{+}92a]. CD66 [KHvdL^{+}92]. CD76 [BZF^{+}92]. CD8 [PTMB94]. CD8-positive [PTMB94]. CD9 [MIU^{+}92]. cdc2 [DFC^{+}92, HTP^{+}92, OHF^{+}94, PLB^{+}91]. cdc25 [GSC^{+}92]. Cdc28 [LR93]. cdc3 [BHBG94, KHP91a]. Cdc31p [BR94b]. CDC42 [AJL^{+}90, JP90]. CDC43 [AJL^{+}90]. CDC48p [FFR^{+}91]. CDK [TSBN94]. CDK-activating [TSBN94]. cdk2 [PPL^{+}93]. cDNA [CH93, FGS^{+}92, FPR90, INY^{+}93, LCMP90, SNE90, TMP^{+}91, UK92, WGD^{+}93]. cDNA-deduced [FPR90]. cDNAs [CFG^{+}94, MMJ^{+}90, MR91, NHH^{+}91, OKMB91, WG92b]. CDw75 [BZF^{+}92]. CE9 [NCM^{+}93, PNB91]. cease [Jas93]. CEF [BIPG91]. CEF-10 [BIPG91]. CEF-4 [MGB90]. Cell [AV91, BS90, Bis90, BKR90, CO91, CGW94, CE94, DKM^{+}92, DFC^{+}94, EG91, FPPD91, FHUY93, GBN^{+}93, HS92, HM91, HKS90, KGK90, KES90, LCD^{+}93, LRSB94, RR92b, SOWDK94, SWM^{+}92, SOY93, TSBN94, VHM^{+}93, WMCW92, ASR91, AJL^{+}90, Ada92, AAM92, ASG^{+}93, AORB90, AMBN91, ARW^{+}94, AG93, AUT^{+}91, AGGS92, AHS^{+}92, ATD^{+}90, AVP92, ASL^{+}94, BP91, BEGN91, BBZ94, BRA^{+}94, BT92, BSSL^{+}91, BMTTG94, BKB^{+}90, BZF^{+}92, BG91a, BG93, BRD94, BDMPJ94, BHS91, BFM93, BGK^{+}94, BLR^{+}92, BJ90, Blo92, Blo93, BS93a, BSHB90, BCG^{+}92, BF93, BT93, BKFT93, BDB90, BD92, BSD94, BL94, BZB^{+}93, BGM^{+}91, BTR92, BML^{+}94, BDZ^{+}92, BFTR93, CSC90, CL91a, CdPA^{+}93, CSMdPSM94, CSCSA94, CWBK90, CHBGL91, CYR^{+}94, CMS^{+}90a, CXS^{+}94, CW90c, COLL^{+}90, CSR^{+}92, CdJC93, CKL^{+}91, CKD90, CMK^{+}90, DMB93, DCY^{+}94, DJ93a]. cell [DG90, DVT92, DJA^{+}90, DCB^{+}92, DY93, DAS^{+}92, DEH^{+}91, DGGR92, EGP^{+}91, EMGS93, EHB^{+}90, EOL^{+}90, EWP^{+}92, EWP^{+}93a, EZF^{+}94, ELS93, ELZ93, FSK^{+}92, FGK^{+}90, FA90, FMPP90, FBL92, FGSBZ93, FCFL94, FRSL90, FAU^{+}93, FP91, FW91, FFB90, FvBuHWS92, FMK^{+}94, FF94, FBS^{+}91, FFR^{+}91, Fuc94, FWF90, FV91, GN92, GHZD90, GAHH94, GBP^{+}92, GSBS92, GL91, GPF94, GSC^{+}92, GSR^{+}93, GAvdM91, GHT92, GFL92, GXHS94, GC90, GC91, GK92a, GMB^{+}91, GFM93, GA91, GVK94, GS94, HTHC93, HRH90, HJI^{+}93, HRC^{+}90, HFN92, HGDA93, HSP^{+}91, HHO^{+}92, HKS^{+}92, HMS^{+}90, HDS90, HAK93, HSG94, HNP94, HAA93, HHY^{+}93, HWW94, HMB^{+}93, HLR92, HFS92, ISF^{+}92, ISN^{+}94, IVBR93, IYN^{+}91, JM92, JvCH^{+}94, JP90, JSG^{+}92, KKAS90b, KES^{+}92, KT92, KSL^{+}94, KAV93, KRdlR^{+}92, KBCM^{+}90, KSB^{+}93, KWW^{+}94, KA90, KDKF93, KHP91a, KRMK93]. cell [KZS94, KCDR91, KYBC94, KKL^{+}91, KHR91, KW93, KB91b, KDN94, KT93, KDT93, LTYY94, LRR^{+}92, LS90, LPA^{+}90, LCC^{+}93, LQNRB90, LFWC92, LSRC93, LTP^{+}92, LYC^{+}94, LDNM92, LBH92, LR93, LML^{+}94, LCT^{+}93, LVB^{+}94, LW92, LMB^{+}94b, LJP^{+}93, LS92, MMG90, MTS^{+}91, MNME^{+}93, MSC92, MK91a, MHT^{+}92, MMJ^{+}90, ML94, MG94a, MKCE92, MMC91, MPM93, MZRM93, MSQ^{+}90, MT94, MRSN93, MBHG91, MWG92, MFS^{+}94, Mil90, MCM92b, MW94, MS92, MDRG92, MVM90, MM92, MPK^{+}94, MCEG90, NIT94, NYA^{+}91, NBT94, NGP^{+}91, NS93, NSWH90, NVR92, NEL^{+}91, NDdC92, NPST94, OHS92, OTT92, OVB^{+}94, OCMB91, OK90, PPL^{+}93, PA91, PWY^{+}92, PLS92, PH94a, PSG^{+}90, PW93, PJS^{+}92, PBB^{+}91, PKPM90, PTMB94, PH91, PSE^{+}90, PPJF91, PRB^{+}91, PFSS92, QvMWV^{+}91, RB94, RWG^{+}92, ROE^{+}90, RZS^{+}94, Ree94, RK94, RCJ^{+}94, RAGC94, RGJ91]. cell [RPH^{+}92, RGT^{+}93, RHR90, RBD92, RWG90, RHG90, RHP^{+}91, RRBG91, RPS^{+}92, SWD92, SHS^{+}93, SEC^{+}93, SOA^{+}93, SH90b, SDW^{+}93, SBN92, SBWV90, SMB^{+}91, SW93a, SNG^{+}91, SWMS90, SCS90, SDCD93, SGN^{+}93, SPKV94, SLWF91, SMCR90, SHBD90, SKM94, STF93, SGP91, SLM^{+}90, SCJ^{+}91, SSA^{+}90, SCF^{+}93, SB91, SSB93, SNRS93, SBB91b, STdL^{+}90, SWFB91, SSM92b, SM92b, TSK^{+}94, TCDH93, TRLG90, TWC93, TM92, TBVS92, THBW94, TFT94, TDT92, TDWT92, TSS^{+}92, TSS91, TM90, TIA93, TRV^{+}90, TSR90, TYF90, TOA^{+}91, TOS^{+}94, TAVC91, TRP^{+}93, UK92, UHFG92, UYA^{+}91, VKR^{+}91, VGRC^{+}92, VFR94, VVR^{+}90, VBS^{+}93, VR92, VM94, VKRD94, VDTT91, WLPB92, WBAP91, WFD92, WSC94, WNTT94, Wat90, WOC91, WS91b, WMvdK94, WS93a, WE92, WG92b, WFCN94, WC93, WBH^{+}90, WCC^{+}90, WKY^{+}93, WDT^{+}92, WHH^{+}93, WKN92, WE90, WTS93]. cell [WP93, YCBM94, YM90b, YED^{+}93, YDS92, YCO^{+}93, ZMV^{+}93, ZSH^{+}93, ZCP^{+}94, ZZL^{+}91, ZBS^{+}94, ZMMP90, ZDR^{+}92, dCV90, dBNI^{+}91, dHRB^{+}93, vKWHF94, GK93a, SBL^{+}91]. Cell- [EG91]. cell-associated [TAVC91, ZMMP90]. cell-basement [SCJ^{+}91]. cell-binding [KSB^{+}93, NYA^{+}91, PTMB94, SLM^{+}90, SSA^{+}90, TRV^{+}90]. Cell-free [WMCW92, LCC^{+}93, MSC92, MMC91, MW94, MPK^{+}94, TM90, UHFG92, YDS92]. cell-matrix [BF93, ELS93, FF94]. cell-mediated [RB94]. Cell-specific [DFC^{+}94, KGK90, LRSB94]. cell-substrate [ASR91, CWBK90, SM92b]. cell-substratum [CL91a]. cell-surface [BZF^{+}92, VM94]. cell-to-cell [TOA^{+}91]. cell-type [SOA^{+}93, SW93a]. cells [AO90, AMG90, ACC^{+}94, AB91, AW92, ABD^{+}94, AM94, ACFFL93, Apg91, AKM94, ABM^{+}93, ATSE90, ASE90, AL91, ASNH93, AVP92, ASL^{+}94, ASK^{+}90, BPK91, BP93, BRA^{+}94, BGGW^{+}94, BS94a, BG91a, BSG^{+}92, BTH^{+}90, BLW92, BS90, BVF^{+}93, BF90, BPZ^{+}92, Bis90, BRF93, BBD^{+}90, BHC^{+}94, BST^{+}90, BG91b, BIPG91, BCB93, BS94c, BPS90, BPH^{+}90, BMM90, BEO90, BM91b, BBGK90, BHLA90, BML^{+}94, CW90a, CW90b, CBRW92, CFW93, CSS^{+}94, CS90, CSCSA94, CM94a, CWBK90, CHBGL91, CvdE91, CSvdE93, CJC93, CRKA92, CNZ^{+}92, CO91, CL94, COTC94, CCG^{+}94, CL93, CAS^{+}92, CML90, Cit92, CMS^{+}90b, CFG^{+}94, CTQ91, CKD90, CCW^{+}92, CM93, CPP^{+}94, CCT94, CSH^{+}92, CC90b, DPW91, DPCP91, DGF^{+}94, DDW90, DR91, DPND90, DvdSM^{+}92, DS90a, DS90b, DG90, DH93, DvHB^{+}91, DDFC94, DP91, DMM^{+}91, DSQ^{+}93, DLSK91]. cells [DS90c, DMB92b, EFTJ^{+}93, EVW91, EP94, ES92, ECAG93, EB92, FB93, FTB94, FA91, FBML90, FGS^{+}92, FGSBZ93, FKS93, FHvZ^{+}94, FCW91, FW93, FRSL90, FAMR92, FAS^{+}93, FAU^{+}93, FGH92, FA92, FFB90, FBS^{+}91, GN92, GCM^{+}94, GCZ^{+}92, GBN^{+}93, GHGP^{+}92, GMGPM94, GHR91, GLQRB91, GFWM92, GSC^{+}92, GT94, GWMP91, GIAS93, GXHS94, GAS90b, GK92a, GRC^{+}91, GEKL^{+}94, GSWW93, GK92b, GZC^{+}91, GFV^{+}91, GHH90, GH91b, HLA^{+}90, HTHC93, HJI^{+}93, HLRBE93, HC94, HSHA90, HG92, HHR^{+}90, HMY92, HBR90, HDS90, HBP^{+}93, HNP94, HW93a, vtHvM90, HSUS90, HMB^{+}93, HFS^{+}93, HH90, HVNC92, INM^{+}93, IK94, IPF^{+}90, INY^{+}93, IDT^{+}92, JPJF91, JT93, Jas93, JGK^{+}94, JZAVP93, JFA^{+}91, JJS92, JH90, JKS93, KYA^{+}94, KD92, KVE92, KSFG91, KGSB92, vdWvKvKdB^{+}92, KWW^{+}94, KAG94, KFW^{+}91, KS92, KMSW91, KDS^{+}91, KBP91, KGA^{+}93]. cells [KCB^{+}92, KTG90, KCR^{+}94, KNR93, KBH91, KJT91, KW92b, KBT92, KCYH92, KKFD90, KCC90, KDT^{+}91, KRB91, KHvdL^{+}92, KHHR92, KA94, LBF92, LSMRB90, LSP^{+}91, LJJC93, LM94, LLIV94, LBL^{+}91, LLG^{+}94, LMIV93, LB92a, LKXS93, LLDJ94, LCT^{+}93, LB92c, LSN^{+}94, LTWH92, LGH91, LJP^{+}93, MSC91, MBW^{+}94, MK93, MMDM93, MHYC90, MCM93, MW93, MMJ^{+}90, MYM94, MMS93, MOMM92, ML91, MWG92, MHV92a, MLS94, MELD94, MMR91, MJM92, MME93, MM91, MIU^{+}92, MMI^{+}91, MHYK92, MOKF92, MHGC90, MSD^{+}92a, MC94, MHP90, MJM91, MAAO92, MYC92, MJ92, MCF93, NHS^{+}94, NNHB^{+}92, NI93, NSWH90, NPH^{+}93, NFL90, NPR^{+}94, OSKN94, OSK^{+}92, OCC94, OBBS90, PMKH91, PT92, PDB^{+}91, PKN91, PZP^{+}91, PC90, PLB^{+}92, PBD93, PBM^{+}90, PSS^{+}93, PBDK90, PMHS92, PBCK93, PYA90, PCG^{+}94, PTMB94, PIS94, PDSB92, PRC^{+}90, PHWM92]. cells [PMD^{+}93, PJS^{+}94b, PHSvD92, PMS^{+}91, QPWG91, QDT^{+}90, RPE^{+}90, RLRR93, RLD^{+}91, RZS^{+}94, RK94, RBR90, RK90a, RA90, RSCS94, RBB^{+}94, RTHB90, Rob93, RBD92, RM94, Rot90, RA94, SWS^{+}92, SR90a, SMRG91, SW90, SPRvD91, SPHvD91, SSTL93, SNE90, SF90a, STL^{+}90, SOA^{+}93, SRV^{+}90, SZB94a, SVR^{+}90, SMC92, SGW91a, SPD^{+}90, SLP91, SHDS93, SH90b, SC90a, SZKM91, SK91, SCMU93, Sch93, SDK^{+}93, SRR90, SW92, SNG^{+}91, SZP^{+}93, SSS93, SSJ^{+}94, SZB^{+}94b, SKK^{+}91, SMJ^{+}94, SBR91, SGM^{+}90a, SG92, SSC90, SSFD93, SdAN^{+}91, SSKT^{+}92, TSP90, TLWA94, TGD93, TTR^{+}90, TND^{+}93, TH91, TH92b, TFP94, TGS92, TGR^{+}92, TSR90, TNT^{+}90, TA91, TAVC91, TRH^{+}92, UK92, VGMS91, VBvG^{+}94, VBV^{+}90, VSVP93a, VVR^{+}90, WG92a, WGR^{+}92, WNBAS90, WHM^{+}92, WLZB93, WK92, WBT^{+}94, WBVB90, WSB93, WFD^{+}91]. cells [WHF^{+}90, WHN^{+}94, WGD^{+}93, WMS^{+}93, WN92, WA93, WWN92, WTK^{+}92, WB93b, XS93, YYM90, YYM94, YFS92, ZJA^{+}94, ZCH90, ZMS^{+}91, ZPS^{+}92, ZLC^{+}93, dAWS^{+}90, dBNI^{+}91, dSBH90, dTF90, eSP91, vGvMS^{+}91]. cellubrevin [GCM^{+}94]. Cellular [EKK94, KHP91a, SLK91, Ada92, AMBN91, BML^{+}90, BST^{+}90, BG91b, BCB93, CNSC90, CCW^{+}93, DPND90, DVT92, EK92, ECM^{+}92, FYDB93, FJRAT91, FFKC94, FKLS92, GTP^{+}92, GT93, HKS^{+}92, HTGN94, HSP93, HWL^{+}90, HNH91, JRK^{+}94, JKLG94, KFC^{+}94, KKL^{+}91, LFR93, LB93, LEM93, MYC92, OR92, PALM92, PPFM90, RHG90, RHB^{+}93, SWD94, SH90a, SBH92, SFF93, SLNN^{+}91, WKM^{+}92, WCM^{+}93, WS92, ZDO^{+}91]. CENP [BCYC94, CBE90, MMY^{+}92, PSGE92, SHM94, TCS^{+}94, YKM^{+}92]. CENP-B [CBE90]. CENP-C [TCS^{+}94]. centers [AL92, DMJ^{+}94, SAWL90]. centractin [CM94b]. central [AST^{+}92, HSSC92, KDTB90, MECE93, MBW^{+}94, NYA^{+}91, PC91c, WBL^{+}91]. Centrin [SS94, TASJ92]. centriole [KB90]. centrioles [BBGK90, NKW^{+}92]. Centromere [MMY^{+}92, CYC91, CBE90, MECE93, PSGE92, SBBS90, SHM94, YKM^{+}92, ZMB91]. centromere-kinetochore [ZMB91]. centromere/kinetochore [CYC91]. centromere/kinetochore-associated [CYC91]. centromeres [FHUY93, PC91c]. centromeric [AS94, MMY^{+}92]. Centrophilin [TZBV91]. centrosomal [CTPP92, KDP^{+}90, MPVB91, RKA93]. Centrosome [FAWM94, GHZD90, BBZ94, ES92, SMCR90, WCR93, YCAB93]. centrosomes [BDK92, TKP^{+}91]. ceramide [PMKH91]. cerebellar [HFS92, KG94, LCT^{+}93, ORD90, SRV^{+}90, SNRS93, WAS^{+}91]. Cerebroglycan [SLL94]. cerevisiae [CM94b, DRJ^{+}91, JKLG94, LBH92, AYM^{+}93, AJL^{+}90, AC92, BZT94, BSF90, BD92, BSD94, BGH93, CYCA90, CSV^{+}92, CF94, DKG92, GS90, GK93b, GSP^{+}94, GFW92, HLA^{+}90, HVK91, HYD93, HHLS92, JP90, JPS91, KCH^{+}94, KAG94, KDKF93, KHP91a, KCY92, KLR^{+}93, KDT93, LPPCF91, LB94, LSA91, MZRM93, MJG^{+}91, MT94, NHO^{+}91, PSHK92, PDC91, RRCR93, RHF91, SP92, SMB^{+}91, SGP91, SCF^{+}93, SH92, VMS^{+}91, VVB^{+}92, VRE^{+}92, WF91, WDP^{+}94, YTN^{+}94, ZHL93]. cGMP [KIV93]. chain [AB90, BHS94, CWB93, CBM^{+}94, DAM^{+}93, DH93, DBC90, DKG92, FBML90, FCW91, GHR^{+}90, GSFS92, HMS^{+}90, Hen90, JH90, KSE^{+}92, KM92c, KB91a, KLBBK91, LPPCF91, LCKW91, LLK91, LFC94, MBC90, MGD^{+}92, Mil90, NNHB^{+}92, OA92, OTX^{+}94, OCC94, PSRS94, PHB^{+}91, PKG^{+}94, RAGC94, RCNO94, SMK91, STK^{+}93, SLW^{+}92, SP92, SLS^{+}90, SWMS90, SLWF91, SWF^{+}92, SJM^{+}94, SBL^{+}91, TWC94, VNS^{+}94, WYM^{+}92, WS90, WBL^{+}91, WW93, WHW^{+}91, YYM94, YKY92, dCTG92, DAMS91, JRCW^{+}90]. chain-2 [WYM^{+}92]. chain-related [PSRS94]. chains [BP91, ECM^{+}92, GTC^{+}90, KA91a, LWB91, MS94]. challenges [SR92c]. change [Ada92, BF93, BR91, Low92, NSH90, SHTL93, SF90b]. Changes [KTG90, Ral93, BTS93, BFR^{+}92, DPCP91, DBV^{+}92, DCY^{+}94, DY93, DWY^{+}93, FGSBZ93, FFB90, GHS^{+}93, GCML94, GPF94, GB90, GXHS94, HKS^{+}92, HFAGM94, KBP91, MMJ^{+}90, NHA^{+}92, OSK^{+}92, PSG^{+}90, PMHS92, RTHB90, SNG^{+}91, SKBD91, SEMH92, SLNN^{+}91, TFT94, TH92b, VCFM94, WGRW91, YM90b]. channel [DPS^{+}93, DFC^{+}94, FKG94, FSL90, FAF^{+}93, HWW^{+}93, HWL^{+}90, JSA^{+}93, LVA^{+}90, MTS^{+}91, MMM^{+}92, PYA90, RRG^{+}94, SSC^{+}94, SVK^{+}92]. channel-mediated [HWL^{+}90]. channel/ryanodine [JSA^{+}93, RRG^{+}94]. channeled [ZKMB93]. channels [ABM^{+}93, BU94, DSR^{+}93, GFWM92, LC91, LC94, MHG^{+}90, NSC^{+}93, PSN^{+}91, SDK^{+}93, SKGU91, VBS^{+}93]. chaperon [KKBK92]. chaperone [HSCN94, KBA92]. Chaperonin [MVCC93, GMW^{+}94, HH94b, LMW^{+}94]. Chaperonin-mediated [MVCC93]. characteristically [KYLV90]. Characteristics [KTR90, RLS94, Fri94, OHC^{+}93, WBVB90, WN92]. Characterization [BEGN91, BOM94, CD91, CMS^{+}90a, CW90c, ET94, EWP^{+}93b, FP91, KLERG94, LSV^{+}94, LB92c, LSA91, MZP91, MMD92a, MKB^{+}94, MGPPH93, ND92, ORS^{+}90b, PSG94, PAFC92, RPPB94, STL^{+}90, SKH90, SSWS92, SF91, SIR^{+}94, TFTH91, VKR^{+}91, WDP^{+}94, WC90b, ZWW90, AC92, BTBO94, BGH90, BGK^{+}94, BWCG93, CSC90, CES^{+}92, CYR^{+}94, CLP^{+}94, CPB94, DPW91, ENS91, EK92, EEC^{+}92, ESL^{+}90, FSL90, FL90, FWBS93, GFM93, HM94, HPH^{+}91, HHF93, HSW90, HAN^{+}93, HMT^{+}92, JG94, JLC93, JP90, JH90, KCH^{+}94, KWS94, LDW^{+}93, LHFV93, LMM^{+}91, MK93, MBB^{+}93, MR91, MLM^{+}93, MSL92, NYA^{+}91, NVWR91, ODPL93, OMBL90, OKMB91, PH92, RHGG92, RB94, RKB^{+}90, RGR90, RKB93, SS93, TWQ^{+}93, TKHM91, TYF90, TSR93, TRP^{+}93, WG92b, WST^{+}90, YCBM94, ZABM94, DJA^{+}90]. characterize [Has93]. characterized [SSB^{+}91]. checkpoint [RSCS94]. Chemical [WSSM91, Ada92]. chemically [BR94a, DS90b]. Chemoattractant [CMZ90, CLZ91, GPF94, MC91b, SY90b, TTR^{+}90]. chemoattractant-induced [GPF94]. Chemoattractant-stimulated [CMZ90, SY90b]. chemotactic [KIV93]. chemotaxis [ASK^{+}90, CCG^{+}94]. chick [FRL93, GBN^{+}93, GDB^{+}94, NVWR91, OH90, SAM^{+}91, TGR^{+}92, dCQTR91, dlRKR^{+}90, dlR91]. Chicken [KA91a, BGLJ^{+}93, BGM^{+}91, CPB94, DCH^{+}92, DBC90, ECM^{+}92, FFW91, HHR^{+}90, HPH^{+}91, KCCG91, LGI^{+}93, NVR92, SEC^{+}93, SMC92, VPC^{+}91, WBAP91, YM90b]. chimera [WIF^{+}90]. chimerae [TBL^{+}92]. chimeras [CCW^{+}93, Rob93]. chimeric [LTYY94, YYY^{+}91]. Chinese [BGGW^{+}94, CSC90, KD92, RHR90, TYF90]. CHIP28 [NSC^{+}93, SSC^{+}94, VBS^{+}93, ZSH^{+}93]. chitin [SMB^{+}91, VMS^{+}91]. Chlamydomonas [BGBWO91, BU94, BBKR94, BS94b, CK91, DA92, DAR93, GOP^{+}94, GSS^{+}93, HSS93, HW93b, HHG^{+}94, HKS90, JD91, JR92, KKM91, KIO^{+}94, MOM^{+}92, MCDY94, MK91b, PRSS90, PR91, PMS92, PMLM94, PPD92, PKG^{+}94, QYC^{+}92, QH94, SSG93, SMK91, STK^{+}93, SS94, SH91, SS92a, TASJ92, TK94, WVH94, YC93, ZS94]. chlorophyll [CL91b]. chloroplast [CLDV90, CL91b, SBP90]. chloroplasts [ET91, SB93a]. CHO [BRA^{+}94, FYDB93, GCM^{+}94, PMD^{+}93, WSSM91, YCBM94]. cholera [LDNM92]. choleretic [WMB^{+}94]. cholestatic [WMB^{+}94]. Cholesterol [MPK93, PK91, RYKA90, CSC90, CRKA92, SYCA94]. chondroblastic [ASNH93]. chondrocyte [BS91, GL91, Knu93, QCCD92]. chondrocytes [AWFP93, BR94a, BCET^{+}92, CGMC92, GBN^{+}93, IBR94, LCLG94, MSL92, QDT^{+}90, THH^{+}90, VIF^{+}93]. Chondrodysplasia [MGD^{+}92]. chondrogenesis [GDB^{+}94, JRSB90]. chondroitin [FSK^{+}92, FCL^{+}94, FMK^{+}94, GFM93, ISF^{+}92, MFS^{+}94, SFSG91, SDH90]. chorioallantoic [YM90b]. chorion [HS90b]. chromaffin [DVT92, EOL^{+}90, MWS^{+}93, SLH92, VDTT91]. chromatid [BB94, GHK94, SM92a]. Chromatin [OHF^{+}94, Woo94, ASR^{+}90, BZT93, DLMS91, DDW94, GCML94, GGGE^{+}92, HODF94, HM91, HASW94, MCG92, PC91c, SNGG94, SAWL90, SBT93, WE90, Wu93, XL91]. chromatin-depleted [XL91]. chromatography [MAA^{+}94]. chromonema [BB94]. chromosomal [FBL92, KSE^{+}92, MAB^{+}90, VNS^{+}94]. chromosomally [GHBA^{+}93]. Chromosome [AR91, GHK94, AS94, BBRE90, CYC91, CLM91, GK93b, HBR90, HAS90, HDP^{+}93, ISN^{+}94, KWK92, OHS92, RA90, RS94, RTHB90, RMG90, SGWE94, SBK94, SW92, SBBS90, SSS93, WKMG92, ZCH90]. chromosome-associated [AS94]. chromosome-specific [OHS92]. chromosome-to-pole [SW92]. chromosomes [AR91, BB94, GG90, GR93, HM93, KLC^{+}92, NKW^{+}92, SNGG94, SMTW94, SGGC^{+}90, TH92a, WSSM91]. Cik1 [PSRS94]. Ciliary [MWBD90, LSR90b, SST92b, TT94]. ciliate [OMBL90]. cingulin [ZSM^{+}93]. circumsporozoite [PHMH92]. Cis [DHS93, ABR^{+}92, AES94, BLR^{+}94, CSR^{+}92, ODK^{+}93, RTC^{+}94, SMKH91, SXM94, WRP^{+}91]. Cis- [DHS93]. cis-acting [CSR^{+}92]. cis-Golgi [AES94, BLR^{+}94, ODK^{+}93, RTC^{+}94, SMKH91]. cis-trans [ABR^{+}92, SXM94]. cis-unsaturated [WRP^{+}91]. cisternae [CP90, PGAR90]. cisternal [CM94a]. Cl-cotransport [PA91]. clam [DIDB^{+}91, HLR92]. clamp [dTF90]. clamped [RSS91, SS90a]. Class [HG92, CNJ^{+}94, DW91, HWW^{+}93, HLWL92, KHHC93, MPR90, MBC90, MB92, NVR92, OTX^{+}94, PL90, PHB^{+}91, QXWN^{+}94, RAGC94, STM91]. classes [JBIM94, TA91]. classic [JM92]. classification [SFC^{+}94]. Clathrin [BCBK92, CVK^{+}94, GSP^{+}94, HSvD93a, MLR^{+}92, OA92, RCNO94, TDSP93, AB90, AU90, CRWS93, CJC93, HSvD93b, LPPCF91, LMSA91, PDA93, Rob90, RPH^{+}92, RYK^{+}90, SP92, SHH94, SWMS90, SCS92, WRA93]. clathrin-associated [AU90]. Clathrin-coated [MLR^{+}92, CRWS93, CJC93, LMSA91, PDA93, Rob90, RYK^{+}90, SHH94]. Clathrin-dependent [GSP^{+}94, SCS92]. clathrin-independent [HSvD93b]. clear [BTH^{+}90]. Cleavage [BTRB93, IKFR91, BG93, CW90a, CL91b, DUVV93, GCM^{+}94, KIS^{+}94, KGA^{+}93, MB93, OHP^{+}94, OK90, PNB91, SYO^{+}91, SKT^{+}90, SN93, VSB^{+}94, WF93, WL91, XS93, YNST93, SWL^{+}92]. Cleavage-dependent [SWL^{+}92]. cleaves [BLT^{+}90]. cleaving [CSMG90, FKB^{+}90a, WF91]. clone [CH93]. cloned [HVNC92]. clones [QDT^{+}90]. Cloning [CLK^{+}91, NNHB^{+}92, NHH^{+}91, Rob90, SS93, WBL^{+}91, ZSH^{+}93, BGK^{+}94, CES^{+}92, DLD^{+}90, DvdSM^{+}92, FCD^{+}93, HHO^{+}92, HSLF^{+}92, ILH^{+}90, INY^{+}93, LCMP90, LGSB93, MBB^{+}93, NVWR91, OOM90, RKB93, TMP^{+}91, TWQ^{+}93, WHD^{+}91]. closely [KRdlR^{+}92, LWV92, MH93b, VKRD94]. closure [BFM93]. clots [WS92]. clover [LS92]. cluster [BP93, CL92]. clustered [HLRBE93, RR92b]. clusterin [FCD^{+}93]. Clustering [PNR^{+}93, SCM^{+}93, BVR^{+}93, Fro91, FHUY93, HMC91, HCS94, PMM91, RYKA90, RPS^{+}92, SF90b, YH93, ZHL93]. Clusters [Blo92, BH94, BRO^{+}91a, PBD93]. CNS [WFD92]. CNTF [SLNN^{+}91]. CNTF-mRNA [SLNN^{+}91]. Co [PPFM90, DMJ^{+}94, FAS^{+}93, PH92, STL^{+}90, SOA^{+}93, THBW94]. co-aggregation [THBW94]. Co-assembly [PPFM90]. co-culture [SOA^{+}93]. co-cultures [FAS^{+}93, STL^{+}90]. co-localized [PH92]. co-localizes [DMJ^{+}94]. CoA [DRJ^{+}91, DKG92, HR94a]. coaligns [SG92]. coat [LKF^{+}94]. coated [AU90, AG93, CRWS93, CJC93, DBWS94, FZRH91, FGS91, HSvD91, HPW94, LMSA91, LSC^{+}93, MLR^{+}92, MMH^{+}92, MW94, PDA93, PMBL^{+}92, PTMB94, PMS^{+}91, Rob90, RYK^{+}90, SS91a, SHH94, TBVS92, WRA93, WB92a, WW91, vdBRD^{+}93, HKKO^{+}94, KHKO^{+}93, TOA^{+}93]. coatomer [EOO^{+}94, HKKO^{+}94, KHKO^{+}93]. coats [WE92]. cochaperonin [GMW^{+}94]. cockroach [WFD92]. cod [BSW91]. coded [DAM^{+}93, KHHC93]. codes [PAC^{+}92, VIF^{+}93]. codistribution [BML^{+}90, TLFS92]. coenzyme [ROBN^{+}92]. coexist [WAS^{+}91]. Coexpression [LPM^{+}91]. cofactor [HH94b]. Cofilin [MJLD93]. cognate [IMK^{+}92]. coil [BSP^{+}94, HW90, LCF92, MCS92, NMPN90, WGP^{+}93, WBL^{+}91, YLS92, dCTG92]. coiled [BSP^{+}94, CFPCL92, CFFL93, FCFL94, HW90, LCF92, MCS92, NMPN90, WGP^{+}93, WBL^{+}91, YLS92, dCTG92]. coiled-coil [LCF92, MCS92, NMPN90, YLS92]. colcemid [MAFR93]. colcemid-treated [MAFR93]. colchicine [BSW91, DA92]. cold [BSW91, CCR93]. cold-labile [BSW91]. cold-sensitive [CCR93]. coli [FPGS91, FSSA92, HPH^{+}91, ORS90a, RLB^{+}91]. collaborate [LSR90b]. collaborations [SK91]. collaborative [BSG^{+}92]. Collagen [MS94, BS91, BMWT93, CMS^{+}90a, CLG^{+}92, CETK^{+}94, CJB^{+}93, DS90b, DG90, DSQ^{+}93, DBC90, DTY92, DWJP93, EUH91, FSK^{+}92, FBLL90, GDKP92, GHH90, GH91b, ILH^{+}90, KLM^{+}91, KWGS92, KDS^{+}91, KBT92, KCCG91, LHT^{+}91, LG93a, LGI^{+}93, LMM^{+}91, MGD^{+}92, NCBS90, NDdC92, ORS^{+}90b, PWY^{+}92, PLB^{+}92, PFSS92, SLKS94, SEBH90, SKH90, SJMK93, SABF^{+}90, SGN^{+}93, SDH90, TMP^{+}91, TRV^{+}90, VIF^{+}93, VKR^{+}91, WOC90, YYY^{+}91, LRSB94]. collagen-binding [CJB^{+}93]. collagen-hGH [LRSB94]. collagen-mediated [FSK^{+}92]. collagenase [LWF^{+}91, RSM^{+}94]. collagenous [YYY^{+}91]. collagens [KLM^{+}91, KHP^{+}91b]. collapsing [FMR^{+}93]. collecting [PRB^{+}91, SWS^{+}92]. colligin [CJB^{+}93]. Colocalization [CSJD91, CFPCL92, GRFB94, SBW90]. colocalizes [AKC92, BSF90, PDWB92, ZJA^{+}94]. colocalizing [INY^{+}93]. colon [ACC^{+}94, JSG^{+}92, WTK^{+}92]. colonic [CSR^{+}92]. colonization [CKD90]. colony [LKXS93, WHN^{+}94]. colony-forming [WHN^{+}94]. colony-stimulating [LKXS93]. colorectal [HHH^{+}93, PCG^{+}94, VKRD94]. colR15 [SH91]. colR4 [SH91]. columnar [BR94a]. combination [FKH^{+}93]. combining [BHLA90]. comitin [WMG^{+}93]. Commitment [DP93, Bis90, MWS^{+}93, VBvG^{+}94]. common [CHVF91, EWP^{+}93a, FM94b, FWD^{+}92, LLIV94, SJGG94, TRD^{+}92]. communication [JFA^{+}91, Lam94, LTP^{+}92, ML91, MCEG90, SJIC92, STdL^{+}90, TND^{+}93]. communication-competent [MCEG90]. companion [HR94b]. Comparative [HIJ90, VBV^{+}90]. compared [FHvZ^{+}94, JETS91]. Comparison [TM92, DJA^{+}90, KFW^{+}91, PYA90, RWG90]. comparisons [KGM92]. compartment [AES94, BWCG93, DUVV93, DP90, GEKL^{+}94, HH94a, HSvD91, HWF92, HH90, HEF92, KD92, KR91, KLERG94, KA92, LTPB92, MC92, MPK^{+}94, OTX^{+}94, PLCB91, PDSB92, PMD^{+}93, RHF91, RTC^{+}94, RNS92, RLWK93, SMKH91, SRHK94, SABF^{+}90, SR90b, SDE^{+}93, SMCA91, VHE93, YCBM94]. compartment-specific [YCBM94]. Compartmental [GE91]. Compartmentalization [RR92a, SCE91, EG91, VSB^{+}94]. Compartmentalized [SK94a]. compartmentalizes [NCM^{+}93]. Compartmentation [CP90]. compartments [BHLA90, FYDB93, KLST93, KDO^{+}94, MCM92b, NDM^{+}94, PCP^{+}91, QXWN^{+}94, RHGG92, TIA^{+}92, WHLW90]. compatibility [WBW^{+}94]. compatible [IKFR91]. compete [HBB94]. competence [BMPW90, PC91b]. competent [MBHG91, MOT90, MCEG90]. Complement [GAHB94, MSD92b, WIF^{+}91]. Complementary [NGB^{+}93]. complementation [GFW92]. complements [OCC94, SEC^{+}93]. complete [MCG92, TRS^{+}90, VNS^{+}94, YYY^{+}91]. completing [GHR^{+}90]. completion [CC93b]. Complex [FJRAT91, AR93, ABR^{+}92, BTSL91, BBMN94, BVF^{+}93, BKWD94, BCKS92, BHA^{+}94, BS93b, BGRW91, BSP^{+}94, CNJ^{+}94, CP90, CFG^{+}94, CRK^{+}93, CTPP92, DLMS91, DIDB^{+}91, EC93, EEC^{+}92, ET91, FS94, FMB^{+}91, FH90, GOP^{+}94, GPH^{+}90, GLAB93, GSP^{+}94, GG92, GEKL^{+}94, HG93a, HP92, HXM^{+}94, HWvF90, HNPN94, HWF93, vtHvM90, HAK94, JGL^{+}93, JGK^{+}94, JCSH93, KD92, KLERG94, LTS93, LTPB92, MAA^{+}94, MTS^{+}91, MG92, MSLS91, MR91, NSH90, NSWH90, OESC91, OTWH92, OK92, PA93, PBM^{+}94, PIRB92, PP91, PHB^{+}91, PMS92, PMLM94, PHSvD92, PSSK93, RBW93, RKA93, RLB^{+}91, Ral93, RHB^{+}90, RB93, SWL^{+}92, SRG^{+}94, SSTL93, SP94, SGC^{+}94, SNFN91, SRK^{+}94, SDR^{+}94, SCJ^{+}91, SMBG92, TBKF^{+}92, VHE93, Wat90, WSM93, WRB92, WB93a, WS93b, WCC^{+}90, WKS^{+}93, ZC92, ZMB91, KGR94]. complex-containing [DLMS91]. complex-stimulated [GLAB93]. complexed [EGP^{+}91]. complexes [BNI^{+}94, CJC93, CG93, DPCP91, DS90b, EMC^{+}90, FGH92, Hen90, HKSL90, HIJ90, KCC90, MMD92b, MWH^{+}91, NHS^{+}94, SR90a, SL90b, SKK^{+}91, SY90b, VL91, dSBH93]. complexity [RHG90]. component [BN91, BBS92, BDGC^{+}93, CMM94, EGW^{+}93, EJ93, GSS^{+}91, HNUO92, HLWL92, JLC93, KBM93, KDP^{+}92, MCS92, MJLD93, PH94b, PLB^{+}91, Rob90, RLKB91, SGC^{+}94, SR92a, SCF^{+}93, SRM94, WGM^{+}93, WMG^{+}93, WHC^{+}93, YLDB91, ZAH^{+}94]. Components [FMS93, GOP^{+}94, RK90b, BR94b, BRO^{+}91a, BMGN92, EK92, FRC^{+}91, KKM91, KCYH92, MPM93, ND92, NR90, RHB^{+}90, RPPB94, SBB^{+}91a, SMHO92, SNRS93, SB90, TW93, dBNI^{+}91]. composed [HASW94]. composition [BB90, PRSS90, TSBN94]. comprises [HSvD91]. concentrated [BSB92, HGS^{+}94, SYO^{+}91, SAF90]. concentrates [BD92]. Concentration [DAW^{+}94, MCM92a, YNST93, GSV93, MHM91, SRV^{+}90, SL90b]. Concentration-dependent [MCM92a]. concentrations [BS91, DP93, GDKP92]. concerning [SGP91]. concert [CNSC90]. concomitant [JS90, LFC94]. condensation [GHK94, OHF^{+}94, RTHB90, WE90]. condensed [SNGG94]. conditional [CBC^{+}94, GVK94, RLRR93]. conditions [GB90, SGW^{+}91b, SNG^{+}91, TBT^{+}92]. conductance [BZB^{+}93, DOW92]. conduction [GSV93]. condyles [SCSE90]. cone [AMMW92, HLB90, LB92b, LF93, MAWM90, OB93, PS94, WG93, ZBH94]. cone-mediated [PS94]. cone-target [LF93]. cones [AC93, FMR^{+}93, LMP91, SOE^{+}91, TK91]. confer [RFF92, SH91]. conferred [PTS^{+}92]. confers [FGSBZ93]. configuration [BTS93, RG94]. confluent [PBDK90]. Confocal [MFOI92, LFF91]. conformation [ATD^{+}90, DWY^{+}93, HWvF90]. Conformational [NSH90, PL90]. congression [KBL^{+}94, SSS93]. connected [GHS^{+}93, MTJM93]. connection [KHF94]. connections [DMB93]. Connective [BIPG91, CQYD94]. connexin [JFA^{+}91, MLRJ92, TND^{+}93, WBW^{+}94]. connexin37 [WHD^{+}91]. connexin40 [HSLF^{+}92]. connexin43 [MCEG90, MG91b]. Connexin46 [PET^{+}91]. connexins [FKG94, WBW^{+}94]. consensus [HW92, PSZ^{+}93]. consequence [FTT^{+}93]. consequences [HK91, MCF91]. consequently [vKWHF94]. conservation [JHK^{+}91, KKG^{+}91]. conserved [BMM^{+}94, FMLD92, GSS^{+}91, KMBR90, RZS91, SS92b, TWV94, VTSS94, WPW92a]. consist [BB90]. consisting [BBD^{+}90]. consists [GK90]. constituent [GH91a]. constitute [CHBGL91]. Constitutive [MK91a, AL91, BMH^{+}93, GK92b, IYN^{+}91, KA92, MM91, TGD93, TRH^{+}92]. constitutive-like [KA92]. constitutively [BZB^{+}93, GSC^{+}92, NSC^{+}93]. Consumption [MCM93]. contact [AUT^{+}91, BMTTG94, BBvH^{+}94, Blo92, CKG^{+}90, CKL^{+}91, GPRB93, MH91, SBP90, SB93a, SDCD93, WST^{+}90, WDT^{+}92]. contact-dependent [AUT^{+}91, WST^{+}90, WDT^{+}92]. contact-mediated [CKL^{+}91]. contactin [ZDR^{+}92]. contactin/F11 [ZDR^{+}92]. contacts [CO91, DI91, LRR^{+}92, RH92b, SLPB93, TWC93]. contain [CMZ90, ET91, KRB91, LVA^{+}90, LB92b, MWBD90, PPJF91, SL90a]. contained [MCDY94]. containing [BTSL91, BLR^{+}92, BNI^{+}94, CPSS92, CLD^{+}91, CRK^{+}93, DLMS91, DIDB^{+}91, GCM^{+}94, HSvD93b, HKSL90, HPYB93, IKL^{+}94, KT92, KBL91, MAA^{+}94, MMF94, MNV93, MC92, NSWH90, NRS93b, PDSB92, QXWN^{+}94, RKA93, SWD94, SGVS^{+}93, SNGG94, TSS^{+}92, Tur91, VCL93, WLWL94, YKM^{+}92, YSK90, CFFL93, MM93a]. contains [ATV^{+}93, BSW^{+}92, DEH^{+}91, FPR90, GWO^{+}93, GFL92, GSMK94, HWB93, HKC92, HSM^{+}93, JKLG94, JGB^{+}93, MLKB92, MGS90, NMPN90, OHC^{+}93, PTK^{+}93, PR91, PRL^{+}94, RBW93, RWL^{+}90, SSW94, SHM94, VJB^{+}94]. content [CRKA92, MSW^{+}94, YYR92]. context [NS91]. contiguous [HMS^{+}90]. continues [GHZD90]. continuous [BB91, BCYC94]. continuously [HASW94]. contracted [HTBF94]. Contractile [SDCD93, CW90a, CW90b, DAIS90, HZC93, MLS94, TASJ92, ZSC91, ZC92]. contractilis [FCF92]. contraction [DAIS90, ELZ93, GT93, Gri94, GHH90, GH91b, JKT91, JT93, KJT91, KW92b, KPAY91, LH92, SGN91, WYM^{+}92, WOC90]. contrary [LVK^{+}91]. contrast [SS90a]. Contrasting [ME91, PH94a]. contribute [WOC91]. contributes [BEO90, CKD90, FS94, WS92]. Control [FH90, IPF^{+}90, IVBR93, KFC^{+}94, SBB91b, SM91d, VDS^{+}92, AV91, BMGN92, BBGK90, CHM^{+}92, CSR^{+}92, CF94, DJF93, EBS91, GMLD92, GA91, HH94a, HGDA93, HM91, HYD93, HW93b, HLR92, KDLS92, Koz91, MMG90, MVM90, MWS^{+}93, NDdC92, RGT^{+}93, SMCR90, SMTW94, SGP91, VPP^{+}93, aUFQ^{+}92, dSBH90]. controlled [EG91, JFA^{+}91, KHS^{+}90, MMH^{+}92, RSCS94]. controlling [DCB^{+}92, PDC91, RRD93, TT94]. controls [HRH90, MSW^{+}94, MHG^{+}90, MR91, RYKA90, SWL^{+}92, SM92b]. conventional [ES91]. converges [THL^{+}90]. Conversion [PL90, MR91, SSB93, TISH94]. converted [MC91d]. converts [KYK^{+}94, Low92, WN92]. COOH [AKC92, BSD^{+}92, CCC^{+}94a, CCC^{+}94b, CCW^{+}93, HCKJ92, JJ92, JBJH^{+}92, KPMG91, LPPCF91, Low92, MMS93, MME93, NNHB^{+}92, NWvH94, OWF^{+}93, SRP90, SSM^{+}92a, TWV94, VJB^{+}94, VSB^{+}94, WGD^{+}93]. COOH-terminal [AKC92, BSD^{+}92, CCC^{+}94a, CCC^{+}94b, CCW^{+}93, HCKJ92, JJ92, JBJH^{+}92, KPMG91, MMS93, NNHB^{+}92, NWvH94, SRP90, SSM^{+}92a, TWV94, VSB^{+}94, WGD^{+}93, MME93, OWF^{+}93]. COOH-terminus [Low92]. cooperate [ZB94]. Cooperation [WN92, KKAS90a]. Coordinate [DY93, ISF^{+}92, WWHA91]. Coordinated [TBW92]. coordinately [WBH^{+}90]. coordination [HAS90, PKN91]. COP [GVK94, HKKO^{+}94, KHKO^{+}93, MW94, ODK^{+}93, PPZ^{+}93, TOA^{+}93]. COP-coated [HKKO^{+}94, KHKO^{+}93, TOA^{+}93, MW94]. Copolymerization [BRG^{+}90]. copper [LIAJS94]. copper-binding [LIAJS94]. corbular [JSA^{+}93]. Core [HNP90, BBJ^{+}91, DDW94, FF93, HFL94, HMY92, KT92, KBL91, LSN^{+}94, PKN91, PC91c]. cornea [FBLL90]. corneal [LGI^{+}93]. coronavirus [SM91c]. coronin [dHRB^{+}93]. Correct [FSSA92, OK90]. corrected [GVK94, KYA^{+}94, SBBS90]. correction [CCT94, NW94]. Correctly [HLRBE93]. correlate [BTS93, GB90, WS93b]. correlated [JPJF91, Woo94]. correlates [AB90, AWFP93, BVF^{+}93, BMPW90, KW93, PC91c, QvMWV^{+}91, WSM93]. Correlation [RHB^{+}90, TSR90, BG91a, DPCP91, KWW91, SPHvD91]. Cortactin [WP93]. cortex [JGLK90, KWS94, ORD90, WP93]. Cortical [HW93a, MMM^{+}92, SMCA91, VDTT91, HHL94, HE91b, MAA^{+}94, MGDS93, MJLD93, TS91, YKY92, dCV90, CW90b]. COS [DMB92b, FGH92, GFV^{+}91, JJS92]. Costameres [DIYSS92]. cotransport [PA91, SZP^{+}93]. counter [DSdF^{+}90, OVB^{+}94]. counter-receptor [DSdF^{+}90, OVB^{+}94]. counteradhesive [PAFC92]. couple [EOO^{+}94]. coupled [BS94c, XL91]. Coupling [TWC94, BCYC94, KJT91]. course [KBP91]. covalent [dSBH93]. covalently [HBS92]. CR3 [BLSR92, GAHB94, RUH^{+}94, ZB94]. CRABP [BG91b]. CRAC [IKL^{+}94]. creatine [EERM^{+}91, SGW^{+}91b]. creatine-deficient [EERM^{+}91]. crest [HS92, LBF92]. cripple [Fuc94]. Crithidia [FTPM^{+}94]. critical [EWP^{+}92, GFH90, GIAS93, HLH^{+}92, SY90a, SWFB91, TKP^{+}91, WW94b]. Cross [CPA^{+}93, BGI^{+}90, FF93, HBS92, KHvdL^{+}92, WDWT90]. cross-bridge [WDWT90]. cross-linked [HBS92]. Cross-linking [CPA^{+}93, BGI^{+}90, FF93]. cross-reacting [KHvdL^{+}92]. crossbridges [HFAGM94]. crosslinked [MWSP91]. crosslinking [CG91, DEH^{+}91]. Crowding [MH94]. Crowding-induced [MH94]. crucial [HNvdK^{+}94]. cruzi [CdJC93]. Cryo [RRG^{+}94, AR93, MMM91]. Cryo-electron [RRG^{+}94, AR93, MMM91]. cryoelectron [DWY^{+}93, MWD94, TU90, YDO^{+}90]. crypt [CSR^{+}92, HKS^{+}92]. crypt-villus [CSR^{+}92]. cryptic [SOY93]. crystallin [KAF^{+}93, IITG94]. crystallization [IBD92]. crystals [BBD^{+}90, WBR91]. CS1 [MH93a]. CS1/fibronectin [MH93a]. CSVTCG [TRP^{+}93]. cues [OB93, dlRKR^{+}90, dlR91]. culture [BCET^{+}92, CGMC92, CCT94, DBTAB90, GL91, IBR94, KCDR91, KW92b, LZBH92, LC91, MHR94, MMI^{+}91, MHYK92, PJS^{+}94b, RHP^{+}91, SSG^{+}90, SOA^{+}93, TND^{+}93, THH^{+}90, ZBH94]. cultured [BP93, BRO^{+}91a, BST^{+}90, CW90a, CW90b, CFW93, CPD^{+}93, DH93, DGGG93, ET94, EERM^{+}91, FNV^{+}92, FW93, FA92, FFB90, GMLD92, HdHD^{+}93, KYA^{+}94, KSFG91, KA91b, LDS^{+}92, MBC^{+}91, MW93, MMJ^{+}90, MTP^{+}92, MJM91, MMD^{+}93b, NR90, PPL^{+}93, PCD^{+}92, PSD92, PBD93, RZS^{+}94, SCO^{+}91, SPD^{+}90, SBK^{+}91, SDCD93, SZP^{+}93, SG92, TSP90, TWMS90, TGR^{+}92, VBA90, WLZB93, WGD^{+}93, YYM94, dAWS^{+}90]. cultures [BS91, BGRW91, DJA^{+}90, DOW92, FA90, FAS^{+}93, KG94, PPFM90, STL^{+}90, SHR93, TISH94]. cumulus [LML^{+}94]. currents [PET^{+}91]. cut9 [SY94a]. cutaneous [KBT92, RSB^{+}90]. cuticle [HDC^{+}91, MPR90]. CV [NNHB^{+}92]. CV-1 [NNHB^{+}92]. cyclase [BMV90, IKL^{+}94, JPJF91, SSG93, ZS94]. cycle [AV91, BEGN91, BBZ94, Bis90, Blo93, FGK^{+}90, FAU^{+}93, FFR^{+}91, FHUY93, GN92, GHZD90, GSC^{+}92, GA91, HFN92, HGDA93, HKS^{+}92, HM91, ISN^{+}94, KDKF93, KHP91a, KDN94, LBH92, LR93, MSC92, MG94a, MT94, PPL^{+}93, PH91, ROE^{+}90, SDW^{+}93, SMB^{+}91, SMCR90, SCF^{+}93, STdL^{+}90, TSBN94, UYA^{+}91]. cycle-dependent [BEGN91, FHUY93, GN92, MSC92, PH91]. cycle-maintaining [SDW^{+}93]. cycle-regulated [KDKF93]. cycle-specific [LBH92]. cycles [AES94, HLR92]. Cyclic [BT93, DES^{+}91, HG94, SAH^{+}92, ZZP94]. Cyclin [BCYC94, DFC^{+}92, GN92, BDK92, DIDB^{+}91, LMB^{+}94b, MMCH91, PLB^{+}91, RSW^{+}91, SLW^{+}92, VDS^{+}92]. Cyclin-like [BCYC94]. cyclin-p34cdc2 [SLW^{+}92]. Cycling [GC90, HH94a, PSE^{+}90]. Cyclins [MCG92, HLR92, KMF^{+}91, LR93, PH91]. cycloheximide [GHZD90]. cycloheximide-treated [GHZD90]. cyclophilin [AKC92]. cylicin [HHF93]. cysteine [BIPG91, CPB94, MPR90, PAL^{+}90]. cysteine-rich [BIPG91, CPB94, MPR90]. cysteine-specific [PAL^{+}90]. Cysteine3 [SSDK^{+}94]. cysteines [SVB94]. cystic [DOW92]. cytoadhesive [KBM93]. cytoarchitecture [SLG92, TFC94]. cytochalasin [DA92, OTT92, PT92]. cytochrome [HSHA90, MC91a, MC93, PAC^{+}92]. Cytokeratin [KMN91, MSF90, OBC^{+}92]. cytokeratins [BMF^{+}91, BRF93]. Cytokine [WHN^{+}94, KHvdL^{+}92, KHHR92]. cytokine-activated [KHvdL^{+}92]. Cytokine-mediated [WHN^{+}94]. cytokine-prestimulated [KHHR92]. Cytokines [NS91]. cytokinesis [BHBG94, COTC94, FCW91, FMS93, FMJ91, FFB90, OCC94, SYO^{+}91, SLW^{+}92, SKBD91, dHRB^{+}93]. cytological [BWLK93]. cytomechanical [LZBH92]. cytomechanics [Ada92]. cytometric [SCS90]. cytopathogenicity [SZB^{+}94b]. cytoplasm [BMF^{+}91, BDMPJ94, EH94, FV91, GHT92, HNH91, KAV93, TSP90, WHW^{+}91]. Cytoplasmic [AEGG93, BHF90, BRO^{+}91a, CTPP92, FWF90, LSK^{+}94, LFC94, PMTB94, SEAB91, VKM93, WRP^{+}91, YPH^{+}92, eSP91, AMG90, ACC^{+}94, BBK^{+}93, BLR^{+}94, BVS^{+}93, BHD^{+}94, BR91, BSP^{+}94, CL93, CLK^{+}91, CB92, DCY^{+}94, DP94, GMW^{+}94, GJJM94, GSS^{+}91, GC90, GSMK94, HHR^{+}90, HW93a, HPYB93, JT93, JSM^{+}90, JLL^{+}94, JK92, KSK^{+}93, KLST93, KGM92, LSP^{+}91, LMSH94, LB93, MY92, MPAF91, MMH^{+}92, MHGJ94, MKF91, NNHB^{+}92, NI93, NRS93b, OKF^{+}94, OTX^{+}94, OHC^{+}93, PMPV92, PKN91, PH94a, RHH92, SFA90, Sch94, SM93b, VBAK91, VM93, VM94, WKS^{+}93, WF90, WF94, YCO^{+}93, YNST93, ZJA^{+}94, dAWL^{+}90]. cytoplasmically [FB93, vZDF^{+}93]. Cytoskeletal [BBGK90, LF93, ATV^{+}93, BFM93, BRF93, BTR92, CKP91, CSCSA94, DAIS90, Fow90, GWO^{+}93, HCS94, KPAY91, MH94, MB92, MKB^{+}94, NSWH90, NHA^{+}92, PMW92, RHH92, SHLS93, SBWV90, SMM90, TFT94, TS93, WS91b, dHRB^{+}93]. cytoskeleton [ATV^{+}93, BTSL91, CPSS92, CSG92, CPB94, EK92, FDS90, GXHS94, GORB91, HD91a, HHF93, HYD93, HBB94, JBJH^{+}92, KKES91, MBC^{+}91, MNME^{+}93, MSF90, MJLD93, MPM^{+}94, NHO^{+}91, NI93, NKW^{+}92, PIRB92, PB91, PvBeHB^{+}91, PMHS92, PALM92, RLB^{+}91, SCMB92, SRM94, TDT92, TDWT92, YFS92, YKY92]. cytoskeleton-related [NHO^{+}91]. cytoskeletons [BB93, KPAY91, TOS^{+}94, YNST93]. Cytosol [SCS92, WS93a, BPM93, CB90, FA92, HSvD93a, HTD90, PMS^{+}91, SMM^{+}91, WHLW90]. Cytosol- [SCS92]. Cytosol-dependent [WS93a]. Cytosolic [JLC^{+}90, PSSK93, AA94, BHFS93, GP91, HM92b, HMB^{+}93, Hur90, IKL^{+}94, JTS^{+}91, JKB^{+}92, JCSH93, KKG^{+}90, LMW^{+}94, MM90, MHM91, MHG^{+}90, NF90, PSE^{+}90, RSC^{+}93, SSSL94, SMBG92, WC90a, WGTCY94]. cytosolic-free [JTS^{+}91]. cytostatic [DPW91, FMLD92]. cytotactin [PJC^{+}90, PAFC92]. cytotoxic [HPM^{+}94]. cytotrophoblasts [LWF^{+}91].

D [BPDM90, DA92, FF93, HWW^{+}93, LKXS93, PT92, ZFF92]. D. [BMTTG94, dSK90]. D1 [LMB^{+}94b]. D3 [BPDM90]. D3-dependent [BPDM90]. DAG [GPF94]. damaging [WW94a]. Darby [BBGK90, NSWH90]. Daughter [KAG94, FCFL94]. day [GBOB90]. Death [RLD^{+}91, BG91a, BG93, DJ93a, EZF^{+}94, GSBS92, MWG92, Ree94, SBG93, ZZL^{+}91]. decapentaplegic [WBWH93]. decarboxylase [SVB94]. decondensed [LSAH90]. decorin [HK91]. decrease [OH90, WGCdL91]. Decreased [LG93a]. dedifferentiated [QDT^{+}90]. deduced [FPR90]. deep [FKH^{+}93, NSH90]. deep-etch [FKH^{+}93, NSH90]. default [DUVV93, RNS92]. defect [CCR93, LBH92, OMRM93]. defective [BEO90, CSC90, COTC94, FDS90, HSUS90, KCH^{+}94, MSTY90, MPAF91, RRCR93, SEC^{+}93, TYF90, WS90, WC90b, dHRB^{+}93, MR94]. defects [BCLM94, BHA^{+}94, CMV^{+}92, CBC^{+}94, CBM^{+}94, GdLMS94, HWL^{+}90, LCD^{+}93, OCC94, RHR90, SHOA92]. defense [SMHO92, ZLG^{+}90]. defensins [SMHO92]. deferens [VPP^{+}93]. deficiencies [MMS93, MHDT94]. deficiency [OGM^{+}93, PBB^{+}91]. deficient [BSG^{+}92, DGF^{+}94, EERM^{+}91, HLA^{+}90, JH90, KB91b, MBW^{+}94, MCEG90, SGW91a, TBT^{+}92, VFR94]. define [MB92, NGB^{+}93, SMRG91]. defined [BR94a, GE91, MY92, OHS92, VWS90]. defines [BKSB94, CSC92, PMM91, SLK93]. Defining [NHS^{+}94, WGBB91]. definition [ES91]. Deflagellation [YC93, QH94]. degeneracy [BSD^{+}92]. degenerating [MHP90]. Degeneration [MBW^{+}94, AMMW92, KHTD92, PBB^{+}91, VGM^{+}93]. degradation [BCYC94, CNU92, DKL^{+}90, EMC^{+}90, GSMK94, HR94a, HLR92, LB92a, MCG92, NFYI93, OHF^{+}94, OTWH92, QvMWV^{+}91, RSC^{+}93, ROBN^{+}92, SWHCW91, SB91, WL91, WCC^{+}90, WKY^{+}93, YL93]. degradative [KP90]. degraded [TIA^{+}92, VGMS91]. degrading [KWS94, TBW92]. Degranulation [HBP^{+}93]. degree [AGB^{+}91, RHG90]. degrees [JK93]. dehydroepiandrosterone [ASG^{+}93]. dehydrogenase [MYT94]. delay [SN93]. delayed [SPB90]. deleted [PCG^{+}94, RPE^{+}90, VKRD94]. Deletion [CdJC93, JO92, JJS92, LSP^{+}91, MGD^{+}92, OMRM93, SSFD93]. Deletions [CCAF90, DCY^{+}94, CB92, SRP90]. delineation [HAA93]. delivered [CVH91, HSvD93b, TLXM90]. Delivery [DM92, BR91, HGB^{+}91, MMF92, NLFRB91, PAL^{+}90, WNBAS90]. delta [CH92, KM92d]. demonstrated [TBT^{+}92, XL91]. demonstrates [WSL91]. Demonstration [TS91, GA92, HKC92, ITSR90]. demyelination [AFHDD90]. demyelinative [KYA^{+}94]. denaturation [LFR93]. denatured [BWV92]. dendrites [MMD^{+}93b]. dendritic [HdHD^{+}93, PSD92]. denervated [CQYD94, LC94]. dense [LSN^{+}94, MOT90]. density [ASG^{+}93, CSC90, DDFC94, GK92a, GRC^{+}91, HPM^{+}94, KT92, MHGC90, MCF93, MTJM93, SAWL90, SF90b]. deoxycytidine [BCG^{+}92]. Deoxyglucose [PT92]. depend [CPB^{+}93, Smi94]. dependence [FRL93, GAS90b, HAK93, VGG^{+}93]. dependency [AM94]. dependent [AEGG93, AUT^{+}91, BEGN91, BAP^{+}90, BPDM90, BSM94, BGK^{+}94, BFLS94, BSW91, BMH^{+}93, CFPCL92, CML90, CF94, DAS^{+}92, DAMS91, DMC^{+}91, FGSBZ93, FKS93, FHUY93, FJN93, FJN94, GN92, GFWM92, GAvdM91, GLAB93, GSP^{+}94, GPRB93, HM92b, HDS90, JPJF91, KKAS90a, KKAS90b, KM92b, KISY91, KCR^{+}94, KCYH92, LGSB93, LHS^{+}92, MCM92a, Mah91, MSC92, MSQ^{+}90, MSPB90, NS93, NHA^{+}92, NFYI93, OR92, OHP^{+}94, PIRB92, PDB^{+}91, PBCK93, PH91, PDC91, RR92a, RZS^{+}94, RGT^{+}93, RBD92, RM94, SWD92, SWL^{+}92, SJIC92, SGW91a, SMM90, SDR^{+}90, SSCD90, SCS92, SSKT^{+}92, SBL^{+}91, TCDH93, TSH^{+}93, TLSW93, TIA^{+}92, VHM^{+}93, VDS^{+}92, VR92, WK92, WHT^{+}93, WFD^{+}91, WS93a, WST^{+}90, WDT^{+}92, WWD94, ZS94, ZZP94, dWLS91, vKWH^{+}91]. depending [GC90]. depends [ABM^{+}93, AGB^{+}91, HS90a, KM92a, MY92, MA90, MJM91, MDKB94, SRHK94, WRHW92]. dephosphorylation [BS94b, DH93, EG91, HHY^{+}93]. depleted [AWH^{+}94, TGS92, XL91]. Depletion [AC93, AG93, MPTW90, HSvD93a, PMA91]. depolymerization [CLZ91, CLM91, ZFF92]. depolymerizing [MLM^{+}93]. Deposition [ORD90]. deprivation [BG91a, RLD^{+}91]. deprived [DJ93a]. derivatized [WS91a, WS91b]. derived [AW92, ASNH93, CHBGL91, CH93, DS90a, DKM^{+}92, FMR^{+}93, FB93, GSFS92, HE93, KBCM^{+}90, KS92, KVF^{+}90, KDP^{+}92, MMW^{+}92, RLS94, RUH^{+}94, SCO^{+}91, SWL90, TVG^{+}93, TRD^{+}92, YRBBP90, ZPS^{+}92, RS91, TOA^{+}93]. dermal [KS94, MLKB92, MRH^{+}91, STF93]. dermal-epidermal [MLKB92]. dermis [ZDZSBT94]. Descemet [FBLL90, SKH90]. desensitization [LASB^{+}93]. desmin [BI90, CK94, LCM^{+}94, RPE^{+}90, SLI^{+}91]. Desmin/ [SLI^{+}91]. desmocollin [LYC^{+}94]. desmocollins [CLK^{+}91]. desmoglein [KAP^{+}93, TTE^{+}94]. desmoplakin [SG92, SBC^{+}93]. desmosomal [BCF^{+}93, CLK^{+}91, KHF94, WRDK92]. desmosome [OS92, PKN91]. desmosomes [CLCC90, KAP^{+}93, SJGG94]. destabilization [FCF92, ZSC91]. destructible [GN92]. destruction [HLR92, KMF^{+}91]. detect [SBR91]. detectable [GHZD90, LD91, SW90]. detected [DAMS91]. Detection [KJ90, SFM^{+}90, SR90b, BTBO94, DBV^{+}92, DMLR^{+}94]. detergent [HNR90, NI93, TCCT91]. detergent-insoluble [NI93]. detergent-solubilized [HNR90]. determinant [GSMK94, MMH^{+}92, WBW^{+}94]. Determinants [CL91b, KAV93, MHGJ94, BZF^{+}92]. Determination [CHW94, KT93, MWD94, KW93, WHN^{+}94]. determine [DVT92, FRSL90, OTX^{+}94]. determined [MAFR93]. determines [BM91b, CC90a, CSvdE93, HWM^{+}93, HSG94, LL91, SAH^{+}92, SSDK^{+}94]. detyrosinated [SBW90]. Detyrosination [WWWB90]. developing [DR91, FB93, FSY^{+}91, Gau90, GDB^{+}94, GJK^{+}91, KYLV90, LOC^{+}90, MTP^{+}92, MMT91, ME91, SHR93, SABF^{+}90, SSA^{+}90, SLL94, TKHM91, YAJ90, YAJ91, dlRKR^{+}90, dlR91]. Development [KCDR91, TM90, AWH^{+}94, BBK90, BLB^{+}91, COTC94, CWB93, CWHH92, DFLK94, FJRAT91, FRL93, FKB^{+}90a, FWF90, GFH90, GB90, JP90, KL90, KGK90, KHaA92, LSR90b, LW92, LWB91, Mah91, NBT94, OA92, OLKD90, PZV^{+}91, PTTA90, PSJ^{+}91, PJC^{+}90, QDT^{+}90, RSM^{+}94, RRH91, SMRG91, SA90, SMWB93, SEMH92, STdL^{+}90, SVD90, SAM^{+}91, Tur91, WC90a, WW94b, WKN92]. Developmental [DBV^{+}92, KVF^{+}90, LSS^{+}90, WMF91, MS94, OCC94, PJS^{+}92, SWD94, SKBD91, SLNN^{+}91, WFD92, WWHA91]. Developmentally [YOWM91, BBD^{+}90, CKB93, CPB94, DAM^{+}93, GF94, SON^{+}94, SWF^{+}92, SR92d, WDRM93]. diabetes [JP94]. diacylglycerol [BGMM^{+}93, HRT^{+}91, PvBeHB^{+}91]. diameter [ASR^{+}90, Woo94]. diameters [ASR^{+}90]. diaphragms [SAF90]. diatom [HSSC92]. dictate [RS94]. Dictyostelium [BCLM94, BBD^{+}90, BGI^{+}90, COTC94, CBC^{+}94, EVW91, FF93, FDS90, FWF90, Ham94, HZC93, HHL94, IKL^{+}94, JH90, KGM92, KPAY91, KIV93, MC91b, ORS90a, OA92, OCC94, RCNO94, SL90a, SVD90, WS90, WE92, YKY92, ZC92, ZFF92, dHRB^{+}93]. differ [BB90, BST^{+}90, CMZ90, KHW93, LB92b]. Differences [BMGC91, ES91, KHV^{+}93, AFA^{+}94, CK94, FTPM^{+}94, FHvZ^{+}94, TGR^{+}92]. Different [FvBuHWS92, MKF91, BJ90, CPB^{+}93, CHVF91, DPS^{+}93, DEH^{+}91, HR90a, HK91, HFS92, JBIM94, KKM91, KA91a, KLST93, KEBD91, LZBH92, LSG92, MNME^{+}93, MMJ^{+}90, MMW^{+}92, MC91b, PA91, PRC^{+}90, RLS92, RA94, SM91a, SMKH91, SNG^{+}91, SLM^{+}90, SAM^{+}91, TLXM90, VVB^{+}92, WMB^{+}94, ZPS^{+}92]. Differential [BBK92, CK94, DvHB^{+}91, DP94, DSJB^{+}94, EK90, FCFL94, FFB^{+}93, Gau90, GKG90, GR93, KG94, KKW^{+}93, ML91, MEM94, MHDT94, MCEG90, OH92, PLB^{+}92, PW93, PHWM92, SKSC94, SLHG93, SLL^{+}90, TOFH94, UKH^{+}91, BKSB94, GHS^{+}93, HWW^{+}93, KA91a, MMD^{+}93b, RSC^{+}93, SMRG91, SS90a, VPC^{+}91, WBE91, WS92]. differentially [BGLB94, EUH91, FTB94, GTC^{+}90, HG93b, KMSW91, PH91, SCO^{+}91, SLG92, TSBS92, WOC91, WWD94]. differentiated [CHBGL91, WBAP91, ZSC91]. differentiating [AW91, CT91, MPR90]. Differentiation [BB91, PRC^{+}90, AWFP93, ACFFL93, AGGS92, ASNH93, BZF^{+}92, BVF^{+}93, BF93, BG91b, CGMC92, CJC93, CMS^{+}90b, CCT94, CLS^{+}92, DBTAB90, DP93, DS90c, EFD^{+}93, FPR90, FCD^{+}93, Fuc90, GBB^{+}94, GBP^{+}92, GBN^{+}93, GFH90, HR90a, HKS^{+}92, HW94, JRCW^{+}90, KRHY93, KYK^{+}94, KS92, KCDR91, LA92, LYC^{+}94, LBL^{+}91, LKXS93, LCT^{+}93, LNS^{+}94, LVTB93, MG94a, MPIDD93, MLS94, MJM91, MDRG92, MWS^{+}93, MMD91, NBT94, NCBS90, OR92, PWY^{+}92, PC91c, PMM90, QPWG91, QCCD92, RWG90, RHG90, RRBG91, SGW91a, SBN92, SSFD93, SLL94, SCSE90, SBB91b, TG94, TNT^{+}90, VGRC^{+}92, VKRD94, WHT^{+}93, WG92b, WBH^{+}90, YRBBP90, YKI^{+}91, YSK^{+}94, YLWS94, vdSKL^{+}93]. differentiation-related [LVTB93]. differentiation-specific [BG91b, KRHY93]. differently [NWvH94, VPP^{+}93]. differing [BGGW^{+}94]. diffusible [BS93a]. diffusion [FMR90, LZI^{+}93, NCM^{+}93, VBA90, ZKMB93, ZCS^{+}91, dBNI^{+}91]. digestion [TCCT91]. digestive [CKGC91]. digital [TLFS92]. digitonin [PDSB92]. digitonin-permeabilized [PDSB92]. Dihydrofolate [LBL^{+}91]. Dihydropyridine [FPP91, YAJ91]. dihydroxybenzoate [NCBS90]. dihydroxyvitamin [BPDM90]. Dilution [WPS91]. dimensional [AR93, DMM93, FW93, FPGS91, HASW94, KDS^{+}91, Lut91, LFF91, MNS90, PBM^{+}94, RRG^{+}94, SAJ^{+}94, TU90, VBV^{+}90, VCL93, VAL^{+}93, WBR91, YDO^{+}90]. dimensions [SLK91]. dimerization [VFC^{+}91]. dimerizing [YKM^{+}92]. dimers [WDB^{+}92]. diminished [AC93]. diminution [GGGE^{+}92]. dimorphic [LNS^{+}94]. dinoflagellate [SGGC^{+}90]. dipeptidyl [HD90, JZAVP93, NRS93b]. diphosphatase [AYM^{+}93]. diphtheria [MIU^{+}92, SMM^{+}91]. diploid [AUT^{+}91]. Direct [Bea91, BR94b, FIH^{+}94, ITSR90, KIN^{+}94, OTT92, SWD92, SBLV92, ATSE90, Bro91b, Car91, HBR90, KHF94, SAH^{+}92]. directed [BS94b, BL94, BSR^{+}90, ESL^{+}90, GKKS90, KYBC94, Knu93, LB92c, ST94, DG90]. direction [ABR^{+}92, SAH^{+}92]. Directional [SSS93, VMB92, KM94]. directly [ASE90, JPJF91, WHF^{+}90]. disappearance [KD92]. disassembles [DIDB^{+}91, HW92]. disassembly [ABR^{+}92, DA92, EBS91, GeKdV^{+}91, VDTT91, WNM^{+}94]. disc [FKB^{+}90b]. discoideum [BCLM94, BMTTG94, BBD^{+}90, CBC^{+}94, EVW91, FWF90, KIV93, MC91b, RCNO94, SL90a, YKY92, ZC92, dSK90]. discovery [SCE91]. Discrete [CTL91, WW93, SYG91]. discs [JG94]. discs-large [JG94]. disease [Fuc94, GK93a, Kos94, LSV^{+}94]. diseases [CHVF91, LWHL93]. disk [AMMW92]. dismutase [PSE^{+}90]. disorder [BMS^{+}91]. disorders [MHDT94]. dispensable [NMPN90, SLI^{+}91]. dispersed [WNTT94]. displacement [GXHS94]. display [EERM^{+}91, JWHPM93, KAG94, KWR94, LSAH90, LBH92, NFL90]. displayed [VSVP93a]. displays [CCR93, LVK^{+}91]. disposition [POGS90]. disrupt [FKB^{+}90b]. disrupted [LB92c]. Disruption [FF94, OMS^{+}94, PB91, AW92, COTC94, CCW^{+}92, DH93, HPM^{+}94]. Disruptions [GVK94]. disruptive [LCF92]. disrupts [AG93, BBRE90, BH94, CdJC93, MPTW90, SG92, WGTCY94, YH93]. dissassembly [FCF92]. dissection [HD90, MRSN93, SY92, TOK93]. dissociated [DBTAB90, Hen90]. Dissociation [DLSB^{+}90, WS93b, BSR^{+}90, Cit92, CSM90, GHS^{+}93, PNPB94]. Dissolution [MH94]. distal [BSB92, DA92, MCM92b, SOY93]. distant [AB91, NYA^{+}91]. distensions [DAW^{+}94]. Distinct [ARW^{+}94, BP91, CKG^{+}90, HR90a, KB91b, KA94, LMB^{+}94a, PM93, RS91, SBH92, WNBAS90, YCO^{+}93, BRG^{+}90, BBZ94, BHFS93, CHW94, DPS^{+}93, DGAB^{+}93, ECO93, GWO^{+}93, HRC^{+}90, HSvD91, Has93, HM92b, HPR93, HPW94, JWHPM93, KWS94, KFN93, KW92a, KYLV90, KBL^{+}94, LS90, LSRC93, LWK^{+}93, LJHC94, LHFV93, MH93a, MSI^{+}94, MGS90, NFL90, NGB^{+}93, PRSS90, PR91, PDWB92, QXWN^{+}94, QH94, RHB^{+}90, SMRG91, SP94, SH90b, SA90, TASJ92, VHSR92, VJB^{+}94, VL91, WGBB91, WB92a, ZSM^{+}93]. distinctive [SWMS90]. distinctly [AVP92]. distinguishable [MG92]. distinguished [AL91]. distinguishes [BT93, CP90]. Distinguishing [MNME^{+}93]. distribute [WOC91]. distributed [BK91, FTB94, PRF^{+}93, PMM90, ZJA^{+}94]. Distribution [FMR^{+}92, GBOB90, LGH91, AGKC92, BJ92, BKSB94, BBZ94, BZT93, BKW91, CLZ91, CSvdE93, CJO92, CWHH92, CGF^{+}93, DvHB^{+}91, DVT92, EVW91, FPP91, FPGS91, Gau90, GB90, HG93a, HAN^{+}93, JG94, JWHPM93, KM92a, KWK92, KB90, LAY92, LCMP90, LC91, MSI^{+}94, MS94, NT90, NPH^{+}93, PRF^{+}93, PSN^{+}91, PMTB94, RH92a, RHB^{+}90, RK94, RAGC94, SLL^{+}90, SWL90, SDK^{+}93, SZP^{+}93, SAWL90, SMJ^{+}94, SEMH92, SLNN^{+}91, TH92b, TGR^{+}92, VHM^{+}93, VPC^{+}91, WBA92, WKM^{+}92, WKN92, XL91, ZPS^{+}92, dTF90, vKWHF94]. distributions [CMZ90, KHV^{+}93, MNME^{+}93, NHS^{+}94, NGB^{+}93]. Disulfide [SBSG92, OPL94, POGS90, VFC^{+}91, YDS92]. disulfide-bonded [OPL94]. disulfide-linked [VFC^{+}91]. Disulfides [SSM92b]. Divalent [DCCH92, EUH91, MH93a]. divergence [CC90b]. divergent [SNGG94]. diverse [AO90, GBOB90]. diversion [NF92]. Diversity [FDT^{+}91]. dividing [ABD^{+}94, CW90a, CW90b, FA90, FCW91, FW93, WN92]. division [FRSL90, GC90, GC91, HHY^{+}93, KSK^{+}93, OSJS94, RK94, RBD92, SCF^{+}93, SLK93, WSC94, WTS93]. divisions [LA92]. DMDL [NEL^{+}91]. DNA [AL92, ASR^{+}90, AS94, BG93, Blo93, CBJER93, CDRL93, DMWJ90, FTPM^{+}94, GSBS92, HODF94, HSCN94, HM91, ISN^{+}94, KT93, LHS^{+}93, LL91, LM94, LMB^{+}94b, MMF94, MMLG94, MMY^{+}92, NDdC92, OHS92, PME93, RDDR93, RG94, RMFA92, SOWDK94, SM92a, SY92, SLS^{+}90, SGGC^{+}90, TN94, WMF91, YKM^{+}92]. dnaJ [CD91, LSA91]. Do [LCF92, OH93, LVA^{+}90, MCG92, RFF92]. docking [MHG93]. does [BMTTG94, BMM^{+}94, DMWJ90, FFB90, FMLD92, HM92a, HKS^{+}92, HM93, KES90, LHH^{+}91, MMCH91, MGW92, NI93, Smi94, WWWB90, ZMV^{+}93]. dog [VSVP93a]. Domain [ENS91, LS92, ACC^{+}94, ATBD90, AM91, AJPB90, BBK^{+}93, BVS^{+}93, BI90, BHD^{+}94, BFL90, BR91, CKB93, CAS^{+}92, CB92, CCC^{+}94a, CCC^{+}94b, CPP^{+}94, DCY^{+}94, DGAB^{+}93, DAR93, EBS91, FL92, FV91, GJJM94, GF94, GSMK94, HFL94, HWB93, HW92, HHR^{+}90, HLH^{+}92, HWM^{+}93, HKC92, HD91b, HSM^{+}93, IKL^{+}94, JJ92, JO92, JJB^{+}93, JSM^{+}90, JLL^{+}94, JBJH^{+}92, JGO^{+}91, KSL^{+}94, KBL91, KPMG91, LBH93, LMH^{+}94, LHT^{+}91, LSK^{+}94, LGI^{+}93, LB93, LFB94, LTWH92, MCF91, MWH^{+}91, MKB^{+}94, MGD^{+}92, MMH^{+}92, MFS^{+}94, MULA^{+}91, NYA^{+}91, NMPN90, NNHB^{+}92, NCM^{+}93, NRS^{+}93a, OWF^{+}93, PKN91, PHSA93, PDA93, PNB91, PKG^{+}94, RHH92, RUH^{+}94, RGT^{+}93, ROBN^{+}92, RWL^{+}90, RHB^{+}93, SCMB92, SMSH92, SCM^{+}93, SM93b, SLWF91, SW93b, SSM^{+}92a, SHM94, SM91c, TTE^{+}94, TRP^{+}93, VLH^{+}93, WHF^{+}90, WBW^{+}94, WBL^{+}91, WGD^{+}93, WLH92, WF94, WSSM91]. domain [WB92c, WBR94, XMW^{+}93, YYY^{+}91, YKM^{+}92, YNST93, YLWS94, ZBW93, CPB94, LWK^{+}93]. domain-binding [JLL^{+}94]. Domain-specific [LS92]. Domains [CLG^{+}92, ATV^{+}93, BB90, BLR^{+}94, BMM^{+}94, BPS90, CPB^{+}93, CTL91, CK94, CSvdE93, CLK^{+}91, DBC90, ES91, FFKL92, FvBuHWS92, GWO^{+}93, GAvdM91, HE91a, HWM^{+}93, HFS92, JO92, KT92, KSST91, LGSB93, LSV^{+}94, LJHC94, MECE93, MNV93, NPST94, NGB^{+}93, NTB90, ORS90a, OKF^{+}94, OVB^{+}94, PH94a, PLB^{+}92, POGS90, PKPM90, PMM91, PTMB94, PDWB92, PM93, PE93, PAFC92, SY90a, SRHK94, SBK^{+}91, SYG91, SEAB91, SBC^{+}93, TVG^{+}93, VHSR92, VJB^{+}94, VRE^{+}92, VTSS94, WNBAS90, WSvdK^{+}93, WLWL94, WIF^{+}91, WPW92b, WCF92, WHK^{+}93, WEB90, XMW^{+}93, YYY^{+}91, YCO^{+}93, ZDR^{+}92, dAWL^{+}90]. dominance [HW93b]. Dominant [ELS93, BRA^{+}94, GWMC90, KRB91, WC90b, ZRW^{+}94]. donor [BFR^{+}92]. dorsal [WS93b]. dorsal-cactus [WS93b]. double [CCC^{+}94a, CCC^{+}94b, ZFA^{+}93]. down [KHHC93]. down-regulated [KHHC93]. downregulate [ZSC91]. downregulated [Mac92, MULA^{+}91, VBV^{+}90]. Downregulation [EFTJ^{+}93, HW94, LTP^{+}92, SGVS^{+}93]. downstream [BMGN92]. DPH [IK94]. dramatic [DWY^{+}93, Lam94]. DRAP27 [MIU^{+}92, MIU^{+}92]. drastic [BF93]. drive [PLB^{+}91]. droplet [OHC^{+}93]. Drosophila [CKMC94, EHB^{+}90, LMSH94, RKA93, AHW90, AS94, BTS93, BF91, BOH^{+}92, BBB91, CBM^{+}94, DBSK90, FJRAT91, FKB^{+}90b, HE91a, HS90b, HPP^{+}93a, HAS90, HDP^{+}93, HR90b, IFLG90, JG94, KM92c, KGK90, LCD^{+}93, MCG92, MMD92a, MMD92b, MMF94, OUS^{+}93, PTTA90, PMG^{+}92, PDTG92, PSG94, PHWM92, PM93, PMW92, RFF92, SSSL94, SRM94, SJM^{+}94, TH92a, UHFG92, VSVP93b, VGM^{+}93, WHM^{+}92, WYM^{+}92, WMF91, WKMG92, WA93, dCTG92, vDS94]. drugs [PSN^{+}91]. Dual [KII^{+}91, BHLA90, GVK^{+}93, KDN94, LCPAM94, SKK^{+}91]. dual-function [BHLA90]. Dual-view [KII^{+}91]. duct [SWS^{+}92]. due [CK94]. duodenal [CSR^{+}92]. duodenal-colonic [CSR^{+}92]. duplication [GHZD90, KB90, WGBB91, WHC^{+}93]. duration [HFN92]. during [AR91, AG93, ACFFL93, AFHDD90, ACSM^{+}94, APP^{+}93, BRG^{+}90, BGBWO91, BFR^{+}92, BKB^{+}90, BBK90, BSF90, BLB^{+}91, BDW92, BS94c, BPH^{+}90, BG94, BGH93, BBGK90, CKMC94, CSCSA94, CJC93, CS91, CHM^{+}92, CWB93, CGF^{+}93, CHW94, CHG94, CC90b, Dav92, DVT92, EP94, EZF^{+}94, ES92, FMD93, FGK^{+}90, FJRAT91, FCFL94, FKS93, FA92, FRL93, FCD^{+}93, FWF90, FFS^{+}93, GCZ^{+}92, GBN^{+}93, GDB^{+}94, GFH90, GB90, GT93, GC90, GC91, GeKdV^{+}91, HKKO^{+}94, HSS93, HS92, HBR90, HS90b, HGW^{+}91, HAG^{+}93, HAS90, HDP^{+}93, HCYK91, HW94, HSSC92, HNH91, HHY^{+}93, HLCB92, JLC^{+}90, JGK^{+}94, JRCW^{+}90, JS92, KL90, KHS^{+}90, KGK90, KMF^{+}91, KTG90, KKW^{+}93, KA93, LWL^{+}92, LFR93, LELB90, LF93, LW92, LWB91, LK93b, LJP^{+}93, Mac92, MECE93, Mah91, MHG93, MRBP^{+}90, MT94, MBS91, MLS94, MS92, MMD91, NBT94]. during [NGP^{+}91, NHTP90, ND92, OHF^{+}94, OH90, PZV^{+}91, PLS92, PTTA90, PSJ^{+}91, PNB91, PRC^{+}90, PJC^{+}90, Ral93, ROE^{+}90, RMK91, RK94, RSM^{+}94, RA90, RGKG90, RRD93, SOE^{+}91, SGVS^{+}93, SMRG91, SS94, SYO^{+}91, SOA^{+}93, SLB^{+}94, SBK94, SBSG92, SST92b, SM92a, SSS93, SA90, SWF^{+}92, SMWB93, SPW93, STH^{+}93, SEMH92, SLL94, SCSE90, STdL^{+}90, SSB^{+}91, SVD90, SAM^{+}91, TBW92, TK91, TTVV^{+}90, TW93, TIA93, TNT^{+}90, Tur91, UNY90, VGRC^{+}92, VCFM94, VKRD94, VDTT91, WSC94, WCR93, WRHW92, WBE91, WIS^{+}94, WBH^{+}90, WKN92, YYM94, YSBM91, ZKMB93, ZBS^{+}94]. DVR [WBWH93]. DVR-6 [WBWH93]. dye [eSP91]. dyes [ZBS^{+}94]. Dynactin [GSS^{+}91, SGC^{+}94]. Dynamic [BTS93, GCML94, GRFB94, MMLG94, OHS92, PTMB94, UNY90, FZRH91, GPS92, HTGN94, LJJC93, PS94, SHLS93]. Dynamics [HNPN94, LFG^{+}90, MKG93, OMH93, RH92b, SBK94, UOKH93, ABD^{+}94, BPH^{+}90, CFW93, DA92, ISN^{+}94, KRCT93, KJT91, MMM91, RLGB94, SW93a, SS90a, SSS^{+}90, TOFH94, TWQ^{+}93, TM92, VDS^{+}92, VLGB92]. dynamin [DBWS94, HBOV93, SM90, VRE^{+}92, vdBRD^{+}93]. Dynein [KGM92, AEGG93, BHS94, CTPP92, DP94, GOP^{+}94, GSS^{+}91, HSYYK90, HHG^{+}94, KKM91, KIO^{+}94, LMSH94, LFC94, MOM^{+}92, MK91b, MSFW92, MGW92, PMPV92, PRSS90, PR91, PMS92, PMLM94, PMTB94, PPD92, PKG^{+}94, SMK91, STK^{+}93, SFA90, Sch94, SFO^{+}91, SS92a, TK94, VKM93, VMB92, VBAK91]. dynein-dependent [AEGG93]. dyneins [PMLM94]. dysfunction [EB92]. dysgenic [FPP91, FAF^{+}93]. dystrophic [NEL^{+}91, TFDS91]. Dystrophin [OC91, OESC91, PDWB92, YLDB91, BKW91, DBSK90, EC93, HKC92, HLCB92, KWK92, MFOI92, NEL^{+}91, NGB^{+}93, PNR^{+}93, SBK^{+}91, SSM^{+}92a, SBLV92]. Dystrophin-associated [OC91]. Dystrophin-glycoprotein [OESC91, EC93]. dystrophin-related [KWK92, NEL^{+}91, PNR^{+}93].

E- [LLIV94]. E-cadherin [HBB94, NSWH90, BCB93, CO91, JFA^{+}91, KW92a, NGP^{+}91, WJ92, BVF^{+}93]. E-cadherin-alpha [NIT94]. E-cadherin-catenin [WNTT94]. E-cadherin-mediated [FBS^{+}91, WHH^{+}93, KWW^{+}94]. E-cadherin/beta-catenin [BVF^{+}93]. E-mediated [SWFB91]. E-receptor [TMI^{+}91]. E-selectin [EWP^{+}92, LMIV93, EWP^{+}93a]. E1 [HWF92, HWF93, LDT^{+}92, SM91c]. E1a [Fri94, MSSD93]. E2 [HWF93]. E3 [GK90, KHHC93, SLM^{+}90]. E3/19K [GK90]. E5 [XMW^{+}93]. E8 [BHS91, ML94, SLM^{+}90, SSA^{+}90]. EA [IRH^{+}91]. EA-1 [IRH^{+}91]. each [DEH^{+}91, WCR93]. ear [RBR90]. Early [DWY^{+}93, GFH90, VM93, YDS92, AEGG93, BMK^{+}90, BLB^{+}91, BIPG91, CC90a, CF90, DFC^{+}92, DLSB^{+}90, DP93, FWF90, GGGE^{+}92, HG93a, HXM^{+}94, HW93a, KGK90, MRSN93, MJM91, MMD^{+}93b, OLKD90, PRGC91, SABF^{+}90, SBG93, STdL^{+}90, THL^{+}90, TH91, WC90a]. Echistatin [SSG^{+}90]. ectodermal [DS90c]. ectodomains [WKY^{+}93]. Ectopic [ARWD93, BRF93, ASDC91, BZB^{+}93]. edge [DI91, EMC^{+}90, KKES91]. EDTA [GHS^{+}93]. EF [CPM^{+}91, DMM^{+}91, TASJ92]. EF-1 [CPM^{+}91, DMM^{+}91]. EF-hand [TASJ92]. Effect [ABD^{+}94, BMM90, JK93, LC91, SWHCW91, WB93b, BRA^{+}94, HRH90, HKS^{+}92, KCDR91, LWK91, MDRG92, OCMB91, PvBeHB^{+}91, PMW92, SPHvD91, WBVB90]. Effects [AGKC92, CBRW92, FKB^{+}90a, HRT^{+}91, HBOV93, LCKW91, PHSvD92, BLR^{+}92, CDRL93, DLSK91, FCW91, GTP^{+}92, HR90a, HSvD93a, JGB^{+}93, Lam94, LSGD^{+}91, LVK^{+}91, MH93a, MJM91, MMD^{+}93b, PCD^{+}92, PW93, RPH^{+}92, SP92, TCDH93, TH92a, WC93, YGEG91]. Efficiency [RGHC91]. efficient [GS90, HSP93, MPAF91, PTS^{+}92, RA94]. efficiently [PL90]. EGF [JS90, TBL^{+}92, DMC^{+}91, FLUS92, HAK93, HSUS90, JS90, MCM93, MNV93, PSZ^{+}93, TBL^{+}92, WGRW91, dHvBeHVB92]. EGF-dependent [DMC^{+}91]. EGF-induced [WGRW91]. EGF-like [HAK93, PSZ^{+}93]. Egg [MGDS93, AL92, BDS92, BMPW90, BDK92, CLP^{+}94, DLMS91, DDW94, FL90, FL92, HN90, HODF94, HM93, JGLK90, KT93, LCMP90, LL91, LML^{+}94, LK93a, ODPL93, OPL94, SM92a, SSJ90, STH^{+}93, TS91, TJ91, TSBS92, VDS^{+}92, YHM^{+}94]. eggs [ABD^{+}94, CSMG90, DPW91, FMJ91, HE91b, KDP^{+}90, MMM^{+}92, SMCR90, VGC94, VBAK91, VL91, WCW91]. egr [OSKN94]. egr-1 [OSKN94]. EGR2 [SDR^{+}90]. egress [BF90]. eight [ROBN^{+}92, WEB90]. EIIIA [JRK^{+}94]. either [JAS^{+}90]. ejection [RS94]. ELAM [SNG^{+}91]. ELAM-1 [SNG^{+}91]. Elastic [FHI90, FKH^{+}93, BDGC^{+}93, ZAH^{+}94, ZDZSBT94]. elasticity [GXHS94]. elastin [BDGC^{+}93, HR94b, MWH^{+}91]. elastin-binding [HR94b]. elastin-microfibrils [BDGC^{+}93]. elastin/laminin [MWH^{+}91]. elastin/laminin-receptor [MWH^{+}91]. Electric [BL94, PBD93, SBK^{+}91]. electrode [KJ90]. Electron [KBL91, AR93, BK90, BZT93, DMLR^{+}94, FKT93, GA92, HASW94, KISY91, MMM91, MAFR93, NSH90, OOBJ92, RRG^{+}94, RSS91, WBR91]. electronmicroscopy [vGvMS^{+}91]. electropermeabilized [DCTG90]. elegans [CMM94, DE93, ECO93, FW91, GW94a, GW94b, HDC^{+}91, HW93a, HWW94, LA92, LH92, SJMK93, WW94b]. element [MCDY94]. Elements [NW94, CSR^{+}92, DAIS90, KT92, MH94, MC91c, OHC^{+}93, SDW^{+}93, VIF^{+}93]. Elevated [BS91, CG90, CETK^{+}94, TOA^{+}91]. elevates [MIU^{+}92]. elevation [JLC^{+}90, Sch93, YM90a, ZZP94]. elevations [JTS^{+}91]. ELH [JKS93]. elicited [CGW94]. elicits [LSG92, PCD^{+}92, SOH^{+}90]. elongation [BGH93, CPM^{+}91, DA92, DMM^{+}91, HSSC92, LSR^{+}90a, LELB90, MHYC90, RMK91, SW92, TK91, WPS91]. elsewhere [SEBH90]. Elucidating [CF90]. embryo [AAM92, BTS93, DFC^{+}92, FSY^{+}91, GDB^{+}94, HLR92, KL90, KSK^{+}93, LOC^{+}90, OLKD90, PSJ^{+}91, RLD^{+}91, TISH94, UHFG92, WFD92, vDS94]. embryogenesis [CHM^{+}92, FCD^{+}93, HDP^{+}93, HLCB92, KHS^{+}90, OH90, PJS^{+}92, SSB^{+}91, WBE91]. Embryonal [KMSW91, RWG90, SEC^{+}93, SBN92, SK91, SSFD93]. embryonic [AW92, BGLJ^{+}93, CWB93, CTQ91, FJRAT91, HW93a, Mah91, MLS94, MJ92, PSJ^{+}91, SHB90, SAM^{+}91, Tur91, WTS93]. embryos [CSMG90, DSJB^{+}94, ECO93, FKB^{+}90a, GKG90, GGGE^{+}92, GC90, GC91, HDDS94, HAS90, LW92, MCG92, MPVL92, MKF91, ORS^{+}90b, PMW92, RLMG92]. emergence [PZB^{+}94]. Emilin [BDGC^{+}93]. enable [IBR94]. enables [LML^{+}94]. enclosed [BBD^{+}90]. encoded [BTG^{+}93, CL92, GTC^{+}90, HE91a, HSCN94, KWK92, NEL^{+}91, PMA91, RR92a]. encodes [BHBG94, BLT^{+}90, BDJ94, EJ93, GW94a, GC93, HP92, JPS91, LXM91, LH92, MCS92, PH94b, PRAK94, SXM94, SNFN91, SB93b, WVH94, WTH^{+}94, WB94, dBBHA94]. encoding [CFG^{+}94, HVK91, MR91, OMRM93, OKMB91, SLK93, WG92b]. end [BCYC94, CC93a, Col94, Fow90, HW92, HWM^{+}93, MKSF93, SYO^{+}91, SRB92, SMM^{+}91, TDWT92, VGC94, MR94]. end-capping [SYO^{+}91]. End2 [PMD^{+}93]. end3 [RRCR93]. end4 [RRCR93, KD92]. endo [LB93]. endoCAM [AMBN91, AORB90]. endoCAM/CD31 [AMBN91]. endochondral [AWH^{+}94, GDB^{+}94, GKY^{+}94]. endocrine [BTH^{+}90, MK91a, MME93]. Endocytic [HH90, LVA^{+}90, BGH90, BWCG93, CNJ^{+}94, DBWS94, DHPG94, HSvD91, HSUS90, KBH91, LGH91, MP94, OTX^{+}94, PSD92, RHGG92, RYK^{+}90, SKSD^{+}94, TMI^{+}91, THL^{+}90, TH92b, WW91, WMCW92]. Endocytosed [TLXM90, SHH94]. Endocytosis [FGS91, HPM^{+}94, KSFG91, MR94, SPRvD91, SSJ^{+}94, BCLM94, CRWS93, CCC^{+}94a, CCC^{+}94b, CVK^{+}94, DHS93, FLUS92, GIAS93, HSvD93b, HK91, HBOV93, Hol93, IK94, JSM^{+}90, MYM94, MMH^{+}92, PMBL^{+}92, PHSvD92, RRCR93, SC90a, SOPA94, SCS92]. endocytotic [MTP^{+}92]. endoderm [DFLK94, RWG90, vdSKL^{+}93]. Endogenous [CMB91, GSS^{+}92, KHR91, LTYY94, LSMRB90, LSG92, NLFRB91, SFF93, WLWL94, WE90, vdSKL^{+}93]. endometrium [KCR^{+}94]. endonuclease [BG91a]. endoplasmic [AES94, BF90, BHLWH91, BKR90, CG91, DCH^{+}92, GK90, GPH^{+}90, GFW92, HR94a, JNP93, KGR94, KBA92, KDO^{+}94, MBC90, MH92, MYT94, NSHN92, NDM^{+}94, PPZ^{+}93, PNPB94, PNM^{+}94, PCP^{+}91, PSSK93, PDC91, RLS94, ROBN^{+}92, SCM90, SSC^{+}94, SDE^{+}93, SPW93, STH^{+}93, SB91, SPB90, SBL^{+}91, TBFP92, TS91, TJ91, TBKF^{+}92, VGMS91, VM94, VPP^{+}93, WF93, WBH^{+}90, YL93, ZABM94, dSBH90]. endoplasmic-sarcoplasmic [VPP^{+}93]. endoprotease [BLT^{+}90]. Endoproteolytic [PNB91, MJM92, XS93]. endosomal [BHD^{+}94, EGW^{+}93, HH90, KSVM92, KHV^{+}93, MPK^{+}94, ST94, TH91, ZBPV92]. endosomal/lysosomal [BHD^{+}94]. endosome [BS94a, BCBK92, HPW94, PLCB91, THL^{+}90, VHE93]. endosome-like [VHE93]. endosomes [AEGG93, AKM94, BHC^{+}94, CvdE91, DAW^{+}94, DKL^{+}90, DM92, GP91, HSvD93b, HG93a, MPM93, MMD^{+}93b, PDB^{+}91, RSS^{+}94, SWS^{+}92, TH92b, WLPB92, WRA93, WB92b]. endothelia [SHB90]. Endothelial [GHH90, HMY92, HMB^{+}93, ZMMP90, AORB90, AVP92, ASL^{+}94, BT92, BSSL^{+}91, BKB^{+}90, BRIA^{+}94, BPZ^{+}92, BIPG91, BDZ^{+}92, CSMdPSM94, CKD90, DCY^{+}94, DvHB^{+}91, DDFC94, EGP^{+}91, FAMR92, FAS^{+}93, GMGPM94, GWMP91, GYE^{+}90, GH91b, GS94, HDS90, ISF^{+}91, IPF^{+}90, JZAVP93, KYA^{+}94, vdWvKvKdB^{+}92, KCB^{+}92, KNR93, KW92b, KCYH92, KCC90, KHvdL^{+}92, KHHR92, LRDM91, LRR^{+}92, LSRC93, LSV^{+}94, LASB^{+}93, MMR91, MCF93, PZP^{+}91, PBM^{+}90, PSS^{+}93, RZS^{+}94, RM94, SR90a, STL^{+}90, SOA^{+}93, SPD^{+}90, SLP91, SH90b, SCMU93, Sch93, SNG^{+}91, STo^{+}92, TCDH93, TRLG90, TSLR90, TSR90, WIF^{+}90, dAWS^{+}90]. endothelial-rich [LSV^{+}94]. endothelial-specific [LRR^{+}92]. Endothelins [GH91b]. endothelioma [HJI^{+}93]. endothelium [KSST91, LPM^{+}91, LKD^{+}94, REW90, SOPA94, dAWS^{+}90, eSP91]. ends [BA92, FSM^{+}93, LCF92, NWvH94, WPBF94]. energy [CRS94, GAS90b]. engagement [LNSSA93]. engulf [ELZ93]. Enhanced [MMW^{+}92, GT94, JS90, OHF^{+}94, WKD^{+}91]. enhancement [DS90b]. enhancers [GFL92]. enhances [BGLJ^{+}93, HR90a, KKAS90b, MIS93]. enhancing [SH91]. enriched [EERM^{+}91, OESC91, TOA^{+}91, WP93]. enrichment [HDDS94, TA91]. entactin [YED^{+}93]. Enteric [SMHO92]. entering [RYK^{+}90]. enterocolitica [YFS92]. enterocytes [DHP93, HKS^{+}92]. enteroendocrine [RHG90]. enteropathogenic [RLB^{+}91]. entry [PMBL^{+}92, SWL^{+}92, STM91, YC93]. envelope [AA94, BAP^{+}90, BDW92, CC93a, FA90, GA92, GMML^{+}92, HWB93, HAS90, MF90, NWD90, ND92, SBP90, SB93a, SW93b, TSBS92, TIA93, UHFG92, VL91, WB93a, WB94, WEB90, WB92c, WBR94, ZPS^{+}92]. environment [Ada92, SHB90]. environmental [dlRKR^{+}90, dlR91]. enzymatic [SFO^{+}91]. enzyme [DBM94, EMC^{+}90, ET91, JK92, KWS94, KGM92, KMN92, LDT^{+}92, MME93, MR91, PDS90, ROBN^{+}92, WWHA91]. enzymes [GK93a, HSHA90, KKG^{+}90, LGH91, RCNO94, SNH^{+}94]. eosin [DMLR^{+}94]. eosinophils [GPF94]. Ep [LVB^{+}94]. Ep-CAM [LVB^{+}94]. Epidermal [CXS^{+}94, Fuc90, AW91, BMK^{+}90, BRF93, BCG^{+}92, BF93, CWBK90, CGW94, CCAF90, DPCP91, FRSL90, FFS^{+}93, HAK94, JS90, JFA^{+}91, KT92, KSL^{+}94, MLKB92, MSSD93, NGP^{+}91, OW90, PvBeHB^{+}91, RLD^{+}91, STF93, SCJ^{+}91, SKK^{+}91, TBL^{+}92, WLPB92, WJ92]. epidermal-dermal [STF93]. epidermis [DY93, EK90, YGEG91, ZDZSBT94]. epidermolysis [CHVF91]. epididymal [OHC^{+}93, PNB91]. epiglycanin [KWW^{+}94]. Epiligrin [WGM^{+}93]. epiphyseal [AWH^{+}94]. epithelia [BO91, DOW92, DFC^{+}94, MNME^{+}93, PRF^{+}93, PYH^{+}90, SMWB93]. Epithelial [RSB^{+}90, RRBG91, TRS^{+}90, AO90, AL91, BVF^{+}93, BML^{+}94, Cit92, CPD^{+}93, CPP^{+}94, DvdSM^{+}92, DS90c, DFC^{+}94, FB93, FTB94, FCD^{+}93, FF94, Fri94, GLQRB91, GIAS93, vtHvM90, IVBR93, INY^{+}93, JPJF91, KSFG91, LQNRB90, LYC^{+}94, LDNM92, LLG^{+}94, LVB^{+}94, LJP^{+}93, MK93, MCDA^{+}93, MRSN93, MELD94, MDRG92, MAAO92, NHS^{+}94, NI93, NSWH90, OB93, OS92, PCD^{+}92, PDB^{+}91, PKN91, PBDK90, PRC^{+}90, RBD92, RNG94, RA94, SZB94a, SVR^{+}90, SMC92, SBN92, SZP^{+}93, SSA^{+}90, SHB90, SdAN^{+}91, SBB91b, TBW92, TRH^{+}92, UK92, VBV^{+}90, WGM^{+}93, WBT^{+}94, WBVB90, WSB93, WG92b, WKN92, ZJA^{+}94, ZPS^{+}92, ZLC^{+}93, vGvMS^{+}91]. Epithelium [SSCD90, CSR^{+}92, GORB91, GPRB93, HKS^{+}92, PPHS^{+}92, RLKB91, SAF90, YLWS94]. Epithelium-dependent [SSCD90]. epithelium-specific [RLKB91]. epitope [CCW^{+}93, ESL^{+}90, FWD^{+}92, GG92, KWW91, KA93, LBH93, RKB^{+}90, RMG90, SEH^{+}93, vKWH^{+}91]. epitope-tagged [CCW^{+}93]. epitopes [FFSW91]. epsilon [GSM93, OSK^{+}92, OBC^{+}92, WCC^{+}90, GVK94]. epsilon- [GSM93]. epsilon-related [OBC^{+}92]. equal [GK92a]. equipped [VPP^{+}93]. equivalent [HR90b]. ER-derived [RS91]. ER-to-Golgi [ON94]. erbA [MWS^{+}93]. erbB [SBG93]. ERM [TSK^{+}94, TOS^{+}94]. error [NW94]. erythrocyte [Fow90, GHT92, KB91b, OKH92, SOH^{+}90]. erythrocytes [EH94, FFW91]. erythroid [BBMN94, GBB^{+}94, KWFM91, PBB^{+}91, PAC^{+}92, PSE^{+}90, TQL^{+}90]. erythroid-specific [PAC^{+}92]. erythroleukemia [ACFFL93]. erythrophores [KM94]. erythropoietin [LWHL93]. Escherichia [FPGS91, FSSA92, HPH^{+}91, ORS90a, RLB^{+}91]. Essential [LOOM93, BHBG94, BW91, BN91, BDJ94, EJ93, FM94a, GW94b, Ham94, HPP^{+}93a, HVK91, HAA93, JPS91, MMWG94, MPEG93, MNB92, MCS92, MJLD93, NMPN90, NCBS90, PA91, PH94b, PPZ^{+}93, PNPB94, RT92, SS94, SM93a, SNFN91, SSA^{+}90, SLK93, UHFG92, WTH^{+}94, dBBHA94]. essentials [Fuc90]. establish [BA93]. established [HVNC92]. establishing [BDJ94, ZCS^{+}91]. establishment [AJL^{+}90, BBGK90, RLMG92, VCFM94]. ester [DCL^{+}92, GBP^{+}92, JGLK90, Lam94]. ester-induced [DCL^{+}92]. esterase [TGR^{+}92]. esterases [BBD^{+}90]. esters [LW92, ZSC91]. estimate [LZI^{+}93]. estimated [CVH91]. estrogen [BPDM90]. estrogen-dependent [BPDM90]. etch [FKH^{+}93, NSH90]. Ethyl [NCBS90]. Ethyl-3 [NCBS90]. ethylmaleimide [GP91, NF90, WRHW92, WRB^{+}94]. eucaryotes [MPVB91]. Euglena [MB92, RMB90]. eukaryote [SGGC^{+}90]. eukaryotes [KMBR90, SSW94]. Eukaryotic [RHB^{+}93, LMW^{+}94, YFS92]. Euplotes [OMBL90]. evaluation [OH93, TLFS92, WTS93]. event [DLSB^{+}90, EP94, HBP^{+}93, SSB93]. events [AWFP93, BKB^{+}90, BBRE90, DJ93a, GE91, HRH90, HAA93, HVNC92, ISN^{+}94, KHP91a, KBL^{+}94, LCC^{+}93, MRSN93, ND92, PRGC91, SZB94a]. every [DEH^{+}91]. Evidence [BGBWO91, BB93, CWB93, CB90, DMB92a, FZRH91, FTT^{+}93, GdLMS94, HW94, HE91b, KDN94, LRAG92, MG94a, MJ92, PKPM90, VKS^{+}90, WPW92b, YGEG91, ZKMB93, dCTG92, ATSE90, ASR^{+}90, BEGN91, BGGW^{+}94, Blo93, BSHB90, CFFL93, CDRL93, DE93, DTY92, EK92, EG91, FHGN91, GSV93, HLH^{+}92, HLWL92, HSSC92, JD91, KA91a, KB91a, KWGS92, KLERG94, MHR94, MCM93, MVM91, NCBS90, POGS90, PB90, RMK^{+}94, RNS92, ROBN^{+}92, STM91, SAWL90, SLM^{+}90, STA^{+}94, TGD93, TDWT92, WPS91, WF93, WHW^{+}91, YMO92, dCQTR91]. evidenced [SVD90]. evoked [SPH^{+}91]. Evolutionary [JHK^{+}91, KKG^{+}91]. evolving [ET91]. examine [SLS^{+}90]. Examining [BLW92]. exchange [FZ90, FMR^{+}92, KSK^{+}93, PB90, SZP^{+}93, SMWB93, WNM^{+}94]. exchangers [RLWK93]. excised [KM92d]. excision [CK91, QYC^{+}92, SS94]. excluded [HTD90]. exclusive [BA92]. excretion [JHGR93]. exemplified [VLH^{+}93]. exhibit [BHA^{+}94, BMH^{+}93, GT94, LWK^{+}93]. exhibited [VDTT91]. exhibits [EOL^{+}90, MMS93, SMKH91, WS90]. existence [FHGN91]. exit [AM94, FBL92, MPK93, MH93b, PMD^{+}93, TGD93, VGMS91]. Exo [MTP^{+}92]. Exo-endocytotic [MTP^{+}92]. exocrine [ODK^{+}93, THL^{+}90]. exocytic [DMB92a, JCSH93, SSRM90]. exocytic-sensitive [SSRM90]. Exocytosis [ATSE90, ASE90, BMGN92, EP94, GCM^{+}94, HRT^{+}91, JGK^{+}94, KC90, KBP91, MHV92a, NSH90, SPH^{+}91, SLH92, TLWA94, VDTT91, WCW91]. exocytotic [BN91, KC90, KTG90, MF92]. Exofacial [CCW^{+}93]. Exogenous [SW90, AAM^{+}93, KA91b, MKG93]. exon [BNI^{+}94, RR92a, SGVS^{+}93, SA90]. exon- [RR92a]. exon-containing [BNI^{+}94]. exons [BBJ^{+}91, HE91a]. expansion [LMP91]. experimental [GB90, dlRKR^{+}90, dlR91]. explanted [WMvdK94]. explants [RRD93]. exploits [TDT92, TDWT92]. explore [KRMG93]. Export [DK92, CB90, ISDM92, JBIM94, Mac92, MCM92b, NDM^{+}94]. exports [EH94]. exposed [BDMPJ94, FMR^{+}93]. exposes [JO92]. exposure [CG90, CPD^{+}93]. express [AVP92, BBJ^{+}91, HTHC93, HNC^{+}92, KGSB92, KA90, PZP^{+}91]. expressed [AO90, BGK^{+}94, CSCSA94, CSMG90, CSH^{+}92, DvdSM^{+}92, DMB92b, FGK^{+}90, GN92, GSS^{+}91, GSC^{+}92, GWMP91, HMC91, HE91a, HHH^{+}93, HSLF^{+}92, HSUS90, KHS^{+}90, KFW^{+}91, KYLV90, LWL^{+}92, MK93, NIK^{+}93, PW93, SJG93, SG92, SLL94, SAM^{+}91, VGRC^{+}92, VKRD94, WBW^{+}94, WBH^{+}90, WHD^{+}91, ZDZSBT94]. expressing [GKY^{+}94, PCG^{+}94, WLZB93, aUFQ^{+}92]. Expression [AW91, AWFP93, BBK^{+}93, BRA^{+}94, BCB93, CH92, DR91, GTP^{+}92, HKS^{+}92, HHR^{+}90, HPH^{+}91, JRK^{+}94, JRCW^{+}90, JKS93, JCO^{+}94, KAF^{+}93, MTS^{+}94, MHR94, MMR91, MJM92, MHGC90, NBT94, ORS90a, OCC94, OBBS90, PPHS^{+}92, RMFA92, RLS92, SWD94, SEC^{+}93, SH90a, SA90, SSC90, SB90, SSB^{+}91, UK92, WBT^{+}94, WGD^{+}93, WTRD92, ZRW^{+}94, ASDC91, ARWD93, ALE^{+}92, AGB^{+}91, BPK91, BGGW^{+}94, BBK90, BLB^{+}91, BG91b, CK91, CETK^{+}94, CWB93, CMS^{+}90b, CSR^{+}92, DDW90, DBV^{+}92, DS90b, DvHB^{+}91, DGGG93, DZdH^{+}93, DY93, DAMS91, DP93, DFC^{+}94, DJF93, DYS^{+}90, ENS91, EMGS93, EFTJ^{+}93, EVL^{+}91, Eri93, EMC^{+}90, EZF^{+}94, ECAG93, FJRAT91, FMPP90, FGSBZ93, FNV^{+}92, FSL90, FPP91, Fri94, FFR^{+}91, FFB^{+}93, GBN^{+}93, GDB^{+}94, GL91, GFH90, GR93, GB90, GFL92, GTH90, HW90, HDC^{+}91, HSLF^{+}92, HW94, JHGR93]. expression [JHK^{+}91, Jas93, JJB^{+}93, KSE^{+}92, KA91a, KCDR91, KCR^{+}94, KBT92, KW93, KKBK92, KVF^{+}90, LWK^{+}93, LYC^{+}94, LCM^{+}94, LRSB94, LFB94, LVTB93, LCLG94, LSS^{+}90, LOC^{+}90, MCM92a, MPR90, MMG90, MRBP^{+}90, MBC^{+}91, MGB90, MSB^{+}91, MHV92a, Mil90, MMT91, MSSD93, MIU^{+}92, MSF90, MR91, MM93b, ME91, NNHB^{+}92, NHH^{+}91, NDdC92, PSZ^{+}93, PTTA90, PC90, PSJ^{+}91, PSS^{+}93, PBCK93, PRC^{+}90, PRB^{+}91, RLRR93, RSM^{+}94, RGKG90, Rob90, RWG90, RDBHG90, SCO^{+}91, SGVS^{+}93, SMRG91, SCM90, SNE90, SLG92, SLHG93, SBH92, SON^{+}94, STF93, SMJ^{+}94, SEMH92, SDPH90, SAVK94, SSFD93, SCSE90, SBB91b, SSK^{+}93, STo^{+}92, SSKT^{+}92, STA^{+}94, TMP^{+}91, TFC94, TBW92, TN94, TFT94, TGD93, TLSW93, TNT^{+}90, TGCT90, VIF^{+}93, VNS^{+}94, WKD^{+}91, WFD^{+}91, WOC90, WBE91, WHD^{+}91, WWHA91, WTK^{+}92, YR93, YOWM91, YH93, ZAH^{+}94, ZSC91, ZZR94]. expression [dCV90]. extend [VFR94]. extended [WKM^{+}92]. extension [CKMC94, PS94]. extensions [AB91]. extensive [WMB^{+}94]. extinguished [GHBA^{+}93]. Extracellular [BJ90, HMC91, JGB^{+}93, LSR^{+}90a, LSR90b, SSK^{+}93, ZZL^{+}91, vKWHF94, AAM92, ATBD90, BKB^{+}90, BS91, BTG^{+}93, BTR92, BFTR93, CSMdPSM94, Cit92, CB90, DP93, FL90, FL92, FvBuHWS92, FSSA92, GBN^{+}93, GL91, GDKP92, HBS92, HR94b, JH93, KSB^{+}93, vdWvKvKdB^{+}92, KBM93, KJ90, MRH^{+}91, MGPPH93, MFS^{+}94, MHG^{+}90, NR90, OLKD90, PSZ^{+}93, PL90, PHSA93, PNB91, PAFC92, RR92a, RPPB94, SSCD90, SF91, SNRS93, SB90, TMHKO94, TBW92, TLSW93, VIF^{+}93, WBW^{+}94, WF94, YMO92]. extract [BDS92, DLMS91, UHFG92]. extraction [EK90, TCCT91]. extracts [AL92, BDK92, DDW94, HODF94, HM93, KT93, LCC^{+}93, LL91, LK93a, MSC92, SM92a, VBAK91, VDS^{+}92, WE90]. extracytoplasmic [JJB^{+}93]. Extragenic [PPD92]. extrinsic [CWB93, ET91]. Ezrin [ATV^{+}93, TWV94].

F [Apg91, BGI^{+}90, CLZ91, CCFZ92, CWHH92, FMR^{+}93, GeKdV^{+}91, JBJH^{+}92, MKB^{+}94, SGC^{+}94, SY90b, VJB^{+}94, WPW92b, WOC90]. F- [VJB^{+}94]. F-actin [Apg91, BGI^{+}90, CLZ91, CCFZ92, CWHH92, FMR^{+}93, GeKdV^{+}91, JBJH^{+}92, SGC^{+}94, SY90b, WPW92b, WOC90]. F-actin-binding [MKB^{+}94]. F11 [ZDR^{+}92]. F3 [DGGR92]. F9 [BG91b, RWG90, SEC^{+}93, SSFD93]. FAA [JKLG94]. FAA1 [DKG92]. Fabpi [CSR^{+}92]. face [GA92, VM94]. facilitate [WIF^{+}91]. Facilitated [PJS^{+}94b, AP92, BHF90]. facilitates [TDSP93]. factor [ALE^{+}92, BWV92, BR94a, BML^{+}90, BG91a, BMK^{+}90, Blo93, BIPG91, BGI^{+}90, BGRW91, BSR^{+}90, BDZ^{+}92, BFTR93, CGW94, CXS^{+}94, CPM^{+}91, CDRL93, DPS^{+}93, DPW91, DPCP91, DDW90, DJ93a, DMM^{+}91, DGGG93, DFC^{+}92, DZdH^{+}93, ET94, EOO^{+}94, Eri93, EB92, FMR^{+}93, FMR90, FAMR92, FMLD92, FSY^{+}91, FJN93, FJN94, FFS^{+}93, GMLD92, GBB^{+}94, GP91, GBOB90, GHH90, GS94, HJI^{+}93, HK94, HMLL94, HAK93, HAK94, JS90, JGB^{+}93, JRSB90, JCO^{+}94, KT92, KSL^{+}94, KS92, KCR^{+}94, KHR91, KNR93, KBT92, KDT^{+}91, KRB91, KDT93, KHHR92, LHS^{+}93, LWV92, LKXS93, LSK^{+}94, LSR90b, LCK^{+}92, LGPM90, MCM92a, MMG90, MPEG93, MGPPH93, MWG92, MMW^{+}92, MMR91, MMT91, MLM^{+}93, NS90, NF90, NHK^{+}94, OSKN94, OH90, OW90, PPHS^{+}92, PC91b, PvBeHB^{+}91, PBM^{+}90, RLD^{+}91, RJJ^{+}94, RNG94, RDBHG90, RHB^{+}93, SR90a, SFM^{+}90, SHR93, SK91]. factor [SCMU93, SST92b, SYG91, SDK^{+}93, SSCD90, SR90b, SKM94, STF93, SMWB93, SKK^{+}91, SR92c, SMBG92, TMHKO94, TK94, TKHM91, TRLG90, TBL^{+}92, TTR^{+}90, TSR90, UK92, VBvG^{+}94, VGRC^{+}92, VFC^{+}91, WLPB92, WRHW92, WBVB90, WSB93, WHT^{+}93, WB93b, YGEG91, YYY^{+}91, YM90a, YM90b, ZMMP90, BCET^{+}92, JK92]. factor-1 [BSR^{+}90, SFM^{+}90]. factor-4 [GS94]. factor-beta [BFTR93, DGGG93, HMLL94, JRSB90, KNR93, KBT92, LCK^{+}92, MMG90, OSKN94, PBM^{+}90, RDBHG90, SCMU93, SSCD90, SR92c, TMHKO94, VBvG^{+}94]. factor-heparan [SR90a]. factor-induced [FAMR92, PBM^{+}90]. factor-inducible [HJI^{+}93, LWV92]. factor-like [KSL^{+}94]. factor-mediated [YGEG91]. factor-receptor [SKK^{+}91]. factor-scatter [JCO^{+}94]. factor/hepatocyte [SMWB93, WSB93]. factor/scatter [UK92]. Factors [SSH94, AMG90, AA94, CG90, CSKZ94, DS90a, DCB^{+}92, GFWM92, HMC91, JBIM94, LSP^{+}93, MTS^{+}91, NBT94, OMS^{+}94, RBD92, RRH91, RDBHG90, SBB^{+}91a, TN94, TWC93, THH^{+}90, VBAK91, WBS91b, YSK^{+}94]. fail [QORM91, SW90]. falciparum [CKGC91, EH94, GRC^{+}91, KCM^{+}93, SOH^{+}90, OKH92]. falciparum-infected [GRC^{+}91]. family [AST^{+}92, CJB^{+}93, CMK^{+}90, FWD^{+}92, FFR^{+}91, FNS^{+}91, JTS^{+}91, JKH94, KCM^{+}93, LDW^{+}93, LWV92, LYC^{+}94, LGSB93, MMS93, MGS90, ME91, OKMB91, PMG^{+}92, RZS91, SGWE94, SSDK^{+}94, SJG93, SLK93, TSK^{+}94, TOA^{+}91, TOS^{+}94, TWV94, WRB92, WKM^{+}92, YSK^{+}94, YLWS94]. fasciclin [EHB^{+}90]. fasciculata [FTPM^{+}94]. fasciculation [KS90]. fascin [CKMC94]. fashion [BMH^{+}93, CPD^{+}93]. fat [BBB91]. fate [FKS93, KRMK93, KW93, LW92, SBN92]. fates [BP91]. Fatty [PGAR90, CSR^{+}92, DKG92, JKLG94, PS94, SY91, WRP^{+}91, YSK^{+}94]. Fc [GeKdV^{+}91, HCYK91, MYM94, MMH^{+}92, OSK^{+}92, PMM90, RGR90, ZB94]. features [KISY91, WFD^{+}91]. Feedback [SMTW94, LSG92, WGRW91]. females [TH92a]. femur [JRSB90]. fertilization [BMPW90, STH^{+}93, TJ91]. fertilized [HE91b]. FET [WTK^{+}92]. Fetal [DFLK94, CVK^{+}94, LVTB93, RGR90, VNS^{+}94]. fetoprotein [TGCT90]. fetus [GBOB90]. FGF [BTRB93, BT93, OR92, SK91, WN92]. FGR [BFLS94]. Fiber [DAMS91, CWB93, DAM^{+}93, DH93, IITG94, KGDE94, KJT91, LWK91, MH91, MBHG91, MS92, SSS^{+}90, SBLV92, Woo94]. fibers [AR91, ASR^{+}90, BDGC^{+}93, DE90, EKK94, Gau90, HSCR94, HASW94, LM91, LC91, MFOI92, SAWL90, TASJ92, VPP^{+}93, WGP^{+}93, Woo94]. fibril [CLG^{+}92, DWJP93]. fibrillar [LGI^{+}93, Sch91]. fibrillarin [JHK^{+}91]. fibrillary [CL94, WSL91]. fibrillin [KS94, ZAH^{+}94]. fibrillin-2 [ZAH^{+}94]. fibrillogenesis [LGI^{+}93, PFSS92]. fibrils [AP92, KLM^{+}91, KDS^{+}91, PPFM90, RSB^{+}90, SLKS94, UNY90, WKS^{+}93]. fibrin [WS92]. fibrinogen [DS93, LFWC92, LSC^{+}93, NVR92]. Fibroblast [OH90, AG93, AUT^{+}91, BL94, BGRW91, BSR^{+}90, DDW90, DS90a, FE90, FMR90, FAMR92, FSY^{+}91, GBOB90, GHBA^{+}93, GDKP92, GA91, GHH90, KDT^{+}91, MH91, MRH^{+}91, MMR91, NRS^{+}93a, PBM^{+}90, PPFM90, PAL^{+}90, PH92, SR90a, SFM^{+}90, TMHKO94, TSR90, VBV^{+}90, VB93, WLWL94, WB93b]. fibroblastic [DvdSM^{+}92, PLB^{+}92]. Fibroblasts [Gri94, AGB^{+}91, BPDM90, BPTS94, BSM90, CETK^{+}94, CGF^{+}93, CVK^{+}94, DLD^{+}90, DI91, DCM^{+}93, DGGG93, EFTJ^{+}93, EOL^{+}90, FYDB93, FE90, GTP^{+}92, GH91b, HE93, HMC91, HG94, HIJ90, HGS^{+}94, HRB92, JS90, KATS90, KSVM92, KS94, KDS^{+}91, KA91b, KW92b, KP90, LSAH90, LSP^{+}93, LG93a, MP94, MHT^{+}92, MMD91, NI93, PPL^{+}93, PA91, PBDK90, PH92, RH92b, RMFA92, SHLS93, SM91d, SM92b, TISH94, TM92, TLSW93, WTRD92, YYM90, ZBPV92]. fibrogenesis [JRK^{+}94]. Fibronectin [DWJP93, BSG^{+}92, BGLB94, CNSC90, DGL^{+}90, DG90, DSQ^{+}93, DKM^{+}92, EUH91, FKHG93, FAYM90, GTH90, HMS^{+}90, HFS92, ISF^{+}92, IPF^{+}90, JJ92, LAY92, LSC^{+}93, MHM91, MH93a, MSQ^{+}90, MIS93, MR92, MM92, NYA^{+}91, PZV^{+}91, RPS^{+}92, SADJGP94, Sch93, Sch91, SGN^{+}93, SGM^{+}90a, TMP^{+}91, TYM^{+}92, TCDH93, WK92, WOC90, YYY^{+}91, ZMV^{+}93, ZCP^{+}94]. fibronectin-coated [LSC^{+}93]. fibronectin-derived [DKM^{+}92]. fibronectin-mediated [SGN^{+}93]. fibronectins [JRK^{+}94, PPFM90, WS92]. fibrosis [DOW92, GKY^{+}94]. fibrous [GA92]. Fibulin [ATBD90, PSZ^{+}93]. fibulin-2 [PSZ^{+}93]. field [BL94]. field-directed [BL94]. fields [PBD93]. filament [BPTS94, Bea91, BMS^{+}91, CLZ91, CCAF90, CF90, DE93, DEH^{+}91, EBS91, ECO93, FH90, GW94a, GMQ^{+}93, HW92, HWM^{+}93, KRHY93, KMN91, KWW91, KO94, MWSP91, MZP91, MH94, MGG93, MVG91, NPR^{+}94, OS92, OD93, SNE90, SGC^{+}94, SAJ^{+}94, SG92, SBC^{+}93, TCP90, TDWT92, TTE^{+}94, VLGB92, WDWT90, WGD^{+}93, WCF92, ZFF92]. filament-associated [GW94a]. filamentous [JETS91, PC91a, PMTB94, VDTT91, WP93]. filaments [AHW90, BMF^{+}91, CW90a, CW90b, CLCC90, CMZ90, CL93, CRK^{+}93, CM93, EK90, EKK94, ECO93, FB93, FW93, FSM^{+}93, Fuc94, FHI90, FKH^{+}93, FKT93, GMNH93, HD91a, HW90, HW92, HNP90, KPMG91, KHF94, KWW91, LM91, LWK91, LCKW91, LB92b, MWSP91, MHG93, MWD94, MA90, MHGJ94, NKW^{+}92, OMH93, RLKB91, SGW^{+}91b, SLI^{+}91, SKT^{+}90, SJGG94, VCL93, WGP^{+}93, WPBF94, WDB^{+}92, ZDP92]. filamin [GYE^{+}90, HLPL90]. Filensin [MBHG91, GMQ^{+}93, MGG93]. Filipin [SOPA94]. Filipin-sensitive [SOPA94]. filopodia [OB93, VFR94, WG93]. Fimbrin [dAWL^{+}90, HSB^{+}94, Mat94]. fimbrin-binding [HSB^{+}94]. finger [SCM^{+}93]. firefly [GKKS90]. first [SB93c]. Fischer [ZLC^{+}93]. fish [BOM94, FMJ91, Lut91, RLGB94]. fission [BWLK93, EPNN91, FHUY93, HRH90, HO94, MHYC90, MSC92, PC91c, SY94a, SBK94, SY92]. Five [HJI^{+}93]. fixed [BPDM90]. FK506 [BZT94]. FK506- [BZT94]. FLA10 [WVH94]. flagella [BU94, Bro91b, GOP^{+}94, HKS90, MOM^{+}92, PRSS90, PR91, PMS92]. Flagellar [DA92, ZS94, BS94b, CK91, DAR93, HSS93, HW93b, HHG^{+}94, HKS90, JR92, KK92, KKM91, KGM92, PMPV92, PPD92, QYC^{+}92, SS94, SS92a, SIR^{+}94, WGP^{+}93]. flagellum [BBKR94, CdJC93, DA92]. flagellum-cell [CdJC93]. flexible [BK91]. Flexural [GMNH93]. FLG [GBP^{+}92]. flight [WYM^{+}92]. Flightin [VSVP93b]. Flow [SCS90, ARW^{+}94, HW93a, KBP91, KES90, KPAY91, LKD^{+}94, MPM93, PDC91, WTSP91]. flower [KHCR91]. flower-like [KHCR91]. fluctuations [GMNH93]. fluid [RRCR93]. fluid-phase [RRCR93]. fluorescein [CVH91]. Fluorescence [DMLR^{+}94, GT93, HKSL90, IK94, OH93, SAH^{+}92]. Fluorescent [GT94, WSvdK^{+}93, WMB^{+}94, KFN93, KA91b, MP94, OH93, PMKH91, THBW94]. fluorescent-labeled [KFN93]. flux [SM91b, TIA93]. Focal [MULA^{+}91, CKG^{+}90, DI91, HWLT94, HSCR94, HSP93, MH91, SLPB93, TWC93, TGB90, YCO^{+}93]. foci [YKY92]. Focusing [FZ93]. fodrin [GORB91, NSWH90]. folate [CRKA92, RYK^{+}90]. fold [SHM94]. folded [BB94, HTBF94]. Folding [BHLWH91, FSSA92, HH94b, MVCC93, SBSG92, WBS^{+}91a, dSBH90, dSBH93]. follicle [FPR90, FMCR93, MSSD93]. follicles [KW93, MPR90]. follicular [SJIC92]. following [MRSN93, RB92]. follows [KCB^{+}92, RMMH93, YC93]. Foot [SAM^{+}91]. footprint [Mat94]. force [AR91, DIYSS92, ELZ93, Gol93, HS90a, HLB90, HFAGM94, SSS^{+}90, WCR93]. force-producing [WCR93]. forces [FE90, LLO^{+}94, RS94]. Forest [GHR^{+}90, MPK93, PK91, SWL^{+}92, WG92a]. form [AJPB90, BO91, BCKS92, CHBGL91, CL94, DGGR92, HMLL94, HBS92, KRHY93, KB91a, KA90, LD91, LVK^{+}91, LSC^{+}93, LNS^{+}94, LWB91, MLM^{+}93, MGW92, OPL94, PMG^{+}92, QORM91, SSTL93, SRHK94, SR92a, TDT92, TDWT92]. Formation [PP91, RGJ91, SGN91, AC93, AV91, AWH^{+}94, AP92, BP93, BGBWO91, BCBK92, CNJ^{+}94, CKMC94, CW90a, CW90b, CC90a, CWB93, DBWS94, Dun90a, DWJP93, FAWM94, FAF^{+}93, FFR^{+}91, GMM^{+}93, GDB^{+}94, GKY^{+}94, HODF94, HWM^{+}93, HNPN94, HAS90, IITG94, KT93, LNSSA93, MWSP91, MH91, MMJ^{+}90, ML94, MKSF93, PBD93, PYA90, PKPM90, PSSK93, SS91a, SGW^{+}91b, SBSG92, SNRS93, TSK^{+}94, VKM93, WRA93, WCR93, WRHW92, WSL91, WOC90, WRP^{+}91, XS93, YTN^{+}94, YDS92, YCO^{+}93, dWLS91, vdBRD^{+}93, TDT92]. formed [BMF^{+}91, MH94, TIA93, WDB^{+}92, ZBS^{+}94]. former [LC94]. forming [CSH^{+}92, FKG94, KGSB92, PYA90, WHN^{+}94]. forms [BMH^{+}93, CH93, DMM^{+}91, DBC90, DMB92b, FGH92, GTH90, KSB^{+}93, LM91, LTS93, MG94a, MJM92, MKF91, PL90, PAC^{+}92, PJC^{+}90, QXWN^{+}94, SC90b, SJM^{+}94, TRS^{+}90, TIA^{+}92, WGP^{+}93, WB93a, YCC92]. formyl [JWHPM93]. Forskolin [LSGD^{+}91]. Forssman [vGvMS^{+}91]. fos [CMS^{+}90b, GMLD92, GSWW93, RMFA92, SDR^{+}90, BMB^{+}92]. found [BZB^{+}93, KHS^{+}90, PLCB91, SF91]. fountain [vDS94]. Four [SNG^{+}91, CHW94, DGAB^{+}93, FRSL90, JKLG94, PH92, SKM94, VWS90, YKM^{+}92, YLWS94]. four-way [FRSL90]. fourth [HSM^{+}93]. FPR3 [BZT94]. fraction [HLWL92, JPJF91, JJS92]. fractionated [HODF94]. fractionation [SMC92, XL91]. fracture [SSM^{+}92a, VBS^{+}93]. fracture-label [SSM^{+}92a]. fragment [BHS91, FCW91, FL90, FL92, KIS^{+}94, MSQ^{+}90, MM92, SLM^{+}90, WLWL94, SRG^{+}94]. fragmentation [GSBS92, MW94, SOWDK94]. fragments [PB91, SHBD90, SLM^{+}90]. fragmin [FH90]. fragmin-actin [FH90]. frames [ZZR94]. free [BF90, BCET^{+}92, CRS94, FSM^{+}93, GC91, HMB^{+}93, JLC^{+}90, JTS^{+}91, KT93, LCC^{+}93, MM90, MHM91, MSC92, MMC91, MW94, MPK^{+}94, QCCD92, RPE^{+}90, RLD^{+}91, Sch93, TM90, UNY90, UHFG92, Wat90, WMCW92, YDS92]. freeze [NSH90, VBS^{+}93, vGvMS^{+}91]. freeze-fracture [VBS^{+}93]. freeze-substitution [vGvMS^{+}91]. freezing [KBP91]. frequency [KK92, WYM^{+}92]. frog [BH94, CQYD94, DP91, HE91b, OH92, SLH92, SCV^{+}91, SM92a]. frozen [ASR^{+}90, Woo94]. frozen-hydrated [ASR^{+}90]. FT210 [HTP^{+}92]. fucosylated [ZMS^{+}91]. full [FFR^{+}91]. full-length [FFR^{+}91]. fully [MC91c]. function [BW91, BLW92, BSF90, BHA^{+}94, BMGN92, BG94, BHLA90, CBRW92, CHVF91, DGL^{+}90, DCCH92, DCBH92, FMD93, FJRAT91, FMB^{+}91, FAF^{+}93, FMLD92, FDS90, HG92, HHR^{+}90, HTGN94, KWS94, LPA^{+}90, LPPCF91, LFB94, LMB^{+}94b, NYA^{+}91, ONN91, OK90, PPHS^{+}92, PMPV92, PNPB94, PRAK94, PC91c, RSS^{+}94, REW90, RHR90, SH90a, SMB^{+}91, SM93b, SEH^{+}93, SZP^{+}93, SFF93, SBG93, TBW92, TSS^{+}92, VMR90, WCM^{+}93, WW94b, WB92b, YGM^{+}94, vKWHF94]. function-associated [TSS^{+}92, vKWHF94]. Functional [AFA^{+}94, BVS^{+}93, FAU^{+}93, Gol93, GFM93, KKAS90a, LS90, NTB90, OW90, RLWK93, SY92, SBC^{+}93, TK94, BBK^{+}93, BO91, BMH^{+}93, CTL91, CHBGL91, CH93, DBM94, DB91, ENS91, FKHG93, FFKL92, FSL90, FFSW91, GdLMS94, GRB^{+}90, HE91a, HSLF^{+}92, HWD94, HKS90, JHGR93, JHK^{+}91, JJB^{+}93, JCO^{+}94, KSST91, LVA^{+}90, MHR94, MC91c, NIT94, ON94, PSN^{+}91, QXWN^{+}94, RG92, RKB^{+}90, SST92a, SYG91, SRP90, SS92a, WPW92b, WHD^{+}91, WW91, ZSH^{+}93]. functionally [BHHG90, HKC92, HR90b, SEAB91, VRE^{+}92]. functions [BLSR92, BGLB94, BT93, BDH^{+}94, CKG^{+}90, CCC^{+}94a, CCC^{+}94b, DHS93, EUH91, FvBuHWS92, GJJM94, GHBA^{+}93, HS92, HO94, LTYY94, MSS^{+}94, RCNO94, SLM^{+}90, SHOA92, VHSR92, YCO^{+}93, ZMV^{+}93]. fungi [PMTB94]. fura [MOKF92]. fura-2 [MOKF92]. furin [BHD^{+}94, VSB^{+}94]. furrow [CW90a, SYO^{+}91, YNST93]. further [CGMC92, ES91]. fuse [MPK^{+}94]. fused [SMM^{+}91]. Fusion [MC91c, WHK90, BGBWO91, BMPW90, BDW92, CH92, CVH91, CCT94, EFD^{+}93, EOO^{+}94, JLC^{+}90, KBL^{+}94, LCM^{+}94, MF92, NIT94, ND92, PW93, PGAR90, PNM^{+}94, RS93, RS91, RKB^{+}90, SWL^{+}92, SCV^{+}91, SRP90, TH92b, TIA93, VHSR92, WG92a, WF93, WRB^{+}94, WDRM93, WWW^{+}92, WMCW92, YM93, ZBS^{+}94]. fusogenic [EGW^{+}93, TIA93]. Fyn [BSM94, LMB^{+}94a]. Fyn-minus [BSM94].

G [HCW94, WCF92, BS94a, BF90, CSNZ92, HSCR94, IKL^{+}94, JGL^{+}93, LTPB92, SBB^{+}91a, SdAN^{+}91, VJB^{+}94, WW94a, dSBH90, dSBH93]. G-actin [CSNZ92, VJB^{+}94]. G-protein [HCW94, JGL^{+}93]. G1 [BB94, BS94c, GA91, HGDA93, LMB^{+}94b]. G1/ [GA91]. G1/S [HGDA93]. G2 [DS90c, Jas93, LOOM93, RTHB90]. G2/M [Jas93, LOOM93]. G4 [KSC^{+}91]. GABAA [ESL^{+}90]. GABAA/benzodiazepine [ESL^{+}90]. GAD65 [CAS^{+}92, SVB94, SDR^{+}94]. Gadus [BSW91]. GAG [LCPAM94]. gain [BVF^{+}93]. Gal [OKH92]. Gal- [OKH92]. galactocerebroside [DB90]. galactosaminoglycan [BLR^{+}92]. galactosaminoglycan-containing [BLR^{+}92]. galactose [CKD90, HSW90]. galactose-specific [HSW90]. galactosidase [IBD92]. galactoside [BZF^{+}92, WLH92]. galactosyltransferase [BS90, BHS91, ELS93, HS92, YHM^{+}94]. galvanotaxis [UNY90]. Gamma [BJ92, BRL^{+}93, DFC^{+}94, FAWM94, HSCR94, HG93b, HO94, JGL^{+}93, LSG92, LG93b, MVCC93, MBB^{+}93, RKA93, Rob90, RPH^{+}92, RELS^{+}91, SLG92, SJG93, TK94, TBFP92, WCC^{+}90, ZB94, CPD^{+}93, YM90a]. gamma-actin [LSG92, LG93b, SLG92]. gamma-adaptin [Rob90]. gamma-subunit [JGL^{+}93]. Gamma-tubulin [BJ92, FAWM94, HO94, MVCC93, RKA93]. ganglion [NFL90, NPST94, SF90a, WMvdK94]. Ganglioside [DAS^{+}92]. gangliosides [OSK^{+}92]. gap [CNZ^{+}92, FKG94, FSL90, GKG90, HSLF^{+}92, HR90c, JFA^{+}91, KGDE94, KKW^{+}93, Lam94, LTP^{+}92, MLRJ92, MCEG90, MG91b, PET^{+}91, RGKG90, SKGU91, TND^{+}93, WHD^{+}91, CSH^{+}92, GB90, MBS91, PBCK93, WC93, dCV90]. GAP-43 [CSH^{+}92, GB90, MBS91, PBCK93, WC93, dCV90]. gap-junction [LTP^{+}92]. GAP43 [DJA^{+}90]. Gas2 [BBS92, BS94c]. gastrointestinal [KEBD91, RHG90, TGCT90, WWHA91]. gastrulation [ORS^{+}90b]. gated [MHG^{+}90, PET^{+}91]. gCap39 [Bea91]. GDI [KSK^{+}93]. GDP [KSK^{+}93, NDM^{+}94]. GDP-bound [NDM^{+}94]. GDP/GTP [KSK^{+}93]. gel [GH91b, VBV^{+}90]. gelatinase [RSM^{+}94, RM94]. gelation [BGI^{+}90, JKT91, dAWL^{+}90]. gelation/solation [JKT91]. gelling [SHTL93]. gels [CETK^{+}94, GHH90]. Gelsolin [HCYS90, Bea91, DPCP91, FFKL92, FHI90, SSSL94, VBV^{+}90, VDTT91, WPW92a, WPW92b]. Gelsolin-actin [HCYS90]. Gene [Eri93, GL91, RWG90, SCSE90, AO90, AW92, BHBG94, BWV92, BVF^{+}93, BZT94, BLB^{+}91, BG91b, BIPG91, BLT^{+}90, BTG^{+}93, COTC94, CK91, CL92, CMM94, CSR^{+}92, CCW^{+}92, CBM^{+}94, CHG94, DDW90, DAM^{+}93, DY93, DMJ^{+}94, DKG92, DYS^{+}90, EZF^{+}94, FMCR93, FCD^{+}93, FFR^{+}91, GS90, GW94a, GK93b, GFH90, GFL92, GSM93, HS90b, HSLF^{+}92, HMS94, HVK91, HTD90, HLCB92, HHLS92, Jas93, JLC93, JP90, JPS91, KHS^{+}90, KM92c, KHP91a, KDT93, LCMP90, LDW^{+}93, LXM91, LWL^{+}92, LYC^{+}94, LBL^{+}91, LH92, LBH92, LYK^{+}94, LG93b, LVTB93, LMB^{+}94b, LSS^{+}90, MCM92a, MBW^{+}94, MRBP^{+}90, MT94, MAB^{+}90, MMT91, MCS92, MK91b, NHO^{+}91, NNHB^{+}92, NEL^{+}91, OMRM93, OGM^{+}93, OOM90, PMG^{+}92, PBCK93, PCG^{+}94, PRC^{+}90, RB94, ROE^{+}90, SBB^{+}91a, SLHG93, SNFN91, SB93b, SJMK93, STF93, SS93, SMJ^{+}94, SSFD93, SBB91b]. gene [SSK^{+}93, SLK93, STA^{+}94, TN94, TQL^{+}90, TFT94, TNT^{+}90, TGCT90, VMS^{+}91, VIF^{+}93, WVH94, WTH^{+}94, WHD^{+}91, WKMG92, ZSC91]. gene-encoded [NEL^{+}91]. general [VGMS91]. generalized [BHF90]. generate [DMB92b, MC91c]. generated [BZF^{+}92, CYC91, DBC90, LLO^{+}94, ZLG^{+}90]. generates [BZB^{+}93, MM92]. generating [MNME^{+}93, RBD92]. Generation [BLSR92, vtHvM90, JH90, Gol93, HRT^{+}91, HFAGM94, KHR91, ZB94]. Genes [WW94b, AJL^{+}90, BS93a, BSHB90, CKL^{+}91, FGK^{+}90, GTC^{+}90, HYD93, JRCW^{+}90, JKLG94, KMSW91, KG94, LWK^{+}93, LSG92, LOC^{+}90, MPR90, MR94, OKMB91, PDC91, SLG92, SJG93, SSB^{+}91, WGBB91, WA93]. Genetic [BKWD94, PDTG92, SJMK93, RMK^{+}94, RPPB94, SGP91, WHK^{+}93]. Genetically [BBB91]. genetics [FKB^{+}90b]. genome [HTGN94, PRB^{+}91]. genotoxic [BRD94]. geometry [FRSL90]. germ [WHM^{+}92, ZWW90]. germinal [RMG90, TSBS92, WMCG91]. GFAP [MJM91]. GFAP- [MJM91]. Gi2 [LHS^{+}93]. giant [SRM94, SJM^{+}94]. gland [LWL^{+}92, SSCD90, STA^{+}94, YSK^{+}94]. GLcNAc [MCJ^{+}93]. GLcNAc-bearing [MCJ^{+}93]. GLFG [WB94]. glia [BGM^{+}91, GAS^{+}90a, MFS^{+}94, MBW^{+}94]. glial [CL94, DJA^{+}90, IITG94, SBWV90, WSL91]. glimepiride [MDKB94]. glioma [CNZ^{+}92]. Globular [AHM94, DMB92b, SLWF91]. globulin [DYS^{+}90]. glomerular [DB91, SAF90]. Glu [GEKL^{+}94]. glucan [BSHB90, RPPB94]. glucocorticoids [EFTJ^{+}93, QCCD92]. Glucose [MB93, CCC^{+}94a, CCC^{+}94b, CCW^{+}93, GCZ^{+}92, HSP^{+}91, HRB92, HFS^{+}93, MDKB94, PTS^{+}92, RPH^{+}92, SGG^{+}91, VHB93]. glucose-induced [GCZ^{+}92]. glucosidase [BKR90]. glucosylasparagine [SSW94]. Glucosylceramide [JKB^{+}92, MP94]. glucosylphosphotransferase [SSRM90]. GLUT [PTS^{+}92, JP94, PTK^{+}93]. GLUT-4 [PTS^{+}92, JP94, PTK^{+}93]. GLUT2 [TGD93]. GLUT4 [HSP^{+}91, HFS^{+}93, RPH^{+}92, CCC^{+}94a, CCC^{+}94b, GJJM94, RPH^{+}92, SLB^{+}94]. glutamic [SVB94]. Gly [BSSL^{+}91, BDB90, DG90, EUH91, YYY^{+}91, YED^{+}93]. glycan [Low92]. glycans [HSM^{+}93, MR91]. glycine [WB94]. glycine-leucine-phenylalanine-glycine [WB94]. glycoconjugates [PHMH92]. glycogen [HODF94]. Glycolipid [Wat90, LFG^{+}90, SPRvD91, SHH94, vGvMS^{+}91]. glycolipid-anchored [LFG^{+}90, SHH94]. glycolipid-binding [SPRvD91]. glycolipidic [MB93]. Glycolipids [DB90, NS93]. glycolithocholate [WMB^{+}94]. glycophorin [OKH92]. glycophosphatidylinositol [MC91c, MC91d, MC92]. glycophospholipid [Car91, RYKA90, RYK^{+}90]. glycophospholipid-anchored [RYKA90]. glycophospholipid-linked [RYK^{+}90]. glycoprotein [AES94, ATBD90, ATD^{+}90, BMTTG94, BS94b, BKR90, BWCG93, CdJC93, EC93, FTT^{+}93, GAHH94, GBS90, GG92, GMB^{+}91, HH94a, HM92a, HDC^{+}91, HWF92, KL90, LLIV94, LMIV93, LB93, LTPB92, MMF92, MGS90, MAB^{+}90, MVM91, MSD^{+}92a, NLFRB91, NCM^{+}93, OESC91, OBBS90, PLCB91, PLS92, PRL^{+}94, PFSS92, SH90b, SL90a, SC90b, SGM^{+}90a, SDPH90, SPB90, SOY93, TOS^{+}94, Wat90, WST^{+}90, YOWM91, ZPS^{+}92, CKGC91]. glycoproteins [AVP92, BS94b, BPS90, CLK^{+}91, CMK^{+}90, HWF93, KS90, KTR90, KKES91, MPVL92, MHV92a, MSLS91, POGS90, SR92b, SMBG92, WF90]. glycosaminoglycan [HMY92, SFSG91]. glycosaminoglycans [BMB^{+}92]. Glycosome [BSD^{+}92]. glycosomes [KKG^{+}91]. Glycosphingolipid [KYA^{+}94]. Glycosyl [VR92, MMC91, MDKB94, SSTL93, SWFB91, VM93]. glycosyl-phosphatidylinositol [MMC91, VM93]. glycosyl-phosphatidylinositol-linked [SSTL93]. glycosyl-phosphatidylinositol-specific [MDKB94]. glycosylate [KMSW91]. glycosylated [KA90]. glycosylation [AYM^{+}93, FT93, FPPD91, IDT^{+}92, SSW94]. Glycosylphosphatidylinositol [ZLC^{+}93, KGA^{+}93, MWLC^{+}94, MC94, ZCS^{+}91]. Glycosylphosphatidylinositol-anchored [ZLC^{+}93]. glycosylphosphatidylinositol-linked [ZCS^{+}91]. glycosyltransferases [IDT^{+}92, NPH^{+}93]. glyoxylate [NFYI93, PTD90]. glyoxysomes [CT91, KKG^{+}91]. GMP [LPM^{+}91, MVM91, PZP^{+}91, ZMS^{+}91]. GMP-140 [LPM^{+}91, MVM91, PZP^{+}91, ZMS^{+}91]. Go [BS94c]. gold [FMR^{+}92, LZI^{+}93, MWH^{+}91]. gold-labeling [FMR^{+}92]. gold-tagged [LZI^{+}93, MWH^{+}91]. goldfish [PLS92]. Golgi [AYM^{+}93, ABR^{+}92, AES94, BLR^{+}94, BBMN94, BHD^{+}94, BKR90, CH91, CP90, CWBB93, CFG^{+}94, CB92, CTPP92, DAW^{+}94, DP90, DP91, DLSB^{+}90, DLSK91, EH94, FB93, FTB94, FRC^{+}91, GtHG^{+}94, GP91, GE91, GSP^{+}94, GAS90b, GK92a, GEKL^{+}94, GRB^{+}90, GVK94, HH94a, HP92, HXM^{+}94, HPR93, HWvF90, HWF92, HWF93, vtHvM90, HPYB93, HEF92, JK93, JKB^{+}92, JGK^{+}94, JK92, JCSH93, KD92, KA91b, KLERG94, KRB91, KDO^{+}94, KDP^{+}92, LKF^{+}94, LSGD^{+}91, LTPB92, MP94, MSS^{+}94, MJG^{+}91, MSW^{+}94, MSPB90, MH93b, MCM92b, MCJ^{+}93, MW94, MSLS91, MR91, NMP^{+}92, NHK^{+}94, NPH^{+}93, NRS93b, NDM^{+}94, ONN91, ON94, OHC^{+}93, ODK^{+}93, OTWH92, PMKH91, PPZ^{+}93, PNPB94, PGAR90, PNM^{+}94, PCP^{+}91, PDSB92, PRAK94, PRL^{+}94, PHSvD92, PDC91, Ral93, RB92, RHF91, RDH^{+}92, RSS^{+}94, RTC^{+}94, Rob90, RLWK93, SPHvD91, SRG^{+}94, SP94]. Golgi [SGBN^{+}93, SPBB92, SMKH91, SRHK94, SP92, SNFN91, SBW90, SNH^{+}94, SYCA94, SDE^{+}93, SDR^{+}94, SFS90, SPW93, SdAN^{+}91, SM91c, TBFP92, TOA^{+}93, TGD93, TBKF^{+}92, TOK93, TIA^{+}92, VYM93, VSB^{+}94, VHE93, WCR92, Wat90, WHLW90, WRHW92, WW94a, WMG^{+}93, WSM93, WL91, WCC^{+}90, WNM^{+}94, WB92a, WLH92, WGTCY94, XS93]. Golgi- [WB92a]. Golgi-associated [CB92, MSPB90]. Golgi-derived [TOA^{+}93, FB93]. Golgi-specific [GE91]. gonadotropin [SSKT^{+}92]. gondii [BDMPJ94]. gonorrhoeae [HRT^{+}91]. governing [CCW^{+}93, ZCS^{+}91]. GP [KFW^{+}91]. gp150 [BLSR92]. gp150/95 [BLSR92]. gp210 [WB92c]. gp74 [AES94]. GP85 [LFB94]. gp90MEL [ITSR90]. GPI [KGA^{+}93, MWLC^{+}94, HLRBE93, KHW93, KGA^{+}93, LTS93, MNV93, MWLC^{+}94, RLS92, VM94]. GPI- [KHW93]. GPI-anchored [HLRBE93, MNV93]. GPI-linked [LTS93, RLS92]. GPIIb [BB93]. Gq [LHS^{+}93]. gracilis [MB92, RMB90]. gradient [GT94]. gradients [HFN92, KJ90]. granular [CDS^{+}91, DY93]. granulation [DGGG93]. granule [CC90b, HFS92, KCB^{+}92, KA92, LCT^{+}93, MEM94, MGDS93, SLH92, SNRS93, TFTH91, dTF90]. granule-associated [MEM94]. granules [BHLA90, FCFL94, GdLMS94, GI93, JGK^{+}94, KA94, MJM92, MOT90, RMG90, ZLG^{+}90]. granulocyte [LKXS93, WHN^{+}94]. granulosa [HSHA90, SSKT^{+}92]. Granzymes [GI93]. greatly [OC91]. group [JD91, KHHC93, PBM^{+}94]. groups [GFW92]. growing [BSB92, DGGG93, OH92, OH93, PLS92, ZZP94]. grown [Smi94]. Growth [HLB90, LSP^{+}93, YTN^{+}94, AC93, AUT^{+}91, ALE^{+}92, APP^{+}93, BML^{+}90, BMS92, BG91a, BMK^{+}90, BGB94, BGLJ^{+}93, BCET^{+}92, BIPG91, BBS92, BS94c, BD92, BSD94, BGRW91, BSR^{+}90, BDZ^{+}92, BFTR93, CG90, CSKZ94, CGW94, CXS^{+}94, CF94, DPS^{+}93, DPCP91, DDW90, DS90a, DJ93a, DGGG93, ET94, Eri93, EB92, FMR^{+}93, FGK^{+}90, FA91, FGSBZ93, FMR90, FAMR92, FSY^{+}91, FFS^{+}93, GMLD92, GBB^{+}94, GFWM92, GBOB90, GA91, HGDA93, HMLL94, HAK93, HAK94, IPF^{+}90, JS90, JK92, JGB^{+}93, JB93, JRSB90, JCO^{+}94, KT92, KSL^{+}94, KS92, KCR^{+}94, KHR91, KNR93, KBT92, KDT^{+}91, KVF^{+}90, LHS^{+}93, LZBH92, LHT^{+}91, LB92b, LB94, LF93, LCK^{+}92, LMP91, LK93b, LGPM90, MCM92a, MMG90, MTS^{+}91, MAWM90, MZRM93, MWG92, MMR91, MMT91, MJM91, MR94, NRS^{+}93a, OB93, OSKN94, OH90, OW90, PPHS^{+}92, PS94]. growth [PvBeHB^{+}91, PBM^{+}90, RLD^{+}91, RRH91, RDBHG90, SOE^{+}91, SR90a, SBB^{+}91a, SFM^{+}90, SH90a, SHR93, SK91, SCMU93, SYG91, SDK^{+}93, SSCD90, SWMS90, STF93, SR92b, SMWB93, SKK^{+}91, SR92c, SKD^{+}91, TMHKO94, TK91, TRLG90, TBL^{+}92, TISH94, TTR^{+}90, TFP94, THH^{+}90, TSR90, UK92, VBvG^{+}94, WLPB92, WNTT94, WSB93, WHT^{+}93, WG92b, WC93, WST^{+}90, WMS^{+}93, WTK^{+}92, WB93b, WG93, YGEG91, YM90b, ZBH94]. growth-arrested [FA91]. growth-associated [AC93, WC93]. growth-related [FGK^{+}90]. GRP78 [VMR90]. Gs [HC94, AJPB90, JJS92, LB92a]. GTP [ATSE90, ASE90, BRL^{+}93, BDW92, CRWS93, CG93, GAS90b, JvCH^{+}94, JGK^{+}94, KSK^{+}93, MSS^{+}94, MJG^{+}91, MPEG93, MM91, NPR^{+}94, ONN91, ON94, RG92, RDDR93, RPH^{+}92, SPBB92, TBFP92, TOA^{+}93, TBKF^{+}92, TSH^{+}93, YTN^{+}94, dWLS91]. GTP-binding [ATSE90, ASE90, CRWS93, GAS90b, JvCH^{+}94, JGK^{+}94, MSS^{+}94, NPR^{+}94, ONN91, RDDR93, SPBB92, YTN^{+}94, dWLS91, MJG^{+}91, TBKF^{+}92]. GTPase [HPP^{+}93b, KFC^{+}94, LJP^{+}93, MPEG93, SM92b, VRE^{+}92, ZJA^{+}94, ZBW93]. GTPase-activating [KFC^{+}94]. GTPases [PZB^{+}94, SKSD^{+}94]. guanethidine [KHTD92]. guanethidine-induced [KHTD92]. Guanine [DLSK91, PNPB94, HBP^{+}93, WNM^{+}94]. Guanosine [AYM^{+}93]. guidance [OB93, SOE^{+}91]. guinea [MFOI92]. gut [CSR^{+}92, CHG94, MDRG92, WG92b]. gut-specific [WG92b].

H [CS91, CMK^{+}90, DMB92b, DBSK90, JHGR93, LVA^{+}90, MR94, SZP^{+}93, TOFH94, VKS^{+}90, VWT^{+}91]. H-ras [VWT^{+}91]. H1 [FGS91, HKSL90, RTHB90, SNGG94, TTR^{+}90]. H2 [HKSL90, YL93]. H3 [RTHB90, SHM94]. H36 [SWF^{+}92]. H36-alpha [SWF^{+}92]. H4 [SBT93]. H5 [AGB^{+}91]. HA2 [HSvD93a]. HaCaT [BCG^{+}92]. haemagglutinin [CC90a]. hair [DEH^{+}91, FPR90, FMCR93, GH91a, KW93, MPR90, MSSD93, RBR90, ZCH90]. half [DA92, SCF^{+}93]. halos [GCML94]. hamster [BGGW^{+}94, CSC90, KD92, RHR90, TYF90]. Han [AGGS92]. hand [TASJ92]. Hansenula [WTH^{+}94]. haptoglobin [WF93]. haptotaxis [ASK^{+}90]. harboring [MGD^{+}92]. having [CLP^{+}94]. HB [BZF^{+}92]. HB-6 [BZF^{+}92]. hCG [GHGP^{+}92, GMGPM94]. hCG/LH [GHGP^{+}92]. HCO3 [SZP^{+}93]. HD1 [HNUO92]. HDEL [TFP94]. head [CK94, HHF93, SLK93, WCF92]. head-rod-tail [SLK93]. heads [FKT93]. healing [CGF^{+}93, EVL^{+}91, JRK^{+}94, KQJ91]. heart [GSV93]. Heat [BWV92, CLDV90, GVK^{+}93, HSCN94, IMK^{+}92, KES^{+}92, KAF^{+}93, LFR93, MVV^{+}91, PTTA90, SSG93, VGG^{+}93, WWS^{+}94]. heat-shock [CLDV90, WWS^{+}94]. heat-stable [KES^{+}92]. heavy [BP91, CWB93, CBM^{+}94, DAM^{+}93, Hen90, JH90, KA91a, KM92c, KB91a, KLBBK91, LPPCF91, LLK91, LFC94, LWB91, Mil90, NNHB^{+}92, OA92, PSRS94, PKG^{+}94, RCNO94, SMK91, STK^{+}93, SP92, SFO^{+}91, SJM^{+}94, WS90, WHW^{+}91, YKY92, dCTG92]. heavy-chain [CBM^{+}94, YKY92]. HeLa [GN92, NPH^{+}93, RK90a, TH92b]. heldup [BF91]. helical [MGD^{+}92, TU90, WBL^{+}91, WDB^{+}92, dCTG92]. helices [YKM^{+}92]. heliozoan [FCF92]. helix [SAJ^{+}94, TN94]. helix-loop-beta [SAJ^{+}94]. helix-loop-helix [TN94]. help [FRSL90]. hemagglutinin [BHLWH91, BR91, FZRH91, KHW93, SBSG92, SC90b, ZBS^{+}94]. hemagglutinin-neuraminidase [SC90b]. hematopoiesis [LWHL93]. hematopoietic [FAU^{+}93, KBM93, RLRR93, WK92]. hemichannels [KGDE94]. hemidesmosomal [HNUO92]. hemidesmosomes [CKG^{+}90, HNUO92, KQJ91, RGJ91, SJGG94, SCJ^{+}91]. hemolytic [DGF^{+}94]. hemopoietic [GBB^{+}94]. heparan [BSEF^{+}93, BBJ^{+}91, BGRW91, CS90, CES^{+}92, DLD^{+}90, DBV^{+}92, DvdSM^{+}92, DKM^{+}92, FMR90, HAK93, KT92, NHK^{+}94, OR92, SR90a, SWM^{+}92, SF91, SLL94, SdAN^{+}91, dAWS^{+}90]. heparan-sulfate [BBJ^{+}91]. Heparin [FMR90, HAK93, BSR^{+}90, DKM^{+}92, HMS^{+}90, HK91, JJ92, LEM93, MSQ^{+}90, SLKS94, SFM^{+}90, SLWF91, SOY93]. Heparin-binding [HAK93, BSR^{+}90, DKM^{+}92, JJ92, LEM93, MSQ^{+}90, SLKS94]. heparin-dependent [MSQ^{+}90]. hepatic [DAW^{+}94, GHBA^{+}93, JRK^{+}94, NBT94, WS91a]. Hepatitis [HEF92, KLERG94, LMW^{+}94]. hepatocarcinogen [BRD94]. Hepatocyte [BDZ^{+}92, CSRS^{+}90, GBB^{+}94, WS91a, WS91b, INM^{+}93, JCO^{+}94, Mac92, SMWB93, UK92, WSB93, WMB^{+}94]. Hepatocytes [DTY92, CKL^{+}91, DCH^{+}92, DYS^{+}90, KKG^{+}90, LFR93, NHK^{+}94, RSC^{+}93, SP94, SMSH92]. hepatoma [CHBGL91, GHBA^{+}93]. hepatoma-derived [CHBGL91]. hepatoma-human [GHBA^{+}93]. herbicides [SH91]. hereditary [DGF^{+}94]. herpes [SWM^{+}92]. HETE [TCDH93]. hetero [LTS93]. hetero-oligomeric [LTS93]. heterochromatin [PE93]. heterochromatin-associated [PE93]. heterodimer [HW90, KRdlR^{+}92]. heterodimeric [PP91]. heterodimers [HRC^{+}90]. Heterogeneity [ND91, TQL^{+}90, FAU^{+}93, RBD92, SEBH90, TISH94]. heterogeneous [CLK^{+}91, WMCG91]. heterokaryons [PB90]. heterologous [BGGW^{+}94, ML91, SDPH90, TKP^{+}91]. heterophilic [KSC^{+}91, MKCE92, MJ92]. heteropolymers [LXWC93]. heterotrimeric [HSCR94, NPR^{+}94, SdAN^{+}91]. heterotypic [HWS90, YDS92]. hexagonal [SKH90]. hexamers [QORM91]. hexokinase [LFF91]. HGF [CPP^{+}94]. hGH [LRSB94]. hierarchy [SNG^{+}91]. High [BMK^{+}90, GH91a, GA92, KK92, RSM^{+}94, SLG92, AHS^{+}92, CE94, DMLR^{+}94, FFW91, GT94, GWMP91, GRC^{+}91, HPM^{+}94, HAS92, HNUO92, HHL94, HNC^{+}92, LMW^{+}94, MAFR93, MR91, NFL90, RSS91, SFM^{+}90, SST92a, SRB92, SGW^{+}91b, VPC^{+}91, WIF^{+}90, RKB^{+}90]. High-affinity [BMK^{+}90]. High-frequency [KK92]. High-purity [GH91a]. high-resolution [SGW^{+}91b]. higher [GCML94, HODF94, HM91, KMBR90]. higher-level [GCML94]. higher-order [HM91]. highly [CHBGL91, CPD^{+}93, EOL^{+}90, KMBR90, KDS^{+}91, KWR94, OESC91, SNGG94, WHD^{+}91]. hinge [JO92]. hippocampal [DZdH^{+}93, FNV^{+}92, MTP^{+}92, MMD^{+}93b]. Histamine [TTR^{+}90, LPM^{+}91]. histochemical [DAMS91]. histocompatibility [CNJ^{+}94, DW91, PHB^{+}91]. histogenesis [ORD90]. histological [KEBD91]. histology [MHP90]. Histone [SBT93, AGB^{+}91, DDW94, RTHB90, SNGG94, SHM94]. histones [RTHB90]. HIV [BBB^{+}94, LCLG94, SZB^{+}94b, VLH^{+}93, WKD^{+}91]. HIV-1 [LCLG94, WKD^{+}91]. HIV-particles [BBB^{+}94]. HLA [LLK91, MBC90]. HLH [TN94]. HMG [HR94a]. HNF1 [GHBA^{+}93]. HNF4 [GHBA^{+}93]. HNK [MGS90, NS93, YOWM91]. HNK-1 [MGS90, NS93, YOWM91]. hnRNP [MMD92a, MMD92b, MMD93a, OOBJ92]. holds [DFLK94]. homeobox [BGLJ^{+}93]. homeostasis [RBB^{+}94, VPP^{+}93]. homing [BRWB91, BFL90, CKP91, ECAG93, ITSR90, ISF^{+}91, IRH^{+}91, TSLR90, WIF^{+}90, WIF^{+}91]. homo [HMLL94, OPL94, SJM^{+}94]. homo-multimer [OPL94]. homo-oligomers [HMLL94, SJM^{+}94]. homodimers [KB91a, LWB91]. homolog [BBMN94, CKMC94, CF94, MZRM93, MB94, VRE^{+}92]. Homologous [LKS94, CSP90, CPM^{+}91, HKC92, HDP^{+}93, JG94, KSE^{+}92, ML91, SAJ^{+}94, WVH94, YYY^{+}91, ZSH^{+}93, dAWL^{+}90]. homologue [CD91, CM94b, CKGC91, GAS^{+}90a, HHH^{+}93, KW92a, LSA91, MIU^{+}92, OUS^{+}93, dAWL^{+}90]. homologues [PS94]. Homology [PMPV92, FFR^{+}91, IKL^{+}94, WPW92b, dBBHA94]. Homophilic [BKFT93, FT93, MKCE92, ZFA^{+}93, EHB^{+}90, LVB^{+}94, RWG^{+}92, VR92]. homotypic [CPU^{+}90, HWS90, RPS^{+}92]. hookworm [RUH^{+}94]. hookworm-derived [RUH^{+}94]. Hormonal [SZP^{+}93]. hormone [AWH^{+}94, MWS^{+}93, QCCD92, SSKT^{+}92, vdSKL^{+}93]. hormone-dependent [SSKT^{+}92]. hormone-related [AWH^{+}94, vdSKL^{+}93]. hormones [GMGPM94]. horseradish [CFG^{+}94]. Host [DCB^{+}92, BDMPJ94, GHT92, SMHO92, TDT92, TDWT92]. HP1 [PE93]. Hsc70 [BHFS93]. hsp [BLW92, BMH^{+}93, KAF^{+}93, ZWW90]. Hsp10 [HH94b]. HSP27 [PTTA90]. HSP47 [NSHN92]. hsp70 [BWV92, RMK^{+}94, MF90]. HSP70-related [MF90]. Hsr [HTSP94]. Hsr-omega-n [HTSP94]. HT [CML90, LBL^{+}91]. HT-1080 [CML90]. HT-29 [LBL^{+}91]. HT29 [KBH91]. huge [EK92]. Human [ASNH93, BBK90, BLT^{+}90, BBJ^{+}91, EFD^{+}93, GYE^{+}90, HTGN94, HO94, KT92, KCC90, PH91, SHM94, VNS^{+}94, AO90, AW91, ATD^{+}90, AS94, ASL^{+}94, BRWB91, BLSR92, BFLS94, BPZ^{+}92, BCG^{+}92, BIPG91, BDB90, BTG^{+}93, BMWT93, BGRW91, BML^{+}94, CPU^{+}90, CWBK90, CSNZ92, CCG^{+}94, CWHH92, CYC91, CBE90, DLD^{+}90, DS90b, DG90, DvHB^{+}91, DSQ^{+}93, EK90, ECAG93, FMPP90, FSL90, FP91, FBS^{+}91, FJN94, GBP^{+}92, GWMP91, GCMDP91, GHBA^{+}93, HPM^{+}94, HD91a, HHH^{+}93, HKSL90, HCYS90, IBD92, JLC^{+}90, JTS^{+}91, JHK^{+}91, JSM^{+}90, JWHPM93, JSG^{+}92, JGO^{+}91, KFW^{+}91, KBM93, KR91, KYLV90, KCYH92, KO94, KHHC93, LQNRB90, LSP^{+}91, LTP^{+}92, LDNM92, LMB^{+}94a, LWF^{+}91, LSV^{+}94, LVB^{+}94, LNSSA93, LCLG94, LK93b, MBS90, MBC^{+}91, MLKB92, MM90, MHM91, MP94, MW93, MRH^{+}91, MCDA^{+}93, MPAF91, MNV93, MIU^{+}92, MVM91, MSD92b, MHDT94]. human [MPVB91, MMY^{+}92, MULA^{+}91, NS90, NNHB^{+}92, NHA^{+}92, OW90, OKMB91, PPL^{+}93, PA91, PWY^{+}92, PZP^{+}91, PYH^{+}90, PPFM90, PAL^{+}90, PSGE92, PPJF91, QvMWV^{+}91, QORM91, RZS^{+}94, RMFA92, RHB^{+}93, SCM90, SPD^{+}90, SST92a, Sch93, SDR^{+}90, SB93b, SSJ^{+}94, TMP^{+}91, TSBN94, TTR^{+}90, TND^{+}93, UK92, VBV^{+}90, VKRD94, WLZB93, WLWL94, WKD^{+}91, WG92b, WTK^{+}92, YGEG91, YKI^{+}91, YKM^{+}92, ZRW^{+}94, aUFQ^{+}92, dBBHA94, vdBRD^{+}93]. humans [MTS^{+}94]. HVEM [LKF^{+}94]. hyalin [AAM92]. Hyaluronan [BT92, HWLT94, Knu93, TAVC91, CNU92, HHO^{+}92, JKH94, KWGS92, SHS^{+}93]. Hyaluronan-binding [BT92, JKH94]. hyaluronate [CMK^{+}90, HLH^{+}92, LWV92, TBVS92]. hyaluronate-coated [TBVS92]. hyaluronic [LFB94, PHSA93]. hyaluronidase [LML^{+}94]. Hybrid [CHBGL91, PC90, PRB94, RH92a, SOY93, VVR^{+}90]. hybridization [BPTS94, DMLR^{+}94, DFC^{+}94, GCML94, PRC^{+}90, PJC^{+}90, SF90a, TLFS92, XL91]. hybridomas [STS91]. hybrids [GHBA^{+}93]. hydra [KHP^{+}91b, DS90c, SBH92]. hydrated [ASR^{+}90, GHH90, Woo94]. hydrogen [NXHMJ93, OSKN94]. hydrolases [EG91]. Hydrolysis [TOA^{+}93, BDW92, CRS94, CG93, HFN92, ON94, WRB^{+}94]. hydrophobic [CCR93, XMW^{+}93]. hydroxy [ROBN^{+}92]. hydroxyacid [ZVF91]. hyperglycosylated [AVP92]. hyperoxaluria [NFYI93, PTD90]. hypertonic [HSvD93a]. Hypertrophic [CGMC92, GBN^{+}93, QDT^{+}90, VIF^{+}93]. hypertrophy [BCET^{+}92, QCCD92, TFT94]. hypothesis [KRMG93].

i-3 [SdAN^{+}91]. I-cell [GK93a]. I-mediated [STZG91]. I2 [PMS92]. Ia [YLWS94]. Ib [YLWS94]. IC [BK90, DP92]. IC1 [MGW92]. ICAM [CPSS92, ACSM^{+}94, CdPA^{+}93, CSMdPSM94, CDS^{+}91, DSdF^{+}90, DS93, LMH^{+}94]. ICAM- [CPU^{+}90]. ICAM-1 [CPSS92, CdPA^{+}93, CDS^{+}91, DSdF^{+}90, DS93, LMH^{+}94]. ICAM-3 [ACSM^{+}94, CdPA^{+}93, CSMdPSM94, LMH^{+}94]. ice [FPGS91]. ICRF [ISN^{+}94]. ICRF-193 [ISN^{+}94]. identical [INY^{+}93, MC94]. Identification [AL92, AA94, AHW90, AM91, BMV90, BDB90, CS90, CYC91, EK92, EWP^{+}92, FL92, GSBS92, GP91, GWMP91, GF94, HDS90, HWW^{+}93, HNUO92, HSP93, HMT^{+}92, ISF^{+}91, JAS^{+}90, KO94, LDW^{+}93, LMIV93, LMM^{+}91, MF90, MK93, MSPB90, MBZL91, MK91b, MSF90, MSD^{+}92a, NMP^{+}92, NSWH90, OUS^{+}93, OMBL90, OS92, PLCB91, PHSA93, PJS^{+}92, PSGE92, PTD90, RWG^{+}92, RHH92, RBR90, RK90a, RKB^{+}90, RKB93, SB93a, Sch91, SYG91, SRP90, SGM^{+}90a, SSSL94, TN94, TTE^{+}94, TRV^{+}90, TSR93, TRP^{+}93, VHB93, WHLW90, WC90a, YM90a, ATD^{+}90, BKSB94, BBMN94, BOH^{+}92, BPS90, BWCG93, DAR93, EBS91, EZF^{+}94, FWF90, HGW^{+}91, HH90, KIS^{+}94, KKBL92, KWS94, KFN93, LKF^{+}94, LM91, LJHC94, LK93a, MPEG93, MR94, OOM90, PSE^{+}90, SJMK93, SOH^{+}90, SKK^{+}91, SOY93, ZVF91]. identified [AU90, CG91, HAS92, RKA93, WW94b]. identifies [CH92, RHF91, SB91, VBV^{+}90]. identify [HYD93, HSB^{+}94, PPD92, PKG^{+}94, SK94b]. IF [MCF91, MKSF93]. IFAP [SJGG94]. Ifi [CL92]. IFN [CPD^{+}93]. IFN-gamma [CPD^{+}93]. Ig [BGM^{+}91, ZFA^{+}93]. IgA [AKM94]. IGF [CSKZ94, JGB^{+}93, LVTB93]. IGF-binding [CSKZ94]. IGF-I [LVTB93]. IgG [CPA^{+}93, WIF^{+}90]. IgM [KYA^{+}94]. II [MGD^{+}92, BK90, BFR^{+}92, BKR90, BPH^{+}90, CNJ^{+}94, CW90b, CM94a, CE94, CGF^{+}93, EKK94, EVW91, EGW^{+}93, Ham94, HG92, HW90, HMLL94, HAG^{+}93, HM93, HSW90, HLWL92, ISN^{+}94, ISDM92, JK92, KISY91, KMBR90, KLC^{+}92, KMN92, KPAY91, KHHC93, MH94, MBC90, MMF94, MKG93, MR91, OTX^{+}94, PL90, PHB^{+}91, QXWN^{+}94, RAGC94, RKB^{+}90, RTHB90, RRH91, RZS91, SWL^{+}92, SMRG91, SPH^{+}91, SLW^{+}92, SM92a, SY92, SRP90, SC90b, SSS^{+}90, SJM^{+}94, SSB^{+}91, TDWT92, VHM^{+}93, VB93, WSvdK^{+}93, WS91b, WS90, WMF91, WGCdL91, WMB^{+}94, WE90, XMW^{+}93, YYM94, YSBM91, YKY92]. II-transcribed [ISDM92]. II2p11 [HG93a]. IIb [FTT^{+}93, GHS^{+}93, GBS90, HLC^{+}93, KFW^{+}91]. IIC9 [HE93]. III [EHB^{+}90, GAvdM91, HMLL94, HLWL92, HFS92, KHCR91, KPMG91, YYY^{+}91]. IIIa [BB93, FTT^{+}93, GBS90, KFW^{+}91]. IL-4-activated [LKD^{+}94]. IL-8 [KHHR92]. ileal [CHG94]. image [DWY^{+}93, MWD94, SGW^{+}91b, TU90, YDO^{+}90]. images [MOKF92]. imaginal [FJRAT91]. Imaging [KRCT93, GT93, GZC^{+}91, KII^{+}91, OOBJ92, SAH^{+}92, TLFS92]. immature [KA92, KA94, TFTH91]. immediate [BIPG91, MRSN93, SP92]. immediately [THL^{+}90]. immobile [HLRBE93]. immobilization [PNR^{+}93]. immobilized [KSC^{+}91]. immortalized [KBCM^{+}90, SSKT^{+}92]. Immune [GLAB93, REW90]. immunization [BLFQ93]. Immuno [SGG^{+}91, vGvMS^{+}91]. immuno-electronmicroscopy [vGvMS^{+}91]. Immuno-localization [SGG^{+}91]. immunochemical [SM90]. immunocolloidal [FMR^{+}92]. Immunocytochemical [GCZ^{+}92, LTPB92, RHGG92, YSBM91, GB92, JSG^{+}92, YAJ90]. immunocytochemistry [DFC^{+}94, NSH90, RSC^{+}93, VPP^{+}93]. Immunocytology [BPDM90]. immunodeficiency [BDB90, RHB^{+}93]. immunodetection [RELS^{+}91]. immunodominant [CdJC93]. Immunoelectron [SLP91, GHGP^{+}92, INY^{+}93]. Immunofluorescence [LB94, AC92, BK90, FMR^{+}92, SMC92]. Immunoglobulin [Hen90, AKK^{+}93, BP91, BHC^{+}94, CPA^{+}93, HSM^{+}93, LGSB93, OWF^{+}93, POGS90, SVR^{+}90, SR92b, VGMS91, VHW^{+}92, ZDR^{+}92]. immunoglobulin-like [HSM^{+}93, POGS90, ZDR^{+}92]. immunoglobulin-stimulated [CPA^{+}93]. immunoglobulins [STS91]. immunogold [BK90, SRV^{+}90]. Immunohistochemical [PSJ^{+}91]. Immunolocalization [DPND90, FFW91, KLBBK91, RHF91, AL92, DMLR^{+}94, YAJ91]. Immunological [ROBN^{+}92, EEC^{+}92, MBB^{+}93]. immunologically [FSK^{+}92, MPVB91]. immunoreactive [dCV90]. immunoselection [TA91]. immunosuppressive [PRB94]. impact [LG93b]. impacts [SLG92]. impairs [GCM^{+}94, SBG93]. impede [BMTTG94]. implicated [HSM^{+}93, LGSB93, RHH92, WWW^{+}92]. implication [VWT^{+}91]. implications [ASDC91, BJ92, CSJD91, FJRAT91, HIJ90, LTS93, LSV^{+}94, ROBN^{+}92, WTSP91, WPW92b, ZDP92]. Import [SSK^{+}90, AMG90, EJ93, GG92, HNvdK^{+}94, HPWN93, IMK^{+}92, JAB^{+}91, MMS93, MPEG93, MG91a, MC91a, MHDT94, NF90, RMK^{+}94, SBP90, SB93a, SMBG92, SS92b, VGG^{+}93, WWS^{+}94, WS93a, vZDF^{+}93]. important [FMS93, GS90, Ham94, SWMS90, VHB93, WPW92a]. imported [CL91b, MG94b, MG92, PC91b]. inactivating [GPS92]. inactivation [AM94, BO91, SWHCW91]. INCENPs [MECE93]. include [TIN^{+}93]. including [XMW^{+}93]. inclusions [EERM^{+}91]. incomplete [REW90]. Incorporation [WE92, DI91, GMML^{+}92, HKKO^{+}94, MVG91, WS92]. increase [ASR^{+}90, FMR90, KKBK92, LCK^{+}92, LASB^{+}93, SDR^{+}90, SY90b, WGCdL91]. Increased [TFC94, TFDS91, YH93, ALE^{+}92, BTRB93, BCB93, MSSD93, MHGC90, MDKB94]. increases [AW92, CLZ91, MM90, SDR^{+}90, WGTCY94]. independent [BSW^{+}92, BL94, CPU^{+}90, CLS^{+}92, GK93a, GK92a, GeKdV^{+}91, GA91, HGDA93, HSP^{+}91, HSvD93b, JM92, KCY92, KLR^{+}93, MSQ^{+}90, MH93b, NWD90, OHP^{+}94, PTMB94, TISH94, TSBS92, UK92, VYM93, WPS91, YCC92, ARWD93, DDW90, LBF92, LSGD^{+}91, RBD92]. independently [DP93, HMS^{+}90, SA90, SSJ^{+}94, VVR^{+}90]. indicates [CLZ91, LB92c, WBL^{+}91, WGD^{+}93]. indirect [BK90, BS94a, Bro91b, SMSH92, WYM^{+}92]. indistinguishable [SSVE91]. Individual [BB90, CMV^{+}92, Gau90, HBP^{+}93, MMD93a, MJM92, SW93a, WPS91, dBNI^{+}91]. individually [KFW^{+}91]. induce [BJ90, BS91, CNJ^{+}94, FMLD92, GKY^{+}94, HMC91, LSP^{+}93, LSR90b, MMJ^{+}90, PZP^{+}91, SW90, TKP^{+}91]. Induced [EVL^{+}91, Apg91, BPK91, BBvH^{+}94, BMW^{+}93, BCG^{+}92, BG94, CG90, CH93, CH91, CXS^{+}94, Cit92, DYV93, DH93, DZdH^{+}93, DCL^{+}92, DSJB^{+}94, FAMR92, FM94b, FTT^{+}93, GHS^{+}93, GCZ^{+}92, GBP^{+}92, GS90, GPF94, GBS90, GSR^{+}93, HN90, HSG94, HCS94, HVNC92, JvCH^{+}94, KVE92, KHTD92, LHS^{+}93, LCT^{+}93, LASB^{+}93, LJP^{+}93, MH94, MGB90, MWG92, MELD94, NR90, PDB^{+}91, PBD93, PBM^{+}90, PRC^{+}90, QPWG91, RBD92, STo^{+}92, SR92e, TTVV^{+}90, UYA^{+}91, VBAK91, Wal94, WRP^{+}91, YM90b, BIPG91, MIS93, PSN^{+}91, QPWG91, TSH^{+}93, WG92a, WGRW91, YGM^{+}94]. inducer [vdSKL^{+}93]. induces [AP92, AGGS92, ASNH93, BCBK92, CCG^{+}94, DA92, DGGG93, ECAG93, FFB90, FF94, Fri94, GBB^{+}94, GPS92, ISN^{+}94, KKL^{+}91, LD91, LCLG94, MB93, PSG^{+}90, PET^{+}91, RTHB90, SADJGP94, SW92, SBB91b, TS93, TH92b, TLSW93, WMS^{+}93]. inducibility [ZWW90]. inducible [BMH^{+}93, CL92, HJI^{+}93, LWV92, SH90b]. inducing [RJJ^{+}94]. Induction [ACSM^{+}94, BCET^{+}92, BML^{+}94, DBWS94, HSHA90, MBG93, SLB^{+}94, WNTT94, WHT^{+}93, BMS92, BSEF^{+}93, BZ90, FKB^{+}90a, GFWM92, KG94, LYK^{+}94, MSC91, OSKN94, RM94, STF93, TBT^{+}92]. Inefficient [TISH94]. inert [HLPL90]. infected [GRC^{+}91, Jas93, SZB^{+}94b, TSP90]. infection [DWY^{+}93, PK91, WKD^{+}91]. inflammatory [SFM^{+}90]. Influence [CLD^{+}91, LG93b, MSB^{+}91, TMI^{+}91, VGG^{+}93]. influenced [EUH91]. influences [NPST94, SWL90, UK92, WCM^{+}93, WF94]. influenza [BHLWH91, BR91, CC90a, FZRH91, GAS90b, KHW93, SBSG92, ZBS^{+}94]. influx [BGB94, CPA^{+}93, DSR^{+}93, DB90, KC90, TFDS91, WDT^{+}92]. information [DFLK94]. infrared [AB91]. inheritance [CSV^{+}92, CHW94, HCW94, MSTY90, MY92, SY94b, SY91]. inherited [VVR^{+}90]. inhibit [BLFQ93, FFB90, GFW92, HCW94, IMK^{+}92, SOH^{+}90, SFY^{+}94, WW94a, WCC^{+}90, WWD94, YSK90]. inhibitable [SSM92b]. inhibited [BSM94, GG92, RGT^{+}93, SSCD90]. inhibiting [SBN92]. Inhibition [AHS^{+}92, AGB^{+}91, DYV93, FAYM90, GS94, HGB^{+}91, ISN^{+}94, JvCH^{+}94, KL90, LCM^{+}94, MPEG93, MLRJ92, ON94, RZS^{+}94, SGN^{+}93, TSH^{+}93, VWT^{+}91, ZFF92, AUT^{+}91, BBvH^{+}94, DFC^{+}92, DMC^{+}91, HWS90, Lam94, LCK^{+}92, LTWH92, MJM91, MW92, SBWV90, SOKS91, WCW91, WST^{+}90, WB93b]. inhibitor [AW92, CML90, CJB^{+}93, EGP^{+}91, EMC^{+}90, KHR91, KVF^{+}90, LK93b, MVV^{+}91, MSL92, PNPB94, PSE^{+}90, QvMWV^{+}91, RUH^{+}94, RTHB90, SSG^{+}90, SPD^{+}90, SLP91, SWL90, UYA^{+}91, WBAP91]. inhibitors [Cit92, LTYY94, SH91, TBW92, TBKF^{+}92, WWD94]. inhibitory [KSK^{+}93, MGPPH93, PS94]. inhibits [AG93, BRL^{+}93, CRKA92, CPA^{+}93, CMB91, ELS93, Fow90, FMK^{+}94, HW92, HPWN93, JAB^{+}91, KHR91, LSGD^{+}91, LASB^{+}93, MCM92b, MM92, MCF93, NCBS90, NXHMJ93, NDM^{+}94, PMBL^{+}92, PMS^{+}91, SKT^{+}90, SFF93, VBvG^{+}94, Wal94, WGCdL91, YKI^{+}91, ZRW^{+}94]. initial [AR91, OMBL90, OBBS90, PKPM90, PTMB94]. initiate [SY94a]. initiates [DKM^{+}92]. Initiation [LM94, BS90, GHR^{+}90, JTS^{+}91, JRSB90, KT93, MG94a, RHB^{+}93, Smi94, WA93]. Injection [BBRE90]. injections [SN93]. injured [GJK^{+}91]. injury [HNH91, LCK^{+}92, MRBP^{+}90, MGB90]. inner [BEGN91, EJ93, KKM91, KIO^{+}94, MECE93, MOM^{+}92, PRSS90, PR91, PMS92, PMLM94, PPD92, PB90, RMK^{+}94, RBR90, SS92a, SB93c, WEB90]. inner-arm [KKM91]. innervated [CG90, LC94]. Inositol [GSV93, HN90, MM93b, QYC^{+}92, ES91, FNM^{+}92, GHR91, HFN92, MIS93, SJIC92, SRV^{+}90, WAS^{+}91, WCM^{+}93, YC93]. Ins [HFN92]. insensitive [PMLM94]. inserted [BDMPJ94]. insertion [DJ93b, FKG94, GFW92, HGW^{+}91, HFD91, HAG^{+}93, MLN93, PC91b, WW93]. insertions [FKB^{+}90b]. inside-out [OKF^{+}94]. insights [KDLS92, LKF^{+}94, Sch94, SL90b, BDH^{+}94]. insoluble [HBS92, HR94b, NI93, SSTL93]. InsP3 [GPF94]. instability [GPS92, SSS93, VMB92]. instructive [AA90]. insufficient [SY90b]. Insulin [BSHW91, JP94, RRH91, BSW^{+}92, BRF93, BCET^{+}92, CG90, CSKZ94, CPB^{+}93, DKL^{+}90, GCZ^{+}92, HSP^{+}91, HFS^{+}93, JGB^{+}93, KCDR91, MDKB94, PTS^{+}92, QCCD92, RPH^{+}92, SGG^{+}91, TBL^{+}92]. insulin-epidermal [TBL^{+}92]. Insulin-like [RRH91, BCET^{+}92, CG90, CSKZ94, JK92, JGB^{+}93]. insulin-regulatable [HSP^{+}91, PTS^{+}92]. insulin-regulated [JP94]. insulin-responsive [HFS^{+}93]. insulinoma [TGD93]. intact [BSHW91, BPZ^{+}92, CM94a, CVH91, GK92b, HN90, IDT^{+}92, KBP91, KGA^{+}93, LFR93, LB92a, MM91, ODPL93, RBB^{+}94, SHDS93, SC90a, SBR91, STA^{+}94]. integral [BWCG93, CSSL93, DvdSM^{+}92, FHI^{+}93, HWB93, HR94a, HP92, HVK91, KBG91, KDP^{+}92, MMF92, MEM94, PL90, PNB91, RB92, SBP90, SL90a, SR92a, SLL94, WCM^{+}93, WBR94, YNST93]. Integration [WFD^{+}91, TAWJ91]. Integrin [ABM^{+}93, BGLB94, BHHG90, KDS^{+}91, LSRC93, LHS^{+}92, OKF^{+}94, SHLS93, SLM^{+}90, SCJ^{+}91, WFCN94, ACC^{+}94, ASMC^{+}92, BBK^{+}93, BRA^{+}94, BKB^{+}90, BVS^{+}93, BBK90, BFLS94, CPU^{+}90, CSMdPSM94, CH93, CTQ91, DGL^{+}90, DS90b, DG90, DvHB^{+}91, DGAB^{+}93, DCCH92, EMGS93, EUH91, FKHG93, FAYM90, FTT^{+}93, FJN94, GHS^{+}93, GMM^{+}93, GSFS92, GSS^{+}92, GAvdM91, GAHB94, HRC^{+}90, HHR^{+}90, HW94, HAN^{+}93, HLC^{+}93, ILH^{+}90, ISF^{+}92, JTS^{+}91, JLL^{+}94, KYBC94, KWR94, KCYH92, KCC90, LTYY94, LBF92, LFWC92, LLSM92, LHT^{+}91, LYK^{+}94, LGSB93, LNSSA93, MHR94, MBC^{+}91, MH91, MH93a, MSD92b, NVR92, OPB90, PRF^{+}93, PH94a, RH92b, RHH92, RUH^{+}94, RA94, RPS^{+}92, SADJGP94, SMM90, SM93b, SEH^{+}93, SLWF91, SEAB91, SWF^{+}92, SSA^{+}90, TMP^{+}91, TYM^{+}92, TRS^{+}90, TWC93, VLH^{+}93, WKD^{+}91, WWD94, YCO^{+}93, ZMV^{+}93, dCQTR91]. Integrin-associated [BHHG90, LGSB93]. Integrin-cytoskeletal [SHLS93]. Integrin-dependent [LHS^{+}92, BFLS94, FJN94, WWD94]. integrin-ligand [ASMC^{+}92]. Integrin-mediated [ABM^{+}93, CSMdPSM94, LYK^{+}94, RH92b]. Integrins [TWC93, WOC91, AW91, BSG^{+}92, BLSR92, BHHG90, CKG^{+}90, CWBK90, DCBH92, GLAB93, HMY92, KFW^{+}91, KYLV90, KQJ91, LTYY94, LRDM91, LPA^{+}90, LKD^{+}94, MHYK92, NS90, SSFD93, STC93, VKR^{+}91]. integrity [BGH93, FBML90, FKB^{+}90b, KDP^{+}90, LRDM91, MULA^{+}91, SHOA92]. inter [GF94]. inter-repeat [GF94]. interact [BCKS92, BMH^{+}93, HMY92, PMS92, SLWF91, SMBG92]. interacting [PSRS94]. Interaction [DCBH92, SDH90, SNRS93, STC93, BEGN91, BR94b, CKP91, CSMdPSM94, CMB92, DVT92, EKK94, FCFL94, GCMDP91, HAK94, HCYS90, HBB94, JCO^{+}94, KWW^{+}94, KSC^{+}91, LCD^{+}93, LHT^{+}91, MVM91, MDRG92, NNHB^{+}92, OCMB91, OPB90, PB90, RMK^{+}94, RNG94, SM90, Sch94, STZG91, STF93, SBC^{+}93, SBB91b, TN94, YNST93]. Interactions [BVR^{+}93, LJHC94, MFS^{+}94, PBM^{+}94, PZB^{+}94, PFSS92, ASR91, ASMC^{+}92, BSG^{+}92, BKWD94, DHPG94, ELS93, FZRH91, FF94, GFM93, HPH^{+}91, HAK93, HNPN94, HYD93, HWD94, KKAS90b, LPA^{+}90, LF93, NHS^{+}94, NHTP90, RPPB94, SHLS93, SCS90, SR92b, SGM^{+}90a, TMI^{+}91, WDWT90, WBW^{+}94, WF94, YHM^{+}94]. interactive [SCMB92]. interacts [CdPA^{+}93, HD91b, KW92a, LTS93]. intercalary [FCF92]. Intercellular [CNZ^{+}92, SJIC92, STdL^{+}90, YYR92, AA90, ACSM^{+}94, BBvH^{+}94, CPSS92, DMB93, EFD^{+}93, JFA^{+}91, STo^{+}92, STC93, WJ92, ZMMP90]. intercentriolar [TKP^{+}91]. interchangeable [SEAB91]. interchromatin [FCFL94]. intercisternal [WCR92]. intercommunications [DDFC94]. intercompartmental [VGE90]. interconversion [CH93]. interface [BDGC^{+}93]. interfere [SBBS90]. Interference [ST94, SFF93, SS90a]. interferes [BRF93, LCM^{+}94, vZDF^{+}93]. interferon [CL92, CL92, ECAG93]. Interferon-alpha [ECAG93]. interferon-inducible [CL92]. Interleukin [VBvG^{+}94, BML^{+}90, PSE^{+}90, LK93b]. Interleukin-6 [LK93b]. Intermediate [BMF^{+}91, Fuc94, AES94, CLCC90, CL93, CCAF90, DSQ^{+}93, GMQ^{+}93, GEKL^{+}94, GRB^{+}90, HH94a, HODF94, HW92, IKFR91, KD92, KPMG91, KHF94, LMW^{+}94, LTPB92, MGG93, MVG91, MHGJ94, ON94, OMH93, OS92, RTC^{+}94, SLI^{+}91, SMKH91, SRHK94, SR90b, SJGG94, SDE^{+}93, SG92, SBC^{+}93, TFTH91, TTE^{+}94, VLGB92, WGD^{+}93, vdBRD^{+}93]. Intermediates [GK92b, BSM90, KBG91, SB93a, VM93]. intermembrane [CVH91]. internal [FM94a, LSP^{+}91, LCF92, MMWG94, NWvH94]. Internalization [MP94, BSW^{+}92, CPB^{+}93, EB92, GJJM94, GHGP^{+}92, JSM^{+}90, KFN93, KTR90, MPAF91, PJS94a, SL90b, SFY^{+}94, SWHCW91, TDSP93, YFS92]. internalization-defective [MPAF91]. internalize [BSHW91]. internalized [HSvD93b, KP90, WHK90]. internalizes [RYK^{+}90]. internally [Car91]. interneurofilament [NPST94]. Internuclear [PB90]. Internucleosomal [BG93]. interphase [BB94, BBRE90, CFFL93, DPND90, GPS92, JS92, LLDJ94, MECE93, PH92, SNGG94, WHW^{+}91, YSBM91]. interphotoreceptor [GFH90, aUFQ^{+}92]. interplay [MG94a]. Interpolar [MMDM93]. interruption [MSB^{+}91]. intersections [BPTS94]. interstitial [AA90]. intestinal [CSR^{+}92, CPD^{+}93, DJF93, DFLK94, FHGN91, GLQRB91, HKS^{+}92, vtHvM90, KRMG93, LQNRB90, LDNM92, MCDA^{+}93, MSF90, MDRG92, RRBG91, SZB94a, SMHO92, WG92b]. intestine [BGK^{+}94, RLB^{+}91, SABF^{+}90]. intimations [Koz91]. intra [CM94a, MSS^{+}94]. intra-cisternal [CM94a]. intra-Golgi [MSS^{+}94]. Intracellular [APH92, GC91, HSP^{+}91, KM92a, KM94, LD92, MCJ^{+}93, NHTP90, PHB^{+}91, QORM91, SAH^{+}92, SC90b, VPC^{+}91, AG93, BP91, BDMPJ94, CSC90, CS91, CCG^{+}94, DAW^{+}94, DCTG90, FKLS92, FHGN91, GJJM94, GC90, HPM^{+}94, HR94b, HNC^{+}92, IK94, IPF^{+}90, IBD92, JJB^{+}93, KZS94, KR91, LMB^{+}94a, LFC94, LVBK92, MHGJ94, MW92, MG91b, NHO^{+}91, PMPV92, PL90, PP91, PTMB94, PTS^{+}92, PTK^{+}93, RAGC94, SPRvD91, Sch93, SPvV94, TLXM90, TGD93, WS90]. intracellularly [AKC92, OTWH92]. Intramitochondrial [MC93]. Intramitotic [HRH90]. Intranuclear [CRK^{+}93, CSC92, HTSP94]. Intrinsic [LLDJ94, BMS^{+}91, BBJ^{+}91, MHV92b, WCR93]. introduces [SES94]. Introduction [SSKT^{+}92]. Introns [KM92d]. invagination [PDA93, PMS^{+}91]. Invariant [MBC90, OTX^{+}94, PHB^{+}91, RAGC94]. invasin [MH93a, TYM^{+}92, YFS92]. invasin-bearing [YFS92]. invasion [AA90, FSK^{+}92, LWF^{+}91, MYC92, OKH92, OCMB91, SOH^{+}90, TSR90]. invasive [AW92, CO91, SRR90]. invasiveness [BVF^{+}93, FBS^{+}91, WBVB90]. investigation [LZBH92]. investment [OBBS90]. involution [TBW92]. involved [AJL^{+}90, ASE90, BFM93, BMPW90, BS94b, BSP^{+}94, CKL^{+}91, CM93, FSK^{+}92, FRC^{+}91, FvBuHWS92, FFR^{+}91, GHR^{+}90, GK93b, HPP^{+}93b, IRH^{+}91, JP90, JCSH93, KHvdL^{+}92, LCMP90, LVTB93, MLR^{+}92, MZRM93, MW94, MR92, MPVB91, NSH90, NS93, PMM91, RWG^{+}92, SR90b, SGM^{+}90a, SS92b, TN94, TDT92, TZBV91, WST^{+}90, ZDO^{+}91]. Involvement [BMW^{+}93, LMH^{+}94, NSHN92, WWS^{+}94, FBL92, FIH^{+}94, GLQRB91, GAS90b, HSSC92, HE91b, HPYB93, KATS90, LB92c, MSC91, MELD94, OSKN94, PFSS92, SCM^{+}93, STM91, SNRS93, TKHM91, WGR^{+}92, WFD92, YTN^{+}94, dSBH93]. involves [BLR^{+}94, BS94a, BS94b, BL94, CS91, GHT92, GS94, HH94a, HMS^{+}90, MSQ^{+}90, SMM90, TISH94, ZFA^{+}93]. involving [LSK^{+}94, SBB^{+}91a]. iodothyronine [LCT^{+}93]. ion [GC93, HWL^{+}90, KSFG91, PSN^{+}91]. ion-transporting [GC93]. Ionic [CVH91]. Ip12 [KFC^{+}94]. Ip12/Bem2 [KFC^{+}94]. IP39 [RMB90]. irradiations [SSS^{+}90]. irreversibly [WW94a]. islet [BRF93]. islets [KA92]. isoenzymes [WKM^{+}92]. Isoform [HRB92, KLST93, BRA^{+}94, BOH^{+}92, CTQ91, DBSK90, EK92, HSP^{+}91, JH90, KDTB90, KOB91, SLHG93, VHB93, WBT^{+}94]. Isoform-specific [HRB92, KLST93, VHB93]. isoforms [ASDC91, AFA^{+}94, BBK92, BF91, CDS^{+}91, FRL93, Gau90, HLH^{+}92, HR90b, KB91b, MTS^{+}94, Mil90, PW93, SGVS^{+}93, SEBH90, SKSC94, SSC90, SAM^{+}91, WBE91, YRBBP90]. isografts [RRBG91]. isolated [BR94a, BMWT93, DJ93b, HNUO92, KGDE94, KKG^{+}90, KLBBK91, KA92, LFR93, LSV^{+}94, MLN93, OESC91, SWS^{+}92, SMKH91, SGM90b, WSSM91].

Isolation [AS94, BGH90, CSC90, DG90, DKG92, FCL^{+}94, GRB^{+}90, HBS92, JLC93, KCH^{+}94, KWR94, MMD92b, MR91, MLM^{+}93, MSL92, OKMB91, RGR90, RB93, SNH^{+}94, SKGU91, TYF90, VVB^{+}92, WST^{+}90, WW91, YCBM94, ZABM94, CHBGL91, GH91a, LHFV93, ODPL93, SFA90, WG92b, HSW90]. isomaltase [FHGN91]. isomerase [SXM94]. Isometric [KW92b]. Isoprenylation [BHS90, KN91, QPWG91]. isotypes [BRG^{+}90]. isozymes [SZB94a, UKH^{+}91]. IV [SJMK93, CMS^{+}90a, CL93, DSQ^{+}93, FBLL90, HD90, JZAVP93, LHT^{+}91, LWF^{+}91, MS94, PWY^{+}92, RSM^{+}94, SEBH90, SABF^{+}90, TRV^{+}90, VKR^{+}91]. IX [YYY^{+}91].

J1 [FWBS93, LVK^{+}91]. J1-160 [FWBS93]. J1-160/180 [FWBS93]. J1/ [LVK^{+}91]. japonicum [HWS90, HSW90]. JNM1 [MT94]. joints [SFM^{+}90]. jun [EZF^{+}94, SDR^{+}90, BMB^{+}92]. Jun/AP [BMB^{+}92]. Jun/AP-1 [BMB^{+}92]. junction [BGMM^{+}93, CSP90, CJO92, Cit92, DE90, FKG94, FSL90, GOP^{+}94, GKG90, Gum93, HSLF^{+}92, INY^{+}93, JG94, KKW^{+}93, LTP^{+}92, MLKB92, MLRJ92, MCEG90, PET^{+}91, PMG^{+}92, RGKG90, RB90, SEBH90, SAF90, SKGU91, TND^{+}93, TTVV^{+}90, WJ92, WHD^{+}91, ZJA^{+}94, ZSM^{+}93]. junction-associated [INY^{+}93, ZSM^{+}93]. junctional [HR90c, JFA^{+}91, JSA^{+}93, KSFG91, Lam94, LTP^{+}92, MTS^{+}91, MCEG90, MG91b, TIN^{+}93]. junctions [ASL^{+}94, CNZ^{+}92, Cit93, FRSL90, FHI^{+}93, FIH^{+}94, KGDE94, ML91, NEL^{+}91, SJGG94, TOA^{+}91, WFD^{+}91]. Jurkat [TSS^{+}92]. justify [LCF92]. juxtacrine [LPM^{+}91]. juxtamembrane [BSHW91, BSW^{+}92].

K-252a [UYA^{+}91]. K-pumps [JJB^{+}93]. K-ras [SZKM91]. K14 [WCF92]. K5 [WCF92]. Kalinin [RLKB91, RA94, KWW^{+}94]. KAR1 [VHSR92]. Kar1p [BR94b]. KAR3 [OMRM93, PSRS94]. KAR3-related [OMRM93]. karyogamy [KBL^{+}94]. karyophilic [IMK^{+}92]. kD [CSM90, AC93, ATD^{+}90, BEGN91, BBZ94, BHS94, BRO^{+}91a, BS94b, BGI^{+}90, BHHG90, BFTR93, CKB93, CJO92, CL92, CG91, DW91, DJ93b, DLSB^{+}90, DLSK91, EOL^{+}90, HFL94, HP92, HD91b, HR94b, IMK^{+}92, IYN^{+}91, INY^{+}93, JS92, KIS^{+}94, KT92, KBG91, KW92a, KOB91, KDP^{+}92, LM91, LWF^{+}91, LPA^{+}93, MWH^{+}91, MGG93, MVV^{+}91, MIU^{+}92, MW92, NMP^{+}92, NHK^{+}94, PC91a, PMPV92, PMM91, PNR^{+}93, RSM^{+}94, RBR90, RM94, RWL^{+}90, RMB90, ST94, SM93a, SCM^{+}93, SNFN91, TKHM91, WLPB92, WWS^{+}94, WCR92, WHLW90, WBL^{+}91, WST^{+}90, WB92a, YH93, YMW^{+}90, ZBW93, dSK90]. KDEL [TWQ^{+}93]. Keratin [MVG91, AO90, BO91, CF90, EK90, HW90, KRHY93, KHF94, KO94, LCF92, MKG93, SBC^{+}93, TFC94, WCF92]. keratinization [TFC94]. keratinocyte [BCG^{+}92, HW94, KAP^{+}93, KCR^{+}94, LPA^{+}90, STC93, WJ92]. keratinocytes [AW91, BBJ^{+}91, CKG^{+}90, DP93, DMC^{+}91, MBC^{+}91, RA94, TG94]. keratins [CCAF90, CHVF91, HW90, KRHY93]. keratocytes [LLO^{+}94]. Kex1p [CB92]. KEX2 [BLT^{+}90, RHF91, WF91]. KEX2-like [BLT^{+}90]. kFGF [FPPD91]. Ki [SDW^{+}93]. Ki-67 [SDW^{+}93]. kidney [BBGK90, KSFG91, KWS94, LGH91, ME91, NSWH90, SWS^{+}92, SSA^{+}90, SBG93, ZSH^{+}93]. KIF3A [KSYN^{+}94]. KIF4 [SON^{+}94]. killer [BHLA90]. killing [HPM^{+}94]. kilodalton [OKH92, WP93]. kinase [Apg91, BAP^{+}90, BFLS94, BDK92, CGW94, CXS^{+}94, Cit92, DYV93, DDW90, DVT92, DJA^{+}90, DFC^{+}92, DY93, DCL^{+}92, DMC^{+}91, EOL^{+}90, EPNN91, EERM^{+}91, FMD93, FCW91, FMLD92, GCZ^{+}92, GPF94, GFWM92, GSR^{+}93, GRFB94, HWLT94, HTP^{+}92, HC94, HPWN93, HSP93, HSUS90, HHG^{+}94, JO92, KISY91, KYBC94, KMN92, KLBBK91, LSP^{+}93, LCKW91, LHS^{+}92, Mah91, MZRM93, MBS91, OHF^{+}94, OBC^{+}92, PPL^{+}93, PIRB92, PvBeHB^{+}91, PMBL^{+}92, PBCK93, PLB^{+}91, PHWM92, PM93, RTHB90, RSW^{+}91, SPH^{+}91, SZB94a, SGW91a, SBN92, SGW^{+}91b, SSDK^{+}94, SLS^{+}90, SRK^{+}94, SFY^{+}94, SMWB93, SWHCW91, SES94, SR92e, TCDH93, TSBN94, UKH^{+}91, UYA^{+}91, VWT^{+}91, VJB^{+}94, WSB93, WRP^{+}91, WKM^{+}92, WHN^{+}94, WWD94, YSBM91, YGM^{+}94]. kinase-defective [HSUS90]. kinase/myosins [PHWM92]. kinases [BSP^{+}94, KO94, KA93, KHHC93, LSP^{+}93, LMB^{+}94a, MSC91, TFT94, TOA^{+}91, VDS^{+}92, ZS94]. kind [SDW^{+}93]. Kinectin [TYSS92]. Kinesin [AST^{+}92, FFKC94, HSYK^{+}91, RMMH93, RMR92, SRB92, BBKR94, FNV^{+}92, HPP^{+}93a, HSSC92, HE91b, HHLS92, KM92b, LB94, NNHB^{+}92, OMRM93, PSRS94, PSG94, RGT^{+}93, SFO^{+}91, TH92a, TYSS92, VMB92, WVH94, WHW^{+}91, WTS93, dCTG92]. kinesin-binding [TYSS92]. kinesin-dependent [KM92b]. kinesin-homologous [WVH94]. kinesin-like [BBKR94, HPP^{+}93a, OMRM93, PSG94, TH92a, WTS93]. Kinesin-mediated [FFKC94]. Kinesin-related [RMR92, HSSC92, HHLS92, LB94]. Kinetic [CLZ91, CFFL93, IK94, MG91a]. kinetically [Wat90]. Kinetics [FLUS92, BST^{+}90, NFL90, WYM^{+}92]. Kinetochore [MOMM92, AR91, BCYC94, CBE90, CBJER93, HS90a, JLC93, MAFR93, MWBD90, MS92, PC91a, RSCS94, SW92, SBBS90, SSS93, TCS^{+}94, WSSM91, ZMB91]. kinetochore-associated [CYC91]. Kinetochores [GGGE^{+}92, HBR90, RA90, GR93, HM90, RS94, SSH94]. Kinetoplast [FTPM^{+}94, PME93, RG94]. kininogen [AHS^{+}92]. Klp1 [BBKR94]. KLP61F [HPP^{+}93a]. KLP68D [PSG94]. klpA [OMRM93]. knockouts [Eri93]. known [BBD^{+}90]. KRE [BSHB90]. Kre6p [RPPB94].

L [GWMC90, WCF92, WKMG92, AFA^{+}94, CO91, DSR^{+}93, ECAG93, GHGP^{+}92, HWM^{+}93, HWW^{+}93, KIS^{+}94, LKD^{+}94, MHGC90, OGM^{+}93, SST92a, TOFH94, ZVF91]. L- [AFA^{+}94]. L-alpha-hydroxyacid [ZVF91]. L-cells [CO91, GHGP^{+}92]. L-selectin [ECAG93, KIS^{+}94, LKD^{+}94, SST92a]. L-type [DSR^{+}93, HWW^{+}93]. L1 [CJC93, FMK^{+}94, HDC^{+}91, HSM^{+}93, KKAS90a, KKAS90b, KRdlR^{+}92, KSC^{+}91, MMD91, RPH^{+}92, WDT^{+}92]. L1-dependent [KKAS90b].

L1-specific [HDC^{+}91]. L1/ [FMK^{+}94]. L16 [MPTW90]. L266D [CCR93]. L6 [VGRC^{+}92]. L929 [IK94]. label [SSM^{+}92a]. labeled [JWHPM93, KFN93]. labeling [FMR^{+}92, GSBS92, IK94, SRV^{+}90, WSvdK^{+}93]. labile [BSW91]. lace [LKF^{+}94]. lace-like [LKF^{+}94]. Lack [DPCP91, SWS^{+}92]. lacking [BGI^{+}90, CF94, dHRB^{+}93]. lacks [FF93, HW93b]. lactase [DJF93]. lactating [TRH^{+}92]. lactolysis [HGB^{+}91]. lactosaminoglycans [ZMS^{+}91]. LacZ [WDRM93]. laevis [ABD^{+}94, BHF90, CSMG90, DK92, DMM^{+}91, DDW94, FHvZ^{+}94, HN90, HFN92, KMF^{+}91, MKF91, RLMG92, SJIC92, SPvV94, SN93, SVK^{+}92, STdL^{+}90, SSB^{+}91, SR92e]. LAM [KSST91]. LAM-1 [KSST91]. lambda [HODF94]. lamella [HGS^{+}94, KKES91]. lamellae [DLMS91]. lamellas [FE90]. lamellipodia [LB92b, VFR94]. lamellipodium [DI91]. lamellum [VB93]. Lamin [BZT93, UHFG92, BHS90, CLCC90, DIDB^{+}91, EPNN91, FKS93, GMML^{+}92, HPH^{+}91, HPWN93, KN91, MMLG94, NWD90, OHF^{+}94, SB93c, SW93b, WEB90]. lamin-independent [NWD90]. lamina [Ang91, AHM94, BEGN91, PB90]. laminae [MS94, SEBH90]. Laminin [BHS91, WHM^{+}92, YCC92, dCQTR91, BS90, BGB94, CL91a, CSRS^{+}90, CMB91, CTQ91, DS90b, EUH91, EC93, GSFS92, GAvdM91, HRC^{+}90, ILH^{+}90, KT92, KSE^{+}92, KWW^{+}94, KCC90, LBF92, LLSM92, MLKB92, ML94, MWH^{+}91, RA94, SEBH90, SF90a, SY90a, SSW94, SMM90, SM93b, SLWF91, SLM^{+}90, SSA^{+}90, SOY93, TMP^{+}91, VNS^{+}94]. laminin-receptor [MWH^{+}91]. laminins [MS94]. Lamins [GG90, LK93a, MMLG94]. lampbrush [RMG90]. Large [THBW94, BSP^{+}94, CDS^{+}91, EERM^{+}91, ELZ93, JG94, OA92, SDW^{+}93, SB93b, ZLG^{+}90, aUFQ^{+}92]. Large-scale [THBW94]. larvae [SRM94, SJM^{+}94]. laser [MFOI92, OH93]. lasting [FFB90]. late [AEGG93, DUVV93, DM92, MTC^{+}91, MPK^{+}94, PDB^{+}91, RHF91, VHE93]. Latent [TMHKO94, FAMR92, FAS^{+}93, KNR93, LGPM90, STL^{+}90, SOA^{+}93, SCMU93]. Lateral [ZCS^{+}91, dBNI^{+}91, BFR^{+}92, BMS^{+}91, CLD^{+}91, ES91, FZRH91, JWHPM93, KPMG91, LZI^{+}93]. latrotoxin [TTVV^{+}90]. Lattice [CMV^{+}92, CRS94, GA92, KIN^{+}94, SKH90]. lattices [KDS^{+}91, RPH^{+}92, WRA93]. layer [LML^{+}94]. layers [HN90, MPR90]. LCb [AB90]. Lck [LMB^{+}94a]. LDL [MYM94, PYH^{+}90, TLXM90, YPH^{+}92]. LDLC [PRAK94]. LDV [WK92]. leading [CCAF90, EMC^{+}90, FE90, HGS^{+}94, KKES91]. leads [AC93, CCW^{+}92, HK91, KKBK92, MSSD93, PRB94, SSKT^{+}92, WMB^{+}94]. leaf [GYE^{+}90]. leaflet [VM93]. least [GWO^{+}93, KA93]. LECAM [BRWB91]. LECAM-1 [BRWB91]. lectin [BFL90, CB90, CMB91, CKD90, HSW90, ITSR90, MWH^{+}91, MVM91, WIF^{+}91, WS91a]. lectin-like [MVM91]. Leishmania [LEM93, MWLC^{+}94, MSD92b, SIR^{+}94]. length [ASR^{+}90, CLZ91, FFR^{+}91, FH90, HRH90, KWW91, MA90, VDS^{+}92, Woo94]. lengthens [HFN92]. lengths [SHBD90, WW93]. lens [IVBR93, KGDE94, MBHG91, MGG93, PET^{+}91, WBW^{+}94]. lens-specific [MGG93]. LEP100 [MMF92]. lesion [SST92b]. lesioned [MMW^{+}92]. less [BMF^{+}91]. lethal [GVK94, HYD93, SRM94, SJM^{+}94, WW94b]. Leu [GEKL^{+}94]. leucine [CCC^{+}94a, CCC^{+}94b, WB94]. leukemia [Apg91, LKXS93, MGPPH93, OSK^{+}92, TNT^{+}90]. leukemic [GBP^{+}92]. leukocyte [BLSR92, CPU^{+}90, CLZ91, DGAB^{+}93, DCCH92, DCBH92, KSST91, MSD92b, vKWHF94]. leukocytes [CCFZ92, CMZ90, CSNZ92, HJI^{+}93, KCYH92, SY90b]. leukoregulin [MRH^{+}91]. leukosialin [NXHMJ93, YNST93]. level [GCML94, GFWM92, LG93a, SK94a, SLG92, TOA^{+}91]. levels [CG90, DPCP91, DJF93, GKG90, HNC^{+}92, OH90, RRD93, SFM^{+}90, SST92a, TSR90, WE90]. Lewis [FWD^{+}92]. Leydig [GHGP^{+}92]. LFA [ACSM^{+}94, BLSR92, CPU^{+}90, CdPA^{+}93, CDS^{+}91, CLD^{+}91, DCCH92, LBH93, LMH^{+}94, SNG^{+}91, TSH^{+}93, vKWH^{+}91]. LFA-1 [BLSR92, CPU^{+}90, CdPA^{+}93, CDS^{+}91, DCCH92, LBH93, LMH^{+}94, SNG^{+}91, vKWH^{+}91]. LFA-1-dependent [TSH^{+}93]. LFA-1-mediated [ACSM^{+}94]. LFA-1/ICAM- [CPU^{+}90]. LFA-1/ICAM-1 [CdPA^{+}93]. LFA-3 [CDS^{+}91, CLD^{+}91]. lgp [HM92a]. lgp120 [HM92a]. LH [GHGP^{+}92]. library [KWR94]. licensing [Blo93, CDRL93]. life [KAG94, MG94a]. Ligand [MWH^{+}91, ASMC^{+}92, BRWB91, DCY^{+}94, DCBH92, EG91, EWP^{+}93a, FTT^{+}93, HWvF90, HLC^{+}93, ISF^{+}91, JPJF91, KSL^{+}94, KWR94, LMIV93, LGSB93, LCPAM94, MRBP^{+}90, MSD^{+}92a, RDH^{+}92, SPRvD91, TSLR90, WGM^{+}93]. ligand-binding [HWvF90]. ligand-dependent [JPJF91]. ligand-specific [EG91]. ligands [ARW^{+}94, CSMdPSM94, DGAB^{+}93, DM92, HCW94, LMH^{+}94, LLIV94, MWH^{+}91, SPKV94, WIF^{+}90]. ligase [EWP^{+}93b]. light [AB90, AB91, BHS94, DMLR^{+}94, DH93, DAMS91, ECM^{+}92, FBML90, FCW91, GTC^{+}90, Hen90, LCKW91, OCC94, SLW^{+}92, SLS^{+}90, SWMS90, SS90a, TWC94, WYM^{+}92, YYM94]. like [AR91, BPK91, BBKR94, BCET^{+}92, BLT^{+}90, BCYC94, CNJ^{+}94, CG90, CSKZ94, CCG^{+}94, CSMG90, COLL^{+}90, CLK^{+}91, FNM^{+}92, GBP^{+}92, GBN^{+}93, GPS92, HWB93, HPP^{+}93a, HAK93, HSM^{+}93, JGK^{+}94, JK92, JGB^{+}93, KSL^{+}94, KHCR91, KLM^{+}91, KCM^{+}93, KA92, LKF^{+}94, LMM^{+}91, MMI^{+}91, MVM91, OMRM93, PSZ^{+}93, POGS90, PSG94, PPJF91, RRH91, SKSD^{+}94, TH92a, TW93, VHE93, WIF^{+}91, WDB^{+}92, WTS93, ZDR^{+}92]. likely [HSB^{+}94]. LIM [CPB94, SCMB92]. LIM-domain [CPB94]. limited [HTSP94]. limulus [OD93, LWK91, LCKW91, SAJ^{+}94]. line [AUT^{+}91, AGGS92, BP91, BCG^{+}92, BT93, COLL^{+}90, FP91, GBP^{+}92, JSG^{+}92, KBCM^{+}90, LQNRB90, LDNM92, MTS^{+}91, MK91a, MWS^{+}93, NVR92, PRB^{+}91, RGJ91, WNTT94, ZWW90]. lineage [AFHDD90, CSR^{+}92, DJA^{+}90, KYK^{+}94, NDdC92, RSM^{+}94, WHN^{+}94]. lineage-specific [CSR^{+}92, NDdC92]. linear [SGW^{+}91b]. lines [CHBGL91, FMPP90, LTP^{+}92, Mil90, MCEG90, NEL^{+}91, QvMWV^{+}91, RAGC94, VGRC^{+}92, VFR94, YCBM94]. link [HHL94, HCS94]. linkage [JD91, POGS90, SSW94, TKP^{+}91]. linked [CHM^{+}92, CKD90, EOL^{+}90, GTC^{+}90, GORB91, HBS92, RLB^{+}91, RYK^{+}90, SSTL93, SMBG92, VFC^{+}91, ZCS^{+}91, HDC^{+}91, LTS93, RLS92]. linker [ATV^{+}93, ASR^{+}90, DDW94, EC93, SJGG94]. linkers [TOS^{+}94]. linking [BGI^{+}90, CPA^{+}93, FF93, SLPB93]. Lipid [GRC^{+}91, CS90, CDS^{+}91, CT91, CHG94, HFN92, vtHvM90, JKLG94, KFN93, KP90, MIS93, THBW94, VM93, WRP^{+}91, ZDO^{+}91, ZBS^{+}94]. lipid-anchored [CS90]. lipid-binding [CHG94]. Lipids [GXHS94, CT91, LZI^{+}93, VPVC91]. lipolytic [MB93]. lipolytica [LNS^{+}94, YMO92]. Lipoprotein [KHR91, CSC90, DDFC94, DCH^{+}92, GK92a, HPM^{+}94, KT92, MCF93, MTJM93]. lipoprotein-cholesterol [CSC90]. lipoproteins [GRC^{+}91]. liposomes [VBS^{+}93]. lipoxygenase [KVE92]. Listeria [HG92, TCP90, TDT92, TDWT92]. live [GXHS94, HBR90, SM91d]. Liver [BGK^{+}94, BRD94, DJ93b, DKL^{+}90, KKW^{+}93, LHFV93, MPK^{+}94, NBT94, SSRM90, SF91, SGM^{+}90a, SEMH92, TSS91, VVR^{+}90, ZABM94]. Liver-intestine [BGK^{+}94]. livers [YPH^{+}92]. living [BM91b, CBRW92, FBML90, GT94, KRCT93, LJJC93, PT92, SW90, TK91, WTSP91]. LLC [JPJF91, SdAN^{+}91]. loaded [MOKF92]. loading [Ada92]. Local [MHV92a, ZZP94]. Localization [AMMW92, BK90, BLR^{+}94, CFW93, CMM94, DP92, FRC^{+}91, FNM^{+}92, GEKL^{+}94, HSCR94, HHY^{+}93, HLCB92, HPYB93, KLC^{+}92, MGB90, NEL^{+}91, PSRS94, PG92, PJC^{+}90, SF90a, SMC92, SBK^{+}91, SZB^{+}94b, SSM^{+}92a, VGM^{+}93, dAWS^{+}90, APH92, AAM^{+}93, AC92, BKSB94, BBK92, BU94, BBK90, BR94b, BHD^{+}94, BKR90, CKB93, CLDV90, CL92, CTPP92, CCW^{+}93, DOW92, DFC^{+}94, EBS91, EKK94, ECM^{+}92, FBL92, FCD^{+}93, FIH^{+}94, FWBS93, GCZ^{+}92, GH91a, GBOB90, GC93, GSP^{+}94, HSHA90, HWW^{+}93, HSP93, IYN^{+}91, JG92, JSM^{+}90, JLL^{+}94, JSG^{+}92, KHP91a, KZS94, KEBD91, LTYY94, LKSR94, LSP^{+}91, LXM91, LMB^{+}94a, LB94, LSN^{+}94, LS92, MFOI92, MMM^{+}92, MB90, MMH^{+}92, MAB^{+}90, NSH90, NI93, PBM^{+}94, PH94a, PSJ^{+}91, PHB^{+}91, PSGE92, PRL^{+}94, PE93, RPH^{+}92, SKSC94, SLP91, SB93a, SST92b, SSDK^{+}94, SPvV94, SGG^{+}91]. localization [SES94, SR92d, SOKS91, SS92b, SLNN^{+}91, TWQ^{+}93, TSBN94, UKH^{+}91, WBA92, WBWH93, WS93b, WC90a, WV93, WHW^{+}91, YTN^{+}94, YSBM91, YAJ90, ZSH^{+}93, vGvMS^{+}91]. localizations [PHWM92]. localize [KLM^{+}91]. localized [BPTS94, BZT94, BBKR94, BSP^{+}94, CKGC91, Fuj93, HNH91, JSA^{+}93, JJS92, KAP^{+}93, KCY92, KM92d, KDTB90, LMSH94, MYT94, MJG^{+}91, MGPPH93, NSC^{+}93, PH92, RZS91]. localizes [DMJ^{+}94, HSYYK90, MM93a, ODK^{+}93, SS90b]. localizing [FHI^{+}93, WKS^{+}93]. located [BDGC^{+}93, CBE90, GOP^{+}94, HG93b, HTD90, PH91, SPD^{+}90, SXM94, SB93b, SDR^{+}94, SCJ^{+}91, TAWJ91, VM93, ZAH^{+}94]. Location [SGGC^{+}90, AA94, DCH^{+}92, GHS^{+}93, HSG94, JD91, PRSS90, WA93]. loci [PPD92]. locomoting [LLO^{+}94, PMM90]. locomotion [BL94, FE90, GPF94, GT93, SHS^{+}93, TAVC91, UNY90]. locus [PKG^{+}94, PMW92]. Long [KKBK92, ZZP94, DKG92, FFB90, KKL^{+}91, PRB94, SOY93, TWMS90, YLS92]. long-lasting [FFB90]. Long-range [ZZP94]. Long-term [KKBK92, PRB94, TWMS90]. longevity [LFG^{+}90]. loop [SAJ^{+}94, TN94, WGRW91]. loops [RMG90, WMCG91]. loses [Dav92, HXM^{+}94]. Loss [BVF^{+}93, VMR90, BCYC94, SWD94, SZKM91]. Low [FV91, GK92a, CSC90, Cit92, CVH91, DDFC94, KT92, KSFG91, MVV^{+}91, MCF93, MTJM93, VPC^{+}91, WG92a]. lower [KMBR90, MPVB91, NS90]. Lowering [CRKA92]. LSP1 [JBJH^{+}92]. LTB4 [GLAB93]. luciferase [GKKS90]. lumen [AYM^{+}93, ET91]. lumenal [GG92, VPC^{+}91]. Lung [JZAVP93, DLD^{+}90, HBR90, HSLF^{+}92, ML94, MS92, RA90, SSS93, SBG93, WHD^{+}91, WHH^{+}93]. lung-metastatic [JZAVP93]. Ly [SWFB91]. Ly-6A [SWFB91]. Ly-6A/E-mediated [SWFB91]. lymph [BRWB91, BFL90, WIF^{+}90]. lymphatic [MSB^{+}91]. lymphoblasts [GK93a]. Lymphocyte [JJ92, ARW^{+}94, CSMdPSM94, FGK^{+}90, ITSR90, ISF^{+}91, JBJH^{+}92, MMI^{+}91, NHH^{+}91, PYA90, RPS^{+}92, TSH^{+}93, TSS^{+}92, TSLR90, vKWH^{+}91]. lymphocyte-specific [JBJH^{+}92]. lymphocytes [BGH90, BZB^{+}93, BMGC91, CDS^{+}91, CPA^{+}93, GRFB94, GI93, IRH^{+}91, vdWvKvKdB^{+}92, LMB^{+}94a, PIRB92, STM91, WGM^{+}93, WKD^{+}91, vKWHF94]. lymphoid [ECAG93]. lymphokine [MRH^{+}91]. lymphoma [MHYK92]. Lys [GEKL^{+}94]. lysates [CLZ91]. lysine [BSR^{+}90, SR92a]. lysis [CCFZ92]. lysophosphatidate [JvCH^{+}94]. lysophosphatidate- [JvCH^{+}94]. lysophosphatidic [ZCP^{+}94]. lysosomal [BHD^{+}94, BHLA90, GAHH94, GK93a, HM92a, HGB^{+}91, JK92, LGH91, NLFRB91, OTWH92, PAL^{+}90, PDS90, RCNO94]. Lysosome [RSS^{+}94, JLC^{+}90]. Lysosomes [MPK^{+}94, GSMK94, HG92, HM92a, IBD92, MMF92, NLFRB91, NF92, PAL^{+}90, RS93, WF90, WB92b]. lytic [BHLA90, FSSA92, GI93].

M [GAHB94, Jas93, JS92, KGSB92, LOOM93, PLB^{+}91, SGBN^{+}93, VNS^{+}94, WC90b, ZB94]. M-phase [JS92]. m5S [AUT^{+}91]. m5S/1M [AUT^{+}91]. MA104 [CRKA92]. mAbs [DBV^{+}92, ESL^{+}90]. Mac [DSdF^{+}90, DGAB^{+}93, DS93, GAHB94, MSD92b, ZB94]. Mac-1 [DSdF^{+}90, DGAB^{+}93, DS93, GAHB94, MSD92b, ZB94]. machinery [Fro91]. macroglia [dCV90]. macroglobulin [ATD^{+}90]. Macromolecular [KCCG91]. macromolecules [HLPL90, SOPA94]. macrophage [HG92, WHN^{+}94, WGCdL91]. macrophages [CKP91, GeKdV^{+}91, HCYK91, MTJM93, RHGG92, RS93, SMM90, TLXM90, TMI^{+}91, WHK90, YR93]. macropinocytic [HPW94]. Macropinosome [RS93]. Madin [BBGK90, NSWH90]. mainly [ODK^{+}93]. maintain [FM94a]. maintaining [BDJ94, SDW^{+}93, TCS^{+}94]. maintains [PC91b]. maintenance [BFM93, BVR^{+}93, GW94b, LRDM91, WWN92]. Major [CNJ^{+}94, BGH90, BS94b, CKB93, CLDV90, CBE90, CTQ91, DGAB^{+}93, EGW^{+}93, HNUO92, HHL94, KO94, KKBK92, KOB91, MPVL92, MB92, MMD92a, MWLC^{+}94, OWF^{+}93, PHB^{+}91, PAL^{+}90, SLM^{+}90, SCJ^{+}91, TYSS92, Tur91, VKR^{+}91, YLWS94]. majority [CSNZ92, SPD^{+}90]. makes [KRB91, MHM91]. Making [KHF94]. Malaria [PHMH92, OKH92, SOH^{+}90]. malarial [GHT92, TRD^{+}92]. male [WHM^{+}92, ZWW90]. malignant [SGVS^{+}93]. maloriented [SMTW94]. Mammalian [GTC^{+}90, SJG93, AMG90, AM94, CH92, CBJER93, DFLK94, FBL92, FMR^{+}92, FFR^{+}91, GSC^{+}92, GJK^{+}91, GKKS90, GEKL^{+}94, GFV^{+}91, GVK94, HTP^{+}92, HM94, HHF93, HNP94, JCO^{+}94, KSVM92, MB94, MVTJ94, MWBD90, MC94, OHS92, PC90, PDWB92, RKB93, TWQ^{+}93, VKM93, WBA92, WGR^{+}92, YLS92, YMW^{+}90, ZBW93]. mammals [LCMP90]. mammary [KVF^{+}90, LTP^{+}92, LWL^{+}92, MELD94, RRD93, SVR^{+}90, SSCD90, SZP^{+}93, SBB91b, STA^{+}94, TBW92, TND^{+}93, TRH^{+}92, YSK^{+}94]. mammary-derived [KVF^{+}90]. man [SGVS^{+}93]. mandibular [SCSE90]. manipulation [CK91]. manipulations [FFKC94]. manner [KHS^{+}90, LBF92]. Mannose [GK93a, EG91, GK92a, GI93, HWvF90, JK92, KHV^{+}93, MH93b, MR91, SGBN^{+}93]. mannose- [GI93]. mannosidase [MR91, VHM^{+}93]. mannosylinositolphosphoceramide [PDC91]. mannosyltransferase [GSP^{+}94]. many [SLI^{+}91]. MAP [MK93, CSR^{+}92, OD93, EOL^{+}90, FMD93, IFLG90, LSP^{+}93, MZRM93, NFL90]. MAP1C [HSYYK90]. MAP2 [BM91b]. MAP2C [UOKH93]. Mapping [CHM^{+}92, CBJER93, HWD94, RHG90, SLKS94, BKSB94, FFSW91, KSST91, LS90, LCPAM94, RKB^{+}90, SEH^{+}93]. MAPs [BGGW^{+}94, GT93, LBH93, BGGW^{+}94, RKA93]. Marfan [KS94]. marginal [ZVF91]. mark [DY93]. marked [RTHB90]. marker [EH94, FPR90, LRR^{+}92, ZJA^{+}94, dBBHA94]. markers [DP93, VBV^{+}90]. marrow [BGRW91, EFTJ^{+}93, MMI^{+}91, PSE^{+}90, VBvG^{+}94]. MAS6 [EJ93]. masked [NKW^{+}92, SPKV94]. mass [HNUO92, MVV^{+}91, RHB^{+}90]. mast [ATSE90, ASE90, HBP^{+}93, KTG90, MOKF92, NPR^{+}94, dTF90]. Mastoparan [GHR91, WW94a]. Maternal [PMW92, KMN91, ORS^{+}90b]. mating [BSF90, GS90, GSS^{+}93, TDSP93]. matrices [LG93a, ZAH^{+}94]. Matrix [NT94, AAM92, ATBD90, AA90, ASK^{+}90, BKB^{+}90, BJ90, BNI^{+}94, BF93, BTG^{+}93, BTR92, BFTR93, CSMdPSM94, CLG^{+}92, CCG^{+}94, CB90, DGL^{+}90, DMJ^{+}94, EGP^{+}91, ELS93, FAYM90, FF94, GBN^{+}93, GL91, GFH90, HNP90, JGB^{+}93, JH93, KWS94, KSB^{+}93, vdWvKvKdB^{+}92, KBM93, Knu93, KCYH92, LTYY94, LSR^{+}90a, LJJC93, LZI^{+}93, LSR90b, LS92, MBC^{+}91, MRH^{+}91, MGPPH93, MR92, NYA^{+}91, NKW^{+}92, NR90, OLKD90, PSZ^{+}93, PAFC92, QvMWV^{+}91, RMK^{+}94, RPPB94, Sch91, SSCD90, SNH^{+}94, SF91, SNRS93, SB90, SSK^{+}93, TMHKO94, TBW92, TVG^{+}93, TLSW93, VIF^{+}93, VGG^{+}93, WTH^{+}94, WOC90, ZMV^{+}93]. matrix-associated [LSR90b]. matrix-binding [BKB^{+}90]. Matrix-bound [NT94]. matrix-degrading [KWS94, TBW92]. matrix-dependent [TLSW93]. Maturation [CS91, VSB^{+}94, DYV93, DFC^{+}92, Dun90b, DM92, FMD93, FGK^{+}90, FMLD92, IK94, JGL^{+}93, KMF^{+}91, KA93, LDT^{+}92, MMCH91, MSSD93, PNB91, PMA91, RS93, YKI^{+}91, YM90a]. maturation-promoting [DFC^{+}92, YM90a]. mature [EH94, HLWL92, IKFR91, VSVP93a, VGG^{+}93]. maxillary [APP^{+}93]. Maximal [DSQ^{+}93]. may [BMGN92, CPM^{+}91, FM94a, ISDM92, QYC^{+}92, RNS92, RRD93]. MCP [KGK90]. MDBK [KSFG91]. MDCK [AKM94, BPS90, BPH^{+}90, BMM90, CSS^{+}94, Cit92, GAS90b, HLRBE93, HC94, LSMRB90, LSP^{+}91, LTWH92, MYM94, NLFRB91, NI93, PKN91, PIS94, PHSvD92, SPRvD91, SSTL93, SZKM91, UK92, WNBAS90, WHF^{+}90, vGvMS^{+}91]. MDCKII [RK94]. MDM1 [MY92]. Mdm10p [SY94b]. mdx [MHP90, OC91]. mean [LCF92]. measured [FV91, GMNH93]. measurement [KPAY91, SAH^{+}92]. Measurements [ZBH94, Bro91b, GXHS94, dTF90]. measuring [LZI^{+}93]. Mechanical [WMS^{+}93, Ada92]. mechanically [LG93a]. Mechanics [FE90]. Mechanism [CW90a, CW90b, ES92, GCMDP91, LDNM92, LGPM90, BML^{+}94, CPU^{+}90, CB90, HIJ90, HVNC92, LSK^{+}94, LSGD^{+}91, PKPM90, RBD92, SJIC92, SOA^{+}93, SLW^{+}92, SCMU93, SSS93, STo^{+}92, TLWA94, WCR93, ZMMP90]. Mechanisms [Ada92, CCFZ92, Har92, JHGR93, BJ90, CCT94, DDW90, Dun90a, Dun90b, HJI^{+}93, KA94, LPK92, LKSR94, LSRC93, MP94, MMW^{+}92]. mechanistic [FTPM^{+}94]. mechanochemical [CSM90]. mechanochemistry [WHFM93]. medaka [FMJ91]. medalist [Blo91]. medial [SNH^{+}94]. mediate [ARW^{+}94, BKFT93, EHB^{+}90, GAvdM91, HJI^{+}93, HRC^{+}90, ISDM92, JO92, LHH^{+}91, MME93, NI93, OKF^{+}94, PRL^{+}94, PE93, SMWB93, SPB90, SVD90, TWC93, TFT94, VHSR92, WNBAS90]. mediated [AA94, AKM94, ABM^{+}93, BMB^{+}92, CSMdPSM94, CSS^{+}94, CRWS93, CRKA92, CO91, CXS^{+}94, CCG^{+}94, CSRS^{+}90, CKL^{+}91, DCM^{+}93, FSK^{+}92, FFKC94, FA91, FBS^{+}91, GCM^{+}94, GtHG^{+}94, GSS^{+}91, GeKdV^{+}91, GDKP92, HM92b, HK94, HWL^{+}90, HCYK91, HVNC92, IMK^{+}92, IKL^{+}94, JGL^{+}93, JAB^{+}91, JBIM94, JFA^{+}91, KGR94, KWW^{+}94, KSC^{+}91, LSRC93, LYK^{+}94, LFB94, LKD^{+}94, MNME^{+}93, MH91, MHT^{+}92, MRSN93, MVCC93, NIT94, NF90, PS94, QH94, QvMWV^{+}91, RLB^{+}91, RB94, RRCR93, RH92b, SR90a, SC90a, SOPA94, SGN^{+}93, SFY^{+}94, TWC94, WRP^{+}91, WHN^{+}94, WFCN94, YGEG91, YGM^{+}94, YKY92, ACSM^{+}94, BTSL91, GMGPM94, PMHS92, STZG91, SWFB91, WHH^{+}93]. mediates [ARWD93, AUT^{+}91, BS90, BHS91, BHD^{+}94, BFTR93, CLCC90, DS93, GSMK94, GRB^{+}90, HHO^{+}92, HAK94, JLC^{+}90, KDS^{+}91, LWF^{+}91, LEM93, MMI^{+}91, MSFW92, MYC92, VR92]. mediating [DG90, Har92, ISF^{+}92, PNM^{+}94, RHB^{+}93]. Mediation [GPF94, LHS^{+}93]. medication [SJIC92]. medium [BR94a, EERM^{+}91, HSvD93a, KSFG91]. medullary [SMJ^{+}94]. mei [CMM94]. mei-1 [CMM94]. Meiosis [TSBS92, FKS93, PPHS^{+}92]. Meiotic [TH92a, BWLK93, CMM94, FMD93, FKB^{+}90a, HLR92, KLC^{+}92, KMF^{+}91, LA92, SBK94, SBBS90]. meiotic-specific [CMM94]. MEL [DGF^{+}94, BFL90]. melanogaster [AS94, BF91, CBM^{+}94, HAS90, HDP^{+}93, KM92c, LCD^{+}93, MMD92a, MMD92b, MMF94, PTTA90, PMW92]. melanoma [CMS^{+}90a, DKM^{+}92, FSK^{+}92, ISF^{+}92, KFW^{+}91, KDS^{+}91, LK93b, QvMWV^{+}91, TTR^{+}90]. melanophores [RLGB94, SAH^{+}92, SR92e, TH90]. melatonin [SR92e]. melatonin-induced [SR92e]. member [CJB^{+}93, CMK^{+}90, FFR^{+}91, FNS^{+}91, JKH94, LDW^{+}93, LWV92, LGSB93, MMS93, SGWE94]. Members [YSK^{+}94, FWD^{+}92, ME91, TSK^{+}94, TOS^{+}94]. Membrane [BBD^{+}90, CAS^{+}92, FKG94, KIS^{+}94, LD91, LK93a, NRS93b, PYA90, RNS92, SWL^{+}92, SP94, TIA93, WG92a, WBH^{+}90, AES94, ATV^{+}93, AV91, AMMW92, AJPB90, BP91, BEGN91, BGBWO91, BTH^{+}90, BGH90, BCBK92, BDMPJ94, BCKS92, BS94b, BMM^{+}94, BPM93, BN91, BMGN92, BPS90, BHHG90, BBJ^{+}91, BDJ94, BWCG93, Car91, CSvdE93, CS91, CT91, CH92, CSSL93, CPA^{+}93, CG91, CB92, CKL^{+}91, CDRL93, CF94, DLD^{+}90, DvdSM^{+}92, DBM94, DB91, DJ93b, DLSB^{+}90, DLSK91, ES91, EJ93, FB93, FKS93, FBLL90, FABS90, FFB90, Fuj93, FHI^{+}93, GVK^{+}93, GAHH94, GdLMS94, GtHG^{+}94, GLQRB91, GBS90, GSR^{+}93, GHT92, GFW92, GVK94, HWB93, HH94a, HR94a, HP92, HM92a, HD91a, Has93, HMY92, HLH^{+}92, HGW^{+}91, HFD91, HAG^{+}93, HLL94, HHL94, vtHvM90, HVK91, HIJ90, HYD93, HNH91, HGS^{+}94, HdHD^{+}93, HPP^{+}93b, HKS90]. membrane [JWHPM93, KT92, KFN93, KBG91, KS90, KN91, KDLS92, KBP91, KGA^{+}93, KCB^{+}92, KP90, KKFD90, KES90, KKES91, KBL^{+}94, KDP^{+}92, LA92, LRR^{+}92, LQNRB90, LSMRB90, LD92, LFG^{+}90, LL91, LM94, LB92a, LML^{+}94, LGSB93, LSV^{+}94, LB93, MHG93, MBC^{+}91, MB92, MNME^{+}93, MMF92, MMC91, MPM93, MMLS92, MMWG94, MNV93, MBHG91, MHV92b, MEM94, MCM92b, MC91b, MC91c, MC91d, MC92, MOT90, MAAO92, MYC92, MPM^{+}94, MB93, NLFRB91, NS93, NCM^{+}93, NSWH90, NHA^{+}92, ND92, NDP^{+}91, PMKH91, PLCB91, PKN91, PL90, PET^{+}91, PC91b, POGS90, PSN^{+}91, PNB91, PMM91, PB90, PMD^{+}93, PRB^{+}91, PMM90, RMK^{+}94, RB92, RK94, RLMG92, ROBN^{+}92, RMB90, RYKA90, RLKB91, RLWK93, SK94a, SLPB93, SKH90, SY90a, SBP90, SBR93, SP92, SL90a, SVB94, SNFN91, SR92a, SR90b, SB93c, SK94b, SY94b, SCJ^{+}91, SLL94]. membrane [SBLV92, SBB91b, SM91c, STA^{+}94, TKHM91, TB90, TLWA94, TAWJ91, TSS^{+}92, TSS91, TM90, VPVC91, VM93, VGM^{+}93, WNBAS90, WF93, WCR92, WSM93, WRB^{+}94, WCM^{+}93, WFCN94, WST^{+}90, WDRM93, WF90, WWW^{+}92, WKN92, WB92a, WEB90, WB92c, WBR94, XMW^{+}93, YM90b, YLDB91, YDS92, YNST93, YSK90, YMW^{+}90, YKY92, ZCS^{+}91, ZRW^{+}94, ZDP92, dWLS91]. Membrane-anchored [LD91, BPM93, LD92]. Membrane-associated [LK93a, BMGN92, FBLL90, MBHG91, VGM^{+}93]. membrane-bound [CG91, TSS91]. membrane-cytoskeletal [ATV^{+}93, NSWH90]. membrane-cytoskeleton [MNME^{+}93, YKY92]. Membrane-enclosed [BBD^{+}90]. membrane-spanning [CH92, LGSB93, NDP^{+}91, SM91c, ZCS^{+}91]. membranes [AMMW92, BHFS93, CLD^{+}91, CWBB93, DLMS91, DP91, FTB94, GAHH94, GBOB90, GAS90b, HPR93, HNC^{+}92, JK93, KSVM92, KLBBK91, LLK91, LMSA91, LB93, MAWM90, MLN93, MSW^{+}94, NMP^{+}92, PDSB92, RSS91, SLH92, SPBB92, SL90a, SS90a, SF91, SdAN^{+}91, TSS91, TOK93, VYM93, VBS^{+}93, VL91, WGM^{+}93, WW94a, WW93, WHW^{+}91, YCC92]. membranous [HSYYK90, HSYK^{+}91]. meprin [KWS94]. meromyosin [SFO^{+}91]. merosin [CL91a, VNS^{+}94]. merozoite [SOH^{+}90]. mesenchymal [AW92, BCG^{+}92, BF93, MELD94, RNG94, SABF^{+}90, SHB90]. mesenchymal/epithelial [RNG94]. mesenchymal/myogenic [BF93]. mesenchyme [SMWB93]. message [ZSC91]. messenger [KLST93, KZS94, PZV^{+}91, SWD92]. messengers [JM92, ZBPV92]. Metabolic [RB90, ALE^{+}92, BLW92, HK91]. metabolism [ASG^{+}93, BMGC91, JKLG94, MRH^{+}91, MMD91, QYC^{+}92]. metabolites [THH^{+}90]. metalloproteinase [KWS94, MSC91]. metalloproteinase-transin [MSC91]. metalloproteinase-transin/stromelysin [MSC91]. metalloproteinases [AW92, TLSW93]. metanephric [KHaA92]. metanephros [RRH91]. metaphase [HS90a, MS92, SMTW94]. metaphase-anaphase [SMTW94]. metastasis [BLFQ93, HHH^{+}93]. metastasis-associated [HHH^{+}93]. Metastatic [QvMWV^{+}91, JZAVP93, MHT^{+}92, eSP91]. metazoans [ZKMB93]. methionine [HD91b, RWL^{+}90]. methionine-rich [HD91b, RWL^{+}90]. method [DMLR^{+}94]. methotrexate [LBL^{+}91]. methotrexate-adapted [LBL^{+}91]. methylation [KN91]. methylglutaryl [ROBN^{+}92]. methylotrophic [SS93]. methyltetrahydrofolate [RYKA90]. mexicana [SIR^{+}94]. Mg2 [GDKP92]. MgATP [HM92b]. MgATP-dependent [HM92b]. MgATPase [WHFM93]. MHC [HG92, OTX^{+}94, QXWN^{+}94, RAGC94, STM91]. mice [AO90, AWH^{+}94, AFHDD90, BSM94, BRF93, CSR^{+}92, DGF^{+}94, GSWW93, GSM93, HKS^{+}92, KB91b, LRSB94, MBW^{+}94, MGD^{+}92, MHGC90, MHP90, NDdC92, OC91, PYH^{+}90, PBB^{+}91, QvMWV^{+}91, RLRR93, RHG90, STS91, SBG93, WDRM93, YPH^{+}92, YHM^{+}94, aUFQ^{+}92]. microbeam [SSS^{+}90]. microbody [BSD^{+}92, KKG^{+}91]. microenvironment [KBM93]. microfibrillar [ZAH^{+}94]. microfibrils [BDGC^{+}93, KWGS92]. microfilament [BBS92, DES^{+}91, FBML90, SLPB93]. microfilamentous [SGN91]. microfilaments [BDH^{+}94, GIAS93, PTMB94, TLFS92, WLWL94]. microfluorimetry [FV91]. Microglial [GJK^{+}91]. microglobulin [LLK91]. Microinjected [MG92, SBBS90, AAM^{+}93, BHF90, CW90b, DK92, GMML^{+}92, WGR^{+}92]. Microinjection [FCW91, FFB90, PSG^{+}90, SW92, YYM90, HWL^{+}90, PB91, PLB^{+}91, SPvV94, WTS93]. Micromanipulation [TSS^{+}92, ZBH94]. micrometer [ES91]. microscopic [FKT93, GHGP^{+}92, MAFR93, SLP91]. microscopy [AR93, BK90, BZT93, DMLR^{+}94, DWY^{+}93, GA92, GT93, INY^{+}93, KISY91, KBL91, KII^{+}91, LFF91, MMM91, MFOI92, MWD94, NSH90, RRG^{+}94, RSS91, SS90a, TLFS92, TU90, WBR91, YDO^{+}90, dBNI^{+}91]. microsequencing [VBV^{+}90]. Microsomal [MYT94, LLK91, WW93]. microspikes [GAHH94]. Microtubular [GLQRB91, SFA90]. Microtubule [AM94, BSW91, Bro91b, CLM91, FCF92, MMM91, RMK91, RGT^{+}93, RLGB94, SOE^{+}91, TK91, VYM93, YCAB93, AV91, ABD^{+}94, BPK91, BRG^{+}90, BTS93, BGGW^{+}94, BSF90, BBKR94, BPH^{+}90, BM91b, BDK92, CRS94, DA92, FFW91, GPS92, GF94, HDDS94, HAS92, HM90, IFLG90, KIN^{+}94, KSYN^{+}94, LM91, LMSH94, LLDJ94, LELB90, MK93, MSC92, MWBD90, NW94, NSKM90, NFL90, PDB^{+}91, PDTG92, RMMH93, RK90a, RA90, SS94, SM91b, SH91, SS91b, SON^{+}94, SW93a, TZBV91, UOKH93, VMB92, VGC94, VBAK91, VDS^{+}92, WRDK92, WGP^{+}93, WDB^{+}92]. Microtubule-associated [BSW91, ABD^{+}94, BGGW^{+}94, BM91b, HAS92, LM91, MK93, NSKM90, NFL90, PDTG92, WDB^{+}92]. microtubule-based [AV91, KSYN^{+}94, SS91b, SON^{+}94]. microtubule-binding [HAS92, IFLG90, RK90a, WRDK92]. microtubule-bound [CRS94]. microtubule-bundling [FFW91]. Microtubule-dependent [RGT^{+}93, PDB^{+}91]. microtubule-nucleating [MSC92]. Microtubules [PBDK90, BB90, BJ92, BA92, BA93, BSW91, BRL^{+}93, BSB92, BM91b, BK91, CMV^{+}92, FMR^{+}93, FCF92, FRSL90, GMNH93, HDDS94, HBR90, HS90a, HE91b, JB93, LSR^{+}90a, MECE93, MMDM93, MOMM92, MSFW92, MPVB91, NNHB^{+}92, OH92, OH93, PC91a, PSHK92, SM90, SRB92, SW92, SBW90, Smi94, SSH94, SH92, UOKH93, VGC94, WPS91, WWWB90]. microvascular [CKD90, KYA^{+}94, KCC90, ND91, REW90]. microvesicles [CAS^{+}92, FABS90]. microvilli [AHW90, TSK^{+}94]. microvillus [RLB^{+}91]. microwave [BPDM90]. microwave-fixed [BPDM90]. mid [LSK^{+}94]. midbodies [NKW^{+}92]. MIF2 [BGH93]. might [TDT92, vDS94]. Migrating [AVP92, EMC^{+}90, HGS^{+}94, PSS^{+}93, SHLS93]. migration [AVP92, BTS93, BML^{+}94, CL91a, CPD^{+}93, DSQ^{+}93, GAHH94, GDKP92, HS92, HAK93, HMB^{+}93, HKS90, HFS92, KHR91, KES90, KHHR92, LTYY94, LFWC92, LSRC93, MM90, MHM91, MT94, MHYK92, MCF93, NYA^{+}91, PSHK92, PCD^{+}92, PH94a, SNRS93, TBVS92, TSR90, YM90b, YED^{+}93, ZMV^{+}93, vDS94]. migration-associated [AVP92]. MIIC [CNJ^{+}94]. MIIC-like [CNJ^{+}94]. Milieu [CH91]. Milieu-induced [CH91]. milk [RRD93]. MIM44 [RMK^{+}94]. mimics [LTYY94, PLB^{+}91, SWD92]. Min [MDRG92]. mineralization [GBN^{+}93, WIS^{+}94]. mineralize [KBCM^{+}90]. mineralizing [GL91]. Mini [KHP^{+}91b]. Mini-collagens [KHP^{+}91b]. minigenes [LRSB94]. Minimal [NDdC92]. minimally [BPZ^{+}92]. minus [BSM94]. Mis [WRA93]. Mis-assembly [WRA93]. misfolded [HH94a]. Misfolding [MH92, dSBH90]. mislocalization [CB92]. missing [BF91, KKM91, SMK91]. mistargeting [PTD90, SBR93]. mitochondria [BDJ94, CVH91, EERM^{+}91, HNvdK^{+}94, HTGN94, HH94b, LFF91, MLN93, SZB^{+}94b, SSK^{+}90, WBS^{+}91a]. Mitochondrial [HSCN94, RMK^{+}94, RBB^{+}94, BDJ94, EJ93, GVK^{+}93, HSHA90, HNvdK^{+}94, HTGN94, IKFR91, KCDR91, MMLS92, MSTY90, MY92, MC91a, SGW^{+}91b, SY94b, SY91, VGG^{+}93]. mitochondrion [PTD90]. Mitogen [TFT94, BIPG91, CXS^{+}94, EOL^{+}90, GHR91, GSR^{+}93, LVTB93]. Mitogen-activated [TFT94, CXS^{+}94, GSR^{+}93]. mitogen-stimulated [EOL^{+}90]. mitogenesis [CGW94, DCM^{+}93, LD91]. mitogenic [BSR^{+}90, RBD92, SK91, WMvdK94, YGM^{+}94]. mitogens [DS90a, GJK^{+}91, PA91]. Mitosis [LE92, MW92, BSF90, BBRE90, BCYC94, CFFL93, CC93a, CSC92, CC93b, Dav92, DBTAB90, DMWJ90, DS90c, FBL92, FCFL94, FCW91, FA92, HTP^{+}92, HPP^{+}93a, MHG93, MSPB90, MM93a, MS92, NKW^{+}92, NW94, SKBD91, UYA^{+}91, VHSR92, YYM94, YSBM91]. Mitosis-specific [LE92]. Mitotic [HGF93, MP91, PMS^{+}91, SM91a, WE90, AM94, BGH93, CM94b, CYC91, DPND90, DMM93, FBL92, FKS93, FKB^{+}90a, GN92, GG90, GPS92, Gol93, HRH90, HBR90, HM93, HHLS92, HLR92, JS92, KATS90, LCC^{+}93, MT94, MW94, NSKM90, OMRM93, PC91a, PH94b, RS94, RTHB90, RMR92, SM91b, SZB94a, SBBS90, SSS93, SSS^{+}90, TZBV91, VKM93, VBAK91, WKMG92, WSSM91, WHW^{+}91]. mitotic-like [GPS92]. mixed [CCT94]. mixing [KBP91]. mixtures [HLPL90]. MK [KHS^{+}90]. MMM1 [BDJ94]. mobile [JPJF91]. mobility [CLD^{+}91, FZRH91, JWHPM93, KAV93, LZI^{+}93, PT92, ZCS^{+}91]. mobilization [LHS^{+}93]. mobilizing [TTR^{+}90]. Mode [KB90, VB93]. model [BB94, BMS^{+}91, CHG94, INM^{+}93, IK94, JKT91, KQJ91, LCC^{+}93, LH93, RHP^{+}91, SCS90, ZMB91]. modeling [SCS90, SL90b]. models [ASR^{+}90, JT93, KRMG93]. modification [CKP91, GE91, IITG94]. modifications [FKS93]. modified [BBB91]. modifier [MDRG92]. modular [FWBS93]. modulate [BTSL91, DLSK91, LS90, PT92, SSM92b, WDWT90]. modulated [CPD^{+}93, HMY92, HR90c, MWH^{+}91, STdL^{+}90]. modulates [BRIA^{+}94, FBML90, GMW^{+}94, HNP94, PBM^{+}90, SBN92, TRLG90, WC93, vKWHF94]. modulating [SYO^{+}91]. Modulation [HM90, JKT91, LFF91, MRH^{+}91, PDS90, RGKG90, RDBHG90, ZCP^{+}94, BCB93, DvHB^{+}91, DAS^{+}92, DES^{+}91, FKHG93, MTS^{+}94, dlRKR^{+}90, dlR91]. modulator [LCPAM94]. Molecular [AMBN91, CES^{+}92, DLD^{+}90, DvdSM^{+}92, FTB94, FSL90, FWBS93, FKB^{+}90b, HHO^{+}92, HLH^{+}92, HSLF^{+}92, HHF93, HD90, ILH^{+}90, JG94, JP90, JD91, KSST91, LGSB93, MECE93, MBB^{+}93, NVWR91, OK92, ROE^{+}90, SEBH90, TMP^{+}91, TWQ^{+}93, AHS^{+}92, BGK^{+}94, BL94, Cit93, CPB94, FFW91, FCD^{+}93, GWMP91, GYE^{+}90, HPH^{+}91, HAS92, HSCN94, HNUO92, HWL^{+}90, HAA93, KBA92, KII^{+}91, KA94, LCMP90, LMW^{+}94, LGI^{+}93, MK93, MA90, MVV^{+}91, NFL90, OOM90, PRSS90, SNG^{+}91, SLS^{+}90, TSH^{+}93, TOS^{+}94, WBVB90]. molecule [AORB90, AMBN91, AUT^{+}91, BLR^{+}94, BTRB93, BGK^{+}94, Blo92, BZB^{+}93, BGM^{+}91, CPSS92, DCY^{+}94, DAS^{+}92, DGGR92, EFD^{+}93, EHB^{+}90, FvBuHWS92, FWBS93, GAS^{+}90a, HDS90, IRH^{+}91, JZAVP93, KKAS90b, KES^{+}92, KA90, LVB^{+}94, MBW^{+}94, MKCE92, MMI^{+}91, NGP^{+}91, NSWH90, NHH^{+}91, OVB^{+}94, PW93, PFSS92, RWG^{+}92, RCJ^{+}94, RLKB91, SCO^{+}91, SWD92, SPKV94, SF90b, STo^{+}92, VKRD94, VHW^{+}92, YYY^{+}91, ZDR^{+}92]. molecule-1 [CPSS92, DCY^{+}94, OVB^{+}94, RCJ^{+}94, STo^{+}92]. molecule-dependent [SWD92]. Molecules [HSvD93b, ASK^{+}90, BJ90, BS93a, BF93, CNJ^{+}94, CLK^{+}91, DMB93, DW91, DS93, FT93, FMK^{+}94, GMB^{+}91, GFM93, HLRBE93, HG92, KKAS90a, KT92, KRdlR^{+}92, KCM^{+}93, LSR90b, LMM^{+}91, MMJ^{+}90, MFS^{+}94, MCJ^{+}93, NIT94, PTMB94, PHB^{+}91, RAGC94, RR92b, SSTL93, SGW^{+}91b, SEMH92, TSS^{+}92, TIN^{+}93, ZLG^{+}90, ZFA^{+}93]. MOM19 [HNvdK^{+}94]. monitor [RSC^{+}93]. monitors [RSCS94]. monkey [MIU^{+}92]. monoamine [LSN^{+}94]. Monoclonal [FFSW91, MBS91, NYA^{+}91, ZSM^{+}93, BMK^{+}90, BLFQ93, CYC91, KYA^{+}94, KCYH92, NEL^{+}91, NSKM90, OSK^{+}92, RKB^{+}90, RMG90, ST94]. Monocyte [LKD^{+}94, SDR^{+}90]. monocytes [EMC^{+}90, FMPP90, LYK^{+}94]. monocytogenes [HG92, TCP90]. monocytoid [FMPP90]. monolayer [LRDM91, PCD^{+}92]. monolayers [CPD^{+}93, HMB^{+}93, KHHR92, MCDA^{+}93]. monomer [CSSL93]. monomeric [HTBF94]. monomers [SLKS94]. monooxygenase [MJM92]. Monospecific [LLIV94]. morhua [BSW91]. Morphogenesis [LR93, AGKC92, AMMW92, BR94a, BT92, BG94, CHG94, DSJB^{+}94, GS90, KHP91a, KFC^{+}94, STA^{+}94, TLSW93, WJ92]. morphogenetic [KYK^{+}94, PWY^{+}92, WBWH93, YKI^{+}91]. morphogenic [WSB93]. Morphological [PDSB92, CSCSA94, GHS^{+}93, SSFD93, SBB91b, WHT^{+}93, WGRW91]. morphologies [VVB^{+}92]. morphology [CSMdPSM94, FGSBZ93, FFB90, GTP^{+}92, LSGD^{+}91, LG93b, PSG^{+}90, RB92, SBWV90, SY94b, TW93, UK92, WS91b, WC93, WB92b]. mos [DPW91, DYV93, FKB^{+}90a]. mos-induced [DYV93]. most [VBV^{+}90, WBH^{+}90]. mosxe [FMLD92]. motheaten [STS91]. mothers [KAG94]. motif [GJJM94, HW92, HAS92, KN91, OWF^{+}93, PTK^{+}93, SR92a, VTSS94]. motifs [LXM91, MYM94, PSZ^{+}93, RBW93, SAJ^{+}94, YYY^{+}91]. Motile [RS94, HIJ90, JKT91, MGW92, RH92b, SFO^{+}91, TM92, dBNI^{+}91]. Motility [BSG^{+}92, CDS^{+}91, WTSP91, BRA^{+}94, BDZ^{+}92, CMS^{+}90a, CXS^{+}94, CCW^{+}92, FSK^{+}92, FGSBZ93, FFB90, GSS^{+}91, HHO^{+}92, KM92b, KYBC94, MBS90, OTT92, SSS93, TRLG90, TWC94, UK92, WSB93, WGCdL91, dHRB^{+}93]. motion [AR91, HLPL90, SK94a, SW92, ZCH90]. motions [FE90]. motoneurons [BGLJ^{+}93]. motor [Ang91, BH94, BCYC94, HLWL92, KSYN^{+}94, KGM92, LH93, LMSH94, RGT^{+}93, SON^{+}94, VMB92]. motors [FTB94, MP91]. Mouse [KRHY93, KW93, WHD^{+}91, AUT^{+}91, BO91, BSM90, BM91a, BLB^{+}91, BKW91, CLP^{+}94, CTQ91, CHG94, CLS^{+}92, DKM^{+}92, GWMC90, GeKdV^{+}91, HTHC93, HJI^{+}93, HHR^{+}90, HSLF^{+}92, HLCB92, HRB92, JFA^{+}91, KHS^{+}90, KBCM^{+}90, KRMG93, KMSW91, KLERG94, KEBD91, LPK92, LWL^{+}92, LWK^{+}93, LMIV93, LSS^{+}90, LOC^{+}90, MLS94, Mil90, MGDS93, MHP90, ME91, MTJM93, NGP^{+}91, NFL90, OLKD90, OSKN94, OH92, PSJ^{+}91, PSE^{+}90, RB94, RLD^{+}91, RSM^{+}94, RRD93, RRBG91, SEC^{+}93, SHR93, SSCD90, SBBS90, SMWB93, SSB93, TLXM90, TMI^{+}91, TFC94, TSBS92, TNT^{+}90, TRH^{+}92, TGCT90, VBvG^{+}94, VGRC^{+}92, WBE91, WC90b, WTRD92, YRBBP90, YED^{+}93, ZWW90]. move [LPK92, OH93, SRB92, WV93]. movement [BBRE90, BBGK90, CW90b, CGW94, CXS^{+}94, CLM91, HAS90, HLWL92, MS92, OOM90, SAH^{+}92, SKT^{+}90, SBBS90, SY91, TDT92, WSC94, WS90, ZBS^{+}94]. movements [BS94b, KM94, SK94a, dBNI^{+}91]. moves [SYCA94, ZDP92]. Moving [CM93, DP94, GR93, HNH91]. MPF [YM90a]. MPK1 [MZRM93]. MPM [KA93]. MPM-2 [KA93]. MPS1 [WGBB91]. MPS2 [WGBB91]. Mr [HWvF90]. mRNA [PSN^{+}91, AAM^{+}93, BML^{+}90, DK92, DE90, FCD^{+}93, FWBS93, HSG94, HNH91, KCH^{+}94, KMN91, KCR^{+}94, Koz91, LKSR94, MCDY94, MMW^{+}92, MKF91, OMS^{+}94, PBCK93, PAC^{+}92, PJC^{+}90, RH92a, SF90a, SHR93, SDR^{+}90, SABF^{+}90, SLNN^{+}91, SS90b, TQL^{+}90, WSL91, WV93]. mRNAs [BPTS94, FFB^{+}93, GKG90, HDDS94, HG93b, KA91a, KEBD91, MVTJ94, SMRG91, WHT^{+}93, WBS91b, ZKMB93]. MRP [LSP^{+}94]. mSP [LWK^{+}93]. mtr [KCH^{+}94]. mu [BKFT93]. mucin [KWW^{+}94]. multi [ECO93]. multi-filament [ECO93]. Multifunctional [BAP^{+}90, SES94, TRV^{+}90, WFD92]. multigene [PMG^{+}92]. Multilayering [SZKM91]. multimer [OPL94]. multimeric [BCKS92, LMW^{+}94]. multimodal [dTF90]. multinucleated [MHV92a, RR92b, Rot90]. Multiple [CRWS93, CH93, DBC90, DJF93, GAvdM91, HS90b, JTS^{+}91, MG91a, SPBB92, VIF^{+}93, YM93, CSRS^{+}90, FM94b, GHT92, GC93, GKKS90, GZC^{+}91, HMS^{+}90, KT92, KKW^{+}93, LGSB93, MSC91, MDRG92, PSZ^{+}93, PL90, PSJ^{+}91, PAFC92, SK94b, SVD90, TMI^{+}91, WBW^{+}94, YHM^{+}94]. multiply [TG94]. multipotent [GBB^{+}94]. multistep [KRMG93]. multisubunit [WWW^{+}92]. multivesicular [FFS^{+}93]. multocida [SBR91]. Murine [FCD^{+}93, AST^{+}92, ACFFL93, BFL90, DUVV93, DW91, DP93, FAU^{+}93, GAvdM91, HLH^{+}92, HCYK91, KAF^{+}93, MBW^{+}94, MMI^{+}91, MR91, NHH^{+}91, RCJ^{+}94, SABF^{+}90, WIF^{+}91]. muscarinic [WHH^{+}93]. Muscle [BF91, FW91, AHM94, AP92, BP93, BW91, BBK90, Bis90, BSM90, BM90, BM91a, BLB^{+}91, BKW91, CG90, CSKZ94, CK94, CYR^{+}94, CH92, CWB93, CQYD94, CB90, CMB91, DAM^{+}93, DGGG93, DSQ^{+}93, DE90, DZdH^{+}93, DAIS90, FAS^{+}93, FPP91, FAF^{+}93, FGH92, FSM^{+}93, FHI90, FKH^{+}93, FKB^{+}90b, Gau90, GW94a, GW94b, GFV^{+}91, GSM93, HAK93, HFAGM94, HTBF94, HWW94, Jas93, JAS^{+}90, JSA^{+}93, KB91a, KHR91, LH92, LWB91, LC91, Lut91, LSS^{+}90, LOC^{+}90, MMG90, MFOI92, MHV92a, Mil90, ND91, NEL^{+}91, NGB^{+}93, OC91, OESC91, OWF^{+}93, PRF^{+}93, PC90, PBD93, PDWB92, RRG^{+}94, Rot90, RB90, RFF92, SBB^{+}91a, STL^{+}90, SOA^{+}93, SBK^{+}91, SHBD90, SKM94, SBLV92, SAM^{+}91, TFT94, TS93, TWC94, TGR^{+}92, TFDS91, VGRC^{+}92, VCL93, VPP^{+}93, WYM^{+}92, WDWT90, WW94b, WMS^{+}93, WHK^{+}93, YRBBP90, YYM90, YAJ90]. muscle [YAJ91]. muscle-affecting [GW94a]. muscles [BOH^{+}92, BOM94, HE91a, KWW91, LC94, MHP90, NEL^{+}91, VSVP93b]. Mutagenesis [MPAF91, PMM91, BSR^{+}90, HW90, HKC92, HWL^{+}90, PM93, SVD90, WC93]. mutant [BGI^{+}90, CBM^{+}94, DBWS94, Dav92, FZRH91, FDS90, FKB^{+}90b, GN92, GE91, GVK94, HTP^{+}92, HBOV93, HW93b, JKS93, KD92, KC90, MOM^{+}92, MHYC90, MMS93, MK91b, NFYI93, OCC94, QORM91, RAGC94, RWG90, SMK91, STK^{+}93, SEC^{+}93, STS91, SLS^{+}90, VGMS91, WB93a, WS90, WGD^{+}93, WNM^{+}94, YCBM94]. Mutants [GFW92, LBH92, BCLM94, BF91, BHA^{+}94, CSC90, CBC^{+}94, CF94, DUVV93, FLUS92, GdLMS94, HHR^{+}90, HDC^{+}91, HSUS90, JJS92, KCH^{+}94, KKM91, KIV93, LCD^{+}93, LKS94, MSTY90, RRCR93, RHR90, TBT^{+}92, TBKF^{+}92, TYF90, VVB^{+}92, ZHL93, dHRB^{+}93]. Mutation [LA92, MH93a, CCR93, CSG92, DRJ^{+}91, ELS93, FGS91, KIO^{+}94, MPAF91, MDRG92, OGM^{+}93, PH94b, PMD^{+}93, SY94a, TYM^{+}92, WYM^{+}92, YMO92]. Mutational [AKK^{+}93, BM90, TASJ92, HGDA93]. Mutations [PMLM94, PKG^{+}94, SK94b, SHOA92, vdBRD^{+}93, AGKC92, FM94b, FHGN91, Fuc94, GWMC90, HSB^{+}94, LCF92, Low92, LMB^{+}94b, PPD92, PMW92, PTD90, SH91, SP92, TH92a, WW94b, WC90b]. Mutual [STF93, HKSL90]. Myb [LE92]. MyBP [OWF^{+}93]. myc [CCT94, HR90a, ME91, LE92]. mycelial [LNS^{+}94]. Myelin [GFL92, KHTD92, AAM^{+}93, ASDC91, FT93, KGSB92, MGS90, MAB^{+}90, MJM91, OBBS90, POGS90, PFSS92, SSWS92, SSC90, WDRM93, WF94]. myelin-associated [OBBS90, POGS90, PFSS92]. myelin-forming [KGSB92]. myelinated [KDTB90]. myelinating [OBBS90]. myelination [EMGS93, MHR94]. myelinogenesis [ASDC91]. myeloblast [NVR92]. myeloid [LLIV94, LMIV93, MSD^{+}92a, TNT^{+}90, ZMS^{+}91]. myeloproliferative [LWHL93]. Myf [BLB^{+}91]. Myf-6 [BLB^{+}91]. Myf4 [SBB^{+}91a]. MYO2 [JPS91]. Myo2p [BSD94, LB94]. myoblast [CWB93, CMB91, JRCW^{+}90, LCM^{+}94, LG93b, NCBS90, PW93, RB94, SLG92]. myoblast-extrinsic [CWB93]. myoblasts [CCT94, CLS^{+}92, EFD^{+}93, HG93b, KYK^{+}94, LSG92, PRB94, YRBBP90]. myocytes [FMR^{+}92, GSV93, LDS^{+}92, MM93b, SDCD93]. MyoD [CLS^{+}92, LCM^{+}94]. MyoD1 [BCG^{+}92, BEO90, Mil90]. myofiber [Ang91]. myofibers [PRB94]. myofibrillar [SLL^{+}90, SDCD93, VSVP93b]. myofibrillogenesis [DE90]. myofibrils [Gau90, SLL^{+}90, SGM90b]. myofibroblasts [DGGG93, Gri94]. myofilament [SHBD90]. myogenesis [Ral93, SLI^{+}91, SWF^{+}92, VWT^{+}91]. Myogenic [Mil90, AGGS92, BF93, BEO90, LCM^{+}94, MTS^{+}91, MHP90, OR92, VGRC^{+}92, YKI^{+}91]. myogenin [AGGS92, CHM^{+}92, CLS^{+}92, LCM^{+}94, Mil90]. Myosin [DE90, ECO93, HE91a, WYM^{+}92, WS90, ZDP92, AP92, BKSB94, BK90, BBK92, BSD94, CWB93, CSM90, CGF^{+}93, CBM^{+}94, DH93, DP92, DAMS91, EBS91, EKK94, EVW91, EEC^{+}92, ECM^{+}92, FBML90, FCW91, FDS90, FKH^{+}93, FKT93, Gau90, GTC^{+}90, Ham94, HM94, HWM90, HCKJ92, HTBF94, JPS91, JH90, KA91a, KM92c, KPAY91, KLBBK91, LCKW91, LB94, LWB91, LSS^{+}90, LOC^{+}90, MBZL91, Mil90, ORS90a, OWF^{+}93, OCC94, PM93, SKT^{+}90, SJG93, SLS^{+}90, SJM^{+}94, TWC94, VB93, WBA92, WDWT90, WGCdL91, WBR91, YYM94, YKY92, ZC92, DAM^{+}93, ECM^{+}92, FB93, RKB^{+}90, RKB93, SLW^{+}92, SRP90, WHFM93]. myosin-binding [OWF^{+}93]. myosin-defective [FDS90]. Myosin-I [ZDP92]. myosin-VI [HM94]. myosins [KB91a, PHWM92]. myotendinous [CSP90, DE90]. myotubes [BRO^{+}91a, FRL93, NR90, RH92a, RR92b, VBA90, YH93, ZSC91]. myr [BKSB94, RKB93]. myristate [CSS^{+}94]. Myristic [DRJ^{+}91]. myristoyl [DRJ^{+}91]. myristoylated [WG92b]. myristoylation [DKG92, JKLG94]. myristoyltransferase [DRJ^{+}91]. myristylated [PAC^{+}92, SR92a]. myristylation [dSK90].

N [FvBuHWS92, JWHPM93, KA90, WRB^{+}94, BTSL91, BZ90, BGB94, CES^{+}92, DAS^{+}92, DSR^{+}93, DSJB^{+}94, DRJ^{+}91, DKG92, FRL93, FMK^{+}94, GP91, GPRB93, JKLG94, KKAS90a, KKAS90b, KW92a, MGDS93, MR91, MJM91, NF90, SSW94, WRHW92, WHF^{+}90, WG92b, HTSP94]. N- [DSR^{+}93]. N-acetylgalactosaminylphosphotransferase [BTSL91]. N-acetylglucosaminidase [MGDS93]. N-cadherin [BTSL91, BZ90, BGB94, FRL93, KW92a]. N-cadherin-dependent [DAS^{+}92]. N-cadherin-mediated [BTSL91]. N-cadherins [DSJB^{+}94]. N-CAM [FvBuHWS92, KA90, FMK^{+}94, GPRB93, KKAS90a, KKAS90b]. N-CAM- [MJM91]. N-ethylmaleimide-sensitive [WRB^{+}94, GP91, NF90, WRHW92]. N-formyl [JWHPM93]. N-glycans [MR91]. N-glycosylation [SSW94]. N-myristoylated [WG92b]. N-myristoylation [DKG92, JKLG94]. N-myristoyltransferase [DRJ^{+}91]. N-syndecan [CES^{+}92]. Na/H [SZP^{+}93]. Na/HCO3 [SZP^{+}93]. NADH [PAC^{+}92]. NADH-cytochrome [PAC^{+}92]. NADPH [SAVK94]. NADPH-oxidase [SAVK94]. nafenopin [BRD94]. naked [HM91]. nanometer [SK94a]. nanometer-level [SK94a]. Nanovid [dBNI^{+}91]. NAP57 [MB94]. narrow [HGS^{+}94]. nascent [AR91, BWV92, DCH^{+}92, ECO93, HAG^{+}93, KBA92, MMD93a, TAWJ91, WSvdK^{+}93, WW93, XS93, YSK90]. native [KGDE94, Woo94]. native-type [KGDE94]. Natural [SPKV94, BHLA90]. Naturally [FHGN91]. nature [HKSL90]. nb [PBB^{+}91]. nb/nb [PBB^{+}91]. NC1 [TRV^{+}90, LHT^{+}91]. NCAM [ASR91, BSM94, FRL93, HNR90, HSM^{+}93, MJ92, RWG^{+}92, RDBHG90, SWD94, SA90, YYR92]. NCAM-dependent [BSM94]. NDC1 [WHC^{+}93]. NDC10 [GK93b]. near [AUT^{+}91, CRS94, KM92d, PC91a]. near-diploid [AUT^{+}91]. Nebulin [KWW91]. necessary [AJPB90, BAP^{+}90, BMK^{+}90, DCBH92, EBS91, EZF^{+}94, HTD90, MBZL91, NF90, PSGE92]. necrosis [FJN93, FJN94, HJI^{+}93, HK94, LWV92, NS90]. Nectadrin [KES^{+}92]. negative [BRA^{+}94, ELS93, MVM90, TA91, VIF^{+}93, WGRW91, WTK^{+}92, MWLC^{+}94]. Neisseria [HRT^{+}91]. neither [RBW93, VGMS91]. nemaline [MNS90]. nematocytes [KHP^{+}91b]. nematode [BW91, HDC^{+}91]. Neogenin [VKRD94]. Neomycin [ASE90]. Neonatal [LWB91, KOB91]. neoplasia [KRMG93, MDRG92]. neoplastic [HKS^{+}92]. neovascularization [MSL92]. nephron [KYLV90, NSC^{+}93]. Nerve [CG90, DDW90, EB92, GMLD92, LB92b, MWG92, SDK^{+}93, Ang91, APP^{+}93, BML^{+}90, BG91a, BH94, CSKZ94, DR91, DJ93a, ET94, FFB^{+}93, KSYN^{+}94, KDT^{+}91, LMP91, MCM92a, MAWM90, MMW^{+}92, MMT91, PPHS^{+}92, SHR93, SST92b, WHT^{+}93, WB93b, YGEG91]. nerves [KDTB90]. nervous [AST^{+}92, CSH^{+}92, DMB93, GMB^{+}91, MBW^{+}94, MKCE92, MLR^{+}92, MMT91, NS93, SSWS92, SR92b, SLL94]. network [AES94, BLR^{+}94, BHD^{+}94, BBS92, CH91, CP90, EH94, GP91, GAS90b, HWM^{+}93, HHL94, HPYB93, JCSH93, KRHY93, KDP^{+}92, LKF^{+}94, MKSF93, MH93b, NHK^{+}94, OS92, PRL^{+}94, RB92, RSS^{+}94, RTC^{+}94, SNE90, SGBN^{+}93, SFS90, SBLV92, TGD93, VSB^{+}94, WB92a, XS93, YKY92, YCC92, ZDZSBT94, vDS94]. network-associated [WB92a]. networks [JETS91, MCF91, MVG91, PME93, SG92, SBC^{+}93, TH91, VLGB92]. Neu5Ac [OKH92]. Neural [AUT^{+}91, GSM93, MVTJ94, WBS91b, Blo92, DAS^{+}92, FCL^{+}94, FvBuHWS92, FMK^{+}94, FWBS93, GMB^{+}91, GFM93, GPRB93, HS92, HMS^{+}90, JM92, KKAS90a, KKAS90b, KT92, KA90, LBF92, MBW^{+}94, MFS^{+}94, OB93, PW93, PFSS92, RWG^{+}92, SCO^{+}91, VHW^{+}92, dlRKR^{+}90, dlR91]. neuraminidase [SC90b]. Neurite [KSC^{+}91, TGS92, ABM^{+}93, BSM94, BS90, BHS91, BZ90, BGB94, DDW90, DAS^{+}92, DSR^{+}93, DGGR92, FCL^{+}94, FMK^{+}94, HSM^{+}93, HFS92, JM92, JvCH^{+}94, KS90, LSR^{+}90a, LZBH92, LELB90, LVK^{+}91, PS94, PCG^{+}94, SWD92, SWD94, SLWF91, Smi94, SKD^{+}91, TKHM91, TWMS90, VFR94, WDT^{+}92, WWD94]. neurites [BJ90, BHC^{+}94]. neuritic [BJ92]. Neuroblast [DBTAB90]. neuroblastoma [ABM^{+}93, DG90, MJ92, WLZB93]. neurocan [FMK^{+}94]. neuroendocrine [BHC^{+}94]. neurofascin [VHW^{+}92]. neurofibroma [SRR90]. Neurofilament [OGM^{+}93, GWMC90, HWM^{+}93, KGSB92, MHGC90, NPST94, WC90b, OGM^{+}93, TOFH94]. Neurofilaments [LXWC93, LPK92]. Neurolin [PLS92]. neuromodulin [DJA^{+}90, DJA^{+}90]. neuromuscular [MVTJ94, NEL^{+}91, RB90, SEBH90, TTVV^{+}90, WFD^{+}91, WBS91b]. Neuron [KS90, BJ92, BGM^{+}91, KB91b, SOE^{+}91, WHT^{+}93]. neuron-glia [BGM^{+}91]. Neuron-specific [KS90, WHT^{+}93]. Neuronal [DPS^{+}93, ZDR^{+}92, BG93, CL91a, CL93, DJA^{+}90, DGGR92, EZF^{+}94, FMK^{+}94, GB90, HR90a, HWW^{+}93, JvCH^{+}94, KS92, MCM92a, MWG92, MHGC90, MWS^{+}93, OMH93, PDB^{+}91, QPWG91, SBN92, SLL94, VCFM94, VKRD94, YCAB93]. neuronal/chromaffin [MWS^{+}93]. Neurons [BS93a, AC93, ARWD93, BG91a, BTH^{+}90, BSM94, BMW^{+}93, CL94, DR91, DJ93a, DZdH^{+}93, DSR^{+}93, ET94, EZF^{+}94, FNV^{+}92, GMLD92, GFM93, HR90a, HdHD^{+}93, KA90, LHT^{+}91, MTP^{+}92, MWG92, MFS^{+}94, MMD^{+}93b, NIK^{+}93, NPST94, OH92, PSD92, RLS92, SWD94, SH90b, SHR93, Smi94, TOFH94, TK91, TWMS90, UKH^{+}91, VPC^{+}91, WAS^{+}91, WSL91, WMvdK94, WDT^{+}92, ZZP94]. neuropathy [KYA^{+}94]. neurosteroids [ASG^{+}93]. Neurothelin [SH90b]. neurotransmitter [TTVV^{+}90]. neurotrophic [LSR90b, MMW^{+}92, SST92b]. neurotrophin [BMW^{+}93, LCT^{+}93]. Neurotrophin-3 [LCT^{+}93]. neurotrophin-induced [BMW^{+}93]. Neurotrophins [APP^{+}93, FFB^{+}93]. neurulation [HS92]. neutralizes [EGP^{+}91]. Neutrophil [CPD^{+}93, KHHR92, PCD^{+}92, BLSR92, DS93, GLAB93, HRT^{+}91, HMB^{+}93, KHvdL^{+}92, MHM91, NXHMJ93, RUH^{+}94, SCS90, STo^{+}92, ZB94]. neutrophils [BLSR92, BFLS94, CWHH92, DCTG90, DES^{+}91, DCL^{+}92, FJN93, FJN94, GAHB94, HCYS90, JLC^{+}90, JTS^{+}91, JWHPM93, KR91, LNSSA93, LPM^{+}91, MBS90, MM90, MCDA^{+}93, MVM91, NS90, PZP^{+}91, PMM90, WRP^{+}91, ZLG^{+}90]. newborn [ASNH93, MHP90]. Newly [BSB92, AU90, BM91a, IDT^{+}92, LGH91, MH92, WCC^{+}90]. newt [GPF94, HBR90, MS92, RA90, SSS93]. nexus [LDS^{+}92]. NF [GWMC90, HWM^{+}93, KGSB92, MHGC90, WC90b]. Ng [BGM^{+}91, FMK^{+}94, KRdlR^{+}92]. Ng-CAM [FMK^{+}94]. Ng-CAM/ [FMK^{+}94]. NG2 [NDP^{+}91, SDH90]. NGF [ARWD93, BLT^{+}90, KG94, LSP^{+}91, MSC91, MMW^{+}92, MJM91, QPWG91, SCO^{+}91]. NGF-independent [ARWD93]. NGF-induced [QPWG91]. NGF-induction [MSC91]. NGF-receptor- [MJM91]. NH2 [CAS^{+}92, HD90, LGI^{+}93, PTS^{+}92, PTK^{+}93, SR92a, SRP90, SC90b, WW94a, YKM^{+}92]. NH2- [SRP90]. NH2-terminal [CAS^{+}92, HD90, LGI^{+}93, SR92a, WW94a]. NH2-terminally [SC90b]. nicotinic [BSM90, BM90, BM91a, CH92, DVT92, Fro91]. nicotinic-receptor [DVT92]. nidulans [MM93a, OMRM93, OOM90, LOOM93]. Niemann [KP90]. NIF [RUH^{+}94]. NIH [BS94c, DPND90, HHR^{+}90, LSAH90]. NIH-3T3 [LSAH90]. NIH3T3 [PCG^{+}94]. NILE [FMK^{+}94]. ninaC [PHWM92, PM93]. nineteen [NEL^{+}91]. nm [LM91, MH91, WGP^{+}93]. no [HKS^{+}92]. NO38 [PG92]. nocodazole [BMM90]. nod [TH92a]. node [BRWB91, BFL90, WIF^{+}90]. nodes [DI91, KDTB90, KB91b]. nodulin [MHV92b]. nodulin-26 [MHV92b]. NOF [TKHM91]. Nomenclature [BCF^{+}93, SFC^{+}94]. Non [KIV93, VB93, AW91, BRD94, GN92, HVNC92, INY^{+}93, LNSSA93, MC91c, NF92, PAC^{+}92, PRL^{+}94, WHK^{+}93, ZJA^{+}94]. non-adherent [LNSSA93]. non-anchored [MC91c]. non-autophagic [NF92]. Non-chemotactic [KIV93]. non-destructible [GN92]. non-differentiating [AW91]. non-epithelial [INY^{+}93]. non-genotoxic [BRD94]. non-muscle [WHK^{+}93]. non-myristylated [PAC^{+}92]. non-overlapping [PRL^{+}94]. non-platelet [HVNC92]. Non-sarcomeric [VB93]. nonadherent [GAHB94]. nonallelic [SSB^{+}91]. noncoated [HSvD91]. Noncoding [LG93b, ZZR94]. noncollagenous [YYY^{+}91]. noncovalent [dSBH93]. noncrystalline [KGDE94]. nondiffusible [KRB91]. nondividing [FA90]. none [HBP^{+}93]. nonendocytic [MP94]. nonerythroid [KWFM91, TQL^{+}90]. nonexchange [TH92a]. nonfilamentous [CFW93]. nonfunctional [Eri93, MC91d]. nonglial [SSC90]. nonglycosylated [KBG91]. nonhelical [HCKJ92]. nonhydrolyzable [MPEG93]. nonintegrin [BKB^{+}90]. nonjunctional [PET^{+}91]. Nonlysosomal [WL91]. nonmuscle [EBS91, FBML90, GYE^{+}90, HCKJ92, KA91a, KJT91, LSG92, RFF92, SLS^{+}90, SJM^{+}94, TNS^{+}94, YYM90]. nonmyelin [CSH^{+}92]. nonmyelin-forming [CSH^{+}92]. nonmyeloid [ZMS^{+}91]. nonneuroendocrine [LLG^{+}94]. nonneuronal [CL93, KKL^{+}91, MHGC90]. nonphototactic [HW93b]. nonreceptor [SF90b]. nonresponsive [EB92]. nonsarcomeric [GTC^{+}90]. Nonselective [KKG^{+}90]. nonsense [OGM^{+}93]. nonspecific [HCYK91, KHvdL^{+}92]. nonstriated [SLL^{+}90]. nontoxic [SRG^{+}94]. nontransformed [GSC^{+}92]. NOP2 [dBBHA94]. NOP3 [RT92]. nor [RBW93, VGMS91]. Normal [MHP90, CCT94, DUVV93, EK90, FPP91, FAF^{+}93, GWMP91, GB90, HSHA90, KP90, LSAH90, LGH91, LHFV93, ML91, PMTB94, PRAK94, RAGC94, RRBG91, WTRD92]. normally [NKW^{+}92]. notch [FJRAT91, KW93]. Novel [LPA^{+}90, PS94, SSW94, ZHL93, AORB90, AMBN91, AM91, BKSB94, BTH^{+}90, BOH^{+}92, BFM93, BZT94, BGK^{+}94, BTG^{+}93, CES^{+}92, CYC91, CSC92, CB90, CTQ91, DvdSM^{+}92, DG90, DW91, DAMS91, DMJ^{+}94, EHB^{+}90, EEC^{+}92, FNS^{+}91, FHI^{+}93, GMW^{+}94, GBP^{+}92, GHR91, GP91, GWMP91, GW94a, GF94, GFW92, HHO^{+}92, HAS92, HNPN94, HWF92, HYD93, HLC^{+}93, IRH^{+}91, JHGR93, JAS^{+}90, KR91, KKFD90, LRR^{+}92, LBH93, LM91, LWV92, LRAG92, LJP^{+}93, MLKB92, MB92, MK93, ML94, MY92, MG91a, MSF90, NVWR91, NDP^{+}91, PC91a, PMKH91, PRF^{+}93, PSZ^{+}93, PH94b, PET^{+}91, POGS90, RBW93, RK90a, SCE91, SCMU93, SON^{+}94, SXM94, SR92a, SSWS92, SWF^{+}92, SF91, SBR91, SGM^{+}90a, TZBV91, TLSW93, VVB^{+}92, VSVP93b, VLH^{+}93, WVH94, WLZB93, WCR92, WB94, WG92b, WKS^{+}93, WGBB91, YOWM91, ZDO^{+}91, ZAH^{+}94, ZSM^{+}93, ZMMP90]. novo [BMF^{+}91]. NPI46 [SXM94]. Nr [KRdlR^{+}92, JS90, GMB^{+}91, MKCE92]. NR-6 [JS90]. Nr-CAM [KRdlR^{+}92]. NSF [GE91, WRHW92]. NSR1 [LXM91, YM93]. nuc2 [MHYC90]. Nuclear [AMG90, ACFFL93, FZ90, FHvZ^{+}94, GG92, HODF94, JBIM94, KBL^{+}94, LCC^{+}93, MY92, AL92, AA94, AR93, AM91, BWLK93, BEGN91, BAP^{+}90, BTS93, BKWD94, BNI^{+}94, BHA^{+}94, BDW92, BMB^{+}92, BSP^{+}94, CFFL93, CLCC90, CTL91, CC93a, CYCA90, CRK^{+}93, CDRL93, DIDB^{+}91, DMJ^{+}94, FA90, FA91, FMB^{+}91, FKS93, FDT^{+}91, GN92, GSBS92, GCML94, GSC^{+}92, GA92, GMML^{+}92, HWB93, HSS93, HNP90, HPWN93, HAS90, HDES94, Hur90, IMK^{+}92, ISDM92, JO92, JAB^{+}91, JS90, KM92a, KKBL92, KN91, KDN94, KMN92, KT93, LXM91, LL91, LM94, LSP^{+}93, LSK^{+}94, LE92, MHG93, MF90, MW93, MT94, MB90, MPEG93, MG91a, MG92, MMLG94, MHGJ94, NF90, NWD90, ND92, NKW^{+}92, OSJS94, OOM90, PA93, PBM^{+}94, PLB^{+}91, PH91, PMTB94, PB90, PE93, PJS^{+}94b, RBW93, RHB^{+}90, RDDR93, RZS91, RB93, SHDS93]. nuclear [SDW^{+}93, SB93c, SK94b, SW93b, SES94, SR92d, SDPH90, SMBG92, SOKS91, SS92b, SLK93, SHOA92, TSBS92, UHFG92, VHSR92, VL91, WLZB93, WRB92, WB93a, WB94, WS93b, WKS^{+}93, WDP^{+}94, WHC^{+}93, WEB90, WB92c, WBR94, WMCG91, Wu93, XL91, ZKMB93, vZDF^{+}93, vDS94]. nucleating [BA92, BDK92, CSSL93, MSC92, SL90a]. nucleation [BPH^{+}90, MPVB91, SW90, TCP90, TZBV91, YCAB93, TDWT92]. nuclei [DK92, DIDB^{+}91, FCFL94, KLC^{+}92, LM94, MG92, OHS92, WE90]. Nucleocytoplasmic [JS90, SLK91, WB93a]. nucleolar [CL92, CYCA90, JHK^{+}91, MB94, PG92, TW93, dBBHA94]. nucleolar-specific [CYCA90]. Nucleolus [TW93, BZT94, MB90, SXM94, YM93]. Nucleolus-like [TW93]. Nucleoplasmic [MW93, CL92, GA92]. nucleoporin [HWB93, WB94]. nucleoporin-like [HWB93]. nucleoskeleton [MCS92]. nucleosomal [HASW94, LSAH90]. nucleosome [Woo94]. Nucleotide [SSVE91, SFO^{+}91, CRS94, PNPB94, RK90a, WNM^{+}94]. nucleotide-sensitive [RK90a]. nucleotides [DLSK91, HBP^{+}93]. Nucleus [FRSL90, Rot90, BMF^{+}91, BHF90, CTL91, GSR^{+}93, HTSP94, HTD90, JS92, RH92a, RR92b, SCE91, SOWDK94, SB93b, SES94, WSvdK^{+}93, YLS92]. Nucleus-associated [FRSL90]. nucleus-limited [HTSP94]. Nucleus-specific [Rot90]. nude [QvMWV^{+}91]. NUF1 [MCS92]. Nuf2 [OSJS94]. null [AGKC92, OCC94, WB93a, WS90]. NuMA [CSC92, CC93b, YLS92]. number [CLZ91, HS90a, MIU^{+}92, TSR90]. numbers [CMV^{+}92]. numerous [SDW^{+}93]. nup1 [BHA^{+}94]. NUP116 [WB93a]. NUP145 [WB94]. Nup155 [RBW93]. Nup180 [WKS^{+}93]. Nup1p [BKWD94]. Nup2p [BKWD94].

O [SPKV94, SMBG92, DMM^{+}91, HDC^{+}91, IDT^{+}92, PIS94, WN92, WWN92]. O-2Aadult [WN92, WWN92]. O-2Aperinatal [WN92]. O-glycosylation [IDT^{+}92]. O-linked [SMBG92, HDC^{+}91]. O-permeabilized [PIS94]. obligate [LXWC93]. Observation [SW93a, Bea91]. observed [WPS91, Woo94, dBNI^{+}91]. obtained [MOKF92, PBCP91]. obvious [RFF92]. Occludin [FHI^{+}93, FIH^{+}94]. occur [DEH^{+}91, DS90c]. occurring [FHGN91]. occurs [DUVV93, DSQ^{+}93, DM92, FGS91, HW94, MPM93, MG91a, MS92, RLWK93, SSJ^{+}94, SFS90, WF93]. octagonal [KHCR91]. octamer [SGW^{+}91b]. octapeptide [IKFR91]. oda6 [MK91b]. often [MSI^{+}94]. OHS [FP91]. OHS-4 [FP91]. oil [TH92c]. Okadaic [BH94, GPS92, PLB^{+}91, LFC94]. old [KWS94, KAG94]. Olfactory [KA90, VSVP93a, CL91a, NIK^{+}93]. oligodendrocyte [AFHDD90, GFL92, MPIDD93, MGS90, MAB^{+}90]. oligodendrocyte-myelin [MGS90, MAB^{+}90]. oligodendrocytes [AAM^{+}93, ASDC91, DJA^{+}90, DB90]. oligodendroglia [MHR94]. oligodeoxynucleotides [SLS^{+}90]. oligomannosidic [HSM^{+}93]. Oligomeric [HKSL90, LTS93, MG94b, PL90]. Oligomerization [WSM93, PMA91]. oligomers [HMLL94, SRHK94, SJM^{+}94]. oligonucleotides [FNV^{+}92, TSK^{+}94]. oligosaccharide [FWD^{+}92]. omega [HTSP94]. oncogene [FKB^{+}90a, LSAH90, LVBK92, RLRR93, SBG93, VWT^{+}91]. oncogene-transformed [LSAH90]. oncogenes [HR90a]. oncogenesis [SHB90]. oncogenic [BSP^{+}94, TOA^{+}91]. oncoprotein [BMB^{+}92, HXM^{+}94, MSSD93, XMW^{+}93]. oncoprotein-mediated [BMB^{+}92]. oncoproteins [LMB^{+}94b, LE92]. one [BMH^{+}93, KLERG94, LWK^{+}93, VBV^{+}90]. only [GAvdM91, HLWL92, IDT^{+}92, KLERG94]. onset [HDP^{+}93, PPHS^{+}92, RSCS94, WKMG92]. onto [TOK93]. Ontogenesis [DB91]. ontogeny [TGCT90]. Oocyte [JGL^{+}93, BBB^{+}94, BHF90, DYV93, DCB^{+}92, FMLD92, JG92, KA93, LMSH94, MF90, MG92, MMCH91, PSN^{+}91, TSBS92]. oocytes [CC90a, CSMG90, CPM^{+}91, DK92, DMM^{+}91, DIDB^{+}91, FMD93, FM94a, FHvZ^{+}94, FKB^{+}90a, MKF91, OW90, PET^{+}91, PG92, PLB^{+}91, SCL91, SJIC92, SLH92, SBBS90, SPvV94]. oogenesis [CKMC94, KMF^{+}91]. ooplasmic [SSJ90, STH^{+}93]. open [KHCR91, ZZR94]. operates [WHLW90]. Opposite [ZPS^{+}92, CSvdE93, VMB92]. opsonic [MSD92b]. optic [LPK92]. optical [BZT93, GXHS94, MOKF92]. order [BRF93, HODF94, HM91]. organ [CKD90]. organelle [BRL^{+}93, FFKC94, HN90, Hol93, KDLS92, LVBK92, SSVE91, SON^{+}94, TH90, TSR93]. organelles [BHLA90, DP94, HSYYK90, HSYK^{+}91, LFR93, Ral93, SM90, SRB92]. Organization [TJ91, ACFFL93, BPK91, BGGW^{+}94, CTL91, CK94, Cit93, CCAF90, DCB^{+}92, ES91, FMR^{+}93, FW93, FBLL90, FAF^{+}93, GCML94, GLQRB91, GE91, HM93, JK93, JGO^{+}91, KCCG91, LSAH90, LB92b, LGI^{+}93, MH94, MW93, MSLS91, NPR^{+}94, OHS92, OK92, PMHS92, SBK94, SLHG93, SBWV90, SDCD93, TFT94, UNY90, VB93, WS91b, WJ92]. Orientation [FW93, RK94, BA93, CM94b, MT94, PSHK92, TT94]. orientations [KII^{+}91]. oriented [FB93]. origin [BGM^{+}91, RSB^{+}90, RR92b, SGVS^{+}93, WWN92]. origins [BJ92, HS90b, SKM94]. orthovanadate [WB93b]. oscillates [GC91]. Osmotic [CSG92, LBH92]. osseous [PLB^{+}92]. osteoblast [GBN^{+}93, KYK^{+}94, TN94]. osteoblast-like [GBN^{+}93]. osteoblast-specific [TN94]. osteoblastic [ASNH93, OSKN94, YKI^{+}91]. Osteoblasts [GSWW93]. Osteocalcin [CCG^{+}94]. Osteoclast [MHG^{+}90, CCG^{+}94, GBP^{+}92, RSM^{+}94]. osteoclast-like [CCG^{+}94, GBP^{+}92]. Osteoclasts [HNC^{+}92, LHH^{+}91, SAVK94, VKS^{+}90]. osteogenesis [JRSB90, WIS^{+}94]. osteogenic [ASNH93, CMS^{+}90b, KBCM^{+}90, SCSE90, VBvG^{+}94]. osteoprogenitor [ASNH93, TA91]. osteosarcoma [FP91, KVE92]. other [ISN^{+}94, JETS91, NEL^{+}91, Ral93, RA94, TH91]. outer [BDJ94, MLN93, MMLS92, MAFR93, MSFW92, MGW92, SMK91, STK^{+}93, SY94b, TK94]. outgrowth [ABM^{+}93, BSM94, BS90, BHS91, BZ90, CK91, DDW90, DAS^{+}92, DSR^{+}93, DGGR92, FCL^{+}94, FMK^{+}94, HSM^{+}93, HFS92, KSC^{+}91, LVK^{+}91, PCG^{+}94, SWD92, SWD94, SLWF91, Smi94, TKHM91, TWMS90, TGS92, VFR94, WDT^{+}92, WWD94]. oval [NBT94]. ovarian [HSHA90]. ovary [BGGW^{+}94, CSC90, KD92, RHR90, TYF90]. over-expressed [PW93]. overcomes [DMWJ90]. overexpresser [DPND90]. Overexpressing [YHM^{+}94]. Overexpression [BG91b, IBD92, KKL^{+}91, WLWL94, AGKC92, CBM^{+}94, DSJB^{+}94]. Overlapping [NPH^{+}93, PE93, BTG^{+}93, PRL^{+}94]. Ovotransferrin [GDB^{+}94, GBN^{+}93]. own [TDT92, TDWT92]. oxidase [MC91a, MC93, SAVK94, WGR^{+}92, ZVF91]. oxidation [SYCA94]. Oxidized [MCF93]. Oxygen [PZP^{+}91, ET91, THH^{+}90]. oxygen-evolving [ET91]. oxysterol [RDH^{+}92].

P [KCB^{+}92, ARW^{+}94, CKGC91, EWP^{+}93a, GSMK94, KSL^{+}94, LWK^{+}93, LLIV94, MSD^{+}92a, RPS^{+}92, STo^{+}92]. P- [EWP^{+}93a]. P-domain [LWK^{+}93]. P-glycoprotein [CKGC91]. P-selectin [KCB^{+}92, GSMK94, KSL^{+}94, LLIV94, MSD^{+}92a, STo^{+}92]. P0 [FT93, MJM91, WF94]. P1 [SLM^{+}90]. p105 [MSSD93]. p105-Rb [MSSD93]. P19 [SBN92]. p190 [EEC^{+}92, ECM^{+}92]. p190-calmodulin [EEC^{+}92]. p21 [BMW^{+}93, KSK^{+}93]. p21rho [PSG^{+}90, APH92]. p23 [TWQ^{+}93]. p24 [WMG^{+}93]. P29 [BTH^{+}90]. P3 [HFN92]. P30 [DR91]. p34cdc2 [AM94, DIDB^{+}91, EPNN91, HTP^{+}92, JAB^{+}91, RSW^{+}91, SLW^{+}92]. p34cdc2-mediated [JAB^{+}91]. p42mapk [LSP^{+}93]. p44mapk [LSP^{+}93]. P450c21B [BTG^{+}93]. P450scc [HSHA90]. p45mapkk [LSP^{+}93]. p55 [PB90]. p56lck [PMBL^{+}92]. P58 [PDSB92]. p60c [KSVM92]. p60c-src [KSVM92]. p62 [GHR^{+}90, JCSH93, SDPH90]. p63 [SRHK94]. p67SRF [VGRC^{+}92]. packing [KPMG91]. pad [SHR93]. PAF [LPM^{+}91]. PAI [EGP^{+}91]. PAI-1 [EGP^{+}91]. pair [FT93]. paired [WDB^{+}92]. pairing [GHK94, HDP^{+}93, LKS94, SBK94]. palmitoylation [CAS^{+}92, MW92, SSDK^{+}94, SVB94]. palmitylation [POGS90]. pancreas [NT90, ODK^{+}93, THL^{+}90]. pancreatic [BRF93, CAS^{+}92, DYS^{+}90, GCZ^{+}92, JGK^{+}94, KA92, KA94, TGD93]. panel [GHBA^{+}93, NEL^{+}91]. papilla [LVA^{+}90]. papilloma [XMW^{+}93]. paracellular [PCD^{+}92]. paracrine [LK93b, OCMB91, YGEG91]. paracrystalline [EERM^{+}91]. paracrystals [LM91]. paradigm [RNG94]. Parafusin [SSRM90]. paragloboside [KYA^{+}94]. Parallel [CJB^{+}93, GHBA^{+}93, SSK^{+}90]. paralysis [RFF92]. paralyzed [CSKZ94, PPD92]. Paramecium [EP94, GdLMS94, KC90, KBP91, SSRM90, SFA90, SMCA91]. paramyosin [BOH^{+}92, DE93, ECO93]. Parascaris [GGGE^{+}92]. parasite [BDMPJ94]. parasitic [MC94]. parasitophorous [BDMPJ94]. parasympathetic [ARWD93]. Parathyroid [AWH^{+}94, vdSKL^{+}93]. parenchymal [SGM^{+}90a]. parietal [vdSKL^{+}93]. parotid [PMHS92]. part [GPH^{+}90, RKA93, VGG^{+}93]. Parthenogenesis [KDP^{+}90, TKP^{+}91]. partial [BWCG93, LMM^{+}91]. partially [BEO90]. participate [CRWS93]. participates [CTPP92, CNU92, SVK^{+}92]. Participation [DW91, SPH^{+}91]. particle [BN91, BMGN92, CG93, CLM91, HD91b, KKBL92, LMW^{+}94, LPA^{+}93, MMF94, MNB92, RG92, RWL^{+}90, TMI^{+}91, WSC94, WS90, WC90a, WWW^{+}92, YMO92, ZBW93]. particle-bound [WC90a]. particles [BBB^{+}94, DWY^{+}93, KHCR91, MKF91, SPvV94]. particular [WF90]. particulate [JJS92]. partitioning [LSP^{+}94]. partner [MGG93, PRF^{+}93]. parva [STM91]. PAS1 [HMS94]. PAS3 [HVK91]. PAS5 [SS93]. pas8 [MMS93, MMS93]. pasin [KKFD90]. passage [BS94a]. passive [KA92]. Pasteurella [SBR91]. pastoris [HMS94, MMS93, SS93, WGR^{+}92]. Patch [dTF90, NHH^{+}91, RSS91, SS90a]. Patch-clamp [dTF90]. patch-clamped [RSS91, SS90a]. pathfinding [WFD92]. pathogen [TCP90]. pathogens [ELZ93]. paths [WMB^{+}94]. Pathway [GMML^{+}92, AB90, BSHW91, BGH90, BS94b, BSHB90, CdPA^{+}93, CM94a, CSC92, DUVV93, EJ93, FM94b, HH94a, HG94, HFS^{+}93, KYK^{+}94, KCY92, KCB^{+}92, KBH91, KA92, KBL^{+}94, LGH91, MPK93, MMF92, MTC^{+}91, MG91a, MC91a, NWD90, NF92, PJS^{+}92, RNS92, RPPB94, RBD92, RYK^{+}90, SB93a, SCV^{+}91, SNG^{+}91, SFS90, TGD93, THL^{+}90, TLSW93, WF91, WV93]. Pathways [GHGP^{+}92, DPS^{+}93, DCM^{+}93, FDT^{+}91, GLQRB91, GHT92, GK92b, HK94, HPW94, HNPN94, HSUS90, HH90, KFN93, KP90, LQNRB90, MWLC^{+}94, MG91a, MG92, MSLS91, PCD^{+}92, PSD92, PRC^{+}90, QH94, SWD92, SBB^{+}91a, SM91a, SNG^{+}91, SKSD^{+}94, SK94b, SKBD91, SKK^{+}91, TRH^{+}92, VVR^{+}90, YL93]. patients [BBB^{+}94, KS94, PBCP91]. pattern [BLB^{+}91, PZV^{+}91, SBH92]. patterning [AA90, CHG94]. patterns [CSR^{+}92, GZC^{+}91, LWK^{+}93, LYC^{+}94, MSF90, ME91, PSJ^{+}91]. Pax [KG94]. Paxillin [TGB90, Tur91, BTR92, FJN94, GAHB94]. PC [COLL^{+}90]. PC-12 [COLL^{+}90]. PC12 [BG91a, BS90, BHS91, BHC^{+}94, BCB93, CM94a, CC90b, DDW90, GFWM92, GK92a, GZC^{+}91, HFS^{+}93, KS92, KDT^{+}91, LSR^{+}90a, LVBK92, LSN^{+}94, MSC91, MWG92, MWS^{+}93, PWY^{+}92, PBCK93, PCG^{+}94, QPWG91, SGW91a, SDK^{+}93, TGS92, VFR94, WB93b]. PCM [BBZ94]. PCM-1 [BBZ94]. PDGF [BRIA^{+}94, JRCW^{+}90, KSB^{+}93, LD91, LG93a, MMG90, MRBP^{+}90, MIS93, OTWH92, PDS90, RR92a, SDR^{+}90, SWHCW91, TS93, WMS^{+}93, WN92, YRBBP90, YGM^{+}94]. PDGF-AA [OTWH92]. PDGF-BB [BRIA^{+}94]. PDGF-induced [MIS93, YGM^{+}94]. PDGF-stimulated [LG93a]. PECAM [AMBN91, ASL^{+}94, DCY^{+}94, NHA^{+}92]. PECAM-1 [AMBN91, ASL^{+}94, DCY^{+}94, NHA^{+}92]. pEL98 [TNS^{+}94]. pemphigoid [CKG^{+}90]. Pemphigus [KAP^{+}93]. penetrate [LML^{+}94]. penetration [WG92a]. peptidase [HD90, IKFR91, JZAVP93]. peptide [AWH^{+}94, BGLJ^{+}93, CMS^{+}90a, DKM^{+}92, EUH91, GSFS92, GFWM92, HW92, JWHPM93, KIS^{+}94, LASB^{+}93, MME93, MR92, RWG^{+}92, SBB^{+}91a, SMHO92, SOH^{+}90, SGN^{+}93, TRV^{+}90, TRD^{+}92, WW94a, vdSKL^{+}93]. peptide-depleted [AWH^{+}94]. Peptides [SMM^{+}91, TVG^{+}93, BI90, FS94, LS90, LIAJS94, LLK91, SLWF91, TRD^{+}92]. peptidylglycine [MJM92]. PER1 [WTH^{+}94]. perfused [dAWS^{+}90]. peribacteroid [MHV92b]. pericellular [Knu93, LJJC93, LZI^{+}93, SWL90]. pericentriolar [HGS^{+}94]. pericytes [ND91, STL^{+}90]. periods [DA92]. peripheral [APP^{+}93, BZT93, BRWB91, BFL90, CSH^{+}92, DR91, DLSB^{+}90, DLSK91, FB93, HSG94, KDTB90, KKFD90, MMW^{+}92, MMT91, NS93, PRAK94, RTC^{+}94, SHR93, SSWS92, WCR92]. Peripherin [DZdH^{+}93, AMMW92, TGS92]. peripherin-depleted [TGS92]. peripherin/rds [AMMW92]. periphery [Hur90, WHC^{+}93]. peritoneal [HCYK91, MTJM93, RHGG92, TLXM90, TMI^{+}91, YR93]. perlecan [HMY92]. permeability [FA90, HR90c, LASB^{+}93, SOPA94]. permeabilization [CDRL93, LM94, WFCN94]. permeabilized [AMG90, BPZ^{+}92, GAS90b, KTG90, PDSB92, RPH^{+}92, SB91, SGN91, SM91d, WS93a, XS93, PIS94]. permeable [NSC^{+}93]. peroxidase [CFG^{+}94]. peroxide [NXHMJ93, OSKN94]. peroxisomal [DJ93b, GKKS90, HVK91, KLR^{+}93, LHFV93, MMWG94, MMS93, WGR^{+}92, WWS^{+}94, WTH^{+}94, WS93a]. peroxisome [BRD94, FFR^{+}91, HMS94, HVK91, MHDT94, PTD90, SS93, VVB^{+}92, WTH^{+}94, ZHL93]. peroxisome-to-mitochondrion [PTD90]. peroxisomes [DJ93b, KKG^{+}91, LHFV93, MMWG94, MG94b, TYF90, ZVF91]. persistently [TSP90]. perspective [JB93]. Perturbation [RB92, TSK^{+}94]. Perturbations [RFF92, DSJB^{+}94]. perturbed [MHT^{+}92]. pertussis [GHR91]. Peyer [NHH^{+}91]. PF [PKG^{+}94]. PH [LCMP90, LML^{+}94, CM94a, CVH91, DVT92, FM94a, FWF90, GC90, HR90c, IPF^{+}90, OHP^{+}94, PDB^{+}91, WG92a]. PH-20 [LCMP90, LML^{+}94]. pH-independent [OHP^{+}94]. pH-induced [WG92a, PDB^{+}91]. phage [KWR94]. phagocytes [ELZ93]. phagocytic [BGLB94, HG92]. Phagocytosis [DH93, CBC^{+}94, GeKdV^{+}91, HCYK91, JLC^{+}90]. phagolysosomal [RHGG92]. phagolysosomes [DHPG94, HG92]. phagosome [JLC^{+}90]. phagosome-lysosome [JLC^{+}90]. phakinin [MGG93]. phase [CVK^{+}94, DS90c, GS94, LCC^{+}93, Mac92, PLB^{+}91, RRCR93, RTHB90, SPD^{+}90, JS92]. phenomenon [VDTT91]. phenotype [BCG^{+}92, BF93, BEO90, CSG92, HAA93, KGSB92, KZS94, MPAF91, MWLC^{+}94, MJM91, SEC^{+}93]. phenotype-specific [CSG92]. phenotypes [CBM^{+}94, RFF92]. phenotypic [CCT94]. phenylalanine [PTK^{+}93, WB94]. phenylalanine-based [PTK^{+}93]. phenylalanines [MPAF91]. pheochromocytoma [COLL^{+}90, EB92, PCG^{+}94]. pheromone [DHS93, GS90, OMBL90]. pheromone-induced [GS90]. pheromones [TDSP93]. Phodopus [KCDR91]. Phorbol [CSS^{+}94, DCL^{+}92, GBP^{+}92, JGLK90, LW92, ZSC91, Lam94]. phosphacan [MFS^{+}94]. phosphatase [ACSM^{+}94, BKFT93, EG91, FBML90, FBL92, Low92, MFS^{+}94, WTRD92]. phosphate [GHR91, GK93a, GK92a, GI93, HWvF90, KHV^{+}93, MH93b, SGBN^{+}93, ZDO^{+}91]. phosphate/insulin [JK92]. phosphatidic [HE93]. phosphatidylcholine [HE93, HRT^{+}91, KFN93, LNS^{+}94, MSW^{+}94]. phosphatidylinositol [BGRW91, DLD^{+}90, DVT92, DKM^{+}92, LNSSA93, LNS^{+}94, Low92, MMC91, MSW^{+}94, MDKB94, SSTL93, SWFB91, VR92, VM93, VGM^{+}93, YOWM91]. phosphatidylinositol-anchored [BGRW91, DLD^{+}90, DKM^{+}92, YOWM91]. phosphatidylinositol-glycan-anchored [Low92]. phosphatidylinositol/phosphatidylcholine [LNS^{+}94]. phosphatidylserine [DVT92, KA91b]. phosphoepitope [GR93]. Phosphoinositide [PvBeHB^{+}91]. Phospholipase [BGRW91, CXS^{+}94, HRT^{+}91, MWLC^{+}94, MDKB94, PvBeHB^{+}91]. phospholipid [BMTTG94, DP92, ES91, HWM90, QYC^{+}92, WCM^{+}93, ZDP92]. phospholipid-anchor [BMTTG94]. phospholipid-anchored [ES91]. phospholipids [BVR^{+}93]. Phosphoprotein [WCW91, BS94b, NIK^{+}93, SSRM90, SB93b, SS92b, dAWL^{+}90]. phosphoproteins [RZS91]. phosphorylate [KA93]. phosphorylated [BTH^{+}90, FA92, FFS^{+}93, JLL^{+}94, MAWM90, RMG90, SY92, VBAK91, WLPB92]. phosphorylates [DIDB^{+}91, DFC^{+}92, OBC^{+}92]. Phosphorylation [BS94c, BM91b, FMLD92, HSS93, HPWN93, MH93b, NPST94, SLW^{+}92, WC93, ACSM^{+}94, AGB^{+}91, BP93, BVF^{+}93, BFLS94, BTR92, CS91, DMB92a, DP94, DDW94, FBML90, FJN93, FJN94, FH90, GBS90, GAHB94, GC90, Ham94, HLC^{+}93, JAB^{+}91, JGLK90, KFC^{+}94, KMN91, KO94, LCKW91, LSK^{+}94, LFC94, LYK^{+}94, LHS^{+}92, LE92, MHT^{+}92, MSPB90, MBS91, MHV92b, MYC92, MCEG90, MG91b, NHA^{+}92, ORS90a, OCC94, PBD93, RTHB90, RELS^{+}91, SMM90, TS93, TOA^{+}91, Wal94, WS93b, WGCdL91, WMCW92, WG93, YYM94, YKY92, ZB94]. Phosphorylation-site [WC93]. phosphotyrosine [Tur91]. phosphotyrosine-containing [Tur91]. photoactivated [OH92]. photobleached [OH93]. photobleaching [HKSL90, OH93]. photoinhibition [BFR^{+}92]. photolysis [FKT93]. photooxidation [DMLR^{+}94]. photoreceptor [DP91, GFH90, PBCP91, PHWM92]. photoreceptors [AHW90, AMMW92]. photosynthetic [ET91, TM90]. photosystem [BFR^{+}92]. Physarum [SKBD91, UNY90]. physical [BKWD94]. physically [BHHG90]. physiologic [KO94]. Physiological [HSSC92, ARW^{+}94]. Physiologically [ET91]. physophilin [TB90]. PI [YGM^{+}94]. PI-3 [YGM^{+}94]. Pichia [HMS94, MMS93, SS93, WGR^{+}92]. Pick [KP90]. picosecond [FV91]. pig [MFOI92]. pigment [GORB91, GPRB93, KM94, SR92e]. pilus [RLB^{+}91]. pinocytosis [OA92]. pioneer [OB93, SOE^{+}91, WFD92]. pit [HPW94, LMSA91, MMH^{+}92, RYK^{+}90, SS91a, WRA93]. pits [AG93, FZRH91, FGS91, PDA93, PMBL^{+}92, PMS^{+}91, SHH94]. pituitary [TLWA94]. PK1 [JPJF91, SdAN^{+}91]. PKC [OBC^{+}92]. PKC1 [LBH92, RPPB94]. place [HWvF90]. placenta [BMWT93]. placental [Low92]. Plakoglobin [KW92a, BCB93, PMG^{+}92, TTE^{+}94]. plakoglobin-binding [TTE^{+}94]. planar [TSS^{+}92]. plane [FRSL90]. plant [FRSL90, GXHS94, TH92c]. plants [KRCT93, MSLS91]. plaque [PKN91, TTE^{+}94, TIN^{+}93]. plaques [DI91, Kos94, MG91b, TYM^{+}92, WRDK92]. Plasma [MAAO92, ATBD90, AJPB90, BN91, BMGN92, BPS90, BHHG90, CSvdE93, CS91, CKL^{+}91, CF94, DBM94, FABS90, Fuj93, FHI90, GLQRB91, GAS90b, HM92a, HHL94, vtHvM90, HNH91, HGS^{+}94, HdHD^{+}93, HPP^{+}93b, HKS90, KFN93, KP90, KLBBK91, LQNRB90, LMSA91, LML^{+}94, MC91b, MOT90, MPM^{+}94, MB93, NCM^{+}93, PKN91, PET^{+}91, PSN^{+}91, PPFM90, PNB91, PMM91, PMM90, RK94, RMB90, RLWK93, SK94a, SST92a, SBR93, SL90a, SR90b, TB90, WNBAS90, WST^{+}90, WS92]. plasmalemma [SP94]. plasmalemmal [FNM^{+}92, LMP91]. plasmid [SSKT^{+}92]. plasmin [BSEF^{+}93, LGPM90]. Plasminogen [CML90, EGP^{+}91, EMC^{+}90, FMPP90, FAMR92, QvMWV^{+}91, SR90a, SPD^{+}90, SLP91, SWL90]. plasmodesmata [TCCT91]. Plasmodium [EH94, UNY90, CKGC91, GRC^{+}91, KCM^{+}93, OKH92, SOH^{+}90]. plasmolysis [TCCT91]. plastin [dAWL^{+}90, AFA^{+}94]. plate [MAFR93]. Platelet [DS90a, TRLG90, DCY^{+}94, FABS90, FTT^{+}93, GHS^{+}93, GBS90, GCMDP91, GS94, HD91a, Har92, HVNC92, KHHR92, NHA^{+}92, ZMMP90]. platelet-activating [KHHR92, ZMMP90]. Platelet-derived [DS90a]. platelet-endothelial [DCY^{+}94]. platelets [BB93, FM94a, KCB^{+}92, LHS^{+}92, LSC^{+}93, MPVL92, MOT90, SGN91]. plates [ZVF91]. play [GIAS93, HM93, HW94, RRD93]. plays [SS94, VKM93]. pleckstrin [IKL^{+}94]. plectin [FFSW91, WBL^{+}91, WGD^{+}93]. pleiotropic [BHA^{+}94]. plug [CCR93]. plus [BA92, SRB92, VGC94]. plus-end [SRB92]. PMA [TSH^{+}93]. PMA-induced [TSH^{+}93]. PMC1 [CF94]. PMP [SSWS92]. PMP-22 [SSWS92]. PMP-22/SR13 [SSWS92]. PMP47 [MMWG94]. pocket [SIR^{+}94]. podosome [MHG^{+}90]. point [DMWJ90, TYM^{+}92, TWC93]. pointed [FSM^{+}93, TDWT92, WPBF94]. poised [AL92]. polar [CWHH92, MCG92, RS94]. Polarity [JR92, STH^{+}93, AJL^{+}90, AL91, BA93, BFM93, BBGK90, GPF94, GB90, GORB91, HW93a, vtHvM90, INM^{+}93, JP90, KDT93, LB92b, PMG^{+}92, SZKM91, SGP91, SBB91b, TDWT92, VMB92, VCFM94, WKN92, WA93, ZPS^{+}92]. polarization [AG93, BPH^{+}90, FV91, GPRB93, LJP^{+}93, SSA^{+}90]. Polarized [BO91, DP91, MBC^{+}91, SVR^{+}90, AGKC92, BS94a, BMM90, BR91, CHBGL91, CvdE91, CPD^{+}93, CPP^{+}94, FTB94, GIAS93, HH90, LDNM92, LB94, MNME^{+}93, MYM94, MZRM93, MAAO92, NHS^{+}94, PYH^{+}90, PSN^{+}91, PHSvD92, RK94, RLMG92, SPRvD91, SZP^{+}93, WNTT94, WBT^{+}94, ZJA^{+}94]. polarizes [EMC^{+}90]. pole [BR94b, MSC92, MM93a, OSJS94, RK90b, SW92, SCF^{+}93, SHOA92, WGBB91, WHC^{+}93]. Poleward [HS90a, MS92, SM91b, RA90]. Poly [TLFS92, KDKF93, CTL91, HDES94, PSN^{+}91]. polyadenylated [HTSP94, WDP^{+}94]. Polyamines [HAA93]. polycephalum [SKBD91, UNY90]. polymer [RMK91, SIR^{+}94, VGC94]. polymerase [ISDM92, RZS91, WSvdK^{+}93]. polymeric [AKK^{+}93, BHC^{+}94, SVR^{+}90]. polymerization [CCR93, CVK^{+}94, FKLS92, HE93, HCYS90, LNSSA93, MBHG91, MVV^{+}91, SW90, SM91d]. polymerization-competent [MBHG91]. polymerized [HLPL90]. polymorpha [WTH^{+}94]. polymorphonuclear [CLZ91, CCFZ92, CMZ90, CSNZ92, CWHH92, SY90b]. polypeptide [FS94, MMC91, MBB^{+}93, NHK^{+}94, PMPV92, WBL^{+}91]. polypeptides [BWV92, BGM^{+}91, HKSL90, KA91a, KMSW91, YSK90, ZLG^{+}90]. polyphosphoinositide [DPCP91]. polyploidization [ISN^{+}94]. polyps [HHH^{+}93]. Polyribosome [HDDS94]. polysaccharides [LS92, MSLS91]. polysialic [ASR91, FRL93, YYR92]. polyspermy [MGDS93]. polytopic [FKG94, WEB90]. POM152 [WBR94]. pombe [BHBG94, DMM93, HRH90, MSC92]. Ponticulin [HLL94, HHL94, CSSL93, SL90a]. pools [ET91]. populate [HZC93]. population [VL91]. populations [CMZ90, CW90c, DE93, HSvD91, HPR93, HPW94, LB92b, LK93a, LHFV93, RHG90, SMRG91]. Porcine [HM94]. pore [AR93, BKWD94, BHA^{+}94, BSP^{+}94, CRK^{+}93, DLMS91, FMB^{+}91, GG92, HWB93, MG92, MF92, NF90, PA93, PBM^{+}94, PYA90, RBW93, RHB^{+}90, RB93, SDPH90, SMBG92, TIA93, VL91, WRB92, WB93a, WKS^{+}93, WB92c, WBR94]. pore-forming [PYA90]. pores [GA92, ZBS^{+}94]. Porters [LH93]. portion [PR91]. position [DAMS91, RS94, WF90]. position-dependent [DAMS91]. positional [DFLK94]. positioned [Car91, NWvH94]. positioning [ES92, FHUY93]. Positive [TA91, CDRL93, VVB^{+}92, WGM^{+}93, PTMB94]. possess [FB93]. possesses [GMQ^{+}93]. possessing [SSA^{+}90]. Possible [BSR^{+}90, BML^{+}90, BCBK92, BVR^{+}93, BML^{+}94, CJO92, FIH^{+}94, KFC^{+}94, PSG94, PNR^{+}93, PAL^{+}90, SJGG94, TKHM91, ZS94]. Post [MCM92b, YR93, BKR90, DP91, MJG^{+}91, NHK^{+}94, SZB94a, HWF92, HEF92, MC92]. Post-Golgi [MCM92b, BKR90, DP91, MJG^{+}91]. post-mitotic [SZB94a]. post-trans-Golgi [NHK^{+}94]. Post-transcriptional [YR93]. posterior [NCM^{+}93, PNB91]. posterior-tail [NCM^{+}93]. postimplantation [WWHA91]. postmitotic [CM93]. postsynaptic [CSP90, PMM91]. Posttranscriptional [PBCK93]. Posttranslational [WF91, dSBH93, CKP91, GFWM92, POGS90, dSK90, dCQTR91]. potassium [AG93, ABM^{+}93, HSvD93a]. potent [ASE90, BZ90, BDZ^{+}92, SSG^{+}90, SKD^{+}91, TBKF^{+}92]. potential [BWV92, CHBGL91, FMCR93, KPMG91, KO94, MCS92, SZB^{+}94b, STC93, TRLG90, TIN^{+}93, WIS^{+}94, WEB90, YKM^{+}92]. potentially [NS93]. potentiates [DFC^{+}92]. potentiation [JGB^{+}93, PWY^{+}92]. potocytosis [SFY^{+}94]. pp125FAK [BTR92, HSP93, HLC^{+}93, LHS^{+}92]. pp170 [WRDK92]. pp60c [HNC^{+}92, LVBK92, MAWM90]. pp60c-src [HNC^{+}92, LVBK92, MAWM90]. pp60src [SBN92, WP93]. pp60v [HAA93, SR92a]. pp60v-src [HAA93, SR92a]. pre [AL92, BHLA90, HWF92, HEF92, JK93, RLWK93, SYG91, TIA^{+}92, WL91, WCC^{+}90, BP91, OMS^{+}94, RT92, ZKMB93]. pre-Golgi [HWF92, HEF92, JK93, RLWK93, TIA^{+}92, WL91, WCC^{+}90]. pre-lysosomal [BHLA90]. pre-mRNA [OMS^{+}94]. pre-mRNAs [ZKMB93]. pre-pro-transforming [SYG91]. pre-replication [AL92]. pre-rRNA [RT92]. preadipocytes [KCDR91]. preblastoderm [BTS93]. precede [VDTT91]. precedes [CMS^{+}90b, PNB91, WG92a]. precipitate [BBB91]. precursor [BHS90, CL91b, DUVV93, MLN93, MC91a, MC93, NRS^{+}93a, PZV^{+}91, PMKH91, SRK^{+}94, ZLG^{+}90]. precursors [IKFR91, MMI^{+}91]. predicted [SGWE94]. Predominance [AB90]. predominantly [KMN92, MK93, NIK^{+}93]. preendosomal [HSvD91]. preexisting [CW90a, CL93, KRHY93, SNE90]. Preferential [KKES91, LSN^{+}94]. preferentially [BDGC^{+}93, GWMP91, HSLF^{+}92, YMO92, ZAH^{+}94, ZLC^{+}93]. preformed [HW92]. prefusion [WHLW90]. pregnancy [RGKG90, RRD93]. pregnenolone [ASG^{+}93]. prehair [WA93]. preliminary [BGH90]. prelysosomal [PLCB91]. prenucleolar [BDS92]. prenylated [FKLS92]. preparation [HDDS94]. preprotein [WBS^{+}91a]. preproteins [GVK^{+}93, VGG^{+}93]. presecretory [WW93]. presence [AB90, NEL^{+}91, OA92, SNE90, SRHK94, VKS^{+}90, VMB92, WRHW92]. present [ABD^{+}94, BTH^{+}90, CSP90, DJA^{+}90, DMM^{+}91, HG92, SST92a, TGB90, TRD^{+}92]. presentation [HG92, REW90, SL90b]. Presequence [VGG^{+}93]. Preservation [XL91]. prestimulated [KHHR92]. prevacuolar [VHE93]. prevent [Cit92]. prevented [PSN^{+}91]. Preventing [Blo93, FABS90, PMBL^{+}92, SFY^{+}94]. prevents [CCT94, FBS^{+}91, KWW^{+}94, LA92, MMH^{+}92, YDS92]. previously [RKA93]. previtellogenic [CPM^{+}91]. primarily [DFLK94, MLR^{+}92, PLCB91, SB93b]. Primary [CSC92, ECM^{+}92, JGO^{+}91, RB94, SOH^{+}90, SDPH90, SVK^{+}92, AC93, AGB^{+}91, DOW92, EP94, GMQ^{+}93, IFLG90, JHGR93, KG94, KCDR91, KGM92, KM92d, MHR94, NDP^{+}91, NFYI93, OTT92, OBBS90, PTD90, RLS92, RA94, SSRM90, TRS^{+}90, TISH94, VNS^{+}94, WC90a, YYY^{+}91, YED^{+}93]. primate [KCR^{+}94]. primitive [AW92, SGGC^{+}90]. primordial [CLS^{+}92]. principal [SWS^{+}92, ZABM94]. principally [BPTS94]. prion [BST^{+}90, SHH94, TSP90]. pro [BML^{+}94, MGD^{+}92, NDdC92, SYG91, BLT^{+}90]. pro-urokinase [BML^{+}94]. probe [HTBF94, WIF^{+}90]. probed [GPH^{+}90]. probes [THBW94]. procedure [VVB^{+}92]. proceeds [DHPG94]. process [BTBO94, BMS92, BHF90, LLK91, MPM93, MVG91, OD93, SWL^{+}92, SNRS93, TIA^{+}92, WG92a]. Processes [BPK91, ABR^{+}92, CL94, FCD^{+}93, KKL^{+}91, MTP^{+}92, OTT92, WSL91]. processing [BHS90, BMPW90, DMJ^{+}94, FRC^{+}91, FFS^{+}93, GSS^{+}92, HG92, HTD90, IKFR91, JKS93, KL90, MJM92, MME93, NSHN92, OTX^{+}94, OTWH92, RAGC94, RT92, SL90b, SC90b, WF91, XS93]. procoagulant [FABS90]. procoagulant-rich [FABS90]. procollagen [CJB^{+}93, NSHN92, SMRG91, SSB^{+}91]. prodermorphin [SCV^{+}91]. produce [GdLMS94, KES90]. produced [GJK^{+}91, GH91b, MWLC^{+}94, PH92, RH92a, RRH91, TW93]. produces [BTRB93, COTC94, OSK^{+}92]. producing [WCR93]. product [BZT94, BIPG91, CMM94, DMJ^{+}94, HBS92, HTD90, KHP91a, KDT93, LVBK92, OOM90, PCG^{+}94, ROE^{+}90, WKMG92]. Production [Has93, OSKN94, AR91, BRF93, GLAB93, HLB90, IITG94, RJJ^{+}94, SBWV90, SSS^{+}90, SAVK94, SSB^{+}91, TSR90]. Products [Hol93, HHLS92, TQL^{+}90, YLWS94]. proenkephalin [SCV^{+}91]. proenzymes [VHE93]. profilactin [CBRW92]. profile [KWW91]. profilin [BHBG94, CBRW92, GT94, GCMDP91, HLA^{+}90, MAA^{+}94, SY90b, VAL^{+}93]. profilin-agarose [MAA^{+}94]. profilin-deficient [HLA^{+}90]. progenitor [IRH^{+}91, MWS^{+}93, PSE^{+}90, WN92, WWN92]. progenitors [GBB^{+}94]. Progesterone [KCR^{+}94, CSMG90, PALM92]. Progesterone-dependent [KCR^{+}94]. progestomedin [KCR^{+}94]. programmed [DJ93a, EZF^{+}94, GSBS92, Ree94, ZZL^{+}91]. programs [HKS^{+}92, Mil90]. progress [SR92c]. progression [GS94, JS92, KATS90, KDN94, LOOM93, LK93b, MG94a, MJM91, NGP^{+}91, NSKM90]. Progressive [BCG^{+}92]. Prohormone [XS93, JKS93, WF91]. prohormone-cleaving [WF91]. prohormones [JKS93]. proinsulin [OHP^{+}94]. proinsulins [QORM91]. proliferating [FA91, NEL^{+}91, ZDZSBT94]. proliferation [ACC^{+}94, AGB^{+}91, BRIA^{+}94, BCET^{+}92, CW90c, DMC^{+}91, GBB^{+}94, GBN^{+}93, GS94, HR90a, LYC^{+}94, LCK^{+}92, LCLG94, MSSD93, MVM90, MM92, PA91, PH94a, SDW^{+}93, SBWV90, VVR^{+}90, WSB93, WMvdK94, YRBBP90, YM90b, ZDO^{+}91, dBBHA94]. proliferation-associated [SDW^{+}93]. proliferative [HKS^{+}92, PRC^{+}90]. proliferator [BRD94]. proline [BZT94, LCF92]. prolyl [SXM94]. promastigotes [MSD92b, SIR^{+}94]. prometaphase [HAS90]. prominent [VBV^{+}90]. promote [CL91a, GH91b, HWLT94, HMS^{+}90, SLWF91, WKY^{+}93]. promoter [GMLD92, NDdC92, PLB^{+}92, PAC^{+}92, aUFQ^{+}92]. promoters [GSM93]. promotes [ACC^{+}94, BSSL^{+}91, CMS^{+}90a, CLM91, DGGR92, EMC^{+}90, GHH90, HFS92, KS92, KDO^{+}94, LHT^{+}91, LLK91, LCT^{+}93, MWSP91, NT94, NRS^{+}93a, OBBS90, PGAR90, VGC94, WFCN94, WNM^{+}94, YED^{+}93, dWLS91]. promoting [CNSC90, DFC^{+}92, FCL^{+}94, KS90, RA94, YM90a]. proof [OTT92]. propeptide [SMRG91, VWS90, VSB^{+}94]. proper [CM94b, CC93b, FPP91, TCS^{+}94]. properdin [PPJF91, TRD^{+}92]. properdin-like [PPJF91]. Properties [BI90, AMBN91, ALE^{+}92, BA93, BSEF^{+}93, BMWT93, DR91, DCY^{+}94, DB91, DES^{+}91, FCL^{+}94, FFW91, GWMC90, GRFB94, HNR90, JETS91, JLL^{+}94, KFW^{+}91, KR91, MMLG94, MSF90, PMKH91, PMG^{+}92, PBM^{+}90, SMKH91, SHTL93, SRR90, SRP90, TH92b, TH92c, WB92a]. prophase [BWLK93, SBK94]. proposed [ATV^{+}93]. proprotein [OTWH92]. prosomatostatin [XS93]. prostate [JZAVP93]. protease [CJB^{+}93, EMC^{+}90, FSSA92, RHF91, TCCT91, VSB^{+}94, WF91, YMO92]. proteasomes [KGK90]. protect [WKY^{+}93]. protected [LSC^{+}93]. protective [HR94b]. Protein [Cit92, FBML90, FBL92, HdHD^{+}93, KA92, LFR93, RG92, SCM90, SMCR90, SFY^{+}94, SR92e, TCDH93, WBS^{+}91a, AYM^{+}93, AL92, AMG90, AAM92, AU90, AC93, AAM^{+}93, ASDC91, AGKC92, AC92, ABD^{+}94, Apg91, AKC92, ASE90, AL91, ASNH93, AGB^{+}91, BEGN91, BAP^{+}90, BT92, BS94a, BTH^{+}90, BCBK92, BDMPJ94, BZT94, BF90, BSF90, BBKR94, BFLS94, BMPW90, BRO^{+}91a, BMM^{+}94, BN91, BKFT93, BDB90, BBS92, BS94c, BGI^{+}90, BTG^{+}93, BHFS93, BS93b, BHHG90, BBJ^{+}91, BM91b, BDK92, BDJ94, BK91, BFTR93, BSP^{+}94, CD91, CSKZ94, CLCC90, CLDV90, CSP90, CXS^{+}94, CL94, CLP^{+}94, CSSL93, CJO92, CSMG90, CL92, CYCA90, CL91b, CJB^{+}93, CSR^{+}92, CG91, CSC92, CBE90, CB92, CRK^{+}93, CKL^{+}91, CPB94, CHG94, DYV93, DDW90, DR91, DPND90, DMB92a, DUVV93, DW91, DVT92, DJA^{+}90, DAR93, DJ93b, DY93]. protein [DLSB^{+}90, DLSK91, DCL^{+}92, DEH^{+}91, DWY^{+}93, DMJ^{+}94, DRJ^{+}91, DKG92, EK92, EOL^{+}90, EJ93, FB93, FF93, FFW91, FT93, FAS^{+}93, Fow90, FMR^{+}92, FRC^{+}91, FMLD92, FFR^{+}91, FNM^{+}92, FJN93, FHI^{+}93, GK90, GVK^{+}93, GCZ^{+}92, GHR^{+}90, GtHG^{+}94, GMGPM94, GPF94, GWO^{+}93, GFWM92, GSC^{+}92, GWMP91, GW94a, GW94b, GFH90, GSR^{+}93, GYE^{+}90, GFL92, GE91, GC90, GG92, GFW92, GRFB94, HCW94, HFL94, HRT^{+}91, HWB93, HR94a, HLRBE93, HC94, HSCR94, HP92, HNvdK^{+}94, HW90, HMY92, HPP^{+}93a, HAS92, Hen90, HPWN93, HSCN94, HHF93, HNUO92, HD91b, HNPN94, HNP94, HR94b, HLL94, HSSC92, HVK91, HH94b, HIJ90, HHG^{+}94, HLC^{+}93, HLR92, Hur90, IMK^{+}92, IKL^{+}94, IKFR91, IYN^{+}91, INY^{+}93, ISDM92, JvCH^{+}94, JO92, JHK^{+}91, JG92, JGK^{+}94, JG94, JGLK90, JLC93, JKLG94, JS92, JGB^{+}93, JBJH^{+}92, KT92]. protein [KM92a, KISY91, KYK^{+}94, KBG91, KGSB92, KWK92, KDKF93, KFC^{+}94, KSK^{+}93, KCDR91, KCB^{+}92, KDN94, KKBK92, KHHC93, KA94, KDP^{+}92, LA92, LRR^{+}92, LCMP90, LM91, LXM91, LD92, LWV92, LFG^{+}90, LH92, LMB^{+}94a, LB94, LML^{+}94, LYK^{+}94, LGSB93, LHS^{+}92, LTS93, LNS^{+}94, LCLG94, LTWH92, Low92, MSC91, Mac92, MWSP91, Mah91, MSS^{+}94, MK93, MJG^{+}91, MMWG94, MMS93, MY92, MSW^{+}94, MSPB90, MG94b, MWH^{+}91, MB90, MB94, MPEG93, MVCC93, MNV93, MKB^{+}94, MBB^{+}93, MBHG91, MGG93, MGPPH93, MHV92b, MNB92, MFS^{+}94, MM93a, MVV^{+}91, MCS92, MIU^{+}92, MSF90, MAAO92, MHDT94, MPVB91, MW92, MMY^{+}92, MCEG90, NMPN90, NSHN92, NHO^{+}91, NMP^{+}92, NF90, NEL^{+}91, NRS^{+}93a, NSKM90, NFL90, NRS93b, NDM^{+}94, OMRM93, OLKD90, OUS^{+}93, ONN91, OWF^{+}93, OSJS94, OS92, PC91a, PPL^{+}93, PSRS94, PSZ^{+}93, PWY^{+}92, PIRB92, PL90, PS94]. protein [PET^{+}91, PTTA90, PC91b, PG92, PMBL^{+}92, PDTG92, PBCK93, PYA90, PP91, PSG94, PPZ^{+}93, PNB91, PMM91, PNR^{+}93, PIS94, PMTB94, PDSB92, PRAK94, PBCP91, PB90, PE93, PAFC92, RBW93, RMK^{+}94, ROE^{+}90, RB92, RDDR93, RBR90, RK90a, RDH^{+}92, RTC^{+}94, REW90, RGKG90, RNS92, RWL^{+}90, RMB90, RT92, SGWE94, SWL^{+}92, SPH^{+}91, SYO^{+}91, SM93a, SZB94a, SMC92, SKSC94, SGC^{+}94, SGW91a, SHDS93, SDW^{+}93, SAF90, SBP90, SB93a, SFC^{+}94, Sch94, SCM^{+}93, SP92, SON^{+}94, SHTL93, SSDK^{+}94, SNFN91, SRK^{+}94, SHH94, SKSD^{+}94, SK94b, SSWS92, SY94b, SCF^{+}93, SMJ^{+}94, SB91, SVK^{+}92, SSC90, SMBG92, SPB90, SOKS91, SLNN^{+}91, SKD^{+}91, SdAN^{+}91, SRM94, SJM^{+}94, SLK93, SS90b, SAM^{+}91, SM91c, TASJ92, TMHKO94, TOFH94, TNS^{+}94, TQL^{+}90, TOA^{+}93, TKHM91, TH92a, TB90, TFT94, TAWJ91, TVG^{+}93, TSH^{+}93]. protein [TZBV91, TYSS92, TLSW93, TIA93, TSR93, TAVC91, TGB90, Tur91, TWV94, UOKH93, UKH^{+}91, UYA^{+}91, VWT^{+}91, VIF^{+}93, VRE^{+}92, VGE90, VKRD94, VSVP93b, VGM^{+}93, VPC^{+}91, VLH^{+}93, VGG^{+}93, WRDK92, WLPB92, WG92a, WBWH93, WVH94, WTH^{+}94, WCR92, WHLW90, WRHW92, WGP^{+}93, WW94a, WMG^{+}93, WSL91, WSM93, WRP^{+}91, WMvdK94, WS93a, WE92, WKM^{+}92, WS93b, WHN^{+}94, WRB^{+}94, WCM^{+}93, WC90a, WC93, WDRM93, WKS^{+}93, WDB^{+}92, WF94, WTRD92, WEB90, WBR94, WP93, WGTCY94, XMW^{+}93, YKI^{+}91, YTN^{+}94, YM93, YLS92, YSK^{+}94, YKM^{+}92, YNST93, YH93, YFS92, YMW^{+}90, YAJ90, ZS94, ZSM^{+}93, ZSC91, ZFF92, ZBW93, aUFQ^{+}92, dWLS91, dSK90, dBBHA94, dHRB^{+}93, dSBH90, dSBH93, dlRKR^{+}90, dlR91, dHvBeHVB92, vZDF^{+}93, BG91b, JGL^{+}93, WDRM93]. protein-1 [ASNH93]. protein-2 [KYK^{+}94, YKI^{+}91]. protein-5 [JGB^{+}93]. protein-6 [WBWH93]. protein-activating [WW94a]. protein-antibody [HIJ90]. protein-mediated [IKL^{+}94]. proteinase [TBT^{+}92, TSR90]. proteinase-deficient [TBT^{+}92]. proteinases [STA^{+}94, TBW92]. Proteins [MC92, TRH^{+}92, APH92, ATSE90, AS94, BWV92, BGGW^{+}94, BKB^{+}90, BKWD94, BCKS92, BSW91, BOM94, BBD^{+}90, BST^{+}90, BBB91, CRWS93, CH91, CH92, CCG^{+}94, CL91b, CYC91, CSC92, CC90b, DMB93, DP90, DAR93, ES91, ENS91, Eri93, ET91, FKG94, FYDB93, FZ90, FKLS92, FMB^{+}91, GLQRB91, GH91a, GBS90, GAvdM91, GHT92, GKKS90, GMQ^{+}93, GAS90b, HSS93, HK91, HM92b, HSCN94, HGW^{+}91, HFD91, HAG^{+}93, IMK^{+}92, JNP93, JAB^{+}91, JKH94, JAS^{+}90, KD92, KKG^{+}91, KAF^{+}93, KGA^{+}93, KHF94, KKFD90, KKW^{+}93, LQNRB90, LWV92, LTP^{+}92, LH93, LRAG92, LSA91, LPA^{+}93, MPR90, MECE93, MH94, MF90, MH92, MB92, MMD92a, MMD93a, MLN93, MGS90, MEM94, MC91c, MPTW90, MPVB91, MYC92, MMD^{+}93b, MKF91, NT90, NPST94, NF92, NPR^{+}94, OC91, PHMH92, PBM^{+}94, PMG^{+}92, PZB^{+}94, PMA91, RSC^{+}93, RK94, RKB^{+}90]. proteins [RMR92, RLS92, SCMB92, SBB^{+}91a, SSTL93, SP94, SDW^{+}93, SLL^{+}90, SPBB92, SRP90, SKM94, SPW93, TSP90, THBW94, TAWJ91, TFP94, TM90, TIN^{+}93, TRD^{+}92, VMB92, VHB93, VM94, VMR90, WWS^{+}94, WNBAS90, WRB92, WBW^{+}94, WG92b, WBH^{+}90, WDP^{+}94, WW93, WKN92, WB92a, WTS93, Wu93, YDS92, YLWS94, ZWW90, ZCS^{+}91, ZLC^{+}93, dAWL^{+}90, vZDF^{+}93]. proteins-dependent [BSW91]. proteoglycan [BBJ^{+}91, BGRW91, CS90, CES^{+}92, CLG^{+}92, DLD^{+}90, DvdSM^{+}92, DKM^{+}92, FSK^{+}92, FCL^{+}94, FMK^{+}94, HK91, ISF^{+}92, JKH94, KT92, MFS^{+}94, MMD91, NDP^{+}91, SF91, SDH90, SLL94, SdAN^{+}91]. proteoglycans [BLR^{+}92, BFTR93, GFM93, LEM93, SWM^{+}92, dAWS^{+}90]. proteolipid [WDRM93]. proteoliposome [BS93b]. Proteolysis [MPVL92, PKPM90, PAL^{+}90]. Proteolytic [BMPW90, GSS^{+}92, WF93, HVNC92, OK90, PB91, PBM^{+}90, SR90a]. proto [LVBK92, TOA^{+}91]. proto-oncogene [LVBK92]. proto-oncogenic [TOA^{+}91]. protofilament [CMV^{+}92, RMMH93]. protofilaments [KM92b]. Proton [HZC93, SWS^{+}92, ST94, VKS^{+}90]. protooncogenes [ME91]. protozoa [MC94]. protozoan [OMBL90]. protruding [KBL91]. protuberances [KKBL92]. provide [BSHB90, DMB93, FHGN91]. provides [PAL^{+}90, STA^{+}94]. proximal [BSB92, DA92, HLH^{+}92, KIS^{+}94, PR91, ZSH^{+}93]. proximity [HGW^{+}91]. pS2 [LWK^{+}93]. psbA [MCDY94]. pseudopodial [RGT^{+}93]. PTK [MMDM93, MOMM92]. PtK1 [KRB91, MAFR93]. PtK2 [CM93]. PTP [BKFT93]. ptx1 [HW93b]. pull [SSS93]. pump [Fuj93, ST94, VKS^{+}90]. pumping [SWS^{+}92]. pumps [HZC93, JJB^{+}93]. pupal [WA93]. Purification [AC92, CYR^{+}94, FL90, HLA^{+}90, MAA^{+}94, PSE^{+}90, SHTL93, TKHM91, ZVF91, BBK90, KWS94, RB94, SKGU91, SOKS91]. Purified [BZ90, BGB94, KLBBK91, ND92, SCS92]. purine [WWHA91]. purity [GH91a]. Purkinje [GSV93, LCT^{+}93, PBB^{+}91, SRV^{+}90, VPC^{+}91, WAS^{+}91]. purse [BFM93]. push [SSS93]. push-pull [SSS93]. putative [BLR^{+}94, BCYC94, CM94b, LA92, LH92, LB94, LGI^{+}93, MH93a, MB94, RHF91, SGWE94, SNFN91, SLK93, TB90]. pyruvate [NFYI93]. pyruvate/alanine [NFYI93].

quail [OGM^{+}93]. Quality [HH94a, dSBH90]. quantification [SW93a]. quantitative [AR91, FE90, GB92, KW92b, LFF91, SRV^{+}90, TLFS92]. quaternary [WG92a]. Quenched [KBP91]. quick [NSH90]. quick-freeze [NSH90]. quiescent [DGGG93, LM94].

r [SGBN^{+}93, WCF92]. R/K [WCF92]. R1 [RLB^{+}91]. R2D5 [NIK^{+}93]. rab [ZJA^{+}94]. rab1 [NDM^{+}94, TBKF^{+}92, PNPB94, PNM^{+}94, WNM^{+}94]. Rab17 [LJP^{+}93]. Rab1b [PCP^{+}91]. rab2 [TBKF^{+}92]. rab3 [JGK^{+}94, WBT^{+}94]. rab3-like [JGK^{+}94]. Rab3A [MTC^{+}91]. rab5 [SKSD^{+}94]. rab5-like [SKSD^{+}94]. rab6 [JCSH93, MSS^{+}94]. Rab8 [HPP^{+}93b]. rab8p [HdHD^{+}93]. Rab9 [RSS^{+}94]. rabbit [JAS^{+}90, PRB^{+}91, RLB^{+}91, SVR^{+}90, YAJ90, YAJ91]. Radial [SAWL90, DAR93, GOP^{+}94, KIO^{+}94]. radicals [PZP^{+}91]. radius [FMR90]. Radixin [FNS^{+}91, SYO^{+}91]. Raf [FMD93]. raikovi [OMBL90]. raised [FL92]. Ran [MPEG93, RDDR93]. Ran/ [RDDR93]. Ran/TC4 [MPEG93]. randomly [PMM90]. range [ZZP94]. Ranvier [KDTB90, KB91b]. RAP1 [KLC^{+}92]. rapamycin [BZT94]. rapamycin-binding [BZT94]. Rapid [DGF^{+}94, GB90, BSHW91, BPDM90, BZ90, FFB90, GSMK94, JM92, KBP91, LLK91, MVG91, PSG^{+}90, VGC94, YC93, dSK90]. rapidly [MKG93, WN92]. ras [BMW^{+}93, KDT^{+}91, QPWG91, SFF93, TAVC91, SZKM91, VWT^{+}91]. ras-induced [QPWG91]. ras-transformed [TAVC91]. rasVal [LSAH90]. Rat [BKSB94, Apg91, AGGS92, APP^{+}93, ASNH93, BML^{+}90, Bis90, BRO^{+}91a, BT93, CVK^{+}94, DIYSS92, DS90a, DJ93b, DKL^{+}90, DFC^{+}94, DJF93, DYS^{+}90, ET94, EFD^{+}93, EERM^{+}91, FSM^{+}93, FSY^{+}91, GCZ^{+}92, GBOB90, GHBA^{+}93, GTH90, HHH^{+}93, HBOV93, HD90, ILH^{+}90, JRK^{+}94, JRCW^{+}90, JZAVP93, JRSB90, KWS94, KTG90, KKG^{+}90, KKW^{+}93, KOB91, LCK^{+}92, LVTB93, LGH91, LHFV93, MDKB94, MPK^{+}94, NCM^{+}93, NHK^{+}94, OHC^{+}93, OSK^{+}92, PZV^{+}91, PPHS^{+}92, PNB91, PCG^{+}94, RGJ91, REW90, RGKG90, RGR90, SSRM90, SM90, SST92b, SGG^{+}91, SGM^{+}90a, SLNN^{+}91, TISH94, VPVC91, VGRC^{+}92, VBA90, WS91a, WMvdK94, WBL^{+}91, WBS91b, YOWM91, ZABM94, ZLC^{+}93, dCV90, dAWS^{+}90, dTF90, PC90]. rate [CC90a, LB92a, SWHCW91, VFR94, WPS91]. rates [GK92a]. ratio [SAH^{+}92, SLK91]. ratiometric [MOKF92]. Rb [MSSD93]. RDEC [RLB^{+}91]. RDEC-1 [RLB^{+}91]. rds [AMMW92]. re [Blo93, OH93]. re-evaluation [OH93]. re-replication [Blo93]. reach [GMGPM94]. reaching [HWvF90]. reacting [KHvdL^{+}92]. reaction [GSS^{+}93, KTG90]. reactions [CSV^{+}92, HCW94]. reactivation [MHYC90]. reactive [NS93, Mac92]. reacts [RBW93, ZSM^{+}93]. reading [ZZR94]. reagents [GPH^{+}90]. real [KII^{+}91, WPS91]. real-time [KII^{+}91]. rearrangement [WG92a]. rearrangements [BS94c, Har92, PMW92]. reassembled [SSH94]. reassembles [SPW93]. reassembly [CC93a, VYM93]. reassociation [TW93]. Receptor [AKM94, EUH91, KYBC94, AKK^{+}93, ATD^{+}90, BSHW91, BSW^{+}92, BP93, BSG^{+}92, BTRB93, BRWB91, BRO^{+}91a, BSM90, BM90, BM91a, BGLB94, BHC^{+}94, BFL90, BKFT93, BSR^{+}90, BML^{+}94, CKP91, CPB^{+}93, CSvdE93, CLD^{+}91, CRKA92, CNSC90, CH92, CGW94, CXS^{+}94, CE94, CG93, CTQ91, CPP^{+}94, CMK^{+}90, CNU92, DS90b, DG90, DDFC94, DVT92, DSdF^{+}90, EMC^{+}90, ECAG93, ESL^{+}90, FLUS92, FMPP90, FPP91, FL90, FL92, FGH92, FWD^{+}92, Fro91, FGS91, FNM^{+}92, FFS^{+}93, GCM^{+}94, GJJM94, GDB^{+}94, GHGP^{+}92, GPH^{+}90, GSV93, GAHB94, GK92a, GeKdV^{+}91, GI93, GEKL^{+}94, GZC^{+}91, GFV^{+}91, GSM93, HWLT94, HFN92, HHO^{+}92, HNvdK^{+}94, HK91, HK94, HKSL90, HWvF90, HCYK91, HSUS90, HAK94, HVNC92, ILH^{+}90, ITSR90, ISF^{+}91, IKL^{+}94, JM92, JS90, JRCW^{+}90, JSM^{+}90, JK92, JSA^{+}93, JCO^{+}94, KT92, KSST91, KHW93, KFW^{+}91, KMSW91, Knu93, KCC90]. receptor [KEBD91, KHHC93, LAY92, LTYY94, LHH^{+}91, LSP^{+}91, LD91, LLSM92, LWV92, LKXS93, LG93a, LWHL93, LASB^{+}93, MRBP^{+}90, MTS^{+}91, MCM93, MPAF91, MWH^{+}91, MH93b, MMH^{+}92, MNB92, MMR91, MFS^{+}94, MMT91, MIU^{+}92, MVM91, MJM91, MM93b, NHTP90, NIK^{+}93, ODPL93, OPL94, OSK^{+}92, OH90, OW90, OKH92, OMBL90, OVB^{+}94, OCMB91, PIRB92, PPHS^{+}92, PYH^{+}90, PRGC91, PBD93, PSS^{+}93, PALM92, RRG^{+}94, RLB^{+}91, RG92, RRCR93, RSS^{+}94, RUH^{+}94, RGR90, RELS^{+}91, RNG94, RGHC91, RYKA90, RYK^{+}90, SK94a, SHS^{+}93, SRV^{+}90, SM93a, SVR^{+}90, SC90a, SBP90, SGBN^{+}93, SCM^{+}93, SBK^{+}91, SM93b, SFY^{+}94, SB93c, SLM^{+}90, SMWB93, SWHCW91, SKK^{+}91, SW93b, SF90b, SSKT^{+}92, SBL^{+}91, TMP^{+}91, TWQ^{+}93, TCDH93, TBL^{+}92, TU90, TSLR90, TGR^{+}92, TSR90, TRP^{+}93, WLPB92, Wal94, WSC94, WIF^{+}91, WSB93, WOC90, WL91, WKY^{+}93, WEB90, YPH^{+}92]. receptor [YH93, YGM^{+}94, YAJ91, YL93, ZMV^{+}93, ZZR94, dHvBeHVB92, GMGPM94, JPJF91, MWS^{+}93, TMI^{+}91, WIF^{+}90]. receptor-associated [MIU^{+}92]. receptor-beta [ZZR94]. receptor-binding [BSR^{+}90]. receptor-containing [GCM^{+}94]. receptor-deficient [BSG^{+}92]. receptor-directed [Knu93]. receptor-like [FNM^{+}92]. Receptor-mediated [AKM94, CRKA92, CXS^{+}94, GeKdV^{+}91, HK94, HCYK91, RRCR93, SC90a, SFY^{+}94, YGM^{+}94, GMGPM94]. receptor-type [BKFT93, MFS^{+}94]. receptor/c [MWS^{+}93]. receptor/c-erbA [MWS^{+}93]. receptor/junctional [MTS^{+}91]. receptors [APP^{+}93, BSHW91, BPDM90, BRIA^{+}94, CSJD91, CDS^{+}91, CvdE91, CPA^{+}93, CSRS^{+}90, DHS93, FFB^{+}93, GAvdM91, HSvD93b, HMC91, HMLL94, HCS94, HGS^{+}94, JTS^{+}91, JWHPM93, KHV^{+}93, LCPAM94, LC91, MBC^{+}91, MYM94, MELD94, MC91b, MIU^{+}92, PNR^{+}93, PMD^{+}93, PRB^{+}91, PMM90, RB90, RA94, SCL91, SJIC92, SWM^{+}92, SSK^{+}90, SVD90, TDSP93, TTR^{+}90, VSVP93a, VBA90, WAS^{+}91, WCC^{+}90, WHH^{+}93, dCQTR91]. Recessive [FM94b, GWMC90, SSB93, WC90b]. reciprocal [MSI^{+}94, ZFA^{+}93]. Recognition [HMS^{+}90, SSA^{+}90, CLP^{+}94, CG93, DGAB^{+}93, EWP^{+}92, EWP^{+}93a, FS94, FWBS93, HLH^{+}92, HD91b, HSM^{+}93, KSL^{+}94, KTR90, LXM91, LPA^{+}93, MCF91, MNB92, NS93, RG92, RWL^{+}90, SCO^{+}91, SADJGP94, SLS^{+}90, SLM^{+}90, VHW^{+}92, WK92, YMO92, YED^{+}93, ZBW93]. recognize [FWD^{+}92, GKKS90]. recognized [PBCP91]. recognizes [LBF92, OKH92, RMG90]. Recombinant [GKY^{+}94, PWY^{+}92, YKI^{+}91, YED^{+}93, GSS^{+}92, HW90, KRCT93, LGPM90, OBBS90, PSG^{+}90, RCJ^{+}94, RKB^{+}90, SRP90, SOY93]. reconstituted [BS93b, CDS^{+}91, HDDS94, JKT91, JS90]. Reconstitution [BHFS93, KGDE94, LMSA91, MM91, ONN91, SHDS93, SS92a, SSKT^{+}92, TK94, VGE90, Wat90]. reconstitutional [OW90]. reconstruction [DMM93, DWY^{+}93, FPGS91, Lut91, MR92, RRG^{+}94, TU90, VCL93, YDO^{+}90]. Recovery [RHR90, HKSL90]. recruitment [FAWM94, TYM^{+}92, TOK93, CW90a]. recycle [CvdE91]. Recycling [DP90, SKK^{+}91, AKM94, CSS^{+}94, FLUS92, HGS^{+}94, KP90, MF90, MCM93, MTP^{+}92, MLR^{+}92, MPM93, MCM92b, MCJ^{+}93, PMD^{+}93, RSS^{+}94]. red [GRC^{+}91, WBAP91]. redistributed [WKMG92]. redistributes [FA92, RTC^{+}94]. Redistribution [CJC93, JGK^{+}94, TTVV^{+}90, BFR^{+}92, LFC94, NHA^{+}92, PDB^{+}91, PNB91, RB92, VDTT91, WMB^{+}94]. reduced [BMK^{+}90, BCG^{+}92, FGS91, GHBA^{+}93, JK93, KAG94, LKS94, OC91, SOPA94, VFR94]. reduces [BMTTG94]. Reducing [HFN92]. reductase [HR94a, LBL^{+}91, PAC^{+}92, ROBN^{+}92]. reduction [BG91b]. redundancy [Gol93]. reentry [HKS^{+}92]. reexpressed [GHBA^{+}93, PLS92]. Reexpression [CL94]. reflecting [KK92, LYC^{+}94]. regenerating [DR91, LHFV93, LC94]. regeneration [HSS93, KKW^{+}93, PLS92, SST92b, SEMH92]. region [ACC^{+}94, BSHW91, BSW^{+}92, BMTTG94, CHM^{+}92, DJF93, DMJ^{+}94, EWP^{+}92, GF94, HWB93, HMS^{+}90, IFLG90, KIO^{+}94, MCDY94, MMWG94, OTX^{+}94, SCE91, SYG91, SB93c, SDR^{+}94, TBFP92, WK92, WS92, YYY^{+}91, ZZR94]. region-specific [DJF93]. Regional [BML^{+}90, SLNN^{+}91, CSR^{+}92, NPST94]. regions [BD92, BSB92, CH92, HNH91, JO92, LB94, LG93b, MCG92, RKB^{+}90, SRP90, YM93]. regular [GA92]. regulatable [HSP^{+}91, PTS^{+}92, SGG^{+}91]. regulate [ASR91, BS93a, DDW90, FMLD92, KM94, SCO^{+}91, SPBB92, SDCD93, TWC93, TAVC91, WBS91b, WA93, ZBPV92]. Regulated [AL91, HR94a, LMP91, Mac92, MHG93, PJS94a, SGVS^{+}93, SNE90, WG93, AB90, BBD^{+}90, CFFL93, CM94a, CKB93, CH91, CPB94, DAM^{+}93, DY93, DP93, DYS^{+}90, GF94, GTC^{+}90, GK92b, HM92b, HTSP94, Hol93, HdHD^{+}93, HFS^{+}93, HKS90, JP94, JGK^{+}94, vdWvKvKdB^{+}92, KDKF93, KCB^{+}92, KHHC93, LKSR94, MPVL92, MTS^{+}91, MHG^{+}90, MLM^{+}93, NPR^{+}94, SBB^{+}91a, SRG^{+}94, SHLS93, SCV^{+}91, SON^{+}94, SA90, SWF^{+}92, SFS90, SR92d, SB90, TRH^{+}92, WDRM93, WMCW92, YOWM91, FF93, SLH92]. regulates [BR94a, BGLB94, CdPA^{+}93, CSMdPSM94, CSM90, DWJP93, HMB^{+}93, LH92, PTK^{+}93, PCP^{+}91, RDDR93, SDK^{+}93, SdAN^{+}91, SSK^{+}93, TBW92, TBL^{+}92, TBVS92]. Regulating [SBL^{+}91, JKLG94, KJT91, KDN94, STL^{+}90, VVR^{+}90]. Regulation [Apg91, ASMC^{+}92, BHS94, BLR^{+}92, BPH^{+}90, BG94, BDK92, FMPP90, FKLS92, FRL93, GMM^{+}93, GDKP92, Ham94, HHG^{+}94, JFA^{+}91, KSK^{+}93, LAY92, LKXS93, MMT91, PPL^{+}93, PKN91, RJJ^{+}94, SM93b, SY94b, SR92b, SMJ^{+}94, SGM90b, TH90, TNT^{+}90, WJ92, WHH^{+}93, YRBBP90, ZCH90, dCV90, ACSM^{+}94, BWV92, BML^{+}90, BLW92, BB91, BEO90, CWB93, CJB^{+}93, CKL^{+}91, DPS^{+}93, DMB92a, DBTAB90, DLSK91, DCCH92, DTY92, EMGS93, FKHG93, FJRAT91, FAMR92, GKG90, GSM93, KKW^{+}93, KVF^{+}90, LR93, LSS^{+}90, MCM92a, Mah91, ML91, MMW^{+}92, MVTJ94, MM91, PBCK93, Ree94, RGHC91, SLHG93, SZP^{+}93, WMF91, WWHA91, WHFM93, YR93, ZS94, ZZR94, dCQTR91, ASG^{+}93, WS91b]. regulator [DOW92, KWW91, LGI^{+}93, WTK^{+}92]. regulators [LCM^{+}94]. regulatory [BLB^{+}91, GOP^{+}94, GTC^{+}90, JO92, LSG92, OCC94, PLB^{+}92, PMS92, PMLM94, PKG^{+}94, SBB^{+}91a, SLW^{+}92, SKM94, TRLG90, TWC94, WRA93, YYM94]. reinhardtii [BGBWO91, BU94, GSS^{+}93, HSS93, HKS90, JD91, KIO^{+}94, MCDY94, PPD92, PKG^{+}94, QYC^{+}92, QH94, SSG93, SS94, SH91, YC93]. reiterated [YYY^{+}91]. related [AWH^{+}94, BIPG91, CLCC90, DMB93, DG90, EOL^{+}90, FSK^{+}92, FGK^{+}90, HSSC92, HHLS92, KRdlR^{+}92, KWK92, KHHC93, LWV92, LB94, LVTB93, MHR94, MVCC93, MNV93, MCS92, MPVB91, NHO^{+}91, NEL^{+}91, OBC^{+}92, OS92, PSRS94, PNR^{+}93, PMTB94, RMR92, SGC^{+}94, SFC^{+}94, Sch94, SHM94, VKRD94, WB92a, YLS92, ZCH90, vdSKL^{+}93, MF90, OMRM93, TNS^{+}94, WBWH93]. relation [BGM^{+}91, CKG^{+}90, SKH90]. Relationship [CL91a, CWHH92, AMMW92, BML^{+}90, CTL91, CW90c, DBTAB90, GMB^{+}91, JKS93, KBA92, KRMK93, MBC^{+}91, NR90, SL90b, VHW^{+}92]. relationships [ASL^{+}94, WOC90]. Relative [CGF^{+}93, CMZ90, SBK^{+}91]. relaxation [HG94]. relaxed [HTBF94, LG93a]. Release [SR90a, YKY92, BPM93, BGRW91, EP94, FAF^{+}93, GHR91, HN90, MMM^{+}92, NXHMJ93, NW94, RRG^{+}94, SIR^{+}94, TS91, TTVV^{+}90, WMvdK94, YCAB93, JSA^{+}93]. released [GK92b, SY90b]. relentlessly [LPK92]. relevant [HAA93]. reliant [TSBS92]. relieve [LMB^{+}94b]. relocated [IDT^{+}92]. relocation [KQJ91, MSPB90]. Remodeling [DDW94, WKN92, CQYD94, LF93, TLSW93]. removal [FHI90, FKH^{+}93, VSB^{+}94]. remyelination [AFHDD90]. renal [JPJF91, LVA^{+}90, ZVF91]. reorganization [BPDM90, BL94, CSCSA94, KDS^{+}91, KO94, NR90, OMS^{+}94, SWL^{+}92, STH^{+}93, TJ91, TS93, YFS92]. reovirus [DWY^{+}93]. repair [MRBP^{+}90]. repeat [FPR90, GF94, JKH94, SGN^{+}93, Woo94, ZMB91]. repeated [ATBD90, DBC90, SDW^{+}93]. repeats [HFS92, PSZ^{+}93, PPJF91]. repetitive [KWFM91, RBW93]. Replacement [BMTTG94]. replica [FKH^{+}93]. replicating [PME93]. replication [AL92, Blo93, CDRL93, DMWJ90, FTPM^{+}94, HS90b, KT93, LL91, LM94, MMLG94, OHS92, RG94]. representing [HW92, PCD^{+}92, SDW^{+}93]. represents [GSFS92, HLWL92, NRS^{+}93a]. repressing [PRB^{+}91]. Repression [OR92, HGF93, ZSC91]. repressor [BMB^{+}92]. reproduction [SMCR90]. reproductive [WWHA91]. require [BMM^{+}94, HM92a, KRHY93, MMCH91, MSD92b]. required [AYM^{+}93, AA94, BHS90, BSF90, BBRE90, BR94b, BDW92, BS93b, BGH93, CKMC94, CXS^{+}94, CM94b, CG91, CC93b, CVK^{+}94, DHS93, DFLK94, FMB^{+}91, FDT^{+}91, FMS93, FPP91, FSSA92, GFWM92, GP91, GAHB94, HP92, HSP93, HWD94, HHLS92, IKL^{+}94, JS92, KN91, KYBC94, KPAY91, KHKO^{+}93, LFB94, LNS^{+}94, MECE93, MM90, MPK93, ML94, MPAF91, MT94, MGDS93, MM93a, MR94, NYA^{+}91, NHO^{+}91, NI93, ORS90a, OA92, OVB^{+}94, OSJS94, OOM90, OK90, PDA93, PK91, PLB^{+}91, PNM^{+}94, PMTB94, PRAK94, PSSK93, QPWG91, RRH91, RMR92, RT92, SC90a, SK91, SOWDK94, Sch91, SVB94, SNFN91, SDR^{+}94, SSH94, SS93, SMBG92, SLK93, SH92, SHM94, TWMS90, TCS^{+}94, TFP94, VMS^{+}91, VBvG^{+}94, VGRC^{+}92, VRE^{+}92, VTSS94, WCR92, WB94, WRB^{+}94, WKMG92, WHC^{+}93, WW93, YM90a, YPH^{+}92, ZBW93]. Requirement [FMD93, HTP^{+}92, KNR93, LFWC92, MHYK92, TSLR90, CL91b, FHvZ^{+}94, LPPCF91, LMB^{+}94b, NF90, RA94, SY94a, STA^{+}94, TK94, VBAK91, WSL91, WCM^{+}93, ZB94]. Requirements [HFD91, MC94, HLR92, KBP91, KGA^{+}93, MYM94, PHWM92, RCJ^{+}94, RS91, VFC^{+}91, WPW92b]. requires [AMG90, BB91, CGW94, DSR^{+}93, FT93, FPPD91, IKFR91, KDP^{+}90, KDT^{+}91, KLERG94, LM94, MCDY94, PSRS94, RG92, RSS^{+}94, SPH^{+}91, SM92a, SL90a, WF90, YC93]. requiring [BSHW91, DS90c]. RER [KLERG94]. RERMS [NRS^{+}93a]. rescue [CBM^{+}94, GPS92, MK91b]. rescues [BG91a, CL94]. resembles [MWG92, SGC^{+}94]. resembling [KT92]. reservoir [CHBGL91, HKS90]. reside [VFC^{+}91]. residence [BMTTG94]. resident [BKR90]. resides [OWF^{+}93]. residue [BSR^{+}90, CL91b, WF90, WLH92]. Residues [VTSS94, FMLD92, NRS93b, OVB^{+}94, SVB94, SR92a]. resistance [JP94, PCD^{+}92, YCBM94]. resistant [KRB91, PYA90]. Resolution [HM92b, DMLR^{+}94, GA92, RKB^{+}90, SGW^{+}91b]. resolved [FKT93, KB91b, MMM91]. resorbing [SAVK94]. resorption [LHH^{+}91, MHG^{+}90, SSG^{+}90]. respectively [MCG92]. respiratory [BLSR92, ZB94]. response [Ada92, ARWD93, Apg91, BGB94, BMW^{+}93, CL94, CSMG90, CGF^{+}93, ES92, FKHG93, GZC^{+}91, HBP^{+}93, HNH91, LFC94, Mac92, OB93, PCD^{+}92, SWD94, SLHG93, SYCA94, VGMS91, VGRC^{+}92, WSL91]. responses [BMK^{+}90, HWL^{+}90, LSG92, SFF93, SVD90]. responsible [CCFZ92, KZS94, MMW^{+}92]. responsive [DFC^{+}94, HFS^{+}93, KHS^{+}90]. Resting [THH^{+}90, CSNZ92, HD91a]. restore [MHP90]. restrict [CO91, VIF^{+}93]. Restricted [CJO92, RH92a, ZBS^{+}94, BOH^{+}92, HTP^{+}92, JKH94, MBS91, dlRKR^{+}90, dlR91]. restriction [DMWJ90]. result [CB92, DCY^{+}94]. results [BRA^{+}94, BG91b, BCB93, BCYC94, EB92, GHS^{+}93, HPM^{+}94, HKS^{+}92, IBD92, IITG94, OMS^{+}94, PMKH91, SBWV90, WKD^{+}91, WMvdK94]. retained [AKC92, MC92, NRS93b, OTWH92]. retaining [Mac92]. retains [FF93, SST92a, XMW^{+}93]. retardation [WNTT94]. Retention [KD92, SRHK94, GK90, GJJM94, HH94a, LD92, LSK^{+}94, NRS93b, PRB94, SCM90, SP92, SPB90, SBL^{+}91, SM91c, WSM93, WLH92]. Reticulomyxa [SSVE91]. reticulum [AES94, BF90, BOM94, BHLWH91, BKR90, CG91, DCH^{+}92, FAF^{+}93, GK90, GPH^{+}90, GFW92, HR94a, JNP93, JSA^{+}93, KGR94, KBA92, KDO^{+}94, MBC90, MH92, MYT94, NSHN92, NDM^{+}94, PPZ^{+}93, PNPB94, PNM^{+}94, PCP^{+}91, PSSK93, PDC91, RLS94, ROBN^{+}92, SCM90, SSC^{+}94, SDE^{+}93, SPW93, STH^{+}93, SB91, SPB90, SBL^{+}91, TBFP92, TS91, TJ91, TBKF^{+}92, VGMS91, VM94, VPP^{+}93, WF93, WBH^{+}90, YL93, ZABM94, dSBH90]. reticulum-derived [RLS94]. retina [BTSL91, GPRB93, TKHM91, dCQTR91]. retinal [AMMW92, DP91, GFH90, GORB91, GPRB93, MJ92, NFL90, NPST94, PLS92, SF90a, UKH^{+}91, VGM^{+}93, aUFQ^{+}92]. retinoblastoma [DMJ^{+}94, LMB^{+}94b]. Retinoic [AGGS92, RELS^{+}91, SGW91a, WGTCY94, BG91b, BMGC91, KHS^{+}90, ML91, ZZR94]. retinoid [aUFQ^{+}92]. retinoid-binding [aUFQ^{+}92]. retinol [BMGC91, GFH90]. retinol-binding [GFH90]. retinopathy [PBCP91]. retinotectal [dlRKR^{+}90, dlR91]. retraction [JM92, JvCH^{+}94]. Retrieval [JNP93, TFP94, GtHG^{+}94]. retrieve [LVA^{+}90]. retrieved [Hol93]. retrieves [TLWA94]. Retrograde [KFN93, SRG^{+}94, TBFP92, Hol93, SSVE91, dBNI^{+}91]. retrogradely [HSYYK90]. Retroviral [GTH90, HR90a]. return [SGBN^{+}93]. Rev [RHB^{+}93]. reveal [CCW^{+}93, GZC^{+}91, OKMB91, dTF90]. revealed [AR93, BZT93, FFKC94, FHI90, GCML94, KISY91, KBP91, LZI^{+}93, NSH90, PM93, SK94a, TCCT91, WTK^{+}92]. reveals [BHFS93, HASW94, KBL91, KIO^{+}94, LCMP90, RHG90, SRV^{+}90, SHBD90, VPP^{+}93, WRA93, WSvdK^{+}93, WFD^{+}91]. reversal [GORB91, SW92]. reverses [LSP^{+}91, LSGD^{+}91]. Reversible [CDRL93, PMHS92]. reversibly [AG93, HFL94]. reverting [MPAF91]. RGD [BDB90, YED^{+}93, GAvdM91, MH91, MSQ^{+}90, SSM92b]. RGD-dependent [GAvdM91]. RGD-independent [MSQ^{+}90]. RGD-inhibitable [SSM92b]. RGDS [SADJGP94]. rhabdomeral [AHW90]. rhabdomyosarcoma [AGGS92]. RHAMM [HWLT94, SHS^{+}93]. rhesus [YDO^{+}90]. rho [KSK^{+}93, TSH^{+}93, JvCH^{+}94, PZB^{+}94]. Rho-type [PZB^{+}94]. Rho1 [YTN^{+}94]. Rho1p [MJG^{+}91]. Rhodobacter [TM90]. rhodopsin [DP91]. rhoptry [BDMPJ94]. ribbon [HASW94]. ribbons [SFA90]. ribonucleoprotein [KHCR91, SPvV94, KMBR90]. ribonucleoproteins [MW93, WMCG91]. ribophorin [SCM90, TIA^{+}92]. ribophorins [IDT^{+}92]. ribosomal [BHF90, FPGS91, MPTW90]. Ribosome [CG91, FPGS91, MPTW90, SM93a]. ribosomes [CG91, KGR94, YSK90]. ribosylation [EOO^{+}94, JvCH^{+}94, NHK^{+}94, TSH^{+}93, SBR91]. rich [BIPG91, CPB94, FABS90, HD91b, LSV^{+}94, MPR90, RWL^{+}90, SSTL93, SLB^{+}94, SBW90]. Ricin [SPHvD91]. riddle [Kos94]. rigid [MWSP91]. rigidity [GMNH93]. rigor [MSFW92, MGW92, SGM90b]. RIII [ZB94]. ring [CW90a, CW90b, OOBJ92, WKS^{+}93]. RMLC [COTC94, OCC94]. RNA [CTL91, DMJ^{+}94, FPGS91, HTSP94, HTD90, HDES94, ISDM92, JBIM94, KBL91, KZS94, LXM91, LSP^{+}94, LPA^{+}93, MCDY94, PZV^{+}91, RWL^{+}90, SZB^{+}94b, TLFS92, WSvdK^{+}93, WV93, WDP^{+}94, XL91, YMO92, ZKMB93]. RNA-binding [WDP^{+}94]. RNA1 [HTD90]. RNAs [CYCA90, ISDM92, KLST93]. RNase [MMF94]. RNase-sensitive [MMF94]. RNP [KBL91, KKBL92]. RNP-antigens [KKBL92]. Ro [SPvV94]. rod [AMMW92, HW92, HWM^{+}93, HCKJ92, LCF92, MKSF93, MNS90, SLK93]. rodent [MS94]. Role [CSKZ94, FAS^{+}93, FABS90, GBS90, HMS94, HCKJ92, HH94b, JJB^{+}93, JSM^{+}90, MMG90, OLKD90, PSHK92, RPS^{+}92, SKSD^{+}94, ATV^{+}93, AA90, BGMM^{+}93, BVS^{+}93, BCBK92, BVR^{+}93, BCET^{+}92, BDK92, BTR92, CM94a, CWBK90, CT91, DDW94, DCTG90, DCL^{+}92, DKG92, EMGS93, EC93, FAWM94, FKS93, FRSL90, GVK^{+}93, GFH90, GSS^{+}93, GIAS93, HNvdK^{+}94, HODF94, HM93, HW94, ISF^{+}92, JKLG94, KSL^{+}94, KFC^{+}94, KJT91, KDN94, KPMG91, KO94, KHHR92, LHH^{+}91, LRDM91, LDNM92, LYK^{+}94, LWHL93, MHR94, MPIDD93, NCBS90, NGP^{+}91, ORD90, ORS^{+}90b, PS94, PBD93, PNR^{+}93, PAL^{+}90, PDSB92, ROE^{+}90, RRD93, SS94, SMTW94, SDE^{+}93, SZB^{+}94b, SCJ^{+}91, SSJ90, SY91, STC93, STA^{+}94, TTE^{+}94, VWT^{+}91, VKM93, WIS^{+}94, WGD^{+}93, YGM^{+}94, ZS94, HTHC93]. Roles [WTS93, BGGW^{+}94, BHFS93, FMCR93, LOOM93, MNME^{+}93, MKSF93, NIT94, PH94a, PSJ^{+}91, PSG94, PM93, SNG^{+}91, WCF92]. rolling [ARW^{+}94, LKD^{+}94]. root [LS92]. roots [WGP^{+}93]. ROP [BDMPJ94]. rotamase [BZT94]. rotation [HE91b]. Rotational [VBA90]. rotavirus [PMA91, SPB90, YDO^{+}90]. rough [DCH^{+}92, KGR94, MBC90, SCM90]. rounding [JM92, JvCH^{+}94]. Rous [MGB90]. route [BHD^{+}94, LRAG92, PAL^{+}90, SMSH92]. routed [HGS^{+}94]. routing [MJM92, MCJ^{+}93, VFC^{+}91, WGTCY94]. rowers [LH93]. rp59 [MPTW90]. rRNA [RT92]. rubella [HWF92, HWF93]. ruffled [GAHH94, VKS^{+}90]. ruffles [Has93]. ruffling [FFB90, GSR^{+}93]. rules [ZCS^{+}91]. Russell [VGMS91]. Ryanodine [WAS^{+}91, JSA^{+}93, MTS^{+}91, RRG^{+}94, YAJ91].

S [TCDH93, YM90a, AKC92, CKD90, DZdH^{+}93, BRL^{+}93, CM94b, DS90c, DRJ^{+}91, GA91, GS94, HGDA93, LBH92, RPH^{+}92, TBFP92]. s-cyclophilin [AKC92]. S100 [SBWV90, TNS^{+}94]. S100-related [TNS^{+}94]. S180 [MMJ^{+}90]. SAC [CKG^{+}90, RGR90]. SAC1p [WCM^{+}93]. Saccharide [HWS90]. Saccharide-specific [HWS90]. Saccharomyces [JKLG94, AYM^{+}93, AJL^{+}90, AC92, BZT94, BSF90, BD92, BSD94, BGH93, CYCA90, CSV^{+}92, CF94, DKG92, GS90, GK93b, GSP^{+}94, GFW92, HLA^{+}90, HVK91, HYD93, HHLS92, JP90, JPS91, KCH^{+}94, KAG94, KDKF93, KHP91a, KCY92, KLR^{+}93, KDT93, LPPCF91, LB94, LSA91, MZRM93, MJG^{+}91, MT94, NHO^{+}91, PSHK92, PDC91, RRCR93, RHF91, SP92, SMB^{+}91, SGP91, SCF^{+}93, SH92, VMS^{+}91, VVB^{+}92, VRE^{+}92, WF91, WDP^{+}94, YTN^{+}94, ZHL93]. saccular [DAW^{+}94, OHC^{+}93]. Salmonella [MCDA^{+}93]. saltatory [AR91]. saltatory-like [AR91]. salts [WMB^{+}94]. same [HK91, HSUS90, KHHC93, LB94, OHP^{+}94]. sandwich [DTY92]. saponin [SGN91]. saponin-permeabilized [SGN91]. Sar1 [KDO^{+}94, ONN91]. Sar1p [ON94, dWLS91]. sarcolemma [JAS^{+}90, NEL^{+}91, NGB^{+}93, OESC91]. sarcolemmal [YAJ90]. sarcoma [MGB90]. sarcomeric [VB93]. sarcomeric-alpha-actinin [LDS^{+}92]. sarcoplasmic [BOM94, FAF^{+}93, JSA^{+}93, VPP^{+}93]. Satellite [WMvdK94, Bis90, MMY^{+}92, OHS92]. saturable [BSHW91]. scaffold [KRHY93, SGWE94]. scaffolding [HM93]. scaffolds [CYC91]. scale [ES91, THBW94]. scanning [GA92]. scapular [MSL92]. Scatter [RNG94, SMWB93, WBVB90, JCO^{+}94, RJJ^{+}94, UK92, WSB93]. scattered [WSvdK^{+}93]. scavenger [SOPA94]. Schizosaccharomyces [BHBG94, DMM93, HRH90, MSC92]. Schwann [KGSB92, MVM90, BS93a, CS90, CSCSA94, CSH^{+}92, DR91, DS90a, EMGS93, GFL92, MJM91, MM92, NS93, OBBS90, SRR90]. Schwannoma [KS92]. Schwannoma-derived [KS92]. sciatic [FFB^{+}93, SST92b]. ScII [SGWE94]. scinderin [DVT92, VDTT91]. Scrapie [BST^{+}90, TSP90]. scruin [OD93, SAJ^{+}94]. sea [AAM92, BMV90, Bro91b, FL90, FL92, HDDS94, JGLK90, KL90, LW92, MPVL92, MMM^{+}92, MNV93, MSFW92, MGW92, ODPL93, OPL94, SMCR90, SMTW94, TS91, TJ91, WCW91, WHW^{+}91, WTS93]. seal [WB93a]. sealing [LHH^{+}91]. seam [KIN^{+}94]. Sec12p [dWLS91]. Sec12p-dependent [dWLS91]. Sec13p [PSSK93]. Sec15 [BN91]. Sec15p [BMGN92]. sec18 [GE91]. Sec2 [NMPN90]. Sec4p [BMGN92]. Sec61p [HAG^{+}93, KGR94]. Sec61p-complex [KGR94]. Sec63p [BS93b]. Sec72p [FS94]. Sec8p [BMGN92]. Second [ZBPV92, GPH^{+}90, HSLF^{+}92, JM92, MPAF91, SWD92, SGN^{+}93, WBW^{+}94]. second-site [MPAF91]. secondary [AKK^{+}93, DJA^{+}90, GMQ^{+}93, MBG93]. seconds [TLWA94]. secretagogue [ASE90]. secretagogues [ATSE90]. secrete [GHH90, KMSW91]. secreted [BIPG91, Low92, PL90, SCMU93, SSSL94, SKD^{+}91, TLSW93, TRH^{+}92]. secretion [AGKC92, BPZ^{+}92, CC90a, CCG^{+}94, DHP93, DUVV93, EOL^{+}90, FFR^{+}91, FSSA92, FPPD91, GCZ^{+}92, HC94, HM92b, IBD92, MK91a, MM91, MOT90, PMHS92, RCNO94, SdAN^{+}91]. secretion-competent [MOT90]. secretogranin [CM94a]. secretory [AB90, AL91, BP91, BHLA90, CM94a, CH91, CB90, CC90b, FS94, GdLMS94, GK92b, HH94a, HBS92, HFS^{+}93, JGK^{+}94, KD92, KCY92, KR91, KCB^{+}92, KA92, KA94, LWV92, Mac92, MTC^{+}91, MJM92, MEM94, MOT90, NHK^{+}94, NF92, OHP^{+}94, PDC91, RNS92, SP94, SCV^{+}91, SFS90, TFTH91, TRH^{+}92, VGMS91, VMR90, WF91, XS93, YSK90]. section [MOKF92]. sectioned [WBR91]. sectioning [BZT93, MOKF92]. sections [Woo94]. SED5 [HP92, BLR^{+}94]. seed [TH92c]. seen [OSK^{+}92]. segment [CE94, JRK^{+}94, LCF92, PMG^{+}92, TDSP93, WPW92a, WB92c]. segmentation [DAIS90]. segmented [WGP^{+}93]. segments [GK90, KYLV90, NSC^{+}93, ORS90a]. segregation [AA90, BKB^{+}90, CSC92, GK93b, ROE^{+}90, SSS93, SSJ90, STH^{+}93, WKMG92]. select [SOPA94]. selected [QDT^{+}90]. selectin [ARW^{+}94, EWP^{+}93a, FWD^{+}92, KCB^{+}92, ZMS^{+}91, EWP^{+}92, ECAG93, GSMK94, KIS^{+}94, KSL^{+}94, LLIV94, LMIV93, LKD^{+}94, MSD^{+}92a, SST92a, STo^{+}92]. selection [BCG^{+}92, FMS93, VVB^{+}92]. Selective [DKL^{+}90, LTWH92, SP92, WBW^{+}94, BB93, CKB93, CH91, FS94, FHI90, FKH^{+}93, IBD92, LMH^{+}94, MMLS92, MCJ^{+}93, OOBJ92]. selectively [WFCN94, ZSC91]. selectivity [ESL^{+}90]. self [MR92, STL^{+}90, SY90a]. self-assembly [MR92, SY90a]. self-regulating [STL^{+}90]. semi [CM94a, GK92b, KRCT93, MM91, SHDS93]. semi-intact [CM94a, GK92b, MM91, SHDS93]. semi-synthetic [KRCT93]. seminiferous [PPHS^{+}92]. semipermeable [DLSK91]. Semliki [GHR^{+}90, MPK93, PK91, SWL^{+}92, WG92a]. senescent [RMFA92]. Sense [IITG94]. sensitive [BCLM94, BS94a, BU94, BVF^{+}93, CCR93, Dav92, GHR91, LVA^{+}90, MHM91, MSTY90, MIU^{+}92, NMP^{+}92, PA91, PC90, RK90a, RHR90, SSRM90, SOPA94, SHTL93, SVK^{+}92, TBFP92, THH^{+}90, WRHW92, WB93a, WRB^{+}94, WL91, GP91, MMF94, NF90, PRAK94]. sensitivities [SH91]. sensitivity [BKSB94, MLS94, PRB^{+}91]. sensitization [KKBK92]. sensory [AC93, ARWD93, BMW^{+}93, GMLD92, RBR90, SHR93, WMvdK94]. separable [CGW94, VRE^{+}92, YKM^{+}92]. Separate [HSUS90, VHSR92, CvdE91, DI91, FYDB93, GK90, GHS^{+}93, GFL92, MECE93, TVG^{+}93]. Separation [QXWN^{+}94, SM92a, WCR93]. Sequence [LPPCF91, PRF^{+}93, AA94, AKC92, AS94, BSSL^{+}91, BBD^{+}90, CLK^{+}91, FPR90, GHR^{+}90, GPH^{+}90, HW92, HD91b, HD90, HPYB93, IFLG90, JSM^{+}90, KDT^{+}91, MYM94, MMLS92, MNB92, MS94, MC91c, MC91d, MMY^{+}92, NRS^{+}93a, RBW93, RWG^{+}92, SADJGP94, SKSC94, SLS^{+}90, SJMK93, SK94b, SDPH90, TN94, TRV^{+}90, WK92, WBE91, WHW^{+}91, YED^{+}93, YPH^{+}92, ZBW93, dCQTR91]. sequence-mediated [AA94]. Sequences [KZS94, SPB90, CCW^{+}93, GCML94, HMS^{+}90, HGW^{+}91, HSP93, KLST93, Koz91, LXM91, MBZL91, NDdC92, NWvH94, OHS92, OKH92, RR92a, RHH92, Sch91, SS92b, TRD^{+}92, WPW92a]. sequencing [WBL^{+}91]. sequential [CHW94, DHPG94, HM92b, LKD^{+}94, PCD^{+}92]. sequentially [WHK90]. sequestered [MBC90, MTJM93]. sequesters [CSNZ92]. sequestration [PTS^{+}92, PTK^{+}93]. series [DHPG94]. serine [CJB^{+}93, FMLD92, NFYI93]. Serum [GSR^{+}93, HSG94, SWL90, VGRC^{+}92, BR94a, BCET^{+}92, GB92, LFC94, QCCD92, RLD^{+}91, RDBHG90, SBB^{+}91a]. Serum-derived [SWL90]. serum-free [BCET^{+}92, QCCD92, RLD^{+}91]. Serum-induced [GSR^{+}93, HSG94]. serve [HPW94]. set [PR91]. sets [RLS92]. seven [TDSP93]. several [HFL94, LSK^{+}94]. severin [ENS91]. severing [Bea91, KMN91, MWSP91, MZP91, SS94, WPW92b]. sexual [ZS94]. SF-assemblin [WGP^{+}93]. SF/HGF [CPP^{+}94]. Sf9 [BPK91]. shadowing [SGW^{+}91b]. shape [Ada92, CWHH92, GPF94, GMNH93, LS90, LCD^{+}93, MMJ^{+}90, RGT^{+}93, SHTL93]. shaped [KBL91]. share [LMB^{+}94b]. sheath [KHTD92]. shedding [FABS90, NXHMJ93]. sheep [FMCR93]. shell [SGN91]. shift [LBL^{+}91]. shifts [GDKP92]. Shiga [SPRvD91, SRG^{+}94]. Shiverer [ASDC91]. shock [BWV92, CLDV90, GVK^{+}93, HSCN94, IMK^{+}92, KAF^{+}93, LFR93, MVV^{+}91, PTTA90, VGG^{+}93, WWS^{+}94]. short [GAvdM91, SY90a, YYY^{+}91, ZZR94]. shortening [DA92]. show [AWH^{+}94, HSB^{+}94, MBS91, TGR^{+}92]. shows [BPDM90, FFR^{+}91, YYY^{+}91]. sialic [SPKV94, TSLR90]. sialoglycoprotein [LSMRB90, RLB^{+}91]. sialophorin [NXHMJ93, YNST93]. sialyl [FWD^{+}92]. sialylated [ZMS^{+}91]. sialyltransferase [BZF^{+}92, WLH92]. side [BFR^{+}92, ODK^{+}93]. sight [Col94]. Signal [ASK^{+}90, FA91, GtHG^{+}94, JH93, KDT^{+}91, SSTL93, AKK^{+}93, BSD^{+}92, BML^{+}94, Car91, CG93, CCC^{+}94a, CCC^{+}94b, DPS^{+}93, FS94, FHvZ^{+}94, FHGN91, GK90, GJJM94, GHR^{+}90, GPH^{+}90, GKKS90, HGW^{+}91, HFD91, HD91b, HAG^{+}93, HSG94, HD90, JSM^{+}90, KKG^{+}91, KTR90, KIV93, LKSR94, LD92, LPA^{+}93, MMF92, MMLS92, MMS93, MB90, MNB92, MC91c, MC91d, MC92, NF90, NWvH94, NRS93b, OKF^{+}94, QH94, RG92, RWL^{+}90, SK94b, SDR^{+}94, SMWB93, SW93b, SES94, SR92d, SM91c, VWS90, WGR^{+}92, WHN^{+}94, XMW^{+}93, YMO92, ZBW93]. signal-anchor [HGW^{+}91, HFD91, HAG^{+}93, MMLS92, NWvH94, XMW^{+}93]. Signal-mediated [FA91, GtHG^{+}94, NF90]. signal/anchor [HD90]. Signaling [DCM^{+}93, BS94b, CNZ^{+}92, CCG^{+}94, DB90, EB92, GZC^{+}91, HG94, HK94, KYBC94, LSRC93, LCPAM94, SBB^{+}91a, VCFM94, YGM^{+}94, ZCH90, ZS94, ZZP94, ZMMP90]. signalling [MCDA^{+}93]. signals [BSW^{+}92, BLSR92, FDT^{+}91, FM94b, GC93, IBR94, JK92, KLR^{+}93, MME93, OTX^{+}94, PRL^{+}94, SOKS91, TWMS90, WTH^{+}94, WSB93]. silver [CYCA90]. silver-binding [CYCA90]. simian [PRB^{+}91, aUFQ^{+}92]. similar [DBSK90, FWBS93, HTHC93, KRdlR^{+}92, KMSW91, MC94, OSK^{+}92, PMG^{+}92, Wat90]. similarities [BBD^{+}90, BMH^{+}93]. similarity [GMQ^{+}93]. simple [Lut91]. simplex [CHVF91, SWM^{+}92]. singed [CKMC94]. Single [BPTS94, LTYY94, WSC94, AO90, BBKR94, Blo93, BR91, BSR^{+}90, CH93, HTGN94, KII^{+}91, Low92, PAC^{+}92, RH92a, RA90, SAJ^{+}94, SGW^{+}91b, TLWA94, WB92c, dBNI^{+}91]. sinusoidal [GWMP91, SP94]. sinusoids [SF91]. sis [HXM^{+}94, LD91, LD92, XMW^{+}93]. sis/PDGF [LD91]. SIS1 [LSA91]. Sister [SM92a, GHK94, RSCS94]. Site [Low92, BMTTG94, BDB90, BSR^{+}90, DGAB^{+}93, DWJP93, FMS93, GSFS92, HW90, HWL^{+}90, HSB^{+}94, HMT^{+}92, KIS^{+}94, KHP91a, KGA^{+}93, KO94, Mac92, ML94, MWD94, MPAF91, MR92, NS93, OTT92, OCC94, PPJF91, RG92, SSA^{+}90, SB91, TWV94, VKR^{+}91, WFD^{+}91, WC93, YYY^{+}91, YTN^{+}94]. site-dependent [WFD^{+}91]. site-directed [BSR^{+}90]. Site-specific [Low92, HW90, HWL^{+}90]. sites [BKSB94, Blo92, BSD94, BK91, DIYSS92, DP92, EWP^{+}93a, FAYM90, HRC^{+}90, HMS^{+}90, HPWN93, IYN^{+}91, KM92d, LS90, LSK^{+}94, LCPAM94, LC94, MH93a, MMLG94, NYA^{+}91, PHSA93, PMLM94, RR92a, RKB^{+}90, RG94, RZS91, SLKS94, SW90, SBP90, SB93a, SOY93, ZCP^{+}94]. situ [BPTS94, DBV^{+}92, DMLR^{+}94, DFC^{+}94, FKH^{+}93, FWBS93, GSBS92, GCML94, HASW94, OB93, PRC^{+}90, PJC^{+}90, SOE^{+}91, SF90a, SZB94a, SAVK94, TLFS92, Woo94, XL91, YAJ90, YAJ91]. six [MR94]. size [DBSK90, HRH90, KK92, KRMK93, KWW91, TMI^{+}91, TCS^{+}94, WPS91]. sizes [dTF90]. skeletal [AWFP93, AHM94, BKW91, CG90, FSM^{+}93, FHI90, Jas93, JAS^{+}90, KB91a, KWW91, LWB91, LC91, LOC^{+}90, MB92, MFOI92, NEL^{+}91, OC91, OESC91, OWF^{+}93, PDWB92, RRG^{+}94, SBK^{+}91, SWF^{+}92, SBLV92, SAM^{+}91, TS93, YRBBP90, YAJ90, YAJ91]. skeleton [FABS90, HD91a, RMB90]. Skin [MSSD93, MLKB92, MP94, PA91, SSB93, TFC94]. Skin-specific [MSSD93]. skinned [LWK91]. Skn1p [RPPB94]. SL50 [JAS^{+}90, YAJ90]. SLA1 [HYD93]. SLA2 [HYD93]. sliding [Bro91b, SKT^{+}90, WDWT90]. slipping [BMS^{+}91]. slit [SAF90]. Slow [FMJ91, LPK92, BPDM90, CWB93, HLWL92]. slowly [WN92]. slows [PMD^{+}93]. SLT2 [MZRM93]. slug [EVW91]. Sm [HFL94]. Small [WMCG91, BTH^{+}90, BLR^{+}92, CYCA90, GWMC90, HKS^{+}92, HK91, HLWL92, HPP^{+}93b, JvCH^{+}94, KAV93, KKBL92, KAF^{+}93, LJP^{+}93, MSS^{+}94, MW93, MJG^{+}91, MPEG93, MBB^{+}93, NPR^{+}94, RLB^{+}91, RDDR93, RMG90, TSH^{+}93, WPS91, WHH^{+}93, YTN^{+}94, ZJA^{+}94, dWLS91]. SMC1 [SLK93]. Smooth [WDWT90, AP92, BBK90, BF90, BKW91, CYR^{+}94, DGGG93, DSQ^{+}93, DAIS90, FAS^{+}93, HAK93, HTBF94, KHR91, MMG90, ND91, NEL^{+}91, NGB^{+}93, STL^{+}90, SOA^{+}93, SHBD90, TWC94, VCL93, VPP^{+}93, WMS^{+}93, YYM90]. Smy1p [LB94]. SNAP [WRHW92]. snR10 [CYCA90]. snR11 [CYCA90]. snRNA [KM92a, vZDF^{+}93]. snRNAs [MG92]. snRNP [CFFL93, FZ90, FHvZ^{+}94, HFL94, JG92, KBL91, MG91a]. snRNP-containing [CFFL93]. snRNP-specific [FZ90]. snRNPs [ACFFL93, CFPCL92, FCFL94, FDT^{+}91, MMD93a]. snurposomes [WMCG91]. Sodium [LC94, DPS^{+}93, DFC^{+}94, FM94a, GFWM92, LC91, SVK^{+}92]. solation [JKT91, KJT91]. solation-contraction [KJT91]. solely [WDT^{+}92]. solid [SAWL90]. solubility [PC91b]. solubilization [KMN91]. solubilized [BHFS93, HNR90]. Soluble [SST92a, AMG90, BN91, DZdH^{+}93, DGGR92, FKHG93, GFW92, LVK^{+}91, MEM94, MKF91, PL90, RCJ^{+}94, RJJ^{+}94, WRHW92]. solute [KAV93]. solution [SPD^{+}90, VAL^{+}93]. solution-phase [SPD^{+}90]. somatic [DMM^{+}91, FHvZ^{+}94, PJS^{+}92, RSCS94, VVR^{+}90, WE90]. some [LBF92, RA94, TH92b]. Somite [SKM94]. somites [CLS^{+}92]. sort [KLST93]. sorted [FYDB93, HLRBE93, SMSH92, VHE93, WHF^{+}90]. Sorting [CC90b, KBH91, KP90, LLG^{+}94, MPM93, MCF91, OTX^{+}94, SFS90, AKK^{+}93, CM94a, DP91, DM92, EHB^{+}90, FHGN91, GE91, GSMK94, vtHvM90, HH94b, JK92, JKS93, KA92, KA94, MP94, MYM94, MMWG94, MC93, MHGJ94, MAAO92, PYH^{+}90, QORM91, ROE^{+}90, RNS92, RCNO94, SKSD^{+}94, SB93c, VRE^{+}92, WB92c]. source [JRK^{+}94, LIAJS94, LSV^{+}94]. sources [AB91]. soybean [MHV92b]. sp56 [CLP^{+}94]. SPA2 [GS90]. space [CVH91, YYR92]. spacer [KDKF93]. spacing [MH91, NPST94]. span [KAG94]. spanning [CH92, LGSB93, NDP^{+}91, SB93c, SM91c, ZCS^{+}91]. spans [ROBN^{+}92]. SPARC [LS90, LIAJS94, TLSW93]. Spatial [ASL^{+}94, EBS91, MRSN93, MSLS91, SBW90, VCFM94, CMZ90, FBLL90, HAS90, KSE^{+}92]. spatially [KHS^{+}90, MBS91, OHS92, YOWM91]. Spatiotemporal [BKB^{+}90]. special [LLG^{+}94]. specialized [BHC^{+}94, VPP^{+}93, WFD^{+}91]. species [HFL94]. Specific [LVBK92, TOA^{+}91, ATD^{+}90, AVP92, BG91b, BKR90, BBS92, BS94c, BCYC94, CLP^{+}94, CSG92, CMM94, CYCA90, CSR^{+}92, CPM^{+}91, CKL^{+}91, DAM^{+}93, DS90b, DEH^{+}91, DFC^{+}94, DJF93, EG91, FFKC94, FHUY93, GSBS92, GCML94, HRT^{+}91, HDDS94, HSP^{+}91, Has93, HE91a, HW90, HWvF90, HWL^{+}90, HSW90, HRB92, IYN^{+}91, JBIM94, JSM^{+}90, JBJH^{+}92, KRHY93, KS90, KLST93, KGK90, KGA^{+}93, KWR94, KJ90, LRR^{+}92, LBH92, LRSB94, Low92, LE92, LJP^{+}93, LS92, Mac92, MGG93, MSSD93, MSD^{+}92a, MDKB94, NDdC92, OHS92, PZV^{+}91, PPHS^{+}92, PYH^{+}90, PTTA90, PAC^{+}92, PAL^{+}90, RELS^{+}91, Rot90, RLKB91, SOA^{+}93, SH90b, SS91a, SGW^{+}91b, SGBN^{+}93, SW93a, SSFD93, SBB91b, SSB^{+}91, STA^{+}94, TN94, VVR^{+}90, VHB93, WFD92, WHT^{+}93, WG92b, WGTCY94, XL91, YRBBP90, YCBM94, ZZR94, dCV90, FZ90, GE91, HDC^{+}91, HWS90]. specific [JG92, RB92]. specifically [BMK^{+}90, BLFQ93, DBWS94, HGDA93, KC90, KPAY91, LXM91, OOM90, PHMH92, SG92, SLL94, SS92b, vZDF^{+}93]. specification [SBC^{+}93]. specificities [SSVE91]. Specificity [TND^{+}93, BBK90, DP92, MMJ^{+}90, SFO^{+}91, VLH^{+}93]. specify [HWM^{+}93]. specifying [JLC93]. spectral [PMKH91]. Spectrin [FA92, BBMN94, DGF^{+}94, DBSK90, GRFB94, KWFM91, LCD^{+}93, PDWB92, STZG91]. spectroscopic [OOBJ92]. Sperm [CLP^{+}94, BMPW90, BMV90, Bro91b, DDW94, FAWM94, FL90, FL92, HHF93, KMSW91, LA92, LCMP90, LML^{+}94, MNV93, MSFW92, MGW92, ODPL93, OPL94, PKPM90, VSVP93a, YHM^{+}94]. Sperm-egg [CLP^{+}94, BMPW90, LCMP90, YHM^{+}94]. spermatozoa [BBB^{+}94, Bro91b, OHC^{+}93]. spermatozoon [NCM^{+}93, PNB91]. sphere [TSR93]. spheres [WMCG91]. sphingolipid [AYM^{+}93, vtHvM90, PMKH91]. sphingolipids [KBH91]. sphingomyelin [EH94]. Sphingosine [ZDO^{+}91]. Sphingosine- [ZDO^{+}91]. sphingosylphosphorylcholine [DCM^{+}93]. sphingosylphosphorylcholine-mediated [DCM^{+}93]. sphinx [Kos94]. spicules [KL90]. spike [HWF93, PP91, SWL^{+}92, WG92a]. spinal [GMLD92]. spindle [AR91, BR94b, BGH93, CMM94, CM94b, HRH90, HBR90, HAS90, HSSC92, HHLS92, KDKF93, MECE93, MMDM93, MHYC90, MSC92, MT94, MM93a, OMRM93, OSJS94, PC91a, PSHK92, PH94b, RK94, RA90, RSCS94, RS94, RMR92, RK90b, SM91a, SCF^{+}93, SSS^{+}90, SSB93, SHOA92, TH92a, TZBV91, VKM93, WCR93, WGBB91, WHC^{+}93, SSS^{+}90]. spindles [DMM93, HSSC92, MCG92, SM91b]. spinous [DY93]. spiralis [Jas93]. spliced [CE94, GTH90, KEBD91, MULA^{+}91, OKMB91, PJC^{+}90, SMRG91, SSB^{+}91, WS92]. spliceosomal [HFL94]. splicing [ACFFL93, BNI^{+}94, DBC90, FCFL94, OMS^{+}94, PZV^{+}91, SJMK93, SES94]. Spoke [PC91a, DAR93]. spokes [GOP^{+}94]. sponge [PMW92]. Spontaneous [DLMS91]. spontaneously [HLPL90, MH94]. spore [WE92]. sporozoites [PHMH92, STM91]. Spreading [Sch93, AC93, AG93, CSCSA94, CNSC90, CMS^{+}90a, CM93, ELZ93, LTYY94, LFWC92, LKD^{+}94, MH91, NXHMJ93, RZS^{+}94, TWC93, WC93, YCO^{+}93]. spring [GYE^{+}90]. sprouting [CG90, CSKZ94]. squamocellular [SGVS^{+}93]. squamous [SSB93]. squid [SRB92]. SR13 [SSWS92]. src [DPND90, Eri93, HR90a, HAA93, HNC^{+}92, HR90c, KSVM92, KDT^{+}91, LVBK92, MHT^{+}92, MAWM90, SR92a, SHB90, TISH94, TOA^{+}91, WTRD92, BVF^{+}93, BIPG91, SSDK^{+}94]. SRC-induced [BIPG91]. Srp1p [BKWD94]. SSB [CYCA90]. SSB-1 [CYCA90]. stabilities [CMZ90]. stability [BB90, BSW91, CLDV90, HDES94, HM90, LBH92, LLDJ94, SH91]. stabilization [CCFZ92, PBCK93, RB90]. stabilize [WWWB90]. stabilized [PBDK90]. stabilizes [CML90, WLWL94]. stabilizing [HNP94, STo^{+}92].

Stable [BGGW^{+}94, BA92, CKG^{+}90, FGH92, KES^{+}92, KB91a, WSL91]. stably [BSM90, HMC91]. stack [JK93, PNPB94, SPW93, WCR92]. stacks [MW94, SDE^{+}93, VYM93]. Stage [SS91a, CPM^{+}91, DJF93, FWF90, SSB^{+}91, WFD92, vdBRD^{+}93]. stage- [DJF93]. Stage-specific [SS91a, CPM^{+}91, SSB^{+}91, WFD92]. stages [BCG^{+}92, CC90a, CHW94, CF90, HM92b, MTC^{+}91, SABF^{+}90]. stain [DAMS91]. staining [OOBJ92, SGW^{+}91b]. stains [RMG90]. stalk [dCTG92]. stamen [ZCH90]. starfish [JGL^{+}93, PLB^{+}91]. starting [QDT^{+}90]. starvation [LFC94]. state [CSvdE93, CW90c, KTG90, PJS^{+}92, SH90a, SHBD90, TSS91, VLGB92, WBAP91]. statin [CW90c]. stationary [CM93]. status [FAU^{+}93]. Staurosporine [Wal94]. STE6 [KDT93]. Steady [VLGB92, CSvdE93]. stem [AW92, CTQ91, DS90c, FAU^{+}93, MLS94]. step [KLERG94, PKPM90, PTMB94, WHLW90]. steps [CPB^{+}93, DWY^{+}93, WGBB91]. Stepwise [CC93a]. steroidogenesis [SSKT^{+}92]. still [FGS91]. stimulate [BLSR92, LIAJS94, MW92, PCG^{+}94]. stimulated [BPZ^{+}92, CMZ90, CPA^{+}93, EOL^{+}90, GLAB93, LPM^{+}91, MGB90, MMR91, SY90b, WCW91, WDT^{+}92, LG93a, TLWA94, WWD94]. stimulates [BDZ^{+}92, GHR91, HR90a, HC94, HLC^{+}93, JGLK90, KBT92, KCYH92, MBS90, MIS93, PWY^{+}92, SGW91a, YKI^{+}91, vKWH^{+}91]. stimulating [LKXS93]. Stimulation [MDKB94, PA91, CLZ91, CSS^{+}94, DVT92, HAK93, MOT90, PIRB92, SHS^{+}93, SCS92, TS93, VDTT91]. stimulator [LK93b]. stochastic [LH93]. stoichiometry [HKSL90]. stomatitis [BF90, LTPB92, dSBH90, dSBH93]. Stopped [KPAY91]. Stopped-flow [KPAY91]. storage [AKC92, GdLMS94, HBS92, LRAG92, SMCA91]. stored [CRS94, ZLG^{+}90]. stores [EP94, VPC^{+}91]. strain [WMS^{+}93]. strategy [WG92b]. stratified [HN90]. streaming [JT93, vDS94]. strength [CVH91, DSQ^{+}93, Fuc94]. strengthening [CLD^{+}91]. streptolysin [PIS94]. Stress [HG94, BLW92, BMH^{+}93, CSG92, DH93, EKK94, HSCR94, IITG94, KAF^{+}93, KJT91, MH91, NSHN92]. stressed [CETK^{+}94]. stretch [VSVP93b]. stretch-activated [VSVP93b]. stretched [DE90]. Striated [LM91, SLL^{+}90, WGP^{+}93]. string [BFM93]. stromal [EFTJ^{+}93, KCR^{+}94, MMI^{+}91, MHYK92, PC91b]. Stromelysin [MM92, LWL^{+}92, MSC91, STA^{+}94]. stromelysin-1 [STA^{+}94]. stromelysin-3 [LWL^{+}92]. strongly [SHB90, VGG^{+}93]. Structural [BFR^{+}92, HFAGM94, KISY91, MYM94, RCJ^{+}94, SS92a, AKK^{+}93, BMS^{+}91, DB91, FMCR93, FFSW91, KKBL92, KWGS92, LKF^{+}94, MSFW92, WGP^{+}93]. Structurally [KHHC93, SNGG94, CPM^{+}91]. Structure [BMWT93, BGM^{+}91, FDS90, GMB^{+}91, KKBL92, MPR90, MAFR93, MAB^{+}90, PSZ^{+}93, SGW^{+}91b, VHW^{+}92, ZAH^{+}94, ATBD90, BB94, BFM93, BDJ94, CBRW92, CFFL93, CKB93, CSC92, DWY^{+}93, ENS91, ECM^{+}92, FWBS93, GMQ^{+}93, GVK94, HD91a, HPH^{+}91, HKSL90, JHGR93, JGO^{+}91, KSE^{+}92, KRdlR^{+}92, KHW93, KDLS92, KBP91, KGM92, KDN94, LCMP90, LJJC93, LWK91, LCKW91, LGI^{+}93, MNS90, NDP^{+}91, PC91a, PME93, PC91c, RHB^{+}90, SGWE94, SK94a, SPHvD91, SAJ^{+}94, SOH^{+}90, SS90a, SVK^{+}92, TRS^{+}90, TCCT91, TU90, TH92c, VAL^{+}93, VNS^{+}94, WG92a, WB94, WBL^{+}91, WC90a, WBR91, Wu93, YYY^{+}91, YDO^{+}90, SEH^{+}93]. Structure/ [SEH^{+}93]. structured [MCDY94]. Structures [SLPB93, BA92, BWLK93, BVR^{+}93, CNJ^{+}94, COLL^{+}90, DSJB^{+}94, ECO93, HWW94, KMBR90, KHCR91, MY92, MWBD90]. STU1 [PH94b]. studied [CFG^{+}94]. Studies [Dun90a, Dun90b, SGP91, BPZ^{+}92, BI90, BDH^{+}94, FZRH91, GHGP^{+}92, INM^{+}93, JD91, NTB90]. Study [DUVV93, AL92, CHBGL91, FMR^{+}92, GB92, KW92b, MMM91, PRC^{+}90, RSS91, SS90a, SBR91, TM90, VBS^{+}93]. studying [CHG94, PMKH91]. Subcellular [LSP^{+}94, NT90, WHW^{+}91, vGvMS^{+}91, BKW91, FJRAT91, GH91a, HG93a, HWW^{+}93, HRB92, KWK92, PH94a, QXWN^{+}94, Ral93, SMJ^{+}94, TWQ^{+}93, TSBN94, VHM^{+}93, VHB93, WBA92, WA93]. subclass [BKSB94]. subcompartment [SRV^{+}90, WF93]. subcompartments [VPP^{+}93]. subdomain [PSGE92]. subdomains [CCR93, SKM94, WSSM91]. subepithelial [MCDA^{+}93]. subfractions [JKB^{+}92]. subfragment [SGM90b, WBR91]. subfragment-1 [WBR91]. subgroup [VHW^{+}92]. subgroups [GMB^{+}91, SR92b]. subjected [BLW92]. submembrane [Fro91, GORB91]. Submembranous [TIN^{+}93, SMCA91]. subnuclear [SCE91]. subplasmalemmal [EP94]. subpopulation [BB93, DS93]. subpopulations [FWF90]. subsarcolemmal [PDWB92]. subsequent [HPM^{+}94, MPM93]. subset [BOH^{+}92]. subsets [OB93, SLL^{+}90]. substituted [GJJM94]. substitution [vGvMS^{+}91]. substrata [LBF92]. substrate [ASR91, BBK90, BLR^{+}92, BZ90, CWBK90, DJA^{+}90, EOL^{+}90, LVK^{+}91, SM92b, TBVS92, VSB^{+}94, WP93, ZDP92]. substrate-bound [LVK^{+}91]. substrates [HLC^{+}93]. Substratum [SK91, CL91a, DIYSS92, MJ92, SWD94, SKD^{+}91, SB90]. substratum-adsorbed [MJ92]. Substratum-growth [SK91]. substructure [OOBJ92, XL91]. substructures [DE93]. subsynaptic [YLDB91]. Subtilisin [SKT^{+}90]. Subtractive [BLFQ93]. Subunit [ESL^{+}90, ASMC^{+}92, BBK^{+}93, BVS^{+}93, BM90, CPB^{+}93, DGL^{+}90, DG90, FHGN91, GSS^{+}92, GWMC90, GAS^{+}90a, GPH^{+}90, GSM93, HSCR94, HHR^{+}90, HAN^{+}93, ILH^{+}90, JGL^{+}93, JJB^{+}93, vdWvKvKdB^{+}92, KB91a, KHKO^{+}93, LTYY94, LFWC92, LB92a, MBW^{+}94, MH93a, MC91a, MC93, MSFW92, MGW92, NHH^{+}91, OPB90, PRF^{+}93, PMPV92, PDA93, RG92, RGHC91, SADJGP94, ST94, SKSC94, SM93b, SEH^{+}93, SLWF91, SSA^{+}90, TMP^{+}91, TYM^{+}92, TSBN94, WKY^{+}93, WBS91b, YCO^{+}93, YL93, ZMB91, ZBW93, dCQTR91]. subunits [BHF90, BM91a, CH92, DBM94, DFC^{+}94, DMB92b, EK90, FPP91, FGH92, GHS^{+}93, HKKO^{+}94, HWW^{+}93, JJB^{+}93, KYLV90, MBZL91, SWS^{+}92, SEAB91, TK94, WL91, WCC^{+}90]. subviral [DWY^{+}93]. successive [PCP^{+}91]. sucrase [FHGN91]. sucrase-isomaltase [FHGN91]. sufficient [CXS^{+}94, DCBH92, MH91, PSGE92, QCCD92, RMFA92, SB93c, WLH92, WB92c, YM93]. suggest [Eri93, PSJ^{+}91]. suggesting [SCJ^{+}91]. suggests [EMGS93, GFH90, ORS^{+}90b, PMPV92, PS94]. sulfate [BSEF^{+}93, BBJ^{+}91, BGRW91, CS90, CES^{+}92, DLD^{+}90, DBV^{+}92, DvdSM^{+}92, DKM^{+}92, FSK^{+}92, FCL^{+}94, FMR90, FMK^{+}94, GFM93, HAK93, ISF^{+}92, KT92, MFS^{+}94, NHK^{+}94, OR92, SR90a, SWM^{+}92, SF91, SFSG91, SDH90, SLL94, SdAN^{+}91, dAWS^{+}90]. sulfate-carrying [NHK^{+}94]. sulfated [PHMH92]. sulfoglucuronosyl [KYA^{+}94]. sulfonylurea [MDKB94]. sungorus [KCDR91]. SUP [PKG^{+}94]. supercoiled [BB94]. Supercontracted [SHBD90]. superfamily [BGM^{+}91, SR92b, VHW^{+}92, ZFA^{+}93]. superfluous [Eri93]. supernatants [PSE^{+}90]. superoxide [PSE^{+}90, SAVK94]. support [ASR^{+}90, IK94, QCCD92, ZMV^{+}93]. supporting [PAL^{+}90]. suppress [BMM^{+}94]. suppresses [RRD93]. Suppression [BRD94, FGS^{+}92, FGSBZ93, FNV^{+}92, OMRM93, WSL91, BG91a, CSKZ94, CSG92, HSB^{+}94, RBD92]. suppressor [JG94, PH94b, PCG^{+}94, SRM94, SJM^{+}94, TIN^{+}93, VKRD94]. suppressors [PPD92, SK94b]. supramolecular [DWY^{+}93]. Surface [AB91, KQJ91, SVD90, TH92c, ARW^{+}94, ATD^{+}90, AVP92, BMTTG94, BZF^{+}92, BS90, BHS91, BL94, BZB^{+}93, BFTR93, BSP^{+}94, CE94, CdJC93, CKD90, CMK^{+}90, CC90b, DKM^{+}92, ELS93, FSK^{+}92, GHS^{+}93, GAHH94, GG90, GSR^{+}93, GAvdM91, GHT92, GIAS93, GAS90b, GK92a, HJI^{+}93, HLRBE93, HS92, HDC^{+}91, HAK93, HEF92, ISF^{+}92, JKB^{+}92, KLM^{+}91, KTR90, LHH^{+}91, LCMP90, MPM93, MCM92b, MTJM93, NHA^{+}92, PLS92, PKPM90, RPH^{+}92, RR92b, SLP91, SH90b, SWM^{+}92, TM92, THBW94, TGD93, TOS^{+}94, VM94, VKRD94, VHW^{+}92, WLPB92, WFD92, WSC94, WOC91, WKN92, YHM^{+}94, ZCP^{+}94, ZLC^{+}93, dBNI^{+}91, dlRKR^{+}90, dlR91, vKWHF94, WS91a]. surface-connected [GHS^{+}93, MTJM93]. surfaces [BMM90, HWD94, HH90, LZBH92, LSC^{+}93, PTMB94, RBR90, WS91a, WS91b, ZBH94]. surrounding [BDMPJ94, LML^{+}94]. survival [ARWD93, BG93, IVBR93, KHTD92, KS92]. survival/death [BG93]. survive [IBR94, THH^{+}90]. susceptible [PYA90]. SV [HKS^{+}92, JAB^{+}91]. SV-40 [HKS^{+}92, JAB^{+}91]. SV2 [FYDB93]. SV40 [PRB^{+}91]. swapping [MCF91]. swimming [Bro91b]. Swiss [DCM^{+}93, GHR91, KATS90, ZBPV92]. switch [WRA93]. switches [MS94]. symmetry [KHCR91]. sympathetic [BG91a, BMW^{+}93, DJ93a, ET94, HR90a, LHT^{+}91, MWG92, Smi94, TWMS90]. synapses [MVTJ94, WBS91b]. Synapsin [STZG91, BVR^{+}93, TTVV^{+}90]. Synaptic [BCKS92, CQYD94, MMD^{+}93b, Ang91, AHM94, BVR^{+}93, BH94, BHC^{+}94, CSJD91, COLL^{+}90, CC90b, FYDB93, HMC91, LVBK92, LC94, MTC^{+}91, MTP^{+}92, MLR^{+}92, STZG91, TB90]. synaptogenesis [VCFM94]. Synaptophysin [TB90, CSJD91, FYDB93, LLG^{+}94, TTVV^{+}90]. synaptotagmin [FYDB93]. synchronized [PA91]. Synchrony [ELZ93]. syncolin [FFW91]. syncytial [MCG92, vDS94]. syndecan [EVL^{+}91, YR93, CSCSA94, CES^{+}92]. syndecan-1 [YR93, CSCSA94]. syndrome [KS94, MMS93]. synergistically [NS90]. synthase [EH94, VMS^{+}91]. synthases [SMB^{+}91]. syntheses [CW90c]. Synthesis [NFL90, SHR93, STS91, SST92b, AL92, AGGS92, BML^{+}90, BRF93, BS91, BST^{+}90, BCYC94, DUVV93, GC90, vtHvM90, HLR92, KMF^{+}91, LHS^{+}93, MIS93, MMW^{+}92, MMCH91, RDDR93, RRD93, RMFA92, SSCD90, SMCR90, SFSG91, SS90b, VGMS91, VBvG^{+}94, YMO92]. synthesized [BM91a, IDT^{+}92, JKB^{+}92, LGH91, MH92, VM93, VM94, WCC^{+}90]. synthetase [DKG92]. Synthetic [HYD93, SLWF91, BI90, CMS^{+}90a, DKM^{+}92, GSFS92, HW92, KRCT93, MR92, WW94a]. Synthetic-lethal [HYD93]. system [AST^{+}92, APP^{+}93, BHD^{+}94, BBS92, CSH^{+}92, DMB93, GHS^{+}93, GL91, GSV93, GMB^{+}91, HSHA90, KHS^{+}90, LCC^{+}93, LPM^{+}91, MBW^{+}94, MKCE92, MMC91, MMT91, MC91b, MW94, MDRG92, MPK^{+}94, NS93, OW90, PLS92, PAL^{+}90, STL^{+}90, SSWS92, SR92b, SB91, SLL94, TM90, WNTT94, Wat90, WS93a, YDS92, dlRKR^{+}90, dlR91]. systematic [PM93].

T [ARW^{+}94, AFA^{+}94, CSMdPSM94, DMB92b, FGK^{+}90, GRFB94, HTHC93, HG92, HKS^{+}92, IRH^{+}91, JAB^{+}91, vdWvKvKdB^{+}92, LMB^{+}94a, PIRB92, RM94, SNG^{+}91, SWFB91, SBL^{+}91, VR92, WGM^{+}93, WKD^{+}91, WCC^{+}90, WKY^{+}93, aUFQ^{+}92, vKWHF94]. T-cadherin [VR92]. T-cell [SBL^{+}91]. T-plastin [AFA^{+}94]. T124 [JAB^{+}91]. TA3 [KWW^{+}94]. TA3/Ha [KWW^{+}94]. tagged [CCW^{+}93, LZI^{+}93, MWH^{+}91]. Tail [SGBN^{+}93, BMF^{+}91, BI90, DP92, JT93, JK92, LSP^{+}91, MKSF93, NI93, NCM^{+}93, ORS90a, OTX^{+}94, PNB91, RKB^{+}90, SRP90, SLK93, TDT92, WF90, WCF92]. tail-less [BMF^{+}91]. Tail-specific [SGBN^{+}93]. tailless [MGG93]. tailpiece [HCKJ92]. tails [SRP90]. take [WMB^{+}94]. takes [HWvF90]. talin [DI91, GWO^{+}93, NTB90, TS93]. tandem [JKH94]. tangles [Kos94]. target [BSD94, GMGPM94, GSWW93, KYA^{+}94, LF93, PYA90, RHB^{+}93, SHR93]. Targeted [AW92, COTC94, CCW^{+}92, RLRR93, SSFD93, STA^{+}94, BO91, GI93, HXM^{+}94, HWF93, NLFRB91, WDRM93, ZLC^{+}93]. Targeting [CPP^{+}94, HFS^{+}93, Hur90, SBR93, SCV^{+}91, AL91, BLR^{+}94, BSD^{+}92, CB92, FPP91, GK93a, GKKS90, HDDS94, HSP^{+}91, HGW^{+}91, HRB92, JO92, KKG^{+}91, KLR^{+}93, KA92, LSMRB90, LTWH92, MMF92, MMS93, MG92, PMPV92, PMM91, PTK^{+}93, RS91, Rob93, SOA^{+}93, SR92a, SDR^{+}94, SW93b, SHM94, TMI^{+}91, VWS90, VHB93, WGR^{+}92, WTH^{+}94, WBT^{+}94, YPH^{+}92, YSK90, ZDP92]. targets [KKG^{+}91, SES94, VL91]. tat [BDB90, LCLG94, VLH^{+}93]. tau [BPK91, GF94, KKL^{+}91, WLZB93, WDB^{+}92, BK91]. Taxol [VBAK91]. Taxol-induced [VBAK91]. TC4 [KDN94, MPEG93, RDDR93]. technique [SSM^{+}92a]. teleost [KM94]. telomeres [FHUY93]. telophase [SPW93]. Temperature [MSTY90, PC90, BCLM94, BVF^{+}93, Dav92, JK93, RHR90, WB93a]. Temperature-sensitive [MSTY90, PC90, BCLM94, BVF^{+}93, Dav92, RHR90, WB93a]. Temporal [DJ93a, HAS90, SEMH92, WOC90, ASL^{+}94, DFLK94, EBS91, MRSN93, SBW90, SGP91]. temporally [KHS^{+}90, MBS91, OHS92]. Tenascin [BTG^{+}93, HFS92, KBM93, EFTJ^{+}93, Eri93, FWBS93, LVK^{+}91, MULA^{+}91, WBE91, ZDR^{+}92, CETK^{+}94, CE94, MSI^{+}94]. Tenascin-C [CETK^{+}94]. Tenascin-X [BTG^{+}93]. tensin [LJHC94]. tension [NW94]. teratocarcinoma [BG91b, KHS^{+}90, KBCM^{+}90]. term [KKBK92, PRB94, TWMS90]. Terminal [DS90c, SY90a, AWFP93, Ang91, AJPB90, CSKZ94, DMJ^{+}94, HW92, HMS^{+}90, HW94, JJS92, MYT94, MSQ^{+}90, MME93, NMPN90, OWF^{+}93, OKH92, SLWF91, SW93b, SMM^{+}91, TG94, VKRD94, VFC^{+}91, WTH^{+}94, AKC92, BSD^{+}92, CAS^{+}92, CCC^{+}94a, CCC^{+}94b, CCW^{+}93, HCKJ92, HD90, JJ92, JBJH^{+}92, KPMG91, LGI^{+}93, MMS93, NNHB^{+}92, NWvH94, SR92a, SRP90, SSM^{+}92a, TWV94, VSB^{+}94, WW94a, WGD^{+}93]. terminally [RPE^{+}90, TIA^{+}92, WBAP91, ZSC91, SC90b]. terminals [BH94]. termini [SY92]. terminus [BSP^{+}94, GK90, GJJM94, GSFS92, LPPCF91, NPST94, PTS^{+}92, PTK^{+}93, SG92, VJB^{+}94, VHB93, YKM^{+}92, Low92]. tertiary [SGWE94]. test [LELB90]. Tetanus [GCM^{+}94]. Tetrameric [VBS^{+}93]. tetranectin [WIS^{+}94]. tetrapeptide [GEKL^{+}94]. tetraurelia [SSRM90]. TGF [VBvG^{+}94, BPM93, FAS^{+}93, HGDA93, LCPAM94, LCLG94, MPIDD93, MELD94, PSJ^{+}91, RRD93, SCO^{+}91, SHS^{+}93, STL^{+}90, SOA^{+}93, SRK^{+}94, SSK^{+}93, TMHKO94, WTK^{+}92]. TGF-alpha [BPM93, SRK^{+}94]. TGF-beta [VBvG^{+}94, FAS^{+}93, LCPAM94, MPIDD93, SCO^{+}91, STL^{+}90, SOA^{+}93, SSK^{+}93, TMHKO94, HGDA93, MELD94, SHS^{+}93, WTK^{+}92]. TGN [HPP^{+}93b, RB92]. TGN-specific [RB92]. TGN38 [HPYB93, JCSH93, PRL^{+}94, RB92]. TGN38/41 [JCSH93]. Theileria [STM91]. their [APP^{+}93, BA93, BTBO94, BBB^{+}94, BST^{+}90, CTL91, CK94, CMZ90, CW90c, FFB^{+}93, GMGPM94, GAS90b, HGW^{+}91, HAG^{+}93, JKLG94, LSP^{+}93, MBC^{+}91, SWD94, SB93a, SLWF91, TBW92, TH92b]. therapy [RB94]. thereby [WB93a]. thermal [GMNH93]. thermodynamically [KB91a]. thermogenic [BOM94]. thermoresistance [IITG94]. these [SJGG94]. Thick [DE93, ECO93, LWK91, LCKW91]. Thin [MOKF92, FSM^{+}93, FHI90, GW94a, KWW91, VCL93]. Thin-section [MOKF92]. thiol [PAL^{+}90]. third [TK94]. thoracic [HE91a]. those [OSK^{+}92]. thread [CHVF91]. Three [DMM93, DEH^{+}91, FPGS91, Lut91, PRSS90, SAJ^{+}94, TU90, VCL93, VAL^{+}93, WBR91, YDO^{+}90, AR93, FW93, FWD^{+}92, GK90, GKG90, GFW92, HASW94, KDS^{+}91, LJHC94, LFF91, MG92, MNS90, ME91, PBM^{+}94, RRG^{+}94, SRHK94, SKSD^{+}94, TASJ92, WBL^{+}91]. Three-dimensional [DMM93, FPGS91, Lut91, SAJ^{+}94, TU90, VCL93, VAL^{+}93, WBR91, YDO^{+}90, AR93, FW93, HASW94, KDS^{+}91, LFF91, MNS90, PBM^{+}94, RRG^{+}94]. three-domain [WBL^{+}91]. Thrombin [BSEF^{+}93, EGP^{+}91, HVNC92, JM92, LASB^{+}93, STo^{+}92, BSSL^{+}91, BB93, BPZ^{+}92, JvCH^{+}94, LPM^{+}91]. thrombin-activated [BB93]. Thrombin-induced [HVNC92, STo^{+}92, JvCH^{+}94, LASB^{+}93]. thrombin-stimulated [BPZ^{+}92]. Thrombospondin [MBS90, SCMU93, LDW^{+}93, NT94, ORD90, OLKD90, PPJF91, SGN^{+}93, SSM92b, TRLG90, TVG^{+}93, TRD^{+}92, TRP^{+}93]. thrombospondin-1 [TVG^{+}93]. thrombospondin-4 [LDW^{+}93]. throughout [GSC^{+}92, TAWJ91, WSvdK^{+}93]. Thy [DSR^{+}93, LFG^{+}90]. Thy-1 [DSR^{+}93, LFG^{+}90]. thylakoid [BGBWO91, ET91, PC91b, SK94b]. thylakoids [BFR^{+}92]. thymocyte [FM94b]. Thymosin [CSNZ92]. thymus [IRH^{+}91]. thyrocytes [MCJ^{+}93]. thyroglobulin [HBS92, KBA92, PRC^{+}90]. Thyroid [MWS^{+}93, QCCD92, AL91, DH93, HBS92, PRC^{+}90, RBD92, VVR^{+}90, ZLC^{+}93]. thyrotropin [DH93, RBD92]. thyrotropin/cAMP [RBD92]. thyrotropin/cAMP-dependent [RBD92]. Thyroxine [BR94a, BCET^{+}92]. tight [BGMM^{+}93, Cit93, FHI^{+}93, FIH^{+}94, Gum93, INY^{+}93, JG94, LHH^{+}91, SAF90, ZJA^{+}94, ZSM^{+}93]. Time [FKT93, BMTTG94, KII^{+}91, KBP91, MMM91, WPS91]. Time-resolved [FKT93, MMM91]. timely [TCS^{+}94]. times [SAM^{+}91]. timing [HRH90, KHP91a, LL91, SLW^{+}92]. tip [LS92]. tips [WG93]. Tissue [Mah91, PZV^{+}91, PYH^{+}90, PTTA90, WBA92, WKM^{+}92, WA93, ZZR94, AWFP93, AW92, AA90, BIPG91, CJO92, CQYD94, CKL^{+}91, DGGG93, GTP^{+}92, HAN^{+}93, JG94, KW92b, KVF^{+}90, MLR^{+}92, PRF^{+}93, PJS^{+}94b, SPD^{+}90, SHR93, SBK^{+}91, SGG^{+}91, SSFD93, SBB91b, STA^{+}94, TLSW93, VVR^{+}90]. Tissue-dependent [Mah91]. Tissue-specific [PZV^{+}91, PYH^{+}90, PTTA90, CKL^{+}91, SSFD93, SBB91b, STA^{+}94, VVR^{+}90]. tissue-type [SPD^{+}90]. tissues [AWFP93, GWMP91, GBOB90, KKFD90, MGB90, SA90, SHB90, TQL^{+}90, VNS^{+}94, ZPS^{+}92]. titin [EK92, EKK94]. TLCK [WL91]. TLCK- [WL91]. TM [PH92]. TM-2 [PH92]. TM-3 [PH92]. TM-5a [PH92]. TM-5b [PH92]. TMA [IK94]. tomography [BZT93, HASW94, LKF^{+}94, MAFR93]. tool [RSC^{+}93]. topography [MWLC^{+}94, WS91a]. Topoisomerase [HM93, ISN^{+}94, KLC^{+}92, MMF94, RTHB90, SM92a, SY92, WMF91, WE90]. topological [GH91a]. Topologically [dlRKR^{+}90, dlR91, BHFS93]. Topology [MHV92b, BST^{+}90, SSC^{+}94]. Torpedo [CSP90, DMB92b, PC90, SCL91]. Torpedo-rat [PC90]. toxic [SPHvD91, THH^{+}90]. toxin [GCM^{+}94, GHR91, LDNM92, MIU^{+}92, SPRvD91, SRG^{+}94, SBR91, SMM^{+}91]. toxin-catalyzed [SBR91]. toxin-mediated [GCM^{+}94]. toxin-sensitive [GHR91, MIU^{+}92]. toxins [WGTCY94]. Toxoplasma [BDMPJ94]. TPCK [WL91]. TPCK-sensitive [WL91]. TPM1 [LB92c]. Tpr [BSP^{+}94, MMS93, MM93a]. TPR-containing [MM93a]. tracer [HLPL90]. tracking [WSC94]. tract [KEBD91, RHG90, TGCT90]. Traction [LLO^{+}94, HLB90]. tracts [WWHA91]. Tradescantia [ZCH90]. traffic [BMM90, GdLMS94, GRC^{+}91, GVK94, vtHvM90, HPP^{+}93b, KFN93, KDLS92, LDNM92, MCM92b, PMKH91, PALM92, RLMG92, WB93a, ZRW^{+}94]. Trafficking [GHT92, CSC90, JP94, MME93, MEM94, OTX^{+}94]. training [KKBK92]. Trans [BF93, BMB^{+}92, ABR^{+}92, BHD^{+}94, CH91, CP90, DHS93, GP91, GAS90b, HPYB93, JCSH93, KDP^{+}92, LKF^{+}94, MH93b, NHK^{+}94, PRL^{+}94, RB92, RSS^{+}94, RHB^{+}93, SGBN^{+}93, SXM94, SFS90, TGD93, VSB^{+}94, WB92a, XS93]. trans-acting [DHS93]. trans-activation [RHB^{+}93]. Trans-differentiation [BF93]. trans-Golgi [BHD^{+}94, CH91, CP90, GP91, GAS90b, HPYB93, JCSH93, LKF^{+}94, MH93b, PRL^{+}94, RB92, RSS^{+}94, SGBN^{+}93, SFS90, TGD93, VSB^{+}94, WB92a, XS93]. trans-Golgi-network-derived [KDP^{+}92]. Trans-repressor [BMB^{+}92]. transactivation [BDB90]. transactivator [LCLG94]. transcellular [MCDA^{+}93, PCD^{+}92]. transcribed [PAC^{+}92, ISDM92]. transcript [HTSP94, KDKF93, KM92d, RGKG90, WLZB93]. Transcription [CFPCL92, SBB^{+}91a, BMB^{+}92, CHM^{+}92, HGF93, KM92d, NBT94, RZS91, SBT93, TN94, WBAP91, WSvdK^{+}93]. Transcription-dependent [CFPCL92]. Transcriptional [LTP^{+}92, DTY92, GMLD92, YR93, ZSC91, dCQTR91]. Transcripts [TGR^{+}92, HLCB92, KKW^{+}93, MMD93a, SSB^{+}91]. Transcytosis [BPS90, AKM94, BS94a, CSS^{+}94, CSvdE93, GMGPM94, HC94, PHSvD92, SOPA94, WMB^{+}94]. transdifferentiation [BCG^{+}92, MELD94]. transduces [BPZ^{+}92, WSB93]. transducing [SSTL93]. transduction [ASK^{+}90, BML^{+}94, DPS^{+}93, HSG94, JH93, KDT^{+}91, KIV93, LKSR94, LD92, OKF^{+}94, QH94]. Transendothelial [GB92]. transfectants [TND^{+}93]. transfected [BSM90, BHC^{+}94, FKS93, FGH92, GMLD92, GHGP^{+}92, GFV^{+}91, JJS92, KMSW91, LSP^{+}91, NNHB^{+}92, Rob93, SVR^{+}90, SLG92, WDT^{+}92]. transfecting [CFG^{+}94]. Transfection [BZB^{+}93, LSG92, BCG^{+}92, FGS^{+}92, FGSBZ93, WTK^{+}92]. Transfer [MMC91, BBB^{+}94, HR90a, LNS^{+}94, LTPB92, MSW^{+}94, VGM^{+}93, WCM^{+}93, eSP91]. transferase [BTSL91]. transferrin [CSJD91, CvdE91, CSvdE93, GCM^{+}94, GJJM94, HSvD93b, JSM^{+}90, MPAF91, NHTP90, PMD^{+}93]. Transformation [CCT94, CBM^{+}94, CSCSA94, FPPD91, GSWW93, HKS^{+}92, HAA93, SGVS^{+}93, SHTL93, TISH94]. transformed [BVF^{+}93, DPW91, DS90b, FGS^{+}92, HSHA90, HAA93, LSAH90, ML91, MYC92, RGJ91, SZKM91, TAVC91, VBV^{+}90, WTRD92]. Transforming [DGGG93, JRSB90, KBT92, LCK^{+}92, PBM^{+}90, SR92c, YM90b, ALE^{+}92, BFTR93, Eri93, FAMR92, HXM^{+}94, HMLL94, KHR91, KNR93, LGPM90, MMG90, OSKN94, RDBHG90, SCMU93, SYG91, SSCD90, TMHKO94, VBvG^{+}94]. transgelin [SHTL93]. transgene [DAMS91, WFD^{+}91]. transgenes [CHM^{+}92]. Transgenic [KRMG93, AO90, BRF93, CSR^{+}92, GSWW93, GSM93, HKS^{+}92, LRSB94, MGD^{+}92, MHGC90, NDdC92, PYH^{+}90, RLRR93, RHG90, RRBG91, SBG93, TFC94, WDRM93, YPH^{+}92, YHM^{+}94, aUFQ^{+}92]. transglutaminase [AWFP93, GTP^{+}92, KNR93]. Transient [KBA92, LB93, MM90, PRB94, GBN^{+}93, HWLT94, SW92]. transiently [IDT^{+}92]. transients [HCYK91, TT94, TWMS90]. transin/stromelysin [MSC91]. transit [HGDA93]. Transition [MLS94, BS94c, DY93, LSK^{+}94, LK93b, SMTW94, TCS^{+}94]. transitional [PJS^{+}92, ZABM94]. transitory [YC93]. translated [LLK91]. Translation [MCDY94, Rot90, TAWJ91, YDS92]. translational [DTY92, KAV93, Koz91, SMCR90, YMO92]. translocase [NWvH94]. translocated [SMM^{+}91]. Translocation [LLK91, MLN93, RDH^{+}92, RPH^{+}92, BHFS93, BS93b, FFKC94, FS94, GN92, GVK^{+}93, GHR^{+}90, GSR^{+}93, GPH^{+}90, GFW92, KVE92, KBG91, LSP^{+}93, LELB90, LB93, MNB92, RG92, SM93a, SPB90, VMR90, WBS^{+}91a, ZBW93]. transmembrane [BLR^{+}94, BMTTG94, BS94b, CES^{+}92, DMB93, DOW92, DB90, EC93, GZC^{+}91, JNP93, JJB^{+}93, KIS^{+}94, KHW93, KTR90, LH92, NCM^{+}93, OTX^{+}94, SRK^{+}94, SSC^{+}94, SEAB91, TDSP93, WLH92, WEB90, WB92c, YLWS94]. transmembrane-anchored [KHW93]. transmission [DIYSS92]. transplantation [RB94]. Transport [AAM^{+}93, CM94a, CFG^{+}94, GAS90b, HM92a, KA91b, RAGC94, WGR^{+}92, AV91, AEGG93, BA93, BHF90, BRL^{+}93, CC90a, CS91, CRKA92, CC90b, DMB92a, FA91, FDT^{+}91, FHvZ^{+}94, FRC^{+}91, GB92, GP91, GRB^{+}90, HP92, HLWL92, Hol93, HDES94, HdHD^{+}93, JJB^{+}93, JS90, JPS91, KCH^{+}94, KM92a, KLERG94, KDP^{+}92, LPK92, LB92c, MP94, MSS^{+}94, MPEG93, MG91a, MC91b, MDKB94, MW92, MG91b, NMPN90, NHO^{+}91, NFL90, NDM^{+}94, ONN91, ON94, PP91, PSG94, PPZ^{+}93, PNPB94, PGAR90, PHB^{+}91, PIS94, PNM^{+}94, PH91, PAL^{+}90, PCP^{+}91, PDSB92, PDS90, PJS^{+}94b, PHSvD92, QORM91, RMK91, RS91, RLS94, SPRvD91, SPHvD91, SRG^{+}94, SVR^{+}90, SHDS93, SSVE91, SOPA94, SFA90, SS91b, SPBB92, SNFN91, SR90b, SK94b, SC90b, TBFP92, TH90, TGD93, TBKF^{+}92, VMB92, VGE90, WNBAS90, WCR92, Wat90, WHLW90, WRHW92]. transport [WW94a, ZKMB93, dWLS91]. transport-defective [KCH^{+}94]. transported [BPS90, GK92a, HSYYK90, HSYK^{+}91, RA90, SP94, SPW93, VHE93]. transporter [CCC^{+}94a, CCC^{+}94b, CCW^{+}93, HSP^{+}91, HFS^{+}93, KDT93, LSN^{+}94, PTS^{+}92, RPH^{+}92, SGG^{+}91, VHB93, ZSH^{+}93]. transporters [HRB92]. transporting [GC93]. transports [SON^{+}94]. Transvection [Wu93]. transverse [JAS^{+}90, YAJ90, YAJ91]. treated [DPCP91, FJN93, FJN94, GHZD90, IDT^{+}92, MAFR93, SPHvD91]. treatment [BCG^{+}92, JGLK90, PRB94, RB92, RTC^{+}94]. tri [LCT^{+}93]. tri-iodothyronine [LCT^{+}93]. Triad [FAF^{+}93]. triadin [FAF^{+}93]. triads [YAJ91]. Trichinella [Jas93]. trichohyalin [FPR90, FMCR93]. trigeminal [APP^{+}93, SHR93, WMvdK94]. trigger [GXHS94, QYC^{+}92, QH94, ZZL^{+}91]. triggered [CPU^{+}90, DSR^{+}93, LSRC93, RDH^{+}92, SF90b, TLWA94]. triggers [LNSSA93, PIRB92, RZS^{+}94, Sch93, TOA^{+}93]. triiodothyronine [KCDR91]. Trimer [CC90a]. trimeric [BS94a, WRB^{+}94]. trimers [MBC90]. tripeptide [MMS93]. triphosphate [CRS94]. triple [MGD^{+}92]. trisphosphate [FNM^{+}92, GSV93, HN90, LNSSA93, MM93b, SJIC92, SRV^{+}90, WAS^{+}91, YC93]. trkA [ARWD93]. trophoblast [YED^{+}93]. tropoelastin [HR94b]. Tropomodulin [Fow90, FSM^{+}93, WPBF94, Col94]. tropomyosin [Fow90, LB92c, SLHG93, TNS^{+}94, WLWL94, WHT^{+}93]. tropomyosin-binding [WLWL94]. tropomyosins [PH92]. troponin [BF91]. truncated [BTRB93, DSJB^{+}94, ENS91, IDT^{+}92, KSE^{+}92, MSSD93, STK^{+}93, SC90b, TIA^{+}92]. truncation [VSB^{+}94]. Trypanosoma [BWCG93, CdJC93, FTPM^{+}94, HAS92, MWLC^{+}94, RG94]. trypanosomes [BSD^{+}92, HPM^{+}94, MG94a]. TS28 [JAS^{+}90, YAJ90, YAJ91]. TSG [LWV92]. TSG-6 [LWV92]. tube [GMM^{+}93]. tubovesicular [EH94]. Tubular [TH91, RS93, TH92b, YAJ90]. Tubulation [CWBB93]. tubule [AV91, PRB^{+}91, ZSH^{+}93]. tubules [HGS^{+}94, JAS^{+}90, MTJM93, SSA^{+}90, YAJ90, YAJ91]. Tubulin [KM92b, MAWM90, BJ92, BRG^{+}90, BSF90, EWP^{+}93b, FAWM94, HLWL92, HO94, HR90b, KK92, MVCC93, NFL90, PH94b, RKA93, SH91, SW92, SBW90, WWWB90]. tubulin-tyrosine [EWP^{+}93b]. Tumor [CKD90, HK94, NS90, ASK^{+}90, FJN93, FJN94, GAHH94, GAvdM91, HJI^{+}93, HHO^{+}92, JG94, LWV92, LTP^{+}92, LMB^{+}94b, NGP^{+}91, OCMB91, PCG^{+}94, SRM94, SJM^{+}94, TBVS92, TIN^{+}93, TRP^{+}93, VKRD94, eSP91]. tumorigenesis [ALE^{+}92]. tumorigenicity [FGS^{+}92]. tumors [GKY^{+}94, MGB90, SGVS^{+}93, aUFQ^{+}92]. tunicamycin [PS94]. tunicate [Bro91b]. turn [BSW^{+}92]. turnover [HWLT94, HTSP94, NFL90, VGC94]. twisted [BB94]. Two [CPB^{+}93, DMM^{+}91, HPR93, HR90b, HHLS92, KKM91, KLM^{+}91, KKFD90, KLR^{+}93, KEBD91, QH94, SS91b, YL93, AJL^{+}90, ABR^{+}92, BSW^{+}92, BRG^{+}90, BMM^{+}94, BT93, CDS^{+}91, CMZ90, CW90c, DPS^{+}93, DE93, EK92, FWF90, GHS^{+}93, GWO^{+}93, GTC^{+}90, Has93, HMS^{+}90, HYD93, JK92, KKBL92, KA91a, KFN93, KS94, KHV^{+}93, KA93, KBL^{+}94, LS90, LXM91, LWK^{+}93, LB92b, LK93a, LHFV93, LMM^{+}91, LS92, MAA^{+}94, MF90, MB92, MPAF91, MWLC^{+}94, NYA^{+}91, NPH^{+}93, NGB^{+}93, PAC^{+}92, PRL^{+}94, PHWM92, RSC^{+}93, RKA93, RRCR93, SCMB92, SM91a, SR92b, SSK^{+}90, SSB^{+}91, TVG^{+}93, TA91, VMB92, VBV^{+}90, VRE^{+}92, WBE91, YLWS94]. two-dimensional [VBV^{+}90]. Type [DAM^{+}93, KWGS92, LGI^{+}93, SABF^{+}90, BS91, BKR90, BKFT93, BDB90, BMWT93, CMS^{+}90a, CLG^{+}92, CL93, CML90, CBM^{+}94, DG90, DSQ^{+}93, DSR^{+}93, DBC90, DAMS91, DWJP93, EMC^{+}90, FSK^{+}92, FBML90, FBL92, FBLL90, GK90, GDKP92, HSP^{+}91, HW90, HWW^{+}93, HFD91, HAG^{+}93, HFS92, KAP^{+}93, KLM^{+}91, KGDE94, KMF^{+}91, KPMG91, KP90, KCCG91, LHT^{+}91, LWF^{+}91, LSR90b, LB93, LMM^{+}91, LS92, MOM^{+}92, MELD94, MFS^{+}94, MKG93, MK91b, MSF90, ND91, NFYI93, ORS^{+}90b, PWY^{+}92, PZB^{+}94, PPJF91, PFSS92, PTD90, QvMWV^{+}91, RSM^{+}94, RWG90, RHB^{+}93, SLKS94, SMRG91, SOA^{+}93, ST94, SKH90, SPD^{+}90, SLP91, SWL90, SW93a, SGN^{+}93, SC90b, SDH90, SSB^{+}91, TN94, TRV^{+}90, VKS^{+}90, VKR^{+}91, VHM^{+}93, WHK^{+}93, XMW^{+}93, YYY^{+}91, dCV90, FBML90]. type- [DAMS91]. type-1 [FBML90, QvMWV^{+}91]. type-2 [LSR90b]. type-2A [FBML90]. type-dependent [VHM^{+}93]. type-specific [BKR90, DAM^{+}93, LS92]. types [BZB^{+}93, CHVF91, DPS^{+}93, DMB92b, Has93, HMLL94, KKM91, LS92]. typhimurium [MCDA^{+}93]. Tyrosine [BP93, BTR92, MYC92, WWD94, ACSM^{+}94, BSW^{+}92, BTH^{+}90, BVF^{+}93, BFLS94, BKFT93, DFC^{+}92, DMC^{+}91, EOL^{+}90, EWP^{+}93b, FGS91, FJN93, FJN94, GBS90, GAHB94, HWLT94, HLC^{+}93, HPYB93, JGLK90, JLL^{+}94, KYBC94, KTR90, KHHC93, LMB^{+}94a, LYK^{+}94, LHS^{+}92, Mah91, MHT^{+}92, MAWM90, MFS^{+}94, PMBL^{+}92, PBD93, SBN92, SSDK^{+}94, SMWB93, TOA^{+}91, VJB^{+}94, WLPB92, WSB93, WF90, WTRD92, WG93, ZB94]. tyrosine-5 [FGS91]. tyrosine-containing [HPYB93]. tyrosine-phosphorylated [BTH^{+}90, JLL^{+}94]. tyrosine/beta [BSW^{+}92]. tyrosine/beta-turn [BSW^{+}92]. tyrphostin [DMC^{+}91].

U [FDT^{+}91, MG92]. U1 [CFPCL92, FZ90, FHvZ^{+}94, JG92, KM92a, KKBL92, vZDF^{+}93]. U1-specific [JG92]. U1A [KM92a]. U2 [CFPCL92, FZ90, MG91a]. U4 [CFPCL92, KBL91]. U4/U6 [CFPCL92, KBL91]. U5 [CFPCL92]. U6 [CFPCL92, KBL91]. Ubiquitin [LDT^{+}92]. Ubiquitin-activating [LDT^{+}92]. ubiquitous [SDW^{+}93, SLK93]. ubiquitously [GSS^{+}91]. UCP [KCDR91]. Ultrastructural [BTBO94, SGC^{+}94, BPTS94, LSP^{+}94]. Ultrastructure [MPM^{+}94, AAM92, CDS^{+}91, GL91, RSS91]. umbilical [BPZ^{+}92, KCYH92]. unable [GdLMS94]. unaffected [PMD^{+}93]. unassembled [ET91, FGH92, WL91]. unc [GW94a, LH92, GW94b]. unc-87 [GW94a, GW94b]. unc-93 [LH92]. uncleaved [MC92]. uncoating [TOA^{+}93]. Unconfined [LZI^{+}93]. unconventional [BSD94, ECM^{+}92, HM94, KM92c, LB94, MAA^{+}94, ZC92]. uncouple [BTSL91]. Uncoupled [UYA^{+}91]. Uncoupling [CK91, SFSG91, ISN^{+}94, KCDR91]. undercoat [IYN^{+}91]. undercoat-constitutive [IYN^{+}91]. undergo [CGMC92, MCG92, PH91, SLH92]. undergoes [GN92, LFC94, LK93b, MSPB90]. unexpected [HODF94, RHG90]. unified [LH93]. unique [ACC^{+}94, AKC92, BZF^{+}92, GS94, Has93, HVNC92, JAS^{+}90, KA90, LSK^{+}94, RLS92, SA90, SLL94]. unit [HTGN94, KWFM91, SSW94, YMW^{+}90]. univalens [GGGE^{+}92]. universal [CL91b]. unlike [MNB92]. unmyelinated [CKB93]. unpolymerized [FZ93]. unsaturated [SY91, WRP^{+}91]. unstable [VFR94]. untranslated [KLST93, MCDY94]. Unusual [BWLK93, KR91, MC91a, MSF90, RMB90, SGWE94]. unusually [YLS92]. upon [RTC^{+}94, RM94]. upregulated [KDS^{+}91]. Upregulation [DDFC94, PSS^{+}93]. uptake [CNU92, TS91]. urchin [AAM92, BMV90, Bro91b, FL90, FL92, HDDS94, JGLK90, KL90, LW92, MPVL92, MMM^{+}92, MNV93, MSFW92, MGW92, ODPL93, OPL94, SMCR90, SMTW94, TS91, TJ91, WCW91, WHW^{+}91, WTS93]. urinary [JHGR93]. urokinase [BML^{+}94, EMC^{+}90, KEBD91, MMR91, OCMB91, PSS^{+}93, QvMWV^{+}91]. urokinase-mediated [QvMWV^{+}91]. urokinase-type [QvMWV^{+}91]. uromodulin [MNV93]. Uroplakin [YMW^{+}90]. Uroplakins [YLWS94]. urothelium [YMW^{+}90]. Use [CSR^{+}92, SLS^{+}90, EWP^{+}93a, Has93, HW90, KCC90, SBR91]. used [HS90b]. uses [LSK^{+}94]. using [BI90, CHM^{+}92, CYC91, FNV^{+}92, FL92, KRCT93, KB91b, MM91, ND92, NEL^{+}91, RKB^{+}90, RSS91, SRP90, SS90a, VVB^{+}92, WE90, WTS93]. uso1 [NHO^{+}91]. UT [BF90]. UT-1 [BF90]. utilization [MP94, PLB^{+}92]. utilizes [SHS^{+}93]. UV [SSS^{+}90]. uvomorulin [CO91, LSMRB90, NSWH90, OK90, OK92]. uvomorulin-catenin [OK92].

V [LGI^{+}93, ACC^{+}94, BGLB94, CNSC90, DG90, FKB^{+}90a, HR90a, HXM^{+}94, HR90c, KYBC94, LD92, LGSB93, MBC^{+}91, MHT^{+}92, SHB90, TISH94, VLH^{+}93, WOC91, WFCN94, WTRD92, XMW^{+}93, ZMV^{+}93, ECM^{+}92, LGI^{+}93, MH91, WK92, WS92, BVF^{+}93, LD91]. v-mos [FKB^{+}90a]. v-myc [HR90a]. v-sis [HXM^{+}94, LD92, XMW^{+}93]. v-src [HR90a, HR90c, MHT^{+}92, SHB90, TISH94]. v-src-transformed [WTRD92]. V2 [JPJF91]. V2-receptor [JPJF91]. v6 [SGVS^{+}93]. Va [MC91a, MC93]. vaccinia [SDE^{+}93]. vacuolar [CB92, GE91, MR94, ROE^{+}90, RHR90, ST94, SKSD^{+}94, VKS^{+}90, VWS90, VRE^{+}92, VHE93]. vacuole [BDMPJ94, CSV^{+}92, CHW94, CKGC91, Dun90a, Dun90b, HCW94, KCY92, RNS92, SR90b, VGE90, VHE93, ZC92]. vacuoles [HZC93, LDT^{+}92, LRAG92, OA92]. Variable [SKBD91, ASR^{+}90, HASW94, WE90]. variant [BBK^{+}93, HHH^{+}93, JRK^{+}94, MTS^{+}94, MLKB92, TRS^{+}90]. variants [HTHC93, HAN^{+}93, SNGG94, SBL^{+}91, VFC^{+}91]. Variations [CKP91, BSD^{+}92, VHM^{+}93]. variety [FDT^{+}91]. various [JKB^{+}92, KKFD90, XMW^{+}93]. vary [CMV^{+}92]. vas [VPP^{+}93]. vascular [AMBN91, BRWB91, BIPG91, LKD^{+}94, MMG90, MSB^{+}91, MMR91, NEL^{+}91, OVB^{+}94, RCJ^{+}94, WMS^{+}93, dAWS^{+}90, eSP91]. VASE [SWD94, SA90]. vasopressin [LVA^{+}90, JPJF91]. vasopressin-sensitive [LVA^{+}90]. vast [SMCA91]. Vault [KHCR91]. Vaults [KMBR90, KHCR91]. VCAM [MH93a, MMI^{+}91, RM94, RPS^{+}92, SNG^{+}91, VTSS94]. VCAM-1 [MH93a, RM94, RPS^{+}92, SNG^{+}91, VTSS94]. VCAM-like [MMI^{+}91]. VCP [FFR^{+}91]. vector [WTK^{+}92]. Vectorial [LSMRB90, JPS91]. vein [BPZ^{+}92, KCYH92]. velocity [WDWT90]. venules [WIF^{+}90]. vero [MIU^{+}92]. Versican [ZDZSBT94]. versus [LH93, SBC^{+}93]. vertebral [SMRG91]. vertebrate [AR91, DBSK90, EEC^{+}92, HCKJ92, Koz91, KWW91, MVCC93, PMG^{+}92, RSCS94, RS94, RB90, SHBD90, SR92b, WBW^{+}94]. vertebrates [SKSC94, WCR93]. very [MTJM93, SDW^{+}93]. Vesicle [SFA90, BLR^{+}94, BVR^{+}93, BCKS92, BDW92, CSJD91, CRWS93, CJC93, COLL^{+}90, CC90b, DBWS94, EOO^{+}94, FYDB93, GSS^{+}91, HSvD91, KHKO^{+}93, KDO^{+}94, LKF^{+}94, LVBK92, LK93a, MLR^{+}92, MMD^{+}93b, OHP^{+}94, PSSK93, RMG90, SAH^{+}92, SS91a, SS91b, TSBS92, VL91, WRHW92, WMCG91, vdBRD^{+}93]. vesicle-like [COLL^{+}90]. vesicle-mediated [CRWS93]. vesicles [AU90, BTH^{+}90, BCBK92, BH94, BHC^{+}94, EGW^{+}93, FB93, GCM^{+}94, HPM^{+}94, HKKO^{+}94, HWM90, JPS91, JCSH93, KDP^{+}92, LVA^{+}90, LLG^{+}94, LSN^{+}94, MHG93, MTC^{+}91, MTP^{+}92, MLR^{+}92, MJG^{+}91, MW94, NHK^{+}94, ON94, PGAR90, PNM^{+}94, PDC91, RS91, RLS94, Rob90, SLB^{+}94, STZG91, TLXM90, TOA^{+}93, VYM93, WGR^{+}92, WNBAS90, WHK90, WRHW92, WB92a, WW91, WMCW92, XS93, ZJA^{+}94, dWLS91]. vesicular [AEGG93, BF90, BMM90, GP91, GRB^{+}90, HP92, vtHvM90, HPP^{+}93b, KLERG94, LDNM92, LB92c, LSN^{+}94, LTPB92, NMPN90, PCP^{+}91, SP94, SPBB92, SNFN91, TBKF^{+}92, WHLW90, dSBH90, dSBH93]. vesiculation [VYM93]. vessel [MSB^{+}91]. Vg [WBWH93]. Vg1 [KMN91]. Vgr [GKY^{+}94]. Vgr-1 [GKY^{+}94]. Vgr-1/BMP- [GKY^{+}94]. VI [DBC90, HM94, KWGS92, ORS^{+}90b, SSJ^{+}94, SDH90]. via [AKM94, AKC92, BRIA^{+}94, CM94a, CNZ^{+}92, DDW90, DFC^{+}92, DSR^{+}93, FGS91, GSBS92, HRT^{+}91, KA92, LKD^{+}94, MYT94, MG92, RJJ^{+}94, SMSH92, SHH94, TMHKO94, TN94, TGD93, VHE93, WMS^{+}93, ZDR^{+}92]. viability [Dav92]. viable [STS91]. vibrating [KJ90]. vibration [KK92]. view [KII^{+}91]. VII [KBT92]. VIII [SKH90]. villin [FFKL92, FFB90]. villus [CSR^{+}92, HKS^{+}92]. villus-associated [HKS^{+}92]. vimentin [CK94, DCB^{+}92, JETS91, MHG93, MKSF93, MBHG91, RPE^{+}90, SNE90, SLI^{+}91, SBC^{+}93, TNT^{+}90]. vimentin-binding [MBHG91]. vimentin-free [RPE^{+}90]. Vinculin [BW91, VFR94, FGS^{+}92, FGSBZ93, GWO^{+}93, LDS^{+}92, MKB^{+}94, NGB^{+}93, SEC^{+}93, TGB90]. vinculin-binding [TGB90]. Vinculin-deficient [VFR94]. vinculin/sarcomeric [LDS^{+}92]. vinculin/sarcomeric-alpha-actinin/alpha-actin [LDS^{+}92]. VIP21 [KDP^{+}92]. viral [HH94a, MHV92a, PP91, SZKM91, TIA93, ZPS^{+}92]. virions [DWY^{+}93]. virus [BF90, BDB90, BR91, CC90a, GHR^{+}90, HWF92, HWF93, KLERG94, LMW^{+}94, LTPB92, LMB^{+}94b, MPK93, MGB90, PK91, PMA91, PRB^{+}91, RHB^{+}93, SWL^{+}92, SBSG92, SWM^{+}92, SDE^{+}93, WG92a, XMW^{+}93, ZPS^{+}92, aUFQ^{+}92, dSBH90, dSBH93]. virus-encoded [PMA91]. virus-induced [MGB90]. visceral [RWG90]. Viscoelastic [JETS91]. viscosity [FV91, LZI^{+}93]. visual [PLS92]. Visualization [BB94, TT94, HBR90, KIN^{+}94, MWH^{+}91, SBLV92]. visualized [BPTS94, OOBJ92]. vitamin [BPDM90]. vitro [AW92, AV91, ACFFL93, AFHDD90, BRG^{+}90, BRIA^{+}94, BRD94, BDS92, BGLJ^{+}93, BDW92, CHBGL91, CMS^{+}90b, CWBB93, CSM90, CSV^{+}92, CHW94, CLM91, CCAF90, DIDB^{+}91, DJ93b, DP93, EKK94, FAWM94, FPP91, FAF^{+}93, GMM^{+}93, GBN^{+}93, GT94, GP91, GK92b, HCW94, HGF93, HW90, HM91, HM90, INM^{+}93, JT93, KBCM^{+}90, KIN^{+}94, KBT92, KQJ91, KCCG91, LCC^{+}93, LM91, LLK91, MHYC90, MSC92, ML94, MLS94, MVM90, MCF93, NT94, OH93, OPB90, PBM^{+}90, PPZ^{+}93, PH92, PSSK93, QDT^{+}90, RGJ91, REW90, RRH91, RLS92, SM91a, SM91b, SM90, SS91a, SKT^{+}90, SBR93, Smi94, SCS92, SSH94, SCSE90, TW93, VGE90, WPS91, WDWT90, WF93, WW94a, WDB^{+}92, WCF92, WE90, WWN92, YKI^{+}91, ZS94, dSBH90]. Vitronectin [LHH^{+}91, BSG^{+}92, CNSC90, CML90, EGP^{+}91, KFW^{+}91, KYBC94, LFWC92, MHM91, NVR92, Sch93, SWL90, TCDH93, VLH^{+}93, WOC91]. vitronectin-dependent [CML90]. vivax [BWCG93]. vivo [BLT^{+}90, CWBB93, CCAF90, DGL^{+}90, DP94, EBS91, FCF92, FFKL92, GKY^{+}94, HS92, Hen90, HSYYK90, HSYK^{+}91, HWD94, HDES94, KBCM^{+}90, KIN^{+}94, LXWC93, MG94b, MBS91, MKG93, NPST94, OMS^{+}94, OCMB91, PH92, PB90, SM90, SW93a, WBWH93, WWWB90, WBS^{+}91a, WTK^{+}92, YYM94, YMO92, ZRW^{+}94]. VLA [TMP^{+}91, ASMC^{+}92, BBK90, CPU^{+}90, CH93, EUH91, vdWvKvKdB^{+}92, KCYH92, MH93a, MHYK92, OVB^{+}94, VTSS94]. VLA-1 [BBK90]. VLA-2 [CH93]. VLA-3 [TMP^{+}91, EUH91]. VLA-4 [CPU^{+}90, MH93a, MHYK92, OVB^{+}94, VTSS94]. VLA-5 [MHYK92]. VLA-dependent [KCYH92]. VLDL [TLXM90, TMI^{+}91]. VM [RTHB90]. VM-26 [RTHB90]. voltage [BU94, GFWM92, MAFR93, MHG^{+}90, PET^{+}91, RSS91]. voltage-dependent [GFWM92]. voltage-gated [MHG^{+}90, PET^{+}91]. voltage-sensitive [BU94]. volume [SLK91]. Volvox [KRMK93]. VP7 [SPB90]. VPS1 [VRE^{+}92]. VPS3 [ROE^{+}90]. vs [TWC93]. vulgaris [KAP^{+}93, SBH92].

wall [BSHB90, HWW94, LS92]. water [LVA^{+}90, NSC^{+}93, SSC^{+}94, VBS^{+}93, ZSH^{+}93]. water-permeable [NSC^{+}93]. wave [HLWL92, SSJ90]. waves [FMJ91]. way [FRSL90]. weak [BK91]. weakness [RFF92]. WEHI [LKXS93]. WEHI-3B [LKXS93]. weight [AHS^{+}92, FFW91, GWMP91, HAS92, LMW^{+}94, NFL90, TSH^{+}93]. well [GHBA^{+}93]. WGA [RBW93]. wheat [LRAG92]. where [TOA^{+}91]. whereas [HR90a, HWM^{+}93, TSBS92]. which [BOH^{+}92, BDZ^{+}92, DY93, FM94a, LVA^{+}90, LEM93, RT92, SOH^{+}90, VGMS91, YKM^{+}92]. whisker [SHR93]. whose [BSP^{+}94, HKS90, KDKF93, SON^{+}94, WRB^{+}94]. wide [LCMP90]. widely [PRF^{+}93, SJG93]. WIF [INM^{+}93]. WIF-B [INM^{+}93]. wild [CBM^{+}94, MOM^{+}92, MK91b, RWG90]. wild-type [CBM^{+}94, MOM^{+}92, MK91b, RWG90]. Willebrand [VFC^{+}91, YYY^{+}91]. Wilson [Blo91]. wing [WYM^{+}92, WA93]. withdrawal [ET94, MWG92]. withdrawal-induced [MWG92]. within [BSSL^{+}91, BHC^{+}94, CMV^{+}92, CBJER93, CCC^{+}94a, CCC^{+}94b, DKL^{+}90, HG93b, HSG94, KAP^{+}93, KA94, Mac92, MCDY94, PLB^{+}92, RG94, RHG90, SCE91, SDR^{+}94, SFS90, SBL^{+}91, VTSS94, VFC^{+}91, WL91, WSSM91, XL91, YM93, YKY92]. without [CLZ91, DS90c, GHR91, LD91, MSSD93, MHGC90, OHF^{+}94, RYK^{+}90, THH^{+}90, UYA^{+}91]. Wnt [BCB93, HNP94]. Wnt-1 [BCB93, HNP94]. wool [MPR90]. wound [BFM93, CGF^{+}93, ES92, Gri94, JRK^{+}94, KQJ91, WOC90]. wounding [SFF93]. wounds [EVL^{+}91].

X [BTG^{+}93, FWD^{+}92, BS91, CLG^{+}92, KCCG91, MSI^{+}94, VIF^{+}93]. XB [BTG^{+}93, DSJB^{+}94]. XB- [DSJB^{+}94]. Xenopus [AL92, AR93, ABD^{+}94, BP93, BHF90, BDS92, BG94, BDK92, CC90a, CSMG90, CPM^{+}91, DPW91, DLMS91, DK92, DCB^{+}92, DMM^{+}91, DFC^{+}92, DDW94, DSJB^{+}94, FMD93, FM94a, FAWM94, FKS93, FHvZ^{+}94, FKB^{+}90a, GHZD90, GKG90, GC90, GC91, HN90, HFN92, HODF94, HM93, KSK^{+}93, KDP^{+}90, KMF^{+}91, KA93, KT93, LL91, LSK^{+}94, LK93a, MF90, MSC92, MBG93, MMCH91, MKF91, OH93, OW90, ORS^{+}90b, PET^{+}91, PBD93, PSN^{+}91, RLMG92, SCL91, SJIC92, SPvV94, SN93, SR92d, SVK^{+}92, STdL^{+}90, SSB^{+}91, SR92e, SM92b, TKP^{+}91, VGC94, VBAK91, VDS^{+}92, VL91]. XII [KLM^{+}91, LMM^{+}91, YYY^{+}91]. XII-like [KLM^{+}91, LMM^{+}91]. XIV [BMWT93]. XIX [JD91]. XMAP [VGC94]. XMAP230 [ABD^{+}94]. XTC [SM92b].

Y-shaped [KBL91]. Yarrowia [LNS^{+}94, YMO92]. YDJ1 [CD91]. Yeast [BSHB90, CCR93, CB92, FFR^{+}91, SXM94, SWMS90, VWS90, VHE93, dBBHA94, AJL^{+}90, AGKC92, BTBO94, BWLK93, BZT94, BR94b, BLT^{+}90, BDH^{+}94, BHFS93, BS93b, BDJ94, CD91, CS91, CSG92, CYCA90, DHS93, DP90, DMM93, EJ93, EPNN91, FMS93, FRC^{+}91, FHUY93, GtHG^{+}94, GE91, GRB^{+}90, GHK94, HRH90, HR94a, HH94b, HTD90, HO94, JHK^{+}91, JLC93, KFN93, KLC^{+}92, KDT93, KBL^{+}94, LR93, LB92c, LKS94, LNS^{+}94, MHYC90, MSC92, MSTY90, MY92, MSW^{+}94, MT94, MB94, MC91a, MC93, MCS92, MJLD93, MPTW90, MPM^{+}94, MB93, MR94, NRS93b, ON94, OSJS94, PSHK92, PH94b, PZB^{+}94, PC91c, ROE^{+}90, RHF91, RNS92, RPPB94, RK90b, RB93, RT92, SY94a, SCM90, SBK94, SHDS93, SNFN91, SY92, SKSD^{+}94, SR90b, SSH94, SS93, SOKS91, SLK93, SHOA92, TBT^{+}92, TDSP93, TFP94, VGE90, VMR90, WHLW90]. yeast [WRB92, WB93a, WB94, WCM^{+}93, WGBB91, WHC^{+}93, WBR94, YMO92]. Yersinia [YFS92]. yields [BLFQ93]. Yolk [FM94a, BBB91, MPVL92, RGR90]. Ypt51p [SKSD^{+}94]. Ypt52p [SKSD^{+}94]. Ypt53p [SKSD^{+}94].

Z [FKB^{+}90b, Lut91, MNS90, SGGC^{+}90]. Z-band [Lut91, MNS90]. zag [HASW94]. Zellweger [MMS93]. zero [CRS94]. zeta [KHKO^{+}93]. zeta-COP [KHKO^{+}93]. zig [HASW94]. zig-zag [HASW94]. zinc [SCM^{+}93]. ZO [FIH^{+}94, INY^{+}93, JG94, SAF90, ZJA^{+}94, ZSM^{+}93]. ZO-1 [FIH^{+}94, INY^{+}93, JG94, SAF90, ZJA^{+}94, ZSM^{+}93]. ZO-2 [JG94, ZSM^{+}93]. zona [MGDS93]. zone [LHH^{+}91, ZDZSBT94]. zones [LSC^{+}93]. ZP3 [CLP^{+}94, KMSW91]. zw10 [WKMG92]. zygote [BGBWO91]. zygotes [SMTW94]. Zyxin [SCMB92, CMB92].

References

Armstrong:1990:IRI

[AA90] P. B. Armstrong and M. T. Armstrong. An instructive role for the interstitial matrix in tissue patterning: tissue segregation and intercellular invasion. Journal of Cell Biology, 110(4):1439–??, April 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/110/4/1439.

Adam:1994:ICF

[AA94] E. J. Adam and S. A. Adam. Identification of cytosolic factors required for nuclear location sequence-mediated binding to the nuclear envelope. Journal of Cell Biology, 125 (3):547–??, May 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/125/3/547.

Adelson:1992:UHC

[AAM92] D. L. Adelson, M. C. Alliegro, and D. R. McClay. On the ultrastructure of hyalin, a cell adhesion protein of the sea urchin embryo extracellular matrix. Journal of Cell Biology, 116(5):1283–??, March 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/116/5/1283.

Ainger:1993:TLE

[AAM^{+}93] K. Ainger, D. Avossa, F. Morgan, S. J. Hill, C. Barry, E. Barbarese, and J. H. Carson. Transport and localization of exogenous myelin basic protein mRNA microinjected into oligodendrocytes. Journal of Cell Biology, 123(2):431–??, October 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/123/2/431.

Acton:1990:PCL

[AB90] S. L. Acton and F. M. Brodsky. Predominance of clathrin light chain LCb correlates with the presence of a regulated secretory pathway. Journal of Cell Biology, 111(4):1419–??, October 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/111/4/1419.

Albrecht-Buehler:1991:SEC

[AB91] G. Albrecht-Buehler. Surface extensions of 3T3 cells towards distant infrared light sources. Journal of Cell Biology, 114(3):493–??, August 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/114/3/493.

Andersen:1994:EMD

[ABD^{+}94] S. S. Andersen, B. Buendia, J. E. Domínguez, A. Sawyer, and E. Karsenti. Effect on microtubule dynamics of XMAP230, a microtubule-associated protein present in Xenopus laevis eggs and dividing cells. Journal of Cell Biology, 127(5):1289–??, December 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/127/5/1289.

Arcangeli:1993:IMN

[ABM^{+}93] A. Arcangeli, A. Becchetti, A. Mannini, G. Mugnai, P. De Filippi, G. Tarone, M. R. Del Bene, E. Barletta, E. Wanke, and M. Olivotto. Integrin-mediated neurite outgrowth in neuroblastoma cells depends on the activation of potassium channels. Journal of Cell Biology, 122(5):1131–??, September 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/122/5/1131.

Alcalde:1992:ADG

[ABR^{+}92] J. Alcalde, P. Bonay, A. Roa, S. Vilaro, and I. V. Sandoval. Assembly and disassembly of the Golgi complex: two processes arranged in a cis-trans direction. Journal of Cell Biology, 116(1):69–??, January 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/116/1/69.

Amatruda:1992:PCI

[AC92] J. F. Amatruda and J. A. Cooper. Purification, characterization, and immunofluorescence localization of Saccharomyces cerevisiae capping protein. Journal of Cell Biology, 117(5):1067–??, June 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/117/5/1067.

Aigner:1993:DKG

[AC93] L. Aigner and P. Caroni. Depletion of 43-kD growthassociated protein in primary sensory neurons leads to diminished formation and spreading of growth cones. Journal of Cell Biology, 123(2):417–??, October 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/123/2/417.

Agrez:1994:AVB

[ACC^{+}94] M. Agrez, A. Chen, R. I. Cone, R. Pytela, and D. Sheppard. The alpha v beta 6 integrin promotes proliferation of colon carcinoma cells through a unique region of the beta 6 cytoplasmic domain. Journal of Cell Biology, 127(2):547–??, October 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/127/2/547.

Antoniou:1993:NOS

[ACFFL93] M. Antoniou, M. Carmo-Fonseca, J. Ferreira, and A. I. Lamond. Nuclear organization of splicing snRNPs during differentiation of murine erythroleukemia cells in vitro. Journal of Cell Biology, 123(5):1055–??, December 1993. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/123/5/
1055.

Arroyo:1994:ITP

[ACSM^{+}94] A. G. Arroyo, M. R. Campanero, P. Sánchez-Mateos, J. M. Zapata, M. A. Ursa, M. A. del Pozo, and F. Sánchez-Madrid. Induction of tyrosine phosphorylation during ICAM-3 and LFA-1-mediated intercellular adhesion, and its regulation by the CD45 tyrosine phosphatase. Journal of Cell Biology, 126(5):1277–??, September 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/126/5/1277.

Adams:1992:MCS

[Ada92] D. S. Adams. Mechanisms of cell shape change: the cytomechanics of cellular response to chemical environment and mechanical loading. Journal of Cell Biology, 117(1): 83–??, April 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/117/1/83.

Aniento:1993:CDD

[AEGG93] F. Aniento, N. Emans, G. Griffiths, and J. Gruenberg. Cytoplasmic dynein-dependent vesicular transport from early to late endosomes. Journal of Cell Biology, 123(6):1373– ??, December 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/123/6/1373.

Alcalde:1994:GMG

[AES94] J. Alcalde, G. Egea, and I. V. Sandoval. gp74 a membrane glycoprotein of the cis-Golgi network that cycles through the endoplasmic reticulum and intermediate compartment. Journal of Cell Biology, 124(5):649–??, March 1994. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/124/5/
649.

Arpin:1994:FDB

[AFA^{+}94] M. Arpin, E. Friederich, M. Algrain, F. Vernel, and D. Louvard. Functional differences between L- and T-plastin isoforms. Journal of Cell Biology, 127(6):1995–??, December 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/127/6/1995.

Armstrong:1990:VAO

[AFHDD90] R. Armstrong, V. L. Friedrich, K. V. Holmes, and M. Dubois-Dalcq. In vitro analysis of the oligodendrocyte lineage in mice during demyelination and remyelination. Journal of Cell Biology, 111(3):1183–??, September 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/111/3/
1183.

Altankov:1993:DIP

[AG93] G. Altankov and F. Grinnell. Depletion of intracellular potassium disrupts coated pits and reversibly inhibits cell polarization during fibroblast spreading. Journal of Cell Biology, 120(6):1449–??, March 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/120/6/1449.

Aubert:1991:IPP

[AGB^{+}91] D. Aubert, M. Garcia, M. Benchaibi, D. Poncet, Y. Chebloune, G. Verdier, V. Nigon, J. Samarut, and C. V. Mura. Inhibition of proliferation of primary avian fibroblasts through expression of histone H5 depends on the degree of phosphorylation of the protein. Journal of Cell Biology, 113 (3):497–??, May 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/113/3/497.

Arnold:1992:RAI

[AGGS92] H. H. Arnold, C. D. Gerharz, H. E. Gabbert, and A. Salminen. Retinoic acid induces myogenin synthesis and myogenic differentiation in the rat rhabdomyosarcoma cell line BA–Han–1C. Journal of Cell Biology, 118(4):877–??, August 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/118/4/877.

Amatruda:1992:ENM

[AGKC92] J. F. Amatruda, D. J. Gattermeir, T. S. Karpova, and J. A. Cooper. Effects of null mutations and overexpression of capping protein on morphogenesis, actin distribution and polarized secretion in yeast. Journal of Cell Biology, 119(5): 1151–??, December 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/119/5/1151.

Anglister:1994:GAA

[AHM94] L. Anglister, B. Haesaert, and U. J. McMahan. Globular and asymmetric acetylcholinesterase in the synaptic basal lamina of skeletal muscle. Journal of Cell Biology, 125(1): 183–??, April 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/125/1/183.

Asakura:1992:ICA

[AHS^{+}92] S. Asakura, R. W. Hurley, K. Skorstengaard, I. Ohkubo, and D. F. Mosher. Inhibition of cell adhesion by high molecular weight kininogen. Journal of Cell Biology, 116(2): 465–??, January 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/116/2/465.

Arikawa:1990:IAF

[AHW90] K. Arikawa, J. L. Hicks, and D. S. Williams. Identification of actin filaments in the rhabdomeral microvilli of Drosophila photoreceptors. Journal of Cell Biology, 110(6):1993–??, June 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/110/6/1993.

Adams:1990:CCT

[AJL^{+}90] A. E. Adams, D. I. Johnson, R. M. Longnecker, B. F. Sloat, and J. R. Pringle. CDC42 and CDC43, two additional genes involved in budding and the establishment of cell polarity in the yeast Saccharomyces cerevisiae. Journal of Cell Biology, 111(1):131–??, July 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/111/1/131.

Audigier:1990:CTD

[AJPB90] Y. Audigier, L. Journot, C. Pantaloni, and J. Bockaert. The carboxy-terminal domain of Gs alpha is necessary for anchorage of the activated form in the plasma membrane. Journal of Cell Biology, 111(4):1427–??, October 1990. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/111/4/
1427.

Arber:1992:CRI

[AKC92] S. Arber, K. H. Krause, and P. Caroni. s-cyclophilin is retained intracellularly via a unique COOH-terminal sequence and colocalizes with the calcium storage protein calreticulin. Journal of Cell Biology, 116(1):113–??, January 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/116/1/
113.

Aroeti:1993:MSS

[AKK^{+}93] B. Aroeti, P. A. Kosen, I. D. Kuntz, F. E. Cohen, and K. E. Mostov. Mutational and secondary structural analysis of the basolateral sorting signal of the polymeric immunoglobulin receptor. Journal of Cell Biology, 123(5):1149–??, December 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/123/5/1149.

Apodaca:1994:RMT

[AKM94] G. Apodaca, L. A. Katz, and K. E. Mostov. Receptormediated transcytosis of IgA in MDCK cells is via apical recycling endosomes. Journal of Cell Biology, 125(1):67–??, April 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/125/1/67.

Arvan:1991:RCP

[AL91] P. Arvan and J. Lee. Regulated and constitutive protein targeting can be distinguished by secretory polarity in thyroid epithelial cells. Journal of Cell Biology, 112(3):365–??, February 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/112/3/365.

Adachi:1992:INP

[AL92] Y. Adachi and U. K. Laemmli. Identification of nuclear pre-replication centers poised for DNA synthesis in Xenopus egg extracts: immunolocalization study of replication protein A. Journal of Cell Biology, 119(1):1–??, October 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/119/1/
1.

Arrick:1992:AMA

[ALE^{+}92] B. A. Arrick, A. R. Lopez, F. Elfman, R. Ebner, C. H. Damsky, and R. Derynck. Altered metabolic and adhesive properties and increased tumorigenesis associated with increased expression of transforming growth factor beta 1. Journal of Cell Biology, 118(3):715–??, August 1992. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/118/3/
715.

Ascoli:1991:INN

[AM91] C. A. Ascoli and G. G. Maul. Identification of a novel nuclear domain. Journal of Cell Biology, 112(5):785–??, March 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/112/5/785.

Andreassen:1994:MDP

[AM94] P. R. Andreassen and R. L. Margolis. Microtubule dependency of p34cdc2 inactivation and mitotic exit in mammalian cells. Journal of Cell Biology, 127(3):789–??, November 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/127/3/789.

Albelda:1991:MCP

[AMBN91] S. M. Albelda, W. A. Muller, C. A. Buck, and P. J. Newman. Molecular and cellular properties of PECAM-1 (endoCAM/ CD31): a novel vascular cell–cell adhesion molecule. Journal of Cell Biology, 114(5):1059–??, September 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/114/5/1059.

Adam:1990:NPI

[AMG90] S. A. Adam, R. S. Marr, and L. Gerace. Nuclear protein import in permeabilized mammalian cells requires soluble cytoplasmic factors. Journal of Cell Biology, 111(3):807– ??, September 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/111/3/807.

Arikawa:1992:LPR

[AMMW92] K. Arikawa, L. L. Molday, R. S. Molday, and D. S. Williams. Localization of peripherin/rds in the disk membranes of cone and rod photoreceptors: relationship to disk membrane morphogenesis and retinal degeneration. Journal of Cell Biology, 116(3):659–??, February 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/116/3/659.

Anglister:1991:AMN

[Ang91] L. Anglister. Acetylcholinesterase from the motor nerve terminal accumulates on the synaptic basal lamina of the myofiber. Journal of Cell Biology, 115(3):755–??, November 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/115/3/755.

Abe:1990:SHK

[AO90] M. Abe and R. G. Oshima. A single human keratin 18 gene is expressed in diverse epithelial cells of transgenic mice. Journal of Cell Biology, 111(3):1197–??, September 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/111/3/
1197.

Albelda:1990:ENE

[AORB90] S. M. Albelda, P. D. Oliver, L. H. Romer, and C. A. Buck. EndoCAM: a novel endothelial cell–cell adhesion molecule. Journal of Cell Biology, 110(4):1227–??, April 1990. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/110/4/
1227.

Applegate:1992:AFA

[AP92] D. Applegate and J. D. Pardee. Actin-facilitated assembly of smooth muscle myosin induces formation of actomyosin fibrils. Journal of Cell Biology, 117(6):1223–??, June 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/117/6/
1223.

Apgar:1991:RAI

[Apg91] J. R. Apgar. Regulation of the antigen-induced F-actin response in rat basophilic leukemia cells by protein kinase C. Journal of Cell Biology, 112(6):1157–??, March 1991. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/112/6/
1157.

Adamson:1992:ILP

[APH92] P. Adamson, H. F. Paterson, and A. Hall. Intracellular localization of the P21rho proteins. Journal of Cell Biology, 119(3):617–??, November 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/119/3/617.

Arumae:1993:NTR

[APP^{+}93] U. Arumäe, U. Pirvola, J. Palgi, T. R. Kiema, K. Palm, M. Moshnyakov, J. Ylikoski, and M. Saarma. Neurotrophins and their receptors in rat peripheral trigeminal system during maxillary nerve growth. Journal of Cell Biology, 122 (5):1053–??, September 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/122/5/1053.

Alexander:1991:CMD

[AR91] S. P. Alexander and C. L. Rieder. Chromosome motion during attachment to the vertebrate spindle: initial saltatorylike behavior of chromosomes and quantitative analysis of force production by nascent kinetochore fibers. Journal of Cell Biology, 113(4):805–??, May 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/113/4/805.

Akey:1993:AXN

[AR93] C. W. Akey and M. Radermacher. Architecture of the Xenopus nuclear pore complex revealed by three-dimensional cryo-electron microscopy. Journal of Cell Biology, 122(1): 1–??, July 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/122/1/1.

Alon:1994:DCS

[ARW^{+}94] R. Alon, H. Rossiter, X. Wang, T. A. Springer, and T. S. Kupper. Distinct cell surface ligands mediate T lymphocyte attachment and rolling on P and E selectin under physiological flow. Journal of Cell Biology, 127(5):1485–??, December 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/127/5/1485.

Allsopp:1993:ETE

[ARWD93] T. E. Allsopp, M. Robinson, S. Wyatt, and A. M. Davies. Ectopic trkA expression mediates a NGF survival response in NGF-independent sensory neurons but not in parasympathetic neurons. Journal of Cell Biology, 123(6):1555–??, December 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/123/6/1555.

Avides:1994:ICA

[AS94] M. C. Avides and C. E. Sunkel. Isolation of chromosomeassociated proteins from Drosophila melanogaster that bind a human centromeric DNA sequence. Journal of Cell Biology, 127(5):1159–??, December 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/127/5/1159.

Allinquant:1991:EEM

[ASDC91] B. Allinquant, S. M. Staugaitis, D. D’Urso, and D. R. Colman. The ectopic expression of myelin basic protein isoforms in Shiverer oligodendrocytes: implications for myelinogenesis. Journal of Cell Biology, 113(2):393–??, April 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/113/2/393.

Aridor:1990:NPS

[ASE90] M. Aridor and R. Sagi-Eisenberg. Neomycin is a potent secretagogue of mast cells that directly activates a GTPbinding protein involved in exocytosis. Journal of Cell Biology, 111(6):2885–??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/111/6/2885.

Akwa:1993:ANM

[ASG^{+}93] Y. Akwa, N. Sananès, M. Gouézou, P. Robel, E. E. Baulieu, and C. Le Goascogne. Astrocytes and neurosteroids: metabolism of pregnenolone and dehydroepiandrosterone. Regulation by cell density. Journal of Cell Biology, 121(1):135–??, April 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/121/1/135.

Aznavoorian:1990:STC

[ASK^{+}90] S. Aznavoorian, M. L. Stracke, H. Krutzsch, E. Schiffmann, and L. A. Liotta. Signal transduction for chemotaxis and haptotaxis by matrix molecules in tumor cells. Journal of Cell Biology, 110(4):1427–??, April 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/110/4/1427.

Ayalon:1994:STR

[ASL^{+}94] O. Ayalon, H. Sabanai, M. G. Lampugnani, E. Dejana, and B. Geiger. Spatial and temporal relationships between cadherins and PECAM-1 in cell–cell junctions of human endothelial cells. Journal of Cell Biology, 126(1):247–??, July 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/126/1/247.

Arroyo:1992:RVI

[ASMC^{+}92] A. G. Arroyo, P. Sánchez-Mateos, M. R. Campanero, I. Martín-Padura, E. Dejana, and F. Sánchez-Madrid. Regulation of the VLA integrin-ligand interactions through the beta 1 subunit. Journal of Cell Biology, 117(3):659–??, May 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/117/3/659.

Asahina:1993:HOP

[ASNH93] I. Asahina, T. K. Sampath, I. Nishimura, and P. V. Hauschka. Human osteogenic protein-1 induces both chondroblastic and osteoblastic differentiation of osteoprogenitor cells derived from newborn rat calvaria. Journal of Cell Biology, 123(4):921–??, November 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/123/4/921.

Athey:1990:DFH

[ASR^{+}90] B. D. Athey, M. F. Smith, D. A. Rankert, S. P. Williams, and J. P. Langmore. The diameters of frozen-hydrated chromatin fibers increase with DNA linker length: evidence in support of variable diameter models for chromatin. Journal of Cell Biology, 111(3):795–??, September 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/111/3/795.

Acheson:1991:NPA

[ASR91] A. Acheson, J. L. Sunshine, and U. Rutishauser. NCAM polysialic acid can regulate both cell–cell and cell-substrate interactions. Journal of Cell Biology, 114(1):143–??, July 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/
114/1/143.

Aizawa:1992:KFM

[AST^{+}92] H. Aizawa, Y. Sekine, R. Takemura, Z. Zhang, M. Nangaku, and N. Hirokawa. Kinesin family in murine central nervous system. Journal of Cell Biology, 119(5):1287–??, December 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/119/5/1287.

Argraves:1990:FEM

[ATBD90] W. S. Argraves, H. Tran, W. H. Burgess, and K. Dickerson. Fibulin is an extracellular matrix and plasma glycoprotein with repeated domain structure. Journal of Cell Biology, 111(6):3155–??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/111/6/3155.

Ashcom:1990:HAM

[ATD^{+}90] J. D. Ashcom, S. E. Tiller, K. Dickerson, J. L. Cravens, W. S. Argraves, and D. K. Strickland. The human alpha 2-macroglobulin receptor: identification of a 420-kD cell surface glycoprotein specific for the activated conformation of alpha 2-macroglobulin. Journal of Cell Biology, 110 (4):1041–??, April 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/110/4/1041.

Aridor:1990:EMC

[ATSE90] M. Aridor, L. M. Traub, and R. Sagi-Eisenberg. Exocytosis in mast cells by basic secretagogues: evidence for direct activation of GTP-binding proteins. Journal of Cell Biology, 111(3):909–??, September 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/111/3/909.

Algrain:1993:ECC

[ATV^{+}93] M. Algrain, O. Turunen, A. Vaheri, D. Louvard, and M. Arpin. Ezrin contains cytoskeleton and membrane binding domains accounting for its proposed role as a membranecytoskeletal linker. Journal of Cell Biology, 120(1):129–??, January 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/120/1/129.

Ahle:1990:ANI

[AU90] S. Ahle and E. Ungewickell. Auxilin, a newly identified clathrin-associated protein in coated vesicles from bovine brain. Journal of Cell Biology, 111(1):19–??, July 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/111/1/
19.

al-Ubaidi:1992:BRB

[aUFQ^{+}92] M. R. al Ubaidi, R. L. Font, A. B. Quiambao, M. J. Keener, G. I. Liou, P. A. Overbeek, and W. Baehr. Bilateral retinal and brain tumors in transgenic mice expressing simian virus 40 large T antigen under control of the human interphotoreceptor retinoid-binding protein promoter. Journal of Cell Biology, 119(6):1681–??, December 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/119/6/1681.

Aoki:1991:NCA

[AUT^{+}91] J. Aoki, M. Umeda, K. Takio, K. Titani, H. Utsumi, M. Sasaki, and K. Inoue. Neural cell adhesion molecule mediates contact-dependent inhibition of growth of neardiploid mouse fibroblast cell line m5S/1M. Journal of Cell Biology, 115(6):1751–??, December 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/115/6/1751.

Allan:1991:CCC

[AV91] V. J. Allan and R. D. Vale. Cell cycle control of microtubulebased membrane transport and tubule formation in vitro. Journal of Cell Biology, 113(2):347–??, April 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/113/2/347.

Augustin-Voss:1992:MEC

[AVP92] H. G. Augustin-Voss and B. U. Pauli. Migrating endothelial cells are distinctly hyperglycosylated and express specific migration-associated cell surface glycoproteins. Journal of Cell Biology, 119(2):483–??, October 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/119/2/483.

Adams:1991:EBB

[AW91] J. C. Adams and F. M. Watt. Expression of beta 1, beta 3, beta 4, and beta 5 integrins by human epidermal keratinocytes and non-differentiating keratinocytes. Journal of Cell Biology, 115(3):829–??, November 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/115/3/829.

Alexander:1992:TDT

[AW92] C. M. Alexander and Z. Werb. Targeted disruption of the tissue inhibitor of metalloproteinases gene increases the invasive behavior of primitive mesenchymal cells derived from embryonic stem cells in vitro. Journal of Cell Biology, 118 (3):727–??, August 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/118/3/727.

Aeschlimann:1993:ETT

[AWFP93] D. Aeschlimann, A. Wetterwald, H. Fleisch, and M. Paulsson. Expression of tissue transglutaminase in skeletal tissues correlates with events of terminal differentiation of chondrocytes. Journal of Cell Biology, 120(6):1461–??, March 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/120/6/
1461.

Amizuka:1994:PHR

[AWH^{+}94] N. Amizuka, H. Warshawsky, J. E. Henderson, D. Goltzman, and A. C. Karaplis. Parathyroid hormone-related peptide-depleted mice show abnormal epiphyseal cartilage development and altered endochondral bone formation. Journal of Cell Biology, 126(6):1611–??, September 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/126/6/
1611.

Abeijon:1993:GDR

[AYM^{+}93] C. Abeijon, K. Yanagisawa, E. C. Mandon, A. Häusler, K. Moremen, C. B. Hirschberg, and P. W. Robbins. Guanosine diphosphatase is required for protein and sphingolipid glycosylation in the Golgi lumen of Saccharomyces cerevisiae. Journal of Cell Biology, 122(2):307–??, July 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/122/2/

Baas:1992:PES

[BA92] P. W. Baas and F. J. Ahmad. The plus ends of stable microtubules are the exclusive nucleating structures for microtubules in the axon. Journal of Cell Biology, 116(5): 1231–??, March 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/116/5/1231.

Baas:1993:TPA

[BA93] P. W. Baas and F. J. Ahmad. The transport properties of axonal microtubules establish their polarity orientation. Journal of Cell Biology, 120(6):1427–??, March 1993. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/120/6/
1427.

Baitinger:1990:MCC

[BAP^{+}90] C. Baitinger, J. Alderton, M. Poenie, H. Schulman, and R. A. Steinhardt. Multifunctional Ca^{2+}/calmodulindependent protein kinase is necessary for nuclear envelope breakdown. Journal of Cell Biology, 111(5):1763–??, November 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/111/5/1763.

Baas:1990:IMA

[BB90] P. W. Baas and M. M. Black. Individual microtubules in the axon consist of domains that differ in both composition and stability. Journal of Cell Biology, 111(2):495–??, August 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/111/2/495.

Blau:1991:DRC

[BB91] H. M. Blau and D. Baltimore. Differentiation requires continuous regulation. Journal of Cell Biology, 112(5):781–??, March 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/112/5/781.

Bertagnolli:1993:ESA

[BB93] M. E. Bertagnolli and M. C. Beckerle. Evidence for the selective association of a subpopulation of GPIIb–IIIa with the actin cytoskeletons of thrombin-activated platelets. Journal of Cell Biology, 121(6):1329–??, June 1993. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/121/6/
1329.

Belmont:1994:VGC

[BB94] A. S. Belmont and K. Bruce. Visualization of G1 chromosomes: a folded, twisted, supercoiled chromonema model of interphase chromatid structure. Journal of Cell Biology, 127(2):287–??, October 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/127/2/287.

Butterworth:1991:GMY

[BBB91] F. M. Butterworth, M. Bownes, and V. S. Burde. Genetically modified yolk proteins precipitate in the adult Drosophila fat body. Journal of Cell Biology, 112(4):727–??, February 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/112/4/727.

Baccetti:1994:HPS

[BBB^{+}94] B. Baccetti, A. Benedetto, A. G. Burrini, G. Collodel, E. C. Ceccarini, N. Crisa, A. Di Caro, M. Estenoz, A. R. Garbuglia, A. Massacesi, P. Piomboni, T. Renieri, and D. Solazzo. HIV-particles in spermatozoa of patients with AIDS and their transfer into the oocyte. Journal of Cell Biology, 127(4):903–??, November 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/127/4/903.

Bomblies:1990:MEC

[BBD^{+}90] L. Bomblies, E. Biegelmann, V. Döring, G. Gerisch, H. Krafft-Czepa, A. A. Noegel, M. Schleicher, and B. M. Humbel. Membrane-enclosed crystals in Dictyostelium discoideum cells, consisting of developmentally regulated proteins with sequence similarities to known esterases. Journal of Cell Biology, 110(3):669–??, March 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/110/3/669.

Buendia:1990:CCC

[BBGK90] B. Buendia, M. H. Bré, G. Griffiths, and E. Karsenti. Cytoskeletal control of centrioles movement during the establishment of polarity in Madin–Darby canine kidney cells. Journal of Cell Biology, 110(4):1123–??, April 1990. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/110/4/
1123.

Brown:1991:HKE

[BBJ^{+}91] T. A. Brown, T. Bouchard, T. St John, E. Wayner, and W. G. Carter. Human keratinocytes express a new CD44 core protein (CD44E) as a heparan-sulfate intrinsic membrane proteoglycan with additional exons. Journal of Cell Biology, 113(1):207–??, April 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/113/1/207.

Belkin:1990:HSM

[BBK90] V. M. Belkin, A. M. Belkin, and V. E. Koteliansky. Human smooth muscle VLA-1 integrin: purification, substrate specificity, localization in aorta, and expression during development. Journal of Cell Biology, 111(5):2159–??, November 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/111/5/2159.

Baines:1992:DLA

[BBK92] I. C. Baines, H. Brzeska, and E. D. Korn. Differential localization of Acanthamoeba myosin I isoforms. Journal of Cell Biology, 119(5):1193–??, December 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/119/5/1193.

Balzac:1993:EFA

[BBK^{+}93] F. Balzac, A. M. Belkin, V. E. Koteliansky, Y. V. Balabanov, F. Altruda, L. Silengo, and G. Tarone. Expression and functional analysis of a cytoplasmic domain variant of the beta 1 integrin subunit. Journal of Cell Biology, 121 (1):171–??, April 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/121/1/171.

Bernstein:1994:NKL

[BBKR94] M. Bernstein, P. L. Beech, S. G. Katz, and J. L. Rosenbaum. A new kinesin-like protein (Klp1) localized to a single microtubule of the Chlamydomonas flagellum. Journal of Cell Biology, 125(6):1313–??, June 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/125/6/1313.

Beck:1994:GSI

[BBMN94] K. A. Beck, J. A. Buchanan, V. Malhotra, and W. J. Nelson. Golgi spectrin: identification of an erythroid beta-spectrin homolog associated with the Golgi complex. Journal of Cell Biology, 127(3):707–??, November 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/127/3/707.

Bernat:1990:IAA

[BBRE90] R. L. Bernat, G. G. Borisy, N. F. Rothfield, and W. C. Earnshaw. Injection of anticentromere antibodies in interphase disrupts events required for chromosome movement at mitosis. Journal of Cell Biology, 111(4):1519–??, October 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/111/4/
1519.

Brancolini:1992:GGA

[BBS92] C. Brancolini, S. Bottega, and C. Schneider. Gas2, a growth arrest-specific protein, is a component of the microfilament network system. Journal of Cell Biology, 117(6):1251–??, June 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/117/6/1251.

Bates:1994:AII

[BBvH^{+}94] R. C. Bates, A. Buret, D. F. van Helden, M. A. Horton, and G. F. Burns. Apoptosis induced by inhibition of intercellular contact. Journal of Cell Biology, 125(2):403–??, April 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/125/2/

Balczon:1994:PKC

[BBZ94] R. Balczon, L. Bao, and W. E. Zimmer. PCM-1, A 228-kD centrosome autoantigen with a distinct cell cycle distribution. Journal of Cell Biology, 124(5):783–??, March 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/124/5/
783.

Bradley:1993:EWP

[BCB93] R. S. Bradley, P. Cowin, and A. M. Brown. Expression of Wnt-1 in PC12 cells results in modulation of plakoglobin and E-cadherin and increased cellular adhesion. Journal of Cell Biology, 123(6):1857–??, December 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/123/6/1857.

Beck:1992:CAP

[BCBK92] K. A. Beck, M. Chang, F. M. Brodsky, and J. H. Keen. Clathrin assembly protein AP-2 induces aggregation of membrane vesicles: a possible role for AP-2 in endosome formation. Journal of Cell Biology, 119(4):787–??, November 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/119/4/787.

Bohme:1992:IPH

[BCET^{+}92] K. Böhme, M. Conscience-Egli, T. Tschan, K. H. Winterhalter, and P. Bruckner. Induction of proliferation or hypertrophy of chondrocytes in serum-free culture: the role of insulin-like growth factor–I, insulin, or thyroxine. Journal of Cell Biology, 116(4):1035–??, February 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/116/4/1035.

Buxton:1993:NDC

[BCF^{+}93] R. S. Buxton, P. Cowin, W. W. Franke, D. R. Garrod, K. J. Green, I. A. King, P. J. Koch, A. I. Magee, D. A. Rees, and J. R. Stanley. Nomenclature of the desmosomal cadherins. Journal of Cell Biology, 121(3):481–??, May 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/121/3/481.

Boukamp:1992:PST

[BCG^{+}92] P. Boukamp, J. Chen, F. Gonzales, P. A. Jones, and N. E. Fusenig. Progressive stages of “transdifferentiation” from epidermal to mesenchymal phenotype induced by MyoD1 transfection, 5-aza-2’-deoxycytidine treatment, and selection for reduced cell attachment in the human keratinocyte line HaCaT. Journal of Cell Biology, 116(5):1257–??, March 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/116/5/1257.

Bennett:1992:SVM

[BCKS92] M. K. Bennett, N. Calakos, T. Kreiner, and R. H. Scheller. Synaptic vesicle membrane proteins interact to form a multimeric complex. Journal of Cell Biology, 116(3):761–??, February 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/116/3/761.

Bacon:1994:DDM

[BCLM94] R. A. Bacon, C. J. Cohen, D. A. Lewin, and I. Mellman. Dictyostelium discoideum mutants with temperaturesensitive defects in endocytosis. Journal of Cell Biology, 127(2):387–??, October 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/127/2/387.

Brown:1994:CLA

[BCYC94] K. D. Brown, R. M. Coulson, T. J. Yen, and D. W. Cleveland. Cyclin-like accumulation and loss of the putative kinetochore motor CENP–E results from coupling continuous synthesis with specific degradation at the end of mitosis. Journal of Cell Biology, 125(6):1303–??, June 1994. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/125/6/
1303.

Brockerhoff:1992:CCR

[BD92] S. E. Brockerhoff and T. N. Davis. Calmodulin concentrates at regions of cell growth in Saccharomyces cerevisiae. Journal of Cell Biology, 118(3):619–??, August 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/118/3/619.

Brake:1990:IAG

[BDB90] D. A. Brake, C. Debouck, and G. Biesecker. Identification of an Arg–Gly–Asp (RGD) cell adhesion site in human immunodeficiency virus type 1 transactivation protein, tat. Journal of Cell Biology, 111(3):1275–??, September 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/111/3/1275.

Bressan:1993:ECE

[BDGC^{+}93] G. M. Bressan, D. Daga-Gordini, A. Colombatti, I. Castellani, V. Marigo, and D. Volpin. Emilin, a component of elastic fibers preferentially located at the elastin-microfibrils interface. Journal of Cell Biology, 121(1):201–??, April 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/121/1/
201.

Bretscher:1994:WBF

[BDH^{+}94] A. Bretscher, B. Drees, E. Harsay, D. Schott, and T. Wang. What are the basic functions of microfilaments? Insights from studies in budding yeast. Journal of Cell Biology, 126 (4):821–??, August 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/126/4/821.

Burgess:1994:MEM

[BDJ94] S. M. Burgess, M. Delannoy, and R. E. Jensen. MMM1 encodes a mitochondrial outer membrane protein essential for establishing and maintaining the structure of yeast mitochondria. Journal of Cell Biology, 126(6):1375–??, September 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/126/6/1375.

Buendia:1992:RMN

[BDK92] B. Buendia, G. Draetta, and E. Karsenti. Regulation of the microtubule nucleating activity of centrosomes in Xenopus egg extracts: role of cyclin A-associated protein kinase. Journal of Cell Biology, 116(6):1431–??, March 1992. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/116/6/

Beckers:1994:TGR

[BDMPJ94] C. J. Beckers, J. F. Dubremetz, O. Mercereau-Puijalon, and K. A. Joiner. The Toxoplasma gondii rhoptry protein ROP 2 is inserted into the parasitophorous vacuole membrane, surrounding the intracellular parasite, and is exposed to the host cell cytoplasm. Journal of Cell Biology, 127(4):947– ??, November 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/127/4/947.

Bell:1992:VAP

[BDS92] P. Bell, M. C. Dabauvalle, and U. Scheer. In vitro assembly of prenucleolar bodies in Xenopus egg extract. Journal of Cell Biology, 118(6):1297–??, September 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/118/6/1297.

Boman:1992:GHR

[BDW92] A. L. Boman, M. R. Delannoy, and K. L. Wilson. GTP hydrolysis is required for vesicle fusion during nuclear envelope assembly in vitro. Journal of Cell Biology, 116(2): 281–??, January 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/116/2/281.

Bussolino:1992:HGF

[BDZ^{+}92] F. Bussolino, M. F. Di Renzo, M. Ziche, E. Bocchietto, M. Olivero, L. Naldini, G. Gaudino, L. Tamagnone, A. Coffer, and P. M. Comoglio. Hepatocyte growth factor is a potent angiogenic factor which stimulates endothelial cell motility and growth. Journal of Cell Biology, 119(3):629– ??, November 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/119/3/629.

Bearer:1991:DOA

[Bea91] E. L. Bearer. Direct observation of actin filament severing by gelsolin and binding by gCap39 and CapZ. Journal of Cell Biology, 115(6):1629–??, December 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/115/6/1629.

Bailer:1991:CKP

[BEGN91] S. M. Bailer, H. M. Eppenberger, G. Griffiths, and E. A. Nigg. Characterization of A 54-kD protein of the inner nuclear membrane: evidence for cell cycle-dependent interaction with the nuclear lamina. Journal of Cell Biology, 114(3):389–??, August 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/114/3/389.

Brennan:1990:ARM

[BEO90] T. J. Brennan, D. G. Edmondson, and E. N. Olson. Aberrant regulation of MyoD1 contributes to the partially defective myogenic phenotype of BC3H1 cells. Journal of Cell Biology, 110(4):929–??, April 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/110/4/929.

Bergmann:1990:GPV

[BF90] J. E. Bergmann and P. J. Fusco. The G protein of vesicular stomatitis virus has free access into and egress from the smooth endoplasmic reticulum of UT-1 cells. Journal of Cell Biology, 110(3):625–??, March 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/110/3/625.

Beall:1991:MAD

[BF91] C. J. Beall and E. Fyrberg. Muscle abnormalities in Drosophila melanogaster heldup mutants are caused by missing or aberrant troponin–I isoforms. Journal of Cell Biology, 114(5):941–??, September 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/114/5/941.

Boukamp:1993:TDE

[BF93] P. Boukamp and N. E. Fusenig. ”trans-differentiation” from epidermal to mesenchymal/myogenic phenotype is associated with a drastic change in cell–cell and cell-matrix adhesion molecules. Journal of Cell Biology, 120(4):981–??, February 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/120/4/981.

Bowen:1990:MAB

[BFL90] B. R. Bowen, C. Fennie, and L. A. Lasky. The Mel 14 antibody binds to the lectin domain of the murine peripheral lymph node homing receptor. Journal of Cell Biology, 110(1):147–??, January 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/110/1/147.

Berton:1994:BID

[BFLS94] G. Berton, L. Fumagalli, C. Laudanna, and C. Sorio. Beta 2 integrin-dependent protein tyrosine phosphorylation and activation of the FGR protein tyrosine kinase in human neutrophils. Journal of Cell Biology, 126(4):1111–??, August 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/126/4/1111.

Bement:1993:NCS

[BFM93] W. M. Bement, P. Forscher, and M. S. Mooseker. A novel cytoskeletal structure involved in purse string wound closure and cell polarity maintenance. Journal of Cell Biology, 121 (3):565–??, May 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/121/3/565.

Barbato:1992:SCL

[BFR^{+}92] R. Barbato, G. Friso, F. Rigoni, F. Dalla Vecchia, and G. M. Giacometti. Structural changes and lateral redistribution of photosystem II during donor side photoinhibition of thylakoids. Journal of Cell Biology, 119(2):325–??, October 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/119/2/325.

Butzow:1993:KPM

[BFTR93] R. Bützow, D. Fukushima, D. R. Twardzik, and E. Ruoslahti. A 60-kD protein mediates the binding of transforming growth factor-beta to cell surface and extracellular matrix proteoglycans. Journal of Cell Biology, 122(3): 721–??, August 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/122/3/721.

Batistatou:1991:AAR

[BG91a] A. Batistatou and L. A. Greene. Aurintricarboxylic acid rescues PC12 cells and sympathetic neurons from cell death caused by nerve growth factor deprivation: correlation with suppression of endonuclease activity. Journal of Cell Biology, 115(2):461–??, October 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/115/2/461.

Boylan:1991:OCR

[BG91b] J. F. Boylan and L. J. Gudas. Overexpression of the cellular retinoic acid binding protein–I (CRABP–I) results in a reduction in differentiation-specific gene expression in F9 teratocarcinoma cells. Journal of Cell Biology, 112(5):965–??, March 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/112/5/965.

Batistatou:1993:IDC

[BG93] A. Batistatou and L. A. Greene. Internucleosomal DNA cleavage and neuronal cell survival/death. Journal of Cell Biology, 122(3):523–??, August 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/122/3/523.

Brieher:1994:RCC

[BG94] W. M. Brieher and B. M. Gumbiner. Regulation of Ccadherin function during activin induced morphogenesis of Xenopus animal caps. Journal of Cell Biology, 126(2):519– ??, July 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/126/2/519.

Bixby:1994:CIN

[BGB94] J. L. Bixby, G. B. Grunwald, and R. J. Bookman. Ca^{2+} influx and neurite growth in response to purified N-cadherin and laminin. Journal of Cell Biology, 127(5):1461–??, December 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/127/5/1461.

Baldan:1991:ETM

[BGBWO91] B. Baldan, J. Girard-Bascou, F. A. Wollman, and J. Olive. Evidence for thylakoid membrane fusion during zygote formation in Chlamydomonas reinhardtii. Journal of Cell Biology, 114(5):905–??, September 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/114/5/905.

Barlow:1994:SEH

[BGGW^{+}94] S. Barlow, M. L. Gonzalez-Garay, R. R. West, J. B. Olmsted, and F. Cabral. Stable expression of heterologous microtubule-associated proteins (MAPs) in Chinese hamster ovary cells: evidence for differing roles of MAPs in microtubule organization. Journal of Cell Biology, 126(4): 1017–??, August 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/126/4/1017.

Beaumelle:1990:IPC

[BGH90] B. D. Beaumelle, A. Gibson, and C. R. Hopkins. Isolation and preliminary characterization of the major membrane boundaries of the endocytic pathway in lymphocytes. Journal of Cell Biology, 111(5):1811–??, November 1990. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/111/5/
1811.

Brown:1993:MRM

[BGH93] M. T. Brown, L. Goetsch, and L. H. Hartwell. MIF2 is required for mitotic spindle integrity during anaphase spindle elongation in Saccharomyces cerevisiae. Journal of Cell Biology, 123(2):387–??, October 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/123/2/387.

Brink:1990:DML

[BGI^{+}90] M. Brink, G. Gerisch, G. Isenberg, A. A. Noegel, J. E. Segall, E. Wallraff, and M. Schleicher. A Dictyostelium mutant lacking an F-actin cross-linking protein, the 120kD gelation factor. Journal of Cell Biology, 111(4):1477–??, October 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/111/4/1477.

Berndorff:1994:LIC

[BGK^{+}94] D. Berndorff, R. Gessner, B. Kreft, N. Schnoy, A. M. Lajous-Petter, N. Loch, W. Reutter, M. Hortsch, and R. Tauber. Liver-intestine cadherin: molecular cloning and characterization of a novel Ca(2+)-dependent cell adhesion molecule expressed in liver and intestine. Journal of Cell Biology, 125(6):1353–??, June 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/125/6/1353.

Blystone:1994:IAV

[BGLB94] S. D. Blystone, I. L. Graham, F. P. Lindberg, and E. J. Brown. Integrin alpha v beta 3 differentially regulates adhesive and phagocytic functions of the fibronectin receptor alpha 5 beta 1. Journal of Cell Biology, 127(4):1129– ??, November 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/127/4/1129.

Bloch-Gallego:1993:AHP

[BGLJ^{+}93] E. Bloch-Gallego, I. Le Roux, A. H. Joliot, M. Volovitch, C. E. Henderson, and A. Prochiantz. Antennapedia homeobox peptide enhances growth and branching of embryonic chicken motoneurons in vitro. Journal of Cell Biology, 120 (2):485–??, January 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/120/2/485.

Burgoon:1991:SCN

[BGM^{+}91] M. P. Burgoon, M. Grumet, V. Mauro, G. M. Edelman, and B. A. Cunningham. Structure of the chicken neuron-glia cell adhesion molecule, Ng–CAM: origin of the polypeptides and relation to the Ig superfamily. Journal of Cell Biology, 112(5):1017–??, March 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/112/5/1017.

Balda:1993:ATJ

[BGMM^{+}93] M. S. Balda, L. Gonzalez-Mariscal, K. Matter, M. Cereijido, and J. M. Anderson. Assembly of the tight junction: the role of diacylglycerol. Journal of Cell Biology, 123(2): 293–??, October 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/123/2/293.

Brunner:1991:PCR

[BGRW91] G. Brunner, J. Gabrilove, D. B. Rifkin, and E. L. Wilson. Phospholipase C release of basic fibroblast growth factor from human bone marrow cultures as a biologically active complex with a phosphatidylinositol-anchored heparan sulfate proteoglycan. Journal of Cell Biology, 114(6):1275– ??, September 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/114/6/1275.

Betz:1994:OAD

[BH94] W. J. Betz and A. W. Henkel. Okadaic acid disrupts clusters of synaptic vesicles in frog motor nerve terminals. Journal of Cell Biology, 124(5):843–??, March 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/124/5/843.

Bogerd:1994:NME

[BHA^{+}94] A. M. Bogerd, J. A. Hoffman, D. C. Amberg, G. R. Fink, and L. I. Davis. nup1 mutants exhibit pleiotropic defects in nuclear pore complex function. Journal of Cell Biology, 127(2):319–??, October 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/127/2/319.

Balasubramanian:1994:SPC

[BHBG94] M. K. Balasubramanian, B. R. Hirani, J. D. Burke, and K. L. Gould. The Schizosaccharomyces pombe cdc3+ gene encodes a profilin essential for cytokinesis. Journal of Cell Biology, 125(6):1289–??, June 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/125/6/1289.

Bonzelius:1994:PIR

[BHC^{+}94] F. Bonzelius, G. A. Herman, M. H. Cardone, K. E. Mostov, and R. B. Kelly. The polymeric immunoglobulin receptor accumulates in specialized endosomes but not synaptic vesicles within the neurites of transfected neuroendocrine PC12 cells. Journal of Cell Biology, 127(6):1603–??, December 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/127/6/1603.

Bosshart:1994:CDM

[BHD^{+}94] H. Bosshart, J. Humphrey, E. Deignan, J. Davidson, J. Drazba, L. Yuan, V. Oorschot, P. J. Peters, and J. S. Bonifacino. The cytoplasmic domain mediates localization of furin to the trans-Golgi network en route to the endosomal/lysosomal system. Journal of Cell Biology, 126(5): 1157–??, September 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/126/5/1157.

Bataille:1990:CTR

[BHF90] N. Bataillé, T. Helser, and H. M. Fried. Cytoplasmic transport of ribosomal subunits microinjected into the Xenopus laevis oocyte nucleus: a generalized, facilitated process. Journal of Cell Biology, 111(4):1571–??, October 1990. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/111/4/
1571.

Brodsky:1993:RPT

[BHFS93] J. L. Brodsky, S. Hamamoto, D. Feldheim, and R. Schekman. Reconstitution of protein translocation from solubilized yeast membranes reveals topologically distinct roles for BiP and cytosolic Hsc70. Journal of Cell Biology, 120(1): 95–??, January 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/120/1/95.

Brown:1990:IAP

[BHHG90] E. Brown, L. Hooper, T. Ho, and H. Gresham. Integrinassociated protein: a 50-kD plasma membrane antigen physically and functionally associated with integrins. Journal of Cell Biology, 111(6):2785–??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/111/6/2785.

Burkhardt:1990:LGN

[BHLA90] J. K. Burkhardt, S. Hester, C. K. Lapham, and Y. Argon. The lytic granules of natural killer cells are dualfunction organelles combining secretory and pre-lysosomal compartments. Journal of Cell Biology, 111(6):2327–??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/111/6/2327.

Braakman:1991:FIH

[BHLWH91] I. Braakman, H. Hoover-Litty, K. R. Wagner, and A. Helenius. Folding of influenza hemagglutinin in the endoplasmic reticulum. Journal of Cell Biology, 114(3):401–??, August 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/114/3/401.

Beck:1990:IRP

[BHS90] L. A. Beck, T. J. Hosick, and M. Sinensky. Isoprenylation is required for the processing of the lamin A precursor. Journal of Cell Biology, 110(5):1489–??, May 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/110/5/1489.

Begovac:1991:LFE

[BHS91] P. C. Begovac, D. E. Hall, and B. D. Shur. Laminin fragment E8 mediates PC12 cell neurite outgrowth by binding to cell surface beta 1,4 galactosyltransferase. Journal of Cell Biology, 113(3):637–??, May 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/113/3/637.

Barkalow:1994:RDK

[BHS94] K. Barkalow, T. Hamasaki, and P. Satir. Regulation of 22S dynein by a 29-kD light chain. Journal of Cell Biology, 126(3):727–??, August 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/126/3/727.

Birkenberger:1990:PDT

[BI90] L. Birkenberger and W. Ip. Properties of the desmin tail domain: studies using synthetic peptides and antipeptide antibodies. Journal of Cell Biology, 111(5):2063–??, November 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/111/5/2063.

Bradham:1991:CTG

[BIPG91] D. M. Bradham, A. Igarashi, R. L. Potter, and G. R. Grotendorst. Connective tissue growth factor: a cysteine-rich mitogen secreted by human vascular endothelial cells is related to the SRC-induced immediate early gene product CEF-10. Journal of Cell Biology, 114(6):1285–??, September 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/114/6/1285.

Bischoff:1990:CCC

[Bis90] R. Bischoff. Cell cycle commitment of rat muscle satellite cells. Journal of Cell Biology, 111(1):201–??, July 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/111/1/
201.

Bixby:1990:EMM

[BJ90] J. L. Bixby and P. Jhabvala. Extracellular matrix molecules and cell adhesion molecules induce neurites through different mechanisms. Journal of Cell Biology, 111(6):2725–??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/111/6/2725.

Baas:1992:GTD

[BJ92] P. W. Baas and H. C. Joshi. Gamma-tubulin distribution in the neuron: implications for the origins of neuritic microtubules. Journal of Cell Biology, 119(1):171–??, October 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/119/1/171.

Baines:1990:LMI

[BK90] I. C. Baines and E. D. Korn. Localization of myosin IC and myosin II in Acanthamoeba castellanii by indirect immunofluorescence and immunogold electron microscopy. Journal of Cell Biology, 111(5):1895–??, November 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/111/5/

Butner:1991:TPB

[BK91] K. A. Butner and M. W. Kirschner. Tau protein binds to microtubules through a flexible array of distributed weak sites. Journal of Cell Biology, 115(3):717–??, November 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/115/3/
717.

Basson:1990:SSE

[BKB^{+}90] C. T. Basson, W. J. Knowles, L. Bell, S. M. Albelda, V. Castronovo, L. A. Liotta, and J. A. Madri. Spatiotemporal segregation of endothelial cell integrin and nonintegrin extracellular matrix-binding proteins during adhesion events. Journal of Cell Biology, 110(3):789–??, March 1990. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/110/3/
789.

Brady-Kalnay:1993:HBP

[BKFT93] S. M. Brady-Kalnay, A. J. Flint, and N. K. Tonks. Homophilic binding of PTP mu, a receptor-type protein tyrosine phosphatase, can mediate cell–cell aggregation. Journal of Cell Biology, 122(4):961–??, August 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/122/4/961.

Brada:1990:CTS

[BKR90] D. Brada, D. Kerjaschki, and J. Roth. Cell type-specific post-Golgi apparatus localization of a “resident” endoplasmic reticulum glycoprotein, glucosidase II. Journal of Cell Biology, 110(2):309–??, February 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/110/2/309.

Bahler:1994:RMD

[BKSB94] M. Bähler, R. Kroschewski, H. E. Stöffler, and T. Behrmann. Rat myr 4 defines a novel subclass of myosin I: identification, distribution, localization, and mapping of calmodulinbinding sites with differential calcium sensitivity. Journal of Cell Biology, 126(2):375–??, July 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/126/2/375.

Byers:1991:SDD

[BKW91] T. J. Byers, L. M. Kunkel, and S. C. Watkins. The subcellular distribution of dystrophin in mouse skeletal, cardiac, and smooth muscle. Journal of Cell Biology, 115(2): 411–??, October 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/115/2/411.

Belanger:1994:GPI

[BKWD94] K. D. Belanger, M. A. Kenna, S. Wei, and L. I. Davis. Genetic and physical interactions between Srp1p and nuclear pore complex proteins Nup1p and Nup2p. Journal of Cell Biology, 126(3):619–??, August 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/126/3/619.

Brown:1994:EFD

[BL94] M. J. Brown and L. M. Loew. Electric field-directed fibroblast locomotion involves cell surface molecular reorganization and is calcium independent. Journal of Cell Biology, 127(1):117–??, October 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/127/1/117.

Bober:1991:MRG

[BLB^{+}91] E. Bober, G. E. Lyons, T. Braun, G. Cossu, M. Buckingham, and H. H. Arnold. The muscle regulatory gene, Myf-6, has a biphasic pattern of expression during early mouse development. Journal of Cell Biology, 113(6):1255– ??, June 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/113/6/1255.

Brooks:1993:SIY

[BLFQ93] P. C. Brooks, J. M. Lin, D. L. French, and J. P. Quigley. Subtractive immunization yields monoclonal antibodies that specifically inhibit metastasis. Journal of Cell Biology, 122(6):1351–??, September 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/122/6/1351.

Blobel:1991:BWM

[Blo91] G. K. Blobel. E. B. Wilson medalist, 1990. Journal of Cell Biology, 113(1):223–??, April 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/113/1/223.

Bloch:1992:CNC

[Blo92] R. J. Bloch. Clusters of neural cell adhesion molecule at sites of cell–cell contact. Journal of Cell Biology, 116(2): 449–??, January 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/116/2/449.

Blow:1993:PRR

[Blo93] J. J. Blow. Preventing re-replication of DNA in a single cell cycle: evidence for a replication licensing factor. Journal of Cell Biology, 122(5):993–??, September 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/122/5/993.

Bidanset:1992:RCS

[BLR^{+}92] D. J. Bidanset, R. LeBaron, L. Rosenberg, J. E. Murphy-Ullrich, and M. Hook. Regulation of cell substrate adhesion: effects of small galactosaminoglycan-containing proteoglycans. Journal of Cell Biology, 118(6):1523–??, September 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/118/6/1523.

Banfield:1994:LSP

[BLR^{+}94] D. K. Banfield, M. J. Lewis, C. Rabouille, G. Warren, and H. R. Pelham. Localization of Sed5, a putative vesicle targeting molecule, to the cis-Golgi network involves both its transmembrane and cytoplasmic domains. Journal of Cell Biology, 127(2):357–??, October 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/127/2/357.

Berton:1992:GSA

[BLSR92] G. Berton, C. Laudanna, C. Sorio, and F. Rossi. Generation of signals activating neutrophil functions by leukocyte integrins: LFA-1 and gp150/95, but not CR3, are able to stimulate the respiratory burst of human neutrophils. Journal of Cell Biology, 116(4):1007–??, February 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/116/4/1007.

Bresnahan:1990:HGE

[BLT^{+}90] P. A. Bresnahan, R. Leduc, L. Thomas, J. Thorner, H. L. Gibson, A. J. Brake, P. J. Barr, and G. Thomas. Human fur gene encodes a yeast KEX2-like endoprotease that cleaves pro–beta–NGF in vivo. Journal of Cell Biology, 111(6): 2851–??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/111/6/2851.

Beckmann:1992:EFR

[BLW92] R. P. Beckmann, M. Lovett, and W. J. Welch. Examining the function and regulation of hsp 70 in cells subjected to metabolic stress. Journal of Cell Biology, 117(6):1137– ??, June 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/117/6/1137.

Blount:1990:MAM

[BM90] P. Blount and J. P. Merlie. Mutational analysis of muscle nicotinic acetylcholine receptor subunit assembly. Journal of Cell Biology, 111(6):2613–??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/111/6/2613.

Blount:1991:BAN

[BM91a] P. Blount and J. P. Merlie. BIP associates with newly synthesized subunits of the mouse muscle nicotinic receptor. Journal of Cell Biology, 113(5):1125–??, June 1991. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/113/5/
1125.

Brugg:1991:PDB

[BM91b] B. Brugg and A. Matus. Phosphorylation determines the binding of microtubule-associated protein 2 (MAP2) to microtubules in living cells. Journal of Cell Biology, 114(4): 735–??, August 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/114/4/735.

Busch:1992:TRA

[BMB^{+}92] S. J. Busch, G. A. Martin, R. L. Barnhart, M. Mano, A. D. Cardin, and R. L. Jackson. Trans-repressor activity of nuclear glycosaminoglycans on Fos and Jun/AP-1 oncoprotein-mediated transcription. Journal of Cell Biology, 116(1):31–??, January 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/116/1/31.

Bader:1991:IFF

[BMF^{+}91] B. L. Bader, T. M. Magin, M. Freudenmann, S. Stumpp, and W. W. Franke. Intermediate filaments formed de novo from tail-less cytokeratins in the cytoplasm and in the nucleus. Journal of Cell Biology, 115(5):1293–??, December 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/115/5/1293.

Buck:1991:DAM

[BMGC91] J. Buck, A. Myc, A. Garbe, and G. Cathomas. Differences in the action and metabolism between retinol and retinoic acid in B lymphocytes. Journal of Cell Biology, 115(3): 851–??, November 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/115/3/851.

Bowser:1992:SSC

[BMGN92] R. Bowser, H. Müller, B. Govindan, and P. Novick. Sec8p and Sec15p are components of a plasma membraneassociated 19.5S particle that may function downstream of Sec4p to control exocytosis. Journal of Cell Biology, 118 (5):1041–??, September 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/118/5/1041.

Brown:1993:CSI

[BMH^{+}93] C. R. Brown, R. L. Martin, W. J. Hansen, R. P. Beckmann, and W. J. Welch. The constitutive and stress inducible forms of hsp 70 exhibit functional similarities and interact with one another in an ATP-dependent fashion. Journal of Cell Biology, 120(5):1101–??, March 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/120/5/1101.

Bellot:1990:HAE

[BMK^{+}90] F. Bellot, W. Moolenaar, R. Kris, B. Mirakhur, I. Verlaan, A. Ullrich, J. Schlessinger, and S. Felder. High-affinity epidermal growth factor binding is specifically reduced by a monoclonal antibody, and appears necessary for early responses. Journal of Cell Biology, 110(2):491–??, February 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/110/2/491.

Bandtlow:1990:RCC

[BML^{+}90] C. E. Bandtlow, M. Meyer, D. Lindholm, M. Spranger, R. Heumann, and H. Thoenen. Regional and cellular codistribution of interleukin 1 beta and nerve growth factor mRNA in the adult rat brain: possible relationship to the regulation of nerve growth factor synthesis. Journal of Cell Biology, 111(4):1701–??, October 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/111/4/1701.

Busso:1994:ICM

[BML^{+}94] N. Busso, S. K. Masur, D. Lazega, S. Waxman, and L. Ossowski. Induction of cell migration by pro-urokinase binding to its receptor: possible mechanism for signal transduction in human epithelial cells. Journal of Cell Biology, 126 (1):259–??, July 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/126/1/259.

Breitfeld:1990:ENV

[BMM90] P. P. Breitfeld, W. C. McKinnon, and K. E. Mostov. Effect of nocodazole on vesicular traffic to the apical and basolateral surfaces of polarized MDCK cells. Journal of Cell Biology, 111(6):2365–??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/111/6/2365.

Borner:1994:PBA

[BMM^{+}94] C. Borner, I. Martinou, C. Mattmann, M. Irmler, E. Schaerer, J. C. Martinou, and J. Tschopp. The protein bcl-2 alpha does not require membrane attachment, but two conserved domains to suppress apoptosis. Journal of Cell Biology, 126(4):1059–??, August 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/126/4/1059.

Blobel:1990:PPP

[BMPW90] C. P. Blobel, D. G. Myles, P. Primakoff, and J. M. White. Proteolytic processing of a protein involved in sperm-egg fusion correlates with acquisition of fertilization competence. Journal of Cell Biology, 111(1):69–??, July 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/111/1/69.

Bremer:1991:SBI

[BMS^{+}91] A. Bremer, R. C. Millonig, R. Sütterlin, A. Engel, T. D. Pollard, and U. Aebi. The structural basis for the intrinsic disorder of the actin filament: the “lateral slipping” model. Journal of Cell Biology, 115(3):689–??, November 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/115/3/
689.

Baorto:1992:APG

[BMS92] D. M. Baorto, W. Mellado, and M. L. Shelanski. Astrocyte process growth induction by actin breakdown. Journal of Cell Biology, 117(2):357–??, April 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/117/2/357.

Barth:1994:RPA

[BMTTG94] A. Barth, A. Müller-Taubenberger, P. Taranto, and G. Gerisch. Replacement of the phospholipid-anchor in the contact site A glycoprotein of D. discoideum by a transmembrane region does not impede cell adhesion but reduces residence time on the cell surface. Journal of Cell Biology, 124(1):205–??, January 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/124/1/205.

Bookbinder:1990:ISU

[BMV90] L. H. Bookbinder, G. W. Moy, and V. D. Vacquier. Identification of sea urchin sperm adenylate cyclase. Journal of Cell Biology, 111(5):1859–??, November 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/111/5/1859.

Borasio:1993:IRP

[BMW^{+}93] G. D. Borasio, A. Markus, A. Wittinghofer, Y. A. Barde, and R. Heumann. Involvement of ras p21 in neurotrophininduced response of sensory, but not sympathetic neurons. Journal of Cell Biology, 121(3):665–??, May 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/121/3/665.

Brown:1993:SBP

[BMWT93] J. C. Brown, K. Mann, H. Wiedemann, and R. Timpl. Structure and binding properties of collagen type XIV isolated from human placenta. Journal of Cell Biology, 120(2): 557–??, January 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/120/2/557.

Bowser:1991:SPE

[BN91] R. Bowser and P. Novick. Sec15 protein, an essential component of the exocytotic apparatus, is associated with the plasma membrane and with a soluble 19.5S particle. Journal of Cell Biology, 112(6):1117–??, March 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/112/6/1117.

Blencowe:1994:ANM

[BNI^{+}94] B. J. Blencowe, J. A. Nickerson, R. Issner, S. Penman, and P. A. Sharp. Association of nuclear matrix antigens with exon-containing splicing complexes. Journal of Cell Biology, 127(3):593–??, November 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/127/3/593.

Baribault:1991:PFE

[BO91] H. Baribault and R. G. Oshima. Polarized and functional epithelia can form after the targeted inactivation of both mouse keratin 8 alleles. Journal of Cell Biology, 115(6): 1675–??, December 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/115/6/1675.

Becker:1992:ADP

[BOH^{+}92] K. D. Becker, P. T. O’Donnell, J. M. Heitz, M. Vito, and S. I. Bernstein. Analysis of Drosophila paramyosin: identification of a novel isoform which is restricted to a subset of adult muscles. Journal of Cell Biology, 116(3):669–??, February 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/116/3/669.

Block:1994:CSR

[BOM94] B. A. Block, J. O’Brien, and G. Meissner. Characterization of the sarcoplasmic reticulum proteins in the thermogenic muscles of fish. Journal of Cell Biology, 127(5):1275–??, December 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/127/5/1275.

Bachhawat:1991:DIF

[BP91] A. K. Bachhawat and S. Pillai. Distinct intracellular fates of membrane and secretory immunoglobulin heavy chains in a pre–B cell line. Journal of Cell Biology, 115(3):619– ??, November 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/115/3/619.

Baker:1993:TPA

[BP93] L. P. Baker and H. B. Peng. Tyrosine phosphorylation and acetylcholine receptor cluster formation in cultured Xenopus muscle cells. Journal of Cell Biology, 120(1):185–??, January 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/120/1/185.

Barsony:1990:IMF

[BPDM90] J. Barsony, J. W. Pike, H. F. DeLuca, and S. J. Marx. Immunocytology with microwave-fixed fibroblasts shows 1 alpha,25-dihydroxyvitamin D3-dependent rapid and estrogen-dependent slow reorganization of vitamin D receptors. Journal of Cell Biology, 111(6):2385–??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/111/6/2385.

Bre:1990:RMD

[BPH^{+}90] M. H. Bré, R. Pepperkok, A. M. Hill, N. Levilliers, W. Ansorge, E. H. Stelzer, and E. Karsenti. Regulation of microtubule dynamics and nucleation during polarization in MDCK II cells. Journal of Cell Biology, 111(6):3013–??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/111/6/3013.

Baas:1991:PIT

[BPK91] P. W. Baas, T. P. Pienkowski, and K. S. Kosik. Processes induced by tau expression in Sf9 cells have an axon-like microtubule organization. Journal of Cell Biology, 115(5):1333– ??, December 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/115/5/1333.

Bosenberg:1993:ARM

[BPM93] M. W. Bosenberg, A. Pandiella, and J. Massagué. Activated release of membrane-anchored TGF-alpha in the absence of cytosol. Journal of Cell Biology, 122(1):95–??, July 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/122/1/
95.

Brandli:1990:TMC

[BPS90] A. W. Brändli, R. G. Parton, and K. Simons. Transcytosis in MDCK cells: identification of glycoproteins transported bidirectionally between both plasma membrane domains. Journal of Cell Biology, 111(6):2909–??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/111/6/
2909.

Bassell:1994:SMV

[BPTS94] G. J. Bassell, C. M. Powers, K. L. Taneja, and R. H. Singer. Single mRNAs visualized by ultrastructural in situ hybridization are principally localized at actin filament intersections in fibroblasts. Journal of Cell Biology, 126(4): 863–??, August 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/126/4/863.

Birch:1992:CCT

[BPZ^{+}92] K. A. Birch, J. S. Pober, G. B. Zavoico, A. R. Means, and B. M. Ewenstein. Calcium/ calmodulin transduces thrombin-stimulated secretion: studies in intact and minimally permeabilized human umbilical vein endothelial cells. Journal of Cell Biology, 118(6):1501–??, September 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/118/6/
1501.

Brewer:1991:SAA

[BR91] C. B. Brewer and M. G. Roth. A single amino acid change in the cytoplasmic domain alters the polarized delivery of influenza virus hemagglutinin. Journal of Cell Biology, 114 (3):413–??, August 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/114/3/413.

Ballock:1994:TSF

[BR94a] R. T. Ballock and A. H. Reddi. Thyroxine is the serum factor that regulates morphogenesis of columnar cartilage from isolated chondrocytes in chemically defined medium. Journal of Cell Biology, 126(5):1311–??, September 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/126/5/
1311.

Biggins:1994:DIB

[BR94b] S. Biggins and M. D. Rose. Direct interaction between yeast spindle pole body components: Kar1p is required for Cdc31p localization to the spindle pole body. Journal of Cell Biology, 125(4):843–??, May 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/125/4/843.

Balzac:1994:EBI

[BRA^{+}94] F. Balzac, S. F. Retta, A. Albini, A. Melchiorri, V. E. Koteliansky, M. Geuna, L. Silengo, and G. Tarone. Expression of beta 1B integrin isoform in CHO cells results in a dominant negative effect on cell adhesion and motility. Journal of Cell Biology, 127(2):557–??, October 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/127/2/557.

Bayly:1994:SLC

[BRD94] A. C. Bayly, R. A. Roberts, and C. Dive. Suppression of liver cell apoptosis in vitro by the non-genotoxic hepatocarcinogen and peroxisome proliferator nafenopin. Journal of Cell Biology, 125(1):197–??, April 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/125/1/197.

Blessing:1993:ESE

[BRF93] M. Blessing, U. Rüther, and W. W. Franke. Ectopic synthesis of epidermal cytokeratins in pancreatic islet cells of transgenic mice interferes with cytoskeletal order and insulin production. Journal of Cell Biology, 120(3):743–??, February 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/120/3/743.

Baker:1990:CTD

[BRG^{+}90] H. N. Baker, S. W. Rothwell, W. A. Grasser, K. T. Wallis, and D. B. Murphy. Copolymerization of two distinct tubulin isotypes during microtubule assembly in vitro. Journal of Cell Biology, 110(1):97–??, January 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/110/1/97.

Battegay:1994:PBM

[BRIA^{+}94] E. J. Battegay, J. Rupp, L. Iruela-Arispe, E. H. Sage, and M. Pech. PDGF-BB modulates endothelial proliferation and angiogenesis in vitro via PDGF beta-receptors. Journal of Cell Biology, 125(4):917–??, May 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/125/4/917.

Bloom:1993:GGI

[BRL^{+}93] G. S. Bloom, B. W. Richards, P. L. Leopold, D. M. Ritchey, and S. T. Brady. GTP gamma S inhibits organelle transport along axonal microtubules. Journal of Cell Biology, 120(2): 467–??, January 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/120/2/467.

Bloch:1991:CCA

[BRO^{+}91a] R. J. Bloch, W. G. Resneck, A. O’Neill, J. Strong, and D. W. Pumplin. Cytoplasmic components of acetylcholine receptor clusters of cultured rat myotubes: the 58-kD protein. Journal of Cell Biology, 115(2):435–??, October 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/115/2/
435.

Brokaw:1991:MSS

[Bro91b] C. J. Brokaw. Microtubule sliding in swimming sperm flagella: direct and indirect measurements on sea urchin and tunicate spermatozoa. Journal of Cell Biology, 114(6):1201– ??, September 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/114/6/1201.

Berg:1991:HPL

[BRWB91] E. L. Berg, M. K. Robinson, R. A. Warnock, and E. C. Butcher. The human peripheral lymph node vascular addressin is a ligand for LECAM-1, the peripheral lymph node homing receptor. Journal of Cell Biology, 114(2):343–??, July 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/114/2/343.

Begovac:1990:CSG

[BS90] P. C. Begovac and B. D. Shur. Cell surface galactosyltransferase mediates the initiation of neurite outgrowth from PC12 cells on laminin. Journal of Cell Biology, 110(2): 461–??, February 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/110/2/461.

Bonen:1991:EEC

[BS91] D. K. Bonen and T. M. Schmid. Elevated extracellular calcium concentrations induce type X collagen synthesis in chondrocyte cultures. Journal of Cell Biology, 115(4):1171– ??, November 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/115/4/1171.

Bolin:1993:NRS

[BS93a] L. M. Bolin and E. M. Shooter. Neurons regulate Schwann cell genes by diffusible molecules. Journal of Cell Biology, 123(1):237–??, October 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/123/1/237.

Brodsky:1993:SBC

[BS93b] J. L. Brodsky and R. Schekman. A Sec63p–BiP complex from yeast is required for protein translocation in a reconstituted proteoliposome. Journal of Cell Biology, 123(6): 1355–??, December 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/123/6/1355.

Barroso:1994:BAT

[BS94a] M. Barroso and E. S. Sztul. Basolateral to apical transcytosis in polarized cells is indirect and involves BFA and trimeric G protein sensitive passage through the apical endosome. Journal of Cell Biology, 124(1):83–??, January 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/124/1/83.

Bloodgood:1994:TSP

[BS94b] R. A. Bloodgood and N. L. Salomonsky. The transmembrane signaling pathway involved in directed movements of Chlamydomonas flagellar membrane glycoproteins involves the dephosphorylation of a 60-kD phosphoprotein that binds to the major flagellar membrane glycoprotein. Journal of Cell Biology, 127(3):803–??, November 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/127/3/
803.

Brancolini:1994:PGA

[BS94c] C. Brancolini and C. Schneider. Phosphorylation of the growth arrest-specific protein Gas2 is coupled to actin rearrangements during Go → G1 transition in NIH 3T3 cells. Journal of Cell Biology, 124(5):743–??, March 1994. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/124/5/

Brown:1992:NAM

[BSB92] A. Brown, T. Slaughter, and M. M. Black. Newly assembled microtubules are concentrated in the proximal and distal regions of growing axons. Journal of Cell Biology, 119(4): 867–??, November 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/119/4/867.

Blattner:1992:GAT

[BSD^{+}92] J. Blattner, B. Swinkels, H. Dörsam, T. Prospero, S. Subramani, and C. Clayton. Glycosome assembly in trypanosomes: variations in the acceptable degeneracy of a COOH-terminal microbody targeting signal. Journal of Cell Biology, 119(5):1129–??, December 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/119/5/1129.

Brockerhoff:1994:UMM

[BSD94] S. E. Brockerhoff, R. C. Stevens, and T. N. Davis. The unconventional myosin, Myo2p, is a calmodulin target at sites of cell growth in Saccharomyces cerevisiae. Journal of Cell Biology, 124(3):315–??, February 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/124/3/315.

Bar-Shavit:1993:TAP

[BSEF^{+}93] R. Bar-Shavit, Y. Eskohjido, J. W. Fenton, J. D. Esko, and I. Vlodavsky. Thrombin adhesive properties: induction by plasmin and heparan sulfate. Journal of Cell Biology, 123(5):1279–??, December 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/123/5/1279.

Berlin:1990:BPR

[BSF90] V. Berlin, C. A. Styles, and G. R. Fink. BIK1, a protein required for microtubule function during mating and mitosis in Saccharomyces cerevisiae, colocalizes with tubulin. Journal of Cell Biology, 111(6):2573–??, December 1990. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/111/6/

Bauer:1992:MFR

[BSG^{+}92] J. S. Bauer, C. L. Schreiner, F. G. Giancotti, E. Ruoslahti, and R. L. Juliano. Motility of fibronectin receptor-deficient cells on fibronectin and vitronectin: collaborative interactions among integrins. Journal of Cell Biology, 116(2): 477–??, January 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/116/2/477.

Boone:1990:YKG

[BSHB90] C. Boone, S. S. Sommer, A. Hensel, and H. Bussey. Yeast KRE genes provide evidence for a pathway of cell wall betaglucan assembly. Journal of Cell Biology, 110(5):1833–??, May 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/110/5/1833.

Backer:1991:IRI

[BSHW91] J. M. Backer, S. E. Shoelson, E. Haring, and M. F. White. Insulin receptors internalize by a rapid, saturable pathway requiring receptor autophosphorylation and an intact juxtamembrane region. Journal of Cell Biology, 115(6):1535– ??, December 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/115/6/1535.

Blount:1990:AIM

[BSM90] P. Blount, M. M. Smith, and J. P. Merlie. Assembly intermediates of the mouse muscle nicotinic acetylcholine receptor in stably transfected fibroblasts. Journal of Cell Biology, 111(6):2601–??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/111/6/2601.

Beggs:1994:NDN

[BSM94] H. E. Beggs, P. Soriano, and P. F. Maness. NCAMdependent neurite outgrowth is inhibited in neurons from Fyn-minus mice. Journal of Cell Biology, 127(3):825–??, November 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/127/3/825.

Byrd:1994:TLC

[BSP^{+}94] D. A. Byrd, D. J. Sweet, N. Panté, K. N. Konstantinov, T. Guan, A. C. Saphire, P. J. Mitchell, C. S. Cooper, U. Aebi, and L. Gerace. Tpr, a large coiled coil protein whose amino terminus is involved in activation of oncogenic kinases, is localized to the cytoplasmic surface of the nuclear pore complex. Journal of Cell Biology, 127(6):1515– ??, December 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/127/6/1515.

Burgess:1990:PDH

[BSR^{+}90] W. H. Burgess, A. M. Shaheen, M. Ravera, M. Jaye, P. J. Donohue, and J. A. Winkles. Possible dissociation of the heparin-binding and mitogenic activities of heparin-binding (acidic fibroblast) growth factor-1 from its receptor-binding activities by site-directed mutagenesis of a single lysine residue. Journal of Cell Biology, 111(5):2129–??, November 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/111/5/2129.

Bar-Shavit:1991:AGA

[BSSL^{+}91] R. Bar-Shavit, V. Sabbah, M. G. Lampugnani, P. C. Marchisio, J. W. Fenton, I. Vlodavsky, and E. Dejana. An Arg–Gly–Asp sequence within thrombin promotes endothelial cell adhesion. Journal of Cell Biology, 112(2):335–??, January 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/112/2/335.

Borchelt:1990:SCP

[BST^{+}90] D. R. Borchelt, M. Scott, A. Taraboulos, N. Stahl, and S. B. Prusiner. Scrapie and cellular prion proteins differ in their kinetics of synthesis and topology in cultured cells. Journal of Cell Biology, 110(3):743–??, March 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/110/3/743.

Billger:1991:MAP

[BSW91] M. Billger, E. Strömberg, and M. Wallin. Microtubuleassociated proteins-dependent colchicine stability of acetylated cold-labile brain microtubules from the Atlantic cod, Gadus morhua. Journal of Cell Biology, 113(2):331–??, April 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/113/2/331.

Backer:1992:IRJ

[BSW^{+}92] J. M. Backer, S. E. Shoelson, M. A. Weiss, Q. X. Hua, R. B. Cheatham, E. Haring, D. C. Cahill, and M. F. White. The insulin receptor juxtamembrane region contains two independent tyrosine/beta-turn internalization signals. Journal of Cell Biology, 118(4):831–??, August 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/118/4/831.

Banerjee:1992:HBP

[BT92] S. D. Banerjee and B. P. Toole. Hyaluronan-binding protein in endothelial cell morphogenesis. Journal of Cell Biology, 119(3):643–??, November 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/119/3/643.

Boyer:1993:CAD

[BT93] B. Boyer and J. P. Thiery. Cyclic AMP distinguishes between two functions of acidic FGF in a rat bladder carcinoma cell line. Journal of Cell Biology, 120(3):767–??, February 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/120/3/767.

Baba:1994:UAA

[BTBO94] M. Baba, K. Takeshige, N. Baba, and Y. Ohsumi. Ultrastructural analysis of the autophagic process in yeast: detection of autophagosomes and their characterization. Journal of Cell Biology, 124(6):903–??, March 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/124/6/903.

Bristow:1993:TXN

[BTG^{+}93] J. Bristow, M. K. Tee, S. E. Gitelman, S. H. Mellon, and W. L. Miller. Tenascin-x: a novel extracellular matrix protein encoded by the human XB gene overlapping P450c21B. Journal of Cell Biology, 122(1):265–??, July 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/122/1/265.

Baumert:1990:PNT

[BTH^{+}90] M. Baumert, K. Takei, J. Hartinger, P. M. Burger, G. Fischer von Mollard, P. R. Maycox, P. De Camilli, and R. Jahn. P29: a novel tyrosine-phosphorylated membrane protein present in small clear vesicles of neurons and endocrine cells. Journal of Cell Biology, 110(4):1285–??, April 1990. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/110/4/
1285.

Burridge:1992:TPP

[BTR92] K. Burridge, C. E. Turner, and L. H. Romer. Tyrosine phosphorylation of paxillin and pp125FAK accompanies cell adhesion to extracellular matrix: a role in cytoskeletal assembly. Journal of Cell Biology, 119(4):893–??, November 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/119/4/
893.

Bellosta:1993:CKF

[BTRB93] P. Bellosta, D. Talarico, D. Rogers, and C. Basilico. Cleavage of K–FGF produces a truncated molecule with increased biological activity and receptor binding affinity. Journal of Cell Biology, 121(3):705–??, May 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/121/3/705.

Baker:1993:DCM

[BTS93] J. Baker, W. E. Theurkauf, and G. Schubiger. Dynamic changes in microtubule configuration correlate with nuclear migration in the preblastoderm Drosophila embryo. Journal of Cell Biology, 122(1):113–??, July 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/122/1/113.

Balsamo:1991:ARA

[BTSL91] J. Balsamo, R. Thiboldeaux, N. Swaminathan, and J. Lilien. Antibodies to the retina N-acetylgalactosaminylphosphotransferase modulate N-cadherin-mediated adhesion and uncouple the N-cadherin transferase complex from the actin-containing cytoskeleton. Journal of Cell Biology, 113(2):429–??, April 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/
113/2/429.

Beck:1994:LVS

[BU94] C. Beck and R. Uhl. On the localization of voltage-sensitive calcium channels in the flagella of Chlamydomonas reinhardtii. Journal of Cell Biology, 125(5):1119–??, June 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/125/5/
1119.

Behrens:1993:LED

[BVF^{+}93] J. Behrens, L. Vakaet, R. Friis, E. Winterhager, F. Van Roy, M. M. Mareel, and W. Birchmeier. Loss of epithelial differentiation and gain of invasiveness correlates with tyrosine phosphorylation of the E-cadherin/beta-catenin complex in cells transformed with a temperature-sensitive v–SRC gene. Journal of Cell Biology, 120(3):757–??, February 1993. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/120/3/
757.

Benfenati:1993:ISP

[BVR^{+}93] F. Benfenati, F. Valtorta, M. C. Rossi, F. Onofri, T. Sihra, and P. Greengard. Interactions of synapsin I with phospholipids: possible role in synaptic vesicle clustering and in the maintenance of bilayer structures. Journal of Cell Biology, 123(6):1845–??, December 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/123/6/1845.

Bauer:1993:FRC

[BVS^{+}93] J. S. Bauer, J. Varner, C. Schreiner, L. Kornberg, R. Nicholas, and R. L. Juliano. Functional role of the cytoplasmic domain of the integrin alpha 5 subunit. Journal of Cell Biology, 122(1):209–??, July 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/122/1/209.

Barstead:1991:VEM

[BW91] R. J. Barstead and R. H. Waterston. Vinculin is essential for muscle function in the nematode. Journal of Cell Biology, 114(4):715–??, August 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/114/4/715.

Burleigh:1993:IMG

[BWCG93] B. A. Burleigh, C. W. Wells, M. W. Clarke, and P. R. Gardiner. An integral membrane glycoprotein associated with an endocytic compartment of Trypanosoma vivax: identification and partial characterization. Journal of Cell Biology, 120(2):339–??, January 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/120/2/339.

Bahler:1993:UNS

[BWLK93] J. Bähler, T. Wyler, J. Loidl, and J. Kohli. Unusual nuclear structures in meiotic prophase of fission yeast: a cytological analysis. Journal of Cell Biology, 121(2):241–??, April 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/121/2/
241.

Baler:1992:HSG

[BWV92] R. Baler, W. J. Welch, and R. Voellmy. Heat shock gene regulation by nascent polypeptides and denatured proteins: hsp70 as a potential autoregulatory factor. Journal of Cell Biology, 117(6):1151–??, June 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/117/6/1151.

Bixby:1990:PCP

[BZ90] J. L. Bixby and R. Zhang. Purified N-cadherin is a potent substrate for the rapid induction of neurite outgrowth. Journal of Cell Biology, 110(4):1253–??, April 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/110/4/1253.

Bubien:1993:TCC

[BZB^{+}93] J. K. Bubien, L. J. Zhou, P. D. Bell, R. A. Frizzell, and T. F. Tedder. Transfection of the CD20 cell surface molecule into ectopic cell types generates a Ca^{2+} conductance found constitutively in B lymphocytes. Journal of Cell Biology, 121 (5):1121–??, June 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/121/5/1121.

Bast:1992:HCC

[BZF^{+}92] B. J. Bast, L. J. Zhou, G. J. Freeman, K. J. Colley, T. J. Ernst, J. M. Munro, and T. F. Tedder. The HB-6, CDw75, and CD76 differentiation antigens are unique cellsurface carbohydrate determinants generated by the betagalactoside alpha 2,6-sialyltransferase. Journal of Cell Biology, 116(2):423–??, January 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/116/2/423.

Belmont:1993:LBD

[BZT93] A. S. Belmont, Y. Zhai, and A. Thilenius. Lamin B distribution and association with peripheral chromatin revealed by optical sectioning and electron microscopy tomography. Journal of Cell Biology, 123(6):1671–??, December 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/123/6/
1671.

Benton:1994:NFR

[BZT94] B. M. Benton, J. H. Zang, and J. Thorner. A novel FK506and rapamycin-binding protein (FPR3 gene product) in the yeast Saccharomyces cerevisiae is a proline rotamase localized to the nucleolus. Journal of Cell Biology, 127(3):623– ??, November 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/127/3/623.

Caras:1991:IPS

[Car91] I. W. Caras. An internally positioned signal can direct attachment of a glycophospholipid membrane anchor. Journal of Cell Biology, 113(1):77–??, April 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/113/1/77.

Christgau:1992:MAA

[CAS^{+}92] S. Christgau, H. J. Aanstoot, H. Schierbeck, K. Begley, S. Tullin, K. Hejnaes, and S. Baekkeskov. Membrane anchoring of the autoantigen GAD65 to microvesicles in pancreatic beta-cells by palmitoylation in the NH2-terminal domain. Journal of Cell Biology, 118(2):309–??, July 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/118/2/
309.

Cooper:1990:EEM

[CB90] D. N. Cooper and S. H. Barondes. Evidence for export of a muscle lectin from cytosol to extracellular matrix and for a novel secretory mechanism. Journal of Cell Biology, 110 (5):1681–??, May 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/110/5/1681.

Cooper:1992:YKG

[CB92] A. Cooper and H. Bussey. Yeast Kex1p is a Golgi-associated membrane protein: deletions in a cytoplasmic targeting domain result in mislocalization to the vacuolar membrane. Journal of Cell Biology, 119(6):1459–??, December 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/119/6/
1459.

Cohen:1994:DDM

[CBC^{+}94] C. J. Cohen, R. Bacon, M. Clarke, K. Joiner, and I. Mellman. Dictyostelium discoideum mutants with conditional defects in phagocytosis. Journal of Cell Biology, 126(4): 955–??, August 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/126/4/955.

Cooke:1990:CBM

[CBE90] C. A. Cooke, R. L. Bernat, and W. C. Earnshaw. CENP-B: a major human centromere protein located beneath the kinetochore. Journal of Cell Biology, 110(5):1475–??, May 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/110/5/1475.

Cooke:1993:MDW

[CBJER93] C. A. Cooke, D. P. Bazett-Jones, W. C. Earnshaw, and J. B. Rattner. Mapping DNA within the mammalian kinetochore. Journal of Cell Biology, 120(5):1083–??, March 1993. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/120/5/
1083.

Cripps:1994:TDM

[CBM^{+}94] R. M. Cripps, K. D. Becker, M. Mardahl, W. A. Kronert, D. Hodges, and S. I. Bernstein. Transformation of Drosophila melanogaster with the wild-type myosin heavychain gene: rescue of mutant phenotypes and analysis of defects caused by overexpression. Journal of Cell Biology, 126(3):689–??, August 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/126/3/689.

Cao:1992:EPP

[CBRW92] L. G. Cao, G. G. Babcock, P. A. Rubenstein, and Y. L. Wang. Effects of profilin and profilactin on actin structure and function in living cells. Journal of Cell Biology, 117(5): 1023–??, June 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/117/5/1023.

Ceriotti:1990:TFD

[CC90a] A. Ceriotti and A. Colman. Trimer formation determines the rate of influenza virus haemagglutinin transport in the early stages of secretion in Xenopus oocytes. Journal of Cell Biology, 111(2):409–??, August 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/111/2/409.

Cutler:1990:SDT

[CC90b] D. F. Cutler and L. P. Cramer. Sorting during transport to the surface of PC12 cells: divergence of synaptic vesicle and secretory granule proteins. Journal of Cell Biology, 110 (3):721–??, March 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/110/3/721.

Chaudhary:1993:SRN

[CC93a] N. Chaudhary and J. C. Courvalin. Stepwise reassembly of the nuclear envelope at the end of mitosis. Journal of Cell Biology, 122(2):295–??, July 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/122/2/295.

Compton:1993:NRP

[CC93b] D. A. Compton and D. W. Cleveland. NuMA is required for the proper completion of mitosis. Journal of Cell Biology, 120(4):947–??, February 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/120/4/947.

Coulombe:1990:DEK

[CCAF90] P. A. Coulombe, Y. M. Chan, K. Albers, and E. Fuchs. Deletions in epidermal keratins leading to alterations in filament organization in vivo and in intermediate filament assembly in vitro. Journal of Cell Biology, 111(6):3049– ??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/111/6/3049.

Corvera:1994:DLWa

[CCC^{+}94a] S. Corvera, A. Chawla, R. Chakrabarti, M. Joly, J. Buxton, and M. P. Czech. A double leucine within the GLUT4 glucose transporter COOH-terminal domain functions as an endocytosis signal. Journal of Cell Biology, 126(4):979–??, August 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/126/4/979.

Corvera:1994:DLWb

[CCC^{+}94b] S. Corvera, A. Chawla, R. Chakrabarti, M. Joly, J. Buxton, and M. P. Czech. A double leucine within the GLUT4 glucose transporter COOH-terminal domain functions as an endocytosis signal. Journal of Cell Biology, 126(6):1625– ??, September 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/126/6/1625.

Cano:1992:MRF

[CCFZ92] M. L. Cano, L. Cassimeris, M. Fechheimer, and S. H. Zigmond. Mechanisms responsible for F-actin stabilization after lysis of polymorphonuclear leukocytes. Journal of Cell Biology, 116(5):1123–??, March 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/116/5/1123.

Chenu:1994:OIC

[CCG^{+}94] C. Chenu, S. Colucci, M. Grano, P. Zigrino, R. Barattolo, G. Zambonin, N. Baldini, P. Vergnaud, P. D. Delmas, and A. Z. Zallone. Osteocalcin induces chemotaxis, secretion of matrix proteins, and calcium-mediated intracellular signaling in human osteoclast-like cells. Journal of Cell Biology, 127(4):1149–??, November 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/127/4/1149.

Chen:1993:YAM

[CCR93] X. Chen, R. K. Cook, and P. A. Rubenstein. Yeast actin with a mutation in the “hydrophobic plug” between subdomains 3 and 4 (L266D) displays a cold-sensitive polymerization defect. Journal of Cell Biology, 123(5):1185–??, December 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/123/5/1185.

Crescenzi:1994:TMP

[CCT94] M. Crescenzi, D. H. Crouch, and F. Tatò. Transformation by myc prevents fusion but not biochemical differentiation of C2C12 myoblasts: mechanisms of phenotypic correction in mixed culture with normal cells. Journal of Cell Biology, 125(5):1137–??, June 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/125/5/1137.

Cox:1992:TDA

[CCW^{+}92] D. Cox, J. Condeelis, D. Wessels, D. Soll, H. Kern, and D. A. Knecht. Targeted disruption of the ABP-120 gene leads to cells with altered motility. Journal of Cell Biology, 116(4): 943–??, February 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/116/4/943.

Czech:1993:EET

[CCW^{+}93] M. P. Czech, A. Chawla, C. W. Woon, J. Buxton, M. Armoni, W. Tang, M. Joly, and S. Corvera. Exofacial epitopetagged glucose transporter chimeras reveal COOH-terminal sequences governing cellular localization. Journal of Cell Biology, 123(1):127–??, October 1993. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/123/1/127.

Caplan:1991:CYY

[CD91] A. J. Caplan and M. G. Douglas. Characterization of YDJ1: a yeast homologue of the bacterial dnaJ protein. Journal of Cell Biology, 114(4):609–??, August 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/114/4/609.

Cooper:1993:DIT

[CdJC93] R. Cooper, A. R. de Jesus, and G. A. Cross. Deletion of an immunodominant Trypanosoma cruzi surface glycoprotein disrupts flagellum-cell adhesion. Journal of Cell Biology, 122(1):149–??, July 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/122/1/149.

Campanero:1993:IIL

[CdPA^{+}93] M. R. Campanero, M. A. del Pozo, A. G. Arroyo, P. Sánchez-Mateos, T. Hernández-Caselles, A. Craig, R. Pulido, and F. Sánchez-Madrid. ICAM-3 interacts with LFA-1 and regulates the LFA-1/ICAM-1 cell adhesion pathway. Journal of Cell Biology, 123(4):1007–??, November 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/123/4/1007.

Coverley:1993:REN

[CDRL93] D. Coverley, C. S. Downes, P. Romanowski, and R. A. Laskey. Reversible effects of nuclear membrane permeabilization on DNA replication: evidence for a positive licensing factor. Journal of Cell Biology, 122(5):985–??, September 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/122/5/985.

Carpen:1991:MUL

[CDS^{+}91] O. Carpén, M. L. Dustin, T. A. Springer, J. A. Swafford, L. A. Beckett, and J. P. Caulfield. Motility and ultrastructure of large granular lymphocytes on lipid bilayers reconstituted with adhesion receptors LFA-1, ICAM-1, and two isoforms of LFA-3. Journal of Cell Biology, 115(3):861– ??, November 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/115/3/861.

Chung:1994:CSA

[CE94] C. Y. Chung and H. P. Erickson. Cell surface annexin II is a high affinity receptor for the alternatively spliced segment of tenascin–C. Journal of Cell Biology, 126(2):539– ??, July 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/126/2/539.

Carey:1992:MCC

[CES^{+}92] D. J. Carey, D. M. Evans, R. C. Stahl, V. K. Asundi, K. J. Conner, P. Garbes, and G. Cizmeci-Smith. Molecular cloning and characterization of N-syndecan, a novel transmembrane heparan sulfate proteoglycan. Journal of Cell Biology, 117(1):191–??, April 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/117/1/191.

Chiquet-Ehrismann:1994:TCE

[CETK^{+}94] R. Chiquet-Ehrismann, M. Tannheimer, M. Koch, A. Brunner, J. Spring, D. Martin, S. Baumgartner, and M. Chiquet. Tenascin-C expression by fibroblasts is elevated in stressed collagen gels. Journal of Cell Biology, 127(6):2093–??, December 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/127/6/2093.

Coulombe:1990:EES

[CF90] P. A. Coulombe and E. Fuchs. Elucidating the early stages of keratin filament assembly. Journal of Cell Biology, 111 (1):153–??, July 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/111/1/153.

Cunningham:1994:CDG

[CF94] K. W. Cunningham and G. R. Fink. Calcineurin-dependent growth control in Saccharomyces cerevisiae mutants lacking PMC1, a homolog of plasma membrane Ca^{2+} ATPases.

Journal of Cell Biology, 124(3):351–??, February 1994. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/124/3/
351.

Carmo-Fonseca:1993:ASC

[CFFL93] M. Carmo-Fonseca, J. Ferreira, and A. I. Lamond. Assembly of snRNP-containing coiled bodies is regulated in interphase and mitosis–evidence that the coiled body is a kinetic nuclear structure. Journal of Cell Biology, 120(4): 841–??, February 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/120/4/841.

Connolly:1994:TGC

[CFG^{+}94] C. N. Connolly, C. E. Futter, A. Gibson, C. R. Hopkins, and D. F. Cutler. Transport into and out of the Golgi complex studied by transfecting cells with cDNAs encoding horseradish peroxidase. Journal of Cell Biology, 127(3): 641–??, November 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/127/3/641.

Carmo-Fonseca:1992:TDC

[CFPCL92] M. Carmo-Fonseca, R. Pepperkok, M. T. Carvalho, and A. I. Lamond. Transcription-dependent colocalization of the U1, U2, U4/U6, and U5 snRNPs in coiled bodies. Journal of Cell Biology, 117(1):1–??, April 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/117/1/1.

Cao:1993:LDN

[CFW93] L. G. Cao, D. J. Fishkind, and Y. L. Wang. Localization and dynamics of nonfilamentous actin in cultured cells. Journal of Cell Biology, 123(1):173–??, October 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/123/1/173.

Caroni:1990:NSI

[CG90] P. Caroni and P. Grandes. Nerve sprouting in innervated adult skeletal muscle induced by exposure to elevated levels of insulin-like growth factors. Journal of Cell Biology, 110 (4):1307–??, April 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/110/4/1307.

Collins:1991:RBE

[CG91] P. G. Collins and R. Gilmore. Ribosome binding to the endoplasmic reticulum: a 180-kD protein identified by crosslinking to membrane-bound ribosomes is not required for ribosome binding activity. Journal of Cell Biology, 114 (4):639–??, August 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/114/4/639.

Connolly:1993:GHC

[CG93] T. Connolly and R. Gilmore. GTP hydrolysis by complexes of the signal recognition particle and the signal recognition particle receptor. Journal of Cell Biology, 123(4):799–??, November 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/123/4/799.

Conrad:1993:RDA

[CGF^{+}93] P. A. Conrad, K. A. Giuliano, G. Fisher, K. Collins, P. T. Matsudaira, and D. L. Taylor. Relative distribution of actin, myosin I, and myosin II during the wound healing response of fibroblasts. Journal of Cell Biology, 120(6): 1381–??, March 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/120/6/1381.

Cancedda:1992:HCU

[CGMC92] F. Descalzi Cancedda, C. Gentili, P. Manduca, and R. Cancedda. Hypertrophic chondrocytes undergo further differentiation in culture. Journal of Cell Biology, 117(2):427–??, April 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/117/2/427.

Chen:1994:CME

[CGW94] P. Chen, K. Gupta, and A. Wells. Cell movement elicited by epidermal growth factor receptor requires kinase and autophosphorylation but is separable from mitogenesis. Journal of Cell Biology, 124(4):547–??, February 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/124/4/547.

Chanat:1991:MIS

[CH91] E. Chanat and W. B. Huttner. Milieu-induced, selective aggregation of regulated secretory proteins in the trans-Golgi network. Journal of Cell Biology, 115(6):1505–??, December 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/115/6/1505.

Chavez:1992:EFP

[CH92] R. A. Chavez and Z. W. Hall. Expression of fusion proteins of the nicotinic acetylcholine receptor from mammalian muscle identifies the membrane-spanning regions in the alpha and delta subunits. Journal of Cell Biology, 116(2): 385–??, January 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/116/2/385.

Chan:1993:MFF

[CH93] B. M. Chan and M. E. Hemler. Multiple functional forms of the integrin VLA-2 can be derived from a single alpha 2 cDNA clone: interconversion of forms induced by an antibeta 1 antibody. Journal of Cell Biology, 120(2):537–??, January 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/120/2/537.

Cassio:1991:HCL

[CHBGL91] D. Cassio, C. Hamon-Benais, M. Guérin, and O. Lecoq. Hybrid cell lines constitute a potential reservoir of polarized cells: isolation and study of highly differentiated hepatomaderived hybrid cells able to form functional bile canaliculi in vitro. Journal of Cell Biology, 115(5):1397–??, December 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/115/5/1397.

Crossman:1994:MIL

[CHG94] M. W. Crossman, S. M. Hauft, and J. I. Gordon. The mouse ileal lipid-binding protein gene: a model for studying axial patterning during gut morphogenesis. Journal of Cell Biology, 126(6):1547–??, September 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/126/6/1547.

Cheng:1992:MMT

[CHM^{+}92] T. C. Cheng, T. A. Hanley, J. Mudd, J. P. Merlie, and E. N. Olson. Mapping of myogenin transcription during embryogenesis using transgenes linked to the myogenin control region. Journal of Cell Biology, 119(6):1649–??, December 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/119/6/1649.

Coulombe:1991:FKC

[CHVF91] P. A. Coulombe, M. E. Hutton, R. Vassar, and E. Fuchs. A function for keratins and a common thread among different types of epidermolysis bullosa simplex diseases. Journal of Cell Biology, 115(6):1661–??, December 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/115/6/1661.

Conradt:1994:DFB

[CHW94] B. Conradt, A. Haas, and W. Wickner. Determination of four biochemically distinct, sequential stages during vacuole inheritance in vitro. Journal of Cell Biology, 126(1):99– ??, July 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/126/1/99.

Citi:1992:PKI

[Cit92] S. Citi. Protein kinase inhibitors prevent junction dissociation induced by low extracellular calcium in MDCK epithelial cells. Journal of Cell Biology, 117(1):169–??, April 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/117/1/
169.

Citi:1993:MOT

[Cit93] S. Citi. The molecular organization of tight junctions. Journal of Cell Biology, 121(3):485–??, May 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/121/3/485.

Clarke:1993:PRP

[CJB^{+}93] E. P. Clarke, N. Jain, A. Brickenden, I. A. Lorimer, and B. D. Sanwal. Parallel regulation of procollagen I and colligin, a collagen-binding protein and a member of the serine protease inhibitor family. Journal of Cell Biology, 121 (1):193–??, April 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/121/1/193.

Chakrabarti:1993:RCC

[CJC93] R. Chakrabarti, M. Joly, and S. Corvera. Redistribution of clathrin-coated vesicle adaptor complexes during adipocytic differentiation of 3T3-L1 cells. Journal of Cell Biology, 123 (1):79–??, October 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/123/1/79.

Chiu:1992:RTD

[CJO92] M. L. Chiu, J. C. Jones, and E. J. O’Keefe. Restricted tissue distribution of a 37-kD possible adherens junction protein. Journal of Cell Biology, 119(6):1689–??, December 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/119/6/
1689.

Cheshire:1991:UCF

[CK91] J. L. Cheshire and L. R. Keller. Uncoupling of Chlamydomonas flagellar gene expression and outgrowth from flagellar excision by manipulation of Ca^{2+}. Journal of Cell Biology, 115(6):1651–??, December 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/115/6/1651.

Cary:1994:DOD

[CK94] R. B. Cary and M. W. Klymkowsky. Differential organization of desmin and vimentin in muscle is due to differences in their head domains. Journal of Cell Biology, 126 (2):445–??, July 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/126/2/445.

Chan:1993:KAS

[CKB93] W. Chan, E. Kordeli, and V. Bennett. 440-kD ankyrinB: structure of the major developmentally regulated domain and selective localization in unmyelinated axons. Journal of Cell Biology, 123(6):1463–??, December 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/123/6/1463.

Cornil:1990:TCS

[CKD90] I. Cornil, R. S. Kerbel, and J. W. Dennis. Tumor cell surface beta 1-4-linked galactose binds to lectin(s) on microvascular endothelial cells and contributes to organ colonization. Journal of Cell Biology, 111(2):773–??, August 1990. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/111/2/
773.

Carter:1990:DFI

[CKG^{+}90] W. G. Carter, P. Kaur, S. G. Gil, P. J. Gahr, and E. A. Wayner. Distinct functions for integrins alpha 3 beta 1 in focal adhesions and alpha 6 beta 4/bullous pemphigoid antigen in a new stable anchoring contact (SAC) of keratinocytes: relation to hemidesmosomes. Journal of Cell Biology, 111(6):3141–??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/111/6/3141.

Cowman:1991:PGH

[CKGC91] A. F. Cowman, S. Karcz, D. Galatis, and J. G. Culvenor. A P-glycoprotein homologue of Plasmodium falciparum is localized on the digestive vacuole. Journal of Cell Biology, 113(5):1033–??, June 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/113/5/1033.

Corlu:1991:PMP

[CKL^{+}91] A. Corlu, B. Kneip, C. Lhadi, G. Leray, D. Glaise, G. Baffet, D. Bourel, and C. Guguen-Guillouzo. A plasma membrane protein is involved in cell contact-mediated regulation of tissue-specific genes in adult hepatocytes. Journal of Cell Biology, 115(2):505–??, October 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/115/2/505.

Cant:1994:DSF

[CKMC94] K. Cant, B. A. Knowles, M. S. Mooseker, and L. Cooley. Drosophila singed, a fascin homolog, is required for actin bundle formation during oogenesis and bristle extension. Journal of Cell Biology, 125(2):369–??, April 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/125/2/369.

Camp:1991:VCI

[CKP91] R. L. Camp, T. A. Kraus, and E. Puré. Variations in the cytoskeletal interaction and posttranslational modification of the CD44 homing receptor in macrophages. Journal of Cell Biology, 115(5):1283–??, December 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/115/5/1283.

Calof:1991:RBN

[CL91a] A. L. Calof and A. D. Lander. Relationship between neuronal migration and cell-substratum adhesion: laminin and merosin promote olfactory neuronal migration but are antiadhesive. Journal of Cell Biology, 115(3):779–??, November 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/115/3/779.

Clark:1991:DCC

[CL91b] S. E. Clark and G. K. Lamppa. Determinants for cleavage of the chlorophyll a/b binding protein precursor: a requirement for a basic residue that is not universal for chloroplast imported proteins. Journal of Cell Biology, 114(4): 681–??, August 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/114/4/681.

Choubey:1992:IAN

[CL92] D. Choubey and P. Lengyel. Interferon action: nucleolar and nucleoplasmic localization of the interferon-inducible 72-kD protein that is encoded by the Ifi 204 gene from the gene 200 cluster. Journal of Cell Biology, 116(6):1333–??, March 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/116/6/1333.

Ching:1993:ATI

[CL93] G. Y. Ching and R. K. Liem. Assembly of type IV neuronal intermediate filaments in nonneuronal cells in the absence of preexisting cytoplasmic intermediate filaments. Journal of Cell Biology, 122(6):1323–??, September 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/122/6/1323.

Chen:1994:RGF

[CL94] W. J. Chen and R. K. Liem. Reexpression of glial fibrillary acidic protein rescues the ability of astrocytoma cells to form processes in response to neurons. Journal of Cell Biology, 127(3):813–??, November 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/127/3/813.

Cartaud:1990:PAR

[CLCC90] A. Cartaud, M. A. Ludosky, J. C. Courvalin, and J. Cartaud. A protein antigenically related to nuclear lamin B mediates the association of intermediate filaments with desmosomes. Journal of Cell Biology, 111(2):581–??, August 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/111/2/
581.

Chan:1991:IRL

[CLD^{+}91] P. Y. Chan, M. B. Lawrence, M. L. Dustin, L. M. Ferguson, D. E. Golan, and T. A. Springer. Influence of receptor lateral mobility on adhesion strengthening between membranes containing LFA-3 and CD2. Journal of Cell Biology, 115(1):245–??, October 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/115/1/245.

Chen:1990:ASL

[CLDV90] Q. Chen, L. M. Lauzon, A. E. DeRocher, and E. Vierling. Accumulation, stability, and localization of a major chloroplast heat-shock protein. Journal of Cell Biology, 110(6): 1873–??, June 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/110/6/1873.

Chen:1992:DTX

[CLG^{+}92] Q. Chen, C. Linsenmayer, H. Gu, T. M. Schmid, and T. F. Linsenmayer. Domains of type X collagen: alteration of cartilage matrix by fibril association and proteoglycan accumulation. Journal of Cell Biology, 117(3):687–??, May 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/117/3/
687.

Collins:1991:CSA

[CLK^{+}91] J. E. Collins, P. K. Legan, T. P. Kenny, J. MacGarvie, J. L. Holton, and D. R. Garrod. Cloning and sequence analysis of desmosomal glycoproteins 2 and 3 (desmocollins): cadherin-like desmosomal adhesion molecules with heterogeneous cytoplasmic domains. Journal of Cell Biology, 113 (2):381–??, April 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/113/2/381.

Coue:1991:MDP

[CLM91] M. Coue, V. A. Lombillo, and J. R. McIntosh. Microtubule depolymerization promotes particle and chromosome movement in vitro. Journal of Cell Biology, 112(6):1165–??, March 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/112/6/1165.

Cheng:1994:SER

[CLP^{+}94] A. Cheng, T. Le, M. Palacios, L. H. Bookbinder, P. M. Wassarman, F. Suzuki, and J. D. Bleil. Sperm-egg recognition in the mouse: characterization of sp56, a sperm protein having specific affinity for ZP3. Journal of Cell Biology, 125 (4):867–??, May 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/125/4/867.

Cusella-DeAngelis:1992:MMI

[CLS^{+}92] M. G. Cusella-De Angelis, G. Lyons, C. Sonnino, L. De Angelis, E. Vivarelli, K. Farmer, W. E. Wright, M. Molinaro, M. Bouchè, and M. Buckingham. MyoD, myogenin independent differentiation of primordial myoblasts in mouse somites. Journal of Cell Biology, 116(5):1243–??, March 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/116/5/1243.

Cano:1991:KAF

[CLZ91] M. L. Cano, D. A. Lauffenburger, and S. H. Zigmond. Kinetic analysis of F-actin depolymerization in polymorphonuclear leukocyte lysates indicates that chemoattractant stimulation increases actin filament number without altering the filament length distribution. Journal of Cell Biology, 115(3):677–??, November 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/115/3/677.

Cramer:1993:MSA

[CM93] L. Cramer and T. J. Mitchison. Moving and stationary actin filaments are involved in spreading of postmitotic PtK2 cells. Journal of Cell Biology, 122(4):833–??, August 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/122/4/
833.

Carnell:1994:TRS

[CM94a] L. Carnell and H. P. Moore. Transport via the regulated secretory pathway in semi-intact PC12 cells: role of intracisternal calcium and pH in the transport and sorting of secretogranin II. Journal of Cell Biology, 127(3):693–??, November 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/127/3/693.

Clark:1994:APC

[CM94b] S. W. Clark and D. I. Meyer. ACT3: a putative centractin homologue in S. cerevisiae is required for proper orientation of the mitotic spindle. Journal of Cell Biology, 127(1): 129–??, October 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/127/1/129.

Cooper:1991:EML

[CMB91] D. N. Cooper, S. M. Massa, and S. H. Barondes. Endogenous muscle lectin inhibits myoblast adhesion to laminin. Journal of Cell Biology, 115(5):1437–??, December 1991.

CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/115/5/
1437.

Crawford:1992:IBZ

[CMB92] A. W. Crawford, J. W. Michelsen, and M. C. Beckerle. An interaction between zyxin and alpha-actinin. Journal of Cell Biology, 116(6):1381–??, March 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/116/6/1381.

Culty:1990:HRM

[CMK^{+}90] M. Culty, K. Miyake, P. W. Kincade, E. Sikorski, E. C. Butcher, C. Underhill, and E. Silorski. The hyaluronate receptor is a member of the CD44 (H–CAM) family of cell surface glycoproteins. Journal of Cell Biology, 111(6):2765– ??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/111/6/2765.

Ciambrone:1990:PAI

[CML90] G. J. Ciambrone and P. J. McKeown-Longo. Plasminogen activator inhibitor type I stabilizes vitronectin-dependent adhesions in HT-1080 cells. Journal of Cell Biology, 111 (5):2183–??, November 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/111/5/2183.

Clark-Maguire:1994:LMG

[CMM94] S. Clark-Maguire and P. E. Mains. Localization of the mei-1 gene product of Caenorhaditis elegans, a meiotic-specific spindle component. Journal of Cell Biology, 126(1):199– ??, July 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/126/1/199.

Chelberg:1990:CSP

[CMS^{+}90a] M. K. Chelberg, J. B. McCarthy, A. P. Skubitz, L. T. Furcht, and E. C. Tsilibary. Characterization of a synthetic peptide from type IV collagen that promotes melanoma cell adhesion, spreading, and motility. Journal of Cell Biology, 111(1):261–??, July 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/111/1/261.

Closs:1990:CFE

[CMS^{+}90b] E. I. Closs, A. B. Murray, J. Schmidt, A. Schön, V. Erfle, and P. G. Strauss. c-fos expression precedes osteogenic differentiation of cartilage cells in vitro. Journal of Cell Biology, 111(3):1313–??, September 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/111/3/1313.

Chretien:1992:LDM

[CMV^{+}92] D. Chrétien, F. Metoz, F. Verde, E. Karsenti, and R. H. Wade. Lattice defects in microtubules: protofilament numbers vary within individual microtubules. Journal of Cell Biology, 117(5):1031–??, June 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/117/5/1031.

Cassimeris:1990:CSP

[CMZ90] L. Cassimeris, H. McNeill, and S. H. Zigmond. Chemoattractantstimulated polymorphonuclear leukocytes contain two populations of actin filaments that differ in their spatial distributions and relative stabilities. Journal of Cell Biology, 110 (4):1067–??, April 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/110/4/1067.

Calafat:1994:MHC

[CNJ^{+}94] J. Calafat, M. Nijenhuis, H. Janssen, A. Tulp, S. Dusseljee, R. Wubbolts, and J. Neefjes. Major histocompatibility complex class II molecules induce the formation of endocytic MIIC-like structures. Journal of Cell Biology, 126(4): 967–??, August 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/126/4/967.

Charo:1990:VRA

[CNSC90] I. F. Charo, L. Nannizzi, J. W. Smith, and D. A. Cheresh. The vitronectin receptor alpha v beta 3 binds fibronectin and acts in concert with alpha 5 beta 1 in promoting cellular attachment and spreading on fibronectin. Journal of Cell Biology, 111(6):2795–??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/111/6/2795.

Culty:1992:HRC

[CNU92] M. Culty, H. A. Nguyen, and C. B. Underhill. The hyaluronan receptor (CD44) participates in the uptake and degradation of hyaluronan. Journal of Cell Biology, 116(4):1055– ??, February 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/116/4/1055.

Charles:1992:ICS

[CNZ^{+}92] A. C. Charles, C. C. Naus, D. Zhu, G. M. Kidder, E. R. Dirksen, and M. J. Sanderson. Intercellular calcium signaling via gap junctions in glioma cells. Journal of Cell Biology, 118(1):195–??, July 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/118/1/195.

Chen:1991:CCC

[CO91] W. C. Chen and B. Obrink. Cell–cell contacts mediated by E-cadherin (uvomorulin) restrict invasive behavior of Lcells. Journal of Cell Biology, 114(2):319–??, July 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/114/2/
319.

Coluccio:1994:EST

[Col94] L. M. Coluccio. An end in sight: tropomodulin. Journal of Cell Biology, 127(6):1497–??, December 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/127/6/1497.

Clift-OGrady:1990:BSV

[COLL^{+}90] L. Clift-O’Grady, A. D. Linstedt, A. W. Lowe, E. Grote, and R. B. Kelly. Biogenesis of synaptic vesicle-like structures in a pheochromocytoma cell line PC-12. Journal of Cell Biology, 110(5):1693–??, May 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/110/5/1693.

Chen:1994:TDD

[COTC94] P. Chen, B. D. Ostrow, S. R. Tafuri, and R. L. Chisholm. Targeted disruption of the Dictyostelium RMLC gene produces cells defective in cytokinesis and development. Journal of Cell Biology, 127(6):1933–??, December 1994. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/127/6/
1933.

Chege:1990:CGC

[CP90] N. W. Chege and S. R. Pfeffer. Compartmentation of the Golgi complex: brefeldin–A distinguishes trans-Golgi cisternae from the trans-Golgi network. Journal of Cell Biology, 111(3):893–??, September 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/111/3/893.

Choquet:1993:CLI

[CPA^{+}93] D. Choquet, M. Partiseti, S. Amigorena, C. Bonnerot, W. H. Fridman, and H. Korn. Cross-linking of IgG receptors inhibits membrane immunoglobulin-stimulated calcium influx in B lymphocytes. Journal of Cell Biology, 121(2): 355–??, April 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/121/2/355.

Carpentier:1993:TSI

[CPB^{+}93] J. L. Carpentier, J. P. Paccaud, J. Backer, A. Gilbert, L. Orci, C. R. Kahn, and J. Backer. Two steps of insulin receptor internalization depend on different domains of the beta-subunit. Journal of Cell Biology, 122(6):1243– ??, September 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/122/6/1243.

Crawford:1994:BMC

[CPB94] A. W. Crawford, J. D. Pino, and M. C. Beckerle. Biochemical and molecular characterization of the chicken cysteinerich protein, a developmentally regulated LIM-domain protein that is associated with the actin cytoskeleton. Journal of Cell Biology, 124(1):117–??, January 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/124/1/117.

Colgan:1993:NMA

[CPD^{+}93] S. P. Colgan, C. A. Parkos, C. Delp, M. A. Arnaout, and J. L. Madara. Neutrophil migration across cultured intestinal epithelial monolayers is modulated by epithelial exposure to IFN-gamma in a highly polarized fashion. Journal of Cell Biology, 120(3):785–??, February 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/120/3/785.

Coppard:1991:PXO

[CPM^{+}91] N. J. Coppard, K. Poulsen, H. O. Madsen, J. Frydenberg, and B. F. Clark. 42sp48 in previtellogenic Xenopus oocytes is structurally homologous to EF-1 alpha and may be a stage-specific elongation factor. Journal of Cell Biology, 112(2):237–??, January 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/112/2/237.

Crepaldi:1994:TSH

[CPP^{+}94] T. Crepaldi, A. L. Pollack, M. Prat, A. Zborek, K. Mostov, and P. M. Comoglio. Targeting of the SF/HGF receptor to the basolateral domain of polarized epithelial cells. Journal of Cell Biology, 125(2):313–??, April 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/125/2/313.

Carpen:1992:AIA

[CPSS92] O. Carpén, P. Pallai, D. E. Staunton, and T. A. Springer. Association of intercellular adhesion molecule-1 (ICAM-1) with actin-containing cytoskeleton and alpha-actinin. Journal of Cell Biology, 118(5):1223–??, September 1992. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/118/5/
1223.

Campanero:1990:ALH

[CPU^{+}90] M. R. Campanero, R. Pulido, M. A. Ursa, M. Rodríguez-Moya, M. O. de Landázuri, and F. Sánchez-Madrid. An alternative leukocyte homotypic adhesion mechanism, LFA-1/ICAM-1-independent, triggered through the human VLA-4 integrin. Journal of Cell Biology, 110(6):2157–??, June 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/110/6/2157.

Connor:1994:SAC

[CQYD94] E. A. Connor, K. Qin, H. Yankelev, and D. DeStefano. Synaptic activity and connective tissue remodeling in denervated frog muscle. Journal of Cell Biology, 127(5):1435– ??, December 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/127/5/1435.

Cordes:1993:IFC

[CRK^{+}93] V. C. Cordes, S. Reidenbach, A. Köhler, N. Stuurman, R. van Driel, and W. W. Franke. Intranuclear filaments containing a nuclear pore complex protein. Journal of Cell Biology, 123(6):1333–??, December 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/123/6/1333.

Chang:1992:LCC

[CRKA92] W. J. Chang, K. G. Rothberg, B. A. Kamen, and R. G. Anderson. Lowering the cholesterol content of MA104 cells inhibits receptor-mediated transport of folate. Journal of Cell Biology, 118(1):63–??, July 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/118/1/63.

Caplow:1994:FEH

[CRS94] M. Caplow, R. L. Ruhlen, and J. Shanks. The free energy for hydrolysis of a microtubule-bound nucleotide triphosphate is near zero: all of the free energy for hydrolysis is stored in the microtubule lattice. Journal of Cell Biology, 127(3): 779–??, November 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/127/3/779.

Carter:1993:MGB

[CRWS93] L. L. Carter, T. E. Redelmeier, L. A. Woollenweber, and S. L. Schmid. Multiple GTP-binding proteins participate in clathrin-coated vesicle-mediated endocytosis. Journal of Cell Biology, 120(1):37–??, January 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/120/1/37.

Carey:1990:ILA

[CS90] D. J. Carey and R. C. Stahl. Identification of a lipidanchored heparan sulfate proteoglycan in Schwann cells. Journal of Cell Biology, 111(5):2053–??, November 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/111/5/
2053.

Chang:1991:MYP

[CS91] A. Chang and C. W. Slayman. Maturation of the yeast plasma membrane [H+]ATPase involves phosphorylation during intracellular transport. Journal of Cell Biology, 115 (2):289–??, October 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/115/2/289.

Cadigan:1990:ICC

[CSC90] K. M. Cadigan, D. M. Spillane, and T. Y. Chang. Isolation and characterization of Chinese hamster ovary cell mutants defective in intracellular low density lipoproteincholesterol trafficking. Journal of Cell Biology, 110(2):295– ??, February 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/110/2/295.

Compton:1992:PSN

[CSC92] D. A. Compton, I. Szilak, and D. W. Cleveland. Primary structure of NuMA, an intranuclear protein that defines a novel pathway for segregation of proteins at mitosis. Journal of Cell Biology, 116(6):1395–??, March 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/116/6/1395.

Carey:1994:SES

[CSCSA94] D. J. Carey, R. C. Stahl, G. Cizmeci-Smith, and V. K. Asundi. Syndecan-1 expressed in Schwann cells causes morphological transformation and cytoskeletal reorganization and associates with actin during cell spreading. Journal of Cell Biology, 124(1):161–??, January 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/124/1/161.

Chowdhury:1992:OSY

[CSG92] S. Chowdhury, K. W. Smith, and M. C. Gustin. Osmotic stress and the yeast cytoskeleton: phenotype-specific suppression of an actin mutation. Journal of Cell Biology, 118 (3):561–??, August 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/118/3/561.

Curtis:1992:GEN

[CSH^{+}92] R. Curtis, H. J. Stewart, S. M. Hall, G. P. Wilkin, R. Mirsky, and K. R. Jessen. GAP-43 is expressed by nonmyelinforming Schwann cells of the peripheral nervous system. Journal of Cell Biology, 116(6):1455–??, March 1992. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/116/6/
1455.

Cameron:1991:CST

[CSJD91] P. L. Cameron, T. C. Südhof, R. Jahn, and P. De Camilli. Colocalization of synaptophysin with transferrin receptors: implications for synaptic vesicle biogenesis. Journal of Cell Biology, 115(1):151–??, October 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/115/1/151.

Caroni:1994:RMI

[CSKZ94] P. Caroni, C. Schneider, M. C. Kiefer, and J. Zapf. Role of muscle insulin-like growth factors in nerve sprouting: suppression of terminal sprouting in paralyzed muscle by IGFbinding protein 4. Journal of Cell Biology, 125(4):893–??, May 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/125/4/893.

Collins:1990:CDR

[CSM90] K. Collins, J. R. Sellers, and P. Matsudaira. Calmodulin dissociation regulates brush border myosin I (110-kD-calmodulin) mechanochemical activity in vitro. Journal of Cell Biology, 110(4):1137–??, April 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/110/4/1137.

Campanero:1994:IRL

[CSMdPSM94] M. R. Campanero, P. Sánchez-Mateos, M. A. del Pozo, and F. Sánchez-Madrid. ICAM-3 regulates lymphocyte morphology and integrin-mediated T cell interaction with endothelial cell and extracellular matrix ligands. Journal of Cell Biology, 127(3):867–??, November 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/127/3/867.

Choi:1990:CLP

[CSMG90] Y. S. Choi, R. Sehgal, P. McCrea, and B. Gumbiner. A cadherin-like protein in eggs and cleaving embryos of Xenopus laevis is expressed in oocytes in response to progesterone. Journal of Cell Biology, 110(5):1575–??, May 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/110/5/
1575.

Cassimeris:1992:TBS

[CSNZ92] L. Cassimeris, D. Safer, V. T. Nachmias, and S. H. Zigmond. Thymosin beta 4 sequesters the majority of G-actin in resting human polymorphonuclear leukocytes. Journal of Cell Biology, 119(5):1261–??, December 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/119/5/1261.

Chen:1990:PHT

[CSP90] Q. Chen, R. Sealock, and H. B. Peng. A protein homologous to the Torpedo postsynaptic 58K protein is present at the myotendinous junction. Journal of Cell Biology, 110(6): 2061–??, June 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/110/6/2061.

Cohn:1992:UTM

[CSR^{+}92] S. M. Cohn, T. C. Simon, K. A. Roth, E. H. Birkenmeier, and J. I. Gordon. Use of transgenic mice to map cis-acting elements in the intestinal fatty acid binding protein gene (Fabpi) that control its cell lineage-specific and regional patterns of expression along the duodenal-colonic and cryptvillus axes of the gut epithelium. Journal of Cell Biology, 119(1):27–??, October 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/119/1/27.

Clement:1990:HAL

[CSRS^{+}90] B. Clément, B. Segui-Real, P. Savagner, H. K. Kleinman, and Y. Yamada. Hepatocyte attachment to laminin is mediated through multiple receptors. Journal of Cell Biology, 110(1):185–??, January 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/110/1/185.

Cardone:1994:PMA

[CSS^{+}94] M. H. Cardone, B. L. Smith, W. Song, D. Mochly-Rosen, and K. E. Mostov. Phorbol myristate acetate-mediated stimulation of transcytosis and apical recycling in MDCK cells. Journal of Cell Biology, 124(5):717–??, March 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/124/5/
717.

Chia:1993:IMP

[CSSL93] C. P. Chia, A. Shariff, S. A. Savage, and E. J. Luna. The integral membrane protein, ponticulin, acts as a monomer in nucleating actin assembly. Journal of Cell Biology, 120(4): 909–??, February 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/120/4/909.

Conradt:1992:VRV

[CSV^{+}92] B. Conradt, J. Shaw, T. Vida, S. Emr, and W. Wickner. In vitro reactions of vacuole inheritance in Saccharomyces cerevisiae. Journal of Cell Biology, 119(6):1469–??, December 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/119/6/1469.

Cerneus:1993:BTD

[CSvdE93] D. P. Cerneus, G. J. Strous, and A. van der Ende. Bidirectional transcytosis determines the steady state distribution of the transferrin receptor at opposite plasma membrane domains of BeWo cells. Journal of Cell Biology, 122(6):1223– ??, September 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/122/6/1223.

Chapman:1991:AML

[CT91] K. D. Chapman and R. N. Trelease. Acquisition of membrane lipids by differentiating glyoxysomes: role of lipid bodies. Journal of Cell Biology, 115(4):995–??, November 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/115/4/995.

Carter:1991:DND

[CTL91] K. C. Carter, K. L. Taneja, and J. B. Lawrence. Discrete nuclear domains of poly(A) RNA and their relationship to the functional organization of the nucleus. Journal of Cell Biology, 115(5):1191–??, December 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/115/5/1191.

Corthesy-Theulaz:1992:CDP

[CTPP92] I. Corthésy-Theulaz, A. Pauloin, and S. R. Pfeffer. Cytoplasmic dynein participates in the centrosomal localization of the Golgi complex. Journal of Cell Biology, 118(6):1333– ??, September 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/118/6/1333.

Cooper:1991:MLR

[CTQ91] H. M. Cooper, R. N. Tamura, and V. Quaranta. The major laminin receptor of mouse embryonic stem cells is a novel isoform of the alpha 6 beta 1 integrin. Journal of Cell Biology, 115(3):843–??, November 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/115/3/843.

Cerneus:1991:ABT

[CvdE91] D. P. Cerneus and A. van der Ende. Apical and basolateral transferrin receptors in polarized BeWo cells recycle through separate endosomes. Journal of Cell Biology, 114(6):1149– ??, September 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/114/6/1149.

Cortese:1991:ISI

[CVH91] J. D. Cortese, A. L. Voglino, and C. R. Hackenbrock. Ionic strength of the intermembrane space of intact mitochondria as estimated with fluorescein–BSA delivered by low pH fusion. Journal of Cell Biology, 113(6):1331–??, June 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/113/6/
1331.

Cupers:1994:CPR

[CVK^{+}94] P. Cupers, A. Veithen, A. Kiss, P. Baudhuin, and P. J. Courtoy. Clathrin polymerization is not required for bulkphase endocytosis in rat fetal fibroblasts. Journal of Cell Biology, 127(3):725–??, November 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/127/3/725.

Cao:1990:MFCa

[CW90a] L. G. Cao and Y. L. Wang. Mechanism of the formation of contractile ring in dividing cultured animal cells. I. Recruitment of preexisting actin filaments into the cleavage furrow. Journal of Cell Biology, 110(4):1089–??, April 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/110/4/
1089.

Cao:1990:MFCb

[CW90b] L. G. Cao and Y. L. Wang. Mechanism of the formation of contractile ring in dividing cultured animal cells. II. Cortical movement of microinjected actin filaments. Journal of Cell Biology, 111(5):1905–??, November 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/111/5/1905.

Ching:1990:CTP

[CW90c] G. Ching and E. Wang. Characterization of two populations of statin and the relationship of their syntheses to the state of cell proliferation. Journal of Cell Biology, 110(2): 255–??, February 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/110/2/255.

Cho:1993:EME

[CWB93] M. Cho, S. G. Webster, and H. M. Blau. Evidence for myoblast-extrinsic regulation of slow myosin heavy chain expression during muscle fiber formation in embryonic development. Journal of Cell Biology, 121(4):795–??, May 1993.

CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/121/4/
795.

Cluett:1993:TGM

[CWBB93] E. B. Cluett, S. A. Wood, M. Banta, and W. J. Brown. Tubulation of Golgi membranes in vivo and in vitro in the absence of brefeldin A. Journal of Cell Biology, 120(1): 15–??, January 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/120/1/15.

Carter:1990:RIA

[CWBK90] W. G. Carter, E. A. Wayner, T. S. Bouchard, and P. Kaur. The role of integrins alpha 2 beta 1 and alpha 3 beta 1 in cell–cell and cell-substrate adhesion of human epidermal cells. Journal of Cell Biology, 110(4):1387–??, April 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/110/4/
1387.

Coates:1992:RFA

[CWHH92] T. D. Coates, R. G. Watts, R. Hartman, and T. H. Howard. Relationship of F-actin distribution to development of polar shape in human polymorphonuclear neutrophils. Journal of Cell Biology, 117(4):765–??, May 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/117/4/765.

Chen:1994:EGF

[CXS^{+}94] P. Chen, H. Xie, M. C. Sekar, K. Gupta, and A. Wells. Epidermal growth factor receptor-mediated cell motility: phospholipase C activity is required, but mitogen-activated protein kinase activity is not sufficient for induced cell movement. Journal of Cell Biology, 127(3):847–??, November 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/127/3/847.

Compton:1991:INC

[CYC91] D. A. Compton, T. J. Yen, and D. W. Cleveland. Identification of novel centromere/kinetochore-associated proteins using monoclonal antibodies generated against human mitotic chromosome scaffolds. Journal of Cell Biology, 112(6):

1083–??, March 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/112/6/1083.

Clark:1990:SYS

[CYCA90] M. W. Clark, M. L. Yip, J. Campbell, and J. Abelson. SSB-1 of the yeast Saccharomyces cerevisiae is a nucleolarspecific, silver-binding protein that is associated with the snR10 and snR11 small nuclear RNAs. Journal of Cell Biology, 111(5):1741–??, November 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/111/5/1741.

Chang:1994:PCS

[CYR^{+}94] W. J. Chang, Y. S. Ying, K. G. Rothberg, N. M. Hooper, A. J. Turner, H. A. Gambliel, J. De Gunzburg, S. M. Mumby, A. G. Gilman, and R. G. Anderson. Purification and characterization of smooth muscle cell caveolae. Journal of Cell Biology, 126(1):127–??, July 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/126/1/127.

Dentler:1992:FMD

[DA92] W. L. Dentler and C. Adams. Flagellar microtubule dynamics in Chlamydomonas: cytochalasin D induces periods of microtubule shortening and elongation; and colchicine induces disassembly of the distal, but not proximal, half of the flagellum. Journal of Cell Biology, 117(6):1289–??, June 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/117/6/1289.

Draeger:1990:CCA

[DAIS90] A. Draeger, W. B. Amos, M. Ikebe, and J. V. Small. The cytoskeletal and contractile apparatus of smooth muscle: contraction bands and segmentation of the contractile elements. Journal of Cell Biology, 111(6):2463–??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/111/6/
2463.

DeNardi:1993:TMH

[DAM^{+}93] C. DeNardi, S. Ausoni, P. Moretti, L. Gorza, M. Velleca, M. Buckingham, and S. Schiaffino. Type 2X-myosin heavy chain is coded by a muscle fiber type-specific and developmentally regulated gene. Journal of Cell Biology, 123(4): 823–??, November 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/123/4/823.

Donoghue:1991:FTP

[DAMS91] M. J. Donoghue, J. D. Alvarez, J. P. Merlie, and J. R. Sanes. Fiber type- and position-dependent expression of a myosin light chain–CAT transgene detected with a novel histochemical stain for CAT. Journal of Cell Biology, 115(2):423–??, October 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/115/2/423.

Diener:1993:AFR

[DAR93] D. R. Diener, L. H. Ang, and J. L. Rosenbaum. Assembly of flagellar radial spoke proteins in Chlamydomonas: identification of the axoneme binding domain of radial spoke protein 3. Journal of Cell Biology, 123(1):183–??, October 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/123/1/183.

Doherty:1992:GMN

[DAS^{+}92] P. Doherty, S. V. Ashton, S. D. Skaper, A. Leon, and F. S. Walsh. Ganglioside modulation of neural cell adhesion molecule and N-cadherin-dependent neurite outgrowth. Journal of Cell Biology, 117(5):1093–??, June 1992. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/117/5/
1093.

Davis:1992:TSC

[Dav92] T. N. Davis. A temperature-sensitive calmodulin mutant loses viability during mitosis. Journal of Cell Biology, 118 (3):607–??, August 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/118/3/607.

Dahan:1994:CIH

[DAW^{+}94] S. Dahan, J. P. Ahluwalia, L. Wong, B. I. Posner, and J. J. Bergeron. Concentration of intracellular hepatic apolipoprotein E in Golgi apparatus saccular distensions and endosomes. Journal of Cell Biology, 127(6):1859–??, December 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/127/6/1859.

deArruda:1990:FHC

[dAWL^{+}90] M. V. de Arruda, S. Watson, C. S. Lin, J. Leavitt, and P. Matsudaira. Fimbrin is a homologue of the cytoplasmic phosphoprotein plastin and has domains homologous with calmodulin and actin gelation proteins. Journal of Cell Biology, 111(3):1069–??, September 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/111/3/1069.

deAgostini:1990:LAA

[dAWS^{+}90] A. I. de Agostini, S. C. Watkins, H. S. Slayter, H. Youssoufian, and R. D. Rosenberg. Localization of anticoagulantly active heparan sulfate proteoglycans in vascular endothelium: antithrombin binding on cultured endothelial cells and perfused rat aorta. Journal of Cell Biology, 111 (3):1293–??, September 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/111/3/1293.

Dyer:1990:GTS

[DB90] C. A. Dyer and J. A. Benjamins. Glycolipids and transmembrane signaling: antibodies to galactocerebroside cause an influx of calcium in oligodendrocytes. Journal of Cell Biology, 111(2):625–??, August 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/111/2/625.

Desjardins:1991:OGB

[DB91] M. Desjardins and M. Bendayan. Ontogenesis of glomerular basement membrane: structural and functional properties. Journal of Cell Biology, 113(3):689–??, May 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/113/3/689.

deBeus:1994:YNE

[dBBHA94] E. de Beus, J. S. Brockenbrough, B. Hong, and J. P. Aris. Yeast NOP2 encodes an essential nucleolar protein with homology to a human proliferation marker. Journal of Cell Biology, 127(6):1799–??, December 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/127/6/1799.

Doliana:1990:MFC

[DBC90] R. Doliana, P. Bonaldo, and A. Colombatti. Multiple forms of chicken alpha 3(VI) collagen chain generated by alternative splicing in type A repeated domains. Journal of Cell Biology, 111(5):2197–??, November 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/111/5/2197.

DeTomaso:1994:ABS

[DBM94] A. W. DeTomaso, G. Blanco, and R. W. Mercer. The alpha and beta subunits of the Na,K–ATPase can assemble at the plasma membrane into functional enzyme. Journal of Cell Biology, 127(1):55–??, October 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/127/1/55.

deBrabander:1991:LDR

[dBNI^{+}91] M. de Brabander, R. Nuydens, A. Ishihara, B. Holifield, K. Jacobson, and H. Geerts. Lateral diffusion and retrograde movements of individual cell surface components on single motile cells observed with Nanovid microscopy. Journal of Cell Biology, 112(1):111–??, January 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/112/1/111.

Dubreuil:1990:BSI

[DBSK90] R. R. Dubreuil, T. J. Byers, C. T. Stewart, and D. P. Kiehart. A beta-spectrin isoform from Drosophila (beta H) is similar in size to vertebrate dystrophin. Journal of Cell Biology, 111(5):1849–??, November 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/111/5/1849.

DiCicco-Bloom:1990:NMD

[DBTAB90] E. DiCicco-Bloom, E. Townes-Anderson, and I. B. Black. Neuroblast mitosis in dissociated culture: regulation and relationship to differentiation. Journal of Cell Biology, 110 (6):2073–??, June 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/110/6/2073.

David:1992:DCH

[DBV^{+}92] G. David, X. M. Bai, B. Van der Schueren, J. J. Cassiman, and H. Van den Berghe. Developmental changes in heparan sulfate expression: in situ detection with mAbs. Journal of Cell Biology, 119(4):961–??, November 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/119/4/961.

Damke:1994:IMD

[DBWS94] H. Damke, T. Baba, D. E. Warnock, and S. L. Schmid. Induction of mutant dynamin specifically blocks endocytic coated vesicle formation. Journal of Cell Biology, 127(4): 915–??, November 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/127/4/915.

Dent:1992:HCF

[DCB^{+}92] J. A. Dent, R. B. Cary, J. B. Bachant, A. Domingo, and M. W. Klymkowsky. Host cell factors controlling vimentin organization in the Xenopus oocyte. Journal of Cell Biology, 119(4):855–??, November 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/119/4/855.

Dransfield:1992:ILI

[DCBH92] I. Dransfield, C. Cabañas, J. Barrett, and N. Hogg. Interaction of leukocyte integrins with ligand is necessary but not sufficient for function. Journal of Cell Biology, 116(6): 1527–??, March 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/116/6/1527.

Dransfield:1992:DCR

[DCCH92] I. Dransfield, C. Cabañas, A. Craig, and N. Hogg. Divalent cation regulation of the function of the leukocyte integrin LFA-1. Journal of Cell Biology, 116(1):219–??, January 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/116/1/219.

Dixon:1992:BLL

[DCH^{+}92] J. L. Dixon, R. Chattapadhyay, T. Huima, C. M. Redman, and D. Banerjee. Biosynthesis of lipoprotein: location of nascent apoAI and apoB in the rough endoplasmic reticulum of chicken hepatocytes. Journal of Cell Biology, 117 (6):1161–??, June 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/117/6/1161.

Downey:1992:PEI

[DCL^{+}92] G. P. Downey, C. K. Chan, P. Lea, A. Takai, and S. Grinstein. Phorbol ester-induced actin assembly in neutrophils: role of protein kinase C. Journal of Cell Biology, 116(3): 695–??, February 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/116/3/695.

Desai:1993:SPS

[DCM^{+}93] N. N. Desai, R. O. Carlson, M. E. Mattie, A. Olivera, N. E. Buckley, T. Seki, G. Brooker, and S. Spiegel. Signaling pathways for sphingosylphosphorylcholine-mediated mitogenesis in Swiss 3T3 fibroblasts. Journal of Cell Biology, 121(6): 1385–??, June 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/121/6/1385.

deCurtis:1991:LRR

[dCQTR91] I. de Curtis, V. Quaranta, R. N. Tamura, and L. F. Reichardt. Laminin receptors in the retina: sequence analysis of the chick integrin alpha 6 subunit. Evidence for transcriptional and posttranslational regulation. Journal of Cell Biology, 113(2):405–??, April 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/113/2/405.

Downey:1990:AAE

[DCTG90] G. P. Downey, C. K. Chan, S. Trudel, and S. Grinstein. Actin assembly in electropermeabilized neutrophils: role of intracellular calcium. Journal of Cell Biology, 110(6):1975– ??, June 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/110/6/1975.

deCuevas:1992:ESD

[dCTG92] M. de Cuevas, T. Tao, and L. S. Goldstein. Evidence that the stalk of Drosophila kinesin heavy chain is an alphahelical coiled coil. Journal of Cell Biology, 116(4):957–??, February 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/116/4/957.

daCunha:1990:RIG

[dCV90] A. da Cunha and L. Vitković. Regulation of immunoreactive GAP-43 expression in rat cortical macroglia is cell type specific. Journal of Cell Biology, 111(1):209–??, July 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/111/1/
209.

DeLisser:1994:DCD

[DCY^{+}94] H. M. DeLisser, J. Chilkotowsky, H. C. Yan, M. L. Daise, C. A. Buck, and S. M. Albelda. Deletions in the cytoplasmic domain of platelet-endothelial cell adhesion molecule-1 (PECAM-1, CD31) result in changes in ligand binding properties. Journal of Cell Biology, 124(1):195–??, January 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/124/1/
195.

Dehouck:1994:ULD

[DDFC94] B. Dehouck, M. P. Dehouck, J. C. Fruchart, and R. Cecchelli. Upregulation of the low density lipoprotein receptor at the blood-brain barrier: intercommunications between brain capillary endothelial cells and astrocytes. Journal of Cell Biology, 126(2):465–??, July 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/126/2/465.

Damon:1990:NGF

[DDW90] D. H. Damon, P. A. D’Amore, and J. A. Wagner. Nerve growth factor and fibroblast growth factor regulate neurite outgrowth and gene expression in PC12 cells via both protein kinase C- and cAMP-independent mechanisms. Journal of Cell Biology, 110(4):1333–??, April 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/110/4/1333.

Dimitrov:1994:RSC

[DDW94] S. Dimitrov, M. C. Dasso, and A. P. Wolffe. Remodeling sperm chromatin in Xenopus laevis egg extracts: the role of core histone phosphorylation and linker histone B4 in chromatin assembly. Journal of Cell Biology, 126(3):591–??, August 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/126/3/591.

Dix:1990:MMA

[DE90] D. J. Dix and B. R. Eisenberg. Myosin mRNA accumulation and myofibrillogenesis at the myotendinous junction of stretched muscle fibers. Journal of Cell Biology, 111(5): 1885–??, November 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/111/5/1885.

Deitiker:1993:TFS

[DE93] P. R. Deitiker and H. F. Epstein. Thick filament substructures in Caenorhabditis elegans: evidence for two populations of paramyosin. Journal of Cell Biology, 123(2):303–??, October 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/123/2/303.

Drenckhahn:1991:TDA

[DEH^{+}91] D. Drenckhahn, K. Engel, D. Höfer, C. Merte, L. Tilney, and M. Tilney. Three different actin filament assemblies occur in every hair cell: each contains a specific actin crosslinking protein. Journal of Cell Biology, 112(4):641– ??, February 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/112/4/641.

Downey:1991:BPM

[DES^{+}91] G. P. Downey, E. L. Elson, B. Schwab, S. C. Erzurum, S. K. Young, and G. S. Worthen. Biophysical properties and microfilament assembly in neutrophils: modulation by cyclic AMP. Journal of Cell Biology, 114(6):1179–??, September 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/114/6/1179.

Devault:1992:CPM

[DFC^{+}92] A. Devault, D. Fesquet, J. C. Cavadore, A. M. Garrigues, J. C. Labbé, T. Lorca, A. Picard, M. Philippe, and M. Dorée. Cyclin A potentiates maturation-promoting factor activation in the early Xenopus embryo via inhibition of the tyrosine kinase that phosphorylates cdc2. Journal of Cell Biology, 118(5):1109–??, September 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/118/5/1109.

Duc:1994:CSE

[DFC^{+}94] C. Duc, N. Farman, C. M. Canessa, J. P. Bonvalet, and B. C. Rossier. Cell-specific expression of epithelial sodium channel alpha, beta, and gamma subunits in aldosteroneresponsive epithelia from the rat: localization by in situ hybridization and immunocytochemistry. Journal of Cell Biology, 127(6):1907–??, December 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/127/6/1907.

Duluc:1994:FEP

[DFLK94] I. Duluc, J. N. Freund, C. Leberquier, and M. Kedinger. Fetal endoderm primarily holds the temporal and positional information required for mammalian intestinal development. Journal of Cell Biology, 126(1):211–??, July 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/126/1/
211.

Dedhar:1990:INI

[DG90] S. Dedhar and V. Gray. Isolation of a novel integrin receptor mediating Arg–Gly–Asp–directed cell adhesion to fibronectin and type I collagen from human neuroblastoma cells. Association of a novel beta 1-related subunit with alpha v. Journal of Cell Biology, 110(6):2185–??, June 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/110/6/
2185.

Diamond:1993:DMR

[DGAB^{+}93] M. S. Diamond, J. Garcia-Aguilar, J. K. Bickford, A. L. Corbi, and T. A. Springer. The I domain is a major recognition site on the leukocyte integrin Mac-1 (CD11b/CD18) for four distinct adhesion ligands. Journal of Cell Biology, 120(4):1031–??, February 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/120/4/1031.

Dahl:1994:RCA

[DGF^{+}94] S. C. Dahl, R. W. Geib, M. T. Fox, M. Edidin, and D. Branton. Rapid capping in alpha-spectrin-deficient MEL cells from mice afflicted with hereditary hemolytic anemia. Journal of Cell Biology, 125(5):1057–??, June 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/125/5/1057.

Desmouliere:1993:TGF

[DGGG93] A. Desmoulière, A. Geinoz, F. Gabbiani, and G. Gabbiani. Transforming growth factor-beta 1 induces alpha-smooth muscle actin expression in granulation tissue myofibroblasts and in quiescent and growing cultured fibroblasts. Journal of Cell Biology, 122(1):103–??, July 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/122/1/103.

Durbec:1992:SFF

[DGGR92] P. Durbec, G. Gennarini, C. Goridis, and G. Rougon. A soluble form of the F3 neuronal cell adhesion molecule promotes neurite outgrowth. Journal of Cell Biology, 117(4): 877–??, May 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/117/4/877.

Darribere:1990:VAI

[DGL^{+}90] T. Darribère, K. Guida, H. Larjava, K. E. Johnson, K. M. Yamada, J. P. Thiery, and J. C. Boucaut. In vivo analyses of integrin beta 1 subunit function in fibronectin matrix assembly. Journal of Cell Biology, 110(5):1813–??, May 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/110/5/
1813.

Deery:1993:PIT

[DH93] W. J. Deery and J. P. Heath. Phagocytosis induced by thyrotropin in cultured thyroid cells is associated with myosin light chain dephosphorylation and stress fiber disruption.

Journal of Cell Biology, 122(1):21–??, July 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/122/1/21.

Danielsen:1993:ASA

[DHP93] E. M. Danielsen, G. H. Hansen, and M. D. Poulsen. Apical secretion of apolipoproteins from enterocytes. Journal of Cell Biology, 120(6):1347–??, March 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/120/6/1347.

Desjardins:1994:BPP

[DHPG94] M. Desjardins, L. A. Huber, R. G. Parton, and G. Griffiths. Biogenesis of phagolysosomes proceeds through a sequential series of interactions with the endocytic apparatus. Journal of Cell Biology, 124(5):677–??, March 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/124/5/677.

deHostos:1993:DML

[dHRB^{+}93] E. L. de Hostos, C. Rehfuess, B. Bradtke, D. R. Waddell, R. Albrecht, J. Murphy, and G. Gerisch. Dictyostelium mutants lacking the cytoskeletal protein coronin are defective in cytokinesis and cell motility. Journal of Cell Biology, 120(1):163–??, January 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/120/1/163.

Davis:1993:CTA

[DHS93] N. G. Davis, J. L. Horecka, and G. F. Sprague. Cis- and trans-acting functions required for endocytosis of the yeast pheromone receptors. Journal of Cell Biology, 122(1):53– ??, July 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/122/1/53.

denHartigh:1992:ERA

[dHvBeHVB92] J. C. den Hartigh, P. M. van Bergen en Henegouwen, A. J. Verkleij, and J. Boonstra. The EGF receptor is an actinbinding protein. Journal of Cell Biology, 119(2):349–??, October 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/119/2/349.

DePasquale:1991:ATN

[DI91] J. A. DePasquale and C. S. Izzard. Accumulation of talin in nodes at the edge of the lamellipodium and separate incorporation into adhesion plaques at focal contacts in fibroblasts. Journal of Cell Biology, 113(6):1351–??, June 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/113/6/
1351.

Dessev:1991:CCP

[DIDB^{+}91] G. Dessev, C. Iovcheva-Dessev, J. R. Bischoff, D. Beach, and R. Goldman. A complex containing p34cdc2 and cyclin B phosphorylates the nuclear lamin and disassembles nuclei of clam oocytes in vitro. Journal of Cell Biology, 112(4): 523–??, February 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/112/4/523.

Danowski:1992:CSF

[DIYSS92] B. A. Danowski, K. Imanaka-Yoshida, J. M. Sanger, and J. W. Sanger. Costameres are sites of force transmission to the substratum in adult rat cardiomyocytes. Journal of Cell Biology, 118(6):1411–??, September 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/118/6/1411.

Deckwerth:1993:TAE

[DJ93a] T. L. Deckwerth and E. M. Johnson. Temporal analysis of events associated with programmed cell death (apoptosis) of sympathetic neurons deprived of nerve growth factor. Journal of Cell Biology, 123(5):1207–??, December 1993. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/123/5/
1207.

Diestelkotter:1993:VIK

[DJ93b] P. Diestelkötter and W. W. Just. In vitro insertion of the 22-kD peroxisomal membrane protein into isolated rat liver peroxisomes. Journal of Cell Biology, 123(6):1717–??, December 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/123/6/1717.

Deloulme:1990:NGN

[DJA^{+}90] J. C. Deloulme, T. Janet, D. Au, D. R. Storm, M. Sensenbrenner, and J. Baudier. Neuromodulin (GAP43): a neuronal protein kinase C substrate is also present in 0-2A glial cell lineage. Characterization of neuromodulin in secondary cultures of oligodendrocytes and comparison with the neuronal antigen. Journal of Cell Biology, 111(4):1559–??, October 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/111/4/1559.

Duluc:1993:MLC

[DJF93] I. Duluc, B. Jost, and J. N. Freund. Multiple levels of control of the stage- and region-specific expression of rat intestinal lactase. Journal of Cell Biology, 123(6):1577–??, December 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/123/6/1577.

Dargemont:1992:EMM

[DK92] C. Dargemont and L. C. Kühn. Export of mRNA from microinjected nuclei of Xenopus laevis oocytes. Journal of Cell Biology, 118(1):1–??, July 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/118/1/1.

Duronio:1992:ISC

[DKG92] R. J. Duronio, L. J. Knoll, and J. I. Gordon. Isolation of a Saccharomyces cerevisiae long chain fatty acyl: CoA synthetase gene (FAA1) and assessment of its role in protein N-myristoylation. Journal of Cell Biology, 117(3):515–??, May 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/117/3/515.

Doherty:1990:SDI

[DKL^{+}90] J. J. Doherty, D. G. Kay, W. H. Lai, B. I. Posner, and J. J. Bergeron. Selective degradation of insulin within rat liver endosomes. Journal of Cell Biology, 110(1):35–??, January 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/110/1/35.

Drake:1992:CSP

[DKM^{+}92] S. L. Drake, D. J. Klein, D. J. Mickelson, T. R. Oegema, L. T. Furcht, and J. B. McCarthy. Cell surface phosphatidylinositol-anchored heparan sulfate proteoglycan initiates mouse melanoma cell adhesion to a fibronectinderived, heparin-binding synthetic peptide. Journal of Cell Biology, 117(6):1331–??, June 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/117/6/1331.

David:1990:MCP

[DLD^{+}90] G. David, V. Lories, B. Decock, P. Marynen, J. J. Cassiman, and H. Van den Berghe. Molecular cloning of a phosphatidylinositol-anchored membrane heparan sulfate proteoglycan from human lung fibroblasts. Journal of Cell Biology, 111(6):3165–??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/111/6/3165.

Dabauvalle:1991:SAP

[DLMS91] M. C. Dabauvalle, K. Loos, H. Merkert, and U. Scheer. Spontaneous assembly of pore complex-containing membranes (”annulate lamellae”) in Xenopus egg extract in the absence of chromatin. Journal of Cell Biology, 112(6):1073– ??, March 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/112/6/1073.

delaRosa:1991:TRA

[dlR91] E. J. de la Rosa. Topologically restricted appearance in the developing chick retinotectal system of Bravo, a neural surface protein: experimental modulation by environmental cues. Journal of Cell Biology, 112(5):1049–??, March 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/112/5/
1049.

delaRosa:1990:TRA

[dlRKR^{+}90] E. J. de la Rosa, J. F. Kayyem, J. M. Roman, Y. D. Stierhof, W. J. Dreyer, and U. Schwarz. Topologically restricted appearance in the developing chick retinotectal system of Bravo, a neural surface protein: experimental modulation by environmental cues. Journal of Cell Biology, 111(6): 3087–??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/111/6/3087.

Donaldson:1990:DKP

[DLSB^{+}90] J. G. Donaldson, J. Lippincott-Schwartz, G. S. Bloom, T. E. Kreis, and R. D. Klausner. Dissociation of a 110-kD peripheral membrane protein from the Golgi apparatus is an early event in brefeldin A action. Journal of Cell Biology, 111(6):2295–??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/111/6/2295.

Donaldson:1991:GNM

[DLSK91] J. G. Donaldson, J. Lippincott-Schwartz, and R. D. Klausner. Guanine nucleotides modulate the effects of brefeldin A in semipermeable cells: regulation of the association of a 110-kD peripheral membrane protein with the Golgi apparatus. Journal of Cell Biology, 112(4):579–??, February 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/112/4/
579.

Dunn:1992:DLS

[DM92] K. W. Dunn and F. R. Maxfield. Delivery of ligands from sorting endosomes to late endosomes occurs by maturation of sorting endosomes. Journal of Cell Biology, 117(2):301– ??, April 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/117/2/301.

Davidson:1992:ERE

[DMB92a] H. W. Davidson, C. H. McGowan, and W. E. Balch. Evidence for the regulation of exocytic transport by protein phosphorylation. Journal of Cell Biology, 116(6):1343–??, March 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/116/6/1343.

Duval:1992:HSA

[DMB92b] N. Duval, J. Massoulié, and S. Bon. H and T subunits of acetylcholinesterase from Torpedo, expressed in COS cells, generate all types of globular forms. Journal of Cell Biology, 118(3):641–??, August 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/118/3/641.

Davis:1993:ABP

[DMB93] J. Q. Davis, T. McLaughlin, and V. Bennett. Ankyrinbinding proteins related to nervous system cell adhesion molecules: candidates to provide transmembrane and intercellular connections in adult brain. Journal of Cell Biology, 121(1):121–??, April 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/121/1/121.

Dvir:1991:IED

[DMC^{+}91] A. Dvir, Y. Milner, O. Chomsky, C. Gilon, A. Gazit, and A. Levitzki. The inhibition of EGF-dependent proliferation of keratinocytes by tyrphostin tyrosine kinase blockers. Journal of Cell Biology, 113(4):857–??, May 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/113/4/857.

Durfee:1994:ATR

[DMJ^{+}94] T. Durfee, M. A. Mancini, D. Jones, S. J. Elledge, and W. H. Lee. The amino-terminal region of the retinoblastoma gene product binds a novel nuclear matrix protein that colocalizes to centers for RNA processing. Journal of Cell Biology, 127(3):609–??, November 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/127/3/609.

Deerinck:1994:FPE

[DMLR^{+}94] T. J. Deerinck, M. E. Martone, V. Lev-Ram, D. P. Green, R. Y. Tsien, D. L. Spector, S. Huang, and M. H. Ellisman. Fluorescence photooxidation with eosin: a method for high resolution immunolocalization and in situ hybridization detection for light and electron microscopy. Journal of Cell Biology, 126(4):901–??, August 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/126/4/901.

Deschamps:1991:TFE

[DMM^{+}91] S. Deschamps, J. Morales, A. Mazabraud, M. le Maire, H. Denis, and D. D. Brown. Two forms of elongation factor 1 alpha (EF-1 alpha O and 42Sp50), present in oocytes, but absent in somatic cells of Xenopus laevis. Journal of Cell Biology, 114(6):1109–??, September 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/114/6/1109.

Ding:1993:TDR

[DMM93] R. Ding, K. L. McDonald, and J. R. McIntosh. Threedimensional reconstruction and analysis of mitotic spindles from the yeast, Schizosaccharomyces pombe. Journal of Cell Biology, 120(1):141–??, January 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/120/1/141.

Downes:1990:COR

[DMWJ90] C. S. Downes, S. R. Musk, J. V. Watson, and R. T. Johnson. Caffeine overcomes a restriction point associated with DNA replication, but does not accelerate mitosis. Journal of Cell Biology, 110(6):1855–??, June 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/110/6/1855.

Denning:1992:ALC

[DOW92] G. M. Denning, L. S. Ostedgaard, and M. J. Welsh. Abnormal localization of cystic fibrosis transmembrane conductance regulator in primary cultures of cystic fibrosis airway epithelia. Journal of Cell Biology, 118(3):551–??, August 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/118/3/551.

Dean:1990:RPG

[DP90] N. Dean and H. R. Pelham. Recycling of proteins from the Golgi compartment to the ER in yeast. Journal of Cell Biology, 111(2):369–??, August 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/111/2/369.

Deretic:1991:PSR

[DP91] D. Deretic and D. S. Papermaster. Polarized sorting of rhodopsin on post-Golgi membranes in frog retinal photoreceptor cells. Journal of Cell Biology, 113(6):1281–??, June 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/113/6/1281.

Doberstein:1992:LSP

[DP92] S. K. Doberstein and T. D. Pollard. Localization and specificity of the phospholipid and actin binding sites on the tail of Acanthamoeba myosin IC. Journal of Cell Biology, 117 (6):1241–??, June 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/117/6/1241.

Drozdoff:1993:CDE

[DP93] V. Drozdoff and W. J. Pledger. Commitment to differentiation and expression of early differentiation markers in murine keratinocytes in vitro are regulated independently of extracellular calcium concentrations. Journal of Cell Biology, 123(4):909–??, November 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/123/4/909.

Dillman:1994:DPV

[DP94] J. F. Dillman and K. K. Pfister. Differential phosphorylation in vivo of cytoplasmic dynein associated with anterogradely moving organelles. Journal of Cell Biology, 127(6): 1671–??, December 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/127/6/1671.

Dadabay:1991:LCB

[DPCP91] C. Y. Dadabay, E. Patton, J. A. Cooper, and L. J. Pike. Lack of correlation between changes in polyphosphoinositide levels and actin/gelsolin complexes in A431 cells treated with epidermal growth factor. Journal of Cell Biology, 112 (6):1151–??, March 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/112/6/1151.

David-Pfeuty:1990:ICS

[DPND90] T. David-Pfeuty and Y. Nouvian-Dooghe. Immunolocalization of the cellular src protein in interphase and mitotic NIH c-src overexpresser cells. Journal of Cell Biology, 111(6): 3097–??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/111/6/3097.

DArcangelo:1993:NGF

[DPS^{+}93] G. D’Arcangelo, K. Paradiso, D. Shepherd, P. Brehm, S. Halegoua, and G. Mandel. Neuronal growth factor regulation of two different sodium channel types through distinct signal transduction pathways. Journal of Cell Biology, 122(4):915–??, August 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/122/4/915.

Daar:1991:CCF

[DPW91] I. Daar, R. S. Paules, and G. F. Vande Woude. A characterization of cytostatic factor activity from Xenopus eggs and c-mos-transformed cells. Journal of Cell Biology, 114 (2):329–??, July 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/114/2/329.

Daston:1991:EPP

[DR91] M. M. Daston and N. Ratner. Expression of P30, a protein with adhesive properties, in Schwann cells and neurons of the developing and regenerating peripheral nerve. Journal of Cell Biology, 112(6):1229–??, March 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/112/6/1229.

Duronio:1991:MAA

[DRJ^{+}91] R. J. Duronio, D. A. Rudnick, R. L. Johnson, D. R. Johnson, and J. I. Gordon. Myristic acid auxotrophy caused by mutation of S. cerevisiae myristoyl–CoA: protein Nmyristoyltransferase. Journal of Cell Biology, 113(6):1313– ??, June 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/113/6/1313.

Davis:1990:PDG

[DS90a] J. B. Davis and P. Stroobant. Platelet-derived growth factors and fibroblast growth factors are mitogens for rat Schwann cells. Journal of Cell Biology, 110(4):1353–??, April 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/110/4/1353.

Dedhar:1990:AIR

[DS90b] S. Dedhar and R. Saulnier. Alterations in integrin receptor expression on chemically transformed human cells: specific enhancement of laminin and collagen receptor complexes. Journal of Cell Biology, 110(2):481–??, February 1990. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/110/2/
481.

Dubel:1990:TDE

[DS90c] S. Dübel and H. C. Schaller. Terminal differentiation of ectodermal epithelial stem cells of Hydra can occur in G2 without requiring mitosis or S phase. Journal of Cell Biology, 110(4):939–??, April 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/110/4/939.

Diamond:1993:SMC

[DS93] M. S. Diamond and T. A. Springer. A subpopulation of Mac-1 (CD11b/CD18) molecules mediates neutrophil adhesion to ICAM-1 and fibrinogen. Journal of Cell Biology, 120(2):545–??, January 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/120/2/545.

deSilva:1990:QCE

[dSBH90] A. M. de Silva, W. E. Balch, and A. Helenius. Quality control in the endoplasmic reticulum: folding and misfolding of vesicular stomatitis virus G protein in cells and in vitro. Journal of Cell Biology, 111(3):857–??, September 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/111/3/

deSilva:1993:PFV

[dSBH93] A. de Silva, I. Braakman, and A. Helenius. Posttranslational folding of vesicular stomatitis virus G protein in the ER: involvement of noncovalent and covalent complexes. Journal of Cell Biology, 120(3):647–??, February 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/120/3/647.

Diamond:1990:ICC

[DSdF^{+}90] M. S. Diamond, D. E. Staunton, A. R. de Fougerolles, S. A. Stacker, J. Garcia-Aguilar, M. L. Hibbs, and T. A. Springer. ICAM-1 (CD54): a counter-receptor for Mac-1 (CD11b/CD18). Journal of Cell Biology, 111(6):3129– ??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/111/6/3129.

Dufour:1994:DPM

[DSJB^{+}94] S. Dufour, J. P. Saint-Jeannet, F. Broders, D. Wedlich, and J. P. Thiery. Differential perturbations in the morphogenesis of anterior structures induced by overexpression of truncated XB- and N-cadherins in Xenopus embryos. Journal of Cell Biology, 127(2):521–??, October 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/127/2/521.

daSilva:1990:RPM

[dSK90] A. M. da Silva and C. Klein. A rapid posttranslational myristylation of a 68-kD protein in D. discoideum. Journal of Cell Biology, 111(2):401–??, August 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/111/2/401.

DiMilla:1993:MMH

[DSQ^{+}93] P. A. DiMilla, J. A. Stone, J. A. Quinn, S. M. Albelda, and D. A. Lauffenburger. Maximal migration of human smooth muscle cells on fibronectin and type IV collagen occurs at an intermediate attachment strength. Journal of Cell Biology, 122(3):729–??, August 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/122/3/729.

Doherty:1993:TAT

[DSR^{+}93] P. Doherty, A. Singh, G. Rimon, S. R. Bolsover, and F. S. Walsh. Thy-1 antibody-triggered neurite outgrowth requires an influx of calcium into neurons via N- and L-type calcium channels. Journal of Cell Biology, 122(1):181–??, July 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/122/1/
181.

deToledo:1990:PCM

[dTF90] G. Alvarez de Toledo and J. M. Fernandez. Patch-clamp measurements reveal multimodal distribution of granule sizes in rat mast cells. Journal of Cell Biology, 110(4): 1033–??, April 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/110/4/1033.

Dunn:1992:HCS

[DTY92] J. C. Dunn, R. G. Tompkins, and M. L. Yarmush. Hepatocytes in collagen sandwich: evidence for transcriptional and translational regulation. Journal of Cell Biology, 116(4): 1043–??, February 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/116/4/1043.

Dunn:1990:SMAa

[Dun90a] W. A. Dunn. Studies on the mechanisms of autophagy: formation of the autophagic vacuole. Journal of Cell Biology, 110(6):1923–??, June 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/110/6/1923.

Dunn:1990:SMAb

[Dun90b] W. A. Dunn. Studies on the mechanisms of autophagy: maturation of the autophagic vacuole. Journal of Cell Biology, 110(6):1935–??, June 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/110/6/1935.

DeStrooper:1993:SSS

[DUVV93] B. De Strooper, L. Umans, F. Van Leuven, and H. Van Den Berghe. Study of the synthesis and secretion of normal and artificial mutants of murine amyloid precursor protein (APP): cleavage of APP occurs in a late compartment of the default secretion pathway. Journal of Cell Biology, 121 (2):295–??, April 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/121/2/295.

David:1992:MCA

[DvdSM^{+}92] G. David, B. van der Schueren, P. Marynen, J. J. Cassiman, and H. van den Berghe. Molecular cloning of amphiglycan, a novel integral membrane heparan sulfate proteoglycan expressed by epithelial and fibroblastic cells. Journal of Cell Biology, 118(4):961–??, August 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/118/4/961.

Defilippi:1991:DDM

[DvHB^{+}91] P. Defilippi, V. van Hinsbergh, A. Bertolotto, P. Rossino, L. Silengo, and G. Tarone. Differential distribution and modulation of expression of alpha 1/beta 1 integrin on human endothelial cells. Journal of Cell Biology, 114(4): 855–??, August 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/114/4/855.

DelCastillo:1992:CPD

[DVT92] A. Rodríguez Del Castillo, M. L. Vitale, and J. M. Trifaró. Ca^{2+} and pH determine the interaction of chromaffin cell scinderin with phosphatidylserine and phosphatidylinositol 4,5,-biphosphate and its cellular distribution during nicotinic-receptor stimulation and protein kinase C activation. Journal of Cell Biology, 119(4):797–??, November 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/119/4/797.

Degen:1991:PNK

[DW91] E. Degen and D. B. Williams. Participation of a novel 88kD protein in the biogenesis of murine class I histocompatibility molecules. Journal of Cell Biology, 112(6):1099–??, March 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/112/6/1099.

Dzamba:1993:FBS

[DWJP93] B. J. Dzamba, H. Wu, R. Jaenisch, and D. M. Peters. Fibronectin binding site in type I collagen regulates fibronectin fibril formation. Journal of Cell Biology, 121(5): 1165–??, June 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/121/5/1165.

dEnfert:1991:SDM

[dWLS91] C. d’Enfert, L. J. Wuestehube, T. Lila, and R. Schekman. Sec12p-dependent membrane binding of the small GTPbinding protein Sar1p promotes formation of transport vesicles from the ER. Journal of Cell Biology, 114(4):663–??, August 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/114/4/663.

Dryden:1993:ESR

[DWY^{+}93] K. A. Dryden, G. Wang, M. Yeager, M. L. Nibert, K. M. Coombs, D. B. Furlong, B. N. Fields, and T. S. Baker. Early steps in reovirus infection are associated with dramatic changes in supramolecular structure and protein conformation: analysis of virions and subviral particles by cryoelectron microscopy and image reconstruction. Journal of Cell Biology, 122(5):1023–??, September 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/122/5/1023.

Dlugosz:1993:CCG

[DY93] A. A. Dlugosz and S. H. Yuspa. Coordinate changes in gene expression which mark the spinous to granular cell transition in epidermis are regulated by protein kinase C. Journal of Cell Biology, 120(1):217–??, January 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/120/1/217.

Dwivedi:1990:ARE

[DYS^{+}90] R. S. Dwivedi, A. V. Yeldandi, V. Subbarao, P. Feigelson, A. K. Roy, J. K. Reddy, and M. S. Rao. Androgen regulated expression of the alpha 2u-globulin gene in pancreatic hepatocytes of rat. Journal of Cell Biology, 110(2): 263–??, February 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/110/2/263.

Daar:1993:IMI

[DYV93] I. Daar, N. Yew, and G. F. Vande Woude. Inhibition of mos-induced oocyte maturation by protein kinase A. Journal of Cell Biology, 120(5):1197–??, March 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/120/5/1197.

Djabali:1993:PEH

[DZdH^{+}93] K. Djabali, A. Zissopoulou, M. J. de Hoop, S. D. Georgatos, and C. G. Dotti. Peripherin expression in hippocampal neurons induced by muscle soluble factor(s). Journal of Cell Biology, 123(5):1197–??, December 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/123/5/1197.

Eveleth:1992:NGF

[EB92] D. D. Eveleth and R. A. Bradshaw. Nerve growth factor nonresponsive pheochromocytoma cells: altered internalization results in signaling dysfunction. Journal of Cell Biology, 117(2):291–??, April 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/117/2/291.

Egelhoff:1991:STC

[EBS91] T. T. Egelhoff, S. S. Brown, and J. A. Spudich. Spatial and temporal control of nonmuscle myosin localization: identification of a domain that is necessary for myosin filament disassembly in vivo. Journal of Cell Biology, 112(4):677– ??, February 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/112/4/677.

Ervasti:1993:RDG

[EC93] J. M. Ervasti and K. P. Campbell. A role for the dystrophinglycoprotein complex as a transmembrane linker between laminin and actin. Journal of Cell Biology, 122(4):809–??, August 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/122/4/809.

Evans:1993:IAI

[ECAG93] S. S. Evans, R. P. Collea, M. M. Appenheimer, and S. O. Gollnick. Interferon-alpha induces the expression of the Lselectin homing receptor in human B lymphoid cells. Journal of Cell Biology, 123(6):1889–??, December 1993. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/123/6/
1889.

Espreafico:1992:PSC

[ECM^{+}92] E. M. Espreafico, R. E. Cheney, M. Matteoli, A. A. Nascimento, P. V. De Camilli, R. E. Larson, and M. S. Mooseker. Primary structure and cellular localization of chicken brain myosin–V (p190), an unconventional myosin with calmodulin light chains. Journal of Cell Biology, 119(6):1541–??, December 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/119/6/1541.

Epstein:1993:MPC

[ECO93] H. F. Epstein, D. L. Casey, and I. Ortiz. Myosin and paramyosin of Caenorhabditis elegans embryos assemble into nascent structures distinct from thick filaments and multi-filament assemblages. Journal of Cell Biology, 122(4): 845–??, August 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/122/4/845.

Espindola:1992:BIC

[EEC^{+}92] F. S. Espindola, E. M. Espreafico, M. V. Coelho, A. R. Martins, F. R. Costa, M. S. Mooseker, and R. E. Larson. Biochemical and immunological characterization of p190-calmodulin complex from vertebrate brain: a novel calmodulin-binding myosin. Journal of Cell Biology, 118 (2):359–??, July 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/118/2/359.

Eppenberger-Eberhardt:1991:ARC

[EERM^{+}91] M. Eppenberger-Eberhardt, I. Riesinger, M. Messerli, P. Schwarb, M. Müller, H. M. Eppenberger, and T. Wallimann. Adult rat cardiomyocytes cultured in creatinedeficient medium display large mitochondria with paracrystalline inclusions, enriched for creatine kinase. Journal of Cell Biology, 113(2):289–??, April 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/113/2/289.

Eidelman:1993:HCA

[EFD^{+}93] F. J. Eidelman, A. Fuks, L. DeMarte, M. Taheri, and C. P. Stanners. Human carcinoembryonic antigen, an intercellular adhesion molecule, blocks fusion and differentiation of rat myoblasts. Journal of Cell Biology, 123(2):467–??, October 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/123/2/467.

Ekblom:1993:DTE

[EFTJ^{+}93] M. Ekblom, R. Fässler, B. Tomasini-Johansson, K. Nilsson, and P. Ekblom. Downregulation of tenascin expression by glucocorticoids in bone marrow stromal cells and in fibroblasts. Journal of Cell Biology, 123(4):1037–??, November 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/123/4/1037.

Einstein:1991:CLS

[EG91] R. Einstein and C. A. Gabel. Cell- and ligand-specific dephosphorylation of acid hydrolases: evidence that the mannose 6-phosphatase is controlled by compartmentalization. Journal of Cell Biology, 112(1):81–??, January 1991. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/112/1/
81.

Ehrlich:1991:TNP

[EGP^{+}91] H. J. Ehrlich, R. K. Gebbink, K. T. Preissner, J. Keijer, N. L. Esmon, K. Mertens, and H. Pannekoek. Thrombin neutralizes plasminogen activator inhibitor 1 (PAI-1) that is complexed with vitronectin in the endothelial cell matrix. Journal of Cell Biology, 115(6):1773–??, December 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/115/6/

Emans:1993:AIM

[EGW^{+}93] N. Emans, J. P. Gorvel, C. Walter, V. Gerke, R. Kellner, G. Griffiths, and J. Gruenberg. Annexin II is a major component of fusogenic endosomal vesicles. Journal of Cell Biology, 120(6):1357–??, March 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/120/6/1357.

Elmendorf:1994:PFE

[EH94] H. G. Elmendorf and K. Haldar. Plasmodium falciparum exports the Golgi marker sphingomyelin synthase into a tubovesicular network in the cytoplasm of mature erythrocytes. Journal of Cell Biology, 124(4):449–??, February 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/124/4/449.

Elkins:1990:DFN

[EHB^{+}90] T. Elkins, M. Hortsch, A. J. Bieber, P. M. Snow, and C. S. Goodman. Drosophila fasciclin I is a novel homophilic adhesion molecule that along with fasciclin III can mediate cell sorting. Journal of Cell Biology, 110(5):1825–??, May 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/110/5/
1825.

Emtage:1993:MEE

[EJ93] J. L. Emtage and R. E. Jensen. MAS6 encodes an essential inner membrane component of the yeast mitochondrial protein import pathway. Journal of Cell Biology, 122(5):1003– ??, September 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/122/5/1003.

Eichner:1990:DEK

[EK90] R. Eichner and M. Kahn. Differential extraction of keratin subunits and filaments from normal human epidermis. Journal of Cell Biology, 110(4):1149–??, April 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/110/4/1149.

Eilertsen:1992:ICT

[EK92] K. J. Eilertsen and T. C. Keller. Identification and characterization of two huge protein components of the brush border cytoskeleton: evidence for a cellular isoform of titin. Journal of Cell Biology, 119(3):549–??, November 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/119/3/
549.

Eilertsen:1994:CTL

[EKK94] K. J. Eilertsen, S. T. Kazmierski, and T. C. Keller. Cellular titin localization in stress fibers and interaction with myosin II filaments in vitro. Journal of Cell Biology, 126(5):1201– ??, September 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/126/5/1201.

Evans:1993:DNM

[ELS93] S. C. Evans, L. C. Lopez, and B. D. Shur. Dominant negative mutation in cell surface beta 1,4-galactosyltransferase inhibits cell–cell and cell-matrix interactions. Journal of Cell Biology, 120(4):1045–??, February 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/120/4/1045.

Evans:1993:SCS

[ELZ93] E. Evans, A. Leung, and D. Zhelev. Synchrony of cell spreading and contraction force as phagocytes engulf large pathogens. Journal of Cell Biology, 122(6):1295– ??, September 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/122/6/1295.

Estreicher:1990:RUT

[EMC^{+}90] A. Estreicher, J. Mühlhauser, J. L. Carpentier, L. Orci, and J. D. Vassalli. The receptor for urokinase type plasminogen activator polarizes expression of the protease to the leading edge of migrating monocytes and promotes degradation of enzyme inhibitor complexes. Journal of Cell Biology, 111(2):783–??, August 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/111/2/783.

Einheber:1993:ARS

[EMGS93] S. Einheber, T. A. Milner, F. Giancotti, and J. L. Salzer. Axonal regulation of Schwann cell integrin expression suggests a role for alpha 6 beta 4 in myelination. Journal of Cell Biology, 123(5):1223–??, December 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/123/5/1223.

Eichinger:1991:DSA

[ENS91] L. Eichinger, A. A. Noegel, and M. Schleicher. Domain structure in actin-binding proteins: expression and functional characterization of truncated severin. Journal of Cell Biology, 112(4):665–??, February 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/112/4/665.

Ely:1990:KTK

[EOL^{+}90] C. M. Ely, K. M. Oddie, J. S. Litz, A. J. Rossomando, S. B. Kanner, T. W. Sturgill, and S. J. Parsons. A 42-kD tyrosine kinase substrate linked to chromaffin cell secretion exhibits an associated MAP kinase activity and is highly related to a 42-kD mitogen-stimulated protein in fibroblasts. Journal of Cell Biology, 110(3):731–??, March 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/110/3/731.

Elazar:1994:ARF

[EOO^{+}94] Z. Elazar, L. Orci, J. Ostermann, M. Amherdt, G. Tanigawa, and J. E. Rothman. ADP-ribosylation factor and coatomer couple fusion to vesicle budding. Journal of Cell Biology, 124(4):415–??, February 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/124/4/415.

Erxleben:1994:CRS

[EP94] C. Erxleben and H. Plattner. Ca^{2+} release from subplasmalemmal stores as a primary event during exocytosis in Paramecium cells. Journal of Cell Biology, 127(4):935–??, November 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/127/4/935.

Enoch:1991:PAL

[EPNN91] T. Enoch, M. Peter, P. Nurse, and E. A. Nigg. p34cdc2 acts as a lamin kinase in fission yeast. Journal of Cell Biology, 112(5):797–??, March 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/112/5/797.

Erickson:1993:GKC

[Eri93] H. P. Erickson. Gene knockouts of c-src, transforming growth factor beta 1, and tenascin suggest superfluous, nonfunctional expression of proteins. Journal of Cell Biology, 120(5):1079–??, March 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/120/5/1079.

Edidin:1991:DBL

[ES91] M. Edidin and I. Stroynowski. Differences between the lateral organization of conventional and inositol phospholipidanchored membrane proteins. A further definition of micrometer scale membrane domains. Journal of Cell Biology, 112(6):1143–??, March 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/112/6/1143.

Euteneuer:1992:MCP

[ES92] U. Euteneuer and M. Schliwa. Mechanism of centrosome positioning during the wound response in BSC-1 cells. Journal of Cell Biology, 116(5):1157–??, March 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/116/5/1157.

Ewert:1990:SSE

[ESL^{+}90] M. Ewert, B. D. Shivers, H. Lüddens, H. Möhler, and P. H. Seeburg. Subunit selectivity and epitope characterization of mAbs directed against the GABAA/benzodiazepine receptor. Journal of Cell Biology, 110(6):2043–??, June 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/110/6/
2043.

el-Sabban:1991:CDT

[eSP91] M. E. el Sabban and B. U. Pauli. Cytoplasmic dye transfer between metastatic tumor cells and vascular endothelium.

Journal of Cell Biology, 115(5):1375–??, December 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/115/5/
1375.

Ettinger:1991:PAC

[ET91] W. F. Ettinger and S. M. Theg. Physiologically active chloroplasts contain pools of unassembled extrinsic proteins of the photosynthetic oxygen-evolving enzyme complex in the thylakoid lumen. Journal of Cell Biology, 115(2):321–??, October 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/115/2/321.

Edwards:1994:CAC

[ET94] S. N. Edwards and A. M. Tolkovsky. Characterization of apoptosis in cultured rat sympathetic neurons after nerve growth factor withdrawal. Journal of Cell Biology, 124(4): 537–??, February 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/124/4/537.

Elices:1991:RFI

[EUH91] M. J. Elices, L. A. Urry, and M. E. Hemler. Receptor functions for the integrin VLA-3: fibronectin, collagen, and laminin binding are differentially influenced by Arg–Gly– Asp peptide and by divalent cations. Journal of Cell Biology, 112(1):169–??, January 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/112/1/169.

Elenius:1991:IES

[EVL^{+}91] K. Elenius, S. Vainio, M. Laato, M. Salmivirta, I. Thesleff, and M. Jalkanen. Induced expression of syndecan in healing wounds. Journal of Cell Biology, 114(3):585–??, August 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/114/3/585.

Eliott:1991:DMI

[EVW91] S. Eliott, P. H. Vardy, and K. L. Williams. The distribution of myosin II in Dictyostelium discoideum slug cells. Journal of Cell Biology, 115(5):1267–??, December 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/115/5/1267.

Erbe:1992:ISR

[EWP^{+}92] D. V. Erbe, B. A. Wolitzky, L. G. Presta, C. R. Norton, R. J. Ramos, D. K. Burns, J. M. Rumberger, B. N. Rao, C. Foxall, and B. K. Brandley. Identification of an E-selectin region critical for carbohydrate recognition and cell adhesion. Journal of Cell Biology, 119(1):215–??, October 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/119/1/
215.

Erbe:1993:PSU

[EWP^{+}93a] D. V. Erbe, S. R. Watson, L. G. Presta, B. A. Wolitzky, C. Foxall, B. K. Brandley, and L. A. Lasky. P- and Eselectin use common sites for carbohydrate ligand recognition and cell adhesion. Journal of Cell Biology, 120(5): 1227–??, March 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/120/5/1227.

Ersfeld:1993:CTT

[EWP^{+}93b] K. Ersfeld, J. Wehland, U. Plessmann, H. Dodemont, V. Gerke, and K. Weber. Characterization of the tubulintyrosine ligase. Journal of Cell Biology, 120(3):725–??, February 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/120/3/725.

Estus:1994:AGE

[EZF^{+}94] S. Estus, W. J. Zaks, R. S. Freeman, M. Gruda, R. Bravo, and E. M. Johnson. Altered gene expression in neurons during programmed cell death: identification of c-jun as necessary for neuronal apoptosis. Journal of Cell Biology, 127(6):1717–??, December 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/127/6/1717.

Feldherr:1990:PNE

[FA90] C. M. Feldherr and D. Akin. The permeability of the nuclear envelope in dividing and nondividing cell cultures. Journal of Cell Biology, 111(1):1–??, July 1990. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/111/1/1.

Feldherr:1991:SMN

[FA91] C. M. Feldherr and D. Akin. Signal-mediated nuclear transport in proliferating and growth-arrested BALB/c 3T3 cells. Journal of Cell Biology, 115(4):933–??, November 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/115/4/
933.

Fowler:1992:SRC

[FA92] V. M. Fowler and E. J. Adam. Spectrin redistributes to the cytosol and is phosphorylated during mitosis in cultured cells. Journal of Cell Biology, 119(6):1559–??, December 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/119/6/1559.

Fox:1990:RMS

[FABS90] J. E. Fox, C. D. Austin, J. K. Boyles, and P. K. Steffen. Role of the membrane skeleton in preventing the shedding of procoagulant-rich microvesicles from the platelet plasma membrane. Journal of Cell Biology, 111(2):483–??, August 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/111/2/483.

Flucher:1993:TFO

[FAF^{+}93] B. E. Flucher, S. B. Andrews, S. Fleischer, A. R. Marks, A. Caswell, and J. A. Powell. Triad formation: organization and function of the sarcoplasmic reticulum calcium release channel and triadin in normal and dysgenic muscle in vitro. Journal of Cell Biology, 123(5):1161–??, December 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/123/5/
1161.

Flaumenhaft:1992:BFG

[FAMR92] R. Flaumenhaft, M. Abe, P. Mignatti, and D. B. Rifkin. Basic fibroblast growth factor-induced activation of latent transforming growth factor beta in endothelial cells: regulation of plasminogen activator activity. Journal of Cell Biology, 118(4):901–??, August 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/118/4/901.

Flaumenhaft:1993:RLT

[FAS^{+}93] R. Flaumenhaft, M. Abe, Y. Sato, K. Miyazono, J. Harpel, C. H. Heldin, and D. B. Rifkin. Role of the latent TGFbeta binding protein in the activation of latent TGF-beta by co-cultures of endothelial and smooth muscle cells. Journal of Cell Biology, 120(4):995–??, February 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/120/4/995.

Fleming:1993:FHA

[FAU^{+}93] W. H. Fleming, E. J. Alpern, N. Uchida, K. Ikuta, G. J. Spangrude, and I. L. Weissman. Functional heterogeneity is associated with the cell cycle status of murine hematopoietic stem cells. Journal of Cell Biology, 122(4):897–??, August 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/122/4/897.

Felix:1994:CAV

[FAWM94] M. A. Félix, C. Antony, M. Wright, and B. Maro. Centrosome assembly in vitro: role of gamma-tubulin recruitment in Xenopus sperm aster formation. Journal of Cell Biology, 124(1):19–??, January 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/124/1/19.

Fogerty:1990:IBF

[FAYM90] F. J. Fogerty, S. K. Akiyama, K. M. Yamada, and D. F. Mosher. Inhibition of binding of fibronectin to matrix assembly sites by anti-integrin (alpha 5 beta 1) antibodies. Journal of Cell Biology, 111(2):699–??, August 1990. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/111/2/
699.

Fath:1993:GDV

[FB93] K. R. Fath and D. R. Burgess. Golgi-derived vesicles from developing epithelial cells bind actin filaments and possess myosin–I as a cytoplasmically oriented peripheral membrane protein. Journal of Cell Biology, 120(1):117–??, January 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/120/1/117.

Fernandez:1992:PPT

[FBL92] A. Fernandez, D. L. Brautigan, and N. J. Lamb. Protein phosphatase type 1 in mammalian cell mitosis: chromosomal localization and involvement in mitotic exit. Journal of Cell Biology, 116(6):1421–??, March 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/116/6/1421.

Fitch:1990:SOD

[FBLL90] J. M. Fitch, D. E. Birk, C. Linsenmayer, and T. F. Linsenmayer. The spatial organization of Descemet’s membraneassociated type IV collagen in the avian cornea. Journal of Cell Biology, 110(4):1457–??, April 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/110/4/1457.

Fernandez:1990:PPT

[FBML90] A. Fernandez, D. L. Brautigan, M. Mumby, and N. J. Lamb. Protein phosphatase type-1, not type-2A, modulates actin microfilament integrity and myosin light chain phosphorylation in living nonmuscle cells. Journal of Cell Biology, 111 (1):103–??, July 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/111/1/103.

Frixen:1991:CMC

[FBS^{+}91] U. H. Frixen, J. Behrens, M. Sachs, G. Eberle, B. Voss, A. Warda, D. Löchner, and W. Birchmeier. E-cadherin-mediated cell–cell adhesion prevents invasiveness of human carcinoma cells. Journal of Cell Biology, 113(1):173–??, April 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/113/1/173.

French:1993:MCM

[FCD^{+}93] L. E. French, A. Chonn, D. Ducrest, B. Baumann, D. Belin, A. Wohlwend, J. Z. Kiss, A. P. Sappino, J. Tschopp, and J. A. Schifferli. Murine clusterin: molecular cloning and mRNA localization of a gene associated with epithelial differentiation processes during embryogenesis. Journal of Cell Biology, 122(5):1119–??, September 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/122/5/1119.

Febvre-Chevalier:1992:MDV

[FCF92] C. Febvre-Chevalier and J. Febvre. Microtubule dissassembly in vivo: intercalary destabilization and breakdown of microtubules in the heliozoan Actinocoryne contractilis. Journal of Cell Biology, 118(3):585–??, August 1992. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/118/3/
585.

Ferreira:1994:DIS

[FCFL94] J. A. Ferreira, M. Carmo-Fonseca, and A. I. Lamond. Differential interaction of splicing snRNPs with coiled bodies and interchromatin granules during mitosis and assembly of daughter cell nuclei. Journal of Cell Biology, 126 (1):11–??, July 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/126/1/11.

Faissner:1994:INC

[FCL^{+}94] A. Faissner, A. Clement, A. Lochter, A. Streit, C. Mandl, and M. Schachner. Isolation of a neural chondroitin sulfate proteoglycan with neurite outgrowth promoting properties. Journal of Cell Biology, 126(3):783–??, August 1994. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/126/3/
783.

Fishkind:1991:MCF

[FCW91] D. J. Fishkind, L. G. Cao, and Y. L. Wang. Microinjection of the catalytic fragment of myosin light chain kinase into dividing cells: effects on mitosis and cytokinesis. Journal of Cell Biology, 114(5):967–??, September 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/114/5/967.

Fukui:1990:SFC

[FDS90] Y. Fukui, A. De Lozanne, and J. A. Spudich. Structure and function of the cytoskeleton of a Dictyostelium myosindefective mutant. Journal of Cell Biology, 110(2):367–??, February 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/110/2/367.

Fischer:1991:DSR

[FDT^{+}91] U. Fischer, E. Darzynkiewicz, S. M. Tahara, N. A. Dathan, R. Lührmann, and I. W. Mattaj. Diversity in the signals required for nuclear accumulation of U snRNPs and variety in the pathways of nuclear transport. Journal of Cell Biology, 113(4):705–??, May 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/113/4/705.

Felder:1990:MFL

[FE90] S. Felder and E. L. Elson. Mechanics of fibroblast locomotion: quantitative analysis of forces and motions at the leading lamellas of fibroblasts. Journal of Cell Biology, 111(6): 2513–??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/111/6/2513.

Fechheimer:1993:DCD

[FF93] M. Fechheimer and R. Furukawa. A 27,000-D core of the Dictyostelium 34,000-D protein retains Ca(2+)-regulated actin cross-linking but lacks bundling activity. Journal of Cell Biology, 120(5):1169–??, March 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/120/5/1169.

Frisch:1994:DEC

[FF94] S. M. Frisch and H. Francis. Disruption of epithelial cellmatrix interactions induces apoptosis. Journal of Cell Biology, 124(4):619–??, February 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/124/4/619.

Franck:1990:MVC

[FFB90] Z. Franck, M. Footer, and A. Bretscher. Microinjection of villin into cultured cells induces rapid and long-lasting changes in cell morphology but does not inhibit cytokinesis, cell motility, or membrane ruffling. Journal of Cell Biology, 111(6):2475–??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/111/6/2475.

Funakoshi:1993:DEM

[FFB^{+}93] H. Funakoshi, J. Frisén, G. Barbany, T. Timmusk, O. Zachrisson, V. M. Verge, and H. Persson. Differential expression of mRNAs for neurotrophins and their receptors after axotomy of the sciatic nerve. Journal of Cell Biology, 123(2):455–??, October 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/123/2/455.

Feiguin:1994:KMO

[FFKC94] F. Feiguin, A. Ferreira, K. S. Kosik, and A. Caceres. Kinesin-mediated organelle translocation revealed by specific cellular manipulations. Journal of Cell Biology, 127 (4):1021–??, November 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/127/4/1021.

Finidori:1992:VAF

[FFKL92] J. Finidori, E. Friederich, D. J. Kwiatkowski, and D. Louvard. In vivo analysis of functional domains from villin and gelsolin. Journal of Cell Biology, 116(5):1145–??, March 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/116/5/1145.

Frohlich:1991:YCC

[FFR^{+}91] K. U. Fröhlich, H. W. Fries, M. Rüdiger, R. Erdmann, D. Botstein, and D. Mecke. Yeast cell cycle protein CDC48p shows full-length homology to the mammalian protein VCP and is a member of a protein family involved in secretion, peroxisome formation, and gene expression. Journal of Cell Biology, 114(3):443–??, August 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/114/3/443.

Futter:1993:APM

[FFS^{+}93] C. E. Futter, S. Felder, J. Schlessinger, A. Ullrich, and C. R. Hopkins. Annexin I is phosphorylated in the multivesicular body during the processing of the epidermal growth factor receptor. Journal of Cell Biology, 120(1):77–??, January 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/120/1/77.

Foisner:1991:MAM

[FFSW91] R. Foisner, B. Feldman, L. Sander, and G. Wiche. Monoclonal antibody mapping of structural and functional plectin epitopes. Journal of Cell Biology, 112(3):397–??, February 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/112/3/397.

Feick:1991:IMP

[FFW91] P. Feick, R. Foisner, and G. Wiche. Immunolocalization and molecular properties of a high molecular weight microtubule-bundling protein (syncolin) from chicken erythrocytes. Journal of Cell Biology, 112(4):689–??, February 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/112/4/689.

Forsayeth:1992:BFS

[FGH92] J. R. Forsayeth, Y. Gu, and Z. W. Hall. BiP forms stable complexes with unassembled subunits of the acetylcholine receptor in transfected COS cells and in C2 muscle cells. Journal of Cell Biology, 117(4):841–??, May 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/117/4/841.

Feder:1990:CCA

[FGK^{+}90] J. N. Feder, C. J. Guidos, B. Kusler, C. Carswell, D. Lewis, and R. T. Schimke. A cell cycle analysis of growth-related genes expressed during T lymphocyte maturation. Journal of Cell Biology, 111(6):2693–??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/111/6/2693.

Fuhrer:1991:EAR

[FGS91] C. Fuhrer, I. Geffen, and M. Spiess. Endocytosis of the ASGP receptor H1 is reduced by mutation of tyrosine-5 but still occurs via coated pits. Journal of Cell Biology, 114(3): 423–??, August 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/114/3/423.

Fernandez:1992:STT

[FGS^{+}92] J. L. Rodríguez Fernández, B. Geiger, D. Salomon, I. Sabanay, M. Zöller, and A. Ben-Ze’ev. Suppression of tumorigenicity in transformed cells after transfection with vinculin cDNA. Journal of Cell Biology, 119(2):427–??, October 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/119/2/427.

Fernandez:1993:SVE

[FGSBZ93] J. L. Rodríguez Fernández, B. Geiger, D. Salomon, and A. Ben-Ze’ev. Suppression of vinculin expression by antisense transfection confers changes in cell morphology, motility, and anchorage-dependent growth of 3T3 cells. Journal of Cell Biology, 122(6):1285–??, September 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/122/6/1285.

Furuhashi:1990:CAF

[FH90] K. Furuhashi and S. Hatano. Control of actin filament length by phosphorylation of fragmin-actin complex. Journal of Cell Biology, 111(3):1081–??, September 1990. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/111/3/
1081.

Fransen:1991:NOM

[FHGN91] J. A. Fransen, H. P. Hauri, L. A. Ginsel, and H. Y. Naim. Naturally occurring mutations in intestinal sucraseisomaltase provide evidence for the existence of an intracellular sorting signal in the isomaltase subunit. Journal of Cell Biology, 115(1):45–??, October 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/115/1/45.

Funatsu:1990:EFS

[FHI90] T. Funatsu, H. Higuchi, and S. Ishiwata. Elastic filaments in skeletal muscle revealed by selective removal of thin filaments with plasma gelsolin. Journal of Cell Biology, 110(1): 53–??, January 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/110/1/53.

Furuse:1993:ONI

[FHI^{+}93] M. Furuse, T. Hirase, M. Itoh, A. Nagafuchi, S. Yonemura, S. Tsukita, and S. Tsukita. Occludin: a novel integral membrane protein localizing at tight junctions. Journal of Cell Biology, 123(6):1777–??, December 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/123/6/1777.

Funabiki:1993:CCD

[FHUY93] H. Funabiki, I. Hagan, S. Uzawa, and M. Yanagida. Cell cycle-dependent specific positioning and clustering of centromeres and telomeres in fission yeast. Journal of Cell Biology, 121(5):961–??, June 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/121/5/961.

Fischer:1994:NTU

[FHvZ^{+}94] U. Fischer, J. Heinrich, K. van Zee, E. Fanning, and R. Lührmann. Nuclear transport of U1 snRNP in somatic cells: differences in signal requirement compared with Xenopus laevis oocytes. Journal of Cell Biology, 125(5):971–??, June 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/125/5/971.

Furuse:1994:DAO

[FIH^{+}94] M. Furuse, M. Itoh, T. Hirase, A. Nagafuchi, S. Yonemura, S. Tsukita, and S. Tsukita. Direct association of occludin with ZO-1 and its possible involvement in the localization of occludin at tight junctions. Journal of Cell Biology, 127(6): 1617–??, December 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/127/6/1617.

Fuortes:1993:ADP

[FJN93] M. Fuortes, W. W. Jin, and C. Nathan. Adhesiondependent protein tyrosine phosphorylation in neutrophils treated with tumor necrosis factor. Journal of Cell Biology, 120(3):777–??, February 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/120/3/777.

Fuortes:1994:BID

[FJN94] M. Fuortes, W. W. Jin, and C. Nathan. Beta 2 integrindependent tyrosine phosphorylation of paxillin in human neutrophils treated with tumor necrosis factor. Journal of Cell Biology, 127(5):1477–??, December 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/127/5/1477.

Fehon:1991:CCS

[FJRAT91] R. G. Fehon, K. Johansen, I. Rebay, and S. Artavanis-Tsakonas. Complex cellular and subcellular regulation of notch expression during embryonic and imaginal development of Drosophila: implications for notch function. Journal of Cell Biology, 113(3):657–??, May 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/113/3/657.

Freeman:1990:EVM

[FKB^{+}90a] R. S. Freeman, J. P. Kanki, S. M. Ballantyne, K. M. Pickham, and D. J. Donoghue. Effects of the v-mos oncogene on Xenopus development: meiotic induction in oocytes and mitotic arrest in cleaving embryos. Journal of Cell Biology, 111(2):533–??, August 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/111/2/533.

Fyrberg:1990:MGD

[FKB^{+}90b] E. Fyrberg, M. Kelly, E. Ball, C. Fyrberg, and M. C. Reedy. Molecular genetics of Drosophila alpha-actinin: mutant alleles disrupt Z disc integrity and muscle insertions. Journal of Cell Biology, 110(6):1999–??, June 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/110/6/1999.

Falk:1994:MIG

[FKG94] M. M. Falk, N. M. Kumar, and N. B. Gilula. Membrane insertion of gap junction connexins: polytopic channel forming membrane proteins. Journal of Cell Biology, 127(2): 343–??, October 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/127/2/343.

Funatsu:1993:EFS

[FKH^{+}93] T. Funatsu, E. Kono, H. Higuchi, S. Kimura, S. Ishiwata, T. Yoshioka, K. Maruyama, and S. Tsukita. Elastic filaments in situ in cardiac muscle: deep-etch replica analysis in combination with selective removal of actin and myosin filaments. Journal of Cell Biology, 120(3):711–??, February 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/120/3/711.

Faull:1993:AMI

[FKHG93] R. J. Faull, N. L. Kovach, J. M. Harlan, and M. H. Ginsberg. Affinity modulation of integrin alpha 5 beta 1: regulation of the functional response by soluble fibronectin. Journal of Cell Biology, 121(1):155–??, April 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/121/1/155.

Fenton:1992:RIA

[FKLS92] R. G. Fenton, H. F. Kung, D. L. Longo, and M. R. Smith. Regulation of intracellular actin polymerization by prenylated cellular proteins. Journal of Cell Biology, 117(2):347– ??, April 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/117/2/347.

Firmbach-Kraft:1993:RCD

[FKS93] I. Firmbach-Kraft and R. Stick. The role of CaaXdependent modifications in membrane association of Xenopus nuclear lamin B3 during meiosis and the fate of B3 in transfected mitotic cells. Journal of Cell Biology, 123(6): 1661–??, December 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/123/6/1661.

Funatsu:1993:TRE

[FKT93] T. Funatsu, E. Kono, and S. Tsukita. Time-resolved electron microscopic analysis of the behavior of myosin heads on actin filaments after photolysis of caged ATP. Journal of Cell Biology, 121(5):1053–??, June 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/121/5/1053.

Foltz:1990:PCE

[FL90] K. R. Foltz and W. J. Lennarz. Purification and characterization of an extracellular fragment of the sea urchin egg receptor for sperm. Journal of Cell Biology, 111(6):2951– ??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/111/6/2951.

Foltz:1992:ISU

[FL92] K. R. Foltz and W. J. Lennarz. Identification of the sea urchin egg receptor for sperm using an antiserum raised against a fragment of its extracellular domain. Journal of Cell Biology, 116(3):647–??, February 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/116/3/647.

Felder:1992:KBE

[FLUS92] S. Felder, J. LaVin, A. Ullrich, and J. Schlessinger. Kinetics of binding, endocytosis, and recycling of EGF receptor mutants. Journal of Cell Biology, 117(1):203–??, April 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/117/1/
203.

Fagotto:1994:YPX

[FM94a] F. Fagotto and F. R. Maxfield. Yolk platelets in Xenopus oocytes maintain an acidic internal pH which may be essential for sodium accumulation. Journal of Cell Biology, 125 (5):1047–??, June 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/125/5/1047.

Flomerfelt:1994:RMC

[FM94b] F. A. Flomerfelt and R. L. Miesfeld. Recessive mutations in a common pathway block thymocyte apoptosis induced by multiple signals. Journal of Cell Biology, 127(6):1729– ??, December 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/127/6/1729.

Finlay:1991:CNP

[FMB^{+}91] D. R. Finlay, E. Meier, P. Bradley, J. Horecka, and D. J. Forbes. A complex of nuclear pore proteins required for pore function. Journal of Cell Biology, 114(1):169–??, July 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/114/1/
169.

Fietz:1993:AST

[FMCR93] M. J. Fietz, C. J. McLaughlan, M. T. Campbell, and G. E. Rogers. Analysis of the sheep trichohyalin gene: potential structural and calcium-binding roles of trichohyalin in the hair follicle. Journal of Cell Biology, 121(4):855–??, May 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/121/4/855.

Fabian:1993:RRM

[FMD93] J. R. Fabian, D. K. Morrison, and I. O. Daar. Requirement for Raf and MAP kinase function during the meiotic maturation of Xenopus oocytes. Journal of Cell Biology, 122(3): 645–??, August 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/122/3/645.

Fluck:1991:SCW

[FMJ91] R. A. Fluck, A. L. Miller, and L. F. Jaffe. Slow calcium waves accompany cytokinesis in medaka fish eggs. Journal of Cell Biology, 115(5):1259–??, December 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/115/5/1259.

Friedlander:1994:NCS

[FMK^{+}94] D. R. Friedlander, P. Milev, L. Karthikeyan, R. K. Margolis, R. U. Margolis, and M. Grumet. The neuronal chondroitin sulfate proteoglycan neurocan binds to the neural cell adhesion molecules Ng-CAM/ L1/ NILE and N-CAM, and inhibits neuronal adhesion and neurite outgrowth. Journal of Cell Biology, 125(3):669–??, May 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/125/3/669.

Freeman:1992:PCS

[FMLD92] R. S. Freeman, A. N. Meyer, J. Li, and D. J. Donoghue. Phosphorylation of conserved serine residues does not regulate the ability of mosxe protein kinase to induce oocyte maturation or function as cytostatic factor. Journal of Cell Biology, 116(3):725–??, February 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/116/3/725.

Felez:1990:RPR

[FMPP90] J. Felez, L. A. Miles, J. Plescia, and E. F. Plow. Regulation of plasminogen receptor expression on human monocytes and monocytoid cell lines. Journal of Cell Biology, 111(4): 1673–??, October 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/111/4/1673.

Flaumenhaft:1990:HHS

[FMR90] R. Flaumenhaft, D. Moscatelli, and D. B. Rifkin. Heparin and heparan sulfate increase the radius of diffusion and action of basic fibroblast growth factor. Journal of Cell Biology, 111(4):1651–??, October 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/111/4/1651.

Frank:1992:DCE

[FMR^{+}92] J. S. Frank, G. Mottino, D. Reid, R. S. Molday, and K. D. Philipson. Distribution of the Na(+)-Ca^{2+} exchange protein in mammalian cardiac myocytes: an immunofluorescence and immunocolloidal gold-labeling study. Journal of Cell Biology, 117(2):337–??, April 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/117/2/337.

Fan:1993:OFA

[FMR^{+}93] J. Fan, S. G. Mansfield, T. Redmond, P. R. Gordon-Weeks, and J. A. Raper. The organization of F-actin and microtubules in growth cones exposed to a brain-derived collapsing factor. Journal of Cell Biology, 121(4):867–??, May 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/121/4/867.

Flescher:1993:CRC

[FMS93] E. G. Flescher, K. Madden, and M. Snyder. Components required for cytokinesis are important for bud site selection in yeast. Journal of Cell Biology, 122(2):373–??, July 1993.

CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/122/2/
373.

Fujimoto:1992:LIT

[FNM^{+}92] T. Fujimoto, S. Nakade, A. Miyawaki, K. Mikoshiba, and K. Ogawa. Localization of inositol 1,4,5-trisphosphate receptor-like protein in plasmalemmal caveolae. Journal of Cell Biology, 119(6):1507–??, December 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/119/6/1507.

Funayama:1991:RNM

[FNS^{+}91] N. Funayama, A. Nagafuchi, N. Sato, S. Tsukita, and S. Tsukita. Radixin is a novel member of the band 4.1 family. Journal of Cell Biology, 115(4):1039–??, November 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/115/4/1039.

Ferreira:1992:SKE

[FNV^{+}92] A. Ferreira, J. Niclas, R. D. Vale, G. Banker, and K. S. Kosik. Suppression of kinesin expression in cultured hippocampal neurons using antisense oligonucleotides. Journal of Cell Biology, 117(3):595–??, May 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/117/3/595.

Fowler:1990:TCP

[Fow90] V. M. Fowler. Tropomodulin: a cytoskeletal protein that binds to the end of erythrocyte tropomyosin and inhibits tropomyosin binding to actin. Journal of Cell Biology, 111 (2):471–??, August 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/111/2/471.

Fournier:1991:CNH

[FP91] B. Fournier and P. A. Price. Characterization of a new human osteosarcoma cell line OHS-4. Journal of Cell Biology, 114(3):577–??, August 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/114/3/577.

Frank:1991:TDR

[FPGS91] J. Frank, P. Penczek, R. Grassucci, and S. Srivastava. Three-dimensional reconstruction of the 70S Escherichia coli ribosome in ice: the distribution of ribosomal RNA. Journal of Cell Biology, 115(3):597–??, November 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/115/3/
597.

Flucher:1991:DRA

[FPP91] B. E. Flucher, J. L. Phillips, and J. A. Powell. Dihydropyridine receptor alpha subunits in normal and dysgenic muscle in vitro: expression of alpha 1 is required for proper targeting and distribution of alpha 2. Journal of Cell Biology, 115(5):1345–??, December 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/115/5/1345.

Fuller-Pace:1991:CTK

[FPPD91] F. Fuller-Pace, G. Peters, and C. Dickson. Cell transformation by kFGF requires secretion but not glycosylation. Journal of Cell Biology, 115(2):547–??, October 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/115/2/547.

Fietz:1990:CDA

[FPR90] M. J. Fietz, R. B. Presland, and G. E. Rogers. The cDNAdeduced amino acid sequence for trichohyalin, a differentiation marker in the hair follicle, contains a 23 amino acid repeat. Journal of Cell Biology, 110(2):427–??, February 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/110/2/
427.

Franzusoff:1991:LCI

[FRC^{+}91] A. Franzusoff, K. Redding, J. Crosby, R. S. Fuller, and R. Schekman. Localization of components involved in protein transport and processing through the yeast Golgi apparatus. Journal of Cell Biology, 112(1):27–??, January 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/112/1/

Frisch:1994:EIE

[Fri94] S. M. Frisch. E1a induces the expression of epithelial characteristics. Journal of Cell Biology, 127(4):1085–??, November 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/127/4/1085.

Fredette:1993:RAD

[FRL93] B. Fredette, U. Rutishauser, and L. Landmesser. Regulation and activity-dependence of N-cadherin, NCAM isoforms, and polysialic acid on chick myotubes during development. Journal of Cell Biology, 123(6):1867–??, December 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/123/6/
1867.

Froehner:1991:SMN

[Fro91] S. C. Froehner. The submembrane machinery for nicotinic acetylcholine receptor clustering. Journal of Cell Biology, 114(1):1–??, July 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/114/1/1.

Flanders:1990:NAM

[FRSL90] D. J. Flanders, D. J. Rawlins, P. J. Shaw, and C. W. Lloyd. Nucleus-associated microtubules help determine the division plane of plant epidermal cells: avoidance of four-way junctions and the role of cell geometry. Journal of Cell Biology, 110(4):1111–??, April 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/110/4/1111.

Feldheim:1994:SCS

[FS94] D. Feldheim and R. Schekman. Sec72p contributes to the selective recognition of signal peptides by the secretory polypeptide translocation complex. Journal of Cell Biology, 126(4):935–??, August 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/126/4/935.

Faassen:1992:CSC

[FSK^{+}92] A. E. Faassen, J. A. Schrager, D. J. Klein, T. R. Oegema, J. R. Couchman, and J. B. McCarthy. A cell surface chondroitin sulfate proteoglycan, immunologically related to CD44, is involved in type I collagen-mediated melanoma cell motility and invasion. Journal of Cell Biology, 116(2): 521–??, January 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/116/2/521.

Fishman:1990:MCF

[FSL90] G. I. Fishman, D. C. Spray, and L. A. Leinwand. Molecular characterization and functional expression of the human cardiac gap junction channel. Journal of Cell Biology, 111(2):589–??, August 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/111/2/589.

Fowler:1993:TAF

[FSM^{+}93] V. M. Fowler, M. A. Sussmann, P. G. Miller, B. E. Flucher, and M. P. Daniels. Tropomodulin is associated with the free (pointed) ends of the thin filaments in rat skeletal muscle. Journal of Cell Biology, 120(2):411–??, January 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/120/2/
411.

Fujishige:1992:CFA

[FSSA92] A. Fujishige, K. R. Smith, J. L. Silen, and D. A. Agard. Correct folding of alpha-lytic protease is required for its extracellular secretion from Escherichia coli. Journal of Cell Biology, 118(1):33–??, July 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/118/1/33.

Fu:1991:AFG

[FSY^{+}91] Y. M. Fu, P. Spirito, Z. X. Yu, S. Biro, J. Sasse, J. Lei, V. J. Ferrans, S. E. Epstein, and W. Casscells. Acidic fibroblast growth factor in the developing rat embryo. Journal of Cell Biology, 114(6):1261–??, September 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/114/6/1261.

Filbin:1993:HAM

[FT93] M. T. Filbin and G. I. Tennekoon. Homophilic adhesion of the myelin P0 protein requires glycosylation of both molecules in the homophilic pair. Journal of Cell Biology, 122(2):451–??, July 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/122/2/451.

Fath:1994:MMD

[FTB94] K. R. Fath, G. M. Trimbur, and D. R. Burgess. Molecular motors are differentially distributed on Golgi membranes from polarized epithelial cells. Journal of Cell Biology, 126 (3):661–??, August 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/126/3/661.

Ferguson:1994:KDR

[FTPM^{+}94] M. L. Ferguson, A. F. Torri, D. Pérez-Morga, D. C. Ward, and P. T. Englund. Kinetoplast DNA replication: mechanistic differences between Trypanosoma brucei and Crithidia fasciculata. Journal of Cell Biology, 126(3):631–??, August 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/126/3/631.

Fox:1993:EAP

[FTT^{+}93] J. E. Fox, R. G. Taylor, M. Taffarel, J. K. Boyles, and D. E. Goll. Evidence that activation of platelet calpain is induced as a consequence of binding of adhesive ligand to the integrin, glycoprotein IIb–IIIa. Journal of Cell Biology, 120(6): 1501–??, March 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/120/6/1501.

Fuchs:1990:EDB

[Fuc90] E. Fuchs. Epidermal differentiation: the bare essentials. Journal of Cell Biology, 111(6):2807–??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/111/6/
2807.

Fuchs:1994:IFD

[Fuc94] E. Fuchs. Intermediate filaments and disease: mutations that cripple cell strength. Journal of Cell Biology, 125(3): 511–??, May 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/125/3/511.

Fujimoto:1993:CPP

[Fuj93] T. Fujimoto. Calcium pump of the plasma membrane is localized in caveolae. Journal of Cell Biology, 120(5):1147–??, March 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/120/5/1147.

Fushimi:1991:LVA

[FV91] K. Fushimi and A. S. Verkman. Low viscosity in the aqueous domain of cell cytoplasm measured by picosecond polarization microfluorimetry. Journal of Cell Biology, 112(4): 719–??, February 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/112/4/719.

Frei:1992:DED

[FvBuHWS92] T. Frei, F. von Bohlen und Halbach, W. Wille, and M. Schachner. Different extracellular domains of the neural cell adhesion molecule (N-CAM) are involved in different functions. Journal of Cell Biology, 118(1):177–??, July 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/118/1/
177.

Francis:1991:MCA

[FW91] R. Francis and R. H. Waterston. Muscle cell attachment in Caenorhabditis elegans. Journal of Cell Biology, 114(3): 465–??, August 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/114/3/465.

Fishkind:1993:OTD

[FW93] D. J. Fishkind and Y. L. Wang. Orientation and threedimensional organization of actin filaments in dividing cultured cells. Journal of Cell Biology, 123(4):837–??, November 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/123/4/837.

Fuss:1993:MCS

[FWBS93] B. Fuss, E. S. Wintergerst, U. Bartsch, and M. Schachner. Molecular characterization and in situ mRNA localization of the neural recognition molecule J1-160/180: a modular structure similar to tenascin. Journal of Cell Biology, 120 (5):1237–??, March 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/120/5/1237.

Foxall:1992:TMS

[FWD^{+}92] C. Foxall, S. R. Watson, D. Dowbenko, C. Fennie, L. A. Lasky, M. Kiso, A. Hasegawa, D. Asa, and B. K. Brandley. The three members of the selectin receptor family recognize a common carbohydrate epitope, the sialyl Lewis(x) oligosaccharide. Journal of Cell Biology, 117(4):895–??, May 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/117/4/895.

Furukawa:1990:CPD

[FWF90] R. Furukawa, J. E. Wampler, and M. Fechheimer. Cytoplasmic pH of Dictyostelium discoideum amebae during early development: identification of two cell subpopulations before the aggregation stage. Journal of Cell Biology, 110(6): 1947–??, June 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/110/6/1947.

Feany:1993:SVP

[FYDB93] M. B. Feany, A. G. Yee, M. L. Delvy, and K. M. Buckley. The synaptic vesicle proteins SV2, synaptotagmin and synaptophysin are sorted to separate cellular compartments in CHO fibroblasts. Journal of Cell Biology, 123(3):575– ??, November 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/123/3/575.

Feeney:1990:NEU

[FZ90] R. J. Feeney and G. W. Zieve. Nuclear exchange of the U1 and U2 snRNP-specific proteins. Journal of Cell Biology, 110(4):871–??, April 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/110/4/871.

Fechheimer:1993:FUA

[FZ93] M. Fechheimer and S. H. Zigmond. Focusing on unpolymerized actin. Journal of Cell Biology, 123(1):1–??, October 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/123/1/1.

Fire:1991:ELM

[FZRH91] E. Fire, D. E. Zwart, M. G. Roth, and Y. I. Henis. Evidence from lateral mobility studies for dynamic interactions of a mutant influenza hemagglutinin with coated pits. Journal of Cell Biology, 115(6):1585–??, December 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/115/6/1585.

Guadagno:1991:GCA

[GA91] T. M. Guadagno and R. K. Assoian. G1/ S control of anchorage-independent growth in the fibroblast cell cycle. Journal of Cell Biology, 115(5):1419–??, December 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/115/5/
1419.

Goldberg:1992:HRS

[GA92] M. W. Goldberg and T. D. Allen. High resolution scanning electron microscopy of the nuclear envelope: demonstration of a new, regular, fibrous lattice attached to the baskets of the nucleoplasmic face of the nuclear pores. Journal of Cell Biology, 119(6):1429–??, December 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/119/6/1429.

Graham:1994:CRC

[GAHB94] I. L. Graham, D. C. Anderson, V. M. Holers, and E. J. Brown. Complement receptor 3 (CR3, Mac-1, integrin alpha M beta 2, CD11b/CD18) is required for tyrosine phosphorylation of paxillin in adherent and nonadherent neutrophils. Journal of Cell Biology, 127(4):1139–??, November 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/127/4/
1139.

Garrigues:1994:ATA

[GAHH94] J. Garrigues, J. Anderson, K. E. Hellström, and I. Hellström. Anti-tumor antibody BR96 blocks cell migration and binds to a lysosomal membrane glycoprotein on cell surface microspikes and ruffled membranes. Journal of Cell Biology, 125(1):129–??, April 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/125/1/129.

Gloor:1990:AMG

[GAS^{+}90a] S. Gloor, H. Antonicek, K. J. Sweadner, S. Pagliusi, R. Frank, M. Moos, and M. Schachner. The adhesion molecule on glia (AMOG) is a homologue of the beta subunit of the Na,K–ATPase. Journal of Cell Biology, 110(1): 165–??, January 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/110/1/165.

Gravotta:1990:TIT

[GAS90b] D. Gravotta, M. Adesnik, and D. D. Sabatini. Transport of influenza HA from the trans-Golgi network to the apical surface of MDCK cells permeabilized in their basolateral plasma membranes: energy dependence and involvement of GTP-binding proteins. Journal of Cell Biology, 111(6):2893–??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/111/6/2893.

Gauthier:1990:DDM

[Gau90] G. F. Gauthier. Differential distribution of myosin isoforms among the myofibrils of individual developing muscle fibers. Journal of Cell Biology, 110(3):693–??, March 1990. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/110/3/
693.

Goodman:1991:MCS

[GAvdM91] S. L. Goodman, M. Aumailley, and H. von der Mark. Multiple cell surface receptors for the short arms of laminin: alpha 1 beta 1 integrin and RGD-dependent proteins mediate cell attachment only to domains III in murine tumor laminin. Journal of Cell Biology, 113(4):931–??, May 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/113/4/931.

Goslin:1990:RCD

[GB90] K. Goslin and G. Banker. Rapid changes in the distribution of GAP-43 correlate with the expression of neuronal polarity during normal development and under experimental conditions. Journal of Cell Biology, 110(4):1319–??, April 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/110/4/
1319.

Ghitescu:1992:TTS

[GB92] L. Ghitescu and M. Bendayan. Transendothelial transport of serum albumin: a quantitative immunocytochemical study. Journal of Cell Biology, 117(4):745–??, May 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/117/4/
745.

Galimi:1994:HGF

[GBB^{+}94] F. Galimi, G. P. Bagnara, L. Bonsi, E. Cottone, A. Follenzi, A. Simeone, and P. M. Comoglio. Hepatocyte growth factor induces proliferation and differentiation of multipotent and erythroid hemopoietic progenitors. Journal of Cell Biology, 127(6):1743–??, December 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/127/6/1743.

Gentili:1993:CPE

[GBN^{+}93] C. Gentili, P. Bianco, M. Neri, M. Malpeli, G. Campanile, P. Castagnola, R. Cancedda, and F. D. Cancedda. Cell proliferation, extracellular matrix mineralization, and ovotransferrin transient expression during in vitro differentiation of chick hypertrophic chondrocytes into osteoblast-like cells. Journal of Cell Biology, 122(3):703–??, August 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/122/3/
703.

Gonzalez:1990:DBF

[GBOB90] A. M. Gonzalez, M. Buscaglia, M. Ong, and A. Baird. Distribution of basic fibroblast growth factor in the 18-day rat fetus: localization in the basement membranes of diverse tissues. Journal of Cell Biology, 110(3):753–??, March 1990.

CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/110/3/
753.

Gattei:1992:PEI

[GBP^{+}92] V. Gattei, P. A. Bernabei, A. Pinto, R. Bezzini, A. Ringressi, L. Formigli, A. Tanini, V. Attadia, and M. L. Brandi. Phorbol ester induced osteoclast-like differentiation of a novel human leukemic cell line (FLG 29.1). Journal of Cell Biology, 116(2):437–??, January 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/116/2/437.

Golden:1990:RPM

[GBS90] A. Golden, J. S. Brugge, and S. J. Shattil. Role of platelet membrane glycoprotein IIb–IIIa in agonist-induced tyrosine phosphorylation of platelet proteins. Journal of Cell Biology, 111(6):3117–??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/111/6/3117.

Grandin:1990:CIP

[GC90] N. Grandin and M. Charbonneau. Cycling of intracellular pH during cell division of Xenopus embryos is a cytoplasmic activity depending on protein synthesis and phosphorylation. Journal of Cell Biology, 111(2):523–??, August 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/111/2/
523.

Grandin:1991:IFC

[GC91] N. Grandin and M. Charbonneau. Intracellular free calcium oscillates during cell division of Xenopus embryos. Journal of Cell Biology, 112(4):711–??, February 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/112/4/711.

Gottardi:1993:ITA

[GC93] C. J. Gottardi and M. J. Caplan. An ion-transporting AT-Pase encodes multiple apical localization signals. Journal of Cell Biology, 121(2):283–??, April 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/121/2/283.

Galli:1994:TTM

[GCM^{+}94] T. Galli, T. Chilcote, O. Mundigl, T. Binz, H. Niemann, and P. De Camilli. Tetanus toxin-mediated cleavage of cellubrevin impairs exocytosis of transferrin receptor-containing vesicles in CHO cells. Journal of Cell Biology, 125(5):1015– ??, June 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/125/5/1015.

Goldschmidt-Clermont:1991:MIH

[GCMDP91] P. J. Goldschmidt-Clermont, L. M. Machesky, S. K. Doberstein, and T. D. Pollard. Mechanism of the interaction of human platelet profilin with actin. Journal of Cell Biology, 113(5):1081–??, June 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/113/5/1081.

Gerdes:1994:DCH

[GCML94] M. G. Gerdes, K. C. Carter, P. T. Moen, and J. B. Lawrence. Dynamic changes in the higher-level chromatin organization of specific sequences revealed by in situ hybridization to nuclear halos. Journal of Cell Biology, 126 (2):289–??, July 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/126/2/289.

Ganesan:1992:ILA

[GCZ^{+}92] S. Ganesan, R. Calle, K. Zawalich, K. Greenawalt, W. Zawalich, G. I. Shulman, and H. Rasmussen. Immunocytochemical localization of alpha-protein kinase C in rat pancreatic beta-cells during glucose-induced insulin secretion. Journal of Cell Biology, 119(2):313–??, October 1992. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/119/2/
313.

Gentili:1994:OOR

[GDB^{+}94] C. Gentili, R. Doliana, P. Bet, G. Campanile, A. Colombatti, F. D. Cancedda, and R. Cancedda. Ovotransferrin and ovotransferrin receptor expression during chondrogenesis and endochondral bone formation in developing chick embryo. Journal of Cell Biology, 124(4):579–??, February 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/124/4/579.

Grzesiak:1992:RAB

[GDKP92] J. J. Grzesiak, G. E. Davis, D. Kirchhofer, and M. D. Pierschbacher. Regulation of alpha 2 beta 1-mediated fibroblast migration on type I collagen by shifts in the concentrations of extracellular Mg2+ and Ca^{2+}. Journal of Cell Biology, 117(5):1109–??, June 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/117/5/1109.

Gautier:1994:EDM

[GdLMS94] M. C. Gautier, N. Garreau de Loubresse, L. Madeddu, and L. Sperling. Evidence for defects in membrane traffic in Paramecium secretory mutants unable to produce functional storage granules. Journal of Cell Biology, 124(6): 893–??, March 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/124/6/893.

Graham:1991:COG

[GE91] T. R. Graham and S. D. Emr. Compartmental organization of Golgi-specific protein modification and vacuolar protein sorting events defined in a yeast sec18 (NSF) mutant. Journal of Cell Biology, 114(2):207–??, July 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/114/2/207.

Greenberg:1991:CIF

[GeKdV^{+}91] S. Greenberg, J. el Khoury, F. di Virgilio, E. M. Kaplan, and S. C. Silverstein. Ca^{(2+)}-independent F-actin assembly and disassembly during Fc receptor-mediated phagocytosis in mouse macrophages. Journal of Cell Biology, 113 (4):757–??, May 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/113/4/757.

Griffiths:1994:LLA

[GEKL^{+}94] G. Griffiths, M. Ericsson, J. Krijnse-Locker, T. Nilsson, B. Goud, H. D. Söling, B. L. Tang, S. H. Wong, and W. Hong. Localization of the Lys, Asp, Glu, Leu tetrapeptide receptor to the Golgi complex and the intermediate compartment in mammalian cells. Journal of Cell Biology, 127(6):1557–??, December 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/127/6/1557.

Goode:1994:INM

[GF94] B. L. Goode and S. C. Feinstein. Identification of a novel microtubule binding and assembly domain in the developmentally regulated inter-repeat region of tau. Journal of Cell Biology, 124(5):769–??, March 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/124/5/769.

Gonzalez-Fernandez:1990:EEG

[GFH90] F. Gonzalez-Fernandez and J. I. Healy. Early expression of the gene for interphotoreceptor retinol-binding protein during photoreceptor differentiation suggests a critical role for the interphotoreceptor matrix in retinal development. Journal of Cell Biology, 111(6):2775–??, December 1990. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/111/6/
2775.

Gow:1992:MBP

[GFL92] A. Gow, V. L. Friedrich, and R. A. Lazzarini. Myelin basic protein gene contains separate enhancers for oligodendrocyte and Schwann cell expression. Journal of Cell Biology, 119(3):605–??, November 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/119/3/605.

Grumet:1993:FCC

[GFM93] M. Grumet, A. Flaccus, and R. U. Margolis. Functional characterization of chondroitin sulfate proteoglycans of brain: interactions with neurons and neural cell adhesion molecules. Journal of Cell Biology, 120(3):815–??, February 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/120/3/815.

Gu:1991:AMM

[GFV^{+}91] Y. Gu, J. R. Forsayeth, S. Verrall, X. M. Yu, and Z. W. Hall. Assembly of the mammalian muscle acetylcholine receptor in transfected COS cells. Journal of Cell Biology, 114(4): 799–??, August 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/114/4/799.

Green:1992:MTN

[GFW92] N. Green, H. Fang, and P. Walter. Mutants in three novel complementation groups inhibit membrane protein insertion into and soluble protein translocation across the endoplasmic reticulum membrane of Saccharomyces cerevisiae. Journal of Cell Biology, 116(3):597–??, February 1992. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/116/3/
597.

Ginty:1992:ACD

[GFWM92] D. D. Ginty, G. R. Fanger, J. A. Wagner, and R. A. Maue. The activity of cAMP-dependent protein kinase is required at a posttranslational level for induction of voltage-dependent sodium channels by peptide growth factors in PC12 cells. Journal of Cell Biology, 116(6):1465–??, March 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/116/6/1465.

Glass:1990:LCB

[GG90] J. R. Glass and L. Gerace. Lamins A and C bind and assemble at the surface of mitotic chromosomes. Journal of Cell Biology, 111(3):1047–??, September 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/111/3/1047.

Greber:1992:NPI

[GG92] U. F. Greber and L. Gerace. Nuclear protein import is inhibited by an antibody to a lumenal epitope of a nuclear pore complex glycoprotein. Journal of Cell Biology, 116(1): 15–??, January 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/116/1/15.

Goday:1992:KCD

[GGGE^{+}92] C. Goday, J. M. González-García, M. R. Esteban, G. Giovinazzo, and S. Pimpinelli. Kinetochores and chromatin diminution in early embryos of Parascaris univalens. Journal of Cell Biology, 118(1):23–??, July 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/118/1/23.

Gillespie:1991:HPI

[GH91a] P. G. Gillespie and A. J. Hudspeth. High-purity isolation of bullfrog hair bundles and subcellular and topological localization of constituent proteins. Journal of Cell Biology, 112(4):625–??, February 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/112/4/625.

Guidry:1991:EPE

[GH91b] C. Guidry and M. Hook. Endothelins produced by endothelial cells promote collagen gel contraction by fibroblasts. Journal of Cell Biology, 115(3):873–??, November 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/115/3/
873.

Griffo:1993:HHW

[GHBA^{+}93] G. Griffo, C. Hamon-Benais, P. O. Angrand, M. Fox, L. West, O. Lecoq, S. Povey, D. Cassio, and M. Weiss. HNF4 and HNF1 as well as a panel of hepatic functions are extinguished and reexpressed in parallel in chromosomally reduced rat hepatoma-human fibroblast hybrids. Journal of Cell Biology, 121(4):887–??, May 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/121/4/887.

Ghinea:1992:PIH

[GHGP^{+}92] N. Ghinea, M. T. Vu Hai, M. T. Groyer-Picard, A. Houllier, D. Schoëvaërt, and E. Milgrom. Pathways of internalization of the hCG/LH receptor: immunoelectron microscopic studies in Leydig cells and transfected L-cells. Journal of Cell Biology, 118(6):1347–??, September 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/118/6/1347.

Guidry:1990:ECS

[GHH90] C. Guidry, S. Hohn, and M. Hook. Endothelial cells secrete a factor that promotes fibroblast contraction of hydrated collagen gels. Journal of Cell Biology, 110(2):519–??, February 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/110/2/519.

Guacci:1994:CCS

[GHK94] V. Guacci, E. Hogan, and D. Koshland. Chromosome condensation and sister chromatid pairing in budding yeast. Journal of Cell Biology, 125(3):517–??, May 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/125/3/517.

Garoff:1990:SSP

[GHR^{+}90] H. Garoff, D. Huylebroeck, A. Robinson, U. Tillman, and P. Liljeström. The signal sequence of the p62 protein of Semliki Forest virus is involved in initiation but not in completing chain translocation. Journal of Cell Biology, 111(3): 867–??, September 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/111/3/867.

Gil:1991:MNM

[GHR91] J. Gil, T. Higgins, and E. Rozengurt. Mastoparan, a novel mitogen for Swiss 3T3 cells, stimulates pertussis toxinsensitive arachidonic acid release without inositol phosphate accumulation. Journal of Cell Biology, 113(4):943– ??, May 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/113/4/943.

Gachet:1993:AIB

[GHS^{+}93] C. Gachet, D. Hanau, D. Spehner, C. Brisson, J. C. Garaud, D. A. Schmitt, P. Ohlmann, and J. P. Cazenave. Alpha IIb beta 3 integrin dissociation induced by EDTA results in morphological changes of the platelet surface-connected canalicular system with differential location of the two separate subunits. Journal of Cell Biology, 120(4):1021–??, February 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/120/4/1021.

Gormley:1992:TMP

[GHT92] J. A. Gormley, R. J. Howard, and T. F. Taraschi. Trafficking of malarial proteins to the host cell cytoplasm and erythrocyte surface membrane involves multiple pathways. Journal of Cell Biology, 119(6):1481–??, December 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/119/6/1481.

Gard:1990:CDC

[GHZD90] D. L. Gard, S. Hafezi, T. Zhang, and S. J. Doxsey. Centrosome duplication continues in cycloheximide-treated Xenopus blastulae in the absence of a detectable cell cycle. Journal of Cell Biology, 110(6):2033–??, June 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/110/6/2033.

Griffiths:1993:GBT

[GI93] G. M. Griffiths and S. Isaaz. Granzymes A and B are targeted to the lytic granules of lymphocytes by the mannose-6-phosphate receptor. Journal of Cell Biology, 120(4):885– ??, February 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/120/4/885.

Gottlieb:1993:AMP

[GIAS93] T. A. Gottlieb, I. E. Ivanov, M. Adesnik, and D. D. Sabatini. Actin microfilaments play a critical role in endocytosis at the apical but not the basolateral surface of polarized epithelial cells. Journal of Cell Biology, 120(3):695–??, February 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/120/3/695.

Garippa:1994:ATG

[GJJM94] R. J. Garippa, T. W. Judge, D. E. James, and T. E. McGraw. The amino terminus of GLUT4 functions as an internalization motif but not an intracellular retention signal when substituted for the transferrin receptor cytoplasmic domain. Journal of Cell Biology, 124(5):705–??, March 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/124/5/705.

Giulian:1991:MMP

[GJK^{+}91] D. Giulian, B. Johnson, J. F. Krebs, J. K. George, and M. Tapscott. Microglial mitogens are produced in the developing and injured mammalian brain. Journal of Cell Biology, 112(2):323–??, January 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/112/2/323.

Gabathuler:1990:ERR

[GK90] R. Gabathuler and S. Kvist. The endoplasmic reticulum retention signal of the E3/19K protein of adenovirus type 2 consists of three separate amino acid segments at the carboxy terminus. Journal of Cell Biology, 111(5):1803– ??, November 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/111/5/1803.

Green:1992:LDL

[GK92a] S. A. Green and R. B. Kelly. Low density lipoprotein receptor and cation-independent mannose 6-phosphate receptor are transported from the cell surface to the Golgi apparatus at equal rates in PC12 cells. Journal of Cell Biology, 117 (1):47–??, April 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/117/1/47.

Grimes:1992:ICR

[GK92b] M. Grimes and R. B. Kelly. Intermediates in the constitutive and regulated secretory pathways released in vitro from semi-intact cells. Journal of Cell Biology, 117(3):539– ??, May 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/117/3/539.

Glickman:1993:MPI

[GK93a] J. N. Glickman and S. Kornfeld. Mannose 6-phosphate-independent targeting of lysosomal enzymes in I-cell disease B lymphoblasts. Journal of Cell Biology, 123(1):99–??, October 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/123/1/99.

Goh:1993:NGI

[GK93b] P. Y. Goh and J. V. Kilmartin. NDC10: a gene involved in chromosome segregation in Saccharomyces cerevisiae. Journal of Cell Biology, 121(3):503–??, May 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/121/3/503.

Gimlich:1990:DRL

[GKG90] R. L. Gimlich, N. M. Kumar, and N. B. Gilula. Differential regulation of the levels of three gap junction mRNAs in Xenopus embryos. Journal of Cell Biology, 110(3):597–??, March 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/110/3/597.

Gould:1990:ADA

[GKKS90] S. J. Gould, S. Krisans, G. A. Keller, and S. Subramani. Antibodies directed against the peroxisomal targeting signal of firefly luciferase recognize multiple mammalian peroxisomal proteins. Journal of Cell Biology, 110(1):27–??, January 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/110/1/27.

Gitelman:1994:RVB

[GKY^{+}94] S. E. Gitelman, M. S. Kobrin, J. Q. Ye, A. R. Lopez, A. Lee, and R. Derynck. Recombinant Vgr-1/BMP-6-expressing tumors induce fibrosis and endochondral bone formation in vivo. Journal of Cell Biology, 126(6):1595–??, September 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/126/6/1595.

Gerstenfeld:1991:GEE

[GL91] L. C. Gerstenfeld and W. J. Landis. Gene expression and extracellular matrix ultrastructure of a mineralizing chondrocyte cell culture system. Journal of Cell Biology, 112(3): 501–??, February 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/112/3/501.

Graham:1993:ICS

[GLAB93] I. L. Graham, J. B. Lefkowith, D. C. Anderson, and E. J. Brown. Immune complex-stimulated neutrophil LTB4 production is dependent on beta 2 integrins. Journal of Cell Biology, 120(6):1509–??, March 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/120/6/1509.

Gilbert:1991:MOI

[GLQRB91] T. Gilbert, A. Le Bivic, A. Quaroni, and E. Rodriguez-Boulan. Microtubular organization and its involvement in the biogenetic pathways of plasma membrane proteins in Caco-2 intestinal epithelial cells. Journal of Cell Biology, 113(2):275–??, April 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/113/2/275.

Grumet:1991:SNN

[GMB^{+}91] M. Grumet, V. Mauro, M. P. Burgoon, G. M. Edelman, and B. A. Cunningham. Structure of a new nervous system glycoprotein, Nr–CAM, and its relationship to subgroups of neural cell adhesion molecules. Journal of Cell Biology, 113 (6):1399–??, June 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/113/6/1399.

Ghinea:1994:HPH

[GMGPM94] N. Ghinea, T. V. Mai, M. T. Groyer-Picard, and E. Milgrom. How protein hormones reach their target cells. Receptor-mediated transcytosis of hCG through endothelial cells. Journal of Cell Biology, 125(1):87–??, April 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/125/1/
87.

Gabellini:1992:NGF

[GMLD92] N. Gabellini, M. C. Minozzi, A. Leon, and R. Dal Toso. Nerve growth factor transcriptional control of c-fos promoter transfected in cultured spinal sensory neurons. Journal of Cell Biology, 118(1):131–??, July 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/118/1/131.

Gamble:1993:RVC

[GMM^{+}93] J. R. Gamble, L. J. Matthias, G. Meyer, P. Kaur, G. Russ, R. Faull, M. C. Berndt, and M. A. Vadas. Regulation of in vitro capillary tube formation by anti-integrin antibodies. Journal of Cell Biology, 121(4):931–??, May 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/121/4/931.

Goldman:1992:PIM

[GMML^{+}92] A. E. Goldman, R. D. Moir, M. Montag-Lowy, M. Stewart, and R. D. Goldman. Pathway of incorporation of microinjected lamin A into the nuclear envelope. Journal of Cell Biology, 119(4):725–??, November 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/119/4/725.

Gittes:1993:FRM

[GMNH93] F. Gittes, B. Mickey, J. Nettleton, and J. Howard. Flexural rigidity of microtubules and actin filaments measured from thermal fluctuations in shape. Journal of Cell Biology, 120(4):923–??, February 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/120/4/923.

Gounari:1993:BFP

[GMQ^{+}93] F. Gounari, A. Merdes, R. Quinlan, J. Hess, P. G. FitzGerald, C. A. Ouzounis, and S. D. Georgatos. Bovine filensin possesses primary and secondary structure similarity to intermediate filament proteins. Journal of Cell Biology, 121 (4):847–??, May 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/121/4/847.

Gao:1994:NCM

[GMW^{+}94] Y. Gao, R. Melki, P. D. Walden, S. A. Lewis, C. Ampe, H. Rommelaere, J. Vandekerckhove, and N. J. Cowan. A novel cochaperonin that modulates the ATPase activity of cytoplasmic chaperonin. Journal of Cell Biology, 125(5): 989–??, June 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/125/5/989.

Gallant:1992:CBU

[GN92] P. Gallant and E. A. Nigg. Cyclin B2 undergoes cell cycledependent nuclear translocation and, when expressed as a non-destructible mutant, causes mitotic arrest in HeLa cells. Journal of Cell Biology, 117(1):213–??, April 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/117/1/213.

Goldstein:1993:FRM

[Gol93] L. S. Goldstein. Functional redundancy in mitotic force generation. Journal of Cell Biology, 120(1):1–??, January 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/120/1/
1.

Gardner:1994:CDR

[GOP^{+}94] L. C. Gardner, E. O’Toole, C. A. Perrone, T. Giddings, and M. E. Porter. Components of a “dynein regulatory complex” are located at the junction between the radial spokes and the dynein arms in Chlamydomonas flagella. Journal of Cell Biology, 127(5):1311–??, December 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/127/5/1311.

Gundersen:1991:APK

[GORB91] D. Gundersen, J. Orlowski, and E. Rodriguez-Boulan. Apical polarity of Na,K–ATPase in retinal pigment epithelium is linked to a reversal of the ankyrin-fodrin submembrane cytoskeleton. Journal of Cell Biology, 112(5):863–??, March 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/112/5/863.

Goda:1991:INE

[GP91] Y. Goda and S. R. Pfeffer. Identification of a novel, Nethylmaleimide-sensitive cytosolic factor required for vesicular transport from endosomes to the trans-Golgi network in vitro. Journal of Cell Biology, 112(5):823–??, March 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/112/5/
823.

Gilbert:1994:MCI

[GPF94] S. H. Gilbert, K. Perry, and F. S. Fay. Mediation of chemoattractant-induced changes in [Ca^{2+}]i and cell shape, polarity, and locomotion by InsP3, DAG, and protein kinase C in newt eosinophils. Journal of Cell Biology, 127(2): 489–??, October 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/127/2/489.

Gorlich:1990:SSR

[GPH^{+}90] D. Görlich, S. Prehn, E. Hartmann, J. Herz, A. Otto, R. Kraft, M. Wiedmann, S. Knespel, B. Dobberstein, and T. A. Rapoport. The signal sequence receptor has a second subunit and is part of a translocation complex in the endoplasmic reticulum as probed by bifunctional reagents. Journal of Cell Biology, 111(6):2283–??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/111/6/
2283.

Gundersen:1993:APC

[GPRB93] D. Gundersen, S. K. Powell, and E. Rodriguez-Boulan. Apical polarization of N-CAM in retinal pigment epithelium is dependent on contact with the neural retina. Journal of Cell Biology, 121(2):335–??, April 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/121/2/335.

Gliksman:1992:OAI

[GPS92] N. R. Gliksman, S. F. Parsons, and E. D. Salmon. Okadaic acid induces interphase to mitotic-like microtubule dynamic instability by inactivating rescue. Journal of Cell Biology, 119(5):1271–??, December 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/119/5/1271.

Gorbsky:1993:DEP

[GR93] G. J. Gorbsky and W. A. Ricketts. Differential expression of a phosphoepitope at the kinetochores of moving chromosomes. Journal of Cell Biology, 122(6):1311–??, September 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/122/6/1311.

Groesch:1990:IFV

[GRB^{+}90] M. E. Groesch, H. Ruohola, R. Bacon, G. Rossi, and S. Ferro-Novick. Isolation of a functional vesicular intermediate that mediates ER to Golgi transport in yeast. Journal of Cell Biology, 111(1):45–??, July 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/111/1/45.

Grellier:1991:LTB

[GRC^{+}91] P. Grellier, D. Rigomier, V. Clavey, J. C. Fruchart, and J. Schrevel. Lipid traffic between high density lipoproteins and Plasmodium falciparum-infected red blood cells. Journal of Cell Biology, 112(2):267–??, January 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/112/2/267.

Gregorio:1994:DPA

[GRFB94] C. C. Gregorio, E. A. Repasky, V. M. Fowler, and J. D. Black. Dynamic properties of ankyrin in T lymphocytes: colocalization with spectrin and protein kinase C beta. Journal of Cell Biology, 125(2):345–??, April 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/125/2/345.

Grinnell:1994:FMW

[Gri94] F. Grinnell. Fibroblasts, myofibroblasts, and wound contraction. Journal of Cell Biology, 124(4):401–??, February 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/124/4/401.

Gehrung:1990:SGS

[GS90] S. Gehrung and M. Snyder. The SPA2 gene of Saccharomyces cerevisiae is important for pheromone-induced morphogenesis and efficient mating. Journal of Cell Biology, 111(4):1451–??, October 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/111/4/1451.

Gupta:1994:IEC

[GS94] S. K. Gupta and J. P. Singh. Inhibition of endothelial cell proliferation by platelet factor-4 involves a unique action on S phase progression. Journal of Cell Biology, 127(4):1121– ??, November 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/127/4/1121.

Gavrieli:1992:IPC

[GSBS92] Y. Gavrieli, Y. Sherman, and S. A. Ben-Sasson. Identification of programmed cell death in situ via specific labeling of nuclear DNA fragmentation. Journal of Cell Biology, 119(3):493–??, November 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/119/3/493.

Girard:1992:CNP

[GSC^{+}92] F. Girard, U. Strausfeld, J. C. Cavadore, P. Russell, A. Fernandez, and N. J. Lamb. cdc25 is a nuclear protein expressed constitutively throughout the cell cycle in nontransformed mammalian cells. Journal of Cell Biology, 118(4): 785–??, August 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/118/4/785.

Gehlsen:1992:SPD

[GSFS92] K. R. Gehlsen, P. Sriramarao, L. T. Furcht, and A. P. Skubitz. A synthetic peptide derived from the carboxy terminus of the laminin A chain represents a binding site for the alpha 3 beta 1 integrin. Journal of Cell Biology, 117 (2):449–??, April 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/117/2/449.

Gundersen:1993:NRM

[GSM93] K. Gundersen, J. R. Sanes, and J. P. Merlie. Neural regulation of muscle acetylcholine receptor epsilon- and alphasubunit gene promoters in transgenic mice. Journal of Cell Biology, 123(6):1535–??, December 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/123/6/1535.

Green:1994:CDP

[GSMK94] S. A. Green, H. Setiadi, R. P. McEver, and R. B. Kelly. The cytoplasmic domain of P-selectin contains a sorting determinant that mediates rapid degradation in lysosomes. Journal of Cell Biology, 124(4):435–??, February 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/124/4/435.

Graham:1994:CDL

[GSP^{+}94] T. R. Graham, M. Seeger, G. S. Payne, V. L. MacKay, and S. D. Emr. Clathrin-dependent localization of alpha 1,3 mannosyltransferase to the Golgi complex of Saccharomyces cerevisiae. Journal of Cell Biology, 127(3):667– ??, November 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/127/3/667.

Gonzalez:1993:SIT

[GSR^{+}93] F. A. Gonzalez, A. Seth, D. L. Raden, D. S. Bowman, F. S. Fay, and R. J. Davis. Serum-induced translocation of mitogen-activated protein kinase to the cell surface ruffling membrane and the nucleus. Journal of Cell Biology, 122(5):1089–??, September 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/122/5/1089.

Gill:1991:DCU

[GSS^{+}91] S. R. Gill, T. A. Schroer, I. Szilak, E. R. Steuer, M. P. Sheetz, and D. W. Cleveland. Dynactin, a conserved, ubiquitously expressed component of an activator of vesicle motility mediated by cytoplasmic dynein. Journal of Cell Biology, 115(6):1639–??, December 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/115/6/1639.

Giancotti:1992:PPE

[GSS^{+}92] F. G. Giancotti, M. A. Stepp, S. Suzuki, E. Engvall, and E. Ruoslahti. Proteolytic processing of endogenous and recombinant beta 4 integrin subunit. Journal of Cell Biology, 118(4):951–??, August 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/118/4/951.

Goodenough:1993:RCC

[GSS^{+}93] U. W. Goodenough, B. Shames, L. Small, T. Saito, R. C. Crain, M. A. Sanders, and J. L. Salisbury. The role of calcium in the Chlamydomonas reinhardtii mating reaction. Journal of Cell Biology, 121(2):365–??, April 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/121/2/365.

Gorza:1993:ITR

[GSV93] L. Gorza, S. Schiaffino, and P. Volpe. Inositol 1,4,5trisphosphate receptor in heart: evidence for its concentration in Purkinje myocytes of the conduction system. Journal of Cell Biology, 121(2):345–??, April 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/121/2/345.

Grigoriadis:1993:OTC

[GSWW93] A. E. Grigoriadis, K. Schellander, Z. Q. Wang, and E. F. Wagner. Osteoblasts are target cells for transformation in cfos transgenic mice. Journal of Cell Biology, 122(3):685–??, August 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/122/3/685.

Gough:1993:FAI

[GT93] A. H. Gough and D. L. Taylor. Fluorescence anisotropy imaging microscopy maps calmodulin binding during cellular contraction and locomotion. Journal of Cell Biology, 121 (5):1095–??, June 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/121/5/1095.

Giuliano:1994:FAA

[GT94] K. A. Giuliano and D. L. Taylor. Fluorescent actin analogs with a high affinity for profilin in vitro exhibit an enhanced gradient of assembly in living cells. Journal of Cell Biology, 124(6):971–??, March 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/124/6/971.

Grant:1990:MNM

[GTC^{+}90] J. W. Grant, M. B. Taubman, S. L. Church, R. L. Johnson, and B. Nadal-Ginard. Mammalian nonsarcomeric myosin regulatory light chains are encoded by two differentially regulated and linked genes. Journal of Cell Biology, 111 (3):1127–??, September 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/111/3/1127.

Guan:1990:REA

[GTH90] J. L. Guan, J. E. Trevithick, and R. O. Hynes. Retroviral expression of alternatively spliced forms of rat fibronectin. Journal of Cell Biology, 110(3):833–??, March 1990. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/110/3/
833.

Gaynor:1994:SMR

[GtHG^{+}94] E. C. Gaynor, S. te Heesen, T. R. Graham, M. Aebi, and S. D. Emr. Signal-mediated retrieval of a membrane protein from the Golgi to the ER in yeast. Journal of Cell Biology, 127(3):653–??, November 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/127/3/653.

Gentile:1992:ETT

[GTP^{+}92] V. Gentile, V. Thomazy, M. Piacentini, L. Fesus, and P. J. Davies. Expression of tissue transglutaminase in Balb–C 3T3 fibroblasts: effects on cellular morphology and adhesion. Journal of Cell Biology, 119(2):463–??, October 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/119/2/
463.

Gumbiner:1993:BTT

[Gum93] B. M. Gumbiner. Breaking through the tight junction barrier. Journal of Cell Biology, 123(6):1631–??, December 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/123/6/1631.

Gambill:1993:DRM

[GVK^{+}93] B. D. Gambill, W. Voos, P. J. Kang, B. Miao, T. Langer, E. A. Craig, and N. Pfanner. A dual role for mitochondrial heat shock protein 70 in membrane translocation of preproteins. Journal of Cell Biology, 123(1):109–??, October 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/123/1/
109.

Guo:1994:DGS

[GVK94] Q. Guo, E. Vasile, and M. Krieger. Disruptions in Golgi structure and membrane traffic in a conditional lethal mammalian cell mutant are corrected by epsilon–COP. Journal of Cell Biology, 125(6):1213–??, June 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/125/6/1213.

Goetinck:1994:CEM

[GW94a] S. Goetinck and R. H. Waterston. The Caenorhabditis elegans muscle-affecting gene unc-87 encodes a novel thin filament-associated protein. Journal of Cell Biology, 127(1): 79–??, October 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/127/1/79.

Goetinck:1994:CEU

[GW94b] S. Goetinck and R. H. Waterston. The Caenorhabditis elegans UNC-87 protein is essential for maintenance, but not assembly, of bodywall muscle. Journal of Cell Biology, 127 (1):71–??, October 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/127/1/71.

Gill:1990:APD

[GWMC90] S. R. Gill, P. C. Wong, M. J. Monteiro, and D. W. Cleveland. Assembly properties of dominant and recessive mutations in the small mouse neurofilament (NF–L) subunit. Journal of Cell Biology, 111(5):2005–??, November 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/111/5/
2005.

Goerdt:1991:INH

[GWMP91] S. Goerdt, L. J. Walsh, G. F. Murphy, and J. S. Pober. Identification of a novel high molecular weight protein preferentially expressed by sinusoidal endothelial cells in normal human tissues. Journal of Cell Biology, 113(6):1425– ??, June 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/113/6/1425.

Gilmore:1993:CPT

[GWO^{+}93] A. P. Gilmore, C. Wood, V. Ohanian, P. Jackson, B. Patel, D. J. Rees, R. O. Hynes, and D. R. Critchley. The cytoskeletal protein talin contains at least two distinct vinculin binding domains. Journal of Cell Biology, 122(2):337– ??, July 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/122/2/337.

Grabski:1994:LTC

[GXHS94] S. Grabski, X. G. Xie, J. F. Holland, and M. Schindler. Lipids trigger changes in the elasticity of the cytoskeleton in plant cells: a cell optical displacement assay for live cell measurements. Journal of Cell Biology, 126(3):713–??, August 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/126/3/713.

Gorlin:1990:HEA

[GYE^{+}90] J. B. Gorlin, R. Yamin, S. Egan, M. Stewart, T. P. Stossel, D. J. Kwiatkowski, and J. H. Hartwig. Human endothelial actin-binding protein (ABP-280, nonmuscle filamin): a molecular leaf spring. Journal of Cell Biology, 111(3):1089– ??, September 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/111/3/1089.

Grohovaz:1991:CIP

[GZC^{+}91] F. Grohovaz, D. Zacchetti, E. Clementi, P. Lorenzon, J. Meldolesi, and G. Fumagalli. [Ca^{2+}]i imaging in PC12 cells: multiple response patterns to receptor activation reveal new aspects of transmembrane signaling. Journal of Cell Biology, 113(6):1341–??, June 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/113/6/1341.

Holtta:1993:PEC

[HAA93] E. Hölttä, M. Auvinen, and L. C. Andersson. Polyamines are essential for cell transformation by pp60v-src: delineation of molecular events relevant for the transformed phenotype. Journal of Cell Biology, 122(4):903–??, August 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/122/4/903.

High:1993:SAN

[HAG^{+}93] S. High, S. S. Andersen, D. Görlich, E. Hartmann, S. Prehn, T. A. Rapoport, and B. Dobberstein. Sec61p is adjacent to nascent type I and type II signal-anchor proteins during their membrane insertion. Journal of Cell Biology, 121 (4):743–??, May 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/121/4/743.

Higashiyama:1993:HBE

[HAK93] S. Higashiyama, J. A. Abraham, and M. Klagsbrun. Heparin-binding EGF-like growth factor stimulation of smooth muscle cell migration: dependence on interactions with cell surface heparan sulfate. Journal of Cell Biology, 122(4):933–??, August 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/122/4/933.

Hoschuetzky:1994:BCM

[HAK94] H. Hoschuetzky, H. Aberle, and R. Kemler. Beta-catenin mediates the interaction of the cadherin-catenin complex with epidermal growth factor receptor. Journal of Cell Biology, 127(5):1375–??, December 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/127/5/1375.

Hammer:1994:RDM

[Ham94] J. A. Hammer. Regulation of Dictyostelium myosin II by phosphorylation: what is essential and what is important? Journal of Cell Biology, 127(6):1779–??, December 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/127/6/
1779.

Hogervorst:1993:BCT

[HAN^{+}93] F. Hogervorst, L. G. Admiraal, C. Niessen, I. Kuikman, H. Janssen, H. Daams, and A. Sonnenberg. Biochemical characterization and tissue distribution of the A and B variants of the integrin alpha 6 subunit. Journal of Cell Biology, 121(1):179–??, April 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/121/1/179.

Hartwig:1992:MAR

[Har92] J. H. Hartwig. Mechanisms of actin rearrangements mediating platelet activation. Journal of Cell Biology, 118 (6):1421–??, September 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/118/6/1421.

Hiraoka:1990:TSC

[HAS90] Y. Hiraoka, D. A. Agard, and J. W. Sedat. Temporal and spatial coordination of chromosome movement, spindle formation, and nuclear envelope breakdown during prometaphase in Drosophila melanogaster embryos. Journal of Cell Biology, 111(6):2815–??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/111/6/2815.

Hemphill:1992:NMB

[HAS92] A. Hemphill, M. Affolter, and T. Seebeck. A novel microtubule-binding motif identified in a high molecular weight microtubule-associated protein from Trypanosoma brucei. Journal of Cell Biology, 117(1):95–??, April 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/117/1/
95.

Hasegawa:1993:PUA

[Has93] T. Hasegawa. Production of a unique antibody specific for membrane ruffles and its use to characterize the behavior of two distinct types of ruffles. Journal of Cell Biology, 120(6): 1439–??, March 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/120/6/1439.

Horowitz:1994:TDA

[HASW94] R. A. Horowitz, D. A. Agard, J. W. Sedat, and C. L. Woodcock. The three-dimensional architecture of chromatin in situ: electron tomography reveals fibers composed of a continuously variable zig-zag nucleosomal ribbon. Journal of Cell Biology, 125(1):1–??, April 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/125/1/1.

Hulsken:1994:CAC

[HBB94] J. Hülsken, W. Birchmeier, and J. Behrens. E-cadherin and APC compete for the interaction with beta-catenin and the cytoskeleton. Journal of Cell Biology, 127(6):2061–??, December 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/127/6/2061.

Herskovits:1993:EMR

[HBOV93] J. S. Herskovits, C. C. Burgess, R. A. Obar, and R. B. Vallee. Effects of mutant rat dynamin on endocytosis. Journal of Cell Biology, 122(3):565–??, August 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/122/3/565.

Hide:1993:DIM

[HBP^{+}93] I. Hide, J. P. Bennett, A. Pizzey, G. Boonen, D. Bar-Sagi, B. D. Gomperts, and P. E. Tatham. Degranulation of individual mast cells in response to Ca^{2+} and guanine nucleotides: an all-or-none event. Journal of Cell Biology, 123(3):585–??, November 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/123/3/585.

Hayden:1990:KCA

[HBR90] J. H. Hayden, S. S. Bowser, and C. L. Rieder. Kinetochores capture astral microtubules during chromosome attachment to the mitotic spindle: direct visualization in live newt lung cells. Journal of Cell Biology, 111(3):1039– ??, September 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/111/3/1039.

Herzog:1992:IIS

[HBS92] V. Herzog, U. Berndorfer, and Y. Saber. Isolation of insoluble secretory product from bovine thyroid: extracellular storage of thyroglobulin in covalently cross-linked form. Journal of Cell Biology, 118(5):1071–??, September 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/118/5/
1071.

Hansen:1994:GAS

[HC94] S. H. Hansen and J. E. Casanova. Gs alpha stimulates transcytosis and apical secretion in MDCK cells through cAMP and protein kinase A. Journal of Cell Biology, 126(3): 677–??, August 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/126/3/677.

Hodge:1992:RCT

[HCKJ92] T. P. Hodge, R. Cross, and J. Kendrick-Jones. Role of the COOH-terminal nonhelical tailpiece in the assembly of a vertebrate nonmuscle myosin rod. Journal of Cell Biology, 118(5):1085–??, September 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/118/5/1085.

Hoch:1994:AIC

[HCS94] W. Hoch, J. T. Campanelli, and R. H. Scheller. Agrininduced clustering of acetylcholine receptors: a cytoskeletal link. Journal of Cell Biology, 126(1):1–??, July 1994. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/126/1/
1.

Haas:1994:GPL

[HCW94] A. Haas, B. Conradt, and W. Wickner. G-protein ligands inhibit in vitro reactions of vacuole inheritance. Journal of Cell Biology, 126(1):87–??, July 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/126/1/87.

Hishikawa:1991:CTD

[HCYK91] T. Hishikawa, J. Y. Cheung, R. V. Yelamarty, and D. W. Knutson. Calcium transients during Fc receptormediated and nonspecific phagocytosis by murine peritoneal macrophages. Journal of Cell Biology, 115(1):59–??, October 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/115/1/59.

Howard:1990:GAI

[HCYS90] T. Howard, C. Chaponnier, H. Yin, and T. Stossel. Gelsolinactin interaction and actin polymerization in human neutrophils. Journal of Cell Biology, 110(6):1983–??, June 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/
110/6/1983.

Hong:1990:MDN

[HD90] W. J. Hong and D. Doyle. Molecular dissection of the NH2terminal signal/anchor sequence of rat dipeptidyl peptidase IV. Journal of Cell Biology, 111(2):323–??, August 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/111/2/
323.

Hartwig:1991:CRH

[HD91a] J. H. Hartwig and M. DeSisto. The cytoskeleton of the resting human blood platelet: structure of the membrane skeleton and its attachment to actin filaments. Journal of Cell Biology, 112(3):407–??, February 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/112/3/407.

High:1991:SSI

[HD91b] S. High and B. Dobberstein. The signal sequence interacts with the methionine-rich domain of the 54-kD protein of signal recognition particle. Journal of Cell Biology, 113(2): 229–??, April 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/113/2/229.

Hemmer:1991:AEL

[HDC^{+}91] R. M. Hemmer, S. G. Donkin, K. J. Chin, D. G. Grenache, H. Bhatt, and S. M. Politz. Altered expression of an L1specific, O-linked cuticle surface glycoprotein in mutants of the nematode Caenorhabditis elegans. Journal of Cell Biology, 115(5):1237–??, December 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/115/5/1237.

Hamill:1994:PTM

[HDDS94] D. Hamill, J. Davis, J. Drawbridge, and K. A. Suprenant. Polyribosome targeting to microtubules: enrichment of specific mRNAs in a reconstituted microtubule preparation from sea urchin embryos. Journal of Cell Biology, 127(4): 973–??, November 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/127/4/973.

Huang:1994:VAS

[HDES94] S. Huang, T. J. Deerinck, M. H. Ellisman, and D. L. Spector. In vivo analysis of the stability and transport of nuclear poly(A)+ RNA. Journal of Cell Biology, 126(4):877–??, August 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/126/4/877.

Huber:1993:PTD

[HdHD^{+}93] L. A. Huber, M. J. de Hoop, P. Dupree, M. Zerial, K. Simons, and C. Dotti. Protein transport to the dendritic plasma membrane of cultured neurons is regulated by rab8p. Journal of Cell Biology, 123(1):47–??, October 1993. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/123/1/
47.

Hiraoka:1993:OHC

[HDP^{+}93] Y. Hiraoka, A. F. Dernburg, S. J. Parmelee, M. C. Rykowski, D. A. Agard, and J. W. Sedat. The onset of homologous chromosome pairing during Drosophila melanogaster embryogenesis. Journal of Cell Biology, 120 (3):591–??, February 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/120/3/591.

Heimark:1990:ICD

[HDS90] R. L. Heimark, M. Degner, and S. M. Schwartz. Identification of a Ca2(+)-dependent cell–cell adhesion molecule in endothelial cells. Journal of Cell Biology, 110(5):1745– ??, May 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/110/5/1745.

Hastings:1991:MFD

[HE91a] G. A. Hastings and C. P. Emerson. Myosin functional domains encoded by alternative exons are expressed in specific thoracic muscles of Drosophila. Journal of Cell Biology, 114 (2):263–??, July 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/114/2/263.

Houliston:1991:EIM

[HE91b] E. Houliston and R. P. Elinson. Evidence for the involvement of microtubules, ER, and kinesin in the cortical rotation of fertilized frog eggs. Journal of Cell Biology, 114 (5):1017–??, September 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/114/5/1017.

Ha:1993:AAP

[HE93] K. S. Ha and J. H. Exton. Activation of actin polymerization by phosphatidic acid derived from phosphatidylcholine in IIC9 fibroblasts. Journal of Cell Biology, 123(6):1789– ??, December 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/123/6/1789.

Huovila:1992:HBS

[HEF92] A. P. Huovila, A. M. Eder, and S. D. Fuller. Hepatitis B surface antigen assembles in a post–ER, pre-Golgi compartment. Journal of Cell Biology, 118(6):1305–??, September 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/118/6/1305.

Hendershot:1990:IHC

[Hen90] L. M. Hendershot. Immunoglobulin heavy chain and binding protein complexes are dissociated in vivo by light chain addition. Journal of Cell Biology, 111(3):829–??, September 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/111/3/829.

Hirose:1994:SCM

[HFAGM94] K. Hirose, C. Franzini-Armstrong, Y. E. Goldman, and J. M. Murray. Structural changes in muscle crossbridges accompanying force generation. Journal of Cell Biology, 127(3):763–??, November 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/127/3/763.

High:1991:RMI

[HFD91] S. High, N. Flint, and B. Dobberstein. Requirements for the membrane insertion of signal-anchor type proteins. Journal of Cell Biology, 113(1):25–??, April 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/113/1/25.

Hackl:1994:KPA

[HFL94] W. Hackl, U. Fischer, and R. Lührmann. A 69-kD protein that associates reversibly with the Sm core domain of several spliceosomal snRNP species. Journal of Cell Biology, 124(3):261–??, February 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/124/3/261.

Han:1992:RIL

[HFN92] J. K. Han, K. Fukami, and R. Nuccitelli. Reducing inositol lipid hydrolysis, Ins(1,4,5)P3 receptor availability, or Ca^{2+} gradients lengthens the duration of the cell cycle in Xenopus laevis blastomeres. Journal of Cell Biology, 116(1): 147–??, January 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/116/1/147.

Husmann:1992:TPC

[HFS92] K. Husmann, A. Faissner, and M. Schachner. Tenascin promotes cerebellar granule cell migration and neurite outgrowth by different domains in the fibronectin type III repeats. Journal of Cell Biology, 116(6):1475–??, March 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/116/6/
1475.

Hudson:1993:TIR

[HFS^{+}93] A. W. Hudson, D. C. Fingar, G. A. Seidner, G. Griffiths, B. Burke, and M. J. Birnbaum. Targeting of the “insulinresponsive” glucose transporter (GLUT4) to the regulated secretory pathway in PC12 cells. Journal of Cell Biology, 122(3):579–??, August 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/122/3/579.

Harding:1992:CIM

[HG92] C. V. Harding and H. J. Geuze. Class II MHC molecules are present in macrophage lysosomes and phagolysosomes that function in the phagocytic processing of Listeria monocytogenes for presentation to T cells. Journal of Cell Biology, 119(3):531–??, November 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/119/3/531.

Harder:1993:SDE

[HG93a] T. Harder and V. Gerke. The subcellular distribution of early endosomes is affected by the annexin II2p11(2) complex. Journal of Cell Biology, 123(5):1119–??, December 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/123/5/1119.

Hill:1993:BGA

[HG93b] M. A. Hill and P. Gunning. Beta and gamma actin mRNAs are differentially located within myoblasts. Journal of Cell Biology, 122(4):825–??, August 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/122/4/825.

He:1994:SRF

[HG94] Y. He and F. Grinnell. Stress relaxation of fibroblasts activates a cyclic AMP signaling pathway. Journal of Cell Biology, 126(2):457–??, July 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/126/2/457.

Hoyvik:1991:IAL

[HGB^{+}91] H. Høyvik, P. B. Gordon, T. O. Berg, P. E. Strømhaug, and P. O. Seglen. Inhibition of autophagic-lysosomal delivery and autophagic lactolysis by asparagine. Journal of Cell Biology, 113(6):1305–??, June 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/113/6/1305.

Han:1993:CCM

[HGDA93] E. K. Han, T. M. Guadagno, S. L. Dalton, and R. K. Assoian. A cell cycle and mutational analysis of anchorageindependent growth: cell adhesion and TGF-beta 1 control G1/S transit specifically. Journal of Cell Biology, 122 (2):461–??, July 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/122/2/461.

Hartl:1993:MRT

[HGF93] P. Hartl, J. Gottesfeld, and D. J. Forbes. Mitotic repression of transcription in vitro. Journal of Cell Biology, 120(3):

613–??, February 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/120/3/613.

Hopkins:1994:MFR

[HGS^{+}94] C. R. Hopkins, A. Gibson, M. Shipman, D. K. Strickland, and I. S. Trowbridge. In migrating fibroblasts, recycling receptors are concentrated in narrow tubules in the pericentriolar area, and then routed to the plasma membrane of the leading lamella. Journal of Cell Biology, 125(6):1265– ??, June 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/125/6/1265.

High:1991:IPP

[HGW^{+}91] S. High, D. Görlich, M. Wiedmann, T. A. Rapoport, and B. Dobberstein. The identification of proteins in the proximity of signal-anchor sequences during their targeting to and insertion into the membrane of the ER. Journal of Cell Biology, 113(1):35–??, April 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/113/1/35.

Hughson:1990:EPP

[HH90] E. J. Hughson and C. R. Hopkins. Endocytic pathways in polarized Caco-2 cells: identification of an endosomal compartment accessible from both apical and basolateral surfaces. Journal of Cell Biology, 110(2):337–??, February 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/110/2/337.

Hammond:1994:QCS

[HH94a] C. Hammond and A. Helenius. Quality control in the secretory pathway: retention of a misfolded viral membrane glycoprotein involves cycling between the ER, intermediate compartment, and Golgi apparatus. Journal of Cell Biology, 126(1):41–??, July 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/126/1/41.

Hohfeld:1994:RCC

[HH94b] J. Höhfeld and F. U. Hartl. Role of the chaperonin cofactor Hsp10 in protein folding and sorting in yeast mitochondria. Journal of Cell Biology, 126(2):305–??, July 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/126/2/305.

Hess:1993:MCM

[HHF93] H. Hess, H. Heid, and W. W. Franke. Molecular characterization of mammalian cylicin, a basic protein of the sperm head cytoskeleton. Journal of Cell Biology, 122(5):1043– ??, September 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/122/5/1043.

Howard:1994:RCF

[HHG^{+}94] D. R. Howard, G. Habermacher, D. B. Glass, E. F. Smith, and W. S. Sale. Regulation of Chlamydomonas flagellar dynein by an axonemal protein kinase. Journal of Cell Biology, 127(6):1683–??, December 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/127/6/1683.

Heider:1993:HHR

[HHH^{+}93] K. H. Heider, M. Hofmann, E. Hors, F. van den Berg, H. Ponta, P. Herrlich, and S. T. Pals. A human homologue of the rat metastasis-associated variant of CD44 is expressed in colorectal carcinomas and adenomatous polyps. Journal of Cell Biology, 120(1):227–??, January 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/120/1/227.

Hitt:1994:PMH

[HHL94] A. L. Hitt, J. H. Hartwig, and E. J. Luna. Ponticulin is the major high affinity link between the plasma membrane and the cortical actin network in Dictyostelium. Journal of Cell Biology, 126(6):1433–??, September 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/126/6/1433.

Hoyt:1992:TSC

[HHLS92] M. A. Hoyt, L. He, K. K. Loo, and W. S. Saunders. Two Saccharomyces cerevisiae kinesin-related gene products required for mitotic spindle assembly. Journal of Cell Biology, 118(1):109–??, July 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/118/1/109.

Hardwick:1992:MCN

[HHO^{+}92] C. Hardwick, K. Hoare, R. Owens, H. P. Hohn, M. Hook, D. Moore, V. Cripps, L. Austen, D. M. Nance, and E. A. Turley. Molecular cloning of a novel hyaluronan receptor that mediates tumor cell motility. Journal of Cell Biology, 117(6):1343–??, June 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/117/6/1343.

Hayashi:1990:EFC

[HHR^{+}90] Y. Hayashi, B. Haimovich, A. Reszka, D. Boettiger, and A. Horwitz. Expression and function of chicken integrin beta 1 subunit and its cytoplasmic domain mutants in mouse NIH 3T3 cells. Journal of Cell Biology, 110(1):175–??, January 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/110/1/175.

Hosoya:1993:LCD

[HHY^{+}93] N. Hosoya, H. Hosoya, S. Yamashiro, H. Mohri, and F. Matsumura. Localization of caldesmon and its dephosphorylation during cell division. Journal of Cell Biology, 121(5): 1075–??, June 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/121/5/1075.

Holifield:1990:CBM

[HIJ90] B. F. Holifield, A. Ishihara, and K. Jacobson. Comparative behavior of membrane protein-antibody complexes on motile fibroblasts: implications for a mechanism of capping. Journal of Cell Biology, 111(6):2499–??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/111/6/
2499.

Hahne:1993:FTN

[HJI^{+}93] M. Hahne, U. Jäger, S. Isenmann, R. Hallmann, and D. Vestweber. Five tumor necrosis factor-inducible cell adhesion mechanisms on the surface of mouse endothelioma cells mediate the binding of leukocytes. Journal of Cell Biology, 121(3):655–??, May 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/121/3/655.

Hausser:1991:BHS

[HK91] H. Hausser and H. Kresse. Binding of heparin and of the small proteoglycan decorin to the same endocytosis receptor proteins leads to different metabolic consequences. Journal of Cell Biology, 114(1):45–??, July 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/114/1/45.

Heller:1994:TNF

[HK94] R. A. Heller and M. Krönke. Tumor necrosis factor receptormediated signaling pathways. Journal of Cell Biology, 126 (1):5–??, July 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/126/1/5.

Hemmings:1992:AAB

[HKC92] L. Hemmings, P. A. Kuhlman, and D. R. Critchley. Analysis of the actin-binding domain of alpha-actinin by mutagenesis and demonstration that dystrophin contains a functionally homologous domain. Journal of Cell Biology, 116(6): 1369–??, March 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/116/6/1369.

Hara-Kuge:1994:BIC

[HKKO^{+}94] S. Hara-Kuge, O. Kuge, L. Orci, M. Amherdt, M. Ravazzola, F. T. Wieland, and J. E. Rothman. En bloc incorporation of coatomer subunits during the assembly of COP-coated vesicles. Journal of Cell Biology, 124(6):883– ??, March 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/124/6/883.

Hunnicutt:1990:CBF

[HKS90] G. R. Hunnicutt, M. G. Kosfiszer, and W. J. Snell. Cell body and flagellar agglutinins in Chlamydomonas reinhardtii: the cell body plasma membrane is a reservoir for agglutinins whose migration to the flagella is regulated by a functional barrier. Journal of Cell Biology, 111(4):1605–??, October 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/111/4/1605.

Hauft:1992:ESA

[HKS^{+}92] S. M. Hauft, S. H. Kim, G. H. Schmidt, S. Pease, S. Rees, S. Harris, K. A. Roth, J. R. Hansbrough, S. M. Cohn, and D. J. Ahnen. Expression of SV-40 T antigen in the small intestinal epithelium of transgenic mice results in proliferative changes in the crypt and reentry of villus-associated enterocytes into the cell cycle but has no apparent effect on cellular differentiation programs and does not cause neoplastic transformation. Journal of Cell Biology, 117(4):825– ??, May 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/117/4/825.

Henis:1990:OSH

[HKSL90] Y. I. Henis, Z. Katzir, M. A. Shia, and H. F. Lodish. Oligomeric structure of the human asialoglycoprotein receptor: nature and stoichiometry of mutual complexes containing H1 and H2 polypeptides assessed by fluorescence photobleaching recovery. Journal of Cell Biology, 111(4): 1409–??, October 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/111/4/1409.

Haarer:1990:PPS

[HLA^{+}90] B. K. Haarer, S. H. Lillie, A. E. Adams, V. Magdolen, W. Bandlow, and S. S. Brown. Purification of profilin from Saccharomyces cerevisiae and analysis of profilin-deficient cells. Journal of Cell Biology, 110(1):105–??, January 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/110/1/
105.

Heidemann:1990:GCB

[HLB90] S. R. Heidemann, P. Lamoureux, and R. E. Buxbaum. Growth cone behavior and production of traction force. Journal of Cell Biology, 111(5):1949–??, November 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/111/5/
1949.

Huang:1993:ALB

[HLC^{+}93] M. M. Huang, L. Lipfert, M. Cunningham, J. S. Brugge, M. H. Ginsberg, and S. J. Shattil. Adhesive ligand binding to integrin alpha IIb beta 3 stimulates tyrosine phosphorylation of novel protein substrates before phosphorylation of pp125FAK. Journal of Cell Biology, 122(2):473– ??, July 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/122/2/473.

Houzelstein:1992:LDG

[HLCB92] D. Houzelstein, G. E. Lyons, J. Chamberlain, and M. E. Buckingham. Localization of dystrophin gene transcripts during mouse embryogenesis. Journal of Cell Biology, 119 (4):811–??, November 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/119/4/811.

He:1992:MIM

[HLH^{+}92] Q. He, J. Lesley, R. Hyman, K. Ishihara, and P. W. Kincade. Molecular isoforms of murine CD44 and evidence that the membrane proximal domain is not critical for hyaluronate recognition. Journal of Cell Biology, 119(6):1711–??, December 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/119/6/1711.

Hitt:1994:PAM

[HLL94] A. L. Hitt, T. H. Lu, and E. J. Luna. Ponticulin is an atypical membrane protein. Journal of Cell Biology, 126 (6):1421–??, September 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/126/6/1421.

Hou:1990:BMI

[HLPL90] L. Hou, K. Luby-Phelps, and F. Lanni. Brownian motion of inert tracer macromolecules in polymerized and spontaneously bundled mixtures of actin and filamin. Journal of Cell Biology, 110(5):1645–??, May 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/110/5/1645.

Hunt:1992:RPS

[HLR92] T. Hunt, F. C. Luca, and J. V. Ruderman. The requirements for protein synthesis and degradation, and the control of destruction of cyclins A and B in the meiotic and mitotic cell cycles of the clam embryo. Journal of Cell Biology, 116(3):707–??, February 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/116/3/707.

Hannan:1993:CSM

[HLRBE93] L. A. Hannan, M. P. Lisanti, E. Rodriguez-Boulan, and M. Edidin. Correctly sorted molecules of a GPI-anchored protein are clustered and immobile when they arrive at the apical surface of MDCK cells. Journal of Cell Biology, 120(2):353–??, January 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/120/2/353.

Hoffman:1992:ATC

[HLWL92] P. N. Hoffman, M. A. Lopata, D. F. Watson, and R. F. Luduena. Axonal transport of class II and III beta-tubulin: evidence that the slow component wave represents the movement of only a small fraction of the tubulin in mature motor axons. Journal of Cell Biology, 119(3):595–??, November 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/119/3/595.

Hyman:1990:MMS

[HM90] A. A. Hyman and T. J. Mitchison. Modulation of microtubule stability by kinetochores in vitro. Journal of Cell Biology, 110(5):1607–??, May 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/110/5/1607.

Hirano:1991:CCC

[HM91] T. Hirano and T. J. Mitchison. Cell cycle control of higherorder chromatin assembly around naked DNA in vitro. Journal of Cell Biology, 115(6):1479–??, December 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/115/6/

Harter:1992:TLM

[HM92a] C. Harter and I. Mellman. Transport of the lysosomal membrane glycoprotein lgp120 (lgp–A) to lysosomes does not require appearance on the plasma membrane. Journal of Cell Biology, 117(2):311–??, April 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/117/2/311.

Hay:1992:RRS

[HM92b] J. C. Hay and T. F. Martin. Resolution of regulated secretion into sequential MgATP-dependent and calciumdependent stages mediated by distinct cytosolic proteins. Journal of Cell Biology, 119(1):139–??, October 1992. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/119/1/
139.

Hirano:1993:TID

[HM93] T. Hirano and T. J. Mitchison. Topoisomerase II does not play a scaffolding role in the organization of mitotic chromosomes assembled in Xenopus egg extracts. Journal of Cell Biology, 120(3):601–??, February 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/120/3/601.

Hasson:1994:PMV

[HM94] T. Hasson and M. S. Mooseker. Porcine myosin-VI: characterization of a new mammalian unconventional myosin. Journal of Cell Biology, 127(2):425–??, October 1994. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/127/2/
425.

Huang:1993:ECC

[HMB^{+}93] A. J. Huang, J. E. Manning, T. M. Bandak, M. C. Ratau, K. R. Hanser, and S. C. Silverstein. Endothelial cell cytosolic free calcium regulates neutrophil migration across monolayers of endothelial cells. Journal of Cell Biology, 120(6): 1371–??, March 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/120/6/1371.

Hartman:1991:ESF

[HMC91] D. S. Hartman, N. S. Millar, and T. Claudio. Extracellular synaptic factors induce clustering of acetylcholine receptors stably expressed in fibroblasts. Journal of Cell Biology, 115(1):165–??, October 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/115/1/165.

Henis:1994:TII

[HMLL94] Y. I. Henis, A. Moustakas, H. Y. Lin, and H. F. Lodish. The types II and III transforming growth factor-beta receptors form homo-oligomers. Journal of Cell Biology, 126 (1):139–??, July 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/126/1/139.

Haugen:1990:RCC

[HMS^{+}90] P. K. Haugen, J. B. McCarthy, A. P. Skubitz, L. T. Furcht, and P. C. Letourneau. Recognition of the A chain carboxyterminal heparin binding region of fibronectin involves multiple sites: two contiguous sequences act independently to promote neural cell adhesion. Journal of Cell Biology, 111 (6):2733–??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/111/6/2733.

Heyman:1994:RPG

[HMS94] J. A. Heyman, E. Monosov, and S. Subramani. Role of the PAS1 gene of Pichia pastoris in peroxisome biogenesis. Journal of Cell Biology, 127(5):1259–??, December 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/127/5/
1259.

Hug:1992:ICA

[HMT^{+}92] C. Hug, T. M. Miller, M. A. Torres, J. F. Casella, and J. A. Cooper. Identification and characterization of an actin-binding site of CapZ. Journal of Cell Biology, 116(4): 923–??, February 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/116/4/923.

Hayashi:1992:ECI

[HMY92] K. Hayashi, J. A. Madri, and P. D. Yurchenco. Endothelial cells interact with the core protein of basement membrane perlecan through beta 1 and beta 3 integrins: an adhesion modulated by glycosaminoglycan. Journal of Cell Biology, 119(4):945–??, November 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/119/4/945.

Han:1990:ITI

[HN90] J. K. Han and R. Nuccitelli. Inositol 1,4,5-trisphosphate-induced calcium release in the organelle layers of the stratified, intact egg of Xenopus laevis. Journal of Cell Biology, 110(4):1103–??, April 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/110/4/1103.

Horne:1992:OEH

[HNC^{+}92] W. C. Horne, L. Neff, D. Chatterjee, A. Lomri, J. B. Levy, and R. Baron. Osteoclasts express high levels of pp60csrc in association with intracellular membranes. Journal of Cell Biology, 119(4):1003–??, November 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/119/4/1003.

Hoock:1991:BAM

[HNH91] T. C. Hoock, P. M. Newcomb, and I. M. Herman. Beta actin and its mRNA are localized at the plasma membrane and the regions of moving cytoplasm during the cellular response to injury. Journal of Cell Biology, 112(4):653–??, February 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/112/4/653.

He:1990:CFN

[HNP90] D. C. He, J. A. Nickerson, and S. Penman. Core filaments of the nuclear matrix. Journal of Cell Biology, 110(3):569–??, March 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/110/3/569.

Hinck:1994:WMC

[HNP94] L. Hinck, W. J. Nelson, and J. Papkoff. Wnt-1 modulates cell–cell adhesion in mammalian cells by stabilizing betacatenin binding to the cell adhesion protein cadherin. Journal of Cell Biology, 124(5):729–??, March 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/124/5/729.

Hinck:1994:DCC

[HNPN94] L. Hinck, I. S. Näthke, J. Papkoff, and W. J. Nelson. Dynamics of cadherin/catenin complex formation: novel protein interactions and pathways of complex assembly. Journal of Cell Biology, 125(6):1327–??, June 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/125/6/1327.

Hall:1990:BPD

[HNR90] A. K. Hall, R. Nelson, and U. Rutishauser. Binding properties of detergent-solubilized NCAM. Journal of Cell Biology, 110(3):817–??, March 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/110/3/817.

Hieda:1992:INH

[HNUO92] Y. Hieda, Y. Nishizawa, J. Uematsu, and K. Owaribe. Identification of a new hemidesmosomal protein, HD1: a major, high molecular mass component of isolated hemidesmosomes. Journal of Cell Biology, 116(6):1497–??, March 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/116/6/
1497.

Harkness:1994:CRM

[HNvdK^{+}94] T. A. Harkness, F. E. Nargang, I. van der Klei, W. Neupert, and R. Lill. A crucial role of the mitochondrial protein import receptor MOM19 for the biogenesis of mitochondria. Journal of Cell Biology, 124(5):637–??, March 1994. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/124/5/

Horio:1994:HGT

[HO94] T. Horio and B. R. Oakley. Human gamma-tubulin functions in fission yeast. Journal of Cell Biology, 126(6):1465– ??, September 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/126/6/1465.

Hartl:1994:NAL

[HODF94] P. Hartl, E. Olson, T. Dang, and D. J. Forbes. Nuclear assembly with lambda DNA in fractionated Xenopus egg extracts: an unexpected role for glycogen in formation of a higher order chromatin intermediate. Journal of Cell Biology, 124(3):235–??, February 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/124/3/235.

Hollenbeck:1993:PEA

[Hol93] P. J. Hollenbeck. Products of endocytosis and autophagy are retrieved from axons by regulated retrograde organelle transport. Journal of Cell Biology, 121(2):305–??, April 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/
121/2/305.

Hardwick:1992:SEK

[HP92] K. G. Hardwick and H. R. Pelham. SED5 encodes a 39-kD integral membrane protein required for vesicular transport between the ER and the Golgi complex. Journal of Cell Biology, 119(3):513–??, November 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/119/3/513.

Heitlinger:1991:ECL

[HPH^{+}91] E. Heitlinger, M. Peter, M. Häner, A. Lustig, U. Aebi, and E. A. Nigg. Expression of chicken lamin B2 in Escherichia coli: characterization of its structure, assembly, and molecular interactions. Journal of Cell Biology, 113(3):485–??, May 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/113/3/485.

Hager:1994:ECH

[HPM^{+}94] K. M. Hager, M. A. Pierce, D. R. Moore, E. M. Tytler, J. D. Esko, and S. L. Hajduk. Endocytosis of a cytotoxic human high density lipoprotein results in disruption of acidic intracellular vesicles and subsequent killing of African trypanosomes. Journal of Cell Biology, 126(1):155– ??, July 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/126/1/155.

Heck:1993:KLP

[HPP^{+}93a] M. M. Heck, A. Pereira, P. Pesavento, Y. Yannoni, A. C. Spradling, and L. S. Goldstein. The kinesin-like protein KLP61F is essential for mitosis in Drosophila. Journal of Cell Biology, 123(3):665–??, November 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/123/3/665.

Huber:1993:RSG

[HPP^{+}93b] L. A. Huber, S. Pimplikar, R. G. Parton, H. Virta, M. Zerial, and K. Simons. Rab8, a small GTPase involved in vesicular traffic between the TGN and the basolateral plasma membrane. Journal of Cell Biology, 123(1):35–??, October 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/123/1/
35.

Helms:1993:TDP

[HPR93] J. B. Helms, D. J. Palmer, and J. E. Rothman. Two distinct populations of ARF bound to Golgi membranes. Journal of Cell Biology, 121(4):751–??, May 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/121/4/751.

Hewlett:1994:CPM

[HPW94] L. J. Hewlett, A. R. Prescott, and C. Watts. The coated pit and macropinocytic pathways serve distinct endosome populations. Journal of Cell Biology, 124(5):689–??, March 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/
124/5/689.

Hennekes:1993:PPK

[HPWN93] H. Hennekes, M. Peter, K. Weber, and E. A. Nigg. Phosphorylation on protein kinase C sites inhibits nuclear import of lamin B2. Journal of Cell Biology, 120(6):1293–??, March 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/120/6/1293.

Humphrey:1993:LTT

[HPYB93] J. S. Humphrey, P. J. Peters, L. C. Yuan, and J. S. Bonifacino. Localization of TGN38 to the trans-Golgi network: involvement of a cytoplasmic tyrosine-containing sequence. Journal of Cell Biology, 120(5):1123–??, March 1993. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/120/5/
1123.

Haltmeier:1990:DDE

[HR90a] H. Haltmeier and H. Rohrer. Distinct and different effects of the oncogenes v-myc and v-src on avian sympathetic neurons: retroviral transfer of v-myc stimulates neuronal proliferation whereas v-src transfer enhances neuronal differentiation. Journal of Cell Biology, 110(6):2087–??, June 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/110/6/
2087.

Hoyle:1990:TDB

[HR90b] H. D. Hoyle and E. C. Raff. Two Drosophila beta tubulin isoforms are not functionally equivalent. Journal of Cell Biology, 111(3):1009–??, September 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/111/3/1009.

Hyrc:1990:AVS

[HR90c] K. Hyrc and B. Rose. The action of v-src on gap junctional permeability is modulated by pH. Journal of Cell Biology, 110(4):1217–??, April 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/110/4/1217.

Hampton:1994:RDH

[HR94a] R. Y. Hampton and J. Rine. Regulated degradation of HMG–CoA reductase, an integral membrane protein of the endoplasmic reticulum, in yeast. Journal of Cell Biology, 125(2):299–??, April 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/125/2/299.

Hinek:1994:KEB

[HR94b] A. Hinek and M. Rabinovitch. 67-kD elastin-binding protein is a protective “companion” of extracellular insoluble elastin and intracellular tropoelastin. Journal of Cell Biology, 126 (2):563–??, July 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/126/2/563.

Hudson:1992:ISS

[HRB92] A. W. Hudson, M. Ruiz, and M. J. Birnbaum. Isoformspecific subcellular targeting of glucose transporters in mouse fibroblasts. Journal of Cell Biology, 116(3):785–??, February 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/116/3/785.

Hall:1990:ABA

[HRC^{+}90] D. E. Hall, L. F. Reichardt, E. Crowley, B. Holley, H. Moezzi, A. Sonnenberg, and C. H. Damsky. The alpha 1/beta 1 and alpha 6/beta 1 integrin heterodimers mediate cell attachment to distinct sites on laminin. Journal of Cell Biology, 110(6):2175–??, June 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/110/6/2175.

Hagan:1990:ICF

[HRH90] I. M. Hagan, P. N. Riddle, and J. S. Hyams. Intramitotic controls in the fission yeast Schizosaccharomyces pombe: the effect of cell size on spindle length and the timing of mitotic events. Journal of Cell Biology, 110(5):1617–??, May 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/110/5/

Haines:1991:EPN

[HRT^{+}91] K. A. Haines, J. Reibman, X. Y. Tang, M. Blake, and G. Weissmann. Effects of protein I of Neisseria gonorrhoeae on neutrophil activation: generation of diacylglycerol from phosphatidylcholine via a specific phospholipase C is associated with exocytosis. Journal of Cell Biology, 114(3): 433–??, August 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/114/3/433.

Hays:1990:PFK

[HS90a] T. S. Hays and E. D. Salmon. Poleward force at the kinetochore in metaphase depends on the number of kinetochore microtubules. Journal of Cell Biology, 110(2):391–??, February 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/110/2/391.

Heck:1990:MRO

[HS90b] M. M. Heck and A. C. Spradling. Multiple replication origins are used during Drosophila chorion gene amplification. Journal of Cell Biology, 110(4):903–??, April 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/110/4/903.

Hathaway:1992:CSB

[HS92] H. J. Hathaway and B. D. Shur. Cell surface beta 1,4galactosyltransferase functions during neural crest cell migration and neurulation in vivo. Journal of Cell Biology, 117 (2):369–??, April 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/117/2/369.

Honts:1994:AMS

[HSB^{+}94] J. E. Honts, T. S. Sandrock, S. M. Brower, J. L. O’Dell, and A. E. Adams. Actin mutations that show suppression with fimbrin mutations identify a likely fimbrin-binding site on actin. Journal of Cell Biology, 126(2):413–??, July 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/126/2/

Herrmann:1994:MHS

[HSCN94] J. M. Herrmann, R. A. Stuart, E. A. Craig, and W. Neupert. Mitochondrial heat shock protein 70, a molecular chaperone for proteins encoded by mitochondrial DNA. Journal of Cell Biology, 127(4):893–??, November 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/127/4/893.

Hansen:1994:LHG

[HSCR94] C. A. Hansen, A. G. Schroering, D. J. Carey, and J. D. Robishaw. Localization of a heterotrimeric G protein gamma subunit to focal adhesions and associated stress fibers. Journal of Cell Biology, 126(3):811–??, August 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/126/3/811.

Hill:1994:SIS

[HSG94] M. A. Hill, L. Schedlich, and P. Gunning. Serum-induced signal transduction determines the peripheral location of beta-actin mRNA within the cell. Journal of Cell Biology, 126(5):1221–??, September 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/126/5/1221.

Hanukoglu:1990:IML

[HSHA90] I. Hanukoglu, B. S. Suh, S. Himmelhoch, and A. Amsterdam. Induction and mitochondrial localization of cytochrome P450scc system enzymes in normal and transformed ovarian granulosa cells. Journal of Cell Biology, 111(4):1373–??, October 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/111/4/1373.

Hennemann:1992:MCF

[HSLF^{+}92] H. Hennemann, T. Suchyna, H. Lichtenberg-Fraté, S. Jungbluth, E. Dahl, J. Schwarz, B. J. Nicholson, and K. Willecke. Molecular cloning and functional expression of mouse connexin40, a second gap junction gene preferentially expressed in lung. Journal of Cell Biology, 117(6):1299–??, June 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/117/6/

Horstkorte:1993:FIL

[HSM^{+}93] R. Horstkorte, M. Schachner, J. P. Magyar, T. Vorherr, and B. Schmitz. The fourth immunoglobulin-like domain of NCAM contains a carbohydrate recognition domain for oligomannosidic glycans implicated in association with L1 and neurite outgrowth. Journal of Cell Biology, 121(6): 1409–??, June 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/121/6/1409.

Haney:1991:ITI

[HSP^{+}91] P. M. Haney, J. W. Slot, R. C. Piper, D. E. James, and M. Mueckler. Intracellular targeting of the insulinregulatable glucose transporter (GLUT4) is isoform specific and independent of cell type. Journal of Cell Biology, 114(4):689–??, August 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/114/4/689.

Hildebrand:1993:ISR

[HSP93] J. D. Hildebrand, M. D. Schaller, and J. T. Parsons. Identification of sequences required for the efficient localization of the focal adhesion kinase, pp125FAK, to cellular focal adhesions. Journal of Cell Biology, 123(4):993–??, November 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/123/4/993.

Harper:1993:PNF

[HSS93] J. D. Harper, M. A. Sanders, and J. L. Salisbury. Phosphorylation of nuclear and flagellar basal apparatus proteins during flagellar regeneration in Chlamydomonas reinhardtii. Journal of Cell Biology, 122(4):877–??, August 1993. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/122/4/
877.

Hogan:1992:PEI

[HSSC92] C. J. Hogan, L. Stephens, T. Shimizu, and W. Z. Cande. Physiological evidence for involvement of a kinesin-related protein during anaphase spindle elongation in diatom central spindles. Journal of Cell Biology, 119(5):1277–??, December 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/119/5/1277.

Honegger:1990:SEP

[HSUS90] A. M. Honegger, A. Schmidt, A. Ullrich, and J. Schlessinger. Separate endocytic pathways of kinase-defective and -active EGF receptor mutants expressed in same cells. Journal of Cell Biology, 110(5):1541–??, May 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/110/5/1541.

Hansen:1991:PCC

[HSvD91] S. H. Hansen, K. Sandvig, and B. van Deurs. The preendosomal compartment comprises distinct coated and noncoated endocytic vesicle populations. Journal of Cell Biology, 113(4):731–??, May 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/113/4/731.

Hansen:1993:CHA

[HSvD93a] S. H. Hansen, K. Sandvig, and B. van Deurs. Clathrin and HA2 adaptors: effects of potassium depletion, hypertonic medium, and cytosol acidification. Journal of Cell Biology, 121(1):61–??, April 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/121/1/61.

Hansen:1993:MIC

[HSvD93b] S. H. Hansen, K. Sandvig, and B. van Deurs. Molecules internalized by clathrin-independent endocytosis are delivered to endosomes containing transferrin receptors. Journal of Cell Biology, 123(1):89–??, October 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/123/1/89.

Ho:1990:CBAb

[HSW90] S. C. Ho, M. Schindler, and J. L. Wang. Carbohydrate binding activities of Bradyrhizobium japonicum. II. Isolation and characterization of a galactose-specific lectin. Journal of Cell Biology, 111(4):1639–??, October 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/111/4/1639.

Hirokawa:1991:KAA

[HSYK^{+}91] N. Hirokawa, R. Sato-Yoshitake, N. Kobayashi, K. K. Pfister, G. S. Bloom, and S. T. Brady. Kinesin associates with anterogradely transported membranous organelles in vivo. Journal of Cell Biology, 114(2):295–??, July 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/114/2/295.

Hirokawa:1990:BDM

[HSYYK90] N. Hirokawa, R. Sato-Yoshitake, T. Yoshida, and T. Kawashima. Brain dynein (MAP1C) localizes on both anterogradely and retrogradely transported membranous organelles in vivo. Journal of Cell Biology, 111(3):1027–??, September 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/111/3/
1027.

Horowitz:1994:APF

[HTBF94] A. Horowitz, K. M. Trybus, D. S. Bowman, and F. S. Fay. Antibodies probe for folded monomeric myosin in relaxed and contracted smooth muscle. Journal of Cell Biology, 126(5):1195–??, September 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/126/5/1195.

Hopper:1990:YRG

[HTD90] A. K. Hopper, H. M. Traglia, and R. W. Dunst. The yeast RNA1 gene product necessary for RNA processing is located in the cytosol and apparently excluded from the nucleus. Journal of Cell Biology, 111(2):309–??, August 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/111/2/
309.

Hayashi:1994:HMM

[HTGN94] J. Hayashi, M. Takemitsu, Y. Goto, and I. Nonaka. Human mitochondria and mitochondrial genome function as a single dynamic cellular unit. Journal of Cell Biology, 125 (1):43–??, April 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/125/1/43.

Haegel:1993:AMA

[HTHC93] H. Haegel, C. Tölg, M. Hofmann, and R. Ceredig. Activated mouse astrocytes and T cells express similar CD44 variants. Role of CD44 in astrocyte/T cell binding. Journal of Cell Biology, 122(5):1067–??, September 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/122/5/1067.

Hamaguchi:1992:RPK

[HTP^{+}92] J. R. Hamaguchi, R. A. Tobey, J. Pines, H. A. Crissman, T. Hunter, and E. M. Bradbury. Requirement for p34cdc2 kinase is restricted to mitosis in the mammalian cdc2 mutant FT210. Journal of Cell Biology, 117(5):1041– ??, June 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/117/5/1041.

Hogan:1994:NLH

[HTSP94] N. C. Hogan, K. L. Traverse, D. E. Sullivan, and M. L. Pardue. The nucleus-limited Hsr-omega-n transcript is a polyadenylated RNA with a regulated intranuclear turnover. Journal of Cell Biology, 125(1):21–??, April 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/125/1/
21.

Hurt:1990:TCP

[Hur90] E. C. Hurt. Targeting of a cytosolic protein to the nuclear periphery. Journal of Cell Biology, 111(6):2829–??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/111/6/2829.

Hohfeld:1991:PSC

[HVK91] J. Höhfeld, M. Veenhuis, and W. H. Kunau. PAS3, a Saccharomyces cerevisiae gene encoding a peroxisomal integral membrane protein essential for peroxisome biogenesis. Journal of Cell Biology, 114(6):1167–??, September 1991. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/114/6/

Hung:1992:TIE

[HVNC92] D. T. Hung, T. H. Vu, N. A. Nelken, and S. R. Coughlin. Thrombin-induced events in non-platelet cells are mediated by the unique proteolytic mechanism established for the cloned platelet thrombin receptor. Journal of Cell Biology, 116(3):827–??, February 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/116/3/827.

Hatzfeld:1990:CCV

[HW90] M. Hatzfeld and K. Weber. The coiled coil of in vitro assembled keratin filaments is a heterodimer of type I and II keratins: use of site-specific mutagenesis and recombinant protein expression. Journal of Cell Biology, 110(4): 1199–??, April 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/110/4/1199.

Hatzfeld:1992:SPR

[HW92] M. Hatzfeld and K. Weber. A synthetic peptide representing the consensus sequence motif at the carboxy-terminal end of the rod domain inhibits intermediate filament assembly and disassembles preformed filaments. Journal of Cell Biology, 116(1):157–??, January 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/116/1/157.

Hird:1993:CCF

[HW93a] S. N. Hird and J. G. White. Cortical and cytoplasmic flow polarity in early embryonic cells of Caenorhabditis elegans. Journal of Cell Biology, 121(6):1343–??, June 1993. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/121/6/
1343.

Horst:1993:PNM

[HW93b] C. J. Horst and G. B. Witman. ptx1, a nonphototactic mutant of Chlamydomonas, lacks control of flagellar dominance. Journal of Cell Biology, 120(3):733–??, February 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/120/3/733.

Hodivala:1994:ECP

[HW94] K. J. Hodivala and F. M. Watt. Evidence that cadherins play a role in the downregulation of integrin expression that occurs during keratinocyte terminal differentiation. Journal of Cell Biology, 124(4):589–??, February 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/124/4/589.

Hallberg:1993:IMP

[HWB93] E. Hallberg, R. W. Wozniak, and G. Blobel. An integral membrane protein of the pore membrane domain of the nuclear envelope contains a nucleoporin-like region. Journal of Cell Biology, 122(3):513–??, August 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/122/3/513.

Holtzman:1994:MAS

[HWD94] D. A. Holtzman, K. F. Wertman, and D. G. Drubin. Mapping actin surfaces required for functional interactions in vivo. Journal of Cell Biology, 126(2):423–??, July 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/126/2/
423.

Hobman:1992:RVE

[HWF92] T. C. Hobman, L. Woodward, and M. G. Farquhar. The rubella virus E1 glycoprotein is arrested in a novel post–ER, pre-Golgi compartment. Journal of Cell Biology, 118(4): 795–??, August 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/118/4/795.

Hobman:1993:RVE

[HWF93] T. C. Hobman, L. Woodward, and M. G. Farquhar. The rubella virus E2 and E1 spike glycoproteins are targeted to the Golgi complex. Journal of Cell Biology, 121(2):269–??, April 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/121/2/269.

Hinrichsen:1990:AMB

[HWL^{+}90] R. Hinrichsen, E. Wilson, T. Lukas, T. Craig, J. Schultz, and D. M. Watterson. Analysis of the molecular basis of calmodulin defects that affect ion channel-mediated cellular responses: site-specific mutagenesis and microinjection. Journal of Cell Biology, 111(6):2537–??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/111/6/
2537.

Hall:1994:HHR

[HWLT94] C. L. Hall, C. Wang, L. A. Lange, and E. A. Turley. Hyaluronan and the hyaluronan receptor RHAMM promote focal adhesion turnover and transient tyrosine kinase activity. Journal of Cell Biology, 126(2):575–??, July 1994. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/126/2/
575.

Hayden:1990:BBB

[HWM90] S. M. Hayden, J. S. Wolenski, and M. S. Mooseker. Binding of brush border myosin I to phospholipid vesicles. Journal of Cell Biology, 111(2):443–??, August 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/111/2/443.

Heins:1993:RDN

[HWM^{+}93] S. Heins, P. C. Wong, S. Müller, K. Goldie, D. W. Cleveland, and U. Aebi. The rod domain of NF–L determines neurofilament architecture, whereas the end domains specify filament assembly and network formation. Journal of Cell Biology, 123(6):1517–??, December 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/123/6/1517.

Ho:1990:CBAa

[HWS90] S. C. Ho, J. L. Wang, and M. Schindler. Carbohydrate binding activities of Bradyrhizobium japonicum. I. Saccharidespecific inhibition of homotypic and heterotypic adhesion. Journal of Cell Biology, 111(4):1631–??, October 1990. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/111/4/
1631.

Hille:1990:ALB

[HWvF90] A. Hille, A. Waheed, and K. von Figura. Assembly of the ligand-binding conformation of Mr 46,000 mannose 6phosphate-specific receptor takes place before reaching the Golgi complex. Journal of Cell Biology, 110(4):963–??, April 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/110/4/963.

Hell:1993:IDS

[HWW^{+}93] J. W. Hell, R. E. Westenbroek, C. Warner, M. K. Ahlijanian, W. Prystay, M. M. Gilbert, T. P. Snutch, and W. A. Catterall. Identification and differential subcellular localization of the neuronal class C and class D L-type calcium channel alpha 1 subunits. Journal of Cell Biology, 123(4): 949–??, November 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/123/4/949.

Hresko:1994:ABW

[HWW94] M. C. Hresko, B. D. Williams, and R. H. Waterston. Assembly of body wall muscle and muscle cell attachment structures in Caenorhabditis elegans. Journal of Cell Biology, 124(4):491–??, February 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/124/4/491.

Hart:1994:VSO

[HXM^{+}94] K. C. Hart, Y. F. Xu, A. N. Meyer, B. A. Lee, and D. J. Donoghue. The v-sis oncoprotein loses transforming activity when targeted to the early Golgi complex. Journal of Cell Biology, 127(6):1843–??, December 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/127/6/1843.

Holtzman:1993:SLI

[HYD93] D. A. Holtzman, S. Yang, and D. G. Drubin. Syntheticlethal interactions identify two novel genes, SLA1 and SLA2, that control membrane cytoskeleton assembly in Saccharomyces cerevisiae. Journal of Cell Biology, 122(3): 635–??, August 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/122/3/635.

Heuser:1993:PPP

[HZC93] J. Heuser, Q. Zhu, and M. Clarke. Proton pumps populate the contractile vacuoles of Dictyostelium amoebae. Journal of Cell Biology, 121(6):1311–??, June 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/121/6/1311.

Ioannou:1992:OHA

[IBD92] Y. A. Ioannou, D. F. Bishop, and R. J. Desnick. Overexpression of human alpha-galactosidase A results in its intracellular aggregation, crystallization in lysosomes, and selective secretion. Journal of Cell Biology, 119(5):1137–??, December 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/119/5/1137.

Ishizaki:1994:ASE

[IBR94] Y. Ishizaki, J. F. Burne, and M. C. Raff. Autocrine signals enable chondrocytes to survive in culture. Journal of Cell Biology, 126(4):1069–??, August 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/126/4/1069.

Ivessa:1992:GIT

[IDT^{+}92] N. E. Ivessa, C. De Lemos-Chiarandini, Y. S. Tsao, A. Takatsuki, M. Adesnik, D. D. Sabatini, and G. Kreibich. O-glycosylation of intact and truncated ribophorins in brefeldin A-treated cells: newly synthesized intact ribophorins are only transiently accessible to the relocated glycosyltransferases. Journal of Cell Biology, 117(5):949– ??, June 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/117/5/949.

Irminger-Finger:1990:APS

[IFLG90] I. Irminger-Finger, R. A. Laymon, and L. S. Goldstein. Analysis of the primary sequence and microtubule-binding region of the Drosophila 205K MAP. Journal of Cell Biology, 111(6):2563–??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/111/6/2563.

Iwaki:1994:SAM

[IITG94] T. Iwaki, A. Iwaki, J. Tateishi, and J. E. Goldman. Sense and antisense modification of glial alpha B-crystallin production results in alterations of stress fiber formation and thermoresistance. Journal of Cell Biology, 125(6):1385–??, June 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/125/6/1385.

Illinger:1994:KAI

[IK94] D. Illinger and J. G. Kuhry. The kinetic aspects of intracellular fluorescence labeling with TMA–DPH support the maturation model for endocytosis in L929 cells. Journal of Cell Biology, 125(4):783–??, May 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/125/4/783.

Isaya:1991:CPM

[IKFR91] G. Isaya, F. Kalousek, W. A. Fenton, and L. E. Rosenberg. Cleavage of precursors by the mitochondrial processing peptidase requires a compatible mature protein or an intermediate octapeptide. Journal of Cell Biology, 113(1):65–??, April 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/113/1/65.

Insall:1994:CCP

[IKL^{+}94] R. Insall, A. Kuspa, P. J. Lilly, G. Shaulsky, L. R. Levin, W. F. Loomis, and P. Devreotes. CRAC, a cytosolic protein containing a pleckstrin homology domain, is required for receptor and G protein-mediated activation of adenylyl cyclase in Dictyostelium. Journal of Cell Biology, 126 (6):1537–??, September 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/126/6/1537.

Ignatius:1990:MCR

[ILH^{+}90] M. J. Ignatius, T. H. Large, M. Houde, J. W. Tawil, A. Barton, F. Esch, S. Carbonetto, and L. F. Reichardt. Molecular cloning of the rat integrin alpha 1-subunit: a receptor for laminin and collagen. Journal of Cell Biology, 111(2): 709–??, August 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/111/2/709.

Imamoto:1992:AAK

[IMK^{+}92] N. Imamoto, Y. Matsuoka, T. Kurihara, K. Kohno, M. Miyagi, F. Sakiyama, Y. Okada, S. Tsunasawa, and Y. Yoneda. Antibodies against 70-kD heat shock cognate protein inhibit mediated nuclear import of karyophilic proteins. Journal of Cell Biology, 119(5):1047–??, December 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/119/5/1047.

Ihrke:1993:WBC

[INM^{+}93] G. Ihrke, E. B. Neufeld, T. Meads, M. R. Shanks, D. Cassio, M. Laurent, T. A. Schroer, R. E. Pagano, and A. L. Hubbard. WIF-b cells: an in vitro model for studies of hepatocyte polarity. Journal of Cell Biology, 123(6):1761– ??, December 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/123/6/1761.

Itoh:1993:KPC

[INY^{+}93] M. Itoh, A. Nagafuchi, S. Yonemura, T. Kitani-Yasuda, S. Tsukita, and S. Tsukita. The 220-kD protein colocalizing with cadherins in non-epithelial cells is identical to ZO-1, a tight junction-associated protein in epithelial cells: cDNA cloning and immunoelectron microscopy. Journal of Cell Biology, 121(3):491–??, May 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/121/3/491.

Ingber:1990:CIP

[IPF^{+}90] D. E. Ingber, D. Prusty, J. V. Frangioni, E. J. Cragoe, C. Lechene, and M. A. Schwartz. Control of intracellular pH and growth by fibronectin in capillary endothelial cells. Journal of Cell Biology, 110(5):1803–??, May 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/110/5/1803.

Imhof:1991:ENA

[IRH^{+}91] B. A. Imhof, P. Ruiz, B. Hesse, R. Palacios, and D. Dunon. EA-1, a novel adhesion molecule involved in the homing of progenitor T lymphocytes to the thymus. Journal of Cell Biology, 114(5):1069–??, September 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/114/5/1069.

Izaurralde:1992:CBP

[ISDM92] E. Izaurralde, J. Stepinski, E. Darzynkiewicz, and I. W. Mattaj. A cap binding protein that may mediate nuclear export of RNA polymerase II-transcribed RNAs. Journal of Cell Biology, 118(6):1287–??, September 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/118/6/1287.

Imai:1991:ICB

[ISF^{+}91] Y. Imai, M. S. Singer, C. Fennie, L. A. Lasky, and S. D. Rosen. Identification of a carbohydrate-based endothelial ligand for a lymphocyte homing receptor. Journal of Cell Biology, 113(5):1213–??, June 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/113/5/1213.

Iida:1992:CRC

[ISF^{+}92] J. Iida, A. P. Skubitz, L. T. Furcht, E. A. Wayner, and J. B. McCarthy. Coordinate role for cell surface chondroitin sulfate proteoglycan and alpha 4 beta 1 integrin in mediating melanoma cell adhesion to fibronectin. Journal of Cell Biology, 118(2):431–??, July 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/118/2/431.

Ishida:1994:IDT

[ISN^{+}94] R. Ishida, M. Sato, T. Narita, K. R. Utsumi, T. Nishimoto, T. Morita, H. Nagata, and T. Andoh. Inhibition of DNA topoisomerase II by ICRF-193 induces polyploidization by uncoupling chromosome dynamics from other cell cycle events. Journal of Cell Biology, 126(6):1341–??, September 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/126/6/1341.

Imai:1990:DDL

[ITSR90] Y. Imai, D. D. True, M. S. Singer, and S. D. Rosen. Direct demonstration of the lectin activity of gp90MEL, a lymphocyte homing receptor. Journal of Cell Biology, 111(3):1225– ??, September 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/111/3/1225.

Ishizaki:1993:CLE

[IVBR93] Y. Ishizaki, J. T. Voyvodic, J. F. Burne, and M. C. Raff. Control of lens epithelial cell survival. Journal of Cell Biology, 121(4):899–??, May 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/121/4/899.

Itoh:1991:KUC

[IYN^{+}91] M. Itoh, S. Yonemura, A. Nagafuchi, S. Tsukita, and S. Tsukita. A 220-kD undercoat-constitutive protein: its specific localization at cadherin-based cell–cell adhesion sites. Journal of Cell Biology, 115(5):1449–??, December 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/115/5/1449.

Jans:1991:PMP

[JAB^{+}91] D. A. Jans, M. J. Ackermann, J. R. Bischoff, D. H. Beach, and R. Peters. p34cdc2-mediated phosphorylation at T124 inhibits nuclear import of SV-40 T antigen proteins. Journal of Cell Biology, 115(5):1203–??, December 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/115/5/1203.

Jorgensen:1990:INP

[JAS^{+}90] A. O. Jorgensen, W. Arnold, A. C. Shen, S. H. Yuan, M. Gaver, and K. P. Campbell. Identification of novel proteins unique to either transverse tubules (TS28) or the sarcolemma (SL50) in rabbit skeletal muscle. Journal of Cell Biology, 110(4):1173–??, April 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/110/4/1173.

Jasmer:1993:TSI

[Jas93] D. P. Jasmer. Trichinella spiralis infected skeletal muscle cells arrest in G2/M and cease muscle gene expression. Journal of Cell Biology, 121(4):785–??, May 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/121/4/785.

Joshi:1993:NPM

[JB93] H. C. Joshi and P. W. Baas. A new perspective on microtubules and axon growth. Journal of Cell Biology, 121 (6):1191–??, June 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/121/6/1191.

Jarmolowski:1994:NED

[JBIM94] A. Jarmolowski, W. C. Boelens, E. Izaurralde, and I. W. Mattaj. Nuclear export of different classes of RNA is mediated by specific factors. Journal of Cell Biology, 124(5): 627–??, March 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/124/5/627.

Jongstra-Bilen:1992:LSP

[JBJH^{+}92] J. Jongstra-Bilen, P. A. Janmey, J. H. Hartwig, S. Galea, and J. Jongstra. The lymphocyte-specific protein LSP1 binds to F-actin and to the cytoskeleton through its COOHterminal basic domain. Journal of Cell Biology, 118(6): 1443–??, September 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/118/6/1443.

Jung:1994:EFI

[JCO^{+}94] W. Jung, E. Castren, M. Odenthal, G. F. Vande Woude, T. Ishii, H. P. Dienes, D. Lindholm, and P. Schirmacher. Expression and functional interaction of hepatocyte growth factor-scatter factor and its receptor c-met in mammalian brain. Journal of Cell Biology, 126(2):485–??, July 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/126/2/
485.

Jones:1993:CCP

[JCSH93] S. M. Jones, J. R. Crosby, J. Salamero, and K. E. Howell. A cytosolic complex of p62 and rab6 associates with TGN38/41 and is involved in budding of exocytic vesicles from the trans-Golgi network. Journal of Cell Biology, 122 (4):775–??, August 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/122/4/775.

Johnson:1991:MSL

[JD91] D. E. Johnson and S. K. Dutcher. Molecular studies of linkage group XIX of Chlamydomonas reinhardtii: evidence against a basal body location. Journal of Cell Biology, 113 (2):339–??, April 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/113/2/339.

Janmey:1991:VPV

[JETS91] P. A. Janmey, U. Euteneuer, P. Traub, and M. Schliwa. Viscoelastic properties of vimentin compared with other filamentous biopolymer networks. Journal of Cell Biology, 113 (1):155–??, April 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/113/1/155.

Jongen:1991:RCM

[JFA^{+}91] W. M. Jongen, D. J. Fitzgerald, M. Asamoto, C. Piccoli, T. J. Slaga, D. Gros, M. Takeichi, and H. Yamasaki. Regulation of connexin 43-mediated gap junctional intercellular communication by Ca^{2+} in mouse epidermal cells is controlled by E-cadherin. Journal of Cell Biology, 114(3): 545–??, August 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/114/3/545.

Jantsch:1992:ALU

[JG92] M. F. Jantsch and J. G. Gall. Assembly and localization of the U1-specific snRNP C protein in the amphibian oocyte. Journal of Cell Biology, 119(5):1037–??, December 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/119/5/
1037.

Jesaitis:1994:MCT

[JG94] L. A. Jesaitis and D. A. Goodenough. Molecular characterization and tissue distribution of ZO-2, a tight junction protein homologous to ZO-1 and the Drosophila discs-large tumor suppressor protein. Journal of Cell Biology, 124(6):

949–??, March 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/124/6/949.

Jones:1993:EMC

[JGB^{+}93] J. I. Jones, A. Gockerman, W. H. Busby, C. Camacho-Hubner, and D. R. Clemmons. Extracellular matrix contains insulin-like growth factor binding protein-5: potentiation of the effects of IGF–I. Journal of Cell Biology, 121 (3):679–??, May 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/121/3/679.

Jena:1994:RRL

[JGK^{+}94] B. P. Jena, F. D. Gumkowski, E. M. Konieczko, G. F. von Mollard, R. Jahn, and J. D. Jamieson. Redistribution of a rab3-like GTP-binding protein from secretory granules to the Golgi complex in pancreatic acinar cells during regulated exocytosis. Journal of Cell Biology, 124(1):43–??, January 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/124/1/43.

Jaffe:1993:OMS

[JGL^{+}93] L. A. Jaffe, C. J. Gallo, R. H. Lee, Y. K. Ho, and T. L. Jones. Oocyte maturation in starfish is mediated by the beta gamma-subunit complex of a G-protein. Journal of Cell Biology, 121(4):775–??, May 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/121/4/775.

Jiang:1990:PET

[JGLK90] W. P. Jiang, R. A. Gottlieb, W. J. Lennarz, and W. H. Kinsey. Phorbol ester treatment stimulates tyrosine phosphorylation of a sea urchin egg cortex protein. Journal of Cell Biology, 110(4):1049–??, April 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/110/4/1049.

Joshi:1991:PSD

[JGO^{+}91] R. Joshi, D. M. Gilligan, E. Otto, T. McLaughlin, and V. Bennett. Primary structure and domain organization of human alpha and beta adducin. Journal of Cell Biology, 115(3):665–??, November 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/115/3/665.

Jung:1990:GCD

[JH90] G. Jung and J. A. Hammer. Generation and characterization of Dictyostelium cells deficient in a myosin I heavy chain isoform. Journal of Cell Biology, 110(6):1955–??, June 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/110/6/1955.

Juliano:1993:STE

[JH93] R. L. Juliano and S. Haskill. Signal transduction from the extracellular matrix. Journal of Cell Biology, 120(3):577– ??, February 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/120/3/577.

Jaisser:1993:MUK

[JHGR93] F. Jaisser, J. D. Horisberger, K. Geering, and B. C. Rossier. Mechanisms of urinary K+ and H+ excretion: primary structure and functional expression of a novel H,K–ATPase. Journal of Cell Biology, 123(6):1421–??, December 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/123/6/
1421.

Jansen:1991:ECH

[JHK^{+}91] R. P. Jansen, E. C. Hurt, H. Kern, H. Lehtonen, M. Carmo-Fonseca, B. Lapeyre, and D. Tollervey. Evolutionary conservation of the human nucleolar protein fibrillarin and its functional expression in yeast. Journal of Cell Biology, 113 (4):715–??, May 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/113/4/715.

Jalkanen:1992:LCB

[JJ92] S. Jalkanen and M. Jalkanen. Lymphocyte CD44 binds the COOH-terminal heparin-binding domain of fibronectin. Journal of Cell Biology, 116(3):817–??, February 1992. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/116/3/

Jaunin:1993:RTE

[JJB^{+}93] P. Jaunin, F. Jaisser, A. T. Beggah, K. Takeyasu, P. Mangeat, B. C. Rossier, J. D. Horisberger, and K. Geering. Role of the transmembrane and extracytoplasmic domain of beta subunits in subunit assembly, intracellular transport, and functional expression of Na,K-pumps. Journal of Cell Biology, 123(6):1751–??, December 1993. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/123/6/
1751.

Juhnn:1992:ACT

[JJS92] Y. S. Juhnn, T. L. Jones, and A. M. Spiegel. Amino- and carboxy-terminal deletion mutants of Gs alpha are localized to the particulate fraction of transfected COS cells. Journal of Cell Biology, 119(3):523–??, November 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/119/3/523.

Johnson:1992:CTM

[JK92] K. F. Johnson and S. Kornfeld. The cytoplasmic tail of the mannose 6-phosphate/insulin-like growth factor–II receptor has two signals for lysosomal enzyme sorting in the Golgi. Journal of Cell Biology, 119(2):249–??, October 1992. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/119/2/
249.

Jantti:1993:ECR

[JK93] J. Jäntti and E. Kuismanen. Effect of caffeine and reduced temperature (20 degrees C) on the organization of the pre-Golgi and the Golgi stack membranes. Journal of Cell Biology, 120(6):1321–??, March 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/120/6/1321.

Jeckel:1992:GSC

[JKB^{+}92] D. Jeckel, A. Karrenbauer, K. N. Burger, G. van Meer, and F. Wieland. Glucosylceramide is synthesized at the cytosolic surface of various Golgi subfractions. Journal of Cell Biology, 117(2):259–??, April 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/117/2/259.

Jaworski:1994:BNM

[JKH94] D. M. Jaworski, G. M. Kelly, and S. Hockfield. BEHAB, a new member of the proteoglycan tandem repeat family of hyaluronan-binding proteins that is restricted to the brain. Journal of Cell Biology, 125(2):495–??, April 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/125/2/495.

Johnson:1994:SCC

[JKLG94] D. R. Johnson, L. J. Knoll, D. E. Levin, and J. I. Gordon. Saccharomyces cerevisiae contains four fatty acid activation (FAA) genes: an assessment of their role in regulating protein N-myristoylation and cellular lipid metabolism. Journal of Cell Biology, 127(3):751–??, November 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/127/3/751.

Jung:1993:EME

[JKS93] L. J. Jung, T. Kreiner, and R. H. Scheller. Expression of mutant ELH prohormones in AtT-20 cells: the relationship between prohormone processing and sorting. Journal of Cell Biology, 121(1):11–??, April 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/121/1/11.

Janson:1991:MCG

[JKT91] L. W. Janson, J. Kolega, and D. L. Taylor. Modulation of contraction by gelation/solation in a reconstituted motile model. Journal of Cell Biology, 114(5):1005–??, September 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/114/5/1005.

Jaconi:1990:CFC

[JLC^{+}90] M. E. Jaconi, D. P. Lew, J. L. Carpentier, K. E. Magnusson, M. Sjögren, and O. Stendahl. Cytosolic free calcium elevation mediates the phagosome-lysosome fusion during phagocytosis in human neutrophils. Journal of Cell Biology, 110(5):1555–??, May 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/110/5/1555.

Jiang:1993:ICG

[JLC93] W. Jiang, J. Lechner, and J. Carbon. Isolation and characterization of a gene (CBF2) specifying a protein component of the budding yeast kinetochore. Journal of Cell Biology, 121(3):513–??, May 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/121/3/513.

Johansson:1994:ALC

[JLL^{+}94] M. W. Johansson, E. Larsson, B. Lüning, E. B. Pasquale, and E. Ruoslahti. Altered localization and cytoplasmic domain-binding properties of tyrosine-phosphorylated beta 1 integrin. Journal of Cell Biology, 126(5):1299–??, September 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/126/5/1299.

Jalink:1992:TRA

[JM92] K. Jalink and W. H. Moolenaar. Thrombin receptor activation causes rapid neural cell rounding and neurite retraction independent of classic second messengers. Journal of Cell Biology, 118(2):411–??, July 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/118/2/411.

Jackson:1993:RTP

[JNP93] M. R. Jackson, T. Nilsson, and P. A. Peterson. Retrieval of transmembrane proteins to the endoplasmic reticulum. Journal of Cell Biology, 121(2):317–??, April 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/121/2/317.

James:1992:DRD

[JO92] G. James and E. Olson. Deletion of the regulatory domain of protein kinase C alpha exposes regions in the hinge and catalytic domains that mediate nuclear targeting. Journal of Cell Biology, 116(4):863–??, February 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/116/4/863.

Johnson:1990:MCC

[JP90] D. I. Johnson and J. R. Pringle. Molecular characterization of CDC42, a Saccharomyces cerevisiae gene involved in the development of cell polarity. Journal of Cell Biology, 111 (1):143–??, July 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/111/1/143.

James:1994:IRD

[JP94] D. E. James and R. C. Piper. Insulin resistance, diabetes, and the insulin-regulated trafficking of GLUT-4. Journal of Cell Biology, 126(5):1123–??, September 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/126/5/1123.

Jans:1991:VVR

[JPJF91] D. A. Jans, R. Peters, P. Jans, and F. Fahrenholz. Vasopressin V2-receptor mobile fraction and ligand-dependent adenylate cyclase activity are directly correlated in LLC– PK1 renal epithelial cells. Journal of Cell Biology, 114 (1):53–??, July 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/114/1/53.

Johnston:1991:SCM

[JPS91] G. C. Johnston, J. A. Prendergast, and R. A. Singer. The Saccharomyces cerevisiae MYO2 gene encodes an essential myosin for vectorial transport of vesicles. Journal of Cell Biology, 113(3):539–??, May 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/113/3/539.

Johnson:1992:PFA

[JR92] K. A. Johnson and J. L. Rosenbaum. Polarity of flagellar assembly in Chlamydomonas. Journal of Cell Biology, 119(6):1605–??, December 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/119/6/1605.

Jin:1990:EPC

[JRCW^{+}90] P. Jin, M. Rahm, L. Claesson-Welsh, C. H. Heldin, and T. Sejersen. Expression of PDGF A-chain and beta-receptor genes during rat myoblast differentiation. Journal of Cell Biology, 110(5):1665–??, May 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/110/5/1665.

Jarnagin:1994:EVF

[JRK^{+}94] W. R. Jarnagin, D. C. Rockey, V. E. Koteliansky, S. S. Wang, and D. M. Bissell. Expression of variant fibronectins in wound healing: cellular source and biological activity of the EIIIA segment in rat hepatic fibrogenesis. Journal of Cell Biology, 127(6):2037–??, December 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/127/6/2037.

Joyce:1990:TGF

[JRSB90] M. E. Joyce, A. B. Roberts, M. B. Sporn, and M. E. Bolander. Transforming growth factor-beta and the initiation of chondrogenesis and osteogenesis in the rat femur. Journal of Cell Biology, 110(6):2195–??, June 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/110/6/2195.

Jiang:1990:NTE

[JS90] L. W. Jiang and M. Schindler. Nucleocytoplasmic transport is enhanced concomitant with nuclear accumulation of epidermal growth factor (EGF) binding activity in both 3T3-1 and EGF receptor reconstituted NR-6 fibroblasts. Journal of Cell Biology, 110(3):559–??, March 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/110/3/559.

Johnston:1992:KPR

[JS92] J. A. Johnston and R. D. Sloboda. A 62-kD protein required for mitotic progression is associated with the mitotic apparatus during M-phase and with the nucleus during interphase. Journal of Cell Biology, 119(4):843–??, November 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/119/4/843.

Jorgensen:1993:CRC

[JSA^{+}93] A. O. Jorgensen, A. C. Shen, W. Arnold, P. S. McPherson, and K. P. Campbell. The Ca^{2+}-release channel/ryanodine receptor is localized in junctional and corbular sarcoplasmic reticulum in cardiac muscle. Journal of Cell Biology, 120(4): 969–??, February 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/120/4/969.

Johnson:1992:AAA

[JSG^{+}92] G. R. Johnson, T. Saeki, A. W. Gordon, M. Shoyab, D. S. Salomon, and K. Stromberg. Autocrine action of amphiregulin in a colon carcinoma cell line and immunocytochemical localization of amphiregulin in human colon. Journal of Cell Biology, 118(3):741–??, August 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/118/3/741.

Jing:1990:RHT

[JSM^{+}90] S. Q. Jing, T. Spencer, K. Miller, C. Hopkins, and I. S. Trowbridge. Role of the human transferrin receptor cytoplasmic domain in endocytosis: localization of a specific signal sequence for internalization. Journal of Cell Biology, 110(2):283–??, February 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/110/2/283.

Janson:1993:VMT

[JT93] L. W. Janson and D. L. Taylor. In vitro models of tail contraction and cytoplasmic streaming in amoeboid cells. Journal of Cell Biology, 123(2):345–??, October 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/123/2/345.

Jaconi:1991:MEC

[JTS^{+}91] M. E. Jaconi, J. M. Theler, W. Schlegel, R. D. Appel, S. D. Wright, and P. D. Lew. Multiple elevations of cytosolic-free Ca^{2+} in human neutrophils: initiation by adherence receptors of the integrin family. Journal of Cell Biology, 112(6): 1249–??, March 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/112/6/1249.

Jalink:1994:ILT

[JvCH^{+}94] K. Jalink, E. J. van Corven, T. Hengeveld, N. Morii, S. Narumiya, and W. H. Moolenaar. Inhibition of lysophosphatidate- and thrombin-induced neurite retraction and neuronal cell rounding by ADP ribosylation of the small GTP-binding protein Rho. Journal of Cell Biology, 126(3):801–??, August 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/126/3/801.

Johansson:1993:FPR

[JWHPM93] B. Johansson, M. P. Wymann, K. Holmgren-Peterson, and K. E. Magnusson. N-formyl peptide receptors in human neutrophils display distinct membrane distribution and lateral mobility when labeled with agonist and antagonist. Journal of Cell Biology, 121(6):1281–??, June 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/121/6/1281.

Johnson:1993:LED

[JZAVP93] R. C. Johnson, D. Zhu, H. G. Augustin-Voss, and B. U. Pauli. Lung endothelial dipeptidyl peptidase IV is an adhesion molecule for lung-metastatic rat breast and prostate carcinoma cells. Journal of Cell Biology, 121(6):1423–??, June 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/121/6/1423.

Key:1990:ONE

[KA90] B. Key and R. A. Akeson. Olfactory neurons express a unique glycosylated form of the neural cell adhesion molecule (N-CAM). Journal of Cell Biology, 110(5):1729– ??, May 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/110/5/1729.

Kawamoto:1991:CNM

[KA91a] S. Kawamoto and R. S. Adelstein. Chicken nonmuscle myosin heavy chains: differential expression of two mR-NAs and evidence for two different polypeptides. Journal of Cell Biology, 112(5):915–??, March 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/112/5/915.

Kobayashi:1991:TEF

[KA91b] T. Kobayashi and Y. Arakawa. Transport of exogenous fluorescent phosphatidylserine analogue to the Golgi apparatus in cultured fibroblasts. Journal of Cell Biology, 113 (2):235–??, April 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/113/2/235.

Kuliawat:1992:PTC

[KA92] R. Kuliawat and P. Arvan. Protein targeting via the “constitutive-like” secretory pathway in isolated pancreatic islets: passive sorting in the immature granule compartment. Journal of Cell Biology, 118(3):521–??, August 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/118/3/
521.

Kuang:1993:LTK

[KA93] J. Kuang and C. L. Ashorn. At least two kinases phosphorylate the MPM-2 epitope during Xenopus oocyte maturation. Journal of Cell Biology, 123(4):859–??, November 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/123/4/
859.

Kuliawat:1994:DMM

[KA94] R. Kuliawat and P. Arvan. Distinct molecular mechanisms for protein sorting within immature secretory granules of pancreatic beta-cells. Journal of Cell Biology, 126(1):77– ??, July 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/126/1/77.

Klemenz:1993:EMS

[KAF^{+}93] R. Klemenz, A. C. Andres, E. Fröhli, R. Schäfer, and A. Aoyama. Expression of the murine small heat shock proteins hsp 25 and alpha B crystallin in the absence of stress. Journal of Cell Biology, 120(3):639–??, February 1993. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/120/3/
639.

Kennedy:1994:DCS

[KAG94] B. K. Kennedy, N. R. Austriaco, and L. Guarente. Daughter cells of Saccharomyces cerevisiae from old mothers display a reduced life span. Journal of Cell Biology, 127(6):1985– ??, December 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/127/6/1985.

Karpati:1993:PVA

[KAP^{+}93] S. Kárpáti, M. Amagai, R. Prussick, K. Cehrs, and J. R. Stanley. Pemphigus vulgaris antigen, a desmoglein type of cadherin, is localized within keratinocyte desmosomes. Journal of Cell Biology, 122(2):409–??, July 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/122/2/409.

Kao:1990:AIC

[KATS90] J. P. Kao, J. M. Alderton, R. Y. Tsien, and R. A. Steinhardt. Active involvement of Ca^{2+} in mitotic progression of Swiss 3T3 fibroblasts. Journal of Cell Biology, 111(1): 183–??, July 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/111/1/183.

Kao:1993:DTM

[KAV93] H. P. Kao, J. R. Abney, and A. S. Verkman. Determinants of the translational mobility of a small solute in cell cytoplasm. Journal of Cell Biology, 120(1):175–??, January 1993. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/120/1/
175.

Kochanski:1990:MCD

[KB90] R. S. Kochanski and G. G. Borisy. Mode of centriole duplication and distribution. Journal of Cell Biology, 110(5): 1599–??, May 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/110/5/1599.

Kerwin:1991:AAS

[KB91a] B. Kerwin and E. Bandman. Assembly of avian skeletal muscle myosins: evidence that homodimers of the heavy chain subunit are the thermodynamically stable form. Journal of Cell Biology, 113(2):311–??, April 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/113/2/311.

Kordeli:1991:DAI

[KB91b] E. Kordeli and V. Bennett. Distinct ankyrin isoforms at neuron cell bodies and nodes of Ranvier resolved using erythrocyte ankyrin-deficient mice. Journal of Cell Biology, 114(6):1243–??, September 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/114/6/1243.

Kim:1992:TAN

[KBA92] P. S. Kim, D. Bole, and P. Arvan. Transient aggregation of nascent thyroglobulin in the endoplasmic reticulum: relationship to the molecular chaperone, BiP. Journal of Cell Biology, 118(3):541–??, August 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/118/3/541.

Kellermann:1990:IOC

[KBCM^{+}90] O. Kellermann, M. H. Buc-Caron, P. J. Marie, D. Lamblin, and F. Jacob. An immortalized osteogenic cell line derived from mouse teratocarcinoma is able to mineralize in vivo and in vitro. Journal of Cell Biology, 110(1):123–??, January 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/110/1/123.

Kellaris:1991:TIA

[KBG91] K. V. Kellaris, S. Bowen, and R. Gilmore. ER translocation intermediates are adjacent to a nonglycosylated 34-kD integral membrane protein. Journal of Cell Biology, 114 (1):21–??, July 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/114/1/21.

Kok:1991:SSE

[KBH91] J. W. Kok, T. Babia, and D. Hoekstra. Sorting of sphingolipids in the endocytic pathway of HT29 cells. Journal of Cell Biology, 114(2):231–??, July 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/114/2/231.

Kastner:1991:EMU

[KBL91] B. Kastner, M. Bach, and R. Lührmann. Electron microscopy of U4/U6 snRNP reveals a Y-shaped U4 and U6 RNA containing domain protruding from the U4 core RNP. Journal of Cell Biology, 112(6):1065–??, March 1991. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/112/6/
1065.

Kurihara:1994:NCM

[KBL^{+}94] L. J. Kurihara, C. T. Beh, M. Latterich, R. Schekman, and M. D. Rose. Nuclear congression and membrane fusion: two distinct events in the yeast karyogamy pathway. Journal of Cell Biology, 126(4):911–??, August 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/126/4/911.

Klein:1993:TCE

[KBM93] G. Klein, S. Beck, and C. A. Müller. Tenascin is a cytoadhesive extracellular matrix component of the human hematopoietic microenvironment. Journal of Cell Biology, 123(4):1027–??, November 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/123/4/1027.

Knoll:1991:QFA

[KBP91] G. Knoll, C. Braun, and H. Plattner. Quenched flow analysis of exocytosis in Paramecium cells: time course, changes in membrane structure, and calcium requirements revealed after rapid mixing and rapid freezing of intact cells. Journal of Cell Biology, 113(6):1295–??, June 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/113/6/1295.

Konig:1992:TGF

[KBT92] A. König and L. Bruckner-Tuderman. Transforming growth factor-beta stimulates collagen VII expression by cutaneous cells in vitro. Journal of Cell Biology, 117(3):679–??, May 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/117/3/679.

Kerboeuf:1990:CIA

[KC90] D. Kerboeuf and J. Cohen. A Ca^{2+} influx associated with exocytosis is specifically abolished in a Paramecium exocytotic mutant. Journal of Cell Biology, 111(6):2527–??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/111/6/2527.

Koedam:1992:PSG

[KCB^{+}92] J. A. Koedam, E. M. Cramer, E. Briend, B. Furie, B. C. Furie, and D. D. Wagner. P-selectin, a granule membrane protein of platelets and endothelial cells, follows the regulated secretory pathway in AtT-20 cells. Journal of Cell Biology, 116(3):617–??, February 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/116/3/617.

Kramer:1990:HME

[KCC90] R. H. Kramer, Y. F. Cheng, and R. Clyman. Human microvascular endothelial cells use beta 1 and beta 3 integrin receptor complexes to attach to laminin. Journal of Cell Biology, 111(3):1233–??, September 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/111/3/1233.

Kwan:1991:MOC

[KCCG91] A. P. Kwan, C. E. Cummings, J. A. Chapman, and M. E. Grant. Macromolecular organization of chicken type X collagen in vitro. Journal of Cell Biology, 114(3):597–??, August 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/114/3/597.

Klaus:1991:DPS

[KCDR91] S. Klaus, A. M. Cassard-Doulcier, and D. Ricquier. Development of Phodopus sungorus brown preadipocytes in primary cell culture: effect of an atypical beta-adrenergic agonist, insulin, and triiodothyronine on differentiation, mitochondrial development, and expression of the uncoupling protein UCP. Journal of Cell Biology, 115(6):1783–??, December 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/115/6/1783.

Kadowaki:1994:ICS

[KCH^{+}94] T. Kadowaki, S. Chen, M. Hitomi, E. Jacobs, C. Kumagai, S. Liang, R. Schneiter, D. Singleton, J. Wisniewska, and A. M. Tartakoff. Isolation and characterization of Saccharomyces cerevisiae mRNA transport-defective (mtr) mutants. Journal of Cell Biology, 126(3):649–??, August 1994.

CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/126/3/
649.

Krishna:1993:FCA

[KCM^{+}93] S. Krishna, G. Cowan, J. C. Meade, R. A. Wells, J. R. Stringer, and K. J. Robson. A family of cation ATPase-like molecules from Plasmodium falciparum. Journal of Cell Biology, 120(2):385–??, January 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/120/2/385.

Koji:1994:PDE

[KCR^{+}94] T. Koji, M. Chedid, J. S. Rubin, O. D. Slayden, K. G. Csaky, S. A. Aaronson, and R. M. Brenner. Progesteronedependent expression of keratinocyte growth factor mRNA in stromal cells of the primate endometrium: keratinocyte growth factor as a progestomedin. Journal of Cell Biology, 125(2):393–??, April 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/125/2/393.

Klionsky:1992:ASC

[KCY92] D. J. Klionsky, R. Cueva, and D. S. Yaver. Aminopeptidase I of Saccharomyces cerevisiae is localized to the vacuole independent of the secretory pathway. Journal of Cell Biology, 119(2):287–??, October 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/119/2/287.

Kovach:1992:MAB

[KCYH92] N. L. Kovach, T. M. Carlos, E. Yee, and J. M. Harlan. A monoclonal antibody to beta 1 integrin (CD29) stimulates VLA-dependent adherence of leukocytes to human umbilical vein endothelial cells and matrix components. Journal of Cell Biology, 116(2):499–??, January 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/116/2/499.

Kao:1992:RSP

[KD92] C. Y. Kao and R. K. Draper. Retention of secretory proteins in an intermediate compartment and disappearance of the Golgi complex in an END4 mutant of Chinese hamster ovary cells. Journal of Cell Biology, 117(4):701–??, May 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/117/4/
701.

Kilmartin:1993:SPS

[KDKF93] J. V. Kilmartin, S. L. Dyos, D. Kershaw, and J. T. Finch. A spacer protein in the Saccharomyces cerevisiae spindle poly body whose transcript is cell cycle-regulated. Journal of Cell Biology, 123(5):1175–??, December 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/123/5/1175.

Klausner:1992:BIC

[KDLS92] R. D. Klausner, J. G. Donaldson, and J. Lippincott-Schwartz. Brefeldin A: insights into the control of membrane traffic and organelle structure. Journal of Cell Biology, 116(5):1071–??, March 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/116/5/1071.

Kornbluth:1994:EDR

[KDN94] S. Kornbluth, M. Dasso, and J. Newport. Evidence for a dual role for TC4 protein in regulating nuclear structure and cell cycle progression. Journal of Cell Biology, 125 (4):705–??, May 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/125/4/705.

Kuge:1994:SPV

[KDO^{+}94] O. Kuge, C. Dascher, L. Orci, T. Rowe, M. Amherdt, H. Plutner, M. Ravazzola, G. Tanigawa, J. E. Rothman, and W. E. Balch. Sar1 promotes vesicle budding from the endoplasmic reticulum but not Golgi compartments. Journal of Cell Biology, 125(1):51–??, April 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/125/1/51.

Klotz:1990:PXE

[KDP^{+}90] C. Klotz, M. C. Dabauvalle, M. Paintrand, T. Weber, M. Bornens, and E. Karsenti. Parthenogenesis in Xenopus eggs requires centrosomal integrity. Journal of Cell Biology, 110(2):405–??, February 1990. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/110/2/405.

Kurzchalia:1992:VKM

[KDP^{+}92] T. V. Kurzchalia, P. Dupree, R. G. Parton, R. Kellner, H. Virta, M. Lehnert, and K. Simons. VIP21, a 21-kD membrane protein is an integral component of trans-Golgi-network-derived transport vesicles. Journal of Cell Biology, 118(5):1003–??, September 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/118/5/1003.

Klein:1991:IAB

[KDS^{+}91] C. E. Klein, D. Dressel, T. Steinmayer, C. Mauch, B. Eckes, T. Krieg, R. B. Bankert, and L. Weber. Integrin alpha 2 beta 1 is upregulated in fibroblasts and highly aggressive melanoma cells in three-dimensional collagen lattices and mediates the reorganization of collagen I fibrils. Journal of Cell Biology, 115(5):1427–??, December 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/115/5/1427.

Kremer:1991:STN

[KDT^{+}91] N. E. Kremer, G. D’Arcangelo, S. M. Thomas, M. DeMarco, J. S. Brugge, and S. Halegoua. Signal transduction by nerve growth factor and fibroblast growth factor in PC12 cells requires a sequence of src and ras actions. Journal of Cell Biology, 115(3):809–??, November 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/115/3/809.

Kuchler:1993:FTS

[KDT93] K. Kuchler, H. G. Dohlman, and J. Thorner. The a-factor transporter (STE6 gene product) and cell polarity in the yeast Saccharomyces cerevisiae. Journal of Cell Biology, 120(5):1203–??, March 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/120/5/1203.

Kordeli:1990:IAL

[KDTB90] E. Kordeli, J. Davis, B. Trapp, and V. Bennett. An isoform of ankyrin is localized at nodes of Ranvier in myelinated axons of central and peripheral nerves. Journal of Cell Biology, 110(4):1341–??, April 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/110/4/1341.

Kristensen:1991:TAS

[KEBD91] P. Kristensen, J. Eriksen, F. Blasi, and K. Danø. Two alternatively spliced mouse urokinase receptor mRNAs with different histological localization in the gastrointestinal tract. Journal of Cell Biology, 115(6):1763–??, December 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/115/6/
1763.

Kucik:1990:CMD

[KES90] D. F. Kucik, E. L. Elson, and M. P. Sheetz. Cell migration does not produce membrane flow. Journal of Cell Biology, 111(4):1617–??, October 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/111/4/1617.

Kadmon:1992:NHS

[KES^{+}92] G. Kadmon, M. Eckert, M. Sammar, M. Schachner, and P. Altevogt. Nectadrin, the heat-stable antigen, is a cell adhesion molecule. Journal of Cell Biology, 118(5):1245– ??, September 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/118/5/1245.

Kim:1994:CCM

[KFC^{+}94] Y. J. Kim, L. Francisco, G. C. Chen, E. Marcotte, and C. S. Chan. Control of cellular morphogenesis by the Ip12/Bem2 GTPase-activating protein: possible role of protein phosphorylation. Journal of Cell Biology, 127(5):1381–??, December 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/127/5/1381.

Kean:1993:RLT

[KFN93] L. S. Kean, R. S. Fuller, and J. W. Nichols. Retrograde lipid traffic in yeast: identification of two distinct pathways for internalization of fluorescent-labeled phosphatidylcholine from the plasma membrane. Journal of Cell Biology, 123(6):1403–??, December 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/123/6/1403.

Kieffer:1991:APB

[KFW^{+}91] N. Kieffer, L. A. Fitzgerald, D. Wolf, D. A. Cheresh, and D. R. Phillips. Adhesive properties of the beta 3 integrins: comparison of GP IIb–IIIa and the vitronectin receptor individually expressed in human melanoma cells. Journal of Cell Biology, 113(2):451–??, April 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/113/2/451.

Kioussi:1994:DIP

[KG94] C. Kioussi and P. Gruss. Differential induction of Pax genes by NGF and BDNF in cerebellar primary cultures. Journal of Cell Biology, 125(2):417–??, April 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/125/2/417.

Kodukula:1993:BGG

[KGA^{+}93] K. Kodukula, L. D. Gerber, R. Amthauer, L. Brink, and S. Udenfriend. Biosynthesis of glycosylphosphatidylinositol (GPI)-anchored membrane proteins in intact cells: specific amino acid requirements adjacent to the site of cleavage and GPI attachment. Journal of Cell Biology, 120(3): 657–??, February 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/120/3/657.

Kistler:1994:RNT

[KGDE94] J. Kistler, K. Goldie, P. Donaldson, and A. Engel. Reconstitution of native-type noncrystalline lens fiber gap junctions from isolated hemichannels. Journal of Cell Biology, 126(4): 1047–??, August 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/126/4/1047.

Klein:1990:CSA

[KGK90] U. Klein, M. Gernold, and P. M. Kloetzel. Cell-specific accumulation of Drosophila proteasomes (MCP) during early development. Journal of Cell Biology, 111(6):2275–??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/111/6/2275.

Koonce:1992:DDP

[KGM92] M. P. Koonce, P. M. Grissom, and J. R. McIntosh. Dynein from Dictyostelium: primary structure comparisons between a cytoplasmic motor enzyme and flagellar dynein. Journal of Cell Biology, 119(6):1597–??, December 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/119/6/
1597.

Kalies:1994:BRR

[KGR94] K. U. Kalies, D. Görlich, and T. A. Rapoport. Binding of ribosomes to the rough endoplasmic reticulum mediated by the Sec61p-complex. Journal of Cell Biology, 126(4): 925–??, August 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/126/4/925.

Kelly:1992:SCM

[KGSB92] B. M. Kelly, C. S. Gillespie, D. L. Sherman, and P. J. Brophy. Schwann cells of the myelin-forming phenotype express neurofilament protein NF–M. Journal of Cell Biology, 118 (2):397–??, July 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/118/2/397.

Koseki:1992:AMD

[KHaA92] C. Koseki, D. Herzlinger, and Q. al Awqati. Apoptosis in metanephric development. Journal of Cell Biology, 119(5): 1327–??, December 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/119/5/1327.

Kedersha:1991:VIV

[KHCR91] N. L. Kedersha, J. E. Heuser, D. C. Chugani, and L. H. Rome. Vaults. III. Vault ribonucleoprotein particles open into flower-like structures with octagonal symmetry. Journal of Cell Biology, 112(2):225–??, January 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/112/2/225.

Kouklis:1994:MCD

[KHF94] P. D. Kouklis, E. Hutton, and E. Fuchs. Making a connection: direct binding between keratin intermediate filaments and desmosomal proteins. Journal of Cell Biology, 127(4):1049–??, November 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/127/4/1049.

Kuivinen:1993:SRC

[KHHC93] E. Kuivinen, B. L. Hoffman, P. A. Hoffman, and C. R. Carlin. Structurally related class I and class II receptor protein tyrosine kinases are down-regulated by the same E3 protein coded for by human group C adenoviruses. Journal of Cell Biology, 120(5):1271–??, March 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/120/5/1271.

Kuijpers:1992:NMA

[KHHR92] T. W. Kuijpers, B. C. Hakkert, M. H. Hart, and D. Roos. Neutrophil migration across monolayers of cytokine-prestimulated endothelial cells: a role for plateletactivating factor and IL-8. Journal of Cell Biology, 117 (3):565–??, May 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/117/3/565.

Kuge:1993:ZCS

[KHKO^{+}93] O. Kuge, S. Hara-Kuge, L. Orci, M. Ravazzola, M. Amherdt, G. Tanigawa, F. T. Wieland, and J. E. Rothman. zeta-COP, a subunit of coatomer, is required for COP-coated vesicle assembly. Journal of Cell Biology, 123(6):1727–??, December 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/123/6/1727.

Kim:1991:CMS

[KHP91a] H. B. Kim, B. K. Haarer, and J. R. Pringle. Cellular morphogenesis in the Saccharomyces cerevisiae cell cycle: localization of the CDC3 gene product and the timing of events at the budding site. Journal of Cell Biology, 112(4):535– ??, February 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/112/4/535.

Kurz:1991:MCH

[KHP^{+}91b] E. M. Kurz, T. W. Holstein, B. M. Petri, J. Engel, and C. N. David. Mini-collagens in hydra nematocytes. Journal of Cell Biology, 115(4):1159–??, November 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/115/4/1159.

Kojima:1991:LIG

[KHR91] S. Kojima, P. C. Harpel, and D. B. Rifkin. Lipoprotein (a) inhibits the generation of transforming growth factor beta: an endogenous inhibitor of smooth muscle cell migration. Journal of Cell Biology, 113(6):1439–??, June 1991. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/113/6/
1439.

Kadomatsu:1990:RAR

[KHS^{+}90] K. Kadomatsu, R. P. Huang, T. Suganuma, F. Murata, and T. Muramatsu. A retinoic acid responsive gene MK found in the teratocarcinoma system is expressed in spatially and temporally controlled manner during mouse embryogenesis. Journal of Cell Biology, 110(3):607–??, March 1990. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/110/3/
607.

Kidd:1992:MSS

[KHTD92] G. J. Kidd, J. W. Heath, B. D. Trapp, and P. R. Dunkley. Myelin sheath survival after guanethidine-induced axonal degeneration. Journal of Cell Biology, 116(2):395–??, January 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/116/2/395.

Klumperman:1993:DED

[KHV^{+}93] J. Klumperman, A. Hille, T. Veenendaal, V. Oorschot, W. Stoorvogel, K. von Figura, and H. J. Geuze. Differences in the endosomal distributions of the two mannose 6-phosphate receptors. Journal of Cell Biology, 121(5):997– ??, June 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/121/5/997.

Kuijpers:1992:CNC

[KHvdL^{+}92] T. W. Kuijpers, M. Hoogerwerf, L. J. van der Laan, G. Nagel, C. E. van der Schoot, F. Grunert, and D. Roos. CD66 nonspecific cross-reacting antigens are involved in neutrophil adherence to cytokine-activated endothelial cells. Journal of Cell Biology, 118(2):457–??, July 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/118/2/457.

Kemble:1993:GTA

[KHW93] G. W. Kemble, Y. I. Henis, and J. M. White. GPI- and transmembrane-anchored influenza hemagglutinin differ in structure and receptor binding activity. Journal of Cell Biology, 122(6):1253–??, September 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/122/6/1253.

Kinosita:1991:DVM

[KII^{+}91] K. Kinosita, H. Itoh, S. Ishiwata, K. Hirano, T. Nishizaka, and T. Hayakawa. Dual-view microscopy with a single camera: real-time imaging of molecular orientations and calcium. Journal of Cell Biology, 115(1):67–??, October 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/115/1/
67.

Kikkawa:1994:DVM

[KIN^{+}94] M. Kikkawa, T. Ishikawa, T. Nakata, T. Wakabayashi, and N. Hirokawa. Direct visualization of the microtubule lattice seam both in vitro and in vivo. Journal of Cell Biology, 127(6):1965–??, December 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/127/6/1965.

King:1994:BMR

[KIO^{+}94] S. J. King, W. B. Inwood, E. T. O’Toole, J. Power, and S. K. Dutcher. The bop2-1 mutation reveals radial asymmetry in the inner dynein arm region of Chlamydomonas reinhardtii. Journal of Cell Biology, 126(5):1255–??, September 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/126/5/

Kahn:1994:MPC

[KIS^{+}94] J. Kahn, R. H. Ingraham, F. Shirley, G. I. Migaki, and T. K. Kishimoto. Membrane proximal cleavage of L-selectin: identification of the cleavage site and a 6-kD transmembrane peptide fragment of L-selectin. Journal of Cell Biology, 125 (2):461–??, April 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/125/2/461.

Kanaseki:1991:SFC

[KISY91] T. Kanaseki, Y. Ikeuchi, H. Sugiura, and T. Yamauchi. Structural features of Ca^{2+}/calmodulin-dependent protein kinase II revealed by electron microscopy. Journal of Cell Biology, 115(4):1049–??, November 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/115/4/1049.

Kuwayama:1993:NCD

[KIV93] H. Kuwayama, S. Ishida, and P. J. Van Haastert. Nonchemotactic Dictyostelium discoideum mutants with altered cGMP signal transduction. Journal of Cell Biology, 123(6): 1453–??, December 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/123/6/1453.

Kuhtreiber:1990:DEC

[KJ90] W. M. Kühtreiber and L. F. Jaffe. Detection of extracellular calcium gradients with a calcium-specific vibrating electrode. Journal of Cell Biology, 110(5):1565–??, May 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/110/5/
1565.

Kolega:1991:RSC

[KJT91] J. Kolega, L. W. Janson, and D. L. Taylor. The role of solation-contraction coupling in regulating stress fiber dynamics in nonmuscle cells. Journal of Cell Biology, 114(5): 993–??, September 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/114/5/993.

Kamimura:1992:HFV

[KK92] S. Kamimura and R. Kamiya. High-frequency vibration in flagellar axonemes with amplitudes reflecting the size of tubulin. Journal of Cell Biology, 116(6):1443–??, March 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/116/6/1443.

Kadmon:1990:FCB

[KKAS90a] G. Kadmon, A. Kowitz, P. Altevogt, and M. Schachner. Functional cooperation between the neural adhesion molecules L1 and N-CAM is carbohydrate dependent. Journal of Cell Biology, 110(1):209–??, January 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/110/1/209.

Kadmon:1990:NCA

[KKAS90b] G. Kadmon, A. Kowitz, P. Altevogt, and M. Schachner. The neural cell adhesion molecule N-CAM enhances L1dependent cell–cell interactions. Journal of Cell Biology, 110(1):193–??, January 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/110/1/193.

Kuhl:1992:LTS

[KKBK92] D. Kuhl, T. E. Kennedy, A. Barzilai, and E. R. Kandel. Long-term sensitization training in Aplysia leads to an increase in the expression of BiP, the major protein chaperon of the ER. Journal of Cell Biology, 119(5):1069–??, December 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/119/5/1069.

Kastner:1992:SSN

[KKBL92] B. Kastner, U. Kornstädt, M. Bach, and R. Lührmann. Structure of the small nuclear RNP particle U1: identification of the two structural protuberances with RNP-antigens A and 70K. Journal of Cell Biology, 116(4):839–??, February 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/116/4/839.

Kucik:1991:PAM

[KKES91] D. F. Kucik, S. C. Kuo, E. L. Elson, and M. P. Sheetz. Preferential attachment of membrane glycoproteins to the cytoskeleton at the leading edge of lamella. Journal of Cell Biology, 114(5):1029–??, September 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/114/5/1029.

Kraemer:1990:TNP

[KKFD90] D. Kraemer, R. Koob, B. Friedrichs, and D. Drenckhahn. Two novel peripheral membrane proteins, pasin 1 and pasin 2, associated with Na+,K(+)-ATPase in various cells and tissues. Journal of Cell Biology, 111(6):2375–??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/111/6/2375.

Kopitz:1990:NAC

[KKG^{+}90] J. Kopitz, G. O. Kisen, P. B. Gordon, P. Bohley, and P. O. Seglen. Nonselective autophagy of cytosolic enzymes by isolated rat hepatocytes. Journal of Cell Biology, 111(3):941– ??, September 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/111/3/941.

Keller:1991:ECM

[KKG^{+}91] G. A. Keller, S. Krisans, S. J. Gould, J. M. Sommer, C. C. Wang, W. Schliebs, W. Kunau, S. Brody, and S. Subramani. Evolutionary conservation of a microbody targeting signal that targets proteins to peroxisomes, glyoxysomes, and glycosomes. Journal of Cell Biology, 114(5):893–??, September 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/114/5/893.

Knops:1991:OTN

[KKL^{+}91] J. Knops, K. S. Kosik, G. Lee, J. D. Pardee, L. Cohen-Gould, and L. McConlogue. Overexpression of tau in a nonneuronal cell induces long cellular processes. Journal of Cell Biology, 114(4):725–??, August 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/114/4/725.

Kamiya:1991:TTC

[KKM91] R. Kamiya, E. Kurimoto, and E. Muto. Two types of Chlamydomonas flagellar mutants missing different components of inner-arm dynein. Journal of Cell Biology, 112(3): 441–??, February 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/112/3/441.

Kren:1993:DRM

[KKW^{+}93] B. T. Kren, N. M. Kumar, S. Q. Wang, N. B. Gilula, and C. J. Steer. Differential regulation of multiple gap junction transcripts and proteins during rat liver regeneration. Journal of Cell Biology, 123(3):707–??, November 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/123/3/707.

Kabakoff:1990:IGP

[KL90] B. Kabakoff and W. J. Lennarz. Inhibition of glycoprotein processing blocks assembly of spicules during development of the sea urchin embryo. Journal of Cell Biology, 111(2): 391–??, August 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/111/2/391.

Kulesza-Lipka:1991:IMH

[KLBBK91] D. Kulesza-Lipka, I. C. Baines, H. Brzeska, and E. D. Korn. Immunolocalization of myosin I heavy chain kinase in Acanthamoeba castellanii and binding of purified kinase to isolated plasma membranes. Journal of Cell Biology, 115(1): 109–??, October 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/115/1/109.

Klein:1992:LRT

[KLC^{+}92] F. Klein, T. Laroche, M. E. Cardenas, J. F. Hofmann, D. Schweizer, and S. M. Gasser. Localization of RAP1 and topoisomerase II in nuclei and meiotic chromosomes of yeast. Journal of Cell Biology, 117(5):935–??, June 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/117/5/

Krijnse-Locker:1994:CBC

[KLERG94] J. Krijnse-Locker, M. Ericsson, P. J. Rottier, and G. Griffiths. Characterization of the budding compartment of mouse hepatitis virus: evidence that transport from the RER to the Golgi complex requires only one vesicular transport step. Journal of Cell Biology, 124(1):55–??, January 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/124/1/55.

Keene:1991:TTX

[KLM^{+}91] D. R. Keene, G. P. Lunstrum, N. P. Morris, D. W. Stoddard, and R. E. Burgeson. Two type XII-like collagens localize to the surface of banded collagen fibrils. Journal of Cell Biology, 113(4):971–??, May 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/113/4/971.

Kragler:1993:TIP

[KLR^{+}93] F. Kragler, A. Langeder, J. Raupachova, M. Binder, and A. Hartig. Two independent peroxisomal targeting signals in catalase A of Saccharomyces cerevisiae. Journal of Cell Biology, 120(3):665–??, February 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/120/3/665.

Kislauskis:1993:ISU

[KLST93] E. H. Kislauskis, Z. Li, R. H. Singer, and K. L. Taneja. Isoform-specific 3’-untranslated sequences sort alpha-cardiac and beta-cytoplasmic actin messenger RNAs to different cytoplasmic compartments. Journal of Cell Biology, 123(1):165–??, October 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/123/1/165.

Kambach:1992:IDU

[KM92a] C. Kambach and I. W. Mattaj. Intracellular distribution of the U1A protein depends on active transport and nuclear binding to U1 snRNA. Journal of Cell Biology, 118 (1):11–??, July 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/118/1/11.

Kamimura:1992:TPK

[KM92b] S. Kamimura and E. Mandelkow. Tubulin protofilaments and kinesin-dependent motility. Journal of Cell Biology, 118 (4):865–??, August 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/118/4/865.

Kellerman:1992:UMH

[KM92c] K. A. Kellerman and K. G. Miller. An unconventional myosin heavy chain gene from Drosophila melanogaster. Journal of Cell Biology, 119(4):823–??, November 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/119/4/
823.

Kopczynski:1992:IED

[KM92d] C. C. Kopczynski and M. A. Muskavitch. Introns excised from the Delta primary transcript are localized near sites of Delta transcription. Journal of Cell Biology, 119(3):503– ??, November 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/119/3/503.

Kotz:1994:ICC

[KM94] K. J. Kotz and M. A. McNiven. Intracellular calcium and cAMP regulate directional pigment movements in teleost erythrophores. Journal of Cell Biology, 124(4):463–??, February 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/124/4/463.

Kedersha:1990:VIR

[KMBR90] N. L. Kedersha, M. C. Miquel, D. Bittner, and L. H. Rome. Vaults. II. Ribonucleoprotein structures are highly conserved among higher and lower eukaryotes. Journal of Cell Biology, 110(4):895–??, April 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/110/4/895.

Kobayashi:1991:SDB

[KMF^{+}91] H. Kobayashi, J. Minshull, C. Ford, R. Golsteyn, R. Poon, and T. Hunt. On the synthesis and destruction of A- and B-type cyclins during oogenesis and meiotic maturation in Xenopus laevis. Journal of Cell Biology, 114(4):755–??, August 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/114/4/755.

Klymkowsky:1991:CPC

[KMN91] M. W. Klymkowsky, L. A. Maynell, and C. Nislow. Cytokeratin phosphorylation, cytokeratin filament severing and the solubilization of the maternal mRNA Vg1. Journal of Cell Biology, 114(4):787–??, August 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/114/4/787.

Krek:1992:CKI

[KMN92] W. Krek, G. Maridor, and E. A. Nigg. Casein kinase II is a predominantly nuclear enzyme. Journal of Cell Biology, 116(1):43–??, January 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/116/1/43.

Kinloch:1991:ECC

[KMSW91] R. A. Kinloch, S. Mortillo, C. L. Stewart, and P. M. Wassarman. Embryonal carcinoma cells transfected with ZP3 genes differentially glycosylate similar polypeptides and secrete active mouse sperm receptor. Journal of Cell Biology, 115(3):655–??, November 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/115/3/655.

Kitten:1991:CMR

[KN91] G. T. Kitten and E. A. Nigg. The CaaX motif is required for isoprenylation, carboxyl methylation, and nuclear membrane association of lamin B2. Journal of Cell Biology, 113 (1):13–??, April 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/113/1/13.

Kojima:1993:RTA

[KNR93] S. Kojima, K. Nara, and D. B. Rifkin. Requirement for transglutaminase in the activation of latent transforming growth factor-beta in bovine endothelial cells. Journal of Cell Biology, 121(2):439–??, April 1993. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/121/2/439.

Knudson:1993:HRD

[Knu93] C. B. Knudson. Hyaluronan receptor-directed assembly of chondrocyte pericellular matrix. Journal of Cell Biology, 120(3):825–??, February 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/120/3/825.

Ku:1994:IMP

[KO94] N. O. Ku and M. B. Omary. Identification of the major physiologic phosphorylation site of human keratin 18: potential kinases and a role in filament reorganization. Journal of Cell Biology, 127(1):161–??, October 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/127/1/161.

Kunimoto:1991:NKI

[KOB91] M. Kunimoto, E. Otto, and V. Bennett. A new 440-kD isoform is the major ankyrin in neonatal rat brain. Journal of Cell Biology, 115(5):1319–??, December 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/115/5/1319.

Kosik:1994:ADS

[Kos94] K. S. Kosik. The Alzheimer’s disease sphinx: a riddle with plaques and tangles. Journal of Cell Biology, 127(6):1501– ??, December 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/127/6/1501.

Kozak:1991:AVM

[Koz91] M. Kozak. An analysis of vertebrate mRNA sequences: intimations of translational control. Journal of Cell Biology, 115(4):887–??, November 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/115/4/887.

Koval:1990:SIP

[KP90] M. Koval and R. E. Pagano. Sorting of an internalized plasma membrane lipid between recycling and degradative pathways in normal and Niemann–Pick, type A fibroblasts.

Journal of Cell Biology, 111(2):429–??, August 1990. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/111/2/
429.

Kuczmarski:1991:SFM

[KPAY91] E. R. Kuczmarski, L. Palivos, C. Aguado, and Z. L. Yao. Stopped-flow measurement of cytoskeletal contraction: Dictyostelium myosin II is specifically required for contraction of amoeba cytoskeletons. Journal of Cell Biology, 114(6): 1191–??, September 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/114/6/1191.

Kouklis:1991:PRC

[KPMG91] P. D. Kouklis, T. Papamarcaki, A. Merdes, and S. D. Georgatos. A potential role for the COOH-terminal domain in the lateral packing of type III intermediate filaments. Journal of Cell Biology, 114(4):773–??, August 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/114/4/773.

Kurpakus:1991:SRA

[KQJ91] M. A. Kurpakus, V. Quaranta, and J. C. Jones. Surface relocation of alpha 6 beta 4 integrins and assembly of hemidesmosomes in an in vitro model of wound healing. Journal of Cell Biology, 115(6):1737–??, December 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/115/6/
1737.

Kobayashi:1991:NIC

[KR91] T. Kobayashi and J. M. Robinson. A novel intracellular compartment with unusual secretory properties in human neutrophils. Journal of Cell Biology, 113(4):743–??, May 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/113/4/743.

Ktistakis:1991:PCC

[KRB91] N. T. Ktistakis, M. G. Roth, and G. S. Bloom. PtK1 cells contain a nondiffusible, dominant factor that makes the Golgi apparatus resistant to brefeldin A. Journal of Cell Biology, 113(5):1009–??, June 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/113/5/1009.

Knight:1993:ICD

[KRCT93] M. R. Knight, N. D. Read, A. K. Campbell, and A. J. Trewavas. Imaging calcium dynamics in living plants using semi-synthetic recombinant aequorins. Journal of Cell Biology, 121(1):83–??, April 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/121/1/83.

Kayyem:1992:BNC

[KRdlR^{+}92] J. F. Kayyem, J. M. Roman, E. J. de la Rosa, U. Schwarz, and W. J. Dreyer. Bravo/ nr-CAM is closely related to the cell adhesion molecules L1 and Ng–CAM and has a similar heterodimer structure. Journal of Cell Biology, 118 (5):1259–??, September 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/118/5/1259.

Kartasova:1993:MDS

[KRHY93] T. Kartasova, D. R. Roop, K. A. Holbrook, and S. H. Yuspa. Mouse differentiation-specific keratins 1 and 10 require a preexisting keratin scaffold to form a filament network. Journal of Cell Biology, 120(5):1251–??, March 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/120/5/
1251.

Kim:1993:TMM

[KRMG93] S. H. Kim, K. A. Roth, A. R. Moser, and J. I. Gordon. Transgenic mouse models that explore the multistep hypothesis of intestinal neoplasia. Journal of Cell Biology, 123(4):877–??, November 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/123/4/877.

Kirk:1993:RBC

[KRMK93] M. M. Kirk, A. Ransick, S. E. McRae, and D. L. Kirk. The relationship between cell size and cell fate in Volvox carteri. Journal of Cell Biology, 123(1):191–??, October 1993.

CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/123/1/
191.

Keller:1990:NSM

[KS90] F. Keller and S. Schacher. Neuron-specific membrane glycoproteins promoting neurite fasciculation in Aplysia californica. Journal of Cell Biology, 111(6):2637–??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/111/6/2637.

Kimura:1992:SDG

[KS92] H. Kimura and D. Schubert. Schwannoma-derived growth factor promotes the neuronal differentiation and survival of PC12 cells. Journal of Cell Biology, 116(3):777–??, February 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/116/3/777.

Kielty:1994:AFA

[KS94] C. M. Kielty and C. A. Shuttleworth. Abnormal fibrillin assembly by dermal fibroblasts from two patients with Marfan syndrome. Journal of Cell Biology, 124(6):997–??, March 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/124/6/997.

Kelly:1993:APB

[KSB^{+}93] J. L. Kelly, A. Sánchez, G. S. Brown, C. N. Chesterman, and M. J. Sleigh. Accumulation of PDGF B and cell-binding forms of PDGF A in the extracellular matrix. Journal of Cell Biology, 121(5):1153–??, June 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/121/5/1153.

Kuhn:1991:NOI

[KSC^{+}91] T. B. Kuhn, E. T. Stoeckli, M. A. Condrau, F. G. Rathjen, and P. Sonderegger. Neurite outgrowth on immobilized axonin-1 is mediated by a heterophilic interaction with L1(G4). Journal of Cell Biology, 115(4):1113–??, November 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/115/4/1113.

Kallunki:1992:TLC

[KSE^{+}92] P. Kallunki, K. Sainio, R. Eddy, M. Byers, T. Kallunki, H. Sariola, K. Beck, H. Hirvonen, T. B. Shows, and K. Tryggvason. A truncated laminin chain homologous to the B2 chain: structure, spatial expression, and chromosomal assignment. Journal of Cell Biology, 119(3):679–??, November 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/119/3/679.

Kartenbeck:1991:EJC

[KSFG91] J. Kartenbeck, M. Schmelz, W. W. Franke, and B. Geiger. Endocytosis of junctional cadherins in bovine kidney epithelial (MDBK) cells cultured in low Ca^{2+} ion medium. Journal of Cell Biology, 113(4):881–??, May 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/113/4/881.

Kishi:1993:RCD

[KSK^{+}93] K. Kishi, T. Sasaki, S. Kuroda, T. Itoh, and Y. Takai. Regulation of cytoplasmic division of Xenopus embryo by rho p21 and its inhibitory GDP/GTP exchange protein (rho GDI). Journal of Cell Biology, 120(5):1187–??, March 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/120/5/
1187.

Kansas:1994:REG

[KSL^{+}94] G. S. Kansas, K. B. Saunders, K. Ley, A. Zakrzewicz, R. M. Gibson, B. C. Furie, B. Furie, and T. F. Tedder. A role for the epidermal growth factor-like domain of P-selectin in ligand recognition and cell adhesion. Journal of Cell Biology, 124(4):609–??, February 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/124/4/609.

Kansas:1991:MMF

[KSST91] G. S. Kansas, O. Spertini, L. M. Stoolman, and T. F. Tedder. Molecular mapping of functional domains of the leukocyte receptor for endothelium, LAM-1. Journal of Cell Biology, 114(2):351–??, July 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/114/2/351.

Kaplan:1992:APS

[KSVM92] K. B. Kaplan, J. R. Swedlow, H. E. Varmus, and D. O. Morgan. Association of p60c-src with endosomal membranes in mammalian fibroblasts. Journal of Cell Biology, 118(2): 321–??, July 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/118/2/321.

Kondo:1994:KNM

[KSYN^{+}94] S. Kondo, R. Sato-Yoshitake, Y. Noda, H. Aizawa, T. Nakata, Y. Matsuura, and N. Hirokawa. KIF3A is a new microtubule-based anterograde motor in the nerve axon. Journal of Cell Biology, 125(5):1095–??, June 1994. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/125/5/
1095.

Kallunki:1992:HBM

[KT92] P. Kallunki and K. Tryggvason. Human basement membrane heparan sulfate proteoglycan core protein: a 467-kD protein containing multiple domains resembling elements of the low density lipoprotein receptor, laminin, neural cell adhesion molecules, and epidermal growth factor. Journal of Cell Biology, 116(2):559–??, January 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/116/2/559.

Kubota:1993:DID

[KT93] Y. Kubota and H. Takisawa. Determination of initiation of DNA replication before and after nuclear formation in Xenopus egg cell free extracts. Journal of Cell Biology, 123(6):1321–??, December 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/123/6/1321.

Koffer:1990:CSA

[KTG90] A. Koffer, P. E. Tatham, and B. D. Gomperts. Changes in the state of actin during the exocytotic reaction of permeabilized rat mast cells. Journal of Cell Biology, 111(3):919– ??, September 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/111/3/919.

Ktistakis:1990:CTR

[KTR90] N. T. Ktistakis, D. Thomas, and M. G. Roth. Characteristics of the tyrosine recognition signal for internalization of transmembrane surface glycoproteins. Journal of Cell Biology, 111(4):1393–??, October 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/111/4/1393.

Kargman:1992:AIT

[KVE92] S. Kargman, P. J. Vickers, and J. F. Evans. A23187-induced translocation of 5-lipoxygenase in osteosarcoma cells. Journal of Cell Biology, 119(6):1701–??, December 1992. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/119/6/
1701.

Kurtz:1990:DRM

[KVF^{+}90] A. Kurtz, F. Vogel, K. Funa, C. H. Heldin, and R. Grosse. Developmental regulation of mammary-derived growth inhibitor expression in bovine mammary tissue. Journal of Cell Biology, 110(5):1779–??, May 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/110/5/1779.

Knudsen:1992:PKH

[KW92a] K. A. Knudsen and M. J. Wheelock. Plakoglobin, or an 83-kD homologue distinct from beta-catenin, interacts with E-cadherin and N-cadherin. Journal of Cell Biology, 118(3): 671–??, August 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/118/3/671.

Kolodney:1992:ICF

[KW92b] M. S. Kolodney and R. B. Wysolmerski. Isometric contraction by fibroblasts and endothelial cells in tissue culture: a quantitative study. Journal of Cell Biology, 117 (1):73–??, April 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/117/1/73.

Kopan:1993:MNE

[KW93] R. Kopan and H. Weintraub. Mouse notch: expression in hair follicles correlates with cell fate determination. Journal of Cell Biology, 121(3):631–??, May 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/121/3/631.

Kennedy:1991:ABR

[KWFM91] S. P. Kennedy, S. L. Warren, B. G. Forget, and J. S. Morrow. Ankyrin binds to the 15th repetitive unit of erythroid and nonerythroid beta-spectrin. Journal of Cell Biology, 115(1): 267–??, October 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/115/1/267.

Kielty:1992:TVC

[KWGS92] C. M. Kielty, S. P. Whittaker, M. E. Grant, and C. A. Shuttleworth. Type VI collagen microfibrils: evidence for a structural association with hyaluronan. Journal of Cell Biology, 118(4):979–??, August 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/118/4/979.

Khurana:1992:SDC

[KWK92] T. S. Khurana, S. C. Watkins, and L. M. Kunkel. The subcellular distribution of chromosome 6-encoded dystrophinrelated protein in the brain. Journal of Cell Biology, 119(2): 357–??, October 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/119/2/357.

Koivunen:1994:IHS

[KWR94] E. Koivunen, B. Wang, and E. Ruoslahti. Isolation of a highly specific ligand for the alpha 5 beta 1 integrin from a phage display library. Journal of Cell Biology, 124(3): 373–??, February 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/124/3/373.

Kaushal:1994:OEN

[KWS94] G. P. Kaushal, P. D. Walker, and S. V. Shah. An old enzyme with a new function: purification and characterization of a distinct matrix-degrading metalloproteinase in rat kidney cortex and its identification as meprin. Journal of Cell Biology, 126(5):1319–??, September 1994. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/126/5/1319.

Kruger:1991:NLR

[KWW91] M. Kruger, J. Wright, and K. Wang. Nebulin as a length regulator of thin filaments of vertebrate skeletal muscles: correlation of thin filament length, nebulin size, and epitope profile. Journal of Cell Biology, 115(1):97–??, October 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/115/1/
97.

Kemperman:1994:MET

[KWW^{+}94] H. Kemperman, Y. Wijnands, J. Wesseling, C. M. Niessen, A. Sonnenberg, and E. Roos. The mucin epiglycanin on TA3/Ha carcinoma cells prevents alpha 6 beta 4-mediated adhesion to laminin and kalinin and E-cadherin-mediated cell–cell interaction. Journal of Cell Biology, 127(6):2071– ??, December 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/127/6/2071.

Kanda:1994:GAC

[KYA^{+}94] T. Kanda, H. Yoshino, T. Ariga, M. Yamawaki, and R. K. Yu. Glycosphingolipid antigens in cultured bovine brain microvascular endothelial cells: sulfoglucuronosyl paragloboside as a target of monoclonal IgM in demyelinative neuropathy [corrected]. Journal of Cell Biology, 126(1):235– ??, July 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/126/1/235.

Klemke:1994:RTK

[KYBC94] R. L. Klemke, M. Yebra, E. M. Bayna, and D. A. Cheresh. Receptor tyrosine kinase signaling required for integrin alpha v beta 5-directed cell motility but not adhesion on vitronectin. Journal of Cell Biology, 127(3):859–??, November 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/127/3/859.

Katagiri:1994:BMP

[KYK^{+}94] T. Katagiri, A. Yamaguchi, M. Komaki, E. Abe, N. Takahashi, T. Ikeda, V. Rosen, J. M. Wozney, A. Fujisawa-Sehara, and T. Suda. Bone morphogenetic protein-2 converts the differentiation pathway of C2C12 myoblasts into the osteoblast lineage. Journal of Cell Biology, 127(6):1755– ??, December 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/127/6/1755.

Korhonen:1990:AAS

[KYLV90] M. Korhonen, J. Ylänne, L. Laitinen, and I. Virtanen. The alpha 1-alpha 6 subunits of integrins are characteristically expressed in distinct segments of developing and adult human nephron. Journal of Cell Biology, 111(3):1245– ??, September 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/111/3/1245.

Kislauskis:1994:SRI

[KZS94] E. H. Kislauskis, X. Zhu, and R. H. Singer. Sequences responsible for intracellular localization of beta-actin messenger RNA also affect cell phenotype. Journal of Cell Biology, 127(2):441–??, October 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/127/2/441.

LHernault:1992:MPS

[LA92] S. W. L’Hernault and P. M. Arduengo. Mutation of a putative sperm membrane protein in Caenorhabditis elegans prevents sperm differentiation but not its associated meiotic divisions. Journal of Cell Biology, 119(1):55–??, October 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/119/1/55.

Lampe:1994:APE

[Lam94] P. D. Lampe. Analyzing phorbol ester effects on gap junctional communication: a dramatic inhibition of assembly. Journal of Cell Biology, 127(6):1895–??, December 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/127/6/
1895.

Lum:1993:TRP

[LASB^{+}93] H. Lum, T. T. Andersen, A. Siflinger-Birnboim, C. Tiruppathi, M. S. Goligorsky, J. W. Fenton, and A. B. Malik.

Thrombin receptor peptide inhibits thrombin-induced increase in endothelial permeability by receptor desensitization. Journal of Cell Biology, 120(6):1491–??, March 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/120/6/
1491.

LaFlamme:1992:RFR

[LAY92] S. E. LaFlamme, S. K. Akiyama, and K. M. Yamada. Regulation of fibronectin receptor distribution. Journal of Cell Biology, 117(2):437–??, April 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/117/2/437.

Levis:1992:AAS

[LB92a] M. J. Levis and H. R. Bourne. Activation of the alpha subunit of Gs in intact cells alters its abundance, rate of degradation, and membrane avidity. Journal of Cell Biology, 119(5):1297–??, December 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/119/5/1297.

Lewis:1992:NGC

[LB92b] A. K. Lewis and P. C. Bridgman. Nerve growth cone lamellipodia contain two populations of actin filaments that differ in organization and polarity. Journal of Cell Biology, 119(5):1219–??, December 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/119/5/1219.

Liu:1992:CTD

[LB92c] H. Liu and A. Bretscher. Characterization of TPM1 disrupted yeast cells indicates an involvement of tropomyosin in directed vesicular transport. Journal of Cell Biology, 118 (2):285–??, July 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/118/2/285.

Liu:1993:TTC

[LB93] N. Liu and D. T. Brown. Transient translocation of the cytoplasmic (endo) domain of a type I membrane glycoprotein into cellular membranes. Journal of Cell Biology, 120(4): 877–??, February 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/120/4/877.

Lillie:1994:ILU

[LB94] S. H. Lillie and S. S. Brown. Immunofluorescence localization of the unconventional myosin, Myo2p, and the putative kinesin-related protein, Smy1p, to the same regions of polarized growth in Saccharomyces cerevisiae. Journal of Cell Biology, 125(4):825–??, May 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/125/4/825.

Lallier:1992:ABI

[LBF92] T. Lallier and M. Bronner-Fraser. Alpha 1 beta 1 integrin on neural crest cells recognizes some laminin substrata in a Ca(2+)-independent manner. Journal of Cell Biology, 119(5):1335–??, December 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/119/5/1335.

Levin:1992:MCP

[LBH92] D. E. Levin and E. Bartlett-Heubusch. Mutants in the S. cerevisiae PKC1 gene display a cell cycle-specific osmotic stability defect. Journal of Cell Biology, 116(5):1221–??, March 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/116/5/1221.

Landis:1993:NLA

[LBH93] R. C. Landis, R. I. Bennett, and N. Hogg. A novel LFA-1 activation epitope maps to the I domain. Journal of Cell Biology, 120(6):1519–??, March 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/120/6/1519.

Lesuffleur:1991:DRG

[LBL^{+}91] T. Lesuffleur, A. Barbat, C. Luccioni, J. Beaumatin, M. Clair, A. Kornowski, E. Dussaulx, B. Dutrillaux, and A. Zweibaum. Dihydrofolate reductase gene amplificationassociated shift of differentiation in methotrexate-adapted HT-29 cells. Journal of Cell Biology, 115(5):1409–??, December 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/115/5/1409.

Lupa:1991:EAD

[LC91] M. T. Lupa and J. H. Caldwell. Effect of agrin on the distribution of acetylcholine receptors and sodium channels on adult skeletal muscle fibers in culture. Journal of Cell Biology, 115(3):765–??, November 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/115/3/765.

Lupa:1994:SCA

[LC94] M. T. Lupa and J. H. Caldwell. Sodium channels aggregate at former synaptic sites in innervated and denervated regenerating muscles. Journal of Cell Biology, 124(1):139–??, January 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/124/1/139.

Lazebnik:1993:NEA

[LCC^{+}93] Y. A. Lazebnik, S. Cole, C. A. Cooke, W. G. Nelson, and W. C. Earnshaw. Nuclear events of apoptosis in vitro in cell-free mitotic extracts: a model system for analysis of the active phase of apoptosis. Journal of Cell Biology, 123 (1):7–??, October 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/123/1/7.

Lee:1993:CSI

[LCD^{+}93] J. K. Lee, R. S. Coyne, R. R. Dubreuil, L. S. Goldstein, and D. Branton. Cell shape and interaction defects in alpha-spectrin mutants of Drosophila melanogaster. Journal of Cell Biology, 123(6):1797–??, December 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/123/6/1797.

Letai:1992:DEJ

[LCF92] A. Letai, P. A. Coulombe, and E. Fuchs. Do the ends justify the mean? Proline mutations at the ends of the keratin coiled-coil rod segment are more disruptive than internal mutations. Journal of Cell Biology, 116(5):1181–??, March 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/116/5/1181.

Lindholm:1992:TGF

[LCK^{+}92] D. Lindholm, E. Castrén, R. Kiefer, F. Zafra, and H. Thoenen. Transforming growth factor-beta 1 in the rat brain: increase after injury and inhibition of astrocyte proliferation. Journal of Cell Biology, 117(2):395–??, April 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/117/2/395.

Levine:1991:EPM

[LCKW91] R. J. Levine, P. D. Chantler, R. W. Kensler, and J. L. Woodhead. Effects of phosphorylation by myosin light chain kinase on the structure of Limulus thick filaments. Journal of Cell Biology, 113(3):563–??, May 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/113/3/563.

Lotz:1994:HTP

[LCLG94] M. Lotz, I. Clark-Lewis, and V. Ganu. HIV-1 transactivator protein Tat induces proliferation and TGF beta expression in human articular chondrocytes. Journal of Cell Biology, 124(3):365–??, February 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/124/3/365.

Li:1994:IDE

[LCM^{+}94] H. Li, S. K. Choudhary, D. J. Milner, M. I. Munir, I. R. Kuisk, and Y. Capetanaki. Inhibition of desmin expression blocks myoblast fusion and interferes with the myogenic regulators MyoD and myogenin. Journal of Cell Biology, 124 (5):827–??, March 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/124/5/827.

Lathrop:1990:CCR

[LCMP90] W. F. Lathrop, E. P. Carmichael, D. G. Myles, and P. Primakoff. cDNA cloning reveals the molecular structure of a sperm surface protein, PH-20, involved in sperm-egg adhesion and the wide distribution of its gene among mammals. Journal of Cell Biology, 111(6):2939–??, December 1990.

CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/111/6/
2939.

Lopez-Casillas:1994:BCA

[LCPAM94] F. López-Casillas, H. M. Payne, J. L. Andres, and J. Massagué. Betaglycan can act as a dual modulator of TGF-beta access to signaling receptors: mapping of ligand binding and GAG attachment sites. Journal of Cell Biology, 124(4): 557–??, February 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/124/4/557.

Lindholm:1993:NIT

[LCT^{+}93] D. Lindholm, E. Castrén, P. Tsoulfas, R. Kolbeck, M. da P. Berzaghi, A. Leingärtner, C. P. Heisenberg, L. Tessarollo, L. F. Parada, H. Thoenen, and L. Tesarollo. Neurotrophin-3 induced by tri-iodothyronine in cerebellar granule cells promotes Purkinje cell differentiation. Journal of Cell Biology, 122(2):443–??, July 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/122/2/443.

Lee:1991:MAF

[LD91] B. A. Lee and D. J. Donoghue. Membrane-anchored form of v–sis/PDGF–B induces mitogenesis without detectable PDGF receptor autophosphorylation. Journal of Cell Biology, 113(2):361–??, April 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/113/2/361.

Lee:1992:IRM

[LD92] B. A. Lee and D. J. Donoghue. Intracellular retention of membrane-anchored v-sis protein abrogates autocrine signal transduction. Journal of Cell Biology, 118(5):1057– ??, September 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/118/5/1057.

Lencer:1992:MCT

[LDNM92] W. I. Lencer, C. Delp, M. R. Neutra, and J. L. Madara. Mechanism of cholera toxin action on a polarized human intestinal epithelial cell line: role of vesicular traffic. Journal of Cell Biology, 117(6):1197–??, June 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/117/6/1197.

Lu:1992:VSA

[LDS^{+}92] M. H. Lu, C. DiLullo, T. Schultheiss, S. Holtzer, J. M. Murray, J. Choi, D. A. Fischman, and H. Holtzer. The vinculin/sarcomeric-alpha-actinin/alpha-actin nexus in cultured cardiac myocytes. Journal of Cell Biology, 117(5): 1007–??, June 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/117/5/1007.

Lenk:1992:UAE

[LDT^{+}92] S. E. Lenk, W. A. Dunn, J. S. Trausch, A. Ciechanover, and A. L. Schwartz. Ubiquitin-activating enzyme, E1, is associated with maturation of autophagic vacuoles. Journal of Cell Biology, 118(2):301–??, July 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/118/2/301.

Lawler:1993:ICT

[LDW^{+}93] J. Lawler, M. Duquette, C. A. Whittaker, J. C. Adams, K. McHenry, and D. W. DeSimone. Identification and characterization of thrombospondin-4, a new member of the thrombospondin gene family. Journal of Cell Biology, 120 (4):1059–??, February 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/120/4/1059.

Luscher:1992:MSP

[LE92] B. Lüscher and R. N. Eisenman. Mitosis-specific phosphorylation of the nuclear oncoproteins Myc and Myb. Journal of Cell Biology, 118(4):775–??, August 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/118/4/775.

Lim:1990:TMT

[LELB90] S. S. Lim, K. J. Edson, P. C. Letourneau, and G. G. Borisy. A test of microtubule translocation during neurite elongation. Journal of Cell Biology, 111(1):123–??, July 1990.

CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/111/1/
123.

Love:1993:HBA

[LEM93] D. C. Love, J. D. Esko, and D. M. Mosser. A heparinbinding activity on Leishmania amastigotes which mediates adhesion to cellular proteoglycans. Journal of Cell Biology, 123(3):759–??, November 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/123/3/759.

Lin:1993:CRD

[LF93] C. H. Lin and P. Forscher. Cytoskeletal remodeling during growth cone-target interactions. Journal of Cell Biology, 121(6):1369–??, June 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/121/6/1369.

Lokeshwar:1994:ABD

[LFB94] V. B. Lokeshwar, N. Fregien, and L. Y. Bourguignon. Ankyrin-binding domain of CD44(GP85) is required for the expression of hyaluronic acid-mediated adhesion function. Journal of Cell Biology, 126(4):1099–??, August 1994. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/126/4/
1099.

Lin:1994:CDU

[LFC94] S. X. Lin, K. L. Ferro, and C. A. Collins. Cytoplasmic dynein undergoes intracellular redistribution concomitant with phosphorylation of the heavy chain in response to serum starvation and okadaic acid. Journal of Cell Biology, 127(4):1009–??, November 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/127/4/1009.

Lynch:1991:MHA

[LFF91] R. M. Lynch, K. E. Fogarty, and F. S. Fay. Modulation of hexokinase association with mitochondria analyzed with quantitative three-dimensional confocal microscopy. Journal of Cell Biology, 112(3):385–??, February 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/112/3/385.

Lemansky:1990:DLG

[LFG^{+}90] P. Lemansky, S. H. Fatemi, B. Gorican, S. Meyale, R. Rossero, and A. M. Tartakoff. Dynamics and longevity of the glycolipid-anchored membrane protein, Thy-1. Journal of Cell Biology, 110(5):1525–??, May 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/110/5/1525.

Lepock:1993:PDI

[LFR93] J. R. Lepock, H. E. Frey, and K. P. Ritchie. Protein denaturation in intact hepatocytes and isolated cellular organelles during heat shock. Journal of Cell Biology, 122(6):1267– ??, September 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/122/6/1267.

Leavesley:1992:RIB

[LFWC92] D. I. Leavesley, G. D. Ferguson, E. A. Wayner, and D. A. Cheresh. Requirement of the integrin beta 3 subunit for carcinoma cell spreading or migration on vitronectin and fibrinogen. Journal of Cell Biology, 117(5):1101–??, June 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/117/5/
1101.

Lin:1993:DLP

[LG93a] Y. C. Lin and F. Grinnell. Decreased level of PDGFstimulated receptor autophosphorylation by fibroblasts in mechanically relaxed collagen matrices. Journal of Cell Biology, 122(3):663–??, August 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/122/3/663.

Lloyd:1993:NRG

[LG93b] C. Lloyd and P. Gunning. Noncoding regions of the gammaactin gene influence the impact of the gene on myoblast morphology. Journal of Cell Biology, 121(1):73–??, April 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/121/1/

Ludwig:1991:DNS

[LGH91] T. Ludwig, G. Griffiths, and B. Hoflack. Distribution of newly synthesized lysosomal enzymes in the endocytic pathway of normal rat kidney cells. Journal of Cell Biology, 115(6):1561–??, December 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/115/6/1561.

Linsenmayer:1993:TVC

[LGI^{+}93] T. F. Linsenmayer, E. Gibney, F. Igoe, M. K. Gordon, J. M. Fitch, L. I. Fessler, and D. E. Birk. Type V collagen: molecular structure and fibrillar organization of the chicken alpha 1(V) NH2-terminal domain, a putative regulator of corneal fibrillogenesis. Journal of Cell Biology, 121 (5):1181–??, June 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/121/5/1181.

Lyons:1990:MAL

[LGPM90] R. M. Lyons, L. E. Gentry, A. F. Purchio, and H. L. Moses. Mechanism of activation of latent recombinant transforming growth factor beta 1 by plasmin. Journal of Cell Biology, 110(4):1361–??, April 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/110/4/1361.

Lindberg:1993:MCI

[LGSB93] F. P. Lindberg, H. D. Gresham, E. Schwarz, and E. J. Brown. Molecular cloning of integrin-associated protein: an immunoglobulin family member with multiple membranespanning domains implicated in alpha v beta 3-dependent ligand binding. Journal of Cell Biology, 123(2):485–??, October 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/123/2/485.

Levin:1992:CEU

[LH92] J. Z. Levin and H. R. Horvitz. The Caenorhabditis elegans unc-93 gene encodes a putative transmembrane protein that regulates muscle contraction. Journal of Cell Biology, 117 (1):143–??, April 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/117/1/143.

Leibler:1993:PVR

[LH93] S. Leibler and D. A. Huse. Porters versus rowers: a unified stochastic model of motor proteins. Journal of Cell Biology, 121(6):1357–??, June 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/121/6/1357.

Luers:1993:BPI

[LHFV93] G. Lüers, T. Hashimoto, H. D. Fahimi, and A. Völkl. Biogenesis of peroxisomes: isolation and characterization of two distinct peroxisomal populations from normal and regenerating rat liver. Journal of Cell Biology, 121(6):1271–??, June 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/121/6/1271.

Lakkakorpi:1991:VRR

[LHH^{+}91] P. T. Lakkakorpi, M. A. Horton, M. H. Helfrich, E. K. Karhukorpi, and H. K. Väänänen. Vitronectin receptor has a role in bone resorption but does not mediate tight sealing zone attachment of osteoclasts to the bone surface. Journal of Cell Biology, 115(4):1179–??, November 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/115/4/1179.

Lipfert:1992:IDP

[LHS^{+}92] L. Lipfert, B. Haimovich, M. D. Schaller, B. S. Cobb, J. T. Parsons, and J. S. Brugge. Integrin-dependent phosphorylation and activation of the protein tyrosine kinase pp125FAK in platelets. Journal of Cell Biology, 119(4):905–??, November 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/119/4/905.

LaMorte:1993:MGF

[LHS^{+}93] V. J. LaMorte, A. T. Harootunian, A. M. Spiegel, R. Y. Tsien, and J. R. Feramisco. Mediation of growth factor induced DNA synthesis and calcium mobilization by Gq and Gi2. Journal of Cell Biology, 121(1):91–??, April 1993. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/121/1/

Lein:1991:NDT

[LHT^{+}91] P. J. Lein, D. Higgins, D. C. Turner, L. A. Flier, and V. P. Terranova. The NC1 domain of type IV collagen promotes axonal growth in sympathetic neurons through interaction with the alpha 1 beta 1 integrin. Journal of Cell Biology, 113 (2):417–??, April 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/113/2/417.

Lane:1994:SSC

[LIAJS94] T. F. Lane, M. L. Iruela-Arispe, R. S. Johnson, and E. H. Sage. SPARC is a source of copper-binding peptides that stimulate angiogenesis. Journal of Cell Biology, 125(4):929– ??, May 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/125/4/929.

Lo:1994:ITA

[LJHC94] S. H. Lo, P. A. Janmey, J. H. Hartwig, and L. B. Chen. Interactions of tensin with actin and identification of its three distinct actin-binding domains. Journal of Cell Biology, 125 (5):1067–??, June 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/125/5/1067.

Lee:1993:DSP

[LJJC93] G. M. Lee, B. Johnstone, K. Jacobson, and B. Caterson. The dynamic structure of the pericellular matrix on living cells. Journal of Cell Biology, 123(6):1899–??, December 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/123/6/1899.

Lutcke:1993:RNS

[LJP^{+}93] A. Lütcke, S. Jansson, R. G. Parton, P. Chavrier, A. Valencia, L. A. Huber, E. Lehtonen, and M. Zerial. Rab17, a novel small GTPase, is specific for epithelial cells and is induced during cell polarization. Journal of Cell Biology, 121 (3):553–??, May 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/121/3/553.

Lourim:1993:MAL

[LK93a] D. Lourim and G. Krohne. Membrane-associated lamins in Xenopus egg extracts: identification of two vesicle populations. Journal of Cell Biology, 123(3):501–??, November 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/123/3/501.

Lu:1993:IUT

[LK93b] C. Lu and R. S. Kerbel. Interleukin-6 undergoes transition from paracrine growth inhibitor to autocrine stimulator during human melanoma progression. Journal of Cell Biology, 120(5):1281–??, March 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/120/5/1281.

Luscinskas:1994:MRA

[LKD^{+}94] F. W. Luscinskas, G. S. Kansas, H. Ding, P. Pizcueta, B. E. Schleiffenbaum, T. F. Tedder, and M. A. Gimbrone. Monocyte rolling, arrest and spreading on IL-4-activated vascular endothelium under flow is mediated via sequential action of L-selectin, beta 1-integrins, and beta 2-integrins. Journal of Cell Biology, 125(6):1417–??, June 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/125/6/1417.

Ladinsky:1994:HTT

[LKF^{+}94] M. S. Ladinsky, J. R. Kremer, P. S. Furcinitti, J. R. McIntosh, and K. E. Howell. HVEM tomography of the trans-Golgi network: structural insights and identification of a lace-like vesicle coat. Journal of Cell Biology, 127(1):29–??, October 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/127/1/29.

Loidl:1994:HPR

[LKS94] J. Loidl, F. Klein, and H. Scherthan. Homologous pairing is reduced but not abolished in asynaptic mutants of yeast. Journal of Cell Biology, 125(6):1191–??, June 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/125/6/

Latham:1994:BAM

[LKSR94] V. M. Latham, E. H. Kislauskis, R. H. Singer, and A. F. Ross. Beta-actin mRNA localization is regulated by signal transduction mechanisms. Journal of Cell Biology, 126 (5):1211–??, September 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/126/5/1211.

Li:1993:RDW

[LKXS93] J. Li, D. C. Koay, H. Xiao, and A. C. Sartorelli. Regulation of the differentiation of WEHI-3B D+ leukemia cells by granulocyte colony-stimulating factor receptor. Journal of Cell Biology, 120(6):1481–??, March 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/120/6/1481.

Leno:1991:NMD

[LL91] G. H. Leno and R. A. Laskey. The nuclear membrane determines the timing of DNA replication in Xenopus egg extracts. Journal of Cell Biology, 112(4):557–??, February 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/112/4/557.

Lieuvin:1994:IMS

[LLDJ94] A. Lieuvin, J. C. Labbé, M. Dorée, and D. Job. Intrinsic microtubule stability in interphase cells. Journal of Cell Biology, 124(6):985–??, March 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/124/6/985.

Leube:1994:SSS

[LLG^{+}94] R. E. Leube, U. Leimer, C. Grund, W. W. Franke, N. Harth, and B. Wiedenmann. Sorting of synaptophysin into special vesicles in nonneuroendocrine epithelial cells. Journal of Cell Biology, 127(6):1589–??, December 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/127/6/1589.

Lenter:1994:MCG

[LLIV94] M. Lenter, A. Levinovitz, S. Isenmann, and D. Vestweber. Monospecific and common glycoprotein ligands for E- and P-selectin on myeloid cells. Journal of Cell Biology, 125 (2):471–??, April 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/125/2/471.

Levy:1991:TPT

[LLK91] F. Lévy, R. Larsson, and S. Kvist. Translocation of peptides through microsomal membranes is a rapid process and promotes assembly of HLA–B27 heavy chain and beta 2microglobulin translated in vitro. Journal of Cell Biology, 115(4):959–??, November 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/115/4/959.

Lee:1994:TFG

[LLO^{+}94] J. Lee, M. Leonard, T. Oliver, A. Ishihara, and K. Jacobson. Traction forces generated by locomoting keratocytes. Journal of Cell Biology, 127(6):1957–??, December 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/127/6/
1957.

Lee:1992:IAB

[LLSM92] E. C. Lee, M. M. Lotz, G. D. Steele, and A. M. Mercurio. The integrin alpha 6 beta 4 is a laminin receptor. Journal of Cell Biology, 117(3):671–??, May 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/117/3/671.

Lechtreck:1991:SMA

[LM91] K. F. Lechtreck and M. Melkonian. Striated microtubuleassociated fibers: identification of assemblin, a novel 34-kD protein that forms paracrystals of 2-nm filaments in vitro. Journal of Cell Biology, 115(3):705–??, November 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/115/3/
705.

Leno:1994:IDR

[LM94] G. H. Leno and R. Munshi. Initiation of DNA replication in nuclei from quiescent cells requires permeabilization of the nuclear membrane. Journal of Cell Biology, 127(1): 5–??, October 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/127/1/5.

Ley:1994:DIL

[LMB^{+}94a] S. C. Ley, M. Marsh, C. R. Bebbington, K. Proudfoot, and P. Jordan. Distinct intracellular localization of Lck and Fyn protein tyrosine kinases in human T lymphocytes. Journal of Cell Biology, 125(3):639–??, May 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/125/3/639.

Lukas:1994:DTV

[LMB^{+}94b] J. Lukas, H. Müller, J. Bartkova, D. Spitkovsky, A. A. Kjerulff, P. Jansen-Dürr, M. Strauss, and J. Bartek. DNA tumor virus oncoproteins and retinoblastoma gene mutations share the ability to relieve the cell’s requirement for cyclin D1 function in G1. Journal of Cell Biology, 125(3): 625–??, May 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/125/3/625.

Landis:1994:IDL

[LMH^{+}94] R. C. Landis, A. McDowall, C. L. Holness, A. J. Littler, D. L. Simmons, and N. Hogg. Involvement of the “I” domain of LFA-1 in selective binding to ligands ICAM-1 and ICAM-3. Journal of Cell Biology, 126(2):529–??, July 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/126/2/
529.

Levinovitz:1993:IGL

[LMIV93] A. Levinovitz, J. Mühlhoff, S. Isenmann, and D. Vestweber. Identification of a glycoprotein ligand for E-selectin on mouse myeloid cells. Journal of Cell Biology, 121(2): 449–??, April 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/121/2/449.

Lin:1994:HAS

[LML^{+}94] Y. Lin, K. Mahan, W. F. Lathrop, D. G. Myles, and P. Primakoff. A hyaluronidase activity of the sperm plasma membrane protein PH-20 enables sperm to penetrate the cumulus cell layer surrounding the egg. Journal of Cell Biology, 125(5):1157–??, June 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/125/5/1157.

Lunstrum:1991:IPC

[LMM^{+}91] G. P. Lunstrum, N. P. Morris, A. M. McDonough, D. R. Keene, and R. E. Burgeson. Identification and partial characterization of two type XII-like collagen molecules. Journal of Cell Biology, 113(4):963–??, May 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/113/4/963.

Lockerbie:1991:RPE

[LMP91] R. O. Lockerbie, V. E. Miller, and K. H. Pfenninger. Regulated plasmalemmal expansion in nerve growth cones. Journal of Cell Biology, 112(6):1215–??, March 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/112/6/1215.

Lin:1991:RCC

[LMSA91] H. C. Lin, M. S. Moore, D. A. Sanan, and R. G. Anderson. Reconstitution of clathrin-coated pit budding from plasma membranes. Journal of Cell Biology, 114(5):881– ??, September 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/114/5/881.

Li:1994:DCD

[LMSH94] M. Li, M. McGrail, M. Serr, and T. S. Hays. Drosophila cytoplasmic dynein, a microtubule motor that is asymmetrically localized in the oocyte. Journal of Cell Biology, 126 (6):1475–??, September 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/126/6/1475.

Lingappa:1994:ECC

[LMW^{+}94] J. R. Lingappa, R. L. Martin, M. L. Wong, D. Ganem, W. J. Welch, and V. R. Lingappa. A eukaryotic cytosolic chaperonin is associated with a high molecular weight intermediate in the assembly of hepatitis B virus capsid, a multimeric particle. Journal of Cell Biology, 125(1):99–??, April 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/125/1/
99.

Lopez:1994:PPT

[LNS^{+}94] M. C. Lopez, J. M. Nicaud, H. B. Skinner, C. Vergnolle, J. C. Kader, V. A. Bankaitis, and C. Gaillardin. A phosphatidylinositol/phosphatidylcholine transfer protein is required for differentiation of the dimorphic yeast Yarrowia lipolytica from the yeast to the mycelial form. Journal of Cell Biology, 125(1):113–??, April 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/125/1/113.

Lofgren:1993:BIE

[LNSSA93] R. Löfgren, J. Ng-Sikorski, A. Sjölander, and T. Andersson. Beta 2 integrin engagement triggers actin polymerization and phosphatidylinositol trisphosphate formation in non-adherent human neutrophils. Journal of Cell Biology, 123(6):1597–??, December 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/123/6/1597.

Lyons:1990:EMG

[LOC^{+}90] G. E. Lyons, M. Ontell, R. Cox, D. Sassoon, and M. Buckingham. The expression of myosin genes in developing skeletal muscle in the mouse embryo. Journal of Cell Biology, 111(4):1465–??, October 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/111/4/1465.

Lu:1993:ERC

[LOOM93] K. P. Lu, S. A. Osmani, A. H. Osmani, and A. R. Means. Essential roles for calcium and calmodulin in G2/M progression in Aspergillus nidulans. Journal of Cell Biology, 121 (3):621–??, May 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/121/3/621.

Lowe:1992:SSM

[Low92] M. E. Lowe. Site-specific mutations in the COOH-terminus of placental alkaline phosphatase: a single amino acid change converts a phosphatidylinositol-glycan-anchored protein to a secreted protein. Journal of Cell Biology, 116 (3):799–??, February 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/116/3/799.

Larjava:1990:NFB

[LPA^{+}90] H. Larjava, J. Peltonen, S. K. Akiyama, S. S. Yamada, H. R. Gralnick, J. Uitto, and K. M. Yamada. Novel function for beta 1 integrins in keratinocyte cell–cell interactions. Journal of Cell Biology, 110(3):803–??, March 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/110/3/803.

Lutcke:1993:AKP

[LPA^{+}93] H. Lütcke, S. Prehn, A. J. Ashford, M. Remus, R. Frank, and B. Dobberstein. Assembly of the 68- and 72-kD proteins of signal recognition particle with 7S RNA. Journal of Cell Biology, 121(5):977–??, June 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/121/5/977.

Lasek:1992:SAT

[LPK92] R. J. Lasek, P. Paggi, and M. J. Katz. Slow axonal transport mechanisms move neurofilaments relentlessly in mouse optic axons. Journal of Cell Biology, 117(3):607–??, May 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/117/3/
607.

Lorant:1991:CGP

[LPM^{+}91] D. E. Lorant, K. D. Patel, T. M. McIntyre, R. P. McEver, S. M. Prescott, and G. A. Zimmerman. Coexpression of GMP-140 and PAF by endothelium stimulated by histamine or thrombin: a juxtacrine system for adhesion and activation of neutrophils. Journal of Cell Biology, 115(1):223–??, October 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/115/1/223.

Lemmon:1991:SCH

[LPPCF91] S. K. Lemmon, A. Pellicena-Palle, K. Conley, and C. L. Freund. Sequence of the clathrin heavy chain from Saccharomyces cerevisiae and requirement of the COOH terminus for clathrin function. Journal of Cell Biology, 112(1):

65–??, January 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/112/1/65.

LeBivic:1990:BPP

[LQNRB90] A. Le Bivic, A. Quaroni, B. Nichols, and E. Rodriguez-Boulan. Biogenetic pathways of plasma membrane proteins in Caco-2, a human intestinal epithelial cell line. Journal of Cell Biology, 111(4):1351–??, October 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/111/4/1351.

Lew:1993:MYC

[LR93] D. J. Lew and S. I. Reed. Morphogenesis in the yeast cell cycle: regulation by Cdc28 and cyclins. Journal of Cell Biology, 120(6):1305–??, March 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/120/6/1305.

Levanony:1992:ENR

[LRAG92] H. Levanony, R. Rubin, Y. Altschuler, and G. Galili. Evidence for a novel route of wheat storage proteins to vacuoles. Journal of Cell Biology, 119(5):1117–??, December 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/119/5/
1117.

Lampugnani:1991:RIM

[LRDM91] M. G. Lampugnani, M. Resnati, E. Dejana, and P. C. Marchisio. The role of integrins in the maintenance of endothelial monolayer integrity. Journal of Cell Biology, 112(3):479–??, February 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/112/3/479.

Lampugnani:1992:NES

[LRR^{+}92] M. G. Lampugnani, M. Resnati, M. Raiteri, R. Pigott, A. Pisacane, G. Houen, L. P. Ruco, and E. Dejana. A novel endothelial-specific membrane protein is a marker of cell–cell contacts. Journal of Cell Biology, 118(6):1511– ??, September 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/118/6/1511.

Liska:1994:CSE

[LRSB94] D. J. Liska, M. J. Reed, E. H. Sage, and P. Bornstein. Cellspecific expression of alpha 1(I) collagen-hGH minigenes in transgenic mice. Journal of Cell Biology, 125(3):695–??, May 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/125/3/695.

Lane:1990:FMS

[LS90] T. F. Lane and E. H. Sage. Functional mapping of SPARC: peptides from two distinct Ca+(+)-binding sites modulate cell shape. Journal of Cell Biology, 111(6):3065–??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/111/6/3065.

Lynch:1992:DSC

[LS92] M. A. Lynch and L. A. Staehelin. Domain-specific and cell type-specific localization of two types of cell wall matrix polysaccharides in the clover root tip. Journal of Cell Biology, 118(2):467–??, July 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/118/2/467.

Luke:1991:CSS

[LSA91] M. M. Luke, A. Sutton, and K. T. Arndt. Characterization of SIS1, a Saccharomyces cerevisiae homologue of bacterial dnaJ proteins. Journal of Cell Biology, 114(4):623–??, August 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/114/4/623.

Laitinen:1990:CRO

[LSAH90] J. Laitinen, L. Sistonen, K. Alitalo, and E. Hölttä. c-ha-rasVal 12 oncogene-transformed NIH-3T3 fibroblasts display more decondensed nucleosomal organization than normal fibroblasts. Journal of Cell Biology, 111(1):9–??, July 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/111/1/9.

Loike:1993:APF

[LSC^{+}93] J. D. Loike, R. Silverstein, L. Cao, L. Solomon, J. Weitz, E. Haber, G. R. Matsueda, M. S. Bernatowicz, and S. C. Silverstein. Activated platelets form protected zones of adhesion on fibrinogen and fibronectin-coated surfaces. Journal of Cell Biology, 121(4):945–??, May 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/121/4/945.

Lloyd:1992:TNB

[LSG92] C. Lloyd, G. Schevzov, and P. Gunning. Transfection of nonmuscle beta- and gamma-actin genes into myoblasts elicits different feedback regulatory responses from endogenous actin genes. Journal of Cell Biology, 117(4):787–??, May 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/117/4/787.

Lippincott-Schwartz:1991:FIR

[LSGD^{+}91] J. Lippincott-Schwartz, J. Glickman, J. G. Donaldson, J. Robbins, T. E. Kreis, K. B. Seamon, M. P. Sheetz, and R. D. Klausner. Forskolin inhibits and reverses the effects of brefeldin A on Golgi morphology by a cAMPindependent mechanism. Journal of Cell Biology, 112(4): 567–??, February 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/112/4/567.

Li:1994:CRX

[LSK^{+}94] X. Li, W. Shou, M. Kloc, B. A. Reddy, and L. D. Etkin. Cytoplasmic retention of Xenopus nuclear factor 7 before the mid blastula transition uses a unique anchoring mechanism involving a retention domain and several phosphorylation sites. Journal of Cell Biology, 124(1):7–??, January 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/124/1/
7.

LeBivic:1990:VTE

[LSMRB90] A. Le Bivic, Y. Sambuy, K. Mostov, and E. Rodriguez-Boulan. Vectorial targeting of an endogenous apical membrane sialoglycoprotein and uvomorulin in MDCK cells.

Journal of Cell Biology, 110(5):1533–??, May 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/110/5/1533.

Liu:1994:PLV

[LSN^{+}94] Y. Liu, E. S. Schweitzer, M. J. Nirenberg, V. M. Pickel, C. J. Evans, and R. H. Edwards. Preferential localization of a vesicular monoamine transporter to dense core vesicles in PC12 cells. Journal of Cell Biology, 127(5):1419– ??, December 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/127/5/1419.

LeBivic:1991:IDC

[LSP^{+}91] A. Le Bivic, Y. Sambuy, A. Patzak, N. Patil, M. Chao, and E. Rodriguez-Boulan. An internal deletion in the cytoplasmic tail reverses the apical localization of human NGF receptor in transfected MDCK cells. Journal of Cell Biology, 115(3):607–??, November 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/115/3/607.

Lenormand:1993:GFI

[LSP^{+}93] P. Lenormand, C. Sardet, G. Pagès, G. L’Allemain, A. Brunet, and J. Pouysségur. Growth factors induce nuclear translocation of MAP kinases (p42mapk and p44mapk) but not of their activator MAP kinase kinase (p45mapkk) in fibroblasts. Journal of Cell Biology, 122 (5):1079–??, September 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/122/5/1079.

Li:1994:SPM

[LSP^{+}94] K. Li, C. S. Smagula, W. J. Parsons, J. A. Richardson, M. Gonzalez, H. K. Hagler, and R. S. Williams. Subcellular partitioning of MRP RNA assessed by ultrastructural and biochemical analysis. Journal of Cell Biology, 124(6): 871–??, March 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/124/6/871.

Lamoureux:1990:EMA

[LSR^{+}90a] P. Lamoureux, V. L. Steel, C. Regal, L. Adgate, R. E. Buxbaum, and S. R. Heidemann. Extracellular matrix allows PC12 neurite elongation in the absence of microtubules. Journal of Cell Biology, 110(1):71–??, January 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/110/1/71.

Lillien:1990:EMA

[LSR90b] L. E. Lillien, M. Sendtner, and M. C. Raff. Extracellular matrix-associated molecules collaborate with ciliary neurotrophic factor to induce type-2 astrocyte development. Journal of Cell Biology, 111(2):635–??, August 1990. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/111/2/
635.

Leavesley:1993:IBB

[LSRC93] D. I. Leavesley, M. A. Schwartz, M. Rosenfeld, and D. A. Cheresh. Integrin beta 1- and beta 3-mediated endothelial cell migration is triggered through distinct signaling mechanisms. Journal of Cell Biology, 121(1):163–??, April 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/121/1/
163.

Lyons:1990:DRM

[LSS^{+}90] G. E. Lyons, S. Schiaffino, D. Sassoon, P. Barton, and M. Buckingham. Developmental regulation of myosin gene expression in mouse cardiac muscle. Journal of Cell Biology, 111(6):2427–??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/111/6/2427.

Lisanti:1994:CCR

[LSV^{+}94] M. P. Lisanti, P. E. Scherer, J. Vidugiriene, Z. Tang, A. Hermanowski-Vosatka, Y. H. Tu, R. F. Cook, and M. Sargiacomo. Characterization of caveolin-rich membrane domains isolated from an endothelial-rich source: implications for human disease. Journal of Cell Biology, 126(1): 111–??, July 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/126/1/111.

Lee:1992:TDG

[LTP^{+}92] S. W. Lee, C. Tomasetto, D. Paul, K. Keyomarsi, and R. Sager. Transcriptional downregulation of gap-junction proteins blocks junctional communication in human mammary tumor cell lines. Journal of Cell Biology, 118(5):1213– ??, September 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/118/5/1213.

Lotti:1992:IAT

[LTPB92] L. V. Lotti, M. R. Torrisi, M. C. Pascale, and S. Bonatti. Immunocytochemical analysis of the transfer of vesicular stomatitis virus G glycoprotein from the intermediate compartment to the Golgi complex. Journal of Cell Biology, 118 (1):43–??, July 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/118/1/43.

Lisanti:1993:CFH

[LTS93] M. P. Lisanti, Z. L. Tang, and M. Sargiacomo. Caveolin forms a hetero-oligomeric protein complex that interacts with an apical GPI-linked protein: implications for the biogenesis of caveolae. Journal of Cell Biology, 123(3):595– ??, November 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/123/3/595.

Low:1992:SIP

[LTWH92] S. H. Low, B. L. Tang, S. H. Wong, and W. Hong. Selective inhibition of protein targeting to the apical domain of MDCK cells by brefeldin A. Journal of Cell Biology, 118 (1):51–??, July 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/118/1/51.

LaFlamme:1994:SSC

[LTYY94] S. E. LaFlamme, L. A. Thomas, S. S. Yamada, and K. M. Yamada. Single subunit chimeric integrins as mimics and inhibitors of endogenous integrin functions in receptor localization, cell spreading and migration, and matrix assembly.

Journal of Cell Biology, 126(5):1287–??, September 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/126/5/
1287.

Luther:1991:TDR

[Lut91] P. K. Luther. Three-dimensional reconstruction of a simple Z-band in fish muscle. Journal of Cell Biology, 113(5): 1043–??, June 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/113/5/1043.

Lencer:1990:EVR

[LVA^{+}90] W. I. Lencer, A. S. Verkman, M. A. Arnaout, D. A. Ausiello, and D. Brown. Endocytic vesicles from renal papilla which retrieve the vasopressin-sensitive water channel do not contain a functional H+ ATPase. Journal of Cell Biology, 111 (2):379–??, August 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/111/2/379.

Litvinov:1994:ECH

[LVB^{+}94] S. V. Litvinov, M. P. Velders, H. A. Bakker, G. J. Fleuren, and S. O. Warnaar. Ep-CAM: a human epithelial antigen is a homophilic cell–cell adhesion molecule. Journal of Cell Biology, 125(2):437–??, April 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/125/2/437.

Linstedt:1992:SAP

[LVBK92] A. D. Linstedt, M. L. Vetter, J. M. Bishop, and R. B. Kelly. Specific association of the proto-oncogene product pp60c-src with an intracellular organelle, the PC12 synaptic vesicle. Journal of Cell Biology, 117(5):1077–??, June 1992. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/117/5/
1077.

Lochter:1991:JTS

[LVK^{+}91] A. Lochter, L. Vaughan, A. Kaplony, A. Prochiantz, M. Schachner, and A. Faissner. J1/ tenascin in substratebound and soluble form displays contrary effects on neurite outgrowth. Journal of Cell Biology, 113(5):1159–??, June 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/113/5/1159.

Lorenzo:1993:IMI

[LVTB93] M. Lorenzo, A. M. Valverde, T. Teruel, and M. Benito. IGF-i is a mitogen involved in differentiation-related gene expression in fetal rat brown adipocytes. Journal of Cell Biology, 123(6):1567–??, December 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/123/6/1567.

Livingston:1992:PEA

[LW92] B. T. Livingston and F. H. Wilt. Phorbol esters alter cell fate during development of sea urchin embryos. Journal of Cell Biology, 119(6):1641–??, December 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/119/6/1641.

Lowey:1991:NAM

[LWB91] S. Lowey, G. S. Waller, and E. Bandman. Neonatal and adult myosin heavy chains form homodimers during avian skeletal muscle development. Journal of Cell Biology, 113 (2):303–??, April 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/113/2/303.

Librach:1991:KTI

[LWF^{+}91] C. L. Librach, Z. Werb, M. L. Fitzgerald, K. Chiu, N. M. Corwin, R. A. Esteves, D. Grobelny, R. Galardy, C. H. Damsky, and S. J. Fisher. 92-kD type IV collagenase mediates invasion of human cytotrophoblasts. Journal of Cell Biology, 113(2):437–??, April 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/113/2/437.

Longmore:1993:ERR

[LWHL93] G. D. Longmore, S. S. Watowich, D. J. Hilton, and H. F. Lodish. The erythropoietin receptor: its role in hematopoiesis and myeloproliferative diseases. Journal of Cell Biology, 123(6):1305–??, December 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/123/6/1305.

Levine:1991:ECA

[LWK91] R. J. Levine, J. L. Woodhead, and H. A. King. The effect of calcium activation of skinned fiber bundles on the structure of Limulus thick filaments. Journal of Cell Biology, 113 (3):573–??, May 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/113/3/573.

Lefebvre:1993:MOP

[LWK^{+}93] O. Lefebvre, C. Wolf, M. Kédinger, M. P. Chenard, C. Tomasetto, P. Chambon, and M. C. Rio. The mouse one P-domain (pS2) and two P-domain (mSP) genes exhibit distinct patterns of expression. Journal of Cell Biology, 122 (1):191–??, July 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/122/1/191.

Lefebvre:1992:BCA

[LWL^{+}92] O. Lefebvre, C. Wolf, J. M. Limacher, P. Hutin, C. Wendling, M. LeMeur, P. Basset, and M. C. Rio. The breast cancerassociated stromelysin-3 gene is expressed during mouse mammary gland apoptosis. Journal of Cell Biology, 119(4): 997–??, November 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/119/4/997.

Lee:1992:NST

[LWV92] T. H. Lee, H. G. Wisniewski, and J. Vilcek. A novel secretory tumor necrosis factor-inducible protein (TSG-6) is a member of the family of hyaluronate binding proteins, closely related to the adhesion receptor CD44. Journal of Cell Biology, 116(2):545–??, January 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/116/2/545.

Lee:1991:NGE

[LXM91] W. C. Lee, Z. X. Xue, and T. Mélèse. The NSR1 gene encodes a protein that specifically binds nuclear localization sequences and has two RNA recognition motifs. Journal of Cell Biology, 113(1):1–??, April 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/113/1/1.

Lee:1993:NOH

[LXWC93] M. K. Lee, Z. Xu, P. C. Wong, and D. W. Cleveland. Neurofilaments are obligate heteropolymers in vivo. Journal of Cell Biology, 122(6):1337–??, September 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/122/6/1337.

Legan:1994:BDF

[LYC^{+}94] P. K. Legan, K. K. Yue, M. A. Chidgey, J. L. Holton, R. W. Wilkinson, and D. R. Garrod. The bovine desmocollin family: a new gene and expression patterns reflecting epithelial cell proliferation and differentiation. Journal of Cell Biology, 126(2):507–??, July 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/126/2/507.

Lin:1994:RPT

[LYK^{+}94] T. H. Lin, A. Yurochko, L. Kornberg, J. Morris, J. J. Walker, S. Haskill, and R. L. Juliano. The role of protein tyrosine phosphorylation in integrin-mediated gene induction in monocytes. Journal of Cell Biology, 126(6):1585– ??, September 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/126/6/1585.

Lamoureux:1992:CIN

[LZBH92] P. Lamoureux, J. Zheng, R. E. Buxbaum, and S. R. Heidemann. A cytomechanical investigation of neurite growth on different culture surfaces. Journal of Cell Biology, 118(3): 655–??, August 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/118/3/655.

Lee:1993:ULD

[LZI^{+}93] G. M. Lee, F. Zhang, A. Ishihara, C. L. McNeil, and K. A. Jacobson. Unconfined lateral diffusion and an estimate of pericellular matrix viscosity revealed by measuring the mobility of gold-tagged lipids. Journal of Cell Biology, 120(1): 25–??, January 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/120/1/25.

Meyer:1990:BAF

[MA90] R. K. Meyer and U. Aebi. Bundling of actin filaments by alpha-actinin depends on its molecular length. Journal of Cell Biology, 110(6):2013–??, June 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/110/6/2013.

Machesky:1994:PCC

[MAA^{+}94] L. M. Machesky, S. J. Atkinson, C. Ampe, J. Vandekerckhove, and T. D. Pollard. Purification of a cortical complex containing two unconventional actins from Acanthamoeba by affinity chromatography on profilin-agarose. Journal of Cell Biology, 127(1):107–??, October 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/127/1/107.

Mostov:1992:PMP

[MAAO92] K. Mostov, G. Apodaca, B. Aroeti, and C. Okamoto. Plasma membrane protein sorting in polarized epithelial cells. Journal of Cell Biology, 116(3):577–??, February 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/116/3/
577.

Mikol:1990:SCL

[MAB^{+}90] D. D. Mikol, M. J. Alexakos, C. A. Bayley, R. S. Lemons, M. M. Le Beau, and K. Stefansson. Structure and chromosomal localization of the gene for the oligodendrocyte-myelin glycoprotein. Journal of Cell Biology, 111(6):2673–??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/111/6/2673.

Macintyre:1992:RES

[Mac92] S. S. Macintyre. Regulated export of a secretory protein from the ER of the hepatocyte: a specific binding site retaining C-reactive protein within the ER is downregulated during the acute phase response. Journal of Cell Biology, 118 (2):253–??, July 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/118/2/253.

McEwen:1993:SCT

[MAFR93] B. F. McEwen, J. T. Arena, J. Frank, and C. L. Rieder. Structure of the colcemid-treated PtK1 kinetochore outer plate as determined by high voltage electron microscopic tomography. Journal of Cell Biology, 120(2):301–??, January 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/120/2/301.

Maher:1991:TDR

[Mah91] P. A. Maher. Tissue-dependent regulation of protein tyrosine kinase activity during embryonic development. Journal of Cell Biology, 112(5):955–??, March 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/112/5/955.

Matsudaira:1994:FAA

[Mat94] P. Matsudaira. The fimbrin and alpha-actinin footprint on actin. Journal of Cell Biology, 126(2):285–??, July 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/126/2/
285.

Matten:1990:TPT

[MAWM90] W. T. Matten, M. Aubry, J. West, and P. F. Maness. Tubulin is phosphorylated at tyrosine by pp60c-src in nerve growth cone membranes. Journal of Cell Biology, 111(5): 1959–??, November 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/111/5/1959.

Meier:1990:NLS

[MB90] U. T. Meier and G. Blobel. A nuclear localization signal binding protein in the nucleolus. Journal of Cell Biology, 111(6):2235–??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/111/6/2235.

Marrs:1992:TMM

[MB92] J. A. Marrs and G. B. Bouck. The two major membrane skeletal proteins (articulins) of Euglena gracilis define a novel class of cytoskeletal proteins. Journal of Cell Biology, 118(6):1465–??, September 1992. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/118/6/1465.

Muller:1993:GIL

[MB93] G. Müller and W. Bandlow. Glucose induces lipolytic cleavage of a glycolipidic plasma membrane anchor in yeast. Journal of Cell Biology, 122(2):325–??, July 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/122/2/325.

Meier:1994:NMN

[MB94] U. T. Meier and G. Blobel. NAP57, a mammalian nucleolar protein with a putative homolog in yeast and bacteria. Journal of Cell Biology, 127(6):1505–??, December 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/127/6/
1505.

Mercer:1993:MCI

[MBB^{+}93] R. W. Mercer, D. Biemesderfer, D. P. Bliss, J. H. Collins, and B. Forbush. Molecular cloning and immunological characterization of the gamma polypeptide, a small protein associated with the Na,K–ATPase. Journal of Cell Biology, 121 (3):579–??, May 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/121/3/579.

Marks:1990:ICT

[MBC90] M. S. Marks, J. S. Blum, and P. Cresswell. Invariant chain trimers are sequestered in the rough endoplasmic reticulum in the absence of association with HLA class II antigens. Journal of Cell Biology, 111(3):839–??, September 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/111/3/
839.

Marchisio:1991:PEI

[MBC^{+}91] P. C. Marchisio, S. Bondanza, O. Cremona, R. Cancedda, and M. De Luca. Polarized expression of integrin receptors (alpha 6 beta 4, alpha 2 beta 1, alpha 3 beta 1, and alpha v beta 5) and their relationship with the cytoskeleton and basement membrane matrix in cultured human keratinocytes. Journal of Cell Biology, 112(4):761–??, February 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/112/4/761.

McCrea:1993:ISB

[MBG93] P. D. McCrea, W. M. Brieher, and B. M. Gumbiner. Induction of a secondary body axis in Xenopus by antibodies to beta-catenin. Journal of Cell Biology, 123(2):477–??, October 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/123/2/477.

Merdes:1991:FNV

[MBHG91] A. Merdes, M. Brunkener, H. Horstmann, and S. D. Georgatos. Filensin: a new vimentin-binding, polymerizationcompetent, and membrane-associated protein of the lens fiber cell. Journal of Cell Biology, 115(2):397–??, October 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/115/2/397.

Mansfield:1990:TSM

[MBS90] P. J. Mansfield, L. A. Boxer, and S. J. Suchard. Thrombospondin stimulates motility of human neutrophils. Journal of Cell Biology, 111(6):3077–??, December 1990. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/111/6/
3077.

Meiri:1991:MAS

[MBS91] K. F. Meiri, L. E. Bickerstaff, and J. E. Schwob. Monoclonal antibodies show that kinase C phosphorylation of GAP-43 during axonogenesis is both spatially and temporally restricted in vivo. Journal of Cell Biology, 112(5): 991–??, March 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/112/5/991.

Magyar:1994:DNC

[MBW^{+}94] J. P. Magyar, U. Bartsch, Z. Q. Wang, N. Howells, A. Aguzzi, E. F. Wagner, and M. Schachner. Degeneration of neural cells in the central nervous system of mice deficient in the gene for the adhesion molecule on Glia, the beta 2 subunit of murine Na,K–ATPase. Journal of Cell Biology, 127(3):835–??, November 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/127/3/835.

McNally:1991:ISN

[MBZL91] E. M. McNally, M. M. Bravo-Zehnder, and L. A. Leinwand. Identification of sequences necessary for the association of cardiac myosin subunits. Journal of Cell Biology, 113(3): 585–??, May 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/113/3/585.

Miller:1991:UMI

[MC91a] B. R. Miller and M. G. Cumsky. An unusual mitochondrial import pathway for the precursor to yeast cytochrome c oxidase subunit Va. Journal of Cell Biology, 112(5):833–??, March 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/112/5/833.

Milne:1991:CTS

[MC91b] J. L. Milne and M. B. Coukell. A Ca^{2+} transport system associated with the plasma membrane of Dictyostelium discoideum is activated by different chemoattractant receptors. Journal of Cell Biology, 112(1):103–??, January 1991. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/112/1/
103.

Moran:1991:FSE

[MC91c] P. Moran and I. W. Caras. Fusion of sequence elements from non-anchored proteins to generate a fully functional signal for glycophosphatidylinositol membrane anchor attachment. Journal of Cell Biology, 115(6):1595–??, December 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/115/6/1595.

Moran:1991:NSC

[MC91d] P. Moran and I. W. Caras. A nonfunctional sequence converted to a signal for glycophosphatidylinositol membrane anchor attachment. Journal of Cell Biology, 115(2):329–??, October 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/115/2/329.

Moran:1992:PCU

[MC92] P. Moran and I. W. Caras. Proteins containing an uncleaved signal for glycophosphatidylinositol membrane anchor attachment are retained in a post–ER compartment. Journal of Cell Biology, 119(4):763–??, November 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/119/4/763.

Miller:1993:ISP

[MC93] B. R. Miller and M. G. Cumsky. Intramitochondrial sorting of the precursor to yeast cytochrome c oxidase subunit Va. Journal of Cell Biology, 121(5):1021–??, June 1993. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/121/5/
1021.

Moran:1994:RGA

[MC94] P. Moran and I. W. Caras. Requirements for glycosylphosphatidylinositol attachment are similar but not identical in mammalian cells and parasitic protozoa. Journal of Cell Biology, 125(2):333–??, April 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/125/2/333.

McCormick:1993:STA

[MCDA^{+}93] B. A. McCormick, S. P. Colgan, C. Delp-Archer, S. I. Miller, and J. L. Madara. Salmonella typhimurium attachment to human intestinal epithelial monolayers: transcellular signalling to subepithelial neutrophils. Journal of Cell Biology, 123(4):895–??, November 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/123/4/895.

Mayfield:1994:TPM

[MCDY94] S. P. Mayfield, A. Cohen, A. Danon, and C. B. Yohn. Translation of the psbA mRNA of Chlamydomonas reinhardtii requires a structured RNA element contained within the 5’ untranslated region. Journal of Cell Biology, 127(6):1537– ??, December 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/127/6/1537.

Musil:1990:DPG

[MCEG90] L. S. Musil, B. A. Cunningham, G. M. Edelman, and D. A. Goodenough. Differential phosphorylation of the gap junction protein connexin43 in junctional communicationcompetent and -deficient cell lines. Journal of Cell Biology, 111(5):2077–??, November 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/111/5/2077.

McCormick:1991:SIN

[MCF91] M. B. McCormick, P. A. Coulombe, and E. Fuchs. Sorting out IF networks: consequences of domain swapping on IF recognition and assembly. Journal of Cell Biology, 113(5): 1111–??, June 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/113/5/1111.

Murugesan:1993:OLD

[MCF93] G. Murugesan, G. M. Chisolm, and P. L. Fox. Oxidized low density lipoprotein inhibits the migration of aortic endothelial cells in vitro. Journal of Cell Biology, 120(4):1011–??, February 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/120/4/1011.

Maldonado-Codina:1992:CBA

[MCG92] G. Maldonado-Codina and D. M. Glover. Cyclins A and B associate with chromatin and the polar regions of spindles, respectively, and do not undergo complete degradation at anaphase in syncytial Drosophila embryos. Journal of Cell Biology, 116(4):967–??, February 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/116/4/967.

Miquelis:1993:IRG

[MCJ^{+}93] R. Miquelis, J. Courageot, A. Jacq, O. Blanck, C. Perrin, and P. Bastiani. Intracellular routing of GLcNAc-bearing molecules in thyrocytes: selective recycling through the Golgi apparatus. Journal of Cell Biology, 123(6):1695–??, December 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/123/6/1695.

Ma:1992:CDR

[MCM92a] Y. Ma, R. B. Campenot, and F. D. Miller. Concentrationdependent regulation of neuronal gene expression by nerve growth factor. Journal of Cell Biology, 117(1):135–??, April 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/117/1/135.

Miller:1992:PGM

[MCM92b] S. G. Miller, L. Carnell, and H. H. Moore. Post-Golgi membrane traffic: brefeldin A inhibits export from distal Golgi compartments to the cell surface but not recycling. Journal of Cell Biology, 118(2):267–??, July 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/118/2/267.

Masui:1993:CEA

[MCM93] H. Masui, L. Castro, and J. Mendelsohn. Consumption of EGF by A431 cells: evidence for receptor recycling. Journal of Cell Biology, 120(1):85–??, January 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/120/1/85.

Mirzayan:1992:NGE

[MCS92] C. Mirzayan, C. S. Copeland, and M. Snyder. The NUF1 gene encodes an essential coiled-coil related protein that is a potential component of the yeast nucleoskeleton. Journal of Cell Biology, 116(6):1319–??, March 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/116/6/1319.

Muller:1994:SGP

[MDKB94] G. Müller, E. A. Dearey, A. Korndörfer, and W. Bandlow. Stimulation of a glycosyl-phosphatidylinositol-specific phospholipase by insulin and the sulfonylurea, glimepiride, in rat adipocytes depends on increased glucose transport. Journal of Cell Biology, 126(5):1267–??, September 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/126/5/1267.

Moser:1992:MMI

[MDRG92] A. R. Moser, W. F. Dove, K. A. Roth, and J. I. Gordon. The Min (multiple intestinal neoplasia) mutation: its effect on gut epithelial cell differentiation and interaction with a modifier system. Journal of Cell Biology, 116(6):1517–??, March 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/116/6/1517.

Mugrauer:1991:CEP

[ME91] G. Mugrauer and P. Ekblom. Contrasting expression patterns of three members of the myc family of protooncogenes in the developing and adult mouse kidney. Journal of Cell Biology, 112(1):13–??, January 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/112/1/13.

Mackay:1993:MAI

[MECE93] A. M. Mackay, D. M. Eckley, C. Chue, and W. C. Earnshaw. Molecular analysis of the INCENPs (inner centromere proteins): separate domains are required for association with microtubules during interphase and with the central spindle during anaphase. Journal of Cell Biology, 123(2):373–??, October 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/123/2/373.

Miettinen:1994:TBI

[MELD94] P. J. Miettinen, R. Ebner, A. R. Lopez, and R. Derynck. TGF-beta induced transdifferentiation of mammary epithelial cells to mesenchymal cells: involvement of type I receptors. Journal of Cell Biology, 127(6):2021–??, December 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/127/6/2021.

Milgram:1994:DTS

[MEM94] S. L. Milgram, B. A. Eipper, and R. E. Mains. Differential trafficking of soluble and integral membrane secretory granule-associated proteins. Journal of Cell Biology, 124(1): 33–??, January 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/124/1/33.

Mandell:1990:ITH

[MF90] R. B. Mandell and C. M. Feldherr. Identification of two HSP70-related Xenopus oocyte proteins that are capable of recycling across the nuclear envelope. Journal of Cell Biology, 111(5):1775–??, November 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/111/5/1775.

Monck:1992:EFP

[MF92] J. R. Monck and J. M. Fernandez. The exocytotic fusion pore. Journal of Cell Biology, 119(6):1395–??, December 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/119/6/1395.

Masuda:1992:CLM

[MFOI92] T. Masuda, N. Fujimaki, E. Ozawa, and H. Ishikawa. Confocal laser microscopy of dystrophin localization in guinea pig skeletal muscle fibers. Journal of Cell Biology, 119(3): 543–??, November 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/119/3/543.

Milev:1994:ICS

[MFS^{+}94] P. Milev, D. R. Friedlander, T. Sakurai, L. Karthikeyan, M. Flad, R. K. Margolis, M. Grumet, and R. U. Margolis. Interactions of the chondroitin sulfate proteoglycan phosphacan, the extracellular domain of a receptor-type protein tyrosine phosphatase, with neurons, glia, and neural cell adhesion molecules. Journal of Cell Biology, 127(6):1703– ??, December 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/127/6/1703.

Michaud:1991:MPN

[MG91a] N. Michaud and D. S. Goldfarb. Multiple pathways in nuclear transport: the import of U2 snRNP occurs by a novel kinetic pathway. Journal of Cell Biology, 112(2):215–??, January 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/112/2/215.

Musil:1991:BAC

[MG91b] L. S. Musil and D. A. Goodenough. Biochemical analysis of connexin43 intracellular transport, phosphorylation, and assembly into gap junctional plaques. Journal of Cell Biology, 115(5):1357–??, December 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/115/5/1357.

Michaud:1992:MUS

[MG92] N. Michaud and D. Goldfarb. Microinjected U snRNAs are imported to oocyte nuclei via the nuclear pore complex by three distinguishable targeting pathways. Journal of Cell Biology, 116(4):851–??, February 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/116/4/851.

Matthews:1994:EIB

[MG94a] K. R. Matthews and K. Gull. Evidence for an interplay between cell cycle progression and the initiation of differentiation between life cycle forms of African trypanosomes. Journal of Cell Biology, 125(5):1147–??, June 1994. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/125/5/
1147.

McNew:1994:OPI

[MG94b] J. A. McNew and J. M. Goodman. An oligomeric protein is imported into peroxisomes in vivo. Journal of Cell Biology, 127(5):1245–??, December 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/127/5/1245.

Martins-Green:1990:LCA

[MGB90] M. Martins-Green and M. J. Bissell. Localization of 9E3/ CEF-4 in avian tissues: expression is absent in Rous sarcoma virus-induced tumors but is stimulated by injury. Journal of Cell Biology, 110(3):581–??, March 1990. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/110/3/
581.

Metsaranta:1992:CTM

[MGD^{+}92] M. Metsäranta, S. Garofalo, G. Decker, M. Rintala, B. de Crombrugghe, and E. Vuorio. Chondrodysplasia in transgenic mice harboring a 15-amino acid deletion in the triple helical domain of pro alpha 1(II) collagen chain. Journal of Cell Biology, 118(1):203–??, July 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/118/1/203.

Miller:1993:ECG

[MGDS93] D. J. Miller, X. Gong, G. Decker, and B. D. Shur. Egg cortical granule N-acetylglucosaminidase is required for the mouse zona block to polyspermy. Journal of Cell Biology, 123(6):1431–??, December 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/123/6/1431.

Merdes:1993:KLS

[MGG93] A. Merdes, F. Gounari, and S. D. Georgatos. The 47-kD lens-specific protein phakinin is a tailless intermediate filament protein and an assembly partner of filensin. Journal of Cell Biology, 123(6):1507–??, December 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/123/6/1507.

Mereau:1993:CBP

[MGPPH93] A. Mereau, L. Grey, C. Piquet-Pellorce, and J. K. Heath. Characterization of a binding protein for leukemia inhibitory factor localized in extracellular matrix. Journal of Cell Biology, 122(3):713–??, August 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/122/3/713.

Mikol:1990:OMG

[MGS90] D. D. Mikol, J. R. Gulcher, and K. Stefansson. The oligodendrocyte-myelin glycoprotein belongs to a distinct family of proteins and contains the HNK-1 carbohydrate. Journal of Cell Biology, 110(2):471–??, February 1990. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/110/2/
471.

Moss:1992:MBI

[MGW92] A. G. Moss, J. L. Gatti, and G. B. Witman. The motile beta/IC1 subunit of sea urchin sperm outer arm dynein does not form a rigor bond. Journal of Cell Biology, 118(5):1177– ??, September 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/118/5/1177.

Massia:1991:RSN

[MH91] S. P. Massia and J. A. Hubbell. An RGD spacing of 440 nm is sufficient for integrin alpha V beta 3-mediated fibroblast spreading and 140 nm for focal contact and stress fiber formation. Journal of Cell Biology, 114(5):1089–??, September 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/114/5/1089.

Marquardt:1992:MAN

[MH92] T. Marquardt and A. Helenius. Misfolding and aggregation of newly synthesized proteins in the endoplasmic reticulum. Journal of Cell Biology, 117(3):505–??, May 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/117/3/505.

Masumoto:1993:MPD

[MH93a] A. Masumoto and M. E. Hemler. Mutation of putative divalent cation sites in the alpha 4 subunit of the integrin VLA-4: distinct effects on adhesion to CS1/fibronectin, VCAM-1, and invasin. Journal of Cell Biology, 123(1):245–??, October 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/123/1/245.

Meresse:1993:PCI

[MH93b] S. Méresse and B. Hoflack. Phosphorylation of the cationindependent mannose 6-phosphate receptor is closely associated with its exit from the trans-Golgi network. Journal of Cell Biology, 120(1):67–??, January 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/120/1/67.

Madden:1994:CIO

[MH94] T. L. Madden and J. Herzfeld. Crowding-induced organization of cytoskeletal elements: II. Dissolution of spontaneously formed filament bundles by capping proteins. Journal of Cell Biology, 126(1):169–??, July 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/126/1/169.

Motley:1994:DPI

[MHDT94] A. Motley, E. Hettema, B. Distel, and H. Tabak. Differential protein import deficiencies in human peroxisome assembly disorders. Journal of Cell Biology, 125(4):755–??, May 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/125/4/
755.

Miyauchi:1990:OCC

[MHG^{+}90] A. Miyauchi, K. A. Hruska, E. M. Greenfield, R. Duncan, J. Alvarez, R. Barattolo, S. Colucci, A. Zambonin-Zallone, S. L. Teitelbaum, and A. Teti. Osteoclast cytosolic calcium, regulated by voltage-gated calcium channels and extracellular calcium, controls podosome assembly and bone resorption. Journal of Cell Biology, 111(6):2543–??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/111/6/2543.

Maison:1993:RDN

[MHG93] C. Maison, H. Horstmann, and S. D. Georgatos. Regulated docking of nuclear membrane vesicles to vimentin filaments during mitosis. Journal of Cell Biology, 123(6):1491–??, December 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/123/6/1491.

Monteiro:1990:ENB

[MHGC90] M. J. Monteiro, P. N. Hoffman, J. D. Gearhart, and D. W. Cleveland. Expression of NF–L in both neuronal and nonneuronal cells of transgenic mice: increased neurofilament density in axons without affecting caliber. Journal of Cell Biology, 111(4):1543–??, October 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/111/4/1543.

Monteiro:1994:DIS

[MHGJ94] M. J. Monteiro, C. Hicks, L. Gu, and S. Janicki. Determinants for intracellular sorting of cytoplasmic and nuclear intermediate filaments. Journal of Cell Biology, 127(5):1327– ??, December 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/127/5/1327.

Marks:1991:AFV

[MHM91] P. W. Marks, B. Hendey, and F. R. Maxfield. Attachment to fibronectin or vitronectin makes human neutrophil migration sensitive to alterations in cytosolic free calcium concentration. Journal of Cell Biology, 112(1):149–??, January 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/112/1/149.

Morgan:1990:NMC

[MHP90] J. E. Morgan, E. P. Hoffman, and T. A. Partridge. Normal myogenic cells from newborn mice restore normal histology to degenerating muscles of the mdx mouse. Journal of Cell Biology, 111(6):2437–??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/111/6/2437.

Malek-Hedayat:1994:EBR

[MHR94] S. Malek-Hedayat and L. H. Rome. Expression of a beta 1-related integrin by oligodendroglia in primary culture: evidence for a functional role in myelination. Journal of Cell Biology, 124(6):1039–??, March 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/124/6/1039.

Matsuyoshi:1992:CMC

[MHT^{+}92] N. Matsuyoshi, M. Hamaguchi, S. Taniguchi, A. Nagafuchi, S. Tsukita, and M. Takeichi. Cadherin-mediated cell–cell adhesion is perturbed by v-src tyrosine phosphorylation in metastatic fibroblasts. Journal of Cell Biology, 118(3): 703–??, August 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/118/3/703.

Metsikko:1992:LEE

[MHV92a] K. Metsikkö, T. Hentunen, and K. Väänänen. Local expression and exocytosis of viral glycoproteins in multinucleated muscle cells. Journal of Cell Biology, 117(5):987–??, June 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/117/5/987.

Miao:1992:TPS

[MHV92b] G. H. Miao, Z. Hong, and D. P. Verma. Topology and phosphorylation of soybean nodulin-26, an intrinsic protein of the peribacteroid membrane. Journal of Cell Biology, 118 (2):481–??, July 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/118/2/481.

Masuda:1990:VRS

[MHYC90] H. Masuda, T. Hirano, M. Yanagida, and W. Z. Cande. In vitro reactivation of spindle elongation in fission yeast nuc2 mutant cells. Journal of Cell Biology, 110(2):417–??, February 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/110/2/417.

Miyake:1992:RVV

[MHYK92] K. Miyake, Y. Hasunuma, H. Yagita, and M. Kimoto. Requirement for VLA-4 and VLA-5 integrins in lymphoma cells binding to and migration beneath stromal cells in culture. Journal of Cell Biology, 119(3):653–??, November 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/119/3/653.

Miller:1990:MPM

[Mil90] J. B. Miller. Myogenic programs of mouse muscle cell lines: expression of myosin heavy chain isoforms, MyoD1, and myogenin. Journal of Cell Biology, 111(3):1149–??, September 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/111/3/1149.

McNamee:1993:AFS

[MIS93] H. P. McNamee, D. E. Ingber, and M. A. Schwartz. Adhesion to fibronectin stimulates inositol lipid synthesis and enhances PDGF-induced inositol lipid breakdown. Journal of Cell Biology, 121(3):673–??, May 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/121/3/673.

Mitamura:1992:KDT

[MIU^{+}92] T. Mitamura, R. Iwamoto, T. Umata, T. Yomo, I. Urabe, M. Tsuneoka, and E. Mekada. The 27-kD diphtheria toxin receptor-associated protein (DRAP27) from vero cells is the monkey homologue of human CD9 antigen: expression of DRAP27 elevates the number of diphtheria toxin receptors on toxin-sensitive cells. Journal of Cell Biology, 118 (6):1389–??, September 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/118/6/1389.

Murray:1992:EHA

[MJ92] B. A. Murray and J. J. Jensen. Evidence for heterophilic adhesion of embryonic retinal cells and neuroblastoma cells to substratum-adsorbed NCAM. Journal of Cell Biology, 117 (6):1311–??, June 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/117/6/1311.

McCaffrey:1991:SGB

[MJG^{+}91] M. McCaffrey, J. S. Johnson, B. Goud, A. M. Myers, J. Rossier, M. R. Popoff, P. Madaule, and P. Boquet. The small GTP-binding protein Rho1p is localized on the Golgi apparatus and post-Golgi vesicles in Saccharomyces cerevisiae. Journal of Cell Biology, 115(2):309–??, October 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/115/2/309.

Moon:1993:CEC

[MJLD93] A. L. Moon, P. A. Janmey, K. A. Louie, and D. G. Drubin. Cofilin is an essential component of the yeast cortical cytoskeleton. Journal of Cell Biology, 120(2):421–??, January 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/120/2/421.

Morgan:1991:ECD

[MJM91] L. Morgan, K. R. Jessen, and R. Mirsky. The effects of cAMP on differentiation of cultured Schwann cells: progression from an early phenotype (04+) to a myelin phenotype (P0+, GFAP-, N-CAM-, NGF-receptor-) depends on growth inhibition. Journal of Cell Biology, 112(3):457–??, February 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/112/3/457.

Milgram:1992:EIF

[MJM92] S. L. Milgram, R. C. Johnson, and R. E. Mains. Expression of individual forms of peptidylglycine alpha-amidating monooxygenase in AtT-20 cells: endoproteolytic processing and routing to secretory granules. Journal of Cell Biology, 117(4):717–??, May 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/117/4/717.

Matsuuchi:1991:CBS

[MK91a] L. Matsuuchi and R. B. Kelly. Constitutive and basal secretion from the endocrine cell line, AtT-20. Journal of Cell Biology, 112(5):843–??, March 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/112/5/843.

Mitchell:1991:IOC

[MK91b] D. R. Mitchell and Y. Kang. Identification of oda6 as a Chlamydomonas dynein mutant by rescue with the wildtype gene. Journal of Cell Biology, 113(4):835–??, May 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/
113/4/835.

Masson:1993:IMC

[MK93] D. Masson and T. E. Kreis. Identification and molecular characterization of E–MAP–115, a novel microtubuleassociated protein predominantly expressed in epithelial cells. Journal of Cell Biology, 123(2):357–??, October 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/123/2/
357.

Menkel:1994:CFA

[MKB^{+}94] A. R. Menkel, M. Kroemker, P. Bubeck, M. Ronsiek, G. Nikolai, and B. M. Jockusch. Characterization of an Factin-binding domain in the cytoskeletal protein vinculin. Journal of Cell Biology, 126(5):1231–??, September 1994.

CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/126/5/
1231.

Mauro:1992:HHB

[MKCE92] V. P. Mauro, L. A. Krushel, B. A. Cunningham, and G. M. Edelman. Homophilic and heterophilic binding activities of Nr–CAM, a nervous system cell adhesion molecule. Journal of Cell Biology, 119(1):191–??, October 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/119/1/191.

Murray:1991:DFS

[MKF91] M. T. Murray, G. Krohne, and W. W. Franke. Different forms of soluble cytoplasmic mRNA binding proteins and particles in Xenopus laevis oocytes and embryos. Journal of Cell Biology, 112(1):1–??, January 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/112/1/1.

Miller:1993:DKA

[MKG93] R. K. Miller, S. Khuon, and R. D. Goldman. Dynamics of keratin assembly: exogenous type I keratin rapidly associates with type II keratin in vivo. Journal of Cell Biology, 122(1):123–??, July 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/122/1/123.

McCormick:1993:RRE

[MKSF93] M. B. McCormick, P. Kouklis, A. Syder, and E. Fuchs. The roles of the rod end and the tail in vimentin IF assembly and IF network formation. Journal of Cell Biology, 122 (2):395–??, July 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/122/2/395.

Mehta:1991:DRC

[ML91] P. P. Mehta and W. R. Loewenstein. Differential regulation of communication by retinoic acid in homologous and heterologous junctions between normal and transformed cells. Journal of Cell Biology, 113(2):371–??, April 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/113/2/371.

Matter:1994:NLE

[ML94] M. L. Matter and G. W. Laurie. A novel laminin E8 cell adhesion site required for lung alveolar formation in vitro. Journal of Cell Biology, 124(6):1083–??, March 1994. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/124/6/
1083.

Marinkovich:1992:DEJ

[MLKB92] M. P. Marinkovich, G. P. Lunstrum, D. R. Keene, and R. E. Burgeson. The dermal-epidermal junction of human skin contains a novel laminin variant. Journal of Cell Biology, 119(3):695–??, November 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/119/3/695.

Morgan:1993:ICR

[MLM^{+}93] T. E. Morgan, R. O. Lockerbie, L. S. Minamide, M. D. Browning, and J. R. Bamburg. Isolation and characterization of a regulated form of actin depolymerizing factor. Journal of Cell Biology, 122(3):623–??, August 1993. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/122/3/
623.

Mayer:1993:TIP

[MLN93] A. Mayer, R. Lill, and W. Neupert. Translocation and insertion of precursor proteins into isolated outer membranes of mitochondria. Journal of Cell Biology, 121(6):1233–??, June 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/121/6/1233.

Maycox:1992:CCV

[MLR^{+}92] P. R. Maycox, E. Link, A. Reetz, S. A. Morris, and R. Jahn. Clathrin-coated vesicles in nervous tissue are involved primarily in synaptic vesicle recycling. Journal of Cell Biology, 118(6):1379–??, September 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/118/6/1379.

Meyer:1992:IGJ

[MLRJ92] R. A. Meyer, D. W. Laird, J. P. Revel, and R. G. Johnson. Inhibition of gap junction and adherens junction assembly by connexin and A–CAM antibodies. Journal of Cell Biology, 119(1):179–??, October 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/119/1/179.

Metzger:1994:TCC

[MLS94] J. M. Metzger, W. I. Lin, and L. C. Samuelson. Transition in cardiac contractile sensitivity to calcium during the in vitro differentiation of mouse embryonic stem cells. Journal of Cell Biology, 126(3):701–??, August 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/126/3/701.

Marks:1990:TIC

[MM90] P. W. Marks and F. R. Maxfield. Transient increases in cytosolic free calcium appear to be required for the migration of adherent human neutrophils. Journal of Cell Biology, 110(1):43–??, January 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/110/1/43.

Miller:1991:RCS

[MM91] S. G. Miller and H. P. Moore. Reconstitution of constitutive secretion using semi-intact cells: regulation by GTP but not calcium. Journal of Cell Biology, 112(1):39–??, January 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/112/1/39.

Muir:1992:SGF

[MM92] D. Muir and M. Manthorpe. Stromelysin generates a fibronectin fragment that inhibits Schwann cell proliferation. Journal of Cell Biology, 116(1):177–??, January 1992. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/116/1/
177.

Mirabito:1993:BTC

[MM93a] P. M. Mirabito and N. R. Morris. BIMA, a TPR-containing protein required for mitosis, localizes to the spindle pole body in Aspergillus nidulans. Journal of Cell Biology, 120 (4):959–??, February 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/120/4/959.

Moschella:1993:ITR

[MM93b] M. C. Moschella and A. R. Marks. Inositol 1,4,5trisphosphate receptor expression in cardiac myocytes. Journal of Cell Biology, 120(5):1137–??, March 1993. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/120/5/
1137.

Mayor:1991:TGP

[MMC91] S. Mayor, A. K. Menon, and G. A. Cross. Transfer of glycosyl-phosphatidylinositol membrane anchors to polypeptide acceptors in a cell-free system. Journal of Cell Biology, 114(1):61–??, July 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/114/1/61.

Minshull:1991:XOM

[MMCH91] J. Minshull, A. Murray, A. Colman, and T. Hunt. Xenopus oocyte maturation does not require new cyclin synthesis. Journal of Cell Biology, 114(4):767–??, August 1991. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/114/4/
767.

Musil:1991:APM

[MMD91] K. J. Musil, A. Malmström, and J. Donnér. Alteration of proteoglycan metabolism during the differentiation of 3T3-L1 fibroblasts into adipocytes. Journal of Cell Biology, 114 (4):821–??, August 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/114/4/821.

Matunis:1992:CMH

[MMD92a] E. L. Matunis, M. J. Matunis, and G. Dreyfuss. Characterization of the major hnRNP proteins from Drosophila melanogaster. Journal of Cell Biology, 116(2):257–??, January 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/116/2/257.

Matunis:1992:IHC

[MMD92b] M. J. Matunis, E. L. Matunis, and G. Dreyfuss. Isolation of hnRNP complexes from Drosophila melanogaster. Journal of Cell Biology, 116(2):245–??, January 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/116/2/245.

Matunis:1993:AIH

[MMD93a] E. L. Matunis, M. J. Matunis, and G. Dreyfuss. Association of individual hnRNP proteins and snRNPs with nascent transcripts. Journal of Cell Biology, 121(2):219– ??, April 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/121/2/219.

Mundigl:1993:SVP

[MMD^{+}93b] O. Mundigl, M. Matteoli, L. Daniell, A. Thomas-Reetz, A. Metcalf, R. Jahn, and P. De Camilli. Synaptic vesicle proteins and early endosomes in cultured hippocampal neurons: differential effects of Brefeldin A in axon and dendrites. Journal of Cell Biology, 122(6):1207–??, September 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/122/6/1207.

Mastronarde:1993:ISM

[MMDM93] D. N. Mastronarde, K. L. McDonald, R. Ding, and J. R. McIntosh. Interpolar spindle microtubules in PTK cells. Journal of Cell Biology, 123(6):1475–??, December 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/123/6/
1475.

Milgram:1993:CTS

[MME93] S. L. Milgram, R. E. Mains, and B. A. Eipper. COOHterminal signals mediate the trafficking of a peptide processing enzyme in endocrine cells. Journal of Cell Biology, 121 (1):23–??, April 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/121/1/23.

Mathews:1992:PTS

[MMF92] P. M. Mathews, J. B. Martinie, and D. M. Fambrough. The pathway and targeting signal for delivery of the integral membrane glycoprotein LEP100 to lysosomes. Journal of Cell Biology, 118(5):1027–??, September 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/118/5/1027.

Meller:1994:RSP

[MMF94] V. H. Meller, M. McConnell, and P. A. Fisher. An RNasesensitive particle containing Drosophila melanogaster DNA topoisomerase II. Journal of Cell Biology, 126(6):1331– ??, September 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/126/6/1331.

Majack:1990:RPE

[MMG90] R. A. Majack, M. W. Majesky, and L. V. Goodman. Role of PDGF–A expression in the control of vascular smooth muscle cell growth by transforming growth factor-beta. Journal of Cell Biology, 111(1):239–??, July 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/111/1/239.

Miettinen:1992:FRE

[MMH^{+}92] H. M. Miettinen, K. Matter, W. Hunziker, J. K. Rose, and I. Mellman. Fc receptor endocytosis is controlled by a cytoplasmic domain determinant that actively prevents coated pit localization. Journal of Cell Biology, 116(4): 875–??, February 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/116/4/875.

Miyake:1991:VLA

[MMI^{+}91] K. Miyake, K. Medina, K. Ishihara, M. Kimoto, R. Auerbach, and P. W. Kincade. A VCAM-like adhesion molecule on murine bone marrow stromal cells mediates binding of lymphocyte precursors in culture. Journal of Cell Biology, 114(3):557–??, August 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/114/3/557.

Matsuzaki:1990:CCA

[MMJ^{+}90] F. Matsuzaki, R. M. Mège, S. H. Jaffe, D. R. Friedlander, W. J. Gallin, J. I. Goldberg, B. A. Cunningham, and G. M. Edelman. cDNAs of cell adhesion molecules of different specificity induce changes in cell shape and border formation in cultured S180 cells. Journal of Cell Biology, 110(4): 1239–??, April 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/110/4/1239.

Moir:1994:DPN

[MMLG94] R. D. Moir, M. Montag-Lowy, and R. D. Goldman. Dynamic properties of nuclear lamins: lamin B is associated with sites of DNA replication. Journal of Cell Biology, 125(6): 1201–??, June 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/125/6/1201.

McBride:1992:SAS

[MMLS92] H. M. McBride, D. G. Millar, J. M. Li, and G. C. Shore. A signal-anchor sequence selective for the mitochondrial outer membrane. Journal of Cell Biology, 119(6):1451–??, December 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/119/6/1451.

Mandelkow:1991:MDM

[MMM91] E. M. Mandelkow, E. Mandelkow, and R. A. Milligan. Microtubule dynamics and microtubule caps: a time-resolved cryo-electron microscopy study. Journal of Cell Biology, 114(5):977–??, September 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/114/5/977.

McPherson:1992:CLC

[MMM^{+}92] S. M. McPherson, P. S. McPherson, L. Mathews, K. P. Campbell, and F. J. Longo. Cortical localization of a calcium release channel in sea urchin eggs. Journal of Cell Biology, 116(5):1111–??, March 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/116/5/1111.

Mignatti:1991:EUR

[MMR91] P. Mignatti, R. Mazzieri, and D. B. Rifkin. Expression of the urokinase receptor in vascular endothelial cells is stimulated by basic fibroblast growth factor. Journal of Cell Biology, 113(5):1193–??, June 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/113/5/1193.

McCollum:1993:PMP

[MMS93] D. McCollum, E. Monosov, and S. Subramani. The pas8 mutant of Pichia pastoris exhibits the peroxisomal protein import deficiencies of Zellweger syndrome cells–the PAS8 protein binds to the COOH-terminal tripeptide peroxisomal targeting signal, and is a member of the TPR protein family. Journal of Cell Biology, 121(4):761–??, May 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/121/4/761.

Miller:1991:RNG

[MMT91] F. D. Miller, T. C. Mathew, and J. G. Toma. Regulation of nerve growth factor receptor gene expression by nerve growth factor in the developing peripheral nervous system. Journal of Cell Biology, 112(2):303–??, January 1991. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/112/2/
303.

Meyer:1992:ESB

[MMW^{+}92] M. Meyer, I. Matsuoka, C. Wetmore, L. Olson, and H. Thoenen. Enhanced synthesis of brain-derived neurotrophic factor in the lesioned peripheral nerve: different mechanisms are responsible for the regulation of BDNF and NGF mRNA. Journal of Cell Biology, 119(1):45–??, October 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/119/1/45.

McCammon:1994:IRP

[MMWG94] M. T. McCammon, J. A. McNew, P. J. Willy, and J. M. Goodman. An internal region of the peroxisomal membrane protein PMP47 is essential for sorting to peroxisomes. Journal of Cell Biology, 124(6):915–??, March 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/124/6/915.

Muro:1992:CPB

[MMY^{+}92] Y. Muro, H. Masumoto, K. Yoda, N. Nozaki, M. Ohashi, and T. Okazaki. Centromere protein B assembles human centromeric alpha-satellite DNA at the 17-bp sequence, CENP–B box. Journal of Cell Biology, 116(3):585–??, February 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/116/3/585.

Migliaccio:1992:SSR

[MNB92] G. Migliaccio, C. V. Nicchitta, and G. Blobel. The signal sequence receptor, unlike the signal recognition particle receptor, is not essential for protein translocation. Journal of Cell Biology, 117(1):15–??, April 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/117/1/15.

Marrs:1993:DRM

[MNME^{+}93] J. A. Marrs, E. W. Napolitano, C. Murphy-Erdosh, R. W. Mays, L. F. Reichardt, and W. J. Nelson. Distinguishing roles of the membrane-cytoskeleton and cadherin mediated cell–cell adhesion in generating different Na+,K(+)-ATPase distributions in polarized epithelia. Journal of Cell Biology, 123(1):149–??, October 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/123/1/149.

Morris:1990:TDS

[MNS90] E. P. Morris, G. Nneji, and J. M. Squire. The threedimensional structure of the nemaline rod Z-band. Journal of Cell Biology, 111(6):2961–??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/111/6/2961.

Mendoza:1993:GAS

[MNV93] L. M. Mendoza, D. Nishioka, and V. D. Vacquier. A GPIanchored sea urchin sperm membrane protein containing EGF domains is related to human uromodulin. Journal of Cell Biology, 121(6):1291–??, June 1993. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/121/6/1291.

Monck:1992:TSR

[MOKF92] J. R. Monck, A. F. Oberhauser, T. J. Keating, and J. M. Fernandez. Thin-section ratiometric Ca^{2+} images obtained by optical sectioning of fura-2 loaded mast cells. Journal of Cell Biology, 116(3):745–??, February 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/116/3/745.

Mastronarde:1992:AID

[MOM^{+}92] D. N. Mastronarde, E. T. O’Toole, K. L. McDonald, J. R. McIntosh, and M. E. Porter. Arrangement of inner dynein arms in wild-type and mutant flagella of Chlamydomonas. Journal of Cell Biology, 118(5):1145–??, September 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/118/5/
1145.

McDonald:1992:KMP

[MOMM92] K. L. McDonald, E. T. O’Toole, D. N. Mastronarde, and J. R. McIntosh. Kinetochore microtubules in PTK cells. Journal of Cell Biology, 118(2):369–??, July 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/118/2/369.

Morimoto:1990:ASC

[MOT90] T. Morimoto, S. Ogihara, and H. Takisawa. Anchorage of secretion-competent dense granules on the plasma membrane of bovine platelets in the absence of secretory stimulation. Journal of Cell Biology, 111(1):79–??, July 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/111/1/
79.

McIntosh:1991:MM

[MP91] J. R. McIntosh and C. M. Pfarr. Mitotic motors. Journal of Cell Biology, 115(3):577–??, November 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/115/3/577.

Martin:1994:ISF

[MP94] O. C. Martin and R. E. Pagano. Internalization and sorting of a fluorescent analogue of glucosylceramide to the Golgi apparatus of human skin fibroblasts: utilization of endocytic and nonendocytic transport mechanisms. Journal of Cell Biology, 125(4):769–??, May 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/125/4/769.

McGraw:1991:MHT

[MPAF91] T. E. McGraw, B. Pytowski, J. Arzt, and C. Ferrone. Mutagenesis of the human transferrin receptor: two cytoplasmic phenylalanines are required for efficient internalization and a second-site mutation is capable of reverting an internalization-defective phenotype. Journal of Cell Biology, 112(5):853–??, March 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/112/5/853.

Melchior:1993:INP

[MPEG93] F. Melchior, B. Paschal, J. Evans, and L. Gerace. Inhibition of nuclear protein import by nonhydrolyzable analogues of GTP and identification of the small GTPase Ran/TC4 as an essential transport factor. Journal of Cell Biology, 123(6):1649–??, December 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/123/6/1649.

McKinnon:1993:RTB

[MPIDD93] R. D. McKinnon, G. Piras, J. A. Ida, and M. Dubois-Dalcq. A role for TGF-beta in oligodendrocyte differentiation. Journal of Cell Biology, 121(6):1397–??, June 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/121/6/
1397.

Marquardt:1993:CRE

[MPK93] M. T. Marquardt, T. Phalen, and M. Kielian. Cholesterol is required in the exit pathway of Semliki Forest virus. Journal of Cell Biology, 123(1):57–??, October 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/123/1/57.

Mullock:1994:LCF

[MPK^{+}94] B. M. Mullock, J. H. Perez, T. Kuwana, S. R. Gray, and J. P. Luzio. Lysosomes can fuse with a late endosomal compartment in a cell-free system from rat liver. Journal of Cell Biology, 126(5):1173–??, September 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/126/5/1173.

Mayor:1993:SMC

[MPM93] S. Mayor, J. F. Presley, and F. R. Maxfield. Sorting of membrane components from endosomes and subsequent recycling to the cell surface occurs by a bulk flow process. Journal of Cell Biology, 121(6):1257–??, June 1993. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/121/6/
1257.

Mulholland:1994:UYA

[MPM^{+}94] J. Mulholland, D. Preuss, A. Moon, A. Wong, D. Drubin, and D. Botstein. Ultrastructure of the yeast actin cytoskeleton and its association with the plasma membrane. Journal of Cell Biology, 125(2):381–??, April 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/125/2/381.

MacKinnon:1990:SEG

[MPR90] P. J. MacKinnon, B. C. Powell, and G. E. Rogers. Structure and expression of genes for a class of cysteine-rich proteins of the cuticle layers of differentiating wool and hair follicles. Journal of Cell Biology, 111(6):2587–??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/111/6/
2587.

Moritz:1990:DYR

[MPTW90] M. Moritz, A. G. Paulovich, Y. F. Tsay, and J. L. Woolford. Depletion of yeast ribosomal proteins L16 or rp59 disrupts ribosome assembly. Journal of Cell Biology, 111(6):2261– ??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/111/6/2261.

Moudjou:1991:HCP

[MPVB91] M. Moudjou, M. Paintrand, B. Vigues, and M. Bornens. A human centrosomal protein is immunologically related to basal body-associated proteins from lower eucaryotes and is involved in the nucleation of microtubules. Journal of Cell Biology, 115(1):129–??, October 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/115/1/129.

Mallya:1992:PMY

[MPVL92] S. K. Mallya, J. S. Partin, M. C. Valdizan, and W. J. Lennarz. Proteolysis of the major yolk glycoproteins is regulated by acidification of the yolk platelets in sea urchin embryos. Journal of Cell Biology, 117(6):1211–??, June 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/117/6/
1211.

Moremen:1991:ICE

[MR91] K. W. Moremen and P. W. Robbins. Isolation, characterization, and expression of cDNAs encoding murine alphamannosidase II, a Golgi enzyme that controls conversion of high mannose to complex N-glycans. Journal of Cell Biology, 115(6):1521–??, December 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/115/6/1521.

Morla:1992:FSA

[MR92] A. Morla and E. Ruoslahti. A fibronectin self-assembly site involved in fibronectin matrix assembly: reconstruction in a synthetic peptide. Journal of Cell Biology, 118(2):421– ??, July 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/118/2/421.

Munn:1994:ERG

[MR94] A. L. Munn and H. Riezman. Endocytosis is required for the growth of vacuolar H(+)-ATPase-defective yeast: identification of six new END genes. Journal of Cell Biology, 127(2): 373–??, October 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/127/2/373.

Majesky:1990:PLR

[MRBP^{+}90] M. W. Majesky, M. A. Reidy, D. F. Bowen-Pope, C. E. Hart, J. N. Wilcox, and S. M. Schwartz. PDGF ligand and receptor gene expression during repair of arterial injury. Journal of Cell Biology, 111(5):2149–??, November 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/111/5/2149.

Mauviel:1991:MHD

[MRH^{+}91] A. Mauviel, F. Rédini, D. J. Hartmann, J. P. Pujol, and C. H. Evans. Modulation of human dermal fibroblast extracellular matrix metabolism by the lymphokine leukoregulin. Journal of Cell Biology, 113(6):1455–??, June 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/113/6/
1455.

McNeill:1993:STD

[MRSN93] H. McNeill, T. A. Ryan, S. J. Smith, and W. J. Nelson. Spatial and temporal dissection of immediate and early events following cadherin-mediated epithelial cell adhesion. Journal of Cell Biology, 120(5):1217–??, March 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/120/5/1217.

Mitchison:1992:PKF

[MS92] T. J. Mitchison and E. D. Salmon. Poleward kinetochore fiber movement occurs during both metaphase and anaphase–A in newt lung cell mitosis. Journal of Cell Biology, 119(3):569–??, November 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/119/3/569.

Miner:1994:CIA

[MS94] J. H. Miner and J. R. Sanes. Collagen IV alpha 3, alpha 4, and alpha 5 chains in rodent basal laminae: sequence, distribution, association with laminins, and developmental switches. Journal of Cell Biology, 127(3):879–??, November 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/127/3/879.

Mebius:1991:IAL

[MSB^{+}91] R. E. Mebius, P. R. Streeter, J. Brevé, A. M. Duijvestijn, and G. Kraal. The influence of afferent lymphatic vessel interruption on vascular addressin expression. Journal of Cell Biology, 115(1):85–??, October 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/115/1/85.

Machida:1991:NIM

[MSC91] C. M. Machida, J. D. Scott, and G. Ciment. NGF-induction of the metalloproteinase-transin/stromelysin in PC12 cells: involvement of multiple protein kinases. Journal of Cell Biology, 114(5):1037–??, September 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/114/5/1037.

Masuda:1992:VMN

[MSC92] H. Masuda, M. Sevik, and W. Z. Cande. In vitro microtubule-nucleating activity of spindle pole bodies in fission yeast Schizosaccharomyces pombe: cell cycle-dependent activation in Xenopus cell-free extracts. Journal of Cell Biology, 117(5):1055–??, June 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/117/5/1055.

Moore:1992:ISG

[MSD^{+}92a] K. L. Moore, N. L. Stults, S. Diaz, D. F. Smith, R. D. Cummings, A. Varki, and R. P. McEver. Identification of a specific glycoprotein ligand for P-selectin (CD62) on myeloid cells. Journal of Cell Biology, 118(2):445–??, July 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/118/2/
445.

Mosser:1992:LPR

[MSD92b] D. M. Mosser, T. A. Springer, and M. S. Diamond. Leishmania promastigotes require opsonic complement to bind to the human leukocyte integrin Mac-1 (CD11b/CD18). Journal of Cell Biology, 116(2):511–??, January 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/116/2/511.

Moll:1990:IPI

[MSF90] R. Moll, D. L. Schiller, and W. W. Franke. Identification of protein IT of the intestinal cytoskeleton as a novel type I cytokeratin with unusual properties and expression patterns. Journal of Cell Biology, 111(2):567–??, August 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/111/2/
567.

Moss:1992:ASS

[MSFW92] A. G. Moss, W. S. Sale, L. A. Fox, and G. B. Witman. The alpha subunit of sea urchin sperm outer arm dynein mediates structural and rigor binding to microtubules. Journal of Cell Biology, 118(5):1189–??, September 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/118/5/1189.

Matsumoto:1994:DTX

[MSI^{+}94] K. Matsumoto, Y. Saga, T. Ikemura, T. Sakakura, and R. Chiquet-Ehrismann. The distribution of tenascin–X is distinct and often reciprocal to that of tenascin–C. Journal of Cell Biology, 125(2):483–??, April 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/125/2/483.

Moses:1992:ICI

[MSL92] M. A. Moses, J. Sudhalter, and R. Langer. Isolation and characterization of an inhibitor of neovascularization from scapular chondrocytes. Journal of Cell Biology, 119(2): 475–??, October 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/119/2/475.

Moore:1991:SOA

[MSLS91] P. J. Moore, K. M. Swords, M. A. Lynch, and L. A. Staehelin. Spatial organization of the assembly pathways of glycoproteins and complex polysaccharides in the Golgi apparatus of plants. Journal of Cell Biology, 112(4):589–??, February 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/112/4/589.

McMorrow:1990:IGA

[MSPB90] I. McMorrow, W. E. Souter, G. Plopper, and B. Burke. Identification of a Golgi-associated protein that undergoes mitosis dependent phosphorylation and relocation. Journal of Cell Biology, 110(5):1513–??, May 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/110/5/1513.

McCarthy:1990:RIC

[MSQ^{+}90] J. B. McCarthy, A. P. Skubitz, Z. Qi, X. Y. Yi, D. J. Mickelson, D. J. Klein, and L. T. Furcht. RGD-independent cell adhesion to the carboxy-terminal heparin-binding fragment of fibronectin involves heparin-dependent and -independent activities. Journal of Cell Biology, 110(3):777–??, March 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/
110/3/777.

Martinez:1994:SGB

[MSS^{+}94] O. Martinez, A. Schmidt, J. Salaméro, B. Hoflack, M. Roa, and B. Goud. The small GTP-binding protein rab6 functions in intra-Golgi transport. Journal of Cell Biology, 127 (6):1575–??, December 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/127/6/1575.

Missero:1993:SSE

[MSSD93] C. Missero, C. Serra, K. Stenn, and G. P. Dotto. Skinspecific expression of a truncated E1a oncoprotein binding to p105-Rb leads to abnormal hair follicle maturation without increased epidermal proliferation. Journal of Cell Biology, 121(5):1109–??, June 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/121/5/1109.

McConnell:1990:TSY

[MSTY90] S. J. McConnell, L. C. Stewart, A. Talin, and M. P. Yaffe. Temperature-sensitive yeast mutants defective in mitochondrial inheritance. Journal of Cell Biology, 111(3):967– ??, September 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/111/3/967.

McGee:1994:PTP

[MSW^{+}94] T. P. McGee, H. B. Skinner, E. A. Whitters, S. A. Henry, and V. A. Bankaitis. A phosphatidylinositol transfer protein controls the phosphatidylcholine content of yeast Golgi membranes. Journal of Cell Biology, 124(3):273–??, February 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/124/3/273.

McMillan:1994:JGY

[MT94] J. N. McMillan and K. Tatchell. The JNM1 gene in the yeast Saccharomyces cerevisiae is required for nuclear migration and spindle orientation during the mitotic cell cycle. Journal of Cell Biology, 125(1):143–??, April 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/125/1/143.

Matteoli:1991:ARS

[MTC^{+}91] M. Matteoli, K. Takei, R. Cameron, P. Hurlbut, P. A. Johnston, T. C. Südhof, R. Jahn, and P. De Camilli. Association of Rab3A with synaptic vesicles at late stages of the secretory pathway. Journal of Cell Biology, 115(3):625–??, November 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/115/3/625.

Myers:1993:BVL

[MTJM93] J. N. Myers, I. Tabas, N. L. Jones, and F. R. Maxfield. Beta-very low density lipoprotein is sequestered in surfaceconnected tubules in mouse peritoneal macrophages. Journal of Cell Biology, 123(6):1389–??, December 1993. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/123/6/
1389.

Matteoli:1992:EER

[MTP^{+}92] M. Matteoli, K. Takei, M. S. Perin, T. C. Südhof, and P. De Camilli. Exo-endocytotic recycling of synaptic vesicles in developing processes of cultured hippocampal neurons. Journal of Cell Biology, 117(4):849–??, May 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/117/4/849.

Marks:1991:RRJ

[MTS^{+}91] A. R. Marks, M. B. Taubman, A. Saito, Y. Dai, and S. Fleischer. The ryanodine receptor/junctional channel complex is regulated by growth factors in a myogenic cell line. Journal of Cell Biology, 114(2):303–??, July 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/114/2/303.

Mackay:1994:EMC

[MTS^{+}94] C. R. Mackay, H. J. Terpe, R. Stauder, W. L. Marston, H. Stark, and U. Günthert. Expression and modulation of CD44 variant isoforms in humans. Journal of Cell Biology, 124(1):71–??, January 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/124/1/71.

Murphy-Ullrich:1991:FAI

[MULA^{+}91] J. E. Murphy-Ullrich, V. A. Lightner, I. Aukhil, Y. Z. Yan, H. P. Erickson, and M. Höök. Focal adhesion integrity is downregulated by the alternatively spliced domain of human tenascin. Journal of Cell Biology, 115(4):1127–??, November 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/115/4/1127.

Melki:1993:CMF

[MVCC93] R. Melki, I. E. Vainberg, R. L. Chow, and N. J. Cowan. Chaperonin-mediated folding of vertebrate actin-related protein and gamma-tubulin. Journal of Cell Biology, 122 (6):1301–??, September 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/122/6/1301.

Miller:1991:KII

[MVG91] R. K. Miller, K. Vikstrom, and R. D. Goldman. Keratin incorporation into intermediate filament networks is a rapid process. Journal of Cell Biology, 113(4):843–??, May 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/113/4/

Muir:1990:SCP

[MVM90] D. Muir, S. Varon, and M. Manthorpe. Schwann cell proliferation in vitro is under negative autocrine control. Journal of Cell Biology, 111(6):2663–??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/111/6/2663.

Moore:1991:GBG

[MVM91] K. L. Moore, A. Varki, and R. P. McEver. GMP-140 binds to a glycoprotein receptor on human neutrophils: evidence for a lectin-like interaction. Journal of Cell Biology, 112(3): 491–??, February 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/112/3/491.

Michel:1994:NRA

[MVTJ94] R. N. Michel, C. Q. Vu, W. Tetzlaff, and B. J. Jasmin. Neural regulation of acetylcholinesterase mRNAs at mammalian neuromuscular synapses. Journal of Cell Biology, 127(4):1061–??, November 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/127/4/1061.

Miron:1991:KIA

[MVV^{+}91] T. Miron, K. Vancompernolle, J. Vandekerckhove, M. Wilchek, and B. Geiger. A 25-kD inhibitor of actin polymerization is a low molecular mass heat shock protein. Journal of Cell Biology, 114(2):255–??, July 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/114/2/255.

Mundy:1992:MII

[MW92] D. I. Mundy and G. Warren. Mitosis and inhibition of intracellular transport stimulate palmitoylation of a 62-kD protein. Journal of Cell Biology, 116(1):135–??, January 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/116/1/
135.

Matera:1993:NOS

[MW93] A. G. Matera and D. C. Ward. Nucleoplasmic organization of small nuclear ribonucleoproteins in cultured human cells.

Journal of Cell Biology, 121(4):715–??, May 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/121/4/715.

Misteli:1994:CCV

[MW94] T. Misteli and G. Warren. COP-coated vesicles are involved in the mitotic fragmentation of Golgi stacks in a cell-free system. Journal of Cell Biology, 125(2):269–??, April 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/125/2/
269.

Miller:1990:CMC

[MWBD90] J. M. Miller, W. Wang, R. Balczon, and W. L. Dentler. Ciliary microtubule capping structures contain a mammalian kinetochore antigen. Journal of Cell Biology, 110(3):703–??, March 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/110/3/703.

McGough:1994:DAA

[MWD94] A. McGough, M. Way, and D. DeRosier. Determination of the alpha-actinin-binding site on actin filaments by cryoelectron microscopy and image analysis. Journal of Cell Biology, 126(2):433–??, July 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/126/2/433.

Mesner:1992:NGF

[MWG92] P. W. Mesner, T. R. Winters, and S. H. Green. Nerve growth factor withdrawal-induced cell death in neuronal PC12 cells resembles that in sympathetic neurons. Journal of Cell Biology, 119(6):1669–??, December 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/119/6/1669.

Mecham:1991:LAK

[MWH^{+}91] R. P. Mecham, L. Whitehouse, M. Hay, A. Hinek, and M. P. Sheetz. Ligand affinity of the 67-kD elastin/laminin binding protein is modulated by the protein’s lectin domain: visualization of elastin/laminin-receptor complexes with gold-tagged ligands. Journal of Cell Biology, 113(1): 187–??, April 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/113/1/187.

Mensa-Wilmot:1994:GGN

[MWLC^{+}94] K. Mensa-Wilmot, J. H. LeBowitz, K. P. Chang, A. al Qahtani, B. S. McGwire, S. Tucker, and J. C. Morris. A glycosylphosphatidylinositol (GPI)-negative phenotype produced in Leishmania major by GPI phospholipase C from Trypanosoma brucei: topography of two GPI pathways. Journal of Cell Biology, 124(6):935–??, March 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/124/6/935.

Munoz:1993:THR

[MWS^{+}93] A. Muñoz, C. Wrighton, B. Seliger, J. Bernal, and H. Beug. Thyroid hormone receptor/c-erbA: control of commitment and differentiation in the neuronal/chromaffin progenitor line PC12. Journal of Cell Biology, 121(2):423–??, April 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/
121/2/423.

Maciver:1991:AFS

[MWSP91] S. K. Maciver, D. H. Wachsstock, W. H. Schwarz, and T. D. Pollard. The actin filament severing protein actophorin promotes the formation of rigid bundles of actin filaments crosslinked with alpha-actinin. Journal of Cell Biology, 115 (6):1621–??, December 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/115/6/1621.

McConnell:1992:NMI

[MY92] S. J. McConnell and M. P. Yaffe. Nuclear and mitochondrial inheritance in yeast depends on novel cytoplasmic structures defined by the MDM1 protein. Journal of Cell Biology, 118 (2):385–??, July 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/118/2/385.

Mueller:1992:TPM

[MYC92] S. C. Mueller, Y. Yeh, and W. T. Chen. Tyrosine phosphorylation of membrane proteins mediates cellular invasion by transformed cells. Journal of Cell Biology, 119(5):1309– ??, December 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/119/5/1309.

Matter:1994:SRS

[MYM94] K. Matter, E. M. Yamamoto, and I. Mellman. Structural requirements and sequence motifs for polarized sorting and endocytosis of LDL and Fc receptors in MDCK cells. Journal of Cell Biology, 126(4):991–??, August 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/126/4/991.

Masaki:1994:MAD

[MYT94] R. Masaki, A. Yamamoto, and Y. Tashiro. Microsomal aldehyde dehydrogenase is localized to the endoplasmic reticulum via its carboxyl-terminal 35 amino acids. Journal of Cell Biology, 126(6):1407–??, September 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/126/6/1407.

Maciver:1991:CAF

[MZP91] S. K. Maciver, H. G. Zot, and T. D. Pollard. Characterization of actin filament severing by actophorin from Acanthamoeba castellanii. Journal of Cell Biology, 115(6):1611– ??, December 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/115/6/1611.

Mazzoni:1993:SMM

[MZRM93] C. Mazzoni, P. Zarov, A. Rambourg, and C. Mann. The SLT2 (MPK1) MAP kinase homolog is involved in polarized cell growth in Saccharomyces cerevisiae. Journal of Cell Biology, 123(6):1821–??, December 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/123/6/1821.

Nagy:1994:EHT

[NBT94] P. Nagy, H. C. Bisgaard, and S. S. Thorgeirsson. Expression of hepatic transcription factors during liver development and oval cell differentiation. Journal of Cell Biology, 126 (1):223–??, July 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/126/1/223.

Nandan:1990:EDI

[NCBS90] D. Nandan, E. P. Clarke, E. H. Ball, and B. D. Sanwal. Ethyl-3,4-dihydroxybenzoate inhibits myoblast differentiation: evidence for an essential role of collagen. Journal of Cell Biology, 110(5):1673–??, May 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/110/5/1673.

Nehme:1993:BDB

[NCM^{+}93] C. L. Nehme, M. M. Cesario, D. G. Myles, D. E. Koppel, and J. R. Bartles. Breaching the diffusion barrier that compartmentalizes the transmembrane glycoprotein CE9 to the posterior-tail plasma membrane domain of the rat spermatozoon. Journal of Cell Biology, 120(3):687–??, February 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/120/3/687.

Nehls:1991:HMP

[ND91] V. Nehls and D. Drenckhahn. Heterogeneity of microvascular pericytes for smooth muscle type alpha-actin. Journal of Cell Biology, 113(1):147–??, April 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/113/1/147.

Newport:1992:CMB

[ND92] J. Newport and W. Dunphy. Characterization of the membrane binding and fusion events during nuclear envelope assembly using purified components. Journal of Cell Biology, 116(2):295–??, January 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/116/2/295.

Niederreither:1992:MDS

[NDdC92] K. Niederreither, R. N. D’Souza, and B. de Crombrugghe. Minimal DNA sequences that control the cell lineagespecific expression of the pro alpha 2(I) collagen promoter in transgenic mice. Journal of Cell Biology, 119(5):1361– ??, December 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/119/5/1361.

Nuoffer:1994:GBR

[NDM^{+}94] C. Nuoffer, H. W. Davidson, J. Matteson, J. Meinkoth, and W. E. Balch. A GDP-bound of rab1 inhibits protein export from the endoplasmic reticulum and transport between Golgi compartments. Journal of Cell Biology, 125 (2):225–??, April 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/125/2/225.

Nishiyama:1991:PSN

[NDP^{+}91] A. Nishiyama, K. J. Dahlin, J. T. Prince, S. R. Johnstone, and W. B. Stallcup. The primary structure of NG2, a novel membrane-spanning proteoglycan. Journal of Cell Biology, 114(2):359–??, July 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/114/2/359.

Nguyen:1991:LDG

[NEL^{+}91] T. M. Nguyen, J. M. Ellis, D. R. Love, K. E. Davies, K. C. Gatter, G. Dickson, and G. E. Morris. Localization of the DMDL gene-encoded dystrophin-related protein using a panel of nineteen monoclonal antibodies: presence at neuromuscular junctions, in the sarcolemma of dystrophic skeletal muscle, in vascular and other smooth muscles, and in proliferating brain cell lines. Journal of Cell Biology, 115(6):1695–??, December 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/115/6/1695.

Newmeyer:1990:ESC

[NF90] D. D. Newmeyer and D. J. Forbes. An N-ethylmaleimide-sensitive cytosolic factor necessary for nuclear protein import: requirement in signal-mediated binding to the nuclear pore. Journal of Cell Biology, 110(3):547–??, March 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/110/3/

Noda:1992:NAP

[NF92] T. Noda and M. G. Farquhar. A non-autophagic pathway for diversion of ER secretory proteins to lysosomes. Journal of Cell Biology, 119(1):85–??, October 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/119/1/85.

Nixon:1990:SAT

[NFL90] R. A. Nixon, I. Fischer, and S. E. Lewis. Synthesis, axonal transport, and turnover of the high molecular weight microtubule-associated protein MAP 1A in mouse retinal ganglion cells: tubulin and MAP 1A display distinct transport kinetics. Journal of Cell Biology, 110(2):437–??, February 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/110/2/437.

Nishiyama:1993:ADD

[NFYI93] K. Nishiyama, T. Funai, S. Yokota, and A. Ichiyama. ATP-dependent degradation of a mutant serine: pyruvate/ alanine: glyoxylate aminotransferase in a primary hyperoxaluria type 1 case. Journal of Cell Biology, 123(5):1237– ??, December 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/123/5/1237.

North:1993:CDV

[NGB^{+}93] A. J. North, B. Galazkiewicz, T. J. Byers, J. R. Glenney, and J. V. Small. Complementary distributions of vinculin and dystrophin define two distinct sarcolemma domains in smooth muscle. Journal of Cell Biology, 120(5):1159–??, March 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/120/5/1159.

Navarro:1991:RCC

[NGP^{+}91] P. Navarro, M. Gómez, A. Pizarro, C. Gamallo, M. Quintanilla, and A. Cano. A role for the E-cadherin cell–cell adhesion molecule during tumor progression of mouse epidermal carcinogenesis. Journal of Cell Biology, 115(2):517–??, October 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/115/2/517.

Newman:1992:ADC

[NHA^{+}92] P. J. Newman, C. A. Hillery, R. Albrecht, L. V. Parise, M. C. Berndt, A. V. Mazurov, L. C. Dunlop, J. Zhang, and S. E. Rittenhouse. Activation-dependent changes in human platelet PECAM-1: phosphorylation, cytoskeletal association, and surface membrane redistribution. Journal of Cell Biology, 119(1):239–??, October 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/119/1/239.

Neuhaus:1991:CEC

[NHH^{+}91] H. Neuhaus, M. C. Hu, M. E. Hemler, Y. Takada, B. Holzmann, and I. L. Weissman. Cloning and expression of cDNAs for the alpha subunit of the murine lymphocyte– Peyer’s patch adhesion molecule. Journal of Cell Biology, 115(4):1149–??, November 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/115/4/1149.

Nickel:1994:ARF

[NHK^{+}94] W. Nickel, L. A. Huber, R. A. Kahn, N. Kipper, A. Barthel, D. Fasshauer, and H. D. Söling. ADP ribosylation factor and a 14-kD polypeptide are associated with heparan sulfate-carrying post-trans-Golgi network secretory vesicles in rat hepatocytes. Journal of Cell Biology, 125(4):721– ??, May 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/125/4/721.

Nakajima:1991:CRG

[NHO^{+}91] H. Nakajima, A. Hirata, Y. Ogawa, T. Yonehara, K. Yoda, and M. Yamasaki. A cytoskeleton-related gene, uso1, is required for intracellular protein transport in Saccharomyces cerevisiae. Journal of Cell Biology, 113(2):245–??, April 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/113/2/245.

Nathke:1994:DID

[NHS^{+}94] I. S. Näthke, L. Hinck, J. R. Swedlow, J. Papkoff, and W. J. Nelson. Defining interactions and distributions of cadherin and catenin complexes in polarized epithelial cells. Journal of Cell Biology, 125(6):1341–??, June 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/125/6/1341.

Neefjes:1990:IIT

[NHTP90] J. J. Neefjes, T. Hengeveld, O. Tol, and H. L. Ploegh. Intracellular interactions of transferrin and its receptor during biosynthesis. Journal of Cell Biology, 111(4):1383–??, October 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/111/4/1383.

Neame:1993:CTC

[NI93] S. J. Neame and C. M. Isacke. The cytoplasmic tail of CD44 is required for basolateral localization in epithelial MDCK cells but does not mediate association with the detergentinsoluble cytoskeleton of fibroblasts. Journal of Cell Biology, 121(6):1299–??, June 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/121/6/1299.

Nemoto:1993:RAC

[NIK^{+}93] Y. Nemoto, J. Ikeda, K. Katoh, H. Koshimoto, Y. Yoshihara, and K. Mori. R2D5 antigen: a calcium-binding phosphoprotein predominantly expressed in olfactory receptor neurons. Journal of Cell Biology, 123(4):963–??, November 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/123/4/963.

Nagafuchi:1994:RCC

[NIT94] A. Nagafuchi, S. Ishihara, and S. Tsukita. The roles of catenins in the cadherin-mediated cell adhesion: functional analysis of E-cadherin-alpha catenin fusion molecules. Journal of Cell Biology, 127(1):235–??, October 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/127/1/235.

Nickerson:1992:NMN

[NKW^{+}92] J. A. Nickerson, G. Krockmalnic, K. M. Wan, C. D. Turner, and S. Penman. A normally masked nuclear matrix antigen that appears at mitosis on cytoskeleton filaments adjoining chromosomes, centrioles, and midbodies. Journal of Cell Biology, 116(4):977–??, February 1992. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/116/4/977.

Nabi:1991:EML

[NLFRB91] I. R. Nabi, A. Le Bivic, D. Fambrough, and E. Rodriguez-Boulan. An endogenous MDCK lysosomal membrane glycoprotein is targeted basolaterally before delivery to lysosomes. Journal of Cell Biology, 115(6):1573–??, December 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/115/6/1573.

Narula:1992:IKB

[NMP^{+}92] N. Narula, I. McMorrow, G. Plopper, J. Doherty, K. S. Matlin, B. Burke, and J. L. Stow. Identification of a 200-kD, brefeldin-sensitive protein on Golgi membranes. Journal of Cell Biology, 117(1):27–??, April 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/117/1/27.

Nair:1990:SPC

[NMPN90] J. Nair, H. Müller, M. Peterson, and P. Novick. Sec2 protein contains a coiled-coil domain essential for vesicular transport and a dispensable carboxy terminal domain. Journal of Cell Biology, 110(6):1897–??, June 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/110/6/1897.

Navone:1992:CEH

[NNHB^{+}92] F. Navone, J. Niclas, N. Hom-Booher, L. Sparks, H. D. Bernstein, G. McCaffrey, and R. D. Vale. Cloning and expression of a human kinesin heavy chain gene: interaction of the COOH-terminal domain with cytoplasmic microtubules in transfected CV-1 cells. Journal of Cell Biology, 117(6): 1263–??, June 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/117/6/1263.

Nilsson:1993:ODT

[NPH^{+}93] T. Nilsson, M. Pypaert, M. H. Hoe, P. Slusarewicz, E. G. Berger, and G. Warren. Overlapping distribution of two glycosyltransferases in the Golgi apparatus of HeLa cells. Journal of Cell Biology, 120(1):5–??, January 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/120/1/5.

Norman:1994:AFO

[NPR^{+}94] J. C. Norman, L. S. Price, A. J. Ridley, A. Hall, and A. Koffer. Actin filament organization in activated mast cells is regulated by heterotrimeric and small GTP-binding proteins. Journal of Cell Biology, 126(4):1005–??, August 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/126/4/
1005.

Nixon:1994:PCT

[NPST94] R. A. Nixon, P. A. Paskevich, R. K. Sihag, and C. Y. Thayer. Phosphorylation on carboxyl terminus domains of neurofilament proteins in retinal ganglion cell neurons in vivo: influences on regional neurofilament accumulation, interneurofilament spacing, and axon caliber. Journal of Cell Biology, 126(4):1031–??, August 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/126/4/1031.

Nitkin:1990:AIR

[NR90] R. M. Nitkin and T. C. Rothschild. Agrin-induced reorganization of extracellular matrix components on cultured myotubes: relationship to AChR aggregation. Journal of Cell Biology, 111(3):1161–??, September 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/111/3/1161.

Ninomiya:1993:AAS

[NRS^{+}93a] H. Ninomiya, J. M. Roch, M. P. Sundsmo, D. A. Otero, and T. Saitoh. Amino acid sequence RERMS represents the active domain of amyloid beta/A4 protein precursor that promotes fibroblast growth. Journal of Cell Biology, 121 (4):879–??, May 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/121/4/879.

Nothwehr:1993:MPR

[NRS93b] S. F. Nothwehr, C. J. Roberts, and T. H. Stevens. Membrane protein retention in the yeast Golgi apparatus: dipeptidyl aminopeptidase A is retained by a cytoplasmic signal containing aromatic residues. Journal of Cell Biology, 121 (6):1197–??, June 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/121/6/1197.

Nathan:1990:TNF

[NS90] C. Nathan and E. Sanchez. Tumor necrosis factor and CD11/CD18 (beta 2) integrins act synergistically to lower cAMP in human neutrophils. Journal of Cell Biology, 111 (5):2171–??, November 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/111/5/2171.

Nathan:1991:CC

[NS91] C. Nathan and M. Sporn. Cytokines in context. Journal of Cell Biology, 113(5):981–??, June 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/113/5/981.

Needham:1993:CRP

[NS93] L. K. Needham and R. L. Schnaar. Carbohydrate recognition in the peripheral nervous system: a calcium-dependent membrane binding site for HNK-1 reactive glycolipids potentially involved in Schwann cell adhesion. Journal of Cell Biology, 121(2):397–??, April 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/121/2/397.

Nielsen:1993:CWC

[NSC^{+}93] S. Nielsen, B. L. Smith, E. I. Christensen, M. A. Knepper, and P. Agre. CHIP28 water channels are localized in constitutively water-permeable segments of the nephron. Journal of Cell Biology, 120(2):371–??, January 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/120/2/371.

Nakata:1990:CCL

[NSH90] T. Nakata, K. Sobue, and N. Hirokawa. Conformational change and localization of calpactin I complex involved in exocytosis as revealed by quick-freeze, deep-etch electron microscopy and immunocytochemistry. Journal of Cell Biology, 110(1):13–??, January 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/110/1/13.

Nakai:1992:ISP

[NSHN92] A. Nakai, M. Satoh, K. Hirayoshi, and K. Nagata. Involvement of the stress protein HSP47 in procollagen processing in the endoplasmic reticulum. Journal of Cell Biology, 117 (4):903–??, May 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/117/4/903.

Nislow:1990:MAM

[NSKM90] C. Nislow, C. Sellitto, R. Kuriyama, and J. R. McIntosh. A monoclonal antibody to a mitotic microtubule-associated protein blocks mitotic progression. Journal of Cell Biology, 111(2):511–??, August 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/111/2/511.

Nelson:1990:IMC

[NSWH90] W. J. Nelson, E. M. Shore, A. Z. Wang, and R. W. Hammerton. Identification of a membrane-cytoskeletal complex containing the cell adhesion molecule uvomorulin (Ecadherin), ankyrin, and fodrin in Madin–Darby canine kidney epithelial cells. Journal of Cell Biology, 110(2):349–??, February 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/110/2/349.

Nigam:1990:SDC

[NT90] S. K. Nigam and T. Towers. Subcellular distribution of calcium-binding proteins and a calcium–ATPase in canine pancreas. Journal of Cell Biology, 111(1):197–??, July 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/111/1/
197.

Nicosia:1994:MBT

[NT94] R. F. Nicosia and G. P. Tuszynski. Matrix-bound thrombospondin promotes angiogenesis in vitro. Journal of Cell Biology, 124(1):183–??, January 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/124/1/183.

Nuckolls:1990:FSD

[NTB90] G. H. Nuckolls, C. E. Turner, and K. Burridge. Functional studies of the domains of talin. Journal of Cell Biology, 110 (5):1635–??, May 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/110/5/1635.

Neugebauer:1992:ACM

[NVR92] K. M. Neugebauer, K. A. Venstrom, and L. F. Reichardt. Adhesion of a chicken myeloblast cell line to fibrinogen and vitronectin through a beta 1-class integrin. Journal of Cell Biology, 116(3):809–??, February 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/116/3/809.

Napolitano:1991:MCC

[NVWR91] E. W. Napolitano, K. Venstrom, E. F. Wheeler, and L. F. Reichardt. Molecular cloning and characterization of Bcadherin, a novel chick cadherin. Journal of Cell Biology, 113(4):893–??, May 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/113/4/893.

Nicklas:1994:EEC

[NW94] R. B. Nicklas and S. C. Ward. Elements of error correction in mitosis: microtubule capture, release, and tension. Journal of Cell Biology, 126(5):1241–??, September 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/126/5/
1241.

Newport:1990:LIP

[NWD90] J. W. Newport, K. L. Wilson, and W. G. Dunphy. A laminindependent pathway for nuclear envelope assembly. Journal of Cell Biology, 111(6):2247–??, December 1990. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/111/6/
2247.

Nilsson:1994:CTE

[NWvH94] I. Nilsson, P. Whitley, and G. von Heijne. The COOHterminal ends of internal signal and signal-anchor sequences are positioned differently in the ER translocase. Journal of Cell Biology, 126(5):1127–??, September 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/126/5/1127.

Nathan:1993:AIN

[NXHMJ93] C. Nathan, Q. W. Xie, L. Halbwachs-Mecarelli, and W. W. Jin. Albumin inhibits neutrophil spreading and hydrogen peroxide release by blocking the shedding of CD43 (sialophorin, leukosialin). Journal of Cell Biology, 122(1): 243–??, July 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/122/1/243.

Nagai:1991:MAC

[NYA^{+}91] T. Nagai, N. Yamakawa, S. Aota, S. S. Yamada, S. K. Akiyama, K. Olden, and K. M. Yamada. Monoclonal antibody characterization of two distant sites required for function of the central cell-binding domain of fibronectin in cell adhesion, cell migration, and matrix assembly. Journal of Cell Biology, 114(6):1295–??, September 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/114/6/1295.

OHalloran:1992:CHC

[OA92] T. J. O’Halloran and R. G. Anderson. Clathrin heavy chain is required for pinocytosis, the presence of large vacuoles, and development in Dictyostelium. Journal of Cell Biology, 118(6):1371–??, September 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/118/6/1371.

OConnor:1993:AAS

[OB93] T. P. O’Connor and D. Bentley. Accumulation of actin in subsets of pioneer growth cone filopodia in response to neural and epithelial guidance cues in situ. Journal of Cell Biology, 123(4):935–??, November 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/123/4/935.

Owens:1990:ERM

[OBBS90] G. C. Owens, C. J. Boyd, R. P. Bunge, and J. L. Salzer. Expression of recombinant myelin-associated glycoprotein in primary Schwann cells promotes the initial investment of axons by myelinating Schwann cells. Journal of Cell Biology, 111(3):1171–??, September 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/111/3/1171.

Omary:1992:PER

[OBC^{+}92] M. B. Omary, G. T. Baxter, C. F. Chou, C. L. Riopel, W. Y. Lin, and B. Strulovici. PKC epsilon-related kinase associates with and phosphorylates cytokeratin 8 and 18. Journal of Cell Biology, 117(3):583–??, May 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/117/3/583.

Ohlendieck:1991:DAP

[OC91] K. Ohlendieck and K. P. Campbell. Dystrophin-associated proteins are greatly reduced in skeletal muscle from mdx mice. Journal of Cell Biology, 115(6):1685–??, December 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/115/6/1685.

Ostrow:1994:EMR

[OCC94] B. D. Ostrow, P. Chen, and R. L. Chisholm. Expression of a myosin regulatory light chain phosphorylation site mutant complements the cytokinesis and developmental defects of Dictyostelium RMLC null cells. Journal of Cell Biology, 127(6):1945–??, December 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/127/6/1945.

Ossowski:1991:VPI

[OCMB91] L. Ossowski, G. Clunie, M. T. Masucci, and F. Blasi. In vivo paracrine interaction between urokinase and its receptor: effect on tumor cell invasion. Journal of Cell Biology, 115(4):1107–??, November 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/115/4/1107.

Owen:1993:MAS

[OD93] C. Owen and D. DeRosier. A 13-A map of the actin-scruin filament from the limulus acrosomal process. Journal of Cell Biology, 123(2):337–??, October 1993. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/123/2/337.

Oprins:1993:BCL

[ODK^{+}93] A. Oprins, R. Duden, T. E. Kreis, H. J. Geuze, and J. W. Slot. Beta-COP localizes mainly to the cis-Golgi side in exocrine pancreas. Journal of Cell Biology, 121(1):49–??, April 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/121/1/49.

Ohlendieck:1993:SUE

[ODPL93] K. Ohlendieck, S. T. Dhume, J. S. Partin, and W. J. Lennarz. The sea urchin egg receptor for sperm: isolation and characterization of the intact, biologically active receptor. Journal of Cell Biology, 122(4):887–??, August 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/122/4/
887.

Ohlendieck:1991:DGC

[OESC91] K. Ohlendieck, J. M. Ervasti, J. B. Snook, and K. P. Campbell. Dystrophin-glycoprotein complex is highly enriched in isolated skeletal muscle sarcolemma. Journal of Cell Biology, 112(1):135–??, January 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/112/1/135.

Ohara:1993:NDQ

[OGM^{+}93] O. Ohara, Y. Gahara, T. Miyake, H. Teraoka, and T. Kitamura. Neurofilament deficiency in quail caused by nonsense mutation in neurofilament–L gene. Journal of Cell Biology, 121(2):387–??, April 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/121/2/387.

Olwin:1990:FGF

[OH90] B. B. Olwin and S. D. Hauschka. Fibroblast growth factor receptor levels decrease during chick embryogenesis. Journal of Cell Biology, 110(2):503–??, February 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/110/2/503.

Okabe:1992:DBP

[OH92] S. Okabe and N. Hirokawa. Differential behavior of photoactivated microtubules in growing axons of mouse and frog neurons. Journal of Cell Biology, 117(1):105–??, April 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/117/1/
105.

Okabe:1993:DPF

[OH93] S. Okabe and N. Hirokawa. Do photobleached fluorescent microtubules move?: re-evaluation of fluorescence laser photobleaching both in vitro and in growing Xenopus axon. Journal of Cell Biology, 120(5):1177–??, March 1993. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/120/5/
1177.

Oko:1993:CDR

[OHC^{+}93] R. Oko, L. Hermo, P. T. Chan, A. Fazel, and J. J. Bergeron. The cytoplasmic droplet of rat epididymal spermatozoa contains saccular elements with Golgi characteristics. Journal of Cell Biology, 123(4):809–??, November 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/123/4/809.

Oberhammer:1994:CCD

[OHF^{+}94] F. A. Oberhammer, K. Hochegger, G. Fröschl, R. Tiefenbacher, and M. Pavelka. Chromatin condensation during apoptosis is accompanied by degradation of lamin A+B, without enhanced activation of cdc2 kinase. Journal of Cell Biology, 126(4):827–??, August 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/126/4/827.

Orci:1994:PID

[OHP^{+}94] L. Orci, P. Halban, A. Perrelet, M. Amherdt, M. Ravazzola, and R. G. Anderson. pH-independent and -dependent cleavage of proinsulin in the same secretory vesicle. Journal of Cell Biology, 126(5):1149–??, September 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/126/5/1149.

OKeefe:1992:DOD

[OHS92] R. T. O’Keefe, S. C. Henderson, and D. L. Spector. Dynamic organization of DNA replication in mammalian cell nuclei: spatially and temporally defined replication of chromosome-specific alpha-satellite DNA sequences. Journal of Cell Biology, 116(5):1095–??, March 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/116/5/1095.

Ozawa:1990:CPC

[OK90] M. Ozawa and R. Kemler. Correct proteolytic cleavage is required for the cell adhesive function of uvomorulin. Journal of Cell Biology, 111(4):1645–??, October 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/111/4/1645.

Ozawa:1992:MOU

[OK92] M. Ozawa and R. Kemler. Molecular organization of the uvomorulin-catenin complex. Journal of Cell Biology, 116 (4):989–??, February 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/116/4/989.

OToole:1994:ICD

[OKF^{+}94] T. E. O’Toole, Y. Katagiri, R. J. Faull, K. Peter, R. Tamura, V. Quaranta, J. C. Loftus, S. J. Shattil, and M. H. Ginsberg. Integrin cytoplasmic domains mediate inside-out signal transduction. Journal of Cell Biology, 124(6):1047–??, March 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/124/6/1047.

Orlandi:1992:MIR

[OKH92] P. A. Orlandi, F. W. Klotz, and J. D. Haynes. A malaria invasion receptor, the 175-kilodalton erythrocyte binding antigen of Plasmodium falciparum recognizes the terminal Neu5Ac(alpha 2-3)Gal- sequences of glycophorin A. Journal of Cell Biology, 116(4):901–??, February 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/116/4/901.

Otto:1991:ICC

[OKMB91] E. Otto, M. Kunimoto, T. McLaughlin, and V. Bennett. Isolation and characterization of cDNAs encoding human brain ankyrins reveal a family of alternatively spliced genes. Journal of Cell Biology, 114(2):241–??, July 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/114/2/241.

OShea:1990:REM

[OLKD90] K. S. O’Shea, L. H. Liu, L. H. Kinnunen, and V. M. Dixit. Role of the extracellular matrix protein thrombospondin in the early development of the mouse embryo. Journal of Cell Biology, 111(6):2713–??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/111/6/2713.

Ortenzi:1990:IIC

[OMBL90] C. Ortenzi, C. Miceli, R. A. Bradshaw, and P. Luporini. Identification and initial characterization of an autocrine pheromone receptor in the protozoan ciliate Euplotes raikovi. Journal of Cell Biology, 111(2):607–??, August 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/111/2/607.

Okabe:1993:DNI

[OMH93] S. Okabe, H. Miyasaka, and N. Hirokawa. Dynamics of the neuronal intermediate filaments. Journal of Cell Biology, 121(2):375–??, April 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/121/2/375.

OConnell:1993:SBM

[OMRM93] M. J. O’Connell, P. B. Meluh, M. D. Rose, and N. R. Morris. Suppression of the bimC4 mitotic spindle defect by deletion of klpA, a gene encoding a KAR3-related kinesin-like protein in Aspergillus nidulans. Journal of Cell Biology, 120(1): 153–??, January 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/120/1/153.

OKeefe:1994:DPM

[OMS^{+}94] R. T. O’Keefe, A. Mayeda, C. L. Sadowski, A. R. Krainer, and D. L. Spector. Disruption of pre-mRNA splicing in vivo results in reorganization of splicing factors. Journal of Cell Biology, 124(3):249–??, February 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/124/3/249.

Oka:1994:IGH

[ON94] T. Oka and A. Nakano. Inhibition of GTP hydrolysis by Sar1p causes accumulation of vesicles that are a functional intermediate of the ER-to-Golgi transport in yeast. Journal of Cell Biology, 124(4):425–??, February 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/124/4/425.

Oka:1991:RGB

[ONN91] T. Oka, S. Nishikawa, and A. Nakano. Reconstitution of GTP-binding Sar1 protein function in ER to Golgi transport. Journal of Cell Biology, 114(4):671–??, August 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/114/4/
671.

Olins:1992:BRH

[OOBJ92] A. L. Olins, D. E. Olins, and D. P. Bazett-Jones. Balbiani ring hnRNP substructure visualized by selective staining and electron spectroscopic imaging. Journal of Cell Biology, 117(3):483–??, May 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/117/3/483.

Osmani:1990:MCI

[OOM90] A. H. Osmani, S. A. Osmani, and N. R. Morris. The molecular cloning and identification of a gene product specifically required for nuclear movement in Aspergillus nidulans. Journal of Cell Biology, 111(2):543–??, August 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/111/2/

Otey:1990:IBA

[OPB90] C. A. Otey, F. M. Pavalko, and K. Burridge. An interaction between alpha-actinin and the beta 1 integrin subunit in vitro. Journal of Cell Biology, 111(2):721–??, August 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/111/2/
721.

Ohlendieck:1994:BAF

[OPL94] K. Ohlendieck, J. S. Partin, and W. J. Lennarz. The biologically active form of the sea urchin egg receptor for sperm is a disulfide-bonded homo-multimer. Journal of Cell Biology, 125(4):817–??, May 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/125/4/817.

Olwin:1992:RMD

[OR92] B. B. Olwin and A. Rapraeger. Repression of myogenic differentiation by aFGF, bFGF, and K–FGF is dependent on cellular heparan sulfate. Journal of Cell Biology, 118(3): 631–??, August 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/118/3/631.

OShea:1990:DRT

[ORD90] K. S. O’Shea, J. S. Rheinheimer, and V. M. Dixit. Deposition and role of thrombospondin in the histogenesis of the cerebellar cortex. Journal of Cell Biology, 110(4):1275–??, April 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/110/4/1275.

OHalloran:1990:EDM

[ORS90a] T. J. O’Halloran, S. Ravid, and J. A. Spudich. Expression of Dictyostelium myosin tail segments in Escherichia coli: domains required for assembly and phosphorylation. Journal of Cell Biology, 110(1):63–??, January 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/110/1/63.

Otte:1990:CMT

[ORS^{+}90b] A. P. Otte, D. Roy, M. Siemerink, C. H. Koster, F. Hochstenbach, A. Timmermans, and A. J. Durston.

Characterization of a maternal type VI collagen in Xenopus embryos suggests a role for collagen in gastrulation. Journal of Cell Biology, 111(1):271–??, July 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/111/1/271.

Ouyang:1992:IEP

[OS92] P. Ouyang and S. P. Sugrue. Identification of an epithelial protein related to the desmosome and intermediate filament network. Journal of Cell Biology, 118(6):1477–??, September 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/118/6/1477.

Osborne:1994:NSP

[OSJS94] M. A. Osborne, G. Schlenstedt, T. Jinks, and P. A. Silver. Nuf2, a spindle pole body-associated protein required for nuclear division in yeast. Journal of Cell Biology, 125 (4):853–??, May 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/125/4/853.

Oliver:1992:BMA

[OSK^{+}92] C. Oliver, N. Sahara, S. Kitani, A. R. Robbins, L. M. Mertz, and R. P. Siraganian. Binding of monoclonal antibody AA4 to gangliosides on rat basophilic leukemia cells produces changes similar to those seen with Fc epsilon receptor activation. Journal of Cell Biology, 116(3):635–??, February 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/116/3/635.

Ohba:1994:PHP

[OSKN94] M. Ohba, M. Shibanuma, T. Kuroki, and K. Nose. Production of hydrogen peroxide by transforming growth factorbeta 1 and its involvement in induction of egr-1 in mouse osteoblastic cells. Journal of Cell Biology, 126(4):1079–??, August 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/126/4/1079.

Ohmori:1992:DPP

[OTT92] H. Ohmori, S. Toyama, and S. Toyama. Direct proof that the primary site of action of cytochalasin on cell motility processes is actin. Journal of Cell Biology, 116(4):933–??, February 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/116/4/933.

Ostman:1992:PAP

[OTWH92] A. Ostman, J. Thyberg, B. Westermark, and C. H. Heldin. PDGF-AA and PDGF–BB biosynthesis: proprotein processing in the Golgi complex and lysosomal degradation of PDGF–BB retained intracellularly. Journal of Cell Biology, 118(3):509–??, August 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/118/3/509.

Odorizzi:1994:SSM

[OTX^{+}94] C. G. Odorizzi, I. S. Trowbridge, L. Xue, C. R. Hopkins, C. D. Davis, and J. F. Collawn. Sorting signals in the MHC class II invariant chain cytoplasmic tail and transmembrane region determine trafficking to an endocytic processing compartment. Journal of Cell Biology, 126(2):317– ??, July 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/126/2/317.

Oda:1993:IDH

[OUS^{+}93] H. Oda, T. Uemura, K. Shiomi, A. Nagafuchi, S. Tsukita, and M. Takeichi. Identification of a Drosophila homologue of alpha-catenin and its association with the armadillo protein. Journal of Cell Biology, 121(5):1133–??, June 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/121/5/
1133.

Osborn:1994:ADA

[OVB^{+}94] L. Osborn, C. Vassallo, B. G. Browning, R. Tizard, D. O. Haskard, C. D. Benjamin, I. Dougas, and T. Kirchhausen. Arrangement of domains, and amino acid residues required for binding of vascular cell adhesion molecule-1 to its counter-receptor VLA-4 (alpha 4 beta 1). Journal of Cell Biology, 124(4):601–??, February 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/124/4/601.

Opresko:1990:FRH

[OW90] L. K. Opresko and H. S. Wiley. Functional reconstitutional of the human epidermal growth factor receptor system in Xenopus oocytes. Journal of Cell Biology, 111(4):1661–??, October 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/111/4/1661.

Okagaki:1993:MMB

[OWF^{+}93] T. Okagaki, F. E. Weber, D. A. Fischman, K. T. Vaughan, T. Mikawa, and F. C. Reinach. The major myosin-binding domain of skeletal muscle MyBP–C (C protein) resides in the COOH-terminal, immunoglobulin C2 motif. Journal of Cell Biology, 123(3):619–??, November 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/123/3/619.

Panet:1991:SBS

[PA91] R. Panet and H. Atlan. Stimulation of bumetanide-sensitive Na+/K+/Cl-cotransport by different mitogens in synchronized human skin fibroblasts is essential for cell proliferation. Journal of Cell Biology, 114(2):337–??, July 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/114/2/
337.

Pante:1993:NPC

[PA93] N. Panté and U. Aebi. The nuclear pore complex. Journal of Cell Biology, 122(5):977–??, September 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/122/5/977.

Pietrini:1992:SMT

[PAC^{+}92] G. Pietrini, D. Aggujaro, P. Carrera, J. Malyszko, A. Vitale, and N. Borgese. A single mRNA, transcribed from an alternative, erythroid-specific, promoter, codes for two non-myristylated forms of NADH-cytochrome b5 reductase. Journal of Cell Biology, 117(5):975–??, June 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/117/5/975.

Prieto:1992:CMA

[PAFC92] A. L. Prieto, C. Andersson-Fisone, and K. L. Crossin. Characterization of multiple adhesive and counteradhesive domains in the extracellular matrix protein cytotactin. Journal of Cell Biology, 119(3):663–??, November 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/119/3/663.

Pisoni:1990:CSL

[PAL^{+}90] R. L. Pisoni, T. L. Acker, K. M. Lisowski, R. M. Lemons, and J. G. Thoene. A cysteine-specific lysosomal transport system provides a major route for the delivery of thiol to human fibroblast lysosomes: possible role in supporting lysosomal proteolysis. Journal of Cell Biology, 110(2):327–??, February 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/110/2/327.

Perrot-Applanat:1992:CCT

[PALM92] M. Perrot-Applanat, P. Lescop, and E. Milgrom. The cytoskeleton and the cellular traffic of the progesterone receptor. Journal of Cell Biology, 119(2):337–??, October 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/119/2/
337.

Powell:1990:IEI

[PB90] L. Powell and B. Burke. Internuclear exchange of an inner nuclear membrane protein (p55) in heterokaryons: in vivo evidence for the interaction of p55 with the nuclear lamina. Journal of Cell Biology, 111(6):2225–??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/111/6/
2225.

Pavalko:1991:DAC

[PB91] F. M. Pavalko and K. Burridge. Disruption of the actin cytoskeleton after microinjection of proteolytic fragments of alpha-actinin. Journal of Cell Biology, 114(3):481–??, August 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/114/3/481.

Peters:1991:PCD

[PBB^{+}91] L. L. Peters, C. S. Birkenmeier, R. T. Bronson, R. A. White, S. E. Lux, E. Otto, V. Bennett, A. Higgins, and J. E. Barker. Purkinje cell degeneration associated with erythroid ankyrin deficiency in nb/nb mice. Journal of Cell Biology, 114(6):1233–??, September 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/114/6/1233.

Perrone-Bizzozero:1993:PRG

[PBCK93] N. I. Perrone-Bizzozero, V. V. Cansino, and D. T. Kohn. Posttranscriptional regulation of GAP-43 gene expression in PC12 cells through protein kinase C-dependent stabilization of the mRNA. Journal of Cell Biology, 120(5):1263–??, March 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/120/5/1263.

Polans:1991:PCB

[PBCP91] A. S. Polans, J. Buczyłko, J. Crabb, and K. Palczewski. A photoreceptor calcium binding protein is recognized by autoantibodies obtained from patients with cancer-associated retinopathy. Journal of Cell Biology, 112(5):981–??, March 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/
112/5/981.

Peng:1993:RTP

[PBD93] H. B. Peng, L. P. Baker, and Z. Dai. A role of tyrosine phosphorylation in the formation of acetylcholine receptor clusters induced by electric fields in cultured Xenopus muscle cells. Journal of Cell Biology, 120(1):197–??, January 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/120/1/
197.

Pepperkok:1990:MSC

[PBDK90] R. Pepperkok, M. H. Bré, J. Davoust, and T. E. Kreis. Microtubules are stabilized in confluent epithelial cells but not in fibroblasts. Journal of Cell Biology, 111(6):3003– ??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/111/6/3003.

Pepper:1990:TGF

[PBM^{+}90] M. S. Pepper, D. Belin, R. Montesano, L. Orci, and J. D. Vassalli. Transforming growth factor-beta 1 modulates basic fibroblast growth factor-induced proteolytic and angiogenic properties of endothelial cells in vitro. Journal of Cell Biology, 111(2):743–??, August 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/111/2/743.

Pante:1994:ITD

[PBM^{+}94] N. Panté, R. Bastos, I. McMorrow, B. Burke, and U. Aebi. Interactions and three-dimensional localization of a group of nuclear pore complex proteins. Journal of Cell Biology, 126(3):603–??, August 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/126/3/603.

Paulson:1990:TSE

[PC90] H. L. Paulson and T. Claudio. Temperature-sensitive expression of all–Torpedo and Torpedo-rat hybrid AChR in mammalian muscle cells. Journal of Cell Biology, 110(5): 1705–??, May 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/110/5/1705.

Paddy:1991:SKP

[PC91a] M. R. Paddy and D. Chelsky. Spoke: a 120-kD protein associated with a novel filamentous structure on or near kinetochore microtubules in the mitotic spindle. Journal of Cell Biology, 113(1):161–??, April 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/113/1/161.

Payan:1991:SPF

[PC91b] L. A. Payan and K. Cline. A stromal protein factor maintains the solubility and insertion competence of an imported thylakoid membrane protein. Journal of Cell Biology, 112 (4):603–??, February 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/112/4/603.

Polizzi:1991:CSC

[PC91c] C. Polizzi and L. Clarke. The chromatin structure of centromeres from fission yeast: differentiation of the central core that correlates with function. Journal of Cell Biology, 112(2):191–??, January 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/112/2/191.

Parkos:1992:NMA

[PCD^{+}92] C. A. Parkos, S. P. Colgan, C. Delp, M. A. Arnaout, and J. L. Madara. Neutrophil migration across a cultured epithelial monolayer elicits a biphasic resistance response representing sequential effects on transcellular and paracellular pathways. Journal of Cell Biology, 117(4):757–??, May 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/117/4/757.

Pierceall:1994:NCE

[PCG^{+}94] W. E. Pierceall, K. R. Cho, R. H. Getzenberg, M. A. Reale, L. Hedrick, B. Vogelstein, and E. R. Fearon. NIH3T3 cells expressing the deleted in colorectal cancer tumor suppressor gene product stimulate neurite outgrowth in rat PC12 pheochromocytoma cells. Journal of Cell Biology, 124(6): 1017–??, March 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/124/6/1017.

Plutner:1991:RRV

[PCP^{+}91] H. Plutner, A. D. Cox, S. Pind, R. Khosravi-Far, J. R. Bourne, R. Schwaninger, C. J. Der, and W. E. Balch. Rab1b regulates vesicular transport between the endoplasmic reticulum and successive Golgi compartments. Journal of Cell Biology, 115(1):31–??, October 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/115/1/31.

Peeler:1993:ADA

[PDA93] J. S. Peeler, W. C. Donzell, and R. G. Anderson. The appendage domain of the AP-2 subunit is not required for assembly or invagination of clathrin-coated pits. Journal of Cell Biology, 120(1):47–??, January 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/120/1/47.

Parton:1991:PIM

[PDB^{+}91] R. G. Parton, C. G. Dotti, R. Bacallao, I. Kurtz, K. Simons, and K. Prydz. pH-induced microtubule-dependent redistribution of late endosomes in neuronal and epithelial cells. Journal of Cell Biology, 113(2):261–??, April 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/113/2/261.

Puoti:1991:BMS

[PDC91] A. Puoti, C. Desponds, and A. Conzelmann. Biosynthesis of mannosylinositolphosphoceramide in Saccharomyces cerevisiae is dependent on genes controlling the flow of secretory vesicles from the endoplasmic reticulum to the Golgi. Journal of Cell Biology, 113(3):515–??, May 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/113/3/515.

Prence:1990:MTL

[PDS90] E. M. Prence, J. M. Dong, and G. G. Sahagian. Modulation of the transport of a lysosomal enzyme by PDGF. Journal of Cell Biology, 110(2):319–??, February 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/110/2/319.

Plutner:1992:MAP

[PDSB92] H. Plutner, H. W. Davidson, J. Saraste, and W. E. Balch. Morphological analysis of protein transport from the ER to Golgi membranes in digitonin-permeabilized cells: role of the P58 containing compartment. Journal of Cell Biology, 119(5):1097–??, December 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/119/5/1097.

Pereira:1992:GAD

[PDTG92] A. Pereira, J. Doshen, E. Tanaka, and L. S. Goldstein. Genetic analysis of a Drosophila microtubule-associated protein. Journal of Cell Biology, 116(2):377–??, January 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/116/2/

Porter:1992:DCB

[PDWB92] G. A. Porter, G. M. Dmytrenko, J. C. Winkelmann, and R. J. Bloch. Dystrophin colocalizes with beta-spectrin in distinct subsarcolemmal domains in mammalian skeletal muscle. Journal of Cell Biology, 117(5):997–??, June 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/117/5/
997.

Powers:1993:ODH

[PE93] J. A. Powers and J. C. Eissenberg. Overlapping domains of the heterochromatin-associated protein HP1 mediate nuclear localization and heterochromatin binding. Journal of Cell Biology, 120(2):291–??, January 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/120/2/291.

Paul:1991:CNL

[PET^{+}91] D. L. Paul, L. Ebihara, L. J. Takemoto, K. I. Swenson, and D. A. Goodenough. Connexin46, a novel lens gap junction protein, induces voltage-gated currents in nonjunctional plasma membrane of Xenopus oocytes. Journal of Cell Biology, 115(4):1077–??, November 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/115/4/1077.

Probstmeier:1992:INC

[PFSS92] R. Probstmeier, T. Fahrig, E. Spiess, and M. Schachner. Interactions of the neural cell adhesion molecule and the myelin-associated glycoprotein with collagen type I: involvement in fibrillogenesis. Journal of Cell Biology, 116(4): 1063–??, February 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/116/4/1063.

Peculis:1992:LNP

[PG92] B. A. Peculis and J. G. Gall. Localization of the nucleolar protein NO38 in amphibian oocytes. Journal of Cell Biology, 116(1):1–??, January 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/116/1/1.

Pfanner:1990:FAP

[PGAR90] N. Pfanner, B. S. Glick, S. R. Arden, and J. E. Rothman. Fatty acylation promotes fusion of transport vesicles with Golgi cisternae. Journal of Cell Biology, 110(4):955–??, April 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/110/4/955.

Pines:1991:HCB

[PH91] J. Pines and T. Hunter. Human cyclins A and B1 are differentially located in the cell and undergo cell cycle-dependent nuclear transport. Journal of Cell Biology, 115(1):1–??, October 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/115/1/1.

Pittenger:1992:VVC

[PH92] M. F. Pittenger and D. M. Helfman. In vitro and in vivo characterization of four fibroblast tropomyosins produced in bacteria: TM-2, TM-3, TM-5a, and TM-5b are co-localized in interphase fibroblasts. Journal of Cell Biology, 118(4): 841–??, August 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/118/4/841.

Pasqualini:1994:CRI

[PH94a] R. Pasqualini and M. E. Hemler. Contrasting roles for integrin beta 1 and beta 5 cytoplasmic domains in subcellular localization, cell proliferation, and cell migration. Journal of Cell Biology, 125(2):447–??, April 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/125/2/447.

Pasqualone:1994:SSB

[PH94b] D. Pasqualone and T. C. Huffaker. STU1, a suppressor of a beta-tubulin mutation, encodes a novel and essential component of the yeast mitotic spindle. Journal of Cell Biology, 127(6):1973–??, December 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/127/6/1973.

Pieters:1991:ITL

[PHB^{+}91] J. Pieters, H. Horstmann, O. Bakke, G. Griffiths, and J. Lipp. Intracellular transport and localization of major histocompatibility complex class II molecules and associated invariant chain. Journal of Cell Biology, 115(5):1213– ??, December 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/115/5/1213.

Pancake:1992:MSC

[PHMH92] S. J. Pancake, G. D. Holt, S. Mellouk, and S. L. Hoffman. Malaria sporozoites and circumsporozoite proteins bind specifically to sulfated glycoconjugates. Journal of Cell Biology, 117(6):1351–??, June 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/117/6/1351.

Peach:1993:IHA

[PHSA93] R. J. Peach, D. Hollenbaugh, I. Stamenkovic, and A. Aruffo. Identification of hyaluronic acid binding sites in the extracellular domain of CD44. Journal of Cell Biology, 122 (1):257–??, July 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/122/1/257.

Prydz:1992:EBE

[PHSvD92] K. Prydz, S. H. Hansen, K. Sandvig, and B. van Deurs. Effects of brefeldin A on endocytosis, transcytosis and transport to the Golgi complex in polarized MDCK cells. Journal of Cell Biology, 119(2):259–??, October 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/119/2/259.

Porter:1992:DLR

[PHWM92] J. A. Porter, J. L. Hicks, D. S. Williams, and C. Montell. Differential localizations of and requirements for the two Drosophila ninaC kinase/myosins in photoreceptor cells. Journal of Cell Biology, 116(3):683–??, February 1992. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/116/3/

Pardi:1992:ARC

[PIRB92] R. Pardi, L. Inverardi, C. Rugarli, and J. R. Bender. Antigen-receptor complex stimulation triggers protein kinase C-dependent CD11a/CD18-cytoskeleton association in T lymphocytes. Journal of Cell Biology, 116(5):1211–??, March 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/116/5/1211.

Pimplikar:1994:BPT

[PIS94] S. W. Pimplikar, E. Ikonen, and K. Simons. Basolateral protein transport in streptolysin O-permeabilized MDCK cells. Journal of Cell Biology, 125(5):1025–??, June 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/125/5/
1025.

Prieto:1990:LDD

[PJC^{+}90] A. L. Prieto, F. S. Jones, B. A. Cunningham, K. L. Crossin, and G. M. Edelman. Localization during development of alternatively spliced forms of cytotactin mRNA by in situ hybridization. Journal of Cell Biology, 111(2):685–??, August 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/111/2/685.

Pennell:1992:ITC

[PJS^{+}92] R. I. Pennell, L. Janniche, G. N. Scofield, H. Booij, S. C. de Vries, and K. Roberts. Identification of a transitional cell state in the developmental pathway to carrot somatic embryogenesis. Journal of Cell Biology, 119(5):1371–??, December 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/119/5/1371.

Parton:1994:RIC

[PJS94a] R. G. Parton, B. Joggerst, and K. Simons. Regulated internalization of caveolae. Journal of Cell Biology, 127(5): 1199–??, December 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/127/5/1199.

Pruschy:1994:FNT

[PJS^{+}94b] M. Pruschy, Y. Ju, L. Spitz, E. Carafoli, and D. S. Goldfarb. Facilitated nuclear transport of calmodulin in tissue culture cells. Journal of Cell Biology, 127(6):1527–??, December 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/127/6/1527.

Phalen:1991:CRI

[PK91] T. Phalen and M. Kielian. Cholesterol is required for infection by Semliki Forest virus. Journal of Cell Biology, 112(4): 615–??, February 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/112/4/615.

Porter:1994:MSP

[PKG^{+}94] M. E. Porter, J. A. Knott, L. C. Gardner, D. R. Mitchell, and S. K. Dutcher. Mutations in the SUP–PF–1 locus of Chlamydomonas reinhardtii identify a regulatory domain in the beta-dynein heavy chain. Journal of Cell Biology, 126(6):1495–??, September 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/126/6/1495.

Pasdar:1991:RDA

[PKN91] M. Pasdar, K. A. Krzeminski, and W. J. Nelson. Regulation of desmosome assembly in MDCK epithelial cells: coordination of membrane core and cytoplasmic plaque domain assembly at the plasma membrane. Journal of Cell Biology, 113(3):645–??, May 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/113/3/645.

Phelps:1990:EPS

[PKPM90] B. M. Phelps, D. E. Koppel, P. Primakoff, and D. G. Myles. Evidence that proteolysis of the surface is an initial step in the mechanism of formation of sperm cell surface domains. Journal of Cell Biology, 111(5):1839–??, November 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/111/5/

Paterson:1990:CCI

[PL90] R. G. Paterson and R. A. Lamb. Conversion of a class II integral membrane protein into a soluble and efficiently secreted protein: multiple intracellular and extracellular oligomeric and conformational forms. Journal of Cell Biology, 110(4):999–??, April 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/110/4/999.

Picard:1991:OAM

[PLB^{+}91] A. Picard, J. C. Labbé, H. Barakat, J. C. Cavadore, and M. Dorée. Okadaic acid mimics a nuclear component required for cyclin B-cdc2 kinase microinjection to drive starfish oocytes into M phase. Journal of Cell Biology, 115(2):337–??, October 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/115/2/337.

Pavlin:1992:DUR

[PLB^{+}92] D. Pavlin, A. C. Lichtler, A. Bedalov, B. E. Kream, J. R. Harrison, H. F. Thomas, G. A. Gronowicz, S. H. Clark, C. O. Woody, and D. W. Rowe. Differential utilization of regulatory domains within the alpha 1(I) collagen promoter in osseous and fibroblastic cells. Journal of Cell Biology, 116(1):227–??, January 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/116/1/227.

Park:1991:IMG

[PLCB91] J. E. Park, J. M. Lopez, E. B. Cluett, and W. J. Brown. Identification of a membrane glycoprotein found primarily in the prelysosomal endosome compartment. Journal of Cell Biology, 112(2):245–??, January 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/112/2/245.

Paschke:1992:NCS

[PLS92] K. A. Paschke, F. Lottspeich, and C. A. Stuermer. Neurolin, a cell surface glycoprotein on growing retinal axons in the goldfish visual system, is reexpressed during retinal axonal regeneration. Journal of Cell Biology, 117(4):863– ??, May 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/117/4/863.

Porter:1993:DRD

[PM93] J. A. Porter and C. Montell. Distinct roles of the Drosophila ninaC kinase and myosin domains revealed by systematic mutagenesis. Journal of Cell Biology, 122(3):601–??, August 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/122/3/601.

Poruchynsky:1991:CDB

[PMA91] M. S. Poruchynsky, D. R. Maass, and P. H. Atkinson. Calcium depletion blocks the maturation of rotavirus by altering the oligomerization of virus-encoded proteins in the ER. Journal of Cell Biology, 114(4):651–??, August 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/114/4/
651.

Pelchen-Matthews:1992:PTK

[PMBL^{+}92] A. Pelchen-Matthews, I. Boulet, D. R. Littman, R. Fagard, and M. Marsh. The protein tyrosine kinase p56lck inhibits CD4 endocytosis by preventing entry of CD4 into coated pits. Journal of Cell Biology, 117(2):279–??, April 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/117/2/
279.

Presley:1993:EMC

[PMD^{+}93] J. F. Presley, S. Mayor, K. W. Dunn, L. S. Johnson, T. E. McGraw, and F. R. Maxfield. The End2 mutation in CHO cells slows the exit of transferrin receptors from the recycling compartment but bulk membrane recycling is unaffected. Journal of Cell Biology, 122(6):1231–??, September 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/122/6/
1231.

Perez-Morga:1993:SRK

[PME93] D. Pérez-Morga and P. T. Englund. The structure of replicating kinetoplast DNA networks. Journal of Cell Biology, 123(5):1069–??, December 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/123/5/1069.

Peifer:1992:VAJ

[PMG^{+}92] M. Peifer, P. D. McCrea, K. J. Green, E. Wieschaus, and B. M. Gumbiner. The vertebrate adhesive junction proteins beta-catenin and plakoglobin and the Drosophila segment polarity gene armadillo form a multigene family with similar properties. Journal of Cell Biology, 118(3):681–??, August 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/118/3/681.

Perrin:1992:CCM

[PMHS92] D. Perrin, K. Möller, K. Hanke, and H. D. Söling. cAMP and Ca(2+)-mediated secretion in parotid acinar cells is associated with reversible changes in the organization of the cytoskeleton. Journal of Cell Biology, 116(1):127–??, January 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/116/1/127.

Pagano:1991:NFC

[PMKH91] R. E. Pagano, O. C. Martin, H. C. Kang, and R. P. Haugland. A novel fluorescent ceramide analogue for studying membrane traffic in animal cells: accumulation at the Golgi apparatus results in altered spectral properties of the sphingolipid precursor. Journal of Cell Biology, 113(6):1267– ??, June 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/113/6/1267.

Piperno:1994:MDR

[PMLM94] G. Piperno, K. Mead, M. LeDizet, and A. Moscatelli. Mutations in the “dynein regulatory complex” alter the ATPinsensitive binding sites for inner arm dyneins in Chlamydomonas axonemes. Journal of Cell Biology, 125(5):1109– ??, June 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/125/5/1109.

Pytowski:1990:FCR

[PMM90] B. Pytowski, F. R. Maxfield, and J. Michl. Fc and C3bi receptors and the differentiation antigen BH2-Ag are randomly distributed in the plasma membrane of locomoting neutrophils. Journal of Cell Biology, 110(3):661–??, March 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/110/3/661.

Phillips:1991:MKP

[PMM91] W. D. Phillips, M. M. Maimone, and J. P. Merlie. Mutagenesis of the 43-kD postsynaptic protein defines domains involved in plasma membrane targeting and AChR clustering. Journal of Cell Biology, 115(6):1713–??, December 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/115/6/
1713.

Paschal:1992:HKC

[PMPV92] B. M. Paschal, A. Mikami, K. K. Pfister, and R. B. Vallee. Homology of the 74-kD cytoplasmic dynein subunit with a flagellar dynein polypeptide suggests an intracellular targeting function. Journal of Cell Biology, 118(5):1133– ??, September 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/118/5/1133.

Pypaert:1991:MCI

[PMS^{+}91] M. Pypaert, D. Mundy, E. Souter, J. C. Labbé, and G. Warren. Mitotic cytosol inhibits invagination of coated pits in broken mitotic cells. Journal of Cell Biology, 114(6):1159– ??, September 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/114/6/1159.

Piperno:1992:IDA

[PMS92] G. Piperno, K. Mead, and W. Shestak. The inner dynein arms I2 interact with a “dynein regulatory complex” in Chlamydomonas flagella. Journal of Cell Biology, 118(6): 1455–??, September 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/118/6/1455.

Plamann:1994:CDA

[PMTB94] M. Plamann, P. F. Minke, J. H. Tinsley, and K. S. Bruno. Cytoplasmic dynein and actin-related protein Arp1 are required for normal nuclear distribution in filamentous fungi.

Journal of Cell Biology, 127(1):139–??, October 1994. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/127/1/
139.

Postner:1992:MEM

[PMW92] M. A. Postner, K. G. Miller, and E. F. Wieschaus. Maternal effect mutations of the sponge locus affect actin cytoskeletal rearrangements in Drosophila melanogaster embryos. Journal of Cell Biology, 119(5):1205–??, December 1992. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/119/5/
1205.

Petruszak:1991:ECE

[PNB91] J. A. Petruszak, C. L. Nehme, and J. R. Bartles. Endoproteolytic cleavage in the extracellular domain of the integral plasma membrane protein CE9 precedes its redistribution from the posterior to the anterior tail of the rat spermatozoon during epididymal maturation. Journal of Cell Biology, 114(5):917–??, September 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/114/5/917.

Pind:1994:RCR

[PNM^{+}94] S. N. Pind, C. Nuoffer, J. M. McCaffery, H. Plutner, H. W. Davidson, M. G. Farquhar, and W. E. Balch. Rab1 and Ca^{2+} are required for the fusion of carrier vesicles mediating endoplasmic reticulum to Golgi transport. Journal of Cell Biology, 125(2):239–??, April 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/125/2/239.

Peter:1994:GND

[PNPB94] F. Peter, C. Nuoffer, S. N. Pind, and W. E. Balch. Guanine nucleotide dissociation inhibitor is essential for Rab1 function in budding from the endoplasmic reticulum and transport through the Golgi stack. Journal of Cell Biology, 126(6):1393–??, September 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/126/6/1393.

Phillips:1993:CIA

[PNR^{+}93] W. D. Phillips, P. G. Noakes, S. L. Roberds, K. P. Campbell, and J. P. Merlie. Clustering and immobilization of acetylcholine receptors by the 43-kD protein: a possible role for dystrophin-related protein. Journal of Cell Biology, 123(3): 729–??, November 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/123/3/729.

Pedraza:1990:MAG

[POGS90] L. Pedraza, G. C. Owens, L. A. Green, and J. L. Salzer. The myelin-associated glycoproteins: membrane disposition, evidence of a novel disulfide linkage between immunoglobulinlike domains, and posttranslational palmitylation. Journal of Cell Biology, 111(6):2651–??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/111/6/2651.

Persson:1991:FIT

[PP91] R. Persson and R. F. Pettersson. Formation and intracellular transport of a heterodimeric viral spike protein complex. Journal of Cell Biology, 112(2):257–??, January 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/112/2/
257.

Porter:1992:ESP

[PPD92] M. E. Porter, J. Power, and S. K. Dutcher. Extragenic suppressors of paralyzed flagellar mutations in Chlamydomonas reinhardtii identify loci that alter the inner dynein arms. Journal of Cell Biology, 118(5):1163–??, September 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/118/5/
1163.

Peters:1990:CAP

[PPFM90] D. M. Peters, L. M. Portz, J. Fullenwider, and D. F. Mosher. Co-assembly of plasma and cellular fibronectins into fibrils in human fibroblast cultures. Journal of Cell Biology, 111 (1):249–??, July 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/111/1/249.

Parvinen:1992:EBN

[PPHS^{+}92] M. Parvinen, M. Pelto-Huikko, O. Söder, R. Schultz, A. Kaipia, P. Mali, J. Toppari, H. Hakovirta, P. Lönnerberg, and E. M. Ritzén. Expression of beta-nerve growth factor and its receptor in rat seminiferous epithelium: specific function at the onset of meiosis. Journal of Cell Biology, 117 (3):629–??, May 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/117/3/629.

Prater:1991:PLT

[PPJF91] C. A. Prater, J. Plotkin, D. Jaye, and W. A. Frazier. The properdin-like type I repeats of human thrombospondin contain a cell attachment site. Journal of Cell Biology, 112 (5):1031–??, March 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/112/5/1031.

Pagano:1993:RCC

[PPL^{+}93] M. Pagano, R. Pepperkok, J. Lukas, V. Baldin, W. Ansorge, J. Bartek, and G. Draetta. Regulation of the cell cycle by the cdk2 protein kinase in cultured human fibroblasts. Journal of Cell Biology, 121(1):101–??, April 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/121/1/101.

Peter:1993:BCE

[PPZ^{+}93] F. Peter, H. Plutner, H. Zhu, T. E. Kreis, and W. E. Balch. Beta-COP is essential for transport of protein from the endoplasmic reticulum to the Golgi in vitro. Journal of Cell Biology, 122(6):1155–??, September 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/122/6/1155.

Piperno:1991:PPC

[PR91] G. Piperno and Z. Ramanis. The proximal portion of Chlamydomonas flagella contains a distinct set of inner dynein arms. Journal of Cell Biology, 112(4):701–??, February 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/112/4/701.

Podos:1994:LEB

[PRAK94] S. D. Podos, P. Reddy, J. Ashkenas, and M. Krieger. LDLC encodes a brefeldin A-sensitive, peripheral Golgi protein required for normal Golgi function. Journal of Cell Biology, 127(3):679–??, November 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/127/3/679.

Prie:1991:ASV

[PRB^{+}91] D. Prié, P. M. Ronco, B. Baudouin, M. Géniteau-Legendre, M. Antoine, R. Piedagnel, S. Estrade, B. Lelongt, P. J. Verroust, and R. Cassingéna. Activation of the simian virus 40 (SV40) genome abrogates sensitivity to AVP in a rabbit collecting tubule cell line by repressing membrane expression of AVP receptors. Journal of Cell Biology, 113 (4):951–??, May 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/113/4/951.

Pavlath:1994:TIT

[PRB94] G. K. Pavlath, T. A. Rando, and H. M. Blau. Transient immunosuppressive treatment leads to long-term retention of allogeneic myoblasts in hybrid myofibers. Journal of Cell Biology, 127(6):1923–??, December 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/127/6/1923.

Pohl:1990:DED

[PRC^{+}90] V. Pohl, P. P. Roger, D. Christophe, G. Pattyn, G. Vassart, and J. E. Dumont. Differentiation expression during proliferative activity induced through different pathways: in situ hybridization study of thyroglobulin gene expression in thyroid epithelial cells. Journal of Cell Biology, 111(2): 663–??, August 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/111/2/663.

Palmer:1993:STD

[PRF^{+}93] E. L. Palmer, C. Rüegg, R. Ferrando, R. Pytela, and D. Sheppard. Sequence and tissue distribution of the integrin alpha 9 subunit, a novel partner of beta 1 that is widely distributed in epithelia and muscle. Journal of Cell Biology, 123(5):1289–??, December 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/123/5/1289.

Paulson:1991:AEE

[PRGC91] H. L. Paulson, A. F. Ross, W. N. Green, and T. Claudio. Analysis of early events in acetylcholine receptor assembly. Journal of Cell Biology, 113(6):1371–??, June 1991. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/113/6/
1371.

Ponnambalam:1994:TGC

[PRL^{+}94] S. Ponnambalam, C. Rabouille, J. P. Luzio, T. Nilsson, and G. Warren. The TGN38 glycoprotein contains two nonoverlapping signals that mediate localization to the trans-Golgi network. Journal of Cell Biology, 125(2):253–??, April 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/125/2/253.

Piperno:1990:TDI

[PRSS90] G. Piperno, Z. Ramanis, E. F. Smith, and W. S. Sale. Three distinct inner dynein arms in Chlamydomonas flagella: molecular composition and location in the axoneme. Journal of Cell Biology, 110(2):379–??, February 1990. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/110/2/
379.

Patterson:1994:NIA

[PS94] S. I. Patterson and J. H. Skene. Novel inhibitory action of tunicamycin homologues suggests a role for dynamic protein fatty acylation in growth cone-mediated neurite extension. Journal of Cell Biology, 124(4):521–??, February 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/124/4/
521.

Parton:1992:ADE

[PSD92] R. G. Parton, K. Simons, and C. G. Dotti. Axonal and dendritic endocytic pathways in cultured neurons. Journal of Cell Biology, 119(1):123–??, October 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/119/1/123.

Pluthero:1990:PIE

[PSE^{+}90] F. G. Pluthero, M. Shreeve, D. Eskinazi, H. van der Gaag, K. S. Huang, J. D. Hulmes, M. Blum, and A. A. Axelrad. Purification of an inhibitor of erythroid progenitor cell cycling and antagonist to interleukin 3 from mouse marrow cell supernatants and its identification as cytosolic superoxide dismutase. Journal of Cell Biology, 111(3):1217– ??, September 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/111/3/1217.

Paterson:1990:MRP

[PSG^{+}90] H. F. Paterson, A. J. Self, M. D. Garrett, I. Just, K. Aktories, and A. Hall. Microinjection of recombinant p21rho induces rapid changes in cell morphology. Journal of Cell Biology, 111(3):1001–??, September 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/111/3/1001.

Pesavento:1994:CKK

[PSG94] P. A. Pesavento, R. J. Stewart, and L. S. Goldstein. Characterization of the KLP68D kinesin-like protein in Drosophila: possible roles in axonal transport. Journal of Cell Biology, 127(4):1041–??, November 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/127/4/1041.

Pluta:1992:ISC

[PSGE92] A. F. Pluta, N. Saitoh, I. Goldberg, and W. C. Earnshaw. Identification of a subdomain of CENP–B that is necessary and sufficient for localization to the human centromere. Journal of Cell Biology, 116(5):1081–??, March 1992. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/116/5/
1081.

Palmer:1992:RAM

[PSHK92] R. E. Palmer, D. S. Sullivan, T. Huffaker, and D. Koshland. Role of astral microtubules and actin in spindle orientation and migration in the budding yeast, Saccharomyces cerevisiae. Journal of Cell Biology, 119(3):583–??, November 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/119/3/583.

Pelton:1991:ILT

[PSJ^{+}91] R. W. Pelton, B. Saxena, M. Jones, H. L. Moses, and L. I. Gold. Immunohistochemical localization of TGF beta 1, TGF beta 2, and TGF beta 3 in the mouse embryo: expression patterns suggest multiple roles during embryonic development. Journal of Cell Biology, 115(4):1091– ??, November 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/115/4/1091.

Peter:1991:PDP

[PSN^{+}91] A. B. Peter, J. C. Schittny, V. Niggli, H. Reuter, and E. Sigel. The polarized distribution of poly(A+)-mRNA-induced functional ion channels in the Xenopus oocyte plasma membrane is prevented by anticytoskeletal drugs. Journal of Cell Biology, 114(3):455–??, August 1991. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/114/3/
455.

Page:1994:LKK

[PSRS94] B. D. Page, L. L. Satterwhite, M. D. Rose, and M. Snyder. Localization of the Kar3 kinesin heavy chain-related protein requires the Cik1 interacting protein. Journal of Cell Biology, 124(4):507–??, February 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/124/4/507.

Pepper:1993:UUR

[PSS^{+}93] M. S. Pepper, A. P. Sappino, R. Stöcklin, R. Montesano, L. Orci, and J. D. Vassalli. Upregulation of urokinase receptor expression on migrating endothelial cells. Journal of Cell Biology, 122(3):673–??, August 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/122/3/673.

Pryer:1993:CSC

[PSSK93] N. K. Pryer, N. R. Salama, R. Schekman, and C. A. Kaiser. Cytosolic Sec13p complex is required for vesicle formation from the endoplasmic reticulum in vitro. Journal of Cell Biology, 120(4):865–??, February 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/120/4/865.

Pan:1993:SEF

[PSZ^{+}93] T. C. Pan, T. Sasaki, R. Z. Zhang, R. Fässler, R. Timpl, and M. L. Chu. Structure and expression of fibulin-2, a novel extracellular matrix protein with multiple EGF-like repeats and consensus motifs for calcium binding. Journal of Cell Biology, 123(5):1269–??, December 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/123/5/1269.

Pagliaro:1992:DCD

[PT92] L. Pagliaro and D. L. Taylor. 2-deoxyglucose and cytochalasin D modulate aldolase mobility in living 3T3 cells. Journal of Cell Biology, 118(4):859–??, August 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/118/4/859.

Purdue:1990:IMA

[PTD90] P. E. Purdue, Y. Takada, and C. J. Danpure. Identification of mutations associated with peroxisome-to-mitochondrion mistargeting of alanine/glyoxylate aminotransferase in primary hyperoxaluria type 1. Journal of Cell Biology, 111(6): 2341–??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/111/6/2341.

Piper:1993:GNT

[PTK^{+}93] R. C. Piper, C. Tai, P. Kulesza, S. Pang, D. Warnock, J. Baenziger, J. W. Slot, H. J. Geuze, C. Puri, and D. E. James. GLUT-4 NH2 terminus contains a phenylalaninebased targeting motif that regulates intracellular sequestration. Journal of Cell Biology, 121(6):1221–??, June 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/121/6/
1221.

Pierres:1994:DAC

[PTMB94] A. Pierres, O. Tissot, B. Malissen, and P. Bongrand. Dynamic adhesion of CD8-positive cells to antibody-coated surfaces: the initial step is independent of microfilaments and intracellular domains of cell-binding molecules. Journal of Cell Biology, 125(4):945–??, May 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/125/4/945.

Piper:1992:EIS

[PTS^{+}92] R. C. Piper, C. Tai, J. W. Slot, C. S. Hahn, C. M. Rice, H. Huang, and D. E. James. The efficient intracellular sequestration of the insulin-regulatable glucose transporter (GLUT-4) is conferred by the NH2 terminus. Journal of Cell Biology, 117(4):729–??, May 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/117/4/729.

Pauli:1990:TSE

[PTTA90] D. Pauli, C. H. Tonka, A. Tissieres, and A. P. Arrigo. Tissue-specific expression of the heat shock protein HSP27 during Drosophila melanogaster development. Journal of Cell Biology, 111(3):817–??, September 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/111/3/817.

Payrastre:1991:PKD

[PvBeHB^{+}91] B. Payrastre, P. M. van Bergen en Henegouwen, M. Breton, J. C. den Hartigh, M. Plantavid, A. J. Verkleij, and J. Boonstra. Phosphoinositide kinase, diacylglycerol kinase, and phospholipase C activities associated to the cytoskeleton: effect of epidermal growth factor. Journal of Cell Biology, 115(1):121–??, October 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/115/1/121.

Peck:1993:DEE

[PW93] D. Peck and F. S. Walsh. Differential effects of overexpressed neural cell adhesion molecule isoforms on myoblast fusion. Journal of Cell Biology, 123(6):1587–??, December 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/123/6/1587.

Paralkar:1992:RHB

[PWY^{+}92] V. M. Paralkar, B. S. Weeks, Y. M. Yu, H. K. Kleinman, and A. H. Reddi. Recombinant human bone morphogenetic protein 2B stimulates PC12 cell differentiation: potentiation and binding to type IV collagen. Journal of Cell Biology, 119(6):1721–??, December 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/119/6/1721.

Persechini:1990:MCF

[PYA90] P. M. Persechini, J. D. Young, and W. Almers. Membrane channel formation by the lymphocyte pore-forming protein: comparison between susceptible and resistant target cells. Journal of Cell Biology, 110(6):2109–??, June 1990. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/110/6/
2109.

Pathak:1990:TSS

[PYH^{+}90] R. K. Pathak, M. Yokode, R. E. Hammer, S. L. Hofmann, M. S. Brown, J. L. Goldstein, and R. G. Anderson. Tissuespecific sorting of the human LDL receptor in polarized epithelia of transgenic mice. Journal of Cell Biology, 111(2): 347–??, August 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/111/2/347.

Peterson:1994:IBB

[PZB^{+}94] J. Peterson, Y. Zheng, L. Bender, A. Myers, R. Cerione, and A. Bender. Interactions between the bud emergence proteins Bem1p and Bem2p and Rho-type GTPases in yeast. Journal of Cell Biology, 127(5):1395–??, December 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/127/5/1395.

Patel:1991:ORI

[PZP^{+}91] K. D. Patel, G. A. Zimmerman, S. M. Prescott, R. P. McEver, and T. M. McIntyre. Oxygen radicals induce human endothelial cells to express GMP-140 and bind neutrophils. Journal of Cell Biology, 112(4):749–??, February 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/112/4/749.

Pagani:1991:TSS

[PZV^{+}91] F. Pagani, L. Zagato, C. Vergani, G. Casari, A. Sidoli, and F. E. Baralle. Tissue-specific splicing pattern of fibronectin messenger RNA precursor during development and aging in rat. Journal of Cell Biology, 113(5):1223–??, June 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/113/5/
1223.

Quarto:1992:THI

[QCCD92] R. Quarto, G. Campanile, R. Cancedda, and B. Dozin. Thyroid hormone, insulin, and glucocorticoids are sufficient to support chondrocyte differentiation to hypertrophy: a serum-free analysis. Journal of Cell Biology, 119(4):989– ??, November 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/119/4/989.

Quarto:1990:VDH

[QDT^{+}90] R. Quarto, B. Dozin, C. Tacchetti, G. Campanile, C. Malfatto, and R. Cancedda. In vitro development of hypertrophic chondrocytes starting from selected clones of dedifferentiated cells. Journal of Cell Biology, 110(4):1379–??, April 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/110/4/1379.

Quarmby:1994:TDC

[QH94] L. M. Quarmby and H. C. Hartzell. Two distinct, calciummediated, signal transduction pathways can trigger deflagellation in Chlamydomonas reinhardtii. Journal of Cell Biology, 124(5):807–??, March 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/124/5/807.

Quinn:1991:ITS

[QORM91] D. Quinn, L. Orci, M. Ravazzola, and H. P. Moore. Intracellular transport and sorting of mutant human proinsulins that fail to form hexamers. Journal of Cell Biology, 113 (5):987–??, June 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/113/5/987.

Qiu:1991:RIR

[QPWG91] M. S. Qiu, A. F. Pitts, T. R. Winters, and S. H. Green. ras isoprenylation is required for ras-induced but not for NGF-induced neuronal differentiation of PC12 cells. Journal of Cell Biology, 115(3):795–??, November 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/115/3/795.

Quax:1991:MBH

[QvMWV^{+}91] P. H. Quax, G. N. van Muijen, E. J. Weening-Verhoeff, L. R. Lund, K. Danø, D. J. Ruiter, and J. H. Verheijen. Metastatic behavior of human melanoma cell lines in nude mice correlates with urokinase-type plasminogen activator, its type-1 inhibitor, and urokinase-mediated matrix degradation. Journal of Cell Biology, 115(1):191–??, October 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/115/1/191.

Qiu:1994:SSC

[QXWN^{+}94] Y. Qiu, X. Xu, A. Wandinger-Ness, D. P. Dalke, and S. K. Pierce. Separation of subcellular compartments containing distinct functional forms of MHC class II. Journal of Cell Biology, 125(3):595–??, May 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/125/3/595.

Quarmby:1992:IPM

[QYC^{+}92] L. M. Quarmby, Y. G. Yueh, J. L. Cheshire, L. R. Keller, W. J. Snell, and R. C. Crain. Inositol phospholipid metabolism may trigger flagellar excision in Chlamydomonas reinhardtii. Journal of Cell Biology, 116(3):737– ??, February 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/116/3/737.

Rieder:1990:KTP

[RA90] C. L. Rieder and S. P. Alexander. Kinetochores are transported poleward along a single astral microtubule during chromosome attachment to the spindle in newt lung cells. Journal of Cell Biology, 110(1):81–??, January 1990. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/110/1/
81.

Rousselle:1994:KME

[RA94] P. Rousselle and M. Aumailley. Kalinin is more efficient than laminin in promoting adhesion of primary keratinocytes and some other epithelial cells and has a different requirement for integrin receptors. Journal of Cell Biology, 125(1):205–??, April 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/125/1/205.

Riberdy:1994:TID

[RAGC94] J. M. Riberdy, R. R. Avva, H. J. Geuze, and P. Cresswell. Transport and intracellular distribution of MHC class II molecules and associated invariant chain in normal and antigen-processing mutant cell lines. Journal of Cell Biology, 125(6):1225–??, June 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/125/6/1225.

Ralston:1993:CAG

[Ral93] E. Ralston. Changes in architecture of the Golgi complex and other subcellular organelles during myogenesis. Journal of Cell Biology, 120(2):399–??, January 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/120/2/399.

Rotzler:1990:MSA

[RB90] S. Rotzler and H. R. Brenner. Metabolic stabilization of acetylcholine receptors in vertebrate neuromuscular junction by muscle activity. Journal of Cell Biology, 111(2): 655–??, August 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/111/2/655.

Reaves:1992:PMT

[RB92] B. Reaves and G. Banting. Perturbation of the morphology of the trans-Golgi network following Brefeldin A treatment: redistribution of a TGN-specific integral membrane protein, TGN38. Journal of Cell Biology, 116(1):85–??, January 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/116/1/85.

Rout:1993:IYN

[RB93] M. P. Rout and G. Blobel. Isolation of the yeast nuclear pore complex. Journal of Cell Biology, 123(4):771– ??, November 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/123/4/771.

Rando:1994:PMM

[RB94] T. A. Rando and H. M. Blau. Primary mouse myoblast purification, characterization, and transplantation for cellmediated gene therapy. Journal of Cell Biology, 125(6): 1275–??, June 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/125/6/1275.

Rizzuto:1994:MCH

[RBB^{+}94] R. Rizzuto, C. Bastianutto, M. Brini, M. Murgia, and T. Pozzan. Mitochondrial Ca^{2+} homeostasis in intact cells. Journal of Cell Biology, 126(5):1183–??, September 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/126/5/
1183.

Roger:1992:MGH

[RBD92] P. P. Roger, M. Baptist, and J. E. Dumont. A mechanism generating heterogeneity in thyroid epithelial cells: suppression of the thyrotropin/cAMP-dependent mitogenic pathway after cell division induced by cAMP-independent factors. Journal of Cell Biology, 117(2):383–??, April 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/117/2/
383.

Richardson:1990:IKP

[RBR90] G. P. Richardson, S. Bartolami, and I. J. Russell. Identification of a 275-kD protein associated with the apical surfaces of sensory hair cells in the avian inner ear. Journal of Cell Biology, 110(4):1055–??, April 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/110/4/1055.

Radu:1993:NNN

[RBW93] A. Radu, G. Blobel, and R. W. Wozniak. Nup155 is a novel nuclear pore complex protein that contains neither repetitive sequence motifs nor reacts with WGA. Journal of Cell Biology, 121(1):1–??, April 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/121/1/1.

Renz:1994:SRA

[RCJ^{+}94] M. E. Renz, H. H. Chiu, S. Jones, J. Fox, K. J. Kim, L. G. Presta, and S. Fong. Structural requirements for adhesion of soluble recombinant murine vascular cell adhesion molecule-1 to alpha 4 beta 1. Journal of Cell Biology, 125 (6):1395–??, June 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/125/6/1395.

Ruscetti:1994:CHC

[RCNO94] T. Ruscetti, J. A. Cardelli, M. L. Niswonger, and T. J. O’Halloran. Clathrin heavy chain functions in sorting and secretion of lysosomal enzymes in Dictyostelium discoideum. Journal of Cell Biology, 126(2):343–??, July 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/126/2/
343.

Roubin:1990:MNE

[RDBHG90] R. Roubin, H. Deagostini-Bazin, M. R. Hirsch, and C. Goridis. Modulation of NCAM expression by transforming growth factor-beta, serum, and autocrine factors. Journal of Cell Biology, 111(2):673–??, August 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/111/2/673.

Ren:1993:RTS

[RDDR93] M. Ren, G. Drivas, P. D’Eustachio, and M. G. Rush. Ran/ TC4: a small nuclear GTP-binding protein that regulates DNA synthesis. Journal of Cell Biology, 120(2):313–??, January 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/120/2/313.

Ridgway:1992:TOB

[RDH^{+}92] N. D. Ridgway, P. A. Dawson, Y. K. Ho, M. S. Brown, and J. L. Goldstein. Translocation of oxysterol binding protein to Golgi apparatus triggered by ligand binding. Journal of Cell Biology, 116(2):307–??, January 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/116/2/307.

Reed:1994:BRP

[Ree94] J. C. Reed. Bcl-2 and the regulation of programmed cell death. Journal of Cell Biology, 124(1):1–??, January 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/124/1/
1.

Rochette-Egly:1991:RAR

[RELS^{+}91] C. Rochette-Egly, Y. Lutz, M. Saunders, I. Scheuer, M. P. Gaub, and P. Chambon. Retinoic acid receptor gamma: specific immunodetection and phosphorylation. Journal of Cell Biology, 115(2):535–??, October 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/115/2/535.

Risau:1990:IFB

[REW90] W. Risau, B. Engelhardt, and H. Wekerle. Immune function of the blood-brain barrier: incomplete presentation of protein (auto-)antigens by rat brain microvascular endothelium in vitro. Journal of Cell Biology, 110(5):1757–??, May 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/110/5/
1757.

Roulier:1992:PDA

[RFF92] E. M. Roulier, C. Fyrberg, and E. Fyrberg. Perturbations of Drosophila alpha-actinin cause muscle paralysis, weakness, and atrophy but do not confer obvious nonmuscle phenotypes. Journal of Cell Biology, 116(4):911–??, February 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/116/4/911.

Rapiejko:1992:PTA

[RG92] P. J. Rapiejko and R. Gilmore. Protein translocation across the ER requires a functional GTP binding site in the alpha subunit of the signal recognition particle receptor. Journal of Cell Biology, 117(3):493–??, May 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/117/3/493.

Robinson:1994:CDR

[RG94] D. R. Robinson and K. Gull. The configuration of DNA replication sites within the Trypanosoma brucei kinetoplast. Journal of Cell Biology, 126(3):641–??, August 1994. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/126/3/
641.

Ross:1991:EAR

[RGHC91] A. F. Ross, W. N. Green, D. S. Hartman, and T. Claudio. Efficiency of acetylcholine receptor subunit assembly and its regulation by cAMP. Journal of Cell Biology, 113 (3):623–??, May 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/113/3/623.

Riddelle:1991:FHV

[RGJ91] K. S. Riddelle, K. J. Green, and J. C. Jones. Formation of hemidesmosomes in vitro by a transformed rat bladder cell line. Journal of Cell Biology, 112(1):159–??, January 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/112/1/
159.

Risek:1990:MGJ

[RGKG90] B. Risek, S. Guthrie, N. Kumar, and N. B. Gilula. Modulation of gap junction transcript and protein expression during pregnancy in the rat. Journal of Cell Biology, 110(2): 269–??, February 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/110/2/269.

Roberts:1990:ICF

[RGR90] D. M. Roberts, M. Guenthert, and R. Rodewald. Isolation and characterization of the Fc receptor from the fetal yolk sac of the rat. Journal of Cell Biology, 111(5):1867– ??, November 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/111/5/1867.

Rodionov:1993:MDC

[RGT^{+}93] V. I. Rodionov, F. K. Gyoeva, E. Tanaka, A. D. Bershadsky, J. M. Vasiliev, and V. I. Gelfand. Microtubule-dependent control of cell shape and pseudopodial activity is inhibited by the antibody to kinesin motor domain. Journal of Cell Biology, 123(6):1811–??, December 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/123/6/1811.

Ralston:1992:RDM

[RH92a] E. Ralston and Z. W. Hall. Restricted distribution of mRNA produced from a single nucleus in hybrid myotubes. Journal of Cell Biology, 119(5):1063–??, December 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/119/5/1063.

Regen:1992:DBI

[RH92b] C. M. Regen and A. F. Horwitz. Dynamics of beta 1 integrin-mediated adhesive contacts in motile fibroblasts. Journal of Cell Biology, 119(5):1347–??, December 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/119/5/
1347.

Reichelt:1990:CBS

[RHB^{+}90] R. Reichelt, A. Holzenburg, E. L. Buhle, M. Jarnik, A. Engel, and U. Aebi. Correlation between structure and mass distribution of the nuclear pore complex and of distinct pore complex components. Journal of Cell Biology, 110(4):883– ??, April 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/110/4/883.

Ruhl:1993:EIF

[RHB^{+}93] M. Ruhl, M. Himmelspach, G. M. Bahr, F. Hammerschmid, H. Jaksche, B. Wolff, H. Aschauer, G. K. Farrington, H. Probst, and D. Bevec. Eukaryotic initiation factor 5A is a cellular target of the human immunodeficiency virus type 1 Rev activation domain mediating trans-activation. Journal of Cell Biology, 123(6):1309–??, December 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/123/6/1309.

Redding:1991:IKP

[RHF91] K. Redding, C. Holcomb, and R. S. Fuller. Immunolocalization of Kex2 protease identifies a putative late Golgi compartment in the yeast Saccharomyces cerevisiae. Journal of Cell Biology, 113(3):527–??, May 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/113/3/527.

Roth:1990:MEC

[RHG90] K. A. Roth, J. M. Hertz, and J. I. Gordon. Mapping enteroendocrine cell populations in transgenic mice reveals an unexpected degree of complexity in cellular differentiation within the gastrointestinal tract. Journal of Cell Biology, 110(5):1791–??, May 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/110/5/1791.

Rabinowitz:1992:ICE

[RHGG92] S. Rabinowitz, H. Horstmann, S. Gordon, and G. Griffiths. Immunocytochemical characterization of the endocytic and phagolysosomal compartments in peritoneal macrophages. Journal of Cell Biology, 116(1):95–??, January 1992. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/116/1/
95.

Reszka:1992:IAA

[RHH92] A. A. Reszka, Y. Hayashi, and A. F. Horwitz. Identification of amino acid sequences in the integrin beta 1 cytoplasmic domain implicated in cytoskeletal association. Journal of Cell Biology, 117(6):1321–??, June 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/117/6/1321.

Rubin:1991:CCM

[RHP^{+}91] L. L. Rubin, D. E. Hall, S. Porter, K. Barbu, C. Cannon, H. C. Horner, M. Janatpour, C. W. Liaw, K. Manning, and J. Morales. A cell culture model of the blood-brain barrier.

Journal of Cell Biology, 115(6):1725–??, December 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/115/6/
1725.

Roff:1990:RFC

[RHR90] C. F. Roff, C. W. Hall, and A. R. Robbins. Recovery of function in Chinese hamster ovary cell mutants with temperature-sensitive defects in vacuolar acidification. Journal of Cell Biology, 110(4):1023–??, April 1990. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/110/4/
1023.

Rosen:1994:RSF

[RJJ^{+}94] E. M. Rosen, A. Joseph, L. Jin, S. Rockwell, J. A. Elias, J. Knesel, J. Wines, J. McClellan, M. J. Kluger, and I. D. Goldberg. Regulation of scatter factor production via a soluble inducing factor. Journal of Cell Biology, 127(1): 225–??, October 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/127/1/225.

Rickard:1990:INN

[RK90a] J. E. Rickard and T. E. Kreis. Identification of a novel nucleotide-sensitive microtubule-binding protein in HeLa cells. Journal of Cell Biology, 110(5):1623–??, May 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/110/5/
1623.

Rout:1990:CYS

[RK90b] M. P. Rout and J. V. Kilmartin. Components of the yeast spindle and spindle pole body. Journal of Cell Biology, 111(5):1913–??, November 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/111/5/1913.

Reinsch:1994:OSA

[RK94] S. Reinsch and E. Karsenti. Orientation of spindle axis and distribution of plasma membrane proteins during cell division in polarized MDCKII cells. Journal of Cell Biology, 126(6):1509–??, September 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/126/6/1509.

Raff:1993:DGT

[RKA93] J. W. Raff, D. R. Kellogg, and B. M. Alberts. Drosophila gamma-tubulin is part of a complex containing two previously identified centrosomal MAPs. Journal of Cell Biology, 121(4):823–??, May 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/121/4/823.

Rimm:1990:IFR

[RKB^{+}90] D. L. Rimm, D. A. Kaiser, D. Bhandari, P. Maupin, D. P. Kiehart, and T. D. Pollard. Identification of functional regions on the tail of Acanthamoeba myosin–II using recombinant fusion proteins. I. High resolution epitope mapping and characterization of monoclonal antibody binding sites. Journal of Cell Biology, 111(6):2405–??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/111/6/
2405.

Ruppert:1993:ICC

[RKB93] C. Ruppert, R. Kroschewski, and M. Bähler. Identification, characterization and cloning of myr 1, a mammalian myosin–I. Journal of Cell Biology, 120(6):1393–??, March 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/120/6/1393.

Rafiee:1991:SCL

[RLB^{+}91] P. Rafiee, H. Leffler, J. C. Byrd, F. J. Cassels, E. C. Boedeker, and Y. S. Kim. A sialoglycoprotein complex linked to the microvillus cytoskeleton acts as a receptor for pilus (AF/R1) mediated adhesion of enteropathogenic Escherichia coli (RDEC-1) in rabbit small intestine. Journal of Cell Biology, 115(4):1021–??, November 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/115/4/1021.

Rawson:1991:DSF

[RLD^{+}91] C. L. Rawson, D. T. Loo, J. R. Duimstra, O. R. Hedstrom, E. E. Schmidt, and D. W. Barnes. Death of serum-free mouse embryo cells caused by epidermal growth factor deprivation. Journal of Cell Biology, 113(3):671–??, May 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/113/3/
671.

Rodionov:1994:MDF

[RLGB94] V. I. Rodionov, S. S. Lim, V. I. Gelfand, and G. G. Borisy. Microtubule dynamics in fish melanophores. Journal of Cell Biology, 126(6):1455–??, September 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/126/6/1455.

Rousselle:1991:KES

[RLKB91] P. Rousselle, G. P. Lunstrum, D. R. Keene, and R. E. Burgeson. Kalinin: an epithelium-specific basement membrane adhesion molecule that is a component of anchoring filaments. Journal of Cell Biology, 114(3):567–??, August 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/114/3/
567.

Roberts:1992:EPM

[RLMG92] S. J. Roberts, D. S. Leaf, H. P. Moore, and J. C. Gerhart. The establishment of polarized membrane traffic in Xenopus laevis embryos. Journal of Cell Biology, 118(6):1359– ??, September 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/118/6/1359.

Ravid:1993:TEC

[RLRR93] K. Ravid, Y. C. Li, H. B. Rayburn, and R. D. Rosenberg. Targeted expression of a conditional oncogene in hematopoietic cells of transgenic mice. Journal of Cell Biology, 123(6): 1545–??, December 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/123/6/1545.

Rosen:1992:EUS

[RLS92] C. L. Rosen, M. P. Lisanti, and J. L. Salzer. Expression of unique sets of GPI-linked proteins by different primary neurons in vitro. Journal of Cell Biology, 117(3):617–??, May 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/117/3/617.

Rexach:1994:CER

[RLS94] M. F. Rexach, M. Latterich, and R. W. Schekman. Characteristics of endoplasmic reticulum-derived transport vesicles. Journal of Cell Biology, 126(5):1133–??, September 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/126/5/1133.

Ruetz:1993:FAP

[RLWK93] S. Ruetz, A. E. Lindsey, C. L. Ward, and R. R. Kopito. Functional activation of plasma membrane anion exchangers occurs in a pre-Golgi compartment. Journal of Cell Biology, 121(1):37–??, April 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/121/1/37.

Romanic:1994:IKG

[RM94] A. M. Romanic and J. A. Madri. The induction of 72kD gelatinase in T cells upon adhesion to endothelial cells is VCAM-1 dependent. Journal of Cell Biology, 125(5): 1165–??, June 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/125/5/1165.

Rosiere:1990:KPM

[RMB90] T. K. Rosiere, J. A. Marrs, and G. B. Bouck. A 39-kD plasma membrane protein (IP39) is an anchor for the unusual membrane skeleton of Euglena gracilis. Journal of Cell Biology, 110(4):1077–??, April 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/110/4/1077.

Rose:1992:ECF

[RMFA92] D. W. Rose, G. McCabe, J. R. Feramisco, and M. Adler. Expression of c-fos and AP-1 activity in senescent human fibroblasts is not sufficient for DNA synthesis. Journal of Cell Biology, 119(6):1405–??, December 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/119/6/1405.

Roth:1990:MAR

[RMG90] M. B. Roth, C. Murphy, and J. G. Gall. A monoclonal antibody that recognizes a phosphorylated epitope stains lampbrush chromosome loops and small granules in the amphibian germinal vesicle. Journal of Cell Biology, 111(6):2217– ??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/111/6/2217.

Reinsch:1991:MPA

[RMK91] S. S. Reinsch, T. J. Mitchison, and M. Kirschner. Microtubule polymer assembly and transport during axonal elongation. Journal of Cell Biology, 115(2):365–??, October 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/115/2/365.

Rassow:1994:MPI

[RMK^{+}94] J. Rassow, A. C. Maarse, E. Krainer, M. Kübrich, H. Müller, M. Meijer, E. A. Craig, and N. Pfanner. Mitochondrial protein import: biochemical and genetic evidence for interaction of matrix hsp70 and the inner membrane protein MIM44. Journal of Cell Biology, 127(6):1547–??, December 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/127/6/1547.

Ray:1993:KFM

[RMMH93] S. Ray, E. Meyhöfer, R. A. Milligan, and J. Howard. Kinesin follows the microtubule’s protofilament axis. Journal of Cell Biology, 121(5):1083–??, June 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/121/5/1083.

Roof:1992:KRP

[RMR92] D. M. Roof, P. B. Meluh, and M. D. Rose. Kinesin-related proteins required for assembly of the mitotic spindle. Journal of Cell Biology, 118(1):95–??, July 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/118/1/95.

Rosen:1994:SFC

[RNG94] E. M. Rosen, S. K. Nigam, and I. D. Goldberg. Scatter factor and the c-met receptor: a paradigm for mesenchymal/ epithelial interaction. Journal of Cell Biology, 127(6):1783– ??, December 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/127/6/1783.

Roberts:1992:MPS

[RNS92] C. J. Roberts, S. F. Nothwehr, and T. H. Stevens. Membrane protein sorting in the yeast secretory pathway: evidence that the vacuole may be the default compartment. Journal of Cell Biology, 119(1):69–??, October 1992. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/119/1/
69.

Robinson:1990:CEG

[Rob90] M. S. Robinson. Cloning and expression of gamma-adaptin, a component of clathrin-coated vesicles associated with the Golgi apparatus. Journal of Cell Biology, 111(6):2319–??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/111/6/2319.

Robinson:1993:ATA

[Rob93] M. S. Robinson. Assembly and targeting of adaptin chimeras in transfected cells. Journal of Cell Biology, 123 (1):67–??, October 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/123/1/67.

Roitelman:1992:IEE

[ROBN^{+}92] J. Roitelman, E. H. Olender, S. Bar-Nun, W. A. Dunn, and R. D. Simoni. Immunological evidence for eight spans in the membrane domain of 3-hydroxy-3-methylglutaryl coenzyme A reductase: implications for enzyme degradation in the endoplasmic reticulum. Journal of Cell Biology, 117 (5):959–??, June 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/117/5/959.

Raymond:1990:MAY

[ROE^{+}90] C. K. Raymond, P. J. O’Hara, G. Eichinger, J. H. Rothman, and T. H. Stevens. Molecular analysis of the yeast VPS3 gene and the role of its product in vacuolar protein sorting and vacuolar segregation during the cell cycle. Journal of Cell Biology, 111(3):877–??, September 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/111/3/877.

Rotundo:1990:NST

[Rot90] R. L. Rotundo. Nucleus-specific translation and assembly of acetylcholinesterase in multinucleated muscle cells. Journal of Cell Biology, 110(3):715–??, March 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/110/3/715.

Raats:1990:AAT

[RPE^{+}90] J. M. Raats, F. R. Pieper, W. T. Vree Egberts, K. N. Verrijp, F. C. Ramaekers, and H. Bloemendal. Assembly of amino-terminally deleted desmin in vimentin-free cells. Journal of Cell Biology, 111(5):1971–??, November 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/111/5/
1971.

Robinson:1992:TGT

[RPH^{+}92] L. J. Robinson, S. Pang, D. S. Harris, J. Heuser, and D. E. James. Translocation of the glucose transporter (GLUT4) to the cell surface in permeabilized 3T3-L1 adipocytes: effects of ATP insulin, and GTP gamma S and localization of GLUT4 to clathrin lattices. Journal of Cell Biology, 117 (6):1181–??, June 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/117/6/1181.

Roemer:1994:CYG

[RPPB94] T. Roemer, G. Paravicini, M. A. Payton, and H. Bussey. Characterization of the yeast (1 → 6)-beta-glucan biosynthetic components, Kre6p and Skn1p, and genetic interactions between the PKC1 pathway and extracellular matrix assembly. Journal of Cell Biology, 127(2):567–??, October 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/127/2/567.

Ruegg:1992:RIA

[RPS^{+}92] C. Rüegg, A. A. Postigo, E. E. Sikorski, E. C. Butcher, R. Pytela, and D. J. Erle. Role of integrin alpha 4 beta 7/alpha 4 beta P in lymphocyte adherence to fibronectin and VCAM-1 and in homotypic cell clustering. Journal of Cell Biology, 117(1):179–??, April 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/117/1/179.

Raines:1992:CPE

[RR92a] E. W. Raines and R. Ross. Compartmentalization of PDGF on extracellular binding sites dependent on exon-6-encoded sequences. Journal of Cell Biology, 116(2):533–??, January 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/116/2/533.

Rossi:1992:CSA

[RR92b] S. G. Rossi and R. L. Rotundo. Cell surface acetylcholinesterase molecules on multinucleated myotubes are clustered over the nucleus of origin. Journal of Cell Biology, 119(6):1657–??, December 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/119/6/1657.

Rubin:1991:ECD

[RRBG91] D. C. Rubin, K. A. Roth, E. H. Birkenmeier, and J. I. Gordon. Epithelial cell differentiation in normal and transgenic mouse intestinal isografts. Journal of Cell Biology, 113(5): 1183–??, June 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/113/5/1183.

Raths:1993:EET

[RRCR93] S. Raths, J. Rohrer, F. Crausaz, and H. Riezman. end3 and end4: two mutants defective in receptor-mediated and fluid-phase endocytosis in Saccharomyces cerevisiae. Journal of Cell Biology, 120(1):55–??, January 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/120/1/55.

Robinson:1993:TBS

[RRD93] S. D. Robinson, A. B. Roberts, and C. W. Daniel. TGF beta suppresses casein synthesis in mouse mammary explants and may play a role in controlling milk levels during pregnancy. Journal of Cell Biology, 120(1):245–??, January 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/120/1/245.

Radermacher:1994:CEM

[RRG^{+}94] M. Radermacher, V. Rao, R. Grassucci, J. Frank, A. P. Timerman, S. Fleischer, and T. Wagenknecht. Cryoelectron microscopy and three-dimensional reconstruction of the calcium release channel/ryanodine receptor from skeletal muscle. Journal of Cell Biology, 127(2):411–??, October 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/127/2/411.

Rogers:1991:ILG

[RRH91] S. A. Rogers, G. Ryan, and M. R. Hammerman. Insulin-like growth factors I and II are produced in the metanephros and are required for growth and development in vitro. Journal of Cell Biology, 113(6):1447–??, June 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/113/6/1447.

Rexach:1991:DBR

[RS91] M. F. Rexach and R. W. Schekman. Distinct biochemical requirements for the budding, targeting, and fusion of ER-derived transport vesicles. Journal of Cell Biology, 114 (2):219–??, July 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/114/2/219.

Racoosin:1993:MMF

[RS93] E. L. Racoosin and J. A. Swanson. Macropinosome maturation and fusion with tubular lysosomes in macrophages. Journal of Cell Biology, 121(5):1011–??, June 1993. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/121/5/

Rieder:1994:MKP

[RS94] C. L. Rieder and E. D. Salmon. Motile kinetochores and polar ejection forces dictate chromosome position on the vertebrate mitotic spindle. Journal of Cell Biology, 124(3): 223–??, February 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/124/3/223.

Regauer:1990:EOC

[RSB^{+}90] S. Regauer, G. R. Seiler, Y. Barrandon, K. W. Easley, and C. C. Compton. Epithelial origin of cutaneous anchoring fibrils. Journal of Cell Biology, 111(5):2109–??, November 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/111/5/2109.

Rabouille:1993:DDT

[RSC^{+}93] C. Rabouille, G. J. Strous, J. D. Crapo, H. J. Geuze, and J. W. Slot. The differential degradation of two cytosolic proteins as a tool to monitor autophagy in hepatocytes by immunocytochemistry. Journal of Cell Biology, 120(4):897– ??, February 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/120/4/897.

Rieder:1994:AOV

[RSCS94] C. L. Rieder, A. Schultz, R. Cole, and G. Sluder. Anaphase onset in vertebrate somatic cells is controlled by a checkpoint that monitors sister kinetochore attachment to the spindle. Journal of Cell Biology, 127(5):1301–??, December 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/127/5/1301.

Reponen:1994:HEK

[RSM^{+}94] P. Reponen, C. Sahlberg, C. Munaut, I. Thesleff, and K. Tryggvason. High expression of 92-kD type IV collagenase (gelatinase B) in the osteoclast lineage during mouse development. Journal of Cell Biology, 124(6):1091– ??, March 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/124/6/1091.

Ruknudin:1991:UPC

[RSS91] A. Ruknudin, M. J. Song, and F. Sachs. The ultrastructure of patch-clamped membranes: a study using high voltage electron microscopy. Journal of Cell Biology, 112(1): 125–??, January 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/112/1/125.

Riederer:1994:LBR

[RSS^{+}94] M. A. Riederer, T. Soldati, A. D. Shapiro, J. Lin, and S. R. Pfeffer. Lysosome biogenesis requires Rab9 function and receptor recycling from endosomes to the trans-Golgi network. Journal of Cell Biology, 125(3):573–??, May 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/125/3/573.

Roy:1991:APK

[RSW^{+}91] L. M. Roy, K. I. Swenson, D. H. Walker, B. G. Gabrielli, R. S. Li, H. Piwnica-Worms, and J. L. Maller. Activation of p34cdc2 kinase by cyclin A. Journal of Cell Biology, 113 (3):507–??, May 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/113/3/507.

Russell:1992:NEY

[RT92] I. D. Russell and D. Tollervey. NOP3 is an essential yeast protein which is required for pre-rRNA processing. Journal of Cell Biology, 119(4):737–??, November 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/119/4/737.

Rios:1994:PPA

[RTC^{+}94] R. M. Rios, A. M. Tassin, C. Celati, C. Antony, M. C. Boissier, J. C. Homberg, and M. Bornens. A peripheral protein associated with the cis-Golgi network redistributes in the intermediate compartment upon brefeldin A treatment. Journal of Cell Biology, 125(5):997–??, June 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/125/5/997.

Roberge:1990:TII

[RTHB90] M. Roberge, J. Th’ng, J. Hamaguchi, and E. M. Bradbury. The topoisomerase II inhibitor VM-26 induces marked changes in histone H1 kinase activity, histones H1 and H3 phosphorylation, and chromosome condensation in G2 phase and mitotic BHK cells. Journal of Cell Biology, 111 (5):1753–??, November 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/111/5/1753.

Rieu:1994:DBI

[RUH^{+}94] P. Rieu, T. Ueda, I. Haruta, C. P. Sharma, and M. A. Arnaout. The A-domain of beta 2 integrin CR3 (CD11b/ CD18) is a receptor for the hookworm-derived neutrophil adhesion inhibitor NIF. Journal of Cell Biology, 127(6): 2081–??, December 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/127/6/2081.

Rogers:1990:GEV

[RWG90] M. B. Rogers, S. C. Watkins, and L. J. Gudas. Gene expression in visceral endoderm: a comparison of mutant and wild-type F9 embryonal carcinoma cell differentiation. Journal of Cell Biology, 110(5):1767–??, May 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/110/5/1767.

Rao:1992:IPS

[RWG^{+}92] Y. Rao, X. F. Wu, J. Gariepy, U. Rutishauser, and C. H. Siu. Identification of a peptide sequence involved in homophilic binding in the neural cell adhesion molecule NCAM. Journal of Cell Biology, 118(4):937–??, August 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/118/4/937.

Romisch:1990:KPS

[RWL^{+}90] K. Römisch, J. Webb, K. Lingelbach, H. Gausepohl, and B. Dobberstein. The 54-kD protein of signal recognition particle contains a methionine-rich RNA binding domain. Journal of Cell Biology, 111(5):1793–??, November 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/111/5/

Rothberg:1990:GLF

[RYK^{+}90] K. G. Rothberg, Y. S. Ying, J. F. Kolhouse, B. A. Kamen, and R. G. Anderson. The glycophospholipid-linked folate receptor internalizes folate without entering the clathrincoated pit endocytic pathway. Journal of Cell Biology, 110 (3):637–??, March 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/110/3/637.

Rothberg:1990:CCC

[RYKA90] K. G. Rothberg, Y. S. Ying, B. A. Kamen, and R. G. Anderson. Cholesterol controls the clustering of the glycophospholipid-anchored membrane receptor for 5methyltetrahydrofolate. Journal of Cell Biology, 111(6): 2931–??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/111/6/2931.

Roth:1991:CFN

[RZS91] M. B. Roth, A. M. Zahler, and J. A. Stolk. A conserved family of nuclear phosphoproteins localized to sites of polymerase II transcription. Journal of Cell Biology, 115(3): 587–??, November 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/115/3/587.

Re:1994:IAD

[RZS^{+}94] F. Re, A. Zanetti, M. Sironi, N. Polentarutti, L. Lanfrancone, E. Dejana, and F. Colotta. Inhibition of anchoragedependent cell spreading triggers apoptosis in cultured human endothelial cells. Journal of Cell Biology, 127(2):537– ??, October 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/127/2/537.

Small:1990:EUN

[SA90] S. J. Small and R. Akeson. Expression of the unique NCAM VASE exon is independently regulated in distinct tissues during development. Journal of Cell Biology, 111(5):2089– ??, November 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/111/5/2089.

Simon-Assmann:1990:TIC

[SABF^{+}90] P. Simon-Assmann, F. Bouziges, J. N. Freund, F. Perrin-Schmitt, and M. Kedinger. Type IV collagen mRNA accumulates in the mesenchymal compartment at early stages of murine developing intestine. Journal of Cell Biology, 110 (3):849–??, March 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/110/3/849.

Sanchez-Aparicio:1994:AAB

[SADJGP94] P. Sánchez-Aparicio, C. Dominguez-Jiménez, and A. Garcia-Pardo. Activation of the alpha 4 beta 1 integrin through the beta 1 subunit induces recognition of the RGDS sequence in fibronectin. Journal of Cell Biology, 126(1):271– ??, July 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/126/1/271.

Schnabel:1990:TJP

[SAF90] E. Schnabel, J. M. Anderson, and M. G. Farquhar. The tight junction protein ZO-1 is concentrated along slit diaphragms of the glomerular epithelium. Journal of Cell Biology, 111(3):1255–??, September 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/111/3/1255.

Sammak:1992:ICA

[SAH^{+}92] P. J. Sammak, S. R. Adams, A. T. Harootunian, M. Schliwa, and R. Y. Tsien. Intracellular cyclic AMP not calcium, determines the direction of vesicle movement in melanophores: direct measurement by fluorescence ratio imaging. Journal of Cell Biology, 117(1):57–??, April 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/117/1/57.

Schmid:1994:TDS

[SAJ^{+}94] M. F. Schmid, J. M. Agris, J. Jakana, P. Matsudaira, and W. Chiu. Three-dimensional structure of a single filament in the Limulus acrosomal bundle: scruin binds to homologous helix-loop-beta motifs in actin. Journal of Cell Biology, 124(3):341–??, February 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/124/3/341.

Sutko:1991:FPI

[SAM^{+}91] J. L. Sutko, J. A. Airey, K. Murakami, M. Takeda, C. Beck, T. Deerinck, and M. H. Ellisman. Foot protein isoforms are expressed at different times during embryonic chick skeletal muscle development. Journal of Cell Biology, 113(4): 793–??, May 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/113/4/793.

Steinbeck:1994:NOE

[SAVK94] M. J. Steinbeck, W. H. Appel, A. J. Verhoeven, and M. J. Karnovsky. NADPH-oxidase expression and in situ production of superoxide by osteoclasts actively resorbing bone. Journal of Cell Biology, 126(3):765–??, August 1994. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/126/3/
765.

Smith:1990:RDD

[SAWL90] M. F. Smith, B. D. Athey, S. P. Williams, and J. P. Langmore. Radial density distribution of chromatin: evidence that chromatin fibers have solid centers. Journal of Cell Biology, 110(2):245–??, February 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/110/2/245.

Streuli:1990:EEM

[SB90] C. H. Streuli and M. J. Bissell. Expression of extracellular matrix components is regulated by substratum. Journal of Cell Biology, 110(4):1405–??, April 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/110/4/1405.

Stafford:1991:PCS

[SB91] F. J. Stafford and J. S. Bonifacino. A permeabilized cell system identifies the endoplasmic reticulum as a site of protein degradation. Journal of Cell Biology, 115(5):1225–??, December 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/115/5/1225.

Schnell:1993:IIP

[SB93a] D. J. Schnell and G. Blobel. Identification of intermediates in the pathway of protein import into chloroplasts and their localization to envelope contact sites. Journal of Cell Biology, 120(1):103–??, January 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/120/1/103.

Shtivelman:1993:HGA

[SB93b] E. Shtivelman and J. M. Bishop. The human gene AH-NAK encodes a large phosphoprotein located primarily in the nucleus. Journal of Cell Biology, 120(3):625–??, February 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/120/3/625.

Smith:1993:FMS

[SB93c] S. Smith and G. Blobel. The first membrane spanning region of the lamin B receptor is sufficient for sorting to the inner nuclear membrane. Journal of Cell Biology, 120(3): 631–??, February 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/120/3/631.

Salminen:1991:TMR

[SBB^{+}91a] A. Salminen, T. Braun, A. Buchberger, S. Jürs, B. Winter, and H. H. Arnold. Transcription of the muscle regulatory gene Myf4 is regulated by serum components, peptide growth factors and signaling pathways involving G proteins. Journal of Cell Biology, 115(4):905–??, November 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/115/4/
905.

Streuli:1991:CME

[SBB91b] C. H. Streuli, N. Bailey, and M. J. Bissell. Control of mammary epithelial differentiation: basement membrane induces tissue-specific gene expression in the absence of cell–cell interaction and morphological polarity. Journal of Cell Biology, 115(5):1383–??, December 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/115/5/1383.

Simerly:1990:MCC

[SBBS90] C. Simerly, R. Balczon, B. R. Brinkley, and G. Schatten. Microinjected centromere [corrected] kinetochore antibodies interfere with chromosome movement in meiotic and mitotic mouse oocytes. Journal of Cell Biology, 111(4): 1491–??, October 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/111/4/1491.

Stappenbeck:1993:FAD

[SBC^{+}93] T. S. Stappenbeck, E. A. Bornslaeger, C. M. Corcoran, H. H. Luu, M. L. Virata, and K. J. Green. Functional analysis of desmoplakin domains: specification of the interaction with keratin versus vimentin intermediate filament networks. Journal of Cell Biology, 123(3):691–??, November 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/123/3/691.

Stocklin:1993:AAC

[SBG93] E. Stöcklin, F. Botteri, and B. Groner. An activated allele of the c-erbB-2 oncogene impairs kidney and lung function and causes early death of transgenic mice. Journal of Cell Biology, 122(1):199–??, July 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/122/1/199.

Schlaepfer:1992:DCE

[SBH92] D. D. Schlaepfer, H. R. Bode, and H. T. Haigler. Distinct cellular expression pattern of annexins in Hydra vulgaris. Journal of Cell Biology, 118(4):911–??, August 1992. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/118/4/
911.

Sealock:1991:LDR

[SBK^{+}91] R. Sealock, M. H. Butler, N. R. Kramarcy, K. X. Gao, A. A. Murnane, K. Douville, and S. C. Froehner. Localization of dystrophin relative to acetylcholine receptor domains in electric tissue and adult and cultured skeletal muscle. Journal of Cell Biology, 113(5):1133–??, June 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/113/5/1133.

Scherthan:1994:DCO

[SBK94] H. Scherthan, J. Bähler, and J. Kohli. Dynamics of chromosome organization and pairing during meiotic prophase in fission yeast. Journal of Cell Biology, 127(2):273–??, October 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/127/2/273.

Suzuki:1991:RRC

[SBL^{+}91] C. K. Suzuki, J. S. Bonifacino, A. Y. Lin, M. M. Davis, and R. D. Klausner. Regulating the retention of T-cell receptor alpha chain variants within the endoplasmic reticulum: Ca(2+)-dependent association with BiP. Journal of Cell Biology, 114(2):189–??, July 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/114/2/189.

Straub:1992:DVD

[SBLV92] V. Straub, R. E. Bittner, J. J. Léger, and T. Voit. Direct visualization of the dystrophin network on skeletal muscle fiber membrane. Journal of Cell Biology, 119(5):1183–??, December 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/119/5/1183.

Schmidt:1992:PTK

[SBN92] J. W. Schmidt, J. S. Brugge, and W. J. Nelson. pp60src tyrosine kinase modulates P19 embryonal carcinoma cell fate by inhibiting neuronal but not epithelial differentiation. Journal of Cell Biology, 116(4):1019–??, February 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/116/4/
1019.

Schnell:1990:CIR

[SBP90] D. J. Schnell, G. Blobel, and D. Pain. The chloroplast import receptor is an integral membrane protein of chloroplast envelope contact sites. Journal of Cell Biology, 111(5):1825– ??, November 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/111/5/1825.

Staddon:1991:NAD

[SBR91] J. M. Staddon, M. M. Bouzyk, and E. Rozengurt. A novel approach to detect toxin-catalyzed ADP-ribosylation in intact cells: its use to study the action of Pasteurella multocida toxin. Journal of Cell Biology, 115(4):949–??, November 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/115/4/949.

Seaman:1993:TMP

[SBR93] M. N. Seaman, C. L. Ball, and M. S. Robinson. Targeting and mistargeting of plasma membrane adaptors in vitro. Journal of Cell Biology, 123(5):1093–??, December 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/123/5/
1093.

Segal:1992:DBF

[SBSG92] M. S. Segal, J. M. Bye, J. F. Sambrook, and M. J. Gething. Disulfide bond formation during the folding of influenza virus hemagglutinin. Journal of Cell Biology, 118(2):227– ??, July 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/118/2/227.

Sommerville:1993:HHA

[SBT93] J. Sommerville, J. Baird, and B. M. Turner. Histone H4 acetylation and transcription in amphibian chromatin. Journal of Cell Biology, 120(2):277–??, January 1993. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/120/2/
277.

Skoufias:1990:STC

[SBW90] D. A. Skoufias, T. L. Burgess, and L. Wilson. Spatial and temporal colocalization of the Golgi apparatus and microtubules rich in detyrosinated tubulin. Journal of Cell Biology, 111(5):1929–??, November 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/111/5/1929.

Selinfreund:1990:AIG

[SBWV90] R. H. Selinfreund, S. W. Barger, M. J. Welsh, and L. J. Van Eldik. Antisense inhibition of glial S100 beta production results in alterations in cell morphology, cytoskeletal organization, and cell proliferation. Journal of Cell Biology, 111(5):2021–??, November 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/111/5/2021.

Schmid:1990:ARR

[SC90a] S. L. Schmid and L. L. Carter. ATP is required for receptormediated endocytosis in intact cells. Journal of Cell Biology, 111(6):2307–??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/111/6/2307.

Spriggs:1990:IPT

[SC90b] M. K. Spriggs and P. L. Collins. Intracellular processing and transport of NH2-terminally truncated forms of a hemagglutinin-neuraminidase type II glycoprotein. Journal of Cell Biology, 111(1):31–??, July 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/111/1/31.

Saunders:1991:CWN

[SCE91] W. S. Saunders, C. A. Cooke, and W. C. Earnshaw. Compartmentalization within the nucleus: discovery of a novel subnuclear region. Journal of Cell Biology, 115(4):919–??, November 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/115/4/919.

Spang:1993:CBP

[SCF^{+}93] A. Spang, I. Courtney, U. Fackler, M. Matzner, and E. Schiebel. The calcium-binding protein cell division cycle 31 of Saccharomyces cerevisiae is a component of the half bridge of the spindle pole body. Journal of Cell Biology, 123(2):405–??, October 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/123/2/405.

Schwarzbauer:1991:IFS

[Sch91] J. E. Schwarzbauer. Identification of the fibronectin sequences required for assembly of a fibrillar matrix. Journal of Cell Biology, 113(6):1463–??, June 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/113/6/1463.

Schwartz:1993:SHE

[Sch93] M. A. Schwartz. Spreading of human endothelial cells on fibronectin or vitronectin triggers elevation of intracellular free calcium. Journal of Cell Biology, 120(4):1003–??, February 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/120/4/1003.

Schroer:1994:NII

[Sch94] T. A. Schroer. New insights into the interaction of cytoplasmic dynein with the actin-related protein, Arp1. Journal of Cell Biology, 127(1):1–??, October 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/127/1/1.

Sonnenberg:1991:IAB

[SCJ^{+}91] A. Sonnenberg, J. Calafat, H. Janssen, H. Daams, L. M. van der Raaij-Helmer, R. Falcioni, S. J. Kennel, J. D. Aplin, J. Baker, and M. Loizidou. Integrin alpha 6/beta 4 complex is located in hemidesmosomes, suggesting a major role in epidermal cell-basement membrane adhesion. Journal of Cell Biology, 113(4):907–??, May 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/113/4/907.

Saedi:1991:ATA

[SCL91] M. S. Saedi, W. G. Conroy, and J. Lindstrom. Assembly of Torpedo acetylcholine receptors in Xenopus oocytes. Journal of Cell Biology, 112(5):1007–??, March 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/112/5/1007.

Sanderson:1990:PRY

[SCM90] C. M. Sanderson, J. S. Crowe, and D. I. Meyer. Protein retention in yeast rough endoplasmic reticulum: expression and assembly of human ribophorin I. Journal of Cell Biology, 111(6):2861–??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/111/6/2861.

Scotland:1993:CAR

[SCM^{+}93] P. B. Scotland, M. Colledge, I. Melnikova, Z. Dai, and S. C. Froehner. Clustering of the acetylcholine receptor by the 43kD protein: involvement of the zinc finger domain. Journal of Cell Biology, 123(3):719–??, November 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/123/3/719.

Sadler:1992:ZCT

[SCMB92] I. Sadler, A. W. Crawford, J. W. Michelsen, and M. C. Beckerle. Zyxin and cCRP: two interactive LIM domain proteins associated with the cytoskeleton. Journal of Cell Biology, 119(6):1573–??, December 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/119/6/1573.

Schultz-Cherry:1993:TCA

[SCMU93] S. Schultz-Cherry and J. E. Murphy-Ullrich. Thrombospondin causes activation of latent transforming growth factor-beta secreted by endothelial cells by a novel mechanism. Journal of Cell Biology, 122(4):923–??, August 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/122/4/
923.

Saad:1991:ADT

[SCO^{+}91] B. Saad, D. B. Constam, R. Ortmann, M. Moos, A. Fontana, and M. Schachner. Astrocyte-derived TGFbeta 2 and NGF differentially regulate neural recognition molecule expression by cultured astrocytes. Journal of Cell Biology, 115(2):473–??, October 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/115/2/473.

Simon:1990:FCA

[SCS90] S. I. Simon, J. D. Chambers, and L. A. Sklar. Flow cytometric analysis and modeling of cell–cell adhesive interactions: the neutrophil as a model. Journal of Cell Biology, 111(6):2747–??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/111/6/2747.

Smythe:1992:CCD

[SCS92] E. Smythe, L. L. Carter, and S. L. Schmid. Cytosol- and clathrin-dependent stimulation of endocytosis in vitro by purified adaptors. Journal of Cell Biology, 119(5):1163– ??, December 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/119/5/1163.

Strauss:1990:GED

[SCSE90] P. G. Strauss, E. I. Closs, J. Schmidt, and V. Erfle. Gene expression during osteogenic differentiation in mandibular condyles in vitro. Journal of Cell Biology, 110(4):1369–??, April 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/110/4/1369.

Seethaler:1991:TFP

[SCV^{+}91] G. Seethaler, M. Chaminade, R. Vlasak, M. Ericsson, G. Griffiths, O. Toffoletto, J. Rossier, H. G. Stunnenberg, and G. Kreil. Targeting of frog prodermorphin to the regulated secretory pathway by fusion to proenkephalin. Journal of Cell Biology, 114(6):1125–??, September 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/114/6/1125.

Stow:1991:HGP

[SdAN^{+}91] J. L. Stow, J. B. de Almeida, N. Narula, E. J. Holtzman, L. Ercolani, and D. A. Ausiello. A heterotrimeric G protein, G alpha i-3, on Golgi membranes regulates the secretion of a heparan sulfate proteoglycan in LLC–PK1 epithelial cells. Journal of Cell Biology, 114(6):1113–??, September 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/114/6/
1113.

Simpson:1993:CAC

[SDCD93] D. G. Simpson, M. L. Decker, W. A. Clark, and R. S. Decker. Contractile activity and cell–cell contact regulate myofibrillar organization in cultured cardiac myocytes.

Journal of Cell Biology, 123(2):323–??, October 1993. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/123/2/
323.

Sodeik:1993:AVV

[SDE^{+}93] B. Sodeik, R. W. Doms, M. Ericsson, G. Hiller, C. E. Machamer, W. van ’t Hof, G. van Meer, B. Moss, and G. Griffiths. Assembly of vaccinia virus: role of the intermediate compartment between the endoplasmic reticulum and the Golgi stacks. Journal of Cell Biology, 121 (3):521–??, May 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/121/3/521.

Stallcup:1990:INC

[SDH90] W. B. Stallcup, K. Dahlin, and P. Healy. Interaction of the NG2 chondroitin sulfate proteoglycan with type VI collagen. Journal of Cell Biology, 111(6):3177–??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/111/6/
3177.

Sharma:1993:NGF

[SDK^{+}93] N. Sharma, G. D’Arcangelo, A. Kleinlaus, S. Halegoua, and J. S. Trimmer. Nerve growth factor regulates the abundance and distribution of K+ channels in PC12 cells. Journal of Cell Biology, 123(6):1835–??, December 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/123/6/1835.

Starr:1990:PSH

[SDPH90] C. M. Starr, M. D’Onofrio, M. K. Park, and J. A. Hanover. Primary sequence and heterologous expression of nuclear pore glycoprotein p62. Journal of Cell Biology, 110(6): 1861–??, June 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/110/6/1861.

Shaw:1990:ADI

[SDR^{+}90] R. J. Shaw, D. E. Doherty, A. G. Ritter, S. H. Benedict, and R. A. Clark. Adherence-dependent increase in human monocyte PDGF(B) mRNA is associated with increases in c-fos, c-jun, and EGR2 mRNA. Journal of Cell Biology, 111(5):2139–??, November 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/111/5/2139.

Solimena:1994:SLW

[SDR^{+}94] M. Solimena, R. Dirkx, M. Radzynski, O. Mundigl, and P. De Camilli. A signal located within amino acids 1–27 of GAD65 is required for its targeting to the Golgi complex region. Journal of Cell Biology, 126(2):331–??, July 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/126/2/
331.

Schluter:1993:CPA

[SDW^{+}93] C. Schlüter, M. Duchrow, C. Wohlenberg, M. H. Becker, G. Key, H. D. Flad, and J. Gerdes. The cell proliferationassociated antigen of antibody Ki-67: a very large, ubiquitous nuclear protein with numerous repeated elements, representing a new kind of cell cycle-maintaining proteins. Journal of Cell Biology, 123(3):513–??, November 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/123/3/
513.

Solowska:1991:CTD

[SEAB91] J. Solowska, J. M. Edelman, S. M. Albelda, and C. A. Buck. Cytoplasmic and transmembrane domains of integrin beta 1 and beta 3 subunits are functionally interchangeable. Journal of Cell Biology, 114(5):1079–??, September 1991. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/114/5/
1079.

Sanes:1990:MHB

[SEBH90] J. R. Sanes, E. Engvall, R. Butkowski, and D. D. Hunter. Molecular heterogeneity of basal laminae: isoforms of laminin and collagen IV at the neuromuscular junction and elsewhere. Journal of Cell Biology, 111(4):1685–??, October 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/111/4/1685.

Samuels:1993:ECV

[SEC^{+}93] M. Samuels, R. M. Ezzell, T. J. Cardozo, D. R. Critchley, J. L. Coll, and E. D. Adamson. Expression of chicken vinculin complements the adhesion-defective phenotype of a mutant mouse F9 embryonal carcinoma cell. Journal of Cell Biology, 121(4):909–??, May 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/121/4/909.

Shih:1993:SFA

[SEH^{+}93] D. T. Shih, J. M. Edelman, A. F. Horwitz, G. B. Grunwald, and C. A. Buck. Structure/ function analysis of the integrin beta 1 subunit by epitope mapping. Journal of Cell Biology, 122(6):1361–??, September 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/122/6/1361.

Stamatoglou:1992:TCE

[SEMH92] S. C. Stamatoglou, C. Enrich, M. M. Manson, and R. C. Hughes. Temporal changes in the expression and distribution of adhesion molecules during liver development and regeneration. Journal of Cell Biology, 116(6):1507–??, March 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/116/6/1507.

Srinivasan:1994:ASI

[SES94] M. Srinivasan, C. F. Edman, and H. Schulman. Alternative splicing introduces a nuclear localization signal that targets multifunctional CaM kinase to the nucleus. Journal of Cell Biology, 126(4):839–??, August 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/126/4/839.

Sarthy:1990:LLB

[SF90a] P. V. Sarthy and M. Fu. Localization of laminin B1 mRNA in retinal ganglion cells by in situ hybridization. Journal of Cell Biology, 110(6):2099–??, June 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/110/6/2099.

Stollberg:1990:ARC

[SF90b] J. Stollberg and S. E. Fraser. Acetylcholine receptor clustering is triggered by a change in the density of a nonreceptor molecule. Journal of Cell Biology, 111(5):2029–??, November 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/111/5/2029.

Soroka:1991:CNH

[SF91] C. J. Soroka and M. G. Farquhar. Characterization of a novel heparan sulfate proteoglycan found in the extracellular matrix of liver sinusoids and basement membranes. Journal of Cell Biology, 113(5):1231–??, June 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/113/5/1231.

Schroeder:1990:VTA

[SFA90] C. C. Schroeder, A. K. Fok, and R. D. Allen. Vesicle transport along microtubular ribbons and isolation of cytoplasmic dynein from Paramecium. Journal of Cell Biology, 111(6):2553–??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/111/6/2553.

Schroer:1994:ARP

[SFC^{+}94] T. A. Schroer, E. Fyrberg, J. A. Cooper, R. H. Waterston, D. Helfman, T. D. Pollard, and D. I. Meyer. Actin-related protein nomenclature and classification. Journal of Cell Biology, 127(6):1777–??, December 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/127/6/1777.

Sosnowski:1993:IER

[SFF93] R. G. Sosnowski, S. Feldman, and J. R. Feramisco. Interference with endogenous ras function inhibits cellular responses to wounding. Journal of Cell Biology, 121(1):113– ??, April 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/121/1/113.

Sano:1990:DHL

[SFM^{+}90] H. Sano, R. Forough, J. A. Maier, J. P. Case, A. Jackson, K. Engleka, T. Maciag, and R. L. Wilder. Detection of high levels of heparin binding growth factor-1 (acidic fibroblast growth factor) in inflammatory arthritic joints. Journal of Cell Biology, 110(4):1417–??, April 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/110/4/1417.

Shimizu:1991:NSE

[SFO^{+}91] T. Shimizu, K. Furusawa, S. Ohashi, Y. Y. Toyoshima, M. Okuno, F. Malik, and R. D. Vale. Nucleotide specificity of the enzymatic and motile activities of dynein, kinesin, and heavy meromyosin. Journal of Cell Biology, 112(6): 1189–??, March 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/112/6/1189.

Sossin:1990:SWR

[SFS90] W. S. Sossin, J. M. Fisher, and R. H. Scheller. Sorting within the regulated secretory pathway occurs in the trans-Golgi network. Journal of Cell Biology, 110(1):1–??, January 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/110/1/1.

Spiro:1991:UCS

[SFSG91] R. C. Spiro, H. H. Freeze, D. Sampath, and J. A. Garcia. Uncoupling of chondroitin sulfate glycosaminoglycan synthesis by brefeldin A. Journal of Cell Biology, 115(5):1463– ??, December 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/115/5/1463.

Smart:1994:PKC

[SFY^{+}94] E. J. Smart, D. C. Foster, Y. S. Ying, B. A. Kamen, and R. G. Anderson. Protein kinase C activators inhibit receptor-mediated potocytosis by preventing internalization of caveolae. Journal of Cell Biology, 124(3):307–??, February 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/124/3/307.

Stappenbeck:1992:DCT

[SG92] T. S. Stappenbeck and K. J. Green. The desmoplakin carboxyl terminus coaligns with and specifically disrupts intermediate filament networks when expressed in cultured cells. Journal of Cell Biology, 116(5):1197–??, March 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/116/5/
1197.

Schulze-Garg:1993:TSA

[SGBN^{+}93] C. Schulze-Garg, C. Böker, S. K. Nadimpalli, K. von Figura, and A. Hille-Rehfeld. Tail-specific antibodies that block return of 46,000 M(r) mannose 6-phosphate receptor to the trans-Golgi network. Journal of Cell Biology, 122(3):541–??, August 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/122/3/541.

Schafer:1994:UAD

[SGC^{+}94] D. A. Schafer, S. R. Gill, J. A. Cooper, J. E. Heuser, and T. A. Schroer. Ultrastructural analysis of the dynactin complex: an actin-related protein is a component of a filament that resembles F-actin. Journal of Cell Biology, 126(2):403– ??, July 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/126/2/403.

Slot:1991:ILI

[SGG^{+}91] J. W. Slot, H. J. Geuze, S. Gigengack, G. E. Lienhard, and D. E. James. Immuno-localization of the insulin regulatable glucose transporter in brown adipose tissue of the rat. Journal of Cell Biology, 113(1):123–??, April 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/113/1/123.

Soyer-Gobillard:1990:LBZ

[SGGC^{+}90] M. O. Soyer-Gobillard, M. L. Géraud, D. Coulaud, M. Barray, B. Théveny, B. Révet, and E. Delain. Location of Band Z–DNA in the chromosomes of a primitive eukaryote dinoflagellate. Journal of Cell Biology, 111(2):293–??, August 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/111/2/293.

Stamatoglou:1990:ING

[SGM^{+}90a] S. C. Stamatoglou, R. C. Ge, G. Mills, T. D. Butters, F. Zaidi, and R. C. Hughes. Identification of a novel glycoprotein (AGp110) involved in interactions of rat liver parenchymal cells with fibronectin. Journal of Cell Biology, 111(5):2117–??, November 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/111/5/2117.

Swartz:1990:RBS

[SGM90b] D. R. Swartz, M. L. Greaser, and B. B. Marsh. Regulation of binding of subfragment 1 in isolated rigor myofibrils. Journal of Cell Biology, 111(6):2989–??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/111/6/
2989.

Stark:1991:FCM

[SGN91] F. Stark, R. Golla, and V. T. Nachmias. Formation and contraction of a microfilamentous shell in saponinpermeabilized platelets. Journal of Cell Biology, 112(5): 903–??, March 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/112/5/903.

Sipes:1993:IFB

[SGN^{+}93] J. M. Sipes, N. Guo, E. Nègre, T. Vogel, H. C. Krutzsch, and D. D. Roberts. Inhibition of fibronectin binding and fibronectin-mediated cell adhesion to collagen by a peptide from the second type I repeat of thrombospondin. Journal of Cell Biology, 121(2):469–??, April 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/121/2/469.

Snyder:1991:SCT

[SGP91] M. Snyder, S. Gehrung, and B. D. Page. Studies concerning the temporal and genetic control of cell polarity in Saccharomyces cerevisiae. Journal of Cell Biology, 114(3):515–??, August 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/114/3/515.

Salmi:1993:REE

[SGVS^{+}93] M. Salmi, K. Grön-Virta, P. Sointu, R. Grenman, H. Kalimo, and S. Jalkanen. Regulated expression of exon v6 containing isoforms of CD44 in man: downregulation during malignant transformation of tumors of squamocellular origin. Journal of Cell Biology, 122(2):431–??, July 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/122/2/
431.

Scheibe:1991:RAS

[SGW91a] R. J. Scheibe, D. D. Ginty, and J. A. Wagner. Retinoic acid stimulates the differentiation of PC12 cells that are deficient in cAMP-dependent protein kinase. Journal of Cell Biology, 113(5):1173–??, June 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/113/5/1173.

Schnyder:1991:SMC

[SGW^{+}91b] T. Schnyder, H. Gross, H. Winkler, H. M. Eppenberger, and T. Wallimann. Structure of the mitochondrial creatine kinase octamer: high-resolution shadowing and image averaging of single molecules and formation of linear filaments under specific staining conditions. Journal of Cell Biology, 112(1):95–??, January 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/112/1/95.

Saitoh:1994:SAC

[SGWE94] N. Saitoh, I. G. Goldberg, E. R. Wood, and W. C. Earnshaw. ScII: an abundant chromosome scaffold protein is a member of a family of putative ATPases with an unusual predicted tertiary structure. Journal of Cell Biology, 127(2): 303–??, October 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/127/2/303.

Schlaepfer:1990:EAF

[SH90a] D. D. Schlaepfer and H. T. Haigler. Expression of annexins as a function of cellular growth state. Journal of Cell Biology, 111(1):229–??, July 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/111/1/229.

Schlosshauer:1990:NIC

[SH90b] B. Schlosshauer and K. H. Herzog. Neurothelin: an inducible cell surface glycoprotein of blood-brain barrierspecific endothelial cells and distinct neurons. Journal of Cell Biology, 110(4):1261–??, April 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/110/4/1261.

Schibler:1991:CCB

[SH91] M. J. Schibler and B. Huang. The colR4 and colR15 betatubulin mutations in Chlamydomonas reinhardtii confer altered sensitivities to microtubule inhibitors and herbicides by enhancing microtubule stability. Journal of Cell Biology, 113(3):605–??, May 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/113/3/605.

Sullivan:1992:AMR

[SH92] D. S. Sullivan and T. C. Huffaker. Astral microtubules are not required for anaphase B in Saccharomyces cerevisiae. Journal of Cell Biology, 119(2):379–??, October 1992. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/119/2/
379.

Stoker:1990:EES

[SHB90] A. W. Stoker, C. Hatier, and M. J. Bissell. The embryonic environment strongly attenuates v-src oncogenesis in mesenchymal and epithelial tissues, but not in endothelia. Journal of Cell Biology, 111(1):217–??, July 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/111/1/217.

Small:1990:SSV

[SHBD90] J. V. Small, M. Herzog, M. Barth, and A. Draeger. Supercontracted state of vertebrate smooth muscle cell fragments reveals myofilament lengths. Journal of Cell Biology, 111(6):2451–??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/111/6/2451.

Schlenstedt:1993:RNP

[SHDS93] G. Schlenstedt, E. Hurt, V. Doye, and P. A. Silver. Reconstitution of nuclear protein transport with semi-intact yeast cells. Journal of Cell Biology, 123(4):785–??, November 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/123/4/785.

Shyng:1994:GAP

[SHH94] S. L. Shyng, J. E. Heuser, and D. A. Harris. A glycolipidanchored prion protein is endocytosed via clathrin-coated pits. Journal of Cell Biology, 125(6):1239–??, June 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/125/6/
1239.

Schmidt:1993:ICI

[SHLS93] C. E. Schmidt, A. F. Horwitz, D. A. Lauffenburger, and M. P. Sheetz. Integrin-cytoskeletal interactions in migrating fibroblasts are dynamic, asymmetric, and regulated. Journal of Cell Biology, 123(4):977–??, November 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/123/4/977.

Sullivan:1994:HCC

[SHM94] K. F. Sullivan, M. Hechenberger, and K. Masri. Human CENP–A contains a histone H3 related histone fold domain that is required for targeting to the centromere. Journal of Cell Biology, 127(3):581–??, November 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/127/3/581.

Sun:1992:MYC

[SHOA92] G. H. Sun, A. Hirata, Y. Ohya, and Y. Anraku. Mutations in yeast calmodulin cause defects in spindle pole body functions and nuclear integrity. Journal of Cell Biology, 119(6):1625–??, December 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/119/6/1625.

Schornig:1993:SNG

[SHR93] M. Schörnig, R. Heumann, and H. Rohrer. Synthesis of nerve growth factor mRNA in cultures of developing mouse whisker pad, a peripheral target tissue of sensory trigeminal neurons. Journal of Cell Biology, 120(6):1471–??, March 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/120/6/1471.

Samuel:1993:TBS

[SHS^{+}93] S. K. Samuel, R. A. Hurta, M. A. Spearman, J. A. Wright, E. A. Turley, and A. H. Greenberg. TGF-beta 1 stimulation of cell locomotion utilizes the hyaluronan receptor RHAMM and hyaluronan. Journal of Cell Biology, 123(3): 749–??, November 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/123/3/749.

Shapland:1993:PPT

[SHTL93] C. Shapland, J. J. Hsuan, N. F. Totty, and D. Lawson. Purification and properties of transgelin: a transformation and shape change sensitive actin-gelling protein. Journal of Cell Biology, 121(5):1065–??, June 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/121/5/1065.

Stierhof:1994:CPR

[SIR^{+}94] Y. D. Stierhof, T. Ilg, D. G. Russell, H. Hohenberg, and P. Overath. Characterization of polymer release from the flagellar pocket of Leishmania mexicana promastigotes. Journal of Cell Biology, 125(2):321–??, April 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/125/2/321.

Sherr:1993:MMA

[SJG93] E. H. Sherr, M. P. Joyce, and L. A. Greene. Mammalian myosin I alpha, I beta, and I gamma: new widely expressed genes of the myosin I family. Journal of Cell Biology, 120(6): 1405–??, March 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/120/6/1405.

Skalli:1994:ICD

[SJGG94] O. Skalli, J. C. Jones, R. Gagescu, and R. D. Goldman. IFAP 300 is common to desmosomes and hemidesmosomes and is a possible linker of intermediate filaments to these junctions. Journal of Cell Biology, 125(1):159–??, April 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/125/1/159.

Sandberg:1992:ICB

[SJIC92] K. Sandberg, H. Ji, T. Iida, and K. J. Catt. Intercellular communication between follicular angiotensin receptors and Xenopus laevis oocytes: medication by an inositol 1,4,5trisphosphate-dependent mechanism. Journal of Cell Biology, 117(1):157–??, April 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/117/1/157.

Strand:1994:DLGb

[SJM^{+}94] D. Strand, R. Jakobs, G. Merdes, B. Neumann, A. Kalmes, H. W. Heid, I. Husmann, and B. M. Mechler. The Drosophila lethal(2)giant larvae tumor suppressor protein forms homo-oligomers and is associated with nonmuscle myosin II heavy chain. Journal of Cell Biology, 127(5):1361– ??, December 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/127/5/1361.

Sibley:1993:GIS

[SJMK93] M. H. Sibley, J. J. Johnson, C. C. Mello, and J. M. Kramer. Genetic identification, sequence, and alternative splicing of the Caenorhabditis elegans alpha 2(IV) collagen gene. Journal of Cell Biology, 123(1):255–??, October 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/123/1/255.

Schubert:1991:SGF

[SK91] D. Schubert and H. Kimura. Substratum-growth factor collaborations are required for the mitogenic activities of activin and FGF on embryonal carcinoma cells. Journal of Cell Biology, 114(4):841–??, August 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/114/4/841.

Sako:1994:CSP

[SK94a] Y. Sako and A. Kusumi. Compartmentalized structure of the plasma membrane for receptor movements as revealed by a nanometer-level motion analysis. Journal of Cell Biology, 125(6):1251–??, June 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/125/6/1251.

Smith:1994:MSS

[SK94b] T. A. Smith and B. D. Kohorn. Mutations in a signal sequence for the thylakoid membrane identify multiple protein transport pathways and nuclear suppressors. Journal of Cell Biology, 126(2):365–??, July 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/126/2/365.

Solnica-Krezel:1991:VPD

[SKBD91] L. Solnica-Krezel, T. G. Burland, and W. F. Dove. Variable pathways for developmental changes of mitosis and cytokinesis in Physarum polycephalum. Journal of Cell Biology, 113(3):591–??, May 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/113/3/591.

Stoeckli:1991:ASP

[SKD^{+}91] E. T. Stoeckli, T. B. Kuhn, C. O. Duc, M. A. Ruegg, and P. Sonderegger. The axonally secreted protein axonin-1 is a potent substratum for neurite growth. Journal of Cell Biology, 112(3):449–??, February 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/112/3/449.

Stauffer:1991:IPG

[SKGU91] K. A. Stauffer, N. M. Kumar, N. B. Gilula, and N. Unwin. Isolation and purification of gap junction channels. Journal of Cell Biology, 115(1):141–??, October 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/115/1/141.

Sawada:1990:CCH

[SKH90] H. Sawada, H. Konomi, and K. Hirosawa. Characterization of the collagen in the hexagonal lattice of Descemet’s membrane: its relation to type VIII collagen. Journal of Cell Biology, 110(1):219–??, January 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/110/1/219.

Sorkin:1991:REG

[SKK^{+}91] A. Sorkin, S. Krolenko, N. Kudrjavtceva, J. Lazebnik, L. Teslenko, A. M. Soderquist, and N. Nikolsky. Recycling of epidermal growth factor-receptor complexes in A431 cells: identification of dual pathways. Journal of Cell Biology, 112(1):55–??, January 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/112/1/55.

Smith:1994:SSM

[SKM94] T. H. Smith, A. M. Kachinsky, and J. B. Miller. Somite subdomains, muscle cell origins, and the four muscle regulatory factor proteins. Journal of Cell Biology, 127(1): 95–??, October 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/127/1/95.

Schafer:1994:DLS

[SKSC94] D. A. Schafer, Y. O. Korshunova, T. A. Schroer, and J. A. Cooper. Differential localization and sequence analysis of capping protein beta-subunit isoforms of vertebrates. Journal of Cell Biology, 127(2):453–??, October 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/127/2/453.

Singer-Kruger:1994:RTR

[SKSD^{+}94] B. Singer-Krüger, H. Stenmark, A. Düsterhöft, P. Philippsen, J. S. Yoo, D. Gallwitz, and M. Zerial. Role of three rab5-like GTPases, Ypt51p, Ypt52p, and Ypt53p, in the endocytic and vacuolar protein sorting pathways of yeast. Journal of Cell Biology, 125(2):283–??, April 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/125/2/283.

Schwyter:1990:SCA

[SKT^{+}90] D. H. Schwyter, S. J. Kron, Y. Y. Toyoshima, J. A. Spudich, and E. Reisler. Subtilisin cleavage of actin inhibits in vitro sliding movement of actin filaments over myosin. Journal of Cell Biology, 111(2):465–??, August 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/111/2/465.

Shariff:1990:DDP

[SL90a] A. Shariff and E. J. Luna. Dictyostelium discoideum plasma membranes contain an actin-nucleating activity that requires ponticulin, an integral membrane glycoprotein. Journal of Cell Biology, 110(3):681–??, March 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/110/3/681.

Singer:1990:RBA

[SL90b] D. F. Singer and J. J. Linderman. The relationship between antigen concentration, antigen internalization, and antigenic complexes: modeling insights into antigen processing and presentation. Journal of Cell Biology, 111(1):55– ??, July 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/111/1/55.

Scherer:1994:ICD

[SLB^{+}94] P. E. Scherer, M. P. Lisanti, G. Baldini, M. Sargiacomo, C. C. Mastick, and H. F. Lodish. Induction of caveolin during adipogenesis and association of GLUT4 with caveolinrich vesicles. Journal of Cell Biology, 127(5):1233–??, December 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/127/5/1233.

Schevzov:1992:HLE

[SLG92] G. Schevzov, C. Lloyd, and P. Gunning. High level expression of transfected beta- and gamma-actin genes differentially impacts on myoblast cytoarchitecture. Journal of Cell Biology, 117(4):775–??, May 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/117/4/775.

Scheuner:1992:BCG

[SLH92] D. Scheuner, C. D. Logsdon, and R. W. Holz. Bovine chromaffin granule membranes undergo Ca(2+)-regulated exocytosis in frog oocytes. Journal of Cell Biology, 116(2): 359–??, January 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/116/2/359.

Schevzov:1993:DRT

[SLHG93] G. Schevzov, C. Lloyd, D. Hailstones, and P. Gunning. Differential regulation of tropomyosin isoform organization and gene expression in response to altered actin gene expression. Journal of Cell Biology, 121(4):811–??, May 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/121/4/811.

Schultheiss:1991:DVI

[SLI^{+}91] T. Schultheiss, Z. X. Lin, H. Ishikawa, I. Zamir, C. J. Stoeckert, and H. Holtzer. Desmin/ vimentin intermediate filaments are dispensable for many aspects of myogenesis. Journal of Cell Biology, 114(5):953–??, September 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/114/5/
953.

Swanson:1991:CDA

[SLK91] J. A. Swanson, M. Lee, and P. E. Knapp. Cellular dimensions affecting the nucleocytoplasmic volume ratio. Journal of Cell Biology, 115(4):941–??, November 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/115/4/941.

Strunnikov:1993:SEY

[SLK93] A. V. Strunnikov, V. L. Larionov, and D. Koshland. SMC1: an essential yeast gene encoding a putative head-rod-tail protein is required for nuclear division and defines a new ubiquitous protein family. Journal of Cell Biology, 123(6): 1635–??, December 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/123/6/1635.

SanAntonio:1994:MHB

[SLKS94] J. D. San Antonio, A. D. Lander, M. J. Karnovsky, and H. S. Slayter. Mapping the heparin-binding sites on type I collagen monomers and fibrils. Journal of Cell Biology, 125 (5):1179–??, June 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/125/5/1179.

Schultheiss:1990:DDS

[SLL^{+}90] T. Schultheiss, Z. X. Lin, M. H. Lu, J. Murray, D. A. Fischman, K. Weber, T. Masaki, M. Imamura, and H. Holtzer. Differential distribution of subsets of myofibrillar proteins in cardiac nonstriated and striated myofibrils. Journal of Cell Biology, 110(4):1159–??, April 1990. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/110/4/1159.

Stipp:1994:CIM

[SLL94] C. S. Stipp, E. D. Litwack, and A. D. Lander. Cerebroglycan: an integral membrane heparan sulfate proteoglycan that is unique to the developing nervous system and expressed specifically during neuronal differentiation. Journal of Cell Biology, 124(1):149–??, January 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/124/1/149.

Sonnenberg:1990:IRD

[SLM^{+}90] A. Sonnenberg, C. J. Linders, P. W. Modderman, C. H. Damsky, M. Aumailley, and R. Timpl. Integrin recognition of different cell-binding fragments of laminin (P1, E3, E8) and evidence that alpha 6 beta 1 but not alpha 6 beta 4 functions as a major receptor for fragment E8. Journal of Cell Biology, 110(6):2145–??, June 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/110/6/2145.

Stockli:1991:RDD

[SLNN^{+}91] K. A. Stöckli, L. E. Lillien, M. Näher-Noé, G. Breitfeld, R. A. Hughes, M. C. Raff, H. Thoenen, and M. Sendtner. Regional distribution, developmental changes, and cellular localization of CNTF-mRNA and protein in the rat brain. Journal of Cell Biology, 115(2):447–??, October 1991. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/115/2/
447.

Schleef:1991:IML

[SLP91] R. R. Schleef, D. J. Loskutoff, and T. J. Podor. Immunoelectron microscopic localization of type 1 plasminogen activator inhibitor on the surface of activated endothelial cells. Journal of Cell Biology, 113(6):1413–??, June 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/113/6/
1413.

Samuelsson:1993:SLM

[SLPB93] S. J. Samuelsson, P. W. Luther, D. W. Pumplin, and R. J. Bloch. Structures linking microfilament bundles to the membrane at focal contacts. Journal of Cell Biology, 122 (2):485–??, July 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/122/2/485.

Shoemaker:1990:UDS

[SLS^{+}90] M. O. Shoemaker, W. Lau, R. L. Shattuck, A. P. Kwiatkowski, P. E. Matrisian, L. Guerra-Santos, E. Wilson, T. J. Lukas, L. J. Van Eldik, and D. M. Watterson. Use of DNA sequence and mutant analyses and antisense oligodeoxynucleotides to examine the molecular basis of nonmuscle myosin light chain kinase autoinhibition, calmodulin recognition, and activity. Journal of Cell Biology, 111(3):1107–??, September 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/111/3/1107.

Satterwhite:1992:PMI

[SLW^{+}92] L. L. Satterwhite, M. J. Lohka, K. L. Wilson, T. Y. Scherson, L. J. Cisek, J. L. Corden, and T. D. Pollard. Phosphorylation of myosin–II regulatory light chain by cyclinp34cdc2: a mechanism for the timing of cytokinesis. Journal of Cell Biology, 118(3):595–??, August 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/118/3/595.

Skubitz:1991:SPC

[SLWF91] A. P. Skubitz, P. C. Letourneau, E. Wayner, and L. T. Furcht. Synthetic peptides from the carboxy-terminal globular domain of the A chain of laminin: their ability to promote cell adhesion and neurite outgrowth, and interact with heparin and the beta 1 integrin subunit. Journal of Cell Biology, 115(4):1137–??, November 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/115/4/1137.

Scaife:1990:BIA

[SM90] R. Scaife and R. L. Margolis. Biochemical and immunochemical analysis of rat brain dynamin interaction with microtubules and organelles in vivo and in vitro. Journal of Cell Biology, 111(6):3023–??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/111/6/3023.

Sawin:1991:MSA

[SM91a] K. E. Sawin and T. J. Mitchison. Mitotic spindle assembly by two different pathways in vitro. Journal of Cell Biology, 112(5):925–??, March 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/112/5/925.

Sawin:1991:PMF

[SM91b] K. E. Sawin and T. J. Mitchison. Poleward microtubule flux mitotic spindles assembled in vitro. Journal of Cell Biology, 112(5):941–??, March 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/112/5/941.

Swift:1991:GRS

[SM91c] A. M. Swift and C. E. Machamer. A Golgi retention signal in a membrane-spanning domain of coronavirus E1 protein. Journal of Cell Biology, 115(1):19–??, October 1991. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/115/1/
19.

Symons:1991:CAP

[SM91d] M. H. Symons and T. J. Mitchison. Control of actin polymerization in live and permeabilized fibroblasts. Journal of Cell Biology, 114(3):503–??, August 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/114/3/503.

Shamu:1992:SCS

[SM92a] C. E. Shamu and A. W. Murray. Sister chromatid separation in frog egg extracts requires DNA topoisomerase II activity during anaphase. Journal of Cell Biology, 117 (5):921–??, June 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/117/5/921.

Symons:1992:GCC

[SM92b] M. H. Symons and T. J. Mitchison. A GTPase controls cellsubstrate adhesion in Xenopus XTC fibroblasts. Journal of Cell Biology, 118(5):1235–??, September 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/118/5/1235.

Savitz:1993:KRR

[SM93a] A. J. Savitz and D. I. Meyer. 180-kD ribosome receptor is essential for both ribosome binding and protein translocation. Journal of Cell Biology, 120(4):853–??, February 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/120/4/
853.

Shaw:1993:RAB

[SM93b] L. M. Shaw and A. M. Mercurio. Regulation of alpha 6 beta 1 integrin laminin receptor function by the cytoplasmic domain of the alpha 6 subunit. Journal of Cell Biology, 123(4):1017–??, November 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/123/4/1017.

Shaw:1991:FCS

[SMB^{+}91] J. A. Shaw, P. C. Mol, B. Bowers, S. J. Silverman, M. H. Valdivieso, A. Durán, and E. Cabib. The function of chitin synthases 2 and 3 in the Saccharomyces cerevisiae cell cycle. Journal of Cell Biology, 114(1):111–??, July 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/114/1/111.

Sterne-Marr:1992:LGN

[SMBG92] R. Sterne-Marr, J. M. Blevitt, and L. Gerace. O-linked glycoproteins of the nuclear pore complex interact with a cytosolic factor required for nuclear protein import. Journal of Cell Biology, 116(2):271–??, January 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/116/2/271.

Schafer:1992:LCP

[SMC92] D. A. Schafer, M. S. Mooseker, and J. A. Cooper. Localization of capping protein in chicken epithelial cells by immunofluorescence and biochemical fractionation. Journal of Cell Biology, 118(2):335–??, July 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/118/2/335.

Stelly:1991:CAP

[SMCA91] N. Stelly, J. P. Mauger, M. Claret, and A. Adoutte. Cortical alveoli of Paramecium: a vast submembranous calcium storage compartment. Journal of Cell Biology, 113 (1):103–??, April 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/113/1/103.

Sluder:1990:PSC

[SMCR90] G. Sluder, F. J. Miller, R. Cole, and C. L. Rieder. Protein synthesis and the cell cycle: centrosome reproduction in sea urchin eggs is not under translational control. Journal of Cell Biology, 110(6):2025–??, June 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/110/6/2025.

Selsted:1992:EDA

[SMHO92] M. E. Selsted, S. I. Miller, A. H. Henschen, and A. J. Ouellette. Enteric defensins: antibiotic peptide components of intestinal host defense. Journal of Cell Biology, 118(4): 929–??, August 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/118/4/929.

Smith:1994:INO

[Smi94] C. L. Smith. The initiation of neurite outgrowth by sympathetic neurons grown in vitro does not depend on assembly of microtubules. Journal of Cell Biology, 127(5):1407– ??, December 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/127/5/1407.

Stachowiak:1994:RBG

[SMJ^{+}94] M. K. Stachowiak, J. Moffett, A. Joy, E. Puchacz, R. Florkiewicz, and E. K. Stachowiak. Regulation of bFGF gene expression and subcellular distribution of bFGF protein in adrenal medullary cells. Journal of Cell Biology, 127(1):203–??, October 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/127/1/203.

Sakakibara:1991:COA

[SMK91] H. Sakakibara, D. R. Mitchell, and R. Kamiya. A Chlamydomonas outer arm dynein mutant missing the alpha heavy chain. Journal of Cell Biology, 113(3):615–??, May 1991.

CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/113/3/
615.

Schweizer:1991:IGI

[SMKH91] A. Schweizer, K. Matter, C. M. Ketcham, and H. P. Hauri. The isolated ER–Golgi intermediate compartment exhibits properties that are different from ER and cis-Golgi. Journal of Cell Biology, 113(1):45–??, April 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/113/1/45.

Shaw:1990:ADA

[SMM90] L. M. Shaw, J. M. Messier, and A. M. Mercurio. The activation dependent adhesion of macrophages to laminin involves cytoskeletal anchoring and phosphorylation of the alpha 6 beta 1 integrin. Journal of Cell Biology, 110(6): 2167–??, June 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/110/6/2167.

Stenmark:1991:PFA

[SMM^{+}91] H. Stenmark, J. O. Moskaug, I. H. Madshus, K. Sandvig, and S. Olsnes. Peptides fused to the amino-terminal end of diphtheria toxin are translocated to the cytosol. Journal of Cell Biology, 113(5):1025–??, June 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/113/5/1025.

Sandell:1991:AST

[SMRG91] L. J. Sandell, N. Morris, J. R. Robbins, and M. B. Goldring. Alternatively spliced type II procollagen mRNAs define distinct populations of cells during vertebral development: differential expression of the amino-propeptide. Journal of Cell Biology, 114(6):1307–??, September 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/114/6/1307.

Schell:1992:SAD

[SMSH92] M. J. Schell, M. Maurice, B. Stieger, and A. L. Hubbard. 5’nucleotidase is sorted to the apical domain of hepatocytes via an indirect route. Journal of Cell Biology, 119(5):1173– ??, December 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/119/5/1173.

Sluder:1994:FCM

[SMTW94] G. Sluder, F. J. Miller, E. A. Thompson, and D. E. Wolf. Feedback control of the metaphase-anaphase transition in sea urchin zygotes: role of maloriented chromosomes. Journal of Cell Biology, 126(1):189–??, July 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/126/1/189.

Sonnenberg:1993:SFH

[SMWB93] E. Sonnenberg, D. Meyer, K. M. Weidner, and C. Birchmeier. Scatter factor/hepatocyte growth factor and its receptor, the c-met tyrosine kinase, can mediate a signal exchange between mesenchyme and epithelia during mouse development. Journal of Cell Biology, 123(1):223–??, October 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/123/1/223.

Snow:1993:CBI

[SN93] P. Snow and R. Nuccitelli. Calcium buffer injections delay cleavage in Xenopus laevis blastomeres. Journal of Cell Biology, 122(2):387–??, July 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/122/2/387.

Sarria:1990:REV

[SNE90] A. J. Sarria, S. K. Nordeen, and R. M. Evans. Regulated expression of vimentin cDNA in cells in the presence and absence of a preexisting vimentin filament network. Journal of Cell Biology, 111(2):553–??, August 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/111/2/553.

Shim:1991:BGE

[SNFN91] J. Shim, A. P. Newman, and S. Ferro-Novick. The BOS1 gene encodes an essential 27-kD putative membrane protein that is required for vesicular transport from the ER to the Golgi complex in yeast. Journal of Cell Biology, 113 (1):55–??, April 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/113/1/55.

Shimizu:1991:FMP

[SNG^{+}91] Y. Shimizu, W. Newman, T. V. Gopal, K. J. Horgan, N. Graber, L. D. Beall, G. A. van Seventer, and S. Shaw. Four molecular pathways of T cell adhesion to endothelial cells: roles of LFA-1, VCAM-1, and ELAM-1 and changes in pathway hierarchy under different activation conditions. Journal of Cell Biology, 113(5):1203–??, June 1991. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/113/5/
1203.

Schulze:1994:SDH

[SNGG94] E. Schulze, S. Nagel, K. Gavenis, and U. Grossbach. Structurally divergent histone H1 variants in chromosomes containing highly condensed interphase chromatin. Journal of Cell Biology, 127(6):1789–??, December 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/127/6/1789.

Slusarewicz:1994:IMB

[SNH^{+}94] P. Slusarewicz, T. Nilsson, N. Hui, R. Watson, and G. Warren. Isolation of a matrix that binds medial Golgi enzymes. Journal of Cell Biology, 124(4):405–??, February 1994. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/124/4/
405.

Streit:1993:IAE

[SNRS93] A. Streit, C. Nolte, T. Rásony, and M. Schachner. Interaction of astrochondrin with extracellular matrix components and its involvement in astrocyte process formation and cerebellar granule cell migration. Journal of Cell Biology, 120(3):799–??, February 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/120/3/799.

Sato:1993:MAL

[SOA^{+}93] Y. Sato, F. Okada, M. Abe, T. Seguchi, M. Kuwano, S. Sato, A. Furuya, N. Hanai, and T. Tamaoki. The mechanism for the activation of latent TGF-beta during co-culture of endothelial cells and smooth muscle cells: cell-type specific targeting of latent TGF-beta to smooth muscle cells. Journal of Cell Biology, 123(5):1249–??, December 1993. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/123/5/
1249.

Sabry:1991:MBD

[SOE^{+}91] J. H. Sabry, T. P. O’Connor, L. Evans, A. Toroian-Raymond, M. Kirschner, and D. Bentley. Microtubule behavior during guidance of pioneer neuron growth cones in situ. Journal of Cell Biology, 115(2):381–??, October 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/115/2/
381.

Sim:1990:PSP

[SOH^{+}90] B. K. Sim, P. A. Orlandi, J. D. Haynes, F. W. Klotz, J. M. Carter, D. Camus, M. E. Zegans, and J. D. Chulay. Primary structure of the 175K Plasmodium falciparum erythrocyte binding antigen and identification of a peptide which elicits antibodies that inhibit malaria merozoite invasion. Journal of Cell Biology, 111(5):1877–??, November 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/111/5/1877.

Stochaj:1991:YPB

[SOKS91] U. Stochaj, M. Osborne, T. Kurihara, and P. Silver. A yeast protein that binds nuclear localization signals: purification localization, and antibody inhibition of binding activity. Journal of Cell Biology, 113(6):1243–??, June 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/113/6/
1243.

Sekine:1994:NMB

[SON^{+}94] Y. Sekine, Y. Okada, Y. Noda, S. Kondo, H. Aizawa, R. Takemura, and N. Hirokawa. A novel microtubule-based motor protein (KIF4) for organelle transports, whose expression is regulated developmentally. Journal of Cell Biology, 127(1):187–??, October 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/127/1/187.

Schnitzer:1994:FSC

[SOPA94] J. E. Schnitzer, P. Oh, E. Pinney, and J. Allard. Filipinsensitive caveolae-mediated transport in endothelium: reduced transcytosis, scavenger endocytosis, and capillary permeability of select macromolecules. Journal of Cell Biology, 127(5):1217–??, December 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/127/5/1217.

Schulze-Osthoff:1994:CND

[SOWDK94] K. Schulze-Osthoff, H. Walczak, W. Dröge, and P. H. Krammer. Cell nucleus and DNA fragmentation are not required for apoptosis. Journal of Cell Biology, 127(1):15–??, October 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/127/1/15.

Sung:1993:CHB

[SOY93] U. Sung, J. J. O’Rear, and P. D. Yurchenco. Cell and heparin binding in the distal long arm of laminin: identification of active and cryptic sites with recombinant and hybrid glycoprotein. Journal of Cell Biology, 123(5):1255– ??, December 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/123/5/1255.

Seeger:1992:SIE

[SP92] M. Seeger and G. S. Payne. Selective and immediate effects of clathrin heavy chain mutations on Golgi membrane protein retention in Saccharomyces cerevisiae. Journal of Cell Biology, 118(3):531–??, August 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/118/3/531.

Saucan:1994:MSP

[SP94] L. Saucan and G. E. Palade. Membrane and secretory proteins are transported from the Golgi complex to the sinusoidal plasmalemma of hepatocytes by distinct vesicular carriers. Journal of Cell Biology, 125(4):733–??, May 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/125/4/

Stirzaker:1990:SRG

[SPB90] S. C. Stirzaker, D. Poncet, and G. W. Both. Sequences in rotavirus glycoprotein VP7 that mediate delayed translocation and retention of the protein in the endoplasmic reticulum. Journal of Cell Biology, 111(4):1343–??, October 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/111/4/
1343.

Schwaninger:1992:MGB

[SPBB92] R. Schwaninger, H. Plutner, G. M. Bokoch, and W. E. Balch. Multiple GTP-binding proteins regulate vesicular transport from the ER to Golgi membranes. Journal of Cell Biology, 119(5):1077–??, December 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/119/5/1077.

Schleef:1990:MTP

[SPD^{+}90] R. R. Schleef, T. J. Podor, E. Dunne, J. Mimuro, and D. J. Loskutoff. The majority of type 1 plasminogen activator inhibitor associated with cultured human endothelial cells is located under the cells and is accessible to solution-phase tissue-type plasminogen activator. Journal of Cell Biology, 110(1):155–??, January 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/110/1/155.

Sarafian:1991:PAI

[SPH^{+}91] T. Sarafian, L. A. Pradel, J. P. Henry, D. Aunis, and M. F. Bader. The participation of annexin II (calpactin I) in calcium-evoked exocytosis requires protein kinase C. Journal of Cell Biology, 114(6):1135–??, September 1991. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/114/6/
1135.

Sandvig:1991:RTB

[SPHvD91] K. Sandvig, K. Prydz, S. H. Hansen, and B. van Deurs. Ricin transport in brefeldin A-treated cells: correlation between Golgi structure and toxic effect. Journal of Cell Biology, 115(4):971–??, November 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/115/4/971.

Sjoberg:1994:NLB

[SPKV94] E. R. Sjoberg, L. D. Powell, A. Klein, and A. Varki. Natural ligands of the B cell adhesion molecule CD22 beta can be masked by 9-O-acetylation of sialic acids. Journal of Cell Biology, 126(2):549–??, July 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/126/2/549.

Sandvig:1991:EIT

[SPRvD91] K. Sandvig, K. Prydz, M. Ryd, and B. van Deurs. Endocytosis and intracellular transport of the glycolipid-binding ligand Shiga toxin in polarized MDCK cells. Journal of Cell Biology, 113(3):553–??, May 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/113/3/553.

Simons:1994:AIL

[SPvV94] F. H. Simons, G. J. Pruijn, and W. J. van Venrooij. Analysis of the intracellular localization and assembly of Ro ribonucleoprotein particles by microinjection into Xenopus laevis oocytes. Journal of Cell Biology, 125(5):981–??, June 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/125/5/
981.

Souter:1993:GSR

[SPW93] E. Souter, M. Pypaert, and G. Warren. The Golgi stack reassembles during telophase before arrival of proteins transported from the endoplasmic reticulum. Journal of Cell Biology, 122(3):533–??, August 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/122/3/533.

Saksela:1990:RBF

[SR90a] O. Saksela and D. B. Rifkin. Release of basic fibroblast growth factor-heparan sulfate complexes from endothelial cells by plasminogen activator-mediated proteolytic activity. Journal of Cell Biology, 110(3):767–??, March 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/110/3/

Singer:1990:DIC

[SR90b] B. Singer and H. Riezman. Detection of an intermediate compartment involved in transport of alpha-factor from the plasma membrane to the vacuole in yeast. Journal of Cell Biology, 110(6):1911–??, June 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/110/6/1911.

Silverman:1992:LRF

[SR92a] L. Silverman and M. D. Resh. Lysine residues form an integral component of a novel NH2-terminal membrane targeting motif for myristylated pp60v-src. Journal of Cell Biology, 119(2):415–??, October 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/119/2/415.

Sonderegger:1992:RAG

[SR92b] P. Sonderegger and F. G. Rathjen. Regulation of axonal growth in the vertebrate nervous system by interactions between glycoproteins belonging to two subgroups of the immunoglobulin superfamily. Journal of Cell Biology, 119(6): 1387–??, December 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/119/6/1387.

Sporn:1992:TGF

[SR92c] M. B. Sporn and A. B. Roberts. Transforming growth factor-beta: recent progress and new challenges. Journal of Cell Biology, 119(5):1017–??, December 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/119/5/1017.

Standiford:1992:ADR

[SR92d] D. M. Standiford and J. D. Richter. Analysis of a developmentally regulated nuclear localization signal in Xenopus. Journal of Cell Biology, 118(5):991–??, September 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/118/5/
991.

Sugden:1992:PKC

[SR92e] D. Sugden and S. J. Rowe. Protein kinase C activation antagonizes melatonin-induced pigment aggregation in Xenopus laevis melanophores. Journal of Cell Biology, 119(6): 1515–??, December 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/119/6/1515.

Schnapp:1992:KBH

[SRB92] B. J. Schnapp, T. S. Reese, and R. Bechtold. Kinesin is bound with high affinity to squid axon organelles that move to the plus-end of microtubules. Journal of Cell Biology, 119(2):389–??, October 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/119/2/389.

Sandvig:1994:RTG

[SRG^{+}94] K. Sandvig, M. Ryd, O. Garred, E. Schweda, P. K. Holm, and B. van Deurs. Retrograde transport from the Golgi complex to the ER of both Shiga toxin and the nontoxic Shiga B-fragment is regulated by butyric acid and cAMP. Journal of Cell Biology, 126(1):53–??, July 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/126/1/53.

Schweizer:1994:RPG

[SRHK94] A. Schweizer, J. Rohrer, H. P. Hauri, and S. Kornfeld. Retention of p63 in an ER–Golgi intermediate compartment depends on the presence of all three of its domains and on its ability to form oligomers. Journal of Cell Biology, 126 (1):25–??, July 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/126/1/25.

Shum:1994:ATT

[SRK^{+}94] L. Shum, S. A. Reeves, A. C. Kuo, E. S. Fromer, and R. Derynck. Association of the transmembrane TGF-alpha precursor with a protein kinase complex. Journal of Cell Biology, 125(4):903–??, May 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/125/4/903.

Strand:1994:DLGa

[SRM94] D. Strand, I. Raska, and B. M. Mechler. The Drosophila lethal(2)giant larvae tumor suppressor protein is a component of the cytoskeleton. Journal of Cell Biology, 127(5):

1345–??, December 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/127/5/1345.

Sinard:1990:IFR

[SRP90] J. H. Sinard, D. L. Rimm, and T. D. Pollard. Identification of functional regions on the tail of Acanthamoeba myosin–II using recombinant fusion proteins. II. Assembly properties of tails with NH2- and COOH-terminal deletions. Journal of Cell Biology, 111(6):2417–??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/111/6/2417.

Sheela:1990:AIP

[SRR90] S. Sheela, V. M. Riccardi, and N. Ratner. Angiogenic and invasive properties of neurofibroma Schwann cells. Journal of Cell Biology, 111(2):645–??, August 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/111/2/645.

Satoh:1990:ITR

[SRV^{+}90] T. Satoh, C. A. Ross, A. Villa, S. Supattapone, T. Pozzan, S. H. Snyder, and J. Meldolesi. The inositol 1,4,5,- trisphosphate receptor in cerebellar Purkinje cells: quantitative immunogold labeling reveals concentration in an ER subcompartment. Journal of Cell Biology, 111(2):615–??, August 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/111/2/615.

Sokabe:1990:SDP

[SS90a] M. Sokabe and F. Sachs. The structure and dynamics of patch-clamped membranes: a study using differential interference contrast light microscopy. Journal of Cell Biology, 111(2):599–??, August 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/111/2/599.

Sundell:1990:AML

[SS90b] C. L. Sundell and R. H. Singer. Actin mRNA localizes in the absence of protein synthesis. Journal of Cell Biology, 111(6):2397–??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/111/6/2397.

Schmid:1991:SSA

[SS91a] S. L. Schmid and E. Smythe. Stage-specific assays for coated pit formation and coated vesicle budding in vitro. Journal of Cell Biology, 114(5):869–??, September 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/114/5/869.

Schroer:1991:TAM

[SS91b] T. A. Schroer and M. P. Sheetz. Two activators of microtubule-based vesicle transport. Journal of Cell Biology, 115(5):1309–??, December 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/115/5/1309.

Smith:1992:SFR

[SS92a] E. F. Smith and W. S. Sale. Structural and functional reconstitution of inner dynein arms in Chlamydomonas flagellar axonemes. Journal of Cell Biology, 117(3):573–??, May 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/117/3/573.

Stochaj:1992:CPS

[SS92b] U. Stochaj and P. A. Silver. A conserved phosphoprotein that specifically binds nuclear localization sequences is involved in nuclear import. Journal of Cell Biology, 117(3): 473–??, May 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/117/3/473.

Spong:1993:CCP

[SS93] A. P. Spong and S. Subramani. Cloning and characterization of PAS5: a gene required for peroxisome biogenesis in the methylotrophic yeast Pichia pastoris. Journal of Cell Biology, 123(3):535–??, November 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/123/3/535.

Sanders:1994:CPE

[SS94] M. A. Sanders and J. L. Salisbury. Centrin plays an essential role in microtubule severing during flagellar excision in Chlamydomonas reinhardtii. Journal of Cell Biology, 124 (5):795–??, March 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/124/5/795.

Sorokin:1990:RLE

[SSA^{+}90] L. Sorokin, A. Sonnenberg, M. Aumailley, R. Timpl, and P. Ekblom. Recognition of the laminin E8 cell-binding site by an integrin possessing the alpha 6 subunit is essential for epithelial polarization in developing kidney tubules. Journal of Cell Biology, 111(3):1265–??, September 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/111/3/1265.

Su:1991:ETN

[SSB^{+}91] M. W. Su, H. R. Suzuki, J. J. Bieker, M. Solursh, and F. Ramirez. Expression of two nonallelic type II procollagen genes during Xenopus laevis embryogenesis is characterized by stage-specific production of alternatively spliced transcripts. Journal of Cell Biology, 115(2):565–??, October 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/115/2/565.

Stoler:1993:CMS

[SSB93] A. B. Stoler, F. Stenback, and A. Balmain. The conversion of mouse skin squamous cell carcinomas to spindle cell carcinomas is a recessive event. Journal of Cell Biology, 122(5):1103–??, September 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/122/5/1103.

Staugaitis:1990:EMB

[SSC90] S. M. Staugaitis, P. R. Smith, and D. R. Colman. Expression of myelin basic protein isoforms in nonglial cells. Journal of Cell Biology, 110(5):1719–??, May 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/110/5/1719.

Skach:1994:BTT

[SSC^{+}94] W. R. Skach, L. B. Shi, M. C. Calayag, A. Frigeri, V. R. Lingappa, and A. S. Verkman. Biogenesis and transmembrane topology of the CHIP28 water channel at the endoplasmic reticulum. Journal of Cell Biology, 125(4):803–??, May 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/125/4/803.

Silberstein:1990:EDE

[SSCD90] G. B. Silberstein, P. Strickland, S. Coleman, and C. W. Daniel. Epithelium-dependent extracellular matrix synthesis in transforming growth factor-beta 1-growth-inhibited mouse mammary gland. Journal of Cell Biology, 110(6): 2209–??, June 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/110/6/2209.

Shenoy-Scaria:1994:CSF

[SSDK^{+}94] A. M. Shenoy-Scaria, D. J. Dietzen, J. Kwong, D. C. Link, and D. M. Lublin. Cysteine3 of Src family protein tyrosine kinase determines palmitoylation and localization in caveolae. Journal of Cell Biology, 126(2):353–??, July 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/126/2/
353.

Stephens:1993:TDB

[SSFD93] L. E. Stephens, J. E. Sonne, M. L. Fitzgerald, and C. H. Damsky. Targeted deletion of beta 1 integrins in F9 embryonal carcinoma cells affects morphological differentiation but not tissue-specific gene expression. Journal of Cell Biology, 123(6):1607–??, December 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/123/6/1607.

Sato:1990:EPI

[SSG^{+}90] M. Sato, M. K. Sardana, W. A. Grasser, V. M. Garsky, J. M. Murray, and R. J. Gould. Echistatin is a potent inhibitor of bone resorption in culture. Journal of Cell Biology, 111(4): 1713–??, October 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/111/4/1713.

Saito:1993:AAC

[SSG93] T. Saito, L. Small, and U. W. Goodenough. Activation of adenylyl cyclase in Chlamydomonas reinhardtii by adhesion and by heat. Journal of Cell Biology, 122(1):137– ??, July 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/122/1/137.

Sorger:1994:FRB

[SSH94] P. K. Sorger, F. F. Severin, and A. A. Hyman. Factors required for the binding of reassembled yeast kinetochores to microtubules in vitro. Journal of Cell Biology, 127(4): 995–??, November 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/127/4/995.

Speksnijder:1990:AWC

[SSJ90] J. E. Speksnijder, C. Sardet, and L. F. Jaffe. The activation wave of calcium in the ascidian egg and its role in ooplasmic segregation. Journal of Cell Biology, 110(5):1589–??, May 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/110/5/1589.

Smythe:1994:EOI

[SSJ^{+}94] E. Smythe, P. D. Smith, S. M. Jacob, J. Theobald, and S. E. Moss. Endocytosis occurs independently of annexin VI in human A431 cells. Journal of Cell Biology, 124(3): 301–??, February 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/124/3/301.

Steger:1990:IAA

[SSK^{+}90] H. F. Steger, T. Söllner, M. Kiebler, K. A. Dietmeier, R. Pfaller, K. S. Trülzsch, M. Tropschug, W. Neupert, and N. Pfanner. Import of ADP/ATP carrier into mitochondria: two receptors act in parallel. Journal of Cell Biology, 111(6):2353–??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/111/6/2353.

Streuli:1993:EMR

[SSK^{+}93] C. H. Streuli, C. Schmidhauser, M. Kobrin, M. J. Bissell, and R. Derynck. Extracellular matrix regulates expression of the TGF-beta 1 gene. Journal of Cell Biology, 120(1): 253–??, January 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/120/1/253.

Suh:1992:IGR

[SSKT^{+}92] B. S. Suh, R. Sprengel, I. Keren-Tal, S. Himmelhoch, and A. Amsterdam. Introduction of a gonadotropin receptor expression plasmid into immortalized granulosa cells leads to reconstitution of hormone-dependent steroidogenesis. Journal of Cell Biology, 119(2):439–??, October 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/119/2/439.

Squarzoni:1992:LDC

[SSM^{+}92a] S. Squarzoni, P. Sabatelli, M. C. Maltarello, A. Cataldi, R. di Primio, and N. M. Maraldi. Localization of dystrophin COOH-terminal domain by the fracture-label technique. Journal of Cell Biology, 118(6):1401–??, September 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/118/6/
1401.

Sun:1992:DMR

[SSM92b] X. Sun, K. Skorstengaard, and D. F. Mosher. Disulfides modulate RGD-inhibitable cell adhesive activity of thrombospondin. Journal of Cell Biology, 118(3):693–??, August 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/118/3/693.

Satir:1990:PES

[SSRM90] B. H. Satir, C. Srisomsap, M. Reichman, and R. B. Marchase. Parafusin, an exocytic-sensitive phosphoprotein, is the primary acceptor for the glucosylphosphotransferase in Paramecium tetraurelia and rat liver. Journal of Cell Biology, 111(3):901–??, September 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/111/3/901.

Spurck:1990:UMI

[SSS^{+}90] T. P. Spurck, O. G. Stonington, J. A. Snyder, J. D. Pickett-Heaps, A. Bajer, and J. Mole-Bajer. UV microbeam irradiations of the mitotic spindle. II. Spindle fiber dynamics and force production. Journal of Cell Biology, 111(4): 1505–??, October 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/111/4/1505.

Skibbens:1993:DIK

[SSS93] R. V. Skibbens, V. P. Skeen, and E. D. Salmon. Directional instability of kinetochore motility during chromosome congression and segregation in mitotic newt lung cells: a pushpull mechanism. Journal of Cell Biology, 122(4):859–??, August 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/122/4/859.

Stella:1994:ISC

[SSSL94] M. C. Stella, H. Schauerte, K. L. Straub, and M. Leptin. Identification of secreted and cytosolic gelsolin in Drosophila. Journal of Cell Biology, 125(3):607–??, May 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/
125/3/607.

Schleiffenbaum:1992:SSP

[SST92a] B. Schleiffenbaum, O. Spertini, and T. F. Tedder. Soluble L-selectin is present in human plasma at high levels and retains functional activity. Journal of Cell Biology, 119(1): 229–??, October 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/119/1/229.

Sendtner:1992:SLC

[SST92b] M. Sendtner, K. A. Stöckli, and H. Thoenen. Synthesis and localization of ciliary neurotrophic factor in the sciatic nerve of the adult rat after lesion and during regeneration. Journal of Cell Biology, 118(1):139–??, July 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/118/1/139.

Sargiacomo:1993:STM

[SSTL93] M. Sargiacomo, M. Sudol, Z. Tang, and M. P. Lisanti. Signal transducing molecules and glycosyl-phosphatidylinositol-linked proteins form a caveolin-rich insoluble complex in MDCK cells. Journal of Cell Biology, 122(4):789–??, August 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/122/4/789.

Schliwa:1991:NSA

[SSVE91] M. Schliwa, T. Shimizu, R. D. Vale, and U. Euteneuer. Nucleotide specificities of anterograde and retrograde organelle transport in Reticulomyxa are indistinguishable. Journal of Cell Biology, 112(6):1199–??, March 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/112/6/1199.

Schreiner:1994:NGE

[SSW94] R. Schreiner, E. Schnabel, and F. Wieland. Novel Nglycosylation in eukaryotes: laminin contains the linkage unit beta-glucosylasparagine. Journal of Cell Biology, 124 (6):1071–??, March 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/124/6/1071.

Snipes:1992:CNP

[SSWS92] G. J. Snipes, U. Suter, A. A. Welcher, and E. M. Shooter. Characterization of a novel peripheral nervous system myelin protein (PMP-22/SR13). Journal of Cell Biology, 117(1):225–??, April 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/117/1/225.

Sato:1994:IEA

[ST94] S. B. Sato and S. Toyama. Interference with the endosomal acidification by a monoclonal antibody directed toward the 116 (100)-kD subunit of the vacuolar type proton pump. Journal of Cell Biology, 127(1):39–??, October 1994. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/127/1/

Sympson:1994:TES

[STA^{+}94] C. J. Sympson, R. S. Talhouk, C. M. Alexander, J. R. Chin, S. M. Clift, M. J. Bissell, and Z. Werb. Targeted expression of stromelysin-1 in mammary gland provides evidence for a role of proteinases in branching morphogenesis and the requirement for an intact basement membrane for tissuespecific gene expression. Journal of Cell Biology, 125(3): 681–??, May 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/125/3/681.

Symington:1993:IIA

[STC93] B. E. Symington, Y. Takada, and W. G. Carter. Interaction of integrins alpha 3 beta 1 and alpha 2 beta 1: potential role in keratinocyte intercellular adhesion. Journal of Cell Biology, 120(2):523–??, January 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/120/2/523.

Su:1990:ICC

[STdL^{+}90] J. W. Su, L. G. Tertoolen, S. W. de Laat, W. J. Hage, and A. J. Durston. Intercellular communication is cell cycle modulated during early Xenopus laevis development. Journal of Cell Biology, 110(1):115–??, January 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/110/1/115.

Smola:1993:MIG

[STF93] H. Smola, G. Thiekötter, and N. E. Fusenig. Mutual induction of growth factor gene expression by epidermal-dermal cell interaction. Journal of Cell Biology, 122(2):417–??, July 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/122/2/417.

Speksnijder:1993:PRE

[STH^{+}93] J. E. Speksnijder, M. Terasaki, W. J. Hage, L. F. Jaffe, and C. Sardet. Polarity and reorganization of the endoplasmic reticulum during fertilization and ooplasmic segregation in the ascidian egg. Journal of Cell Biology, 120(6): 1337–??, March 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/120/6/1337.

Sakakibara:1993:COA

[STK^{+}93] H. Sakakibara, S. Takada, S. M. King, G. B. Witman, and R. Kamiya. A Chlamydomonas outer arm dynein mutant with a truncated beta heavy chain. Journal of Cell Biology, 122(3):653–??, August 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/122/3/653.

Sato:1990:CAL

[STL^{+}90] Y. Sato, R. Tsuboi, R. Lyons, H. Moses, and D. B. Rifkin. Characterization of the activation of latent TGF-beta by cocultures of endothelial cells and pericytes or smooth muscle cells: a self-regulating system. Journal of Cell Biology, 111(2):757–??, August 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/111/2/757.

Shaw:1991:ETP

[STM91] M. K. Shaw, L. G. Tilney, and A. J. Musoke. The entry of Theileria parva sporozoites into bovine lymphocytes: evidence for MHC class I involvement. Journal of Cell Biology, 113(1):87–??, April 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/113/1/87.

Sugama:1992:TIE

[STo^{+}92] Y. Sugama, C. Tiruppathi, K. offakidevi, T. T. Andersen, J. W. Fenton, and A. B. Malik. Thrombin-induced expression of endothelial P-selectin and intercellular adhesion molecule-1: a mechanism for stabilizing neutrophil adhesion. Journal of Cell Biology, 119(4):935–??, November 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/119/4/935.

Schweitzer:1991:SAI

[STS91] P. A. Schweitzer, S. E. Taylor, and L. D. Shultz. Synthesis of abnormal immunoglobulins by hybridomas from autoimmune “viable motheaten” mutant mice. Journal of Cell Biology, 114(1):35–??, July 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/114/1/35.

Sikorski:1991:SMI

[STZG91] A. F. Sikorski, G. Terlecki, I. S. Zagon, and S. R. Goodman. Synapsin I-mediated interaction of brain spectrin with synaptic vesicles. Journal of Cell Biology, 114(2):313–??, July 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/114/2/313.

Shi:1994:AAR

[SVB94] Y. Shi, B. Veit, and S. Baekkeskov. Amino acid residues 24–31 but not palmitoylation of cysteines 30 and 45 are required for membrane anchoring of glutamic acid decarboxylase, GAD65. Journal of Cell Biology, 124(6):927–??, March 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/124/6/927.

Sun:1990:SCR

[SVD90] T. J. Sun, P. J. Van Haastert, and P. N. Devreotes. Surface cAMP receptors mediate multiple responses during development in Dictyostelium: evidenced by antisense mutagenesis. Journal of Cell Biology, 110(5):1549–??, May 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/110/5/1549.

Staub:1992:PSA

[SVK^{+}92] O. Staub, F. Verrey, T. R. Kleyman, D. J. Benos, B. C. Rossier, and J. P. Kraehenbuhl. Primary structure of an apical protein from Xenopus laevis that participates in amiloride-sensitive sodium channel activity. Journal of Cell Biology, 119(6):1497–??, December 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/119/6/1497.

Schaerer:1990:PTP

[SVR^{+}90] E. Schaerer, F. Verrey, L. Racine, C. Tallichet, M. Reinhardt, and J. P. Kraehenbuhl. Polarized transport of the polymeric immunoglobulin receptor in transfected rabbit mammary epithelial cells. Journal of Cell Biology, 110(4): 987–??, April 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/110/4/987.

Sanders:1990:ENS

[SW90] M. C. Sanders and Y. L. Wang. Exogenous nucleation sites fail to induce detectable polymerization of actin in living cells. Journal of Cell Biology, 110(2):359–??, February 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/110/2/
359.

Shelden:1992:MBT

[SW92] E. Shelden and P. Wadsworth. Microinjection of biotintubulin into anaphase cells induces transient elongation of kinetochore microtubules and reversal of chromosometo-pole motion. Journal of Cell Biology, 116(6):1409–??, March 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/116/6/1409.

Shelden:1993:OQI

[SW93a] E. Shelden and P. Wadsworth. Observation and quantification of individual microtubule behavior in vivo: microtubule dynamics are cell-type specific. Journal of Cell Biology, 120(4):935–??, February 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/120/4/935.

Soullam:1993:ATD

[SW93b] B. Soullam and H. J. Worman. The amino-terminal domain of the lamin B receptor is a nuclear envelope targeting signal. Journal of Cell Biology, 120(5):1093–??, March 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/120/5/
1093.

Saffell:1992:DAS

[SWD92] J. L. Saffell, F. S. Walsh, and P. Doherty. Direct activation of second messenger pathways mimics cell adhesion molecule-dependent neurite outgrowth. Journal of Cell Biology, 118(3):663–??, August 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/118/3/663.

Saffell:1994:ENC

[SWD94] J. L. Saffell, F. S. Walsh, and P. Doherty. Expression of NCAM containing VASE in neurons can account for a developmental loss in their neurite outgrowth response to NCAM in a cellular substratum. Journal of Cell Biology, 125(2): 427–??, April 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/125/2/427.

Song:1992:HAN

[SWF^{+}92] W. K. Song, W. Wang, R. F. Foster, D. A. Bielser, and S. J. Kaufman. H36-alpha 7 is a novel integrin alpha chain that is developmentally regulated during skeletal myogenesis. Journal of Cell Biology, 117(3):643–??, May 1992. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/117/3/
643.

Su:1991:GPA

[SWFB91] B. Su, G. L. Waneck, R. A. Flavell, and A. L. Bothwell. The glycosyl phosphatidylinositol anchor is critical for Ly-6A/E-mediated T cell activation. Journal of Cell Biology, 112(3): 377–??, February 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/112/3/377.

Sorkin:1991:ERK

[SWHCW91] A. Sorkin, B. Westermark, C. H. Heldin, and L. Claesson-Welsh. Effect of receptor kinase inactivation on the rate of internalization and degradation of PDGF and the PDGF beta-receptor. Journal of Cell Biology, 112(3):469–??, February 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/112/3/469.

Seiffert:1990:SDV

[SWL90] D. Seiffert, N. N. Wagner, and D. J. Loskutoff. Serumderived vitronectin influences the pericellular distribution of type 1 plasminogen activator inhibitor. Journal of Cell Biology, 111(3):1283–??, September 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/111/3/1283.

Salminen:1992:MFP

[SWL^{+}92] A. Salminen, J. M. Wahlberg, M. Lobigs, P. Liljeström, and H. Garoff. Membrane fusion process of Semliki Forest virus. II: Cleavage-dependent reorganization of the spike protein complex controls virus entry. Journal of Cell Biology, 116(2):349–??, January 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/116/2/349.

Shieh:1992:CSR

[SWM^{+}92] M. T. Shieh, D. WuDunn, R. I. Montgomery, J. D. Esko, and P. G. Spear. Cell surface receptors for herpes simplex virus are heparan sulfate proteoglycans. Journal of Cell Biology, 116(5):1273–??, March 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/116/5/1273.

Silveira:1990:YCD

[SWMS90] L. A. Silveira, D. H. Wong, F. R. Masiarz, and R. Schekman. Yeast clathrin has a distinctive light chain that is important for cell growth. Journal of Cell Biology, 111(4): 1437–??, October 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/111/4/1437.

Sabolic:1992:AEI

[SWS^{+}92] I. Sabolic, F. Wuarin, L. B. Shi, A. S. Verkman, D. A. Ausiello, S. Gluck, and D. Brown. Apical endosomes isolated from kidney collecting duct principal cells lack subunits of the proton pumping ATPase. Journal of Cell Biology, 119(1):111–??, October 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/119/1/111.

Shan:1994:YNE

[SXM94] X. Shan, Z. Xue, and T. Mélèse. Yeast NPI46 encodes a novel prolyl cis-trans isomerase that is located in the nucleolus. Journal of Cell Biology, 126(4):853–??, August 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/126/4/

Schittny:1990:TSA

[SY90a] J. C. Schittny and P. D. Yurchenco. Terminal short arm domains of basement membrane laminin are critical for its self-assembly. Journal of Cell Biology, 110(3):825–??, March 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/110/3/825.

Southwick:1990:ARP

[SY90b] F. S. Southwick and C. L. Young. The actin released from profilin–actin complexes is insufficient to account for the increase in F-actin in chemoattractant-stimulated polymorphonuclear leukocytes. Journal of Cell Biology, 110(6): 1965–??, June 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/110/6/1965.

Stewart:1991:RUF

[SY91] L. C. Stewart and M. P. Yaffe. A role for unsaturated fatty acids in mitochondrial movement and inheritance. Journal of Cell Biology, 115(5):1249–??, December 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/115/5/1249.

Shiozaki:1992:FDP

[SY92] K. Shiozaki and M. Yanagida. Functional dissection of the phosphorylated termini of fission yeast DNA topoisomerase II. Journal of Cell Biology, 119(5):1023–??, December 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/119/5/
1023.

Samejima:1994:BAF

[SY94a] I. Samejima and M. Yanagida. Bypassing anaphase by fission yeast cut9 mutation: requirement of cut9+ to initiate anaphase. Journal of Cell Biology, 127(6):1655–??, December 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/127/6/1655.

Sogo:1994:RMM

[SY94b] L. F. Sogo and M. P. Yaffe. Regulation of mitochondrial morphology and inheritance by Mdm10p, a protein of the mitochondrial outer membrane. Journal of Cell Biology, 126(6):1361–??, September 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/126/6/1361.

Smart:1994:CMC

[SYCA94] E. J. Smart, Y. S. Ying, P. A. Conrad, and R. G. Anderson. Caveolin moves from caveolae to the Golgi apparatus in response to cholesterol oxidation. Journal of Cell Biology, 127(5):1185–??, December 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/127/5/1185.

Sha:1991:IAD

[SYG91] X. Sha, L. Yang, and L. E. Gentry. Identification and analysis of discrete functional domains in the pro region of prepro-transforming growth factor beta 1. Journal of Cell Biology, 114(4):827–??, August 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/114/4/827.

Sato:1991:RBE

[SYO^{+}91] N. Sato, S. Yonemura, T. Obinata, S. Tsukita, and S. Tsukita. Radixin, a barbed end-capping actinmodulating protein, is concentrated at the cleavage furrow during cytokinesis. Journal of Cell Biology, 113(2):321–??, April 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/113/2/321.

Saxon:1994:APK

[SZB94a] M. L. Saxon, X. Zhao, and J. D. Black. Activation of protein kinase C isozymes is associated with post-mitotic events in intestinal epithelial cells in situ. Journal of Cell Biology, 126(3):747–??, August 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/126/3/747.

Somasundaran:1994:LHR

[SZB^{+}94b] M. Somasundaran, M. L. Zapp, L. K. Beattie, L. Pang, K. S. Byron, G. J. Bassell, J. L. Sullivan, and R. H. Singer. Localization of HIV RNA in mitochondria of infected cells: potential role in cytopathogenicity. Journal of Cell Biology, 126(6):1353–??, September 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/126/6/1353.

Schoenenberger:1991:MLA

[SZKM91] C. A. Schoenenberger, A. Zuk, D. Kendall, and K. S. Matlin. Multilayering and loss of apical polarity in MDCK cells transformed with viral K-ras. Journal of Cell Biology, 112 (5):873–??, March 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/112/5/873.

Sjaastad:1993:HRP

[SZP^{+}93] M. D. Sjaastad, K. S. Zettl, G. Parry, G. L. Firestone, and T. E. Machen. Hormonal regulation of the polarized function and distribution of Na/H exchange and Na/HCO3 cotransport in cultured mammary epithelial cells. Journal of Cell Biology, 122(3):589–??, August 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/122/3/589.

Turksen:1991:PNI

[TA91] K. Turksen and J. E. Aubin. Positive and negative immunoselection for enrichment of two classes of osteoprogenitor cells. Journal of Cell Biology, 114(2):373–??, July 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/114/2/
373.

Taillon:1992:MAC

[TASJ92] B. E. Taillon, S. A. Adler, J. P. Suhan, and J. W. Jarvik. Mutational analysis of centrin: an EF-hand protein associated with three distinct contractile fibers in the basal body apparatus of Chlamydomonas. Journal of Cell Biology, 119(6):1613–??, December 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/119/6/1613.

Turley:1991:HCA

[TAVC91] E. A. Turley, L. Austen, K. Vandeligt, and C. Clary. Hyaluronan and a cell-associated hyaluronan binding protein regulate the locomotion of ras-transformed cells. Journal of Cell Biology, 112(5):1041–??, March 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/112/5/1041.

Thrift:1991:NMP

[TAWJ91] R. N. Thrift, D. W. Andrews, P. Walter, and A. E. Johnson. A nascent membrane protein is located adjacent to ER membrane proteins throughout its integration and translation. Journal of Cell Biology, 112(5):809–??, March 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/112/5/
809.

Thomas:1990:SBP

[TB90] L. Thomas and H. Betz. Synaptophysin binds to physophilin, a putative synaptic plasma membrane protein. Journal of Cell Biology, 111(5):2041–??, November 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/111/5/
2041.

Tan:1992:RTG

[TBFP92] A. Tan, J. Bolscher, C. Feltkamp, and H. Ploegh. Retrograde transport from the Golgi region to the endoplasmic reticulum is sensitive to GTP gamma S. Journal of Cell Biology, 116(6):1357–??, March 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/116/6/1357.

Tisdale:1992:GBM

[TBKF^{+}92] E. J. Tisdale, J. R. Bourne, R. Khosravi-Far, C. J. Der, and W. E. Balch. GTP-binding mutants of rab1 and rab2 are potent inhibitors of vesicular transport from the endoplasmic reticulum to the Golgi complex. Journal of Cell Biology, 119(4):749–??, November 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/119/4/749.

Tartare:1992:AIE

[TBL^{+}92] S. Tartare, R. Ballotti, R. Lammers, C. Filloux, A. Chauvel, J. Schlessinger, A. Ullrich, and E. Van Obberghen. Activation of insulin-epidermal growth factor (EGF) receptor chimerae regulates EGF receptor binding affinity. Journal of Cell Biology, 116(3):627–??, February 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/116/3/627.

Takeshige:1992:AYD

[TBT^{+}92] K. Takeshige, M. Baba, S. Tsuboi, T. Noda, and Y. Ohsumi. Autophagy in yeast demonstrated with proteinase-deficient mutants and conditions for its induction. Journal of Cell Biology, 119(2):301–??, October 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/119/2/301.

Thomas:1992:CRT

[TBVS92] L. Thomas, H. R. Byers, J. Vink, and I. Stamenkovic. CD44H regulates tumor cell migration on hyaluronatecoated substrate. Journal of Cell Biology, 118(4):971–??, August 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/118/4/971.

Talhouk:1992:CEE

[TBW92] R. S. Talhouk, M. J. Bissell, and Z. Werb. Coordinated expression of extracellular matrix-degrading proteinases and their inhibitors regulates mammary epithelial function during involution. Journal of Cell Biology, 118(5):1271–??, September 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/118/5/1271.

Tilney:1991:SPR

[TCCT91] L. G. Tilney, T. J. Cooke, P. S. Connelly, and M. S. Tilney. The structure of plasmodesmata as revealed by plasmolysis, detergent extraction, and protease digestion. Journal of Cell Biology, 112(4):739–??, February 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/112/4/739.

Tang:1993:PKC

[TCDH93] D. G. Tang, Y. Q. Chen, C. A. Diglio, and K. V. Honn. Protein kinase C-dependent effects of 12(S)-HETE on endothelial cell vitronectin receptor and fibronectin receptor. Journal of Cell Biology, 121(3):689–??, May 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/121/3/689.

Tilney:1990:AFN

[TCP90] L. G. Tilney, P. S. Connelly, and D. A. Portnoy. Actin filament nucleation by the bacterial pathogen, Listeria monocytogenes. Journal of Cell Biology, 111(6):2979–??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/111/6/2979.

Tomkiel:1994:CCR

[TCS^{+}94] J. Tomkiel, C. A. Cooke, H. Saitoh, R. L. Bernat, and W. C. Earnshaw. CENP-c is required for maintaining proper kinetochore size and for a timely transition to anaphase. Journal of Cell Biology, 125(3):531–??, May 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/125/3/531.

Tan:1993:CFI

[TDSP93] P. K. Tan, N. G. Davis, G. F. Sprague, and G. S. Payne. Clathrin facilitates the internalization of seven transmembrane segment receptors for mating pheromones in yeast. Journal of Cell Biology, 123(6):1707–??, December 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/123/6/
1707.

Tilney:1992:HLEa

[TDT92] L. G. Tilney, D. J. DeRosier, and M. S. Tilney. How Listeria exploits host cell actin to form its own cytoskeleton. I. Formation of a tail and how that tail might be involved in movement. Journal of Cell Biology, 118(1):71–??, July 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/118/1/
71.

Tilney:1992:HLEb

[TDWT92] L. G. Tilney, D. J. DeRosier, A. Weber, and M. S. Tilney. How Listeria exploits host cell actin to form its own cytoskeleton. II. Nucleation, actin filament polarity, filament assembly, and evidence for a pointed end capper. Journal of Cell Biology, 118(1):83–??, July 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/118/1/83.

Takahashi:1994:IEK

[TFC94] K. Takahashi, J. Folmer, and P. A. Coulombe. Increased expression of keratin 16 causes anomalies in cytoarchitecture and keratinization in transgenic mouse skin. Journal of Cell Biology, 127(2):505–??, October 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/127/2/505.

Turner:1991:ICI

[TFDS91] P. R. Turner, P. Y. Fong, W. F. Denetclaw, and R. A. Steinhardt. Increased calcium influx in dystrophic muscle. Journal of Cell Biology, 115(6):1701–??, December 1991. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/115/6/
1701.

Townsley:1994:RHP

[TFP94] F. M. Townsley, G. Frigerio, and H. R. Pelham. Retrieval of HDEL proteins is required for growth of yeast cells. Journal of Cell Biology, 127(1):21–??, October 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/127/1/21.

Thorburn:1994:MAP

[TFT94] J. Thorburn, J. A. Frost, and A. Thorburn. Mitogenactivated protein kinases mediate changes in gene expression, but not cytoskeletal organization associated with cardiac muscle cell hypertrophy. Journal of Cell Biology, 126 (6):1565–??, September 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/126/6/1565.

Tooze:1991:CIS

[TFTH91] S. A. Tooze, T. Flatmark, J. Tooze, and W. B. Huttner. Characterization of the immature secretory granule, an intermediate in granule biogenesis. Journal of Cell Biology, 115(6):1491–??, December 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/115/6/1491.

Tseng:1994:ABA

[TG94] H. Tseng and H. Green. Association of basonuclin with ability of keratinocytes to multiply and with absence of terminal differentiation. Journal of Cell Biology, 126(2):495– ??, July 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/126/2/495.

Turner:1990:PNV

[TGB90] C. E. Turner, J. R. Glenney, and K. Burridge. Paxillin: a new vinculin-binding protein present in focal adhesions. Journal of Cell Biology, 111(3):1059–??, September 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/111/3/
1059.

Tyner:1990:OAF

[TGCT90] A. L. Tyner, R. Godbout, R. S. Compton, and S. M. Tilghman. The ontogeny of alpha-fetoprotein gene expression in the mouse gastrointestinal tract. Journal of Cell Biology, 110(4):915–??, April 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/110/4/915.

Thorens:1993:GSE

[TGD93] B. Thorens, N. Gérard, and N. Dériaz. GLUT2 surface expression and intracellular transport via the constitutive pathway in pancreatic beta cells and insulinoma: evidence for a block in trans-Golgi network exit by brefeldin A. Journal of Cell Biology, 123(6):1687–??, December 1993. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/123/6/
1687.

Tsim:1992:TAR

[TGR^{+}92] K. W. Tsim, I. Greenberg, M. Rimer, W. R. Randall, and M. M. Salpeter. Transcripts for the acetylcholine receptor and acetylcholine esterase show distribution differences in cultured chick muscle cells. Journal of Cell Biology, 118 (5):1201–??, September 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/118/5/1201.

Troy:1992:NOP

[TGS92] C. M. Troy, L. A. Greene, and M. L. Shelanski. Neurite outgrowth in peripherin-depleted PC12 cells. Journal of Cell Biology, 117(5):1085–??, June 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/117/5/1085.

Thaler:1990:ROT

[TH90] C. D. Thaler and L. T. Haimo. Regulation of organelle transport in melanophores by calcineurin. Journal of Cell Biology, 111(5):1939–??, November 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/111/5/1939.

Tooze:1991:TEE

[TH91] J. Tooze and M. Hollinshead. Tubular early endosomal networks in AtT20 and other cells. Journal of Cell Biology, 115(3):635–??, November 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/115/3/635.

Theurkauf:1992:MSA

[TH92a] W. E. Theurkauf and R. S. Hawley. Meiotic spindle assembly in Drosophila females: behavior of nonexchange chromosomes and the effects of mutations in the nod kinesinlike protein. Journal of Cell Biology, 116(5):1167–??, March 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/116/5/1167.

Tooze:1992:AHC

[TH92b] J. Tooze and M. Hollinshead. In AtT20 and HeLa cells brefeldin A induces the fusion of tubular endosomes and changes their distribution and some of their endocytic properties. Journal of Cell Biology, 118(4):813–??, August 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/118/4/
813.

Tzen:1992:SSP

[TH92c] J. T. Tzen and A. H. Huang. Surface structure and properties of plant seed oil bodies. Journal of Cell Biology, 117 (2):327–??, April 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/117/2/327.

Thomas:1994:LSC

[THBW94] J. L. Thomas, D. Holowka, B. Baird, and W. W. Webb. Large-scale co-aggregation of fluorescent lipid probes with cell surface proteins. Journal of Cell Biology, 125(4):795– ??, May 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/125/4/795.

Tschan:1990:RCC

[THH^{+}90] T. Tschan, I. Höerler, Y. Houze, K. H. Winterhalter, C. Richter, and P. Bruckner. Resting chondrocytes in culture survive without growth factors, but are sensitive to toxic oxygen metabolites. Journal of Cell Biology, 111(1): 257–??, July 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/111/1/257.

Tooze:1990:EPB

[THL^{+}90] J. Tooze, M. Hollinshead, T. Ludwig, K. Howell, B. Hoflack, and H. Kern. In exocrine pancreas, the basolateral endocytic pathway converges with the autophagic pathway immediately after the early endosome. Journal of Cell Biology, 111(2):329–??, August 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/111/2/329.

Tsao:1992:CTT

[TIA^{+}92] Y. S. Tsao, N. E. Ivessa, M. Adesnik, D. D. Sabatini, and G. Kreibich. Carboxy terminally truncated forms of ribophorin I are degraded in pre-Golgi compartments by a calcium-dependent process. Journal of Cell Biology, 116(1): 57–??, January 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/116/1/57.

Tse:1993:MFT

[TIA93] F. W. Tse, A. Iwata, and W. Almers. Membrane flux through the pore formed by a fusogenic viral envelope protein during cell fusion. Journal of Cell Biology, 121(3):543– ??, May 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/121/3/543.

Tsukita:1993:SJP

[TIN^{+}93] S. Tsukita, M. Itoh, A. Nagafuchi, S. Yonemura, and S. Tsukita. Submembranous junctional plaque proteins include potential tumor suppressor molecules. Journal of Cell Biology, 123(5):1049–??, December 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/123/5/1049.

Tavoloni:1994:VST

[TISH94] N. Tavoloni, H. Inoue, H. Sabe, and H. Hanafusa. v-src transformation of rat embryo fibroblasts. Inefficient conversion to anchorage-independent growth involves heterogeneity of primary cultures. Journal of Cell Biology, 126 (2):475–??, July 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/126/2/475.

Terasaki:1991:OSU

[TJ91] M. Terasaki and L. A. Jaffe. Organization of the sea urchin egg endoplasmic reticulum and its reorganization at fertilization. Journal of Cell Biology, 114(5):929–??, September 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/114/5/929.

Tanaka:1991:MBG

[TK91] E. M. Tanaka and M. W. Kirschner. Microtubule behavior in the growth cones of living neurons during axon elongation. Journal of Cell Biology, 115(2):345–??, October 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/115/2/
345.

Takada:1994:FRC

[TK94] S. Takada and R. Kamiya. Functional reconstitution of Chlamydomonas outer dynein arms from alpha-beta and gamma subunits: requirement of a third factor. Journal of Cell Biology, 126(3):737–??, August 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/126/3/737.

Taniura:1991:PCK

[TKHM91] H. Taniura, C. H. Kuo, Y. Hayashi, and N. Miki. Purification and characterization of an 82-kD membrane protein as a neurite outgrowth factor binding protein: possible involvement of NOF binding protein in axonal outgrowth in developing retina. Journal of Cell Biology, 112(2):313–??, January 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/112/2/313.

Tournier:1991:ILC

[TKP^{+}91] F. Tournier, S. Komesli, M. Paintrand, D. Job, and M. Bornens. The intercentriolar linkage is critical for the ability of heterologous centrosomes to induce parthenogenesis in Xenopus. Journal of Cell Biology, 113(6):1361–??, June 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/113/6/1361.

Taneja:1992:PRC

[TLFS92] K. L. Taneja, L. M. Lifshitz, F. S. Fay, and R. H. Singer. Poly(a) RNA codistribution with microfilaments: evaluation by in situ hybridization and quantitative digital imaging microscopy. Journal of Cell Biology, 119(5):1245–??, December 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/119/5/1245.

Tremble:1993:SSP

[TLSW93] P. M. Tremble, T. F. Lane, E. H. Sage, and Z. Werb. SPARC, a secreted protein associated with morphogenesis and tissue remodeling, induces expression of metalloproteinases in fibroblasts through a novel extracellular matrixdependent pathway. Journal of Cell Biology, 121(6):1433– ??, June 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/121/6/1433.

Thomas:1994:TMR

[TLWA94] P. Thomas, A. K. Lee, J. G. Wong, and W. Almers. A triggered mechanism retrieves membrane in seconds after Ca(2+)-stimulated exocytosis in single pituitary cells. Journal of Cell Biology, 124(5):667–??, March 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/124/5/667.

Tabas:1990:EBV

[TLXM90] I. Tabas, S. Lim, X. X. Xu, and F. R. Maxfield. Endocytosed beta–VLDL and LDL are delivered to different intracellular vesicles in mouse peritoneal macrophages. Journal of Cell Biology, 111(3):929–??, September 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/111/3/929.

Troschel:1990:DCF

[TM90] D. Troschel and M. Müller. Development of a cell-free system to study the membrane assembly of photosynthetic proteins of Rhodobacter capsulatus. Journal of Cell Biology, 111(1):87–??, July 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/111/1/87.

Theriot:1992:CAC

[TM92] J. A. Theriot and T. J. Mitchison. Comparison of actin and cell surface dynamics in motile fibroblasts. Journal of Cell Biology, 119(2):367–??, October 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/119/2/367.

Taipale:1994:LTG

[TMHKO94] J. Taipale, K. Miyazono, C. H. Heldin, and J. Keski-Oja. Latent transforming growth factor-beta 1 associates to fibroblast extracellular matrix via latent TGF-beta binding protein. Journal of Cell Biology, 124(1):171–??, January 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/124/1/
171.

Tabas:1991:IPS

[TMI^{+}91] I. Tabas, J. N. Myers, T. L. Innerarity, X. X. Xu, K. Arnold, J. Boyles, and F. R. Maxfield. The influence of particle size and multiple apoprotein E-receptor interactions on the endocytic targeting of beta–VLDL in mouse peritoneal macrophages. Journal of Cell Biology, 115(6):1547–??, December 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/115/6/1547.

Takada:1991:MCE

[TMP^{+}91] Y. Takada, E. Murphy, P. Pil, C. Chen, M. H. Ginsberg, and M. E. Hemler. Molecular cloning and expression of the cDNA for alpha 3 subunit of human alpha 3 beta 1 (VLA-3), an integrin receptor for fibronectin, laminin, and collagen. Journal of Cell Biology, 115(1):257–??, October 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/115/1/
257.

Tamura:1994:IDS

[TN94] M. Tamura and M. Noda. Identification of a DNA sequence involved in osteoblast-specific gene expression via interaction with helix-loop-helix (HLH)-type transcription factors. Journal of Cell Biology, 126(3):773–??, August 1994. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/126/3/
773.

Tomasetto:1993:SGJ

[TND^{+}93] C. Tomasetto, M. J. Neveu, J. Daley, P. K. Horan, and R. Sager. Specificity of gap junction communication among human mammary cells and connexin transfectants in culture. Journal of Cell Biology, 122(1):157–??, July 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/122/1/
157.

Takenaga:1994:BPP

[TNS^{+}94] K. Takenaga, Y. Nakamura, S. Sakiyama, Y. Hasegawa, K. Sato, and H. Endo. Binding of pEL98 protein, an S100related calcium-binding protein, to nonmuscle tropomyosin. Journal of Cell Biology, 124(5):757–??, March 1994. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/124/5/
757.

Tsuru:1990:REV

[TNT^{+}90] A. Tsuru, N. Nakamura, E. Takayama, Y. Suzuki, K. Hirayoshi, and K. Nagata. Regulation of the expression of vimentin gene during the differentiation of mouse myeloid leukemia cells. Journal of Cell Biology, 110(5):1655–??, May 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/110/5/1655.

Tsukita:1991:SPO

[TOA^{+}91] S. Tsukita, K. Oishi, T. Akiyama, Y. Yamanashi, T. Yamamoto, and S. Tsukita. Specific proto-oncogenic tyrosine kinases of src family are enriched in cell-to-cell adherens junctions where the level of tyrosine phosphorylation is elevated. Journal of Cell Biology, 113(4):867–??, May 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/113/4/
867.

Tanigawa:1993:HBG

[TOA^{+}93] G. Tanigawa, L. Orci, M. Amherdt, M. Ravazzola, J. B. Helms, and J. E. Rothman. Hydrolysis of bound GTP by ARF protein triggers uncoating of Golgi-derived COPcoated vesicles. Journal of Cell Biology, 123(6):1365–??, December 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/123/6/1365.

Takeda:1994:DDN

[TOFH94] S. Takeda, S. Okabe, T. Funakoshi, and N. Hirokawa. Differential dynamics of neurofilament–H protein and neurofilament–L protein in neurons. Journal of Cell Biology, 127(1):173–??, October 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/127/1/173.

Traub:1993:BDA

[TOK93] L. M. Traub, J. A. Ostrom, and S. Kornfeld. Biochemical dissection of AP-1 recruitment onto Golgi membranes. Journal of Cell Biology, 123(3):561–??, November 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/123/3/
561.

Tsukita:1994:EFM

[TOS^{+}94] S. Tsukita, K. Oishi, N. Sato, J. Sagara, A. Kawai, and S. Tsukita. ERM family members as molecular linkers between the cell surface glycoprotein CD44 and actinbased cytoskeletons. Journal of Cell Biology, 126(2):391– ??, July 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/126/2/391.

Tang:1990:HMP

[TQL^{+}90] T. K. Tang, Z. Qin, T. Leto, V. T. Marchesi, and E. J. Benz. Heterogeneity of mRNA and protein products arising from the protein 4.1 gene in erythroid and nonerythroid tissues. Journal of Cell Biology, 110(3):617–??, March 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/110/3/
617.

Tuszynski:1992:BAP

[TRD^{+}92] G. P. Tuszynski, V. L. Rothman, A. H. Deutch, B. K. Hamilton, and J. Eyal. Biological activities of peptides and peptide analogues derived from common sequences present in thrombospondin, properdin, and malarial proteins. Journal of Cell Biology, 116(1):209–??, January 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/116/1/209.

Turner:1992:PSB

[TRH^{+}92] M. D. Turner, M. E. Rennison, S. E. Handel, C. J. Wilde, and R. D. Burgoyne. Proteins are secreted by both constitutive and regulated secretory pathways in lactating mouse mammary epithelial cells. Journal of Cell Biology, 117(2): 269–??, April 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/117/2/269.

Taraboletti:1990:PTM

[TRLG90] G. Taraboletti, D. Roberts, L. A. Liotta, and R. Giavazzi. Platelet thrombospondin modulates endothelial cell adhesion, motility, and growth: a potential angiogenesis regulatory factor. Journal of Cell Biology, 111(2):765–??, August 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/111/2/765.

Tuszynski:1993:ICT

[TRP^{+}93] G. P. Tuszynski, V. L. Rothman, M. Papale, B. K. Hamilton, and J. Eyal. Identification and characterization of a tumor cell receptor for CSVTCG, a thrombospondin adhesive domain. Journal of Cell Biology, 120(2):513–??, January 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/120/2/513.

Tamura:1990:EIA

[TRS^{+}90] R. N. Tamura, C. Rozzo, L. Starr, J. Chambers, L. F. Reichardt, H. M. Cooper, and V. Quaranta. Epithelial integrin alpha 6 beta 4: complete primary structure of alpha 6 and variant forms of beta 4. Journal of Cell Biology, 111(4): 1593–??, October 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/111/4/1593.

Tsilibary:1990:IMC

[TRV^{+}90] E. C. Tsilibary, L. A. Reger, A. M. Vogel, G. G. Koliakos, S. S. Anderson, A. S. Charonis, J. N. Alegre, and L. T. Furcht. Identification of a multifunctional, cell-binding peptide sequence from the a1(NC1) of type IV collagen. Journal of Cell Biology, 111(4):1583–??, October 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/111/4/1583.

Terasaki:1991:DCU

[TS91] M. Terasaki and C. Sardet. Demonstration of calcium uptake and release by sea urchin egg cortical endoplasmic reticulum. Journal of Cell Biology, 115(4):1031–??, November 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/115/4/1031.

Tidball:1993:PSI

[TS93] J. G. Tidball and M. J. Spencer. PDGF stimulation induces phosphorylation of talin and cytoskeletal reorganization in skeletal muscle. Journal of Cell Biology, 123(3): 627–??, November 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/123/3/627.

Tassan:1994:CCA

[TSBN94] J. P. Tassan, S. J. Schultz, J. Bartek, and E. A. Nigg. Cell cycle analysis of the activity, subcellular localization, and subunit composition of human CAK (CDK-activating kinase). Journal of Cell Biology, 127(2):467–??, October 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/127/2/
467.

Tombes:1992:MEA

[TSBS92] R. M. Tombes, C. Simerly, G. G. Borisy, and G. Schatten. Meiosis, egg activation, and nuclear envelope breakdown are differentially reliant on Ca^{2+}, whereas germinal vesicle breakdown is Ca^{2+} independent in the mouse oocyte. Journal of Cell Biology, 117(4):799–??, May 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/117/4/799.

Tominaga:1993:IPI

[TSH^{+}93] T. Tominaga, K. Sugie, M. Hirata, N. Morii, J. Fukata, A. Uchida, H. Imura, and S. Narumiya. Inhibition of PMA-induced, LFA-1-dependent lymphocyte aggregation by ADP ribosylation of the small molecular weight GTP binding protein, rho. Journal of Cell Biology, 120(6):1529– ??, March 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/120/6/1529.

Takeuchi:1994:PCA

[TSK^{+}94] K. Takeuchi, N. Sato, H. Kasahara, N. Funayama, A. Nagafuchi, S. Yonemura, S. Tsukita, and S. Tsukita. Perturbation of cell adhesion and microvilli formation by antisense oligonucleotides to ERM family members. Journal of Cell Biology, 125(6):1371–??, June 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/125/6/1371.

True:1990:RSA

[TSLR90] D. D. True, M. S. Singer, L. A. Lasky, and S. D. Rosen. Requirement for sialic acid on the endothelial ligand of a lymphocyte homing receptor. Journal of Cell Biology, 111(6): 2757–??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/111/6/2757.

Taraboulos:1990:SPP

[TSP90] A. Taraboulos, D. Serban, and S. B. Prusiner. Scrapie prion proteins accumulate in the cytoplasm of persistently infected cultured cells. Journal of Cell Biology, 110(6):2117– ??, June 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/110/6/2117.

Tsuboi:1990:CCM

[TSR90] R. Tsuboi, Y. Sato, and D. B. Rifkin. Correlation of cell migration, cell invasion, receptor number, proteinase production, and basic fibroblast growth factor levels in endothelial cells. Journal of Cell Biology, 110(2):511–??, February 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/110/2/
511.

Tuma:1993:ICS

[TSR93] R. S. Tuma, J. A. Stolk, and M. B. Roth. Identification and characterization of a sphere organelle protein. Journal of Cell Biology, 122(4):767–??, August 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/122/4/767.

Tranter:1991:BAL

[TSS91] M. P. Tranter, S. P. Sugrue, and M. A. Schwartz. Binding of actin to liver cell membranes: the state of membrane-bound actin. Journal of Cell Biology, 112(5):891–??, March 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/112/5/
891.

Tozeren:1992:MAJ

[TSS^{+}92] A. Tözeren, K. L. Sung, L. A. Sung, M. L. Dustin, P. Y. Chan, T. A. Springer, and S. Chien. Micromanipulation of adhesion of a Jurkat cell to a planar bilayer membrane containing lymphocyte function-associated antigen 3 molecules. Journal of Cell Biology, 116(4):997–??, February 1992. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/116/4/

Tamm:1994:VCT

[TT94] S. L. Tamm and M. Terasaki. Visualization of calcium transients controlling orientation of ciliary beat. Journal of Cell Biology, 125(5):1127–??, June 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/125/5/1127.

Troyanovsky:1994:IPB

[TTE^{+}94] S. M. Troyanovsky, R. B. Troyanovsky, L. G. Eshkind, V. A. Krutovskikh, R. E. Leube, and W. W. Franke. Identification of the plakoglobin-binding domain in desmoglein and its role in plaque assembly and intermediate filament anchorage. Journal of Cell Biology, 127(1):151–??, October 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/127/1/
151.

Tilly:1990:HGF

[TTR^{+}90] B. C. Tilly, L. G. Tertoolen, R. Remorie, A. Ladoux, I. Verlaan, S. W. de Laat, and W. H. Moolenaar. Histamine as a growth factor and chemoattractant for human carcinoma and melanoma cells: action through Ca2(+)- mobilizing H1 receptors. Journal of Cell Biology, 110(4): 1211–??, April 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/110/4/1211.

Torri-Tarelli:1990:RSS

[TTVV^{+}90] F. Torri-Tarelli, A. Villa, F. Valtorta, P. De Camilli, P. Greengard, and B. Ceccarelli. Redistribution of synaptophysin and synapsin I during alpha-latrotoxin-induced release of neurotransmitter at the neuromuscular junction. Journal of Cell Biology, 110(2):449–??, February 1990. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/110/2/
449.

Toyoshima:1990:TDS

[TU90] C. Toyoshima and N. Unwin. Three-dimensional structure of the acetylcholine receptor by cryoelectron microscopy and helical image reconstruction. Journal of Cell Biology, 111 (6):2623–??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/111/6/2623.

Turner:1991:PMP

[Tur91] C. E. Turner. Paxillin is a major phosphotyrosinecontaining protein during embryonic development. Journal of Cell Biology, 115(1):201–??, October 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/115/1/201.

Tolsma:1993:PDT

[TVG^{+}93] S. S. Tolsma, O. V. Volpert, D. J. Good, W. A. Frazier, P. J. Polverini, and N. Bouck. Peptides derived from two separate domains of the matrix protein thrombospondin-1 have anti-angiogenic activity. Journal of Cell Biology, 122 (2):497–??, July 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/122/2/497.

Trimbur:1993:NLM

[TW93] G. M. Trimbur and C. J. Walsh. Nucleolus-like morphology produced during the in vitro reassociation of nucleolar components. Journal of Cell Biology, 122(4):753–??, August 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/122/4/753.

Tawil:1993:IPC

[TWC93] N. Tawil, P. Wilson, and S. Carbonetto. Integrins in point contacts mediate cell spreading: factors that regulate integrin accumulation in point contacts vs. focal contacts. Journal of Cell Biology, 120(1):261–??, January 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/120/1/261.

Trybus:1994:CAA

[TWC94] K. M. Trybus, G. S. Waller, and T. A. Chatman. Coupling of ATPase activity and motility in smooth muscle myosin is mediated by the regulatory light chain. Journal of Cell Biology, 124(6):963–??, March 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/124/6/963.

Tolkovsky:1990:CTR

[TWMS90] A. M. Tolkovsky, A. E. Walker, R. D. Murrell, and H. S. Suidan. Ca^{2+} transients are not required as signals for longterm neurite outgrowth from cultured sympathetic neurons. Journal of Cell Biology, 110(4):1295–??, April 1990. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/110/4/
1295.

Tang:1993:MCC

[TWQ^{+}93] B. L. Tang, S. H. Wong, X. L. Qi, S. H. Low, and W. Hong. Molecular cloning, characterization, subcellular localization and dynamics of p23, the mammalian KDEL receptor. Journal of Cell Biology, 120(2):325–??, January 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/120/2/325.

Turunen:1994:ECT

[TWV94] O. Turunen, T. Wahlström, and A. Vaheri. Ezrin has a COOH-terminal actin-binding site that is conserved in the ezrin protein family. Journal of Cell Biology, 126(6):1445– ??, September 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/126/6/1445.

Tsukamoto:1990:ICC

[TYF90] T. Tsukamoto, S. Yokota, and Y. Fujiki. Isolation and characterization of Chinese hamster ovary cell mutants defective in assembly of peroxisomes. Journal of Cell Biology, 110(3): 651–??, March 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/110/3/651.

Takada:1992:PMI

[TYM^{+}92] Y. Takada, J. Ylänne, D. Mandelman, W. Puzon, and M. H. Ginsberg. A point mutation of integrin beta 1 subunit blocks binding of alpha 5 beta 1 to fibronectin and invasin but not recruitment to adhesion plaques. Journal of Cell Biology, 119(4):913–??, November 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/119/4/913.

Toyoshima:1992:KMK

[TYSS92] I. Toyoshima, H. Yu, E. R. Steuer, and M. P. Sheetz. Kinectin, a major kinesin-binding protein on ER. Journal of Cell Biology, 118(5):1121–??, September 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/118/5/1121.

Tousson:1991:CNM

[TZBV91] A. Tousson, C. Zeng, B. R. Brinkley, and M. M. Valdivia. Centrophilin: a novel mitotic spindle protein involved in microtubule nucleation. Journal of Cell Biology, 112(3): 427–??, February 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/112/3/427.

Ulitzur:1992:LAE

[UHFG92] N. Ulitzur, A. Harel, N. Feinstein, and Y. Gruenbaum. Lamin activity is essential for nuclear envelope assembly in a Drosophila embryo cell-free extract. Journal of Cell Biology, 119(1):17–??, October 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/119/1/17.

Uehara:1992:EHH

[UK92] Y. Uehara and N. Kitamura. Expression of a human hepatocyte growth factor/scatter factor cDNA in MDCK epithelial cells influences cell morphology, motility, and anchorageindependent growth. Journal of Cell Biology, 117(4):889– ??, May 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/117/4/889.

Usuda:1991:DLP

[UKH^{+}91] N. Usuda, Y. Kong, M. Hagiwara, C. Uchida, M. Terasawa, T. Nagata, and H. Hidaka. Differential localization of protein kinase C isozymes in retinal neurons. Journal of Cell Biology, 112(6):1241–??, March 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/112/6/1241.

Ueda:1990:DOA

[UNY90] T. Ueda, T. Nakagaki, and T. Yamada. Dynamic organization of ATP and birefringent fibrils during free locomotion and galvanotaxis in the plasmodium of Physarum polycephalum. Journal of Cell Biology, 110(4):1097–??, April 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/110/4/1097.

Umeyama:1993:DMB

[UOKH93] T. Umeyama, S. Okabe, Y. Kanai, and N. Hirokawa. Dynamics of microtubules bundled by microtubule associated protein 2C (MAP2C). Journal of Cell Biology, 120(2): 451–??, January 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/120/2/451.

Usui:1991:UCC

[UYA^{+}91] T. Usui, M. Yoshida, K. Abe, H. Osada, K. Isono, and T. Beppu. Uncoupled cell cycle without mitosis induced by a protein kinase inhibitor, K-252a. Journal of Cell Biology, 115(5):1275–??, December 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/115/5/1275.

Vinson:1993:TDS

[VAL^{+}93] V. K. Vinson, S. J. Archer, E. E. Lattman, T. D. Pollard, and D. A. Torchia. Three-dimensional solution structure of Acanthamoeba profilin–I. Journal of Cell Biology, 122 (6):1277–??, September 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/122/6/1277.

Verkhovsky:1993:NSM

[VB93] A. B. Verkhovsky and G. G. Borisy. Non-sarcomeric mode of myosin II organization in the fibroblast lamellum. Journal of Cell Biology, 123(3):637–??, November 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/123/3/637.

Velez:1990:RDA

[VBA90] M. Velez, K. F. Barald, and D. Axelrod. Rotational diffusion of acetylcholine receptors on cultured rat myotubes. Journal of Cell Biology, 110(6):2049–??, June 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/110/6/2049.

Verde:1991:TIM

[VBAK91] F. Verde, J. M. Berrez, C. Antony, and E. Karsenti. Taxolinduced microtubule asters in mitotic extracts of Xenopus eggs: requirement for phosphorylated factors and cytoplasmic dynein. Journal of Cell Biology, 112(6):1177–??, March 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/112/6/1177.

Verbavatz:1993:TAC

[VBS^{+}93] J. M. Verbavatz, D. Brown, I. Sabolić, G. Valenti, D. A. Ausiello, A. N. Van Hoek, T. Ma, and A. S. Verkman. Tetrameric assembly of CHIP28 water channels in liposomes and cell membranes: a freeze-fracture study. Journal of Cell Biology, 123(3):605–??, November 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/123/3/605.

Vandekerckhove:1990:CTD

[VBV^{+}90] J. Vandekerckhove, G. Bauw, K. Vancompernolle, B. Honoré, and J. Celis. Comparative two-dimensional gel analysis and microsequencing identifies gelsolin as one of the most prominent downregulated markers of transformed human fibroblast and epithelial cells. Journal of Cell Biology, 111(1):95– ??, July 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/111/1/95.

VanVlasselaer:1994:IIT

[VBvG^{+}94] P. Van Vlasselaer, B. Borremans, U. van Gorp, J. R. Dasch, and R. De Waal-Malefyt. Interleukin 10 inhibits transforming growth factor-beta (TGF-beta) synthesis required for osteogenic commitment of mouse bone marrow cells. Journal of Cell Biology, 124(4):569–??, February 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/124/4/569.

Verderio:1994:SCC

[VCFM94] C. Verderio, S. Coco, G. Fumagalli, and M. Matteoli. Spatial changes in calcium signaling during the establishment of neuronal polarity and synaptogenesis. Journal of Cell Biology, 126(6):1527–??, September 1994. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/126/6/1527.

Vibert:1993:TDR

[VCL93] P. Vibert, R. Craig, and W. Lehman. Three-dimensional reconstruction of caldesmon-containing smooth muscle thin filaments. Journal of Cell Biology, 123(2):313–??, October 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/123/2/313.

vanderBliek:1993:MHD

[vdBRD^{+}93] A. M. van der Bliek, T. E. Redelmeier, H. Damke, E. J. Tisdale, E. M. Meyerowitz, and S. L. Schmid. Mutations in human dynamin block an intermediate stage in coated vesicle formation. Journal of Cell Biology, 122(3):553–??, August 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/122/3/553.

Verde:1992:CMD

[VDS^{+}92] F. Verde, M. Dogterom, E. Stelzer, E. Karsenti, and S. Leibler. Control of microtubule dynamics and length by cyclin A- and cyclin B-dependent kinases in Xenopus egg extracts. Journal of Cell Biology, 118(5):1097–??, September 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/118/5/1097.

vonDassow:1994:HAN

[vDS94] G. von Dassow and G. Schubiger. How an actin network might cause fountain streaming and nuclear migration in the syncytial Drosophila embryo. Journal of Cell Biology, 127(6):1637–??, December 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/127/6/1637.

vandeStolpe:1993:PHR

[vdSKL^{+}93] A. van de Stolpe, M. Karperien, C. W. Löwik, H. Jüppner, G. V. Segre, A. B. Abou-Samra, S. W. de Laat, and L. H. Defize. Parathyroid hormone-related peptide as an endogenous inducer of parietal endoderm differentiation. Journal of Cell Biology, 120(1):235–??, January 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/120/1/235.

Vitale:1991:CFA

[VDTT91] M. L. Vitale, A. Rodríguez Del Castillo, L. Tchakarov, and J. M. Trifaró. Cortical filamentous actin disassembly and scinderin redistribution during chromaffin cell stimulation precede exocytosis, a phenomenon not exhibited by gelsolin. Journal of Cell Biology, 113(5):1057–??, June 1991. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/113/5/
1057.

Kemenade:1992:ABL

[vdWvKvKdB^{+}92] E. van de Wiel-van Kemenade, Y. van Kooyk, A. J. de Boer, R. J. Huijbens, P. Weder, W. van de Kasteele, C. J. Melief, and C. G. Figdor. Adhesion of T and B lymphocytes to extracellular matrix and endothelial cells can be regulated through the beta subunit of VLA. Journal of Cell Biology, 117(2):461–??, April 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/117/2/461.

Voorberg:1991:ARW

[VFC^{+}91] J. Voorberg, R. Fontijn, J. Calafat, H. Janssen, J. A. van Mourik, and H. Pannekoek. Assembly and routing of von Willebrand factor variants: the requirements for disulfide-linked dimerization reside within the carboxyterminal 151 amino acids. Journal of Cell Biology, 113(1): 195–??, April 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/113/1/195.

Varnum-Finney:1994:VDP

[VFR94] B. Varnum-Finney and L. F. Reichardt. Vinculin-deficient PC12 cell lines extend unstable lamellipodia and filopodia and have a reduced rate of neurite outgrowth. Journal of Cell Biology, 127(4):1071–??, November 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/127/4/1071.

Vasquez:1994:XXE

[VGC94] R. J. Vasquez, D. L. Gard, and L. Cassimeris. XMAP from Xenopus eggs promotes rapid plus end assembly of microtubules and rapid microtubule polymer turnover. Journal of Cell Biology, 127(4):985–??, November 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/127/4/985.

Vida:1990:VRI

[VGE90] T. A. Vida, T. R. Graham, and S. D. Emr. In vitro reconstitution of intercompartmental protein transport to the yeast vacuole. Journal of Cell Biology, 111(6):2871–??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/111/6/2871.

Voos:1993:PMP

[VGG^{+}93] W. Voos, B. D. Gambill, B. Guiard, N. Pfanner, and E. A. Craig. Presequence and mature part of preproteins strongly influence the dependence of mitochondrial protein import on heat shock protein 70 in the matrix. Journal of Cell Biology, 123(1):119–??, October 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/123/1/119.

Vihtelic:1993:LDR

[VGM^{+}93] T. S. Vihtelic, M. Goebl, S. Milligan, J. E. O’Tousa, and D. R. Hyde. Localization of Drosophila retinal degeneration B, a membrane-associated phosphatidylinositol transfer protein. Journal of Cell Biology, 122(5):1013– ??, September 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/122/5/1013.

Valetti:1991:RBG

[VGMS91] C. Valetti, C. E. Grossi, C. Milstein, and R. Sitia. Russell bodies: a general response of secretory cells to synthesis of a mutant immunoglobulin which can neither exit from, nor be degraded in, the endoplasmic reticulum. Journal of Cell Biology, 115(4):983–??, November 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/115/4/983.

Vandromme:1992:SRF

[VGRC^{+}92] M. Vandromme, C. Gauthier-Rouvière, G. Carnac, N. Lamb, and A. Fernandez. Serum response factor p67SRF is expressed and required during myogenic differentiation of both mouse C2 and rat L6 muscle cell lines. Journal of Cell Biology, 118(6):1489–??, September 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/118/6/1489.

vanGenderen:1991:SLF

[vGvMS^{+}91] I. L. van Genderen, G. van Meer, J. W. Slot, H. J. Geuze, and W. F. Voorhout. Subcellular localization of Forssman glycolipid in epithelial MDCK cells by immunoelectronmicroscopy after freeze-substitution. Journal of Cell Biology, 115(4):1009–??, November 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/115/4/1009.

Verhey:1993:ICT

[VHB93] K. J. Verhey, S. F. Hausdorff, and M. J. Birnbaum. Identification of the carboxy terminus as important for the isoformspecific subcellular targeting of glucose transporter proteins. Journal of Cell Biology, 123(1):137–??, October 1993. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/123/1/
137.

Vida:1993:YVP

[VHE93] T. A. Vida, G. Huyer, and S. D. Emr. Yeast vacuolar proenzymes are sorted in the late Golgi complex and transported to the vacuole via a prevacuolar endosome-like compartment. Journal of Cell Biology, 121(6):1245–??, June 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/121/6/
1245.

Velasco:1993:CTD

[VHM^{+}93] A. Velasco, L. Hendricks, K. W. Moremen, D. R. Tulsiani, O. Touster, and M. G. Farquhar. Cell type-dependent variations in the subcellular distribution of alpha-mannosidase I and II. Journal of Cell Biology, 122(1):39–??, July 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/122/1/
39.

Vallen:1992:SDK

[VHSR92] E. A. Vallen, M. A. Hiller, T. Y. Scherson, and M. D. Rose. Separate domains of KAR1 mediate distinct functions in mitosis and nuclear fusion. Journal of Cell Biology, 117(6): 1277–??, June 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/117/6/1277.

Volkmer:1992:SAS

[VHW^{+}92] H. Volkmer, B. Hassel, J. M. Wolff, R. Frank, and F. G. Rathjen. Structure of the axonal surface recognition molecule neurofascin and its relationship to a neural subgroup of the immunoglobulin superfamily. Journal of Cell Biology, 118(1):149–??, July 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/118/1/149.

Valle:1993:MNE

[VIF^{+}93] P. Lu Valle, M. Iwamoto, P. Fanning, M. Pacifici, and B. R. Olsen. Multiple negative elements in a gene that codes for an extracellular matrix protein, collagen X, restrict expression to hypertrophic chondrocytes. Journal of Cell Biology, 121 (5):1173–??, June 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/121/5/1173.

VanEtten:1994:CTC

[VJB^{+}94] R. A. Van Etten, P. K. Jackson, D. Baltimore, M. C. Sanders, P. T. Matsudaira, and P. A. Janmey. The COOH terminus of the c–Abl tyrosine kinase contains distinct Fand G-actin binding domains with bundling activity. Journal of Cell Biology, 124(3):325–??, February 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/124/3/325.

Vaisberg:1993:CDP

[VKM93] E. A. Vaisberg, M. P. Koonce, and J. R. McIntosh. Cytoplasmic dynein plays a role in mammalian mitotic spindle formation. Journal of Cell Biology, 123(4):849–??, November 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/123/4/849.

Vandenberg:1991:CTI

[VKR^{+}91] P. Vandenberg, A. Kern, A. Ries, L. Luckenbill-Edds, K. Mann, and K. Kühn. Characterization of a type IV collagen major cell binding site with affinity to the alpha 1 beta 1 and the alpha 2 beta 1 integrins. Journal of Cell Biology, 113(6):1475–??, June 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/113/6/1475.

Vielmetter:1994:NAC

[VKRD94] J. Vielmetter, J. F. Kayyem, J. M. Roman, and W. J. Dreyer. Neogenin, an avian cell surface protein expressed during terminal neuronal differentiation, is closely related to the human tumor suppressor molecule deleted in colorectal cancer. Journal of Cell Biology, 127(6):2009–??, December 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/127/6/2009.

Vaananen:1990:EPP

[VKS^{+}90] H. K. Väänänen, E. K. Karhukorpi, K. Sundquist, B. Wallmark, I. Roininen, T. Hentunen, J. Tuukkanen, and P. Lakkakorpi. Evidence for the presence of a proton pump of the vacuolar H(+)-ATPase type in the ruffled borders of osteoclasts. Journal of Cell Biology, 111(3):1305– ??, September 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/111/3/1305.

vanKooyk:1991:ALT

[vKWH^{+}91] Y. van Kooyk, P. Weder, F. Hogervorst, A. J. Verhoeven, G. van Seventer, A. A. te Velde, J. Borst, G. D. Keizer, and C. G. Figdor. Activation of LFA-1 through a Ca2(+)- dependent epitope stimulates lymphocyte adhesion. Journal of Cell Biology, 112(2):345–??, January 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/112/2/345.

vanKooyk:1994:ECM

[vKWHF94] Y. van Kooyk, P. Weder, K. Heije, and C. G. Figdor. Extracellular Ca^{2+} modulates leukocyte function-associated antigen-1 cell surface distribution on T lymphocytes and consequently affects cell adhesion. Journal of Cell Biology, 124(6):1061–??, March 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/124/6/1061.

Vigers:1991:DVP

[VL91] G. P. Vigers and M. J. Lohka. A distinct vesicle population targets membranes and pore complexes to the nuclear envelope in Xenopus eggs. Journal of Cell Biology, 112(4): 545–??, February 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/112/4/545.

Vikstrom:1992:SSD

[VLGB92] K. L. Vikstrom, S. S. Lim, R. D. Goldman, and G. G. Borisy. Steady state dynamics of intermediate filament networks. Journal of Cell Biology, 118(1):121–??, July 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/118/1/121.

Vogel:1993:NIS

[VLH^{+}93] B. E. Vogel, S. J. Lee, A. Hildebrand, W. Craig, M. D. Pierschbacher, F. Wong-Staal, and E. Ruoslahti. A novel integrin specificity exemplified by binding of the alpha v beta 5 integrin to the basic domain of the HIV Tat protein and vitronectin. Journal of Cell Biology, 121(2):461–??, April 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/121/2/
461.

Vidugiriene:1993:ELI

[VM93] J. Vidugiriene and A. K. Menon. Early lipid intermediates in glycosyl-phosphatidylinositol anchor assembly are synthesized in the ER and located in the cytoplasmic leaflet of the ER membrane bilayer. Journal of Cell Biology, 121 (5):987–??, June 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/121/5/987.

Vidugiriene:1994:GAC

[VM94] J. Vidugiriene and A. K. Menon. The GPI anchor of cellsurface proteins is synthesized on the cytoplasmic face of the endoplasmic reticulum. Journal of Cell Biology, 127(2): 333–??, October 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/127/2/333.

Vale:1992:DIM

[VMB92] R. D. Vale, F. Malik, and D. Brown. Directional instability of microtubule transport in the presence of kinesin and dynein, two opposite polarity motor proteins. Journal of Cell Biology, 119(6):1589–??, December 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/119/6/1589.

Vogel:1990:LBG

[VMR90] J. P. Vogel, L. M. Misra, and M. D. Rose. Loss of BiP/ GRP78 function blocks translocation of secretory proteins in yeast. Journal of Cell Biology, 110(6):1885–??, June 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/110/6/
1885.

Valdivieso:1991:CGR

[VMS^{+}91] M. H. Valdivieso, P. C. Mol, J. A. Shaw, E. Cabib, and A. Durán. CAL1, a gene required for activity of chitin synthase 3 in Saccharomyces cerevisiae. Journal of Cell Biology, 114(1):101–??, July 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/114/1/101.

Vuolteenaho:1994:HLM

[VNS^{+}94] R. Vuolteenaho, M. Nissinen, K. Sainio, M. Byers, R. Eddy, H. Hirvonen, T. B. Shows, H. Sariola, E. Engvall, and K. Tryggvason. Human laminin M chain (merosin): complete primary structure, chromosomal assignment, and expression of the M and A chain in human fetal tissues. Journal of Cell Biology, 124(3):381–??, February 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/124/3/381.

Villa:1991:ICS

[VPC^{+}91] A. Villa, P. Podini, D. O. Clegg, T. Pozzan, and J. Meldolesi. Intracellular Ca^{2+} stores in chicken Purkinje neurons: differential distribution of the low affinity-high capacity Ca^{2+} binding protein, calsequestrin, of Ca^{2+} ATPase and of the ER lumenal protein, Bip. Journal of Cell Biology, 113(4):779–??, May 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/113/4/779.

Villa:1993:ESR

[VPP^{+}93] A. Villa, P. Podini, M. C. Panzeri, H. D. Söling, P. Volpe, and J. Meldolesi. The endoplasmic-sarcoplasmic reticulum of smooth muscle: immunocytochemistry of vas deferens fibers reveals specialized subcompartments differently equipped for the control of Ca^{2+} homeostasis. Journal of Cell Biology, 121(5):1041–??, June 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/121/5/1041.

Vance:1991:BML

[VPVC91] J. E. Vance, D. Pan, D. E. Vance, and R. B. Campenot. Biosynthesis of membrane lipids in rat axons. Journal of Cell Biology, 115(4):1061–??, November 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/115/4/1061.

Vestal:1992:GPA

[VR92] D. J. Vestal and B. Ranscht. Glycosyl phosphatidylinositol– anchored T-cadherin mediates calcium-dependent, homophilic cell adhesion. Journal of Cell Biology, 119(2): 451–??, October 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/119/2/451.

Vater:1992:VPH

[VRE^{+}92] C. A. Vater, C. K. Raymond, K. Ekena, I. Howald-Stevenson, and T. H. Stevens. The VPS1 protein, a homolog of dynamin required for vacuolar protein sorting in Saccharomyces cerevisiae, is a GTPase with two functionally separable domains. Journal of Cell Biology, 119(4): 773–??, November 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/119/4/773.

Vey:1994:MTG

[VSB^{+}94] M. Vey, W. Schäfer, S. Berghöfer, H. D. Klenk, and W. Garten. Maturation of the trans-Golgi network protease furin: compartmentalization of propeptide removal, substrate cleavage, and COOH-terminal truncation. Journal of Cell Biology, 127(6):1829–??, December 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/127/6/1829.

Vanderhaeghen:1993:ORD

[VSVP93a] P. Vanderhaeghen, S. Schurmans, G. Vassart, and M. Parmentier. Olfactory receptors are displayed on dog mature sperm cells. Journal of Cell Biology, 123(6):1441–??, December 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/123/6/1441.

Vigoreaux:1993:FNM

[VSVP93b] J. O. Vigoreaux, J. D. Saide, K. Valgeirsdottir, and M. L. Pardue. Flightin, a novel myofibrillar protein of Drosophila stretch-activated muscles. Journal of Cell Biology, 121(3): 587–??, May 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/121/3/587.

Hof:1990:GLP

[vtHvM90] W. van ’t Hof and G. van Meer. Generation of lipid polarity in intestinal epithelial (Caco-2) cells: sphingolipid synthesis in the Golgi complex and sorting before vesicular traffic to the plasma membrane. Journal of Cell Biology, 111(3):977– ??, September 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/111/3/977.

Vonderheide:1994:RWC

[VTSS94] R. H. Vonderheide, T. F. Tedder, T. A. Springer, and D. E. Staunton. Residues within a conserved amino acid motif of domains 1 and 4 of VCAM-1 are required for binding to VLA-4. Journal of Cell Biology, 125(1):215–??, April 1994.

CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/125/1/
215.

VanderLeij:1992:IPA

[VVB^{+}92] I. Van der Leij, M. Van den Berg, R. Boot, M. Franse, B. Distel, and H. F. Tabak. Isolation of peroxisome assembly mutants from Saccharomyces cerevisiae with different morphologies using a novel positive selection procedure. Journal of Cell Biology, 119(1):153–??, October 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/119/1/153.

Veneziani:1990:TSP

[VVR^{+}90] B. M. Veneziani, G. Villone, R. Romano, A. Di Carlo, C. Garbi, and D. Tramontano. The tissue-specific pathways regulating cell proliferation are inherited independently in somatic hybrid between thyroid and liver cells. Journal of Cell Biology, 111(6):2703–??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/111/6/2703.

Valls:1990:YCV

[VWS90] L. A. Valls, J. R. Winther, and T. H. Stevens. Yeast carboxypeptidase Y vacuolar targeting signal is defined by four propeptide amino acids. Journal of Cell Biology, 111(2): 361–??, August 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/111/2/361.

Vaidya:1991:IMH

[VWT^{+}91] T. B. Vaidya, C. M. Weyman, D. Teegarden, C. L. Ashendel, and E. J. Taparowsky. Inhibition of myogenesis by the H-ras oncogene: implication of a role for protein kinase C. Journal of Cell Biology, 114(4):809–??, August 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/114/4/809.

Veit:1993:MIV

[VYM93] B. Veit, J. K. Yucel, and V. Malhotra. Microtubule independent vesiculation of Golgi membranes and the reassembly of vesicles into Golgi stacks. Journal of Cell Biology, 122(6):1197–??, September 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/122/6/1197.

vanZee:1993:CAN

[vZDF^{+}93] K. van Zee, A. Dickmanns, U. Fischer, R. Lührmann, and E. Fanning. A cytoplasmically anchored nuclear protein interferes specifically with the import of nuclear proteins but not U1 snRNA. Journal of Cell Biology, 121(2):229–??, April 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/121/2/229.

Wong:1993:TPG

[WA93] L. L. Wong and P. N. Adler. Tissue polarity genes of Drosophila regulate the subcellular location for prehair initiation in pupal wing cells. Journal of Cell Biology, 123(1): 209–??, October 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/123/1/209.

Wallace:1994:SIA

[Wal94] B. G. Wallace. Staurosporine inhibits agrin-induced acetylcholine receptor phosphorylation and aggregation. Journal of Cell Biology, 125(3):661–??, May 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/125/3/661.

Walton:1991:RIT

[WAS^{+}91] P. D. Walton, J. A. Airey, J. L. Sutko, C. F. Beck, G. A. Mignery, T. C. Südhof, T. J. Deerinck, and M. H. Ellisman. Ryanodine and inositol trisphosphate receptors coexist in avian cerebellar Purkinje neurons. Journal of Cell Biology, 113(5):1145–??, June 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/113/5/1145.

Wattenberg:1990:GGT

[Wat90] B. W. Wattenberg. Glycolipid and glycoprotein transport through the Golgi complex are similar biochemically and kinetically. Reconstitution of glycolipid transport in a cell free system. Journal of Cell Biology, 111(2):421–??, August 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/111/2/421.

Wong:1992:KPG

[WB92a] D. H. Wong and F. M. Brodsky. 100-kD proteins of Golgiand trans-Golgi network-associated coated vesicles have related but distinct membrane binding properties. Journal of Cell Biology, 117(6):1171–??, June 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/117/6/1171.

Wood:1992:MFE

[WB92b] S. A. Wood and W. J. Brown. The morphology but not the function of endosomes and lysosomes is altered by brefeldin A. Journal of Cell Biology, 119(2):273–??, October 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/119/2/
273.

Wozniak:1992:STS

[WB92c] R. W. Wozniak and G. Blobel. The single transmembrane segment of gp210 is sufficient for sorting to the pore membrane domain of the nuclear envelope. Journal of Cell Biology, 119(6):1441–??, December 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/119/6/1441.

Wente:1993:TSN

[WB93a] S. R. Wente and G. Blobel. A temperature-sensitive NUP116 null mutant forms a nuclear envelope seal over the yeast nuclear pore complex thereby blocking nucleocytoplasmic traffic. Journal of Cell Biology, 123(2):275–??, October 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/123/2/275.

Wu:1993:ENG

[WB93b] Y. Y. Wu and R. A. Bradshaw. Effect of nerve growth factor and fibroblast growth factor on PC12 cells: inhibition by orthovanadate. Journal of Cell Biology, 121(2):409–??, April 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/121/2/409.

Wente:1994:NEN

[WB94] S. R. Wente and G. Blobel. NUP145 encodes a novel yeast glycine-leucine-phenylalanine-glycine (GLFG) nucleoporin required for nuclear envelope structure. Journal of Cell Biology, 125(5):955–??, June 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/125/5/955.

Wagner:1992:TDS

[WBA92] M. C. Wagner, B. Barylko, and J. P. Albanesi. Tissue distribution and subcellular localization of mammalian myosin I. Journal of Cell Biology, 119(1):163–??, October 1992. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/119/1/
163.

Walmsley:1991:CRC

[WBAP91] M. E. Walmsley, R. S. Buckle, J. Allan, and R. K. Patient. A chicken red cell inhibitor of transcription associated with the terminally differentiated state. Journal of Cell Biology, 114(1):9–??, July 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/114/1/9.

Weller:1991:AAS

[WBE91] A. Weller, S. Beck, and P. Ekblom. Amino acid sequence of mouse tenascin and differential expression of two tenascin isoforms during embryogenesis. Journal of Cell Biology, 112 (2):355–??, January 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/112/2/355.

Wiest:1990:MBD

[WBH^{+}90] D. L. Wiest, J. K. Burkhardt, S. Hester, M. Hortsch, D. I. Meyer, and Y. Argon. Membrane biogenesis during B cell differentiation: most endoplasmic reticulum proteins are expressed coordinately. Journal of Cell Biology, 110(5):1501– ??, May 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/110/5/1501.

Wiche:1991:CSR

[WBL^{+}91] G. Wiche, B. Becker, K. Luber, G. Weitzer, M. J. Castañon, R. Hauptmann, C. Stratowa, and M. Stewart. Cloning and sequencing of rat plectin indicates a 466-kD polypeptide chain with a three-domain structure based on a central alpha-helical coiled coil. Journal of Cell Biology, 114 (1):83–??, July 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/114/1/83.

Winkelmann:1991:TDS

[WBR91] D. A. Winkelmann, T. S. Baker, and I. Rayment. Threedimensional structure of myosin subfragment-1 from electron microscopy of sectioned crystals. Journal of Cell Biology, 114(4):701–??, August 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/114/4/701.

Wozniak:1994:PIP

[WBR94] R. W. Wozniak, G. Blobel, and M. P. Rout. POM152 is an integral protein of the pore membrane domain of the yeast nuclear envelope. Journal of Cell Biology, 125(1):31– ??, April 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/125/1/31.

Wienhues:1991:PFC

[WBS^{+}91a] U. Wienhues, K. Becker, M. Schleyer, B. Guiard, M. Tropschug, A. L. Horwich, N. Pfanner, and W. Neupert. Protein folding causes an arrest of preprotein translocation into mitochondria in vivo. Journal of Cell Biology, 115(6):1601– ??, December 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/115/6/1601.

Witzemann:1991:NFR

[WBS91b] V. Witzemann, H. R. Brenner, and B. Sakmann. Neural factors regulate AChR subunit mRNAs at rat neuromuscular synapses. Journal of Cell Biology, 114(1):125–??, July 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/114/1/

Weber:1994:EPT

[WBT^{+}94] E. Weber, G. Berta, A. Tousson, P. St John, M. W. Green, U. Gopalokrishnan, T. Jilling, E. J. Sorscher, T. S. Elton, and D. R. Abrahamson. Expression and polarized targeting of a rab3 isoform in epithelial cells. Journal of Cell Biology, 125(3):583–??, May 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/125/3/583.

Weidner:1990:SFM

[WBVB90] K. M. Weidner, J. Behrens, J. Vandekerckhove, and W. Birchmeier. Scatter factor: molecular characteristics and effect on the invasiveness of epithelial cells. Journal of Cell Biology, 111(5):2097–??, November 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/111/5/2097.

White:1994:SIA

[WBW^{+}94] T. W. White, R. Bruzzone, S. Wolfram, D. L. Paul, and D. A. Goodenough. Selective interactions among the multiple connexin proteins expressed in the vertebrate lens: the second extracellular domain is a determinant of compatibility between connexins. Journal of Cell Biology, 125(4): 879–??, May 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/125/4/879.

Wall:1993:BVL

[WBWH93] N. A. Wall, M. Blessing, C. V. Wright, and B. L. Hogan. Biosynthesis and in vivo localization of the decapentaplegic– Vg–related protein, DVR-6 (bone morphogenetic protein-6). Journal of Cell Biology, 120(2):493–??, January 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/120/2/
493.

Widmer:1990:ILP

[WC90a] F. Widmer and P. Caroni. Identification, localization, and primary structure of CAP-23, a particle-bound cytosolic protein of early development. Journal of Cell Biology, 111 (6):3035–??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/111/6/3035.

Wong:1990:CDR

[WC90b] P. C. Wong and D. W. Cleveland. Characterization of dominant and recessive assembly-defective mutations in mouse neurofilament NF–M. Journal of Cell Biology, 111(5):1987– ??, November 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/111/5/1987.

Widmer:1993:PSM

[WC93] F. Widmer and P. Caroni. Phosphorylation-site mutagenesis of the growth-associated protein GAP-43 modulates its effects on cell spreading and morphology. Journal of Cell Biology, 120(2):503–??, January 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/120/2/503.

Wileman:1990:GES

[WCC^{+}90] T. Wileman, G. R. Carson, M. Concino, A. Ahmed, and C. Terhorst. The gamma and epsilon subunits of the CD3 complex inhibit pre-Golgi degradation of newly synthesized T cell antigen receptors. Journal of Cell Biology, 110(4): 973–??, April 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/110/4/973.

Wilson:1992:RKK

[WCF92] A. K. Wilson, P. A. Coulombe, and E. Fuchs. The roles of K5 and K14 head, tail, and R/K L L E G E domains in keratin filament assembly in vitro. Journal of Cell Biology, 119(2):401–??, October 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/119/2/401.

Whitters:1993:SIM

[WCM^{+}93] E. A. Whitters, A. E. Cleves, T. P. McGee, H. B. Skinner, and V. A. Bankaitis. SAC1p is an integral membrane protein that influences the cellular requirement for phospholipid transfer protein function and inositol in yeast. Journal of Cell Biology, 122(1):79–??, July 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/122/1/79.

Waters:1992:NKP

[WCR92] M. G. Waters, D. O. Clary, and J. E. Rothman. A novel 115-kD peripheral membrane protein is required for intercisternal transport in the Golgi stack. Journal of Cell Biology, 118(5):1015–??, September 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/118/5/1015.

Waters:1993:FPM

[WCR93] J. C. Waters, R. W. Cole, and C. L. Rieder. The forceproducing mechanism for centrosome separation during spindle formation in vertebrates is intrinsic to each aster. Journal of Cell Biology, 122(2):361–??, July 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/122/2/361.

Whalley:1991:PIC

[WCW91] T. Whalley, I. Crossley, and M. Whitaker. Phosphoprotein inhibition of calcium-stimulated exocytosis in sea urchin eggs. Journal of Cell Biology, 113(4):769–??, May 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/113/4/
769.

Wille:1992:ALP

[WDB^{+}92] H. Wille, G. Drewes, J. Biernat, E. M. Mandelkow, and E. Mandelkow. Alzheimer-like paired helical filaments and antiparallel dimers formed from microtubule-associated protein tau in vitro. Journal of Cell Biology, 118(3):573–??, August 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/118/3/573.

Wilson:1994:CNP

[WDP^{+}94] S. M. Wilson, K. V. Datar, M. R. Paddy, J. R. Swedlow, and M. S. Swanson. Characterization of nuclear polyadenylated RNA-binding proteins in Saccharomyces cerevisiae. Journal of Cell Biology, 127(5):1173–??, December 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/127/5/1173.

Wight:1993:MPP

[WDRM93] P. A. Wight, C. S. Duchala, C. Readhead, and W. B. Macklin. A myelin proteolipid protein–LacZ fusion protein is developmentally regulated and targeted to the myelin membrane in transgenic mice. Journal of Cell Biology, 123(2): 443–??, October 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/123/2/443.

Williams:1992:CIN

[WDT^{+}92] E. J. Williams, P. Doherty, G. Turner, R. A. Reid, J. J. Hemperly, and F. S. Walsh. Calcium influx into neurons can solely account for cell contact-dependent neurite outgrowth stimulated by transfected L1. Journal of Cell Biology, 119(4):883–??, November 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/119/4/883.

Warshaw:1990:SMM

[WDWT90] D. M. Warshaw, J. M. Desrosiers, S. S. Work, and K. M. Trybus. Smooth muscle myosin cross-bridge interactions modulate actin filament sliding velocity in vitro. Journal of Cell Biology, 111(2):453–??, August 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/111/2/453.

Wood:1990:MCC

[WE90] E. R. Wood and W. C. Earnshaw. Mitotic chromatin condensation in vitro using somatic cell extracts and nuclei with variable levels of endogenous topoisomerase II. Journal of Cell Biology, 111(6):2839–??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/111/6/2839.

West:1992:IPS

[WE92] C. M. West and G. W. Erdos. Incorporation of protein into spore coats is not cell autonomous in Dictyostelium. Journal of Cell Biology, 116(5):1291–??, March 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/116/5/1291.

Worman:1990:LBR

[WEB90] H. J. Worman, C. D. Evans, and G. Blobel. The lamin B receptor of the nuclear envelope inner membrane: a polytopic protein with eight potential transmembrane domains. Journal of Cell Biology, 111(4):1535–??, October 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/111/4/1535.

Williams:1990:AMG

[WF90] M. A. Williams and M. Fukuda. Accumulation of membrane glycoproteins in lysosomes requires a tyrosine residue at a particular position in the cytoplasmic tail. Journal of Cell Biology, 111(3):955–??, September 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/111/3/955.

Wilcox:1991:PPP

[WF91] C. A. Wilcox and R. S. Fuller. Posttranslational processing of the prohormone-cleaving Kex2 protease in the Saccharomyces cerevisiae secretory pathway. Journal of Cell Biology, 115(2):297–??, October 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/115/2/297.

Wassler:1993:PCH

[WF93] M. Wassler and E. Fries. Proteolytic cleavage of haptoglobin occurs in a subcompartment of the endoplasmic reticulum: evidence from membrane fusion in vitro. Journal of Cell Biology, 123(2):285–??, October 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/123/2/285.

Wong:1994:CDM

[WF94] M. H. Wong and M. T. Filbin. The cytoplasmic domain of the myelin P0 protein influences the adhesive interactions of its extracellular domain. Journal of Cell Biology, 126(4): 1089–??, August 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/126/4/1089.

Wickham:1994:IAV

[WFCN94] T. J. Wickham, E. J. Filardo, D. A. Cheresh, and G. R. Nemerow. Integrin alpha v beta 5 selectively promotes adenovirus mediated cell membrane permeabilization. Journal of Cell Biology, 127(1):257–??, October 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/127/1/257.

Weis:1991:ISD

[WFD^{+}91] J. Weis, S. M. Fine, C. David, S. Savarirayan, and J. R. Sanes. Integration site-dependent expression of a transgene reveals specialized features of cells associated with neuromuscular junctions. Journal of Cell Biology, 113(6):1385– ??, June 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/113/6/1385.

Wang:1992:MCS

[WFD92] L. Wang, Y. Feng, and J. L. Denburg. A multifunctional cell surface developmental stage-specific antigen in the cockroach embryo: involvement in pathfinding by CNS pioneer axons. Journal of Cell Biology, 118(1):163–??, July 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/118/1/
163.

Wahlberg:1992:MFP

[WG92a] J. M. Wahlberg and H. Garoff. Membrane fusion process of Semliki Forest virus. I: Low pH-induced rearrangement in spike protein quaternary structure precedes virus penetration into cells. Journal of Cell Biology, 116(2):339–??, January 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/116/2/339.

Wice:1992:SIC

[WG92b] B. M. Wice and J. I. Gordon. A strategy for isolation of cD-NAs encoding proteins affecting human intestinal epithelial cell growth and differentiation: characterization of a novel gut-specific N-myristoylated annexin. Journal of Cell Biology, 116(2):405–??, January 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/116/2/405.

Wu:1993:RTP

[WG93] D. Y. Wu and D. J. Goldberg. Regulated tyrosine phosphorylation at the tips of growth cone filopodia. Journal of Cell Biology, 123(3):653–??, November 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/123/3/653.

Winey:1991:MMN

[WGBB91] M. Winey, L. Goetsch, P. Baum, and B. Byers. MPS1 and MPS2: novel yeast genes defining distinct steps of spindle pole body duplication. Journal of Cell Biology, 114(4): 745–??, August 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/114/4/745.

Wilson:1991:IDM

[WGCdL91] A. K. Wilson, G. Gorgas, W. D. Claypool, and P. de Lanerolle. An increase or a decrease in myosin II phosphorylation inhibits macrophage motility. Journal of Cell Biology, 114 (2):277–??, July 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/114/2/277.

Wiche:1993:EPM

[WGD^{+}93] G. Wiche, D. Gromov, A. Donovan, M. J. Castañón, and E. Fuchs. Expression of plectin mutant cDNA in cultured cells indicates a role of COOH-terminal domain in intermediate filament association. Journal of Cell Biology, 121 (3):607–??, May 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/121/3/607.

Wayner:1993:ECE

[WGM^{+}93] E. A. Wayner, S. G. Gil, G. F. Murphy, M. S. Wilke, and W. G. Carter. Epiligrin, a component of epithelial basement membranes, is an adhesive ligand for alpha 3 beta 1 positive T lymphocytes. Journal of Cell Biology, 121(5): 1141–??, June 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/121/5/1141.

Weber:1993:SAS

[WGP^{+}93] K. Weber, N. Geisler, U. Plessmann, A. Bremerich, K. F. Lechtreck, and M. Melkonian. SF-assemblin, the structural protein of the 2-nm filaments from striated microtubule associated fibers of algal flagellar roots, forms a segmented coiled coil. Journal of Cell Biology, 121(4):837–??, May 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/121/4/837.

Walton:1992:TMA

[WGR^{+}92] P. A. Walton, S. J. Gould, R. A. Rachubinski, S. Subramani, and J. R. Feramisco. Transport of microinjected alcohol oxidase from Pichia pastoris into vesicles in mammalian cells: involvement of the peroxisomal targeting signal. Journal of Cell Biology, 118(3):499–??, August 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/118/3/499.

Welsh:1991:NFL

[WGRW91] J. B. Welsh, G. N. Gill, M. G. Rosenfeld, and A. Wells. A negative feedback loop attenuates EGF-induced morphological changes. Journal of Cell Biology, 114(3):533–??, August 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/114/3/533.

Wu:1994:RAD

[WGTCY94] Y. N. Wu, M. Gadina, J. H. Tao-Cheng, and R. J. Youle. Retinoic acid disrupts the Golgi apparatus and increases the cytosolic routing of specific protein toxins. Journal of Cell Biology, 125(4):743–??, May 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/125/4/743.

Winey:1993:NNP

[WHC^{+}93] M. Winey, M. A. Hoyt, C. Chan, L. Goetsch, D. Botstein, and B. Byers. NDC1: a nuclear periphery component required for yeast spindle pole body duplication. Journal of Cell Biology, 122(4):743–??, August 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/122/4/743.

Willecke:1991:MCC

[WHD^{+}91] K. Willecke, R. Heynkes, E. Dahl, R. Stutenkemper, H. Hennemann, S. Jungbluth, T. Suchyna, and B. J. Nicholson. Mouse connexin37: cloning and functional expression of a gap junction gene highly expressed in lung. Journal of Cell Biology, 114(5):1049–??, September 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/114/5/1049.

Wessels:1990:ADS

[WHF^{+}90] H. P. Wessels, G. H. Hansen, C. Fuhrer, A. T. Look, H. Sjöström, O. Norén, and M. Spiess. Aminopeptidase N is directly sorted to the apical domain in MDCK cells. Journal of Cell Biology, 111(6):2923–??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/111/6/2923.

Wolenski:1993:CCR

[WHFM93] J. S. Wolenski, S. M. Hayden, P. Forscher, and M. S. Mooseker. Calcium-calmodulin and regulation of brush border myosin–I MgATPase and mechanochemistry. Journal of Cell Biology, 122(3):613–??, August 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/122/3/613.

Williams:1993:RCM

[WHH^{+}93] C. L. Williams, V. Y. Hayes, A. M. Hummel, J. E. Tarara, and T. J. Halsey. Regulation of E-cadherin-mediated adhesion by muscarinic acetylcholine receptors in small cell lung carcinoma. Journal of Cell Biology, 121(3):643–??, May 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/121/3/643.

Ward:1990:FSI

[WHK90] D. M. Ward, D. P. Hackenyos, and J. Kaplan. Fusion of sequentially internalized vesicles in alveolar macrophages. Journal of Cell Biology, 110(4):1013–??, April 1990. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/110/4/
1013.

Witke:1993:CBD

[WHK^{+}93] W. Witke, A. Hofmann, B. Köppel, M. Schleicher, and A. A. Noegel. The Ca(2+)-binding domains in non-muscle type alpha-actinin: biochemical and genetic analysis. Journal of Cell Biology, 121(3):599–??, May 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/121/3/599.

Wattenberg:1990:IKP

[WHLW90] B. W. Wattenberg, R. R. Hiebsch, L. W. LeCureux, and M. P. White. Identification of a 25-kD protein from yeast cytosol that operates in a prefusion step of vesicular transport between compartments of the Golgi. Journal of Cell Biology, 110(4):947–??, April 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/110/4/947.

Wang:1992:LMG

[WHM^{+}92] F. Wang, M. Hanske, K. Miedema, G. Klein, P. Ekblom, and W. Hennig. Laminin in the male germ cells of Drosophila. Journal of Cell Biology, 119(4):977–??, November 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/119/4/
977.

Whetton:1994:CMP

[WHN^{+}94] A. D. Whetton, C. M. Heyworth, S. E. Nicholls, C. A. Evans, J. M. Lord, T. M. Dexter, and P. J. Owen-Lynch. Cytokine-mediated protein kinase C activation is a signal for lineage determination in bipotential granulocyte macrophage colony-forming cells. Journal of Cell Biology, 125(3):651–??, May 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/125/3/651.

Weinberger:1993:INS

[WHT^{+}93] R. P. Weinberger, R. C. Henke, O. Tolhurst, P. L. Jeffrey, and P. Gunning. Induction of neuron-specific tropomyosin mRNAs by nerve growth factor is dependent on morphological differentiation. Journal of Cell Biology, 120(1):205–??, January 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/120/1/205.

Wright:1991:SLS

[WHW^{+}91] B. D. Wright, J. H. Henson, K. P. Wedaman, P. J. Willy, J. N. Morand, and J. M. Scholey. Subcellular localization and sequence of sea urchin kinesin heavy chain: evidence for its association with membranes in the mitotic apparatus and interphase cytoplasm. Journal of Cell Biology, 113 (4):817–??, May 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/113/4/817.

Watson:1990:HRI

[WIF^{+}90] S. R. Watson, Y. Imai, C. Fennie, J. S. Geoffroy, S. D. Rosen, and L. A. Lasky. A homing receptor–IgG chimera as a probe for adhesive ligands of lymph node high endothelial venules. Journal of Cell Biology, 110(6):2221–??, June 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/110/6/
2221.

Watson:1991:CBL

[WIF^{+}91] S. R. Watson, Y. Imai, C. Fennie, J. Geoffrey, M. Singer, S. D. Rosen, and L. A. Lasky. The complement binding-like domains of the murine homing receptor facilitate lectin activity. Journal of Cell Biology, 115(1):235–??, October 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/115/1/
235.

Wewer:1994:PRT

[WIS^{+}94] U. M. Wewer, K. Ibaraki, P. Schjørring, M. E. Durkin, M. F. Young, and R. Albrechtsen. A potential role for tetranectin in mineralization during osteogenesis. Journal of Cell Biology, 127(6):1767–??, December 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/127/6/1767.

Wheelock:1992:RKI

[WJ92] M. J. Wheelock and P. J. Jensen. Regulation of keratinocyte intercellular junction organization and epidermal morphogenesis by E-cadherin. Journal of Cell Biology, 117(2):415– ??, April 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/117/2/415.

Wayner:1992:ADR

[WK92] E. A. Wayner and N. L. Kovach. Activation-dependent recognition by hematopoietic cells of the LDV sequence in the V region of fibronectin. Journal of Cell Biology, 116(2): 489–??, January 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/116/2/489.

Weeks:1991:HIH

[WKD^{+}91] B. S. Weeks, M. E. Klotman, S. Dhawan, M. Kibbey, J. Rappaport, H. K. Kleinman, K. M. Yamada, and P. E. Klotman. HIV-1 infection of human T lymphocytes results in enhanced alpha 5 beta 1 integrin expression. Journal of Cell Biology, 114(4):847–??, August 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/114/4/847.

Wetsel:1992:TCD

[WKM^{+}92] W. C. Wetsel, W. A. Khan, I. Merchenthaler, H. Rivera, A. E. Halpern, H. M. Phung, A. Negro-Vilar, and Y. A. Hannun. Tissue and cellular distribution of the extended family of protein kinase C isoenzymes. Journal of Cell Biology, 117(1):121–??, April 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/117/1/121.

Williams:1992:DZG

[WKMG92] B. C. Williams, T. L. Karr, J. M. Montgomery, and M. L. Goldberg. The Drosophila l(1)zw10 gene product, required for accurate mitotic chromosome segregation, is redistributed at anaphase onset. Journal of Cell Biology, 118 (4):759–??, August 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/118/4/759.

Wollner:1992:RCS

[WKN92] D. A. Wollner, K. A. Krzeminski, and W. J. Nelson. Remodeling the cell surface distribution of membrane proteins during the development of epithelial cell polarity. Journal of Cell Biology, 116(4):889–??, February 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/116/4/889.

Wilken:1993:NNN

[WKS^{+}93] N. Wilken, U. Kossner, J. L. Senécal, U. Scheer, and M. C. Dabauvalle. Nup180, a novel nuclear pore complex protein localizing to the cytoplasmic ring and associated fibrils. Journal of Cell Biology, 123(6):1345–??, December 1993.

CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/123/6/
1345.

Wileman:1993:ABS

[WKY^{+}93] T. Wileman, L. P. Kane, J. Young, G. R. Carson, and C. Terhorst. Associations between subunit ectodomains promote T cell antigen receptor assembly and protect against degradation in the ER. Journal of Cell Biology, 122(1):67– ??, July 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/122/1/67.

Wikstrom:1991:NPG

[WL91] L. Wikström and H. F. Lodish. Nonlysosomal, pre-Golgi degradation of unassembled asialoglycoprotein receptor subunits: a TLCK- and TPCK-sensitive cleavage within the ER. Journal of Cell Biology, 113(5):997–??, June 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/113/5/
997.

Wong:1992:RTD

[WLH92] S. H. Wong, S. H. Low, and W. Hong. The 17residue transmembrane domain of beta-galactoside alpha 2,6-sialyltransferase is sufficient for Golgi retention. Journal of Cell Biology, 117(2):245–??, April 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/117/2/245.

Wada:1992:ATP

[WLPB92] I. Wada, W. H. Lai, B. I. Posner, and J. J. Bergeron. Association of the tyrosine phosphorylated epidermal growth factor receptor with a 55-kD tyrosine phosphorylated protein at the cell surface and in endosomes. Journal of Cell Biology, 116(2):321–??, January 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/116/2/321.

Warren:1994:OHF

[WLWL94] K. S. Warren, J. L. Lin, D. D. Wamboldt, and J. J. Lin. Overexpression of human fibroblast caldesmon fragment containing actin-, Ca++/calmodulin-, and tropomyosinbinding domains stabilizes endogenous tropomyosin and microfilaments. Journal of Cell Biology, 125(2):359–??, April 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/125/2/359.

Wang:1993:NTT

[WLZB93] Y. Wang, P. A. Loomis, R. P. Zinkowski, and L. I. Binder. A novel tau transcript in cultured human neuroblastoma cells expressing nuclear tau. Journal of Cell Biology, 121 (2):257–??, April 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/121/2/257.

Wilton:1994:FCC

[WMB^{+}94] J. C. Wilton, G. M. Matthews, R. D. Burgoyne, C. O. Mills, J. K. Chipman, and R. Coleman. Fluorescent choleretic and cholestatic bile salts take different paths across the hepatocyte: transcytosis of glycolithocholate leads to an extensive redistribution of annexin II. Journal of Cell Biology, 127(2): 401–??, October 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/127/2/401.

Wu:1991:SNR

[WMCG91] Z. A. Wu, C. Murphy, H. G. Callan, and J. G. Gall. Small nuclear ribonucleoproteins and heterogeneous nuclear ribonucleoproteins in the amphibian germinal vesicle: loops, spheres, and snurposomes. Journal of Cell Biology, 113 (3):465–??, May 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/113/3/465.

Woodman:1992:CFF

[WMCW92] P. G. Woodman, D. I. Mundy, P. Cohen, and G. Warren. Cell-free fusion of endocytic vesicles is regulated by phosphorylation. Journal of Cell Biology, 116(2):331–??, January 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/116/2/331.

Whalen:1991:DRD

[WMF91] A. M. Whalen, M. McConnell, and P. A. Fisher. Developmental regulation of Drosophila DNA topoisomerase II.

Journal of Cell Biology, 112(2):203–??, January 1991. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/112/2/
203.

Weiner:1993:ABP

[WMG^{+}93] O. H. Weiner, J. Murphy, G. Griffiths, M. Schleicher, and A. A. Noegel. The actin-binding protein comitin (p24) is a component of the Golgi apparatus. Journal of Cell Biology, 123(1):23–??, October 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/123/1/23.

Wilson:1993:MSI

[WMS^{+}93] E. Wilson, Q. Mai, K. Sudhir, R. H. Weiss, and H. E. Ives. Mechanical strain induces growth of vascular smooth muscle cells via autocrine action of PDGF. Journal of Cell Biology, 123(3):741–??, November 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/123/3/741.

Wen:1994:SCP

[WMvdK94] J. Y. Wen, C. M. Morshead, and D. van der Kooy. Satellite cell proliferation in the adult rat trigeminal ganglion results from the release of a mitogenic protein from explanted sensory neurons. Journal of Cell Biology, 124(6):1005–??, March 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/124/6/1005.

Wolswijk:1992:CBP

[WN92] G. Wolswijk and M. Noble. Cooperation between PDGF and FGF converts slowly dividing O-2Aadult progenitor cells to rapidly dividing cells with characteristics of O-2Aperinatal progenitor cells. Journal of Cell Biology, 118 (4):889–??, August 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/118/4/889.

Wandinger-Ness:1990:DTV

[WNBAS90] A. Wandinger-Ness, M. K. Bennett, C. Antony, and K. Simons. Distinct transport vesicles mediate the delivery of plasma membrane proteins to the apical and basolateral domains of MDCK cells. Journal of Cell Biology, 111(3):987– ??, September 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/111/3/987.

Wilson:1994:RMA

[WNM^{+}94] B. S. Wilson, C. Nuoffer, J. L. Meinkoth, M. McCaffery, J. R. Feramisco, W. E. Balch, and M. G. Farquhar. A Rab1 mutant affecting guanine nucleotide exchange promotes disassembly of the Golgi apparatus. Journal of Cell Biology, 125(3):557–??, May 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/125/3/557.

Watabe:1994:IPC

[WNTT94] M. Watabe, A. Nagafuchi, S. Tsukita, and M. Takeichi. Induction of polarized cell–cell association and retardation of growth by activation of the E-cadherin-catenin adhesion system in a dispersed carcinoma line. Journal of Cell Biology, 127(1):247–??, October 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/127/1/247.

Welch:1990:TRF

[WOC90] M. P. Welch, G. F. Odland, and R. A. Clark. Temporal relationships of F-actin bundle formation, collagen and fibronectin matrix assembly, and fibronectin receptor expression to wound contraction. Journal of Cell Biology, 110(1): 133–??, January 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/110/1/133.

Wayner:1991:IAV

[WOC91] E. A. Wayner, R. A. Orlando, and D. A. Cheresh. Integrins alpha v beta 3 and alpha v beta 5 contribute to cell attachment to vitronectin but differentially distribute on the cell surface. Journal of Cell Biology, 113(4):919–??, May 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/113/4/

Woodcock:1994:CFO

[Woo94] C. L. Woodcock. Chromatin fibers observed in situ in frozen hydrated sections. Native fiber diameter is not correlated with nucleosome repeat length. Journal of Cell Biology, 125 (1):11–??, April 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/125/1/11.

Wu:1993:CKP

[WP93] H. Wu and J. T. Parsons. Cortactin, an 80/85-kilodalton pp60src substrate, is a filamentous actin-binding protein enriched in the cell cortex. Journal of Cell Biology, 120(6): 1417–??, March 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/120/6/1417.

Weber:1994:TCP

[WPBF94] A. Weber, C. R. Pennise, G. G. Babcock, and V. M. Fowler. Tropomodulin caps the pointed ends of actin filaments. Journal of Cell Biology, 127(6):1627–??, December 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/127/6/
1627.

Walker:1991:DIM

[WPS91] R. A. Walker, N. K. Pryer, and E. D. Salmon. Dilution of individual microtubules observed in real time in vitro: evidence that cap size is small and independent of elongation rate. Journal of Cell Biology, 114(1):73–??, July 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/114/1/
73.

Way:1992:CSS

[WPW92a] M. Way, B. Pope, and A. G. Weeds. Are the conserved sequences in segment 1 of gelsolin important for binding actin? Journal of Cell Biology, 116(5):1135–??, March 1992. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/116/5/

Way:1992:EFH

[WPW92b] M. Way, B. Pope, and A. G. Weeds. Evidence for functional homology in the F-actin binding domains of gelsolin and alpha-actinin: implications for the requirements of severing and capping. Journal of Cell Biology, 119(4):835– ??, November 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/119/4/835.

Wang:1993:MAC

[WRA93] L. H. Wang, K. G. Rothberg, and R. G. Anderson. Misassembly of clathrin lattices on endosomes reveals a regulatory switch for coated pit formation. Journal of Cell Biology, 123(5):1107–??, December 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/123/5/1107.

Wente:1992:NFY

[WRB92] S. R. Wente, M. P. Rout, and G. Blobel. A new family of yeast nuclear pore complex proteins. Journal of Cell Biology, 119(4):705–??, November 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/119/4/705.

Whiteheart:1994:ESF

[WRB^{+}94] S. W. Whiteheart, K. Rossnagel, S. A. Buhrow, M. Brunner, R. Jaenicke, and J. E. Rothman. N-ethylmaleimide-sensitive fusion protein: a trimeric ATPase whose hydrolysis of ATP is required for membrane fusion. Journal of Cell Biology, 126(4):945–??, August 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/126/4/945.

Wacker:1992:AMB

[WRDK92] I. U. Wacker, J. E. Rickard, J. R. De Mey, and T. E. Kreis. Accumulation of a microtubule-binding protein, pp170, at desmosomal plaques. Journal of Cell Biology, 117(4):813– ??, May 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/117/4/813.

Wattenberg:1992:AGT

[WRHW92] B. W. Wattenberg, T. J. Raub, R. R. Hiebsch, and P. J. Weidman. The activity of Golgi transport vesicles depends on the presence of the N-ethylmaleimide-sensitive factor (NSF) and a soluble NSF attachment protein (alpha SNAP) during vesicle formation. Journal of Cell Biology, 118(6): 1321–??, September 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/118/6/1321.

Weller:1991:CLB

[WRP^{+}91] P. F. Weller, S. W. Ryeom, S. T. Picard, S. J. Ackerman, and A. M. Dvorak. Cytoplasmic lipid bodies of neutrophils: formation induced by cis-unsaturated fatty acids and mediated by protein kinase C. Journal of Cell Biology, 113 (1):137–??, April 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/113/1/137.

Wessels:1990:MIH

[WS90] D. Wessels and D. R. Soll. Myosin II heavy chain null mutant of Dictyostelium exhibits defective intracellular particle movement. Journal of Cell Biology, 111(3):1137– ??, September 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/111/3/1137.

Weisz:1991:HACa

[WS91a] O. A. Weisz and R. L. Schnaar. Hepatocyte adhesion to carbohydrate-derivatized surfaces. I. Surface topography of the rat hepatic lectin. Journal of Cell Biology, 115(2): 485–??, October 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/115/2/485.

Weisz:1991:HACb

[WS91b] O. A. Weisz and R. L. Schnaar. Hepatocyte adhesion to carbohydrate-derivatized surfaces. II. Regulation of cytoskeletal organization and cell morphology. Journal of Cell Biology, 115(2):495–??, October 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/115/2/495.

Wilson:1992:ASV

[WS92] C. L. Wilson and J. E. Schwarzbauer. The alternatively spliced V region contributes to the differential incorporation of plasma and cellular fibronectins into fibrin clots. Journal of Cell Biology, 119(4):923–??, November 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/119/4/923.

Wendland:1993:CDP

[WS93a] M. Wendland and S. Subramani. Cytosol-dependent peroxisomal protein import in a permeabilized cell system. Journal of Cell Biology, 120(3):675–??, February 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/120/3/675.

Whalen:1993:DDC

[WS93b] A. M. Whalen and R. Steward. Dissociation of the dorsalcactus complex and phosphorylation of the dorsal protein correlate with the nuclear localization of dorsal. Journal of Cell Biology, 123(3):523–??, November 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/123/3/523.

Weidner:1993:MRT

[WSB93] K. M. Weidner, M. Sachs, and W. Birchmeier. The Met receptor tyrosine kinase transduces motility, proliferation, and morphogenic signals of scatter factor/hepatocyte growth factor in epithelial cells. Journal of Cell Biology, 121(1): 145–??, April 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/121/1/145.

Wang:1994:SPT

[WSC94] Y. L. Wang, J. D. Silverman, and L. G. Cao. Single particle tracking of surface receptor movement during cell division. Journal of Cell Biology, 127(4):963–??, November 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/127/4/
963.

Weinstein:1991:SAM

[WSL91] D. E. Weinstein, M. L. Shelanski, and R. K. Liem. Suppression by antisense mRNA demonstrates a requirement for the glial fibrillary acidic protein in the formation of stable astrocytic processes in response to neurons. Journal of Cell Biology, 112(6):1205–??, March 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/112/6/1205.

Weisz:1993:OMP

[WSM93] O. A. Weisz, A. M. Swift, and C. E. Machamer. Oligomerization of a membrane protein correlates with its retention in the Golgi complex. Journal of Cell Biology, 122(6):1185– ??, September 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/122/6/1185.

Wordeman:1991:CSW

[WSSM91] L. Wordeman, E. R. Steuer, M. P. Sheetz, and T. Mitchison. Chemical subdomains within the kinetochore domain of isolated CHO mitotic chromosomes. Journal of Cell Biology, 114(2):285–??, July 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/114/2/285.

Wieser:1990:ICK

[WST^{+}90] R. J. Wieser, S. Schütz, G. Tschank, H. Thomas, H. P. Dienes, and F. Oesch. Isolation and characterization of a 60-70-kD plasma membrane glycoprotein involved in the contact-dependent inhibition of growth. Journal of Cell Biology, 111(6):2681–??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/111/6/2681.

Wansink:1993:FLN

[WSvdK^{+}93] D. G. Wansink, W. Schul, I. van der Kraan, B. van Steensel, R. van Driel, and L. de Jong. Fluorescent labeling of nascent RNA reveals transcription by RNA polymerase II in domains scattered throughout the nucleus. Journal of Cell Biology, 122(2):283–??, July 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/122/2/283.

Waterham:1994:HPP

[WTH^{+}94] H. R. Waterham, V. I. Titorenko, P. Haima, J. M. Cregg, W. Harder, and M. Veenhuis. The Hansenula polymorpha PER1 gene is essential for peroxisome biogenesis and encodes a peroxisomal matrix protein with both carboxy- and amino-terminal targeting signals. Journal of Cell Biology, 127(3):737–??, November 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/127/3/737.

Wu:1992:TBA

[WTK^{+}92] S. P. Wu, D. Theodorescu, R. S. Kerbel, J. K. Willson, K. M. Mulder, L. E. Humphrey, and M. G. Brattain. TGFbeta 1 is an autocrine-negative growth regulator of human colon carcinoma FET cells in vivo as revealed by transfection of an antisense expression vector. Journal of Cell Biology, 116(1):187–??, January 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/116/1/187.

Woodford-Thomas:1992:EPT

[WTRD92] T. A. Woodford-Thomas, J. D. Rhodes, and J. E. Dixon. Expression of a protein tyrosine phosphatase in normal and v-src-transformed mouse 3T3 fibroblasts. Journal of Cell Biology, 117(2):401–??, April 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/117/2/401.

Wright:1993:RKK

[WTS93] B. D. Wright, M. Terasaki, and J. M. Scholey. Roles of kinesin and kinesin-like proteins in sea urchin embryonic cell division: evaluation using antibody microinjection. Journal of Cell Biology, 123(3):681–??, November 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/123/3/681.

Watanabe:1991:MBC

[WTSP91] N. Watanabe, N. Tsukada, C. R. Smith, and M. J. Phillips. Motility of bile canaliculi in the living animal: implications for bile flow. Journal of Cell Biology, 113(5):1069– ??, June 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/113/5/1069.

Wu:1993:TNS

[Wu93] C. T. Wu. Transvection, nuclear structure, and chromatin proteins. Journal of Cell Biology, 120(3):587–??, February 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/120/3/587.

Wilhelm:1993:RMM

[WV93] J. E. Wilhelm and R. D. Vale. RNA on the move: the mRNA localization pathway. Journal of Cell Biology, 123 (2):269–??, October 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/123/2/269.

Walther:1994:CFG

[WVH94] Z. Walther, M. Vashishtha, and J. L. Hall. The Chlamydomonas FLA10 gene encodes a novel kinesin-homologous protein. Journal of Cell Biology, 126(1):175–??, July 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/126/1/
175.

Woodman:1991:IFC

[WW91] P. G. Woodman and G. Warren. Isolation of functional, coated, endocytic vesicles. Journal of Cell Biology, 112(6): 1133–??, March 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/112/6/1133.

Wolin:1993:DNC

[WW93] S. L. Wolin and P. Walter. Discrete nascent chain lengths are required for the insertion of presecretory proteins into microsomal membranes. Journal of Cell Biology, 121(6): 1211–??, June 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/121/6/1211.

Weidman:1994:GPA

[WW94a] P. J. Weidman and W. M. Winter. The G protein-activating peptide, mastoparan, and the synthetic NH2-terminal ARF peptide, ARFp13, inhibit in vitro Golgi transport by irreversibly damaging membranes. Journal of Cell Biology, 127(6):1815–??, December 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/127/6/1815.

Williams:1994:GCM

[WW94b] B. D. Williams and R. H. Waterston. Genes critical for muscle development and function in Caenorhabditis elegans identified through lethal mutations. Journal of Cell Biology, 124(4):475–??, February 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/124/4/475.

Williams:1994:TKI

[WWD94] E. J. Williams, F. S. Walsh, and P. Doherty. Tyrosine kinase inhibitors can differentially inhibit integrin-dependent and CAM-stimulated neurite outgrowth. Journal of Cell Biology, 124(6):1029–??, March 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/124/6/1029.

Witte:1991:CDR

[WWHA91] D. P. Witte, D. A. Wiginton, J. J. Hutton, and B. J. Aronow. Coordinate developmental regulation of purine catabolic enzyme expression in gastrointestinal and postimplantation reproductive tracts. Journal of Cell Biology, 115 (1):179–??, October 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/115/1/179.

Wren:1992:VAO

[WWN92] D. Wren, G. Wolswijk, and M. Noble. In vitro analysis of the origin and maintenance of O-2Aadult progenitor cells. Journal of Cell Biology, 116(1):167–??, January 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/116/1/167.

Walton:1994:IKH

[WWS^{+}94] P. A. Walton, M. Wendland, S. Subramani, R. A. Rachubinski, and W. J. Welch. Involvement of 70-kD heat-shock proteins in peroxisomal import. Journal of Cell Biology, 125 (5):1037–??, June 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/125/5/1037.

Wilson:1992:MPI

[WWW^{+}92] D. W. Wilson, S. W. Whiteheart, M. Wiedmann, M. Brunner, and J. E. Rothman. A multisubunit particle implicated in membrane fusion. Journal of Cell Biology, 117 (3):531–??, May 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/117/3/531.

Webster:1990:DAT

[WWWB90] D. R. Webster, J. Wehland, K. Weber, and G. G. Borisy. Detyrosination of alpha tubulin does not stabilize microtubules in vivo. Journal of Cell Biology, 111(1):113–??, July 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/111/1/
113.

Warmke:1992:MLC

[WYM^{+}92] J. Warmke, M. Yamakawa, J. Molloy, S. Falkenthal, and D. Maughan. Myosin light chain-2 mutation affects flight, wing beat frequency, and indirect flight muscle contraction kinetics in Drosophila. Journal of Cell Biology, 119(6):1523– ??, December 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/119/6/1523.

Xing:1991:PSR

[XL91] Y. G. Xing and J. B. Lawrence. Preservation of specific RNA distribution within the chromatin-depleted nuclear substructure demonstrated by in situ hybridization coupled with biochemical fractionation. Journal of Cell Biology, 112 (6):1055–??, March 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/112/6/1055.

Xu:1993:VSP

[XMW^{+}93] Y. F. Xu, A. N. Meyer, M. K. Webster, B. A. Lee, and D. J. Donoghue. The v-sis protein retains biological activity as a type II membrane protein when anchored by various signalanchor domains, including the hydrophobic domain of the bovine papilloma virus E5 oncoprotein. Journal of Cell Biology, 123(3):549–??, November 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/123/3/549.

Xu:1993:PPT

[XS93] H. Xu and D. Shields. Prohormone processing in the trans-Golgi network: endoproteolytic cleavage of prosomatostatin and formation of nascent secretory vesicles in permeabilized cells. Journal of Cell Biology, 122(6):1169–??, September 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/122/6/1169.

Yuan:1990:BTT

[YAJ90] S. Yuan, W. Arnold, and A. O. Jorgensen. Biogenesis of transverse tubules: immunocytochemical localization of a transverse tubular protein (TS28) and a sarcolemmal protein (SL50) in rabbit skeletal muscle developing in situ. Journal of Cell Biology, 110(4):1187–??, April 1990. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/110/4/
1187.

Yuan:1991:BTT

[YAJ91] S. H. Yuan, W. Arnold, and A. O. Jorgensen. Biogenesis of transverse tubules and triads: immunolocalization of the 1,4-dihydropyridine receptor, TS28, and the ryanodine receptor in rabbit skeletal muscle developing in situ. Journal of Cell Biology, 112(2):289–??, January 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/112/2/289.

Yueh:1993:DCR

[YC93] Y. G. Yueh and R. C. Crain. Deflagellation of Chlamydomonas reinhardtii follows a rapid transitory accumulation of inositol 1,4,5-trisphosphate and requires Ca^{2+} entry. Journal of Cell Biology, 123(4):869–??, November 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/123/4/
869.

Yu:1993:MNR

[YCAB93] W. Yu, V. E. Centonze, F. J. Ahmad, and P. W. Baas. Microtubule nucleation and release from the neuronal centrosome. Journal of Cell Biology, 122(2):349–??, July 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/122/2/
349.

Yan:1994:ICM

[YCBM94] J. P. Yan, M. E. Colon, L. A. Beebe, and P. Melançon. Isolation and characterization of mutant CHO cell lines with compartment-specific resistance to brefeldin A. Journal of Cell Biology, 126(1):65–??, July 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/126/1/65.

Yurchenco:1992:LFI

[YCC92] P. D. Yurchenco, Y. S. Cheng, and H. Colognato. Laminin forms an independent network in basement membranes. Journal of Cell Biology, 117(5):1119–??, June 1992. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/117/5/
1119.

Ylanne:1993:DFI

[YCO^{+}93] J. Ylänne, Y. Chen, T. E. O’Toole, J. C. Loftus, Y. Takada, and M. H. Ginsberg. Distinct functions of integrin alpha and beta subunit cytoplasmic domains in cell spreading and formation of focal adhesions. Journal of Cell Biology, 122 (1):223–??, July 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/122/1/223.

Yeager:1990:TDS

[YDO^{+}90] M. Yeager, K. A. Dryden, N. H. Olson, H. B. Greenberg, and T. S. Baker. Three-dimensional structure of rhesus rotavirus by cryoelectron microscopy and image reconstruction. Journal of Cell Biology, 110(6):2133–??, June 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/110/6/
2133.

Yilla:1992:EDB

[YDS92] M. Yilla, D. Doyle, and J. T. Sawyer. Early disulfide bond formation prevents heterotypic aggregation of membrane proteins in a cell-free translation system. Journal of Cell Biology, 118(2):245–??, July 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/118/2/245.

Yelian:1993:REP

[YED^{+}93] F. D. Yelian, N. A. Edgeworth, L. J. Dong, A. E. Chung, and D. R. Armant. Recombinant entactin promotes mouse primary trophoblast cell adhesion and migration through the Arg–Gly–Asp (RGD) recognition sequence. Journal of Cell Biology, 121(4):923–??, May 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/121/4/923.

Young:1992:IPY

[YFS92] V. B. Young, S. Falkow, and G. K. Schoolnik. The invasin protein of Yersinia enterocolitica: internalization of invasinbearing bacteria by eukaryotic cells is associated with reorganization of the cytoskeleton. Journal of Cell Biology, 116 (1):197–??, January 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/116/1/197.

Yaar:1991:ENG

[YGEG91] M. Yaar, K. Grossman, M. Eller, and B. A. Gilchrest. Evidence for nerve growth factor-mediated paracrine effects in human epidermis. Journal of Cell Biology, 115(3):821– ??, November 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/115/3/821.

Yu:1994:BFP

[YGM^{+}94] J. C. Yu, J. S. Gutkind, D. Mahadevan, W. Li, K. A. Meyers, J. H. Pierce, and M. A. Heidaran. Biological function of PDGF-induced PI-3 kinase activity: its role in alpha PDGF receptor-mediated mitogenic signaling. Journal of Cell Biology, 127(2):479–??, October 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/127/2/479.

Yoshihara:1993:IEK

[YH93] C. M. Yoshihara and Z. W. Hall. Increased expression of the 43-kD protein disrupts acetylcholine receptor clustering in myotubes. Journal of Cell Biology, 122(1):169–??, July 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/122/1/169.

Youakim:1994:OSS

[YHM^{+}94] A. Youakim, H. J. Hathaway, D. J. Miller, X. Gong, and B. D. Shur. Overexpressing sperm surface beta 1,4galactosyltransferase in transgenic mice affects multiple aspects of sperm-egg interactions. Journal of Cell Biology, 126(6):1573–??, September 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/126/6/1573.

Yamaguchi:1991:RHB

[YKI^{+}91] A. Yamaguchi, T. Katagiri, T. Ikeda, J. M. Wozney, V. Rosen, E. A. Wang, A. J. Kahn, T. Suda, and S. Yoshiki. Recombinant human bone morphogenetic protein-2 stimulates osteoblastic maturation and inhibits myogenic differentiation in vitro. Journal of Cell Biology, 113(3):681–??, May 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/113/3/681.

Yoda:1992:HCP

[YKM^{+}92] K. Yoda, K. Kitagawa, H. Masumoto, Y. Muro, and T. Okazaki. A human centromere protein, CENP–B, has a DNA binding domain containing four potential alpha helices at the NH2 terminus, which is separable from dimerizing activity. Journal of Cell Biology, 119(6):1413–??, December 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/119/6/1413.

Yumura:1992:RMI

[YKY92] S. Yumura and T. Kitanishi-Yumura. Release of myosin II from the membrane-cytoskeleton of Dictyostelium discoideum mediated by heavy-chain phosphorylation at the foci within the cortical actin network. Journal of Cell Biology, 117(6):1231–??, June 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/117/6/1231.

Yuk:1993:TPD

[YL93] M. H. Yuk and H. F. Lodish. Two pathways for the degradation of the H2 subunit of the asialoglycoprotein receptor in the endoplasmic reticulum. Journal of Cell Biology, 123(6):1735–??, December 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/123/6/1735.

Yeadon:1991:DCS

[YLDB91] J. E. Yeadon, H. Lin, S. M. Dyer, and S. J. Burden. Dystrophin is a component of the subsynaptic membrane. Journal of Cell Biology, 115(4):1069–??, November 1991. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/115/4/
1069.

Yang:1992:NUL

[YLS92] C. H. Yang, E. J. Lambie, and M. Snyder. NuMA: an unusually long coiled-coil related protein in the mammalian nucleus. Journal of Cell Biology, 116(6):1303–??, March 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/116/6/
1303.

Yu:1994:UII

[YLWS94] J. Yu, J. H. Lin, X. R. Wu, and T. T. Sun. Uroplakins Ia and Ib, two major differentiation products of bladder epithelium, belong to a family of four transmembrane domain (4TM) proteins. Journal of Cell Biology, 125(1):171–??, April 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/125/1/171.

Yamashita:1990:IAR

[YM90a] S. Yamashita and J. L. Maller. Identification of an activator required for elevation of maturation-promoting factor (MPF) activity by gamma–S–ATP. Journal of Cell Biology, 110(5):1583–??, May 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/110/5/1583.

Yang:1990:TGF

[YM90b] E. Y. Yang and H. L. Moses. Transforming growth factor beta 1-induced changes in cell migration, proliferation, and angiogenesis in the chicken chorioallantoic membrane. Journal of Cell Biology, 111(2):731–??, August 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/111/2/731.

Yan:1993:MRN

[YM93] C. Yan and T. Mélèse. Multiple regions of NSR1 are sufficient for accumulation of a fusion protein within the nucleolus. Journal of Cell Biology, 123(5):1081–??, December 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/123/5/1081.

Yaver:1992:MSR

[YMO92] D. S. Yaver, S. Matoba, and D. M. Ogrydziak. A mutation in the signal recognition particle 7S RNA of the yeast Yarrowia lipolytica preferentially affects synthesis of the alkaline extracellular protease: in vivo evidence for translational arrest. Journal of Cell Biology, 116(3):605–??, February 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/116/3/605.

Yu:1990:UKP

[YMW^{+}90] J. Yu, M. Manabe, X. R. Wu, C. Xu, B. Surya, and T. T. Sun. Uroplakin I: a 27-kD protein associated with the asymmetric unit membrane of mammalian urothelium. Journal of Cell Biology, 111(3):1207–??, September 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/111/3/1207.

Yonemura:1993:CIM

[YNST93] S. Yonemura, A. Nagafuchi, N. Sato, and S. Tsukita. Concentration of an integral membrane protein, CD43 (leukosialin, sialophorin), in the cleavage furrow through the interaction of its cytoplasmic domain with actin-based cytoskeletons. Journal of Cell Biology, 120(2):437–??, January 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/120/2/437.

Yoshihara:1991:DSR

[YOWM91] Y. Yoshihara, S. Oka, Y. Watanabe, and K. Mori. Developmentally and spatially regulated expression of HNK-1 carbohydrate antigen on a novel phosphatidylinositol-anchored glycoprotein in rat brain. Journal of Cell Biology, 115(3): 731–??, November 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/115/3/731.

Yokode:1992:CSR

[YPH^{+}92] M. Yokode, R. K. Pathak, R. E. Hammer, M. S. Brown, J. L. Goldstein, and R. G. Anderson. Cytoplasmic sequence required for basolateral targeting of LDL receptor in livers of transgenic mice. Journal of Cell Biology, 117(1):39–??, April 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/117/1/39.

Yeaman:1993:PTR

[YR93] C. Yeaman and A. C. Rapraeger. Post-transcriptional regulation of syndecan-1 expression by cAMP in peritoneal macrophages. Journal of Cell Biology, 122(4):941–??, August 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/122/4/941.

Yablonka-Reuveni:1990:RPD

[YRBBP90] Z. Yablonka-Reuveni, T. M. Balestreri, and D. F. Bowen-Pope. Regulation of proliferation and differentiation of myoblasts derived from adult mouse skeletal muscle by specific isoforms of PDGF. Journal of Cell Biology, 111(4): 1623–??, October 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/111/4/1623.

Yu:1991:ILC

[YSBM91] I. J. Yu, D. L. Spector, Y. S. Bae, and D. R. Marshak. Immunocytochemical localization of casein kinase II during interphase and mitosis. Journal of Cell Biology, 114 (6):1217–??, September 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/114/6/1217.

Yu:1990:AAI

[YSK90] Y. H. Yu, D. D. Sabatini, and G. Kreibich. Antiribophorin antibodies inhibit the targeting to the ER membrane of ribosomes containing nascent secretory polypeptides. Journal of Cell Biology, 111(4):1335–??, October 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/111/4/1335.

Yang:1994:MFA

[YSK^{+}94] Y. Yang, E. Spitzer, N. Kenney, W. Zschiesche, M. Li, A. Kromminga, T. Müller, F. Spener, A. Lezius, and J. H. Veerkamp. Members of the fatty acid binding protein family are differentiation factors for the mammary gland. Journal of Cell Biology, 127(4):1097–??, November 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/127/4/1097.

Yamochi:1994:GSL

[YTN^{+}94] W. Yamochi, K. Tanaka, H. Nonaka, A. Maeda, T. Musha, and Y. Takai. Growth site localization of Rho1 small GTPbinding protein and its involvement in bud formation in Saccharomyces cerevisiae. Journal of Cell Biology, 125(5): 1077–??, June 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/125/5/1077.

Yamakita:1990:MNS

[YYM90] Y. Yamakita, S. Yamashiro, and F. Matsumura. Microinjection of nonmuscle and smooth muscle caldesmon into fibroblasts and muscle cells. Journal of Cell Biology, 111(6): 2487–??, December 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/111/6/2487.

Yamakita:1994:VPR

[YYM94] Y. Yamakita, S. Yamashiro, and F. Matsumura. In vivo phosphorylation of regulatory light chain of myosin II during mitosis of cultured cells. Journal of Cell Biology, 124(1): 129–??, January 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/124/1/129.

Yang:1992:ISA

[YYR92] P. Yang, X. Yin, and U. Rutishauser. Intercellular space is affected by the polysialic acid content of NCAM. Journal of Cell Biology, 116(6):1487–??, March 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/116/6/1487.

Yamagata:1991:CPS

[YYY^{+}91] M. Yamagata, K. M. Yamada, S. S. Yamada, T. Shinomura, H. Tanaka, Y. Nishida, M. Obara, and K. Kimata. The complete primary structure of type XII collagen shows a chimeric molecule with reiterated fibronectin type III motifs, von Willebrand factor A motifs, a domain homologous to a noncollagenous region of type IX collagen, and short collagenous domains with an Arg–Gly–Asp site. Journal of Cell Biology, 115(1):209–??, October 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/115/1/209.

Zhang:1994:ICP

[ZABM94] L. Zhang, C. L. Ashendel, G. W. Becker, and D. J. Morré. Isolation and characterization of the principal ATPase associated with transitional endoplasmic reticulum of rat liver. Journal of Cell Biology, 127(6):1871–??, December 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/127/6/
1871.

Zhang:1994:SEF

[ZAH^{+}94] H. Zhang, S. D. Apfelroth, W. Hu, E. C. Davis, C. Sanguineti, J. Bonadio, R. P. Mecham, and F. Ramirez. Structure and expression of fibrillin-2, a novel microfibrillar component preferentially located in elastic matrices. Journal of Cell Biology, 124(5):855–??, March 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/124/5/855.

Zhou:1994:CMA

[ZB94] M. J. Zhou and E. J. Brown. CR3 (Mac-1, alpha M beta 2, CD11b/CD18) and Fc gamma RIII cooperate in generation of a neutrophil respiratory burst: requirement for Fc gamma RIII and tyrosine phosphorylation. Journal of Cell Biology, 125(6):1407–??, June 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/125/6/1407.

Zheng:1994:MGC

[ZBH94] J. Zheng, R. E. Buxbaum, and S. R. Heidemann. Measurements of growth cone adhesion to culture surfaces by micromanipulation. Journal of Cell Biology, 127(6):2049– ??, December 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/127/6/2049.

Zen:1992:SMR

[ZBPV92] K. Zen, J. Biwersi, N. Periasamy, and A. S. Verkman. Second messengers regulate endosomal acidification in Swiss 3T3 fibroblasts. Journal of Cell Biology, 119(1):99–??, October 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/119/1/99.

Zimmerberg:1994:RML

[ZBS^{+}94] J. Zimmerberg, R. Blumenthal, D. P. Sarkar, M. Curran, and S. J. Morris. Restricted movement of lipid and aqueous dyes through pores formed by influenza hemagglutinin during cell fusion. Journal of Cell Biology, 127(6):1885– ??, December 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/127/6/1885.

Zopf:1993:GDK

[ZBW93] D. Zopf, H. D. Bernstein, and P. Walter. GTPase domain of the 54-kD subunit of the mammalian signal recognition particle is required for protein translocation but not for signal sequence binding. Journal of Cell Biology, 120(5):1113–??, March 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/120/5/1113.

Zhu:1992:ACU

[ZC92] Q. Zhu and M. Clarke. Association of calmodulin and an unconventional myosin with the contractile vacuole complex of Dictyostelium discoideum. Journal of Cell Biology, 118 (2):347–??, July 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/118/2/347.

Zhang:1990:RAC

[ZCH90] D. H. Zhang, D. A. Callaham, and P. K. Hepler. Regulation of anaphase chromosome motion in Tradescantia stamen hair cells by calcium and related signaling agents. Journal of Cell Biology, 111(1):171–??, July 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/111/1/171.

Zhang:1994:MCS

[ZCP^{+}94] Q. Zhang, W. J. Checovich, D. M. Peters, R. M. Albrecht, and D. F. Mosher. Modulation of cell surface fibronectin assembly sites by lysophosphatidic acid. Journal of Cell Biology, 127(5):1447–??, December 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/127/5/1447.

Zhang:1991:LDM

[ZCS^{+}91] F. Zhang, B. Crise, B. Su, Y. Hou, J. K. Rose, A. Bothwell, and K. Jacobson. Lateral diffusion of membranespanning and glycosylphosphatidylinositol-linked proteins: toward establishing rules governing the lateral mobility of membrane proteins. Journal of Cell Biology, 115(1):75–??, October 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/115/1/75.

Zhang:1991:SPN

[ZDO^{+}91] H. Zhang, N. N. Desai, A. Olivera, T. Seki, G. Brooker, and S. Spiegel. Sphingosine-1-phosphate, a novel lipid, involved in cellular proliferation. Journal of Cell Biology, 114 (1):155–??, July 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/114/1/155.

Zot:1992:MMA

[ZDP92] H. G. Zot, S. K. Doberstein, and T. D. Pollard. Myosini moves actin filaments on a phospholipid substrate: implications for membrane targeting. Journal of Cell Biology, 116(2):367–??, January 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/116/2/367.

Zisch:1992:NCA

[ZDR^{+}92] A. H. Zisch, L. D’Alessandri, B. Ranscht, R. Falchetto, K. H. Winterhalter, and L. Vaughan. Neuronal cell adhesion molecule contactin/F11 binds to tenascin via its immunoglobulin-like domains. Journal of Cell Biology, 119 (1):203–??, October 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/119/1/203.

Zimmermann:1994:VEP

[ZDZSBT94] D. R. Zimmermann, M. T. Dours-Zimmermann, M. Schubert, and L. Bruckner-Tuderman. Versican is expressed in the proliferating zone in the epidermis and in association with the elastic network of the dermis. Journal of Cell Biology, 124(5):817–??, March 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/124/5/817.

Zhou:1993:HAB

[ZFA^{+}93] H. Zhou, A. Fuks, G. Alcaraz, T. J. Bolling, and C. P. Stanners. Homophilic adhesion between Ig superfamily carcinoembryonic antigen molecules involves double reciprocal bonds. Journal of Cell Biology, 122(4):951–??, August 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/122/4/
951.

Zigmond:1992:IAF

[ZFF92] S. H. Zigmond, R. Furukawa, and M. Fechheimer. Inhibition of actin filament depolymerization by the Dictyostelium 30,000-D actin-bundling protein. Journal of Cell Biology, 119(3):559–??, November 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/119/3/559.

Zhang:1993:NPC

[ZHL93] J. W. Zhang, Y. Han, and P. B. Lazarow. Novel peroxisome clustering mutants and peroxisome biogenesis mutants of Saccharomyces cerevisiae. Journal of Cell Biology, 123(5):1133–??, December 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/123/5/1133.

Zahraoui:1994:SRG

[ZJA^{+}94] A. Zahraoui, G. Joberty, M. Arpin, J. J. Fontaine, R. Hellio, A. Tavitian, and D. Louvard. A small rab GTPase is distributed in cytoplasmic vesicles in non polarized cells but colocalizes with the tight junction marker ZO-1 in polarized epithelial cells. Journal of Cell Biology, 124(1):101–??, January 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/124/1/101.

Zachar:1993:ECD

[ZKMB93] Z. Zachar, J. Kramer, I. P. Mims, and P. M. Bingham. Evidence for channeled diffusion of pre-mRNAs during nuclear RNA transport in metazoans. Journal of Cell Biology, 121 (4):729–??, May 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/121/4/729.

Zurzolo:1993:GAP

[ZLC^{+}93] C. Zurzolo, M. P. Lisanti, I. W. Caras, L. Nitsch, and E. Rodriguez-Boulan. Glycosylphosphatidylinositolanchored proteins are preferentially targeted to the basolateral surface in Fischer rat thyroid epithelial cells. Journal of Cell Biology, 121(5):1031–??, June 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/121/5/1031.

Zanetti:1990:BDP

[ZLG^{+}90] M. Zanetti, L. Litteri, R. Gennaro, H. Horstmann, and D. Romeo. Bactenecins, defense polypeptides of bovine neutrophils, are generated from precursor molecules stored in the large granules. Journal of Cell Biology, 111(4):1363–??, October 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/111/4/1363.

Zinkowski:1991:CKC

[ZMB91] R. P. Zinkowski, J. Meyne, and B. R. Brinkley. The centromere-kinetochore complex: a repeat subunit model. Journal of Cell Biology, 113(5):1091–??, June 1991. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/113/5/
1091.

Zimmerman:1990:ECA

[ZMMP90] G. A. Zimmerman, T. M. McIntyre, M. Mehra, and S. M. Prescott. Endothelial cell-associated platelet-activating factor: a novel mechanism for signaling intercellular adhesion.

Journal of Cell Biology, 110(2):529–??, February 1990. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/110/2/
529.

Zhou:1991:SGB

[ZMS^{+}91] Q. Zhou, K. L. Moore, D. F. Smith, A. Varki, R. P. McEver, and R. D. Cummings. The selectin GMP-140 binds to sialylated, fucosylated lactosaminoglycans on both myeloid and nonmyeloid cells. Journal of Cell Biology, 115(2):557–??, October 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/115/2/557.

Zhang:1993:AVB

[ZMV^{+}93] Z. Zhang, A. O. Morla, K. Vuori, J. S. Bauer, R. L. Juliano, and E. Ruoslahti. The alpha v beta 1 integrin functions as a fibronectin receptor but does not support fibronectin matrix assembly and cell migration on fibronectin. Journal of Cell Biology, 122(1):235–??, July 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/122/1/235.

Zurzolo:1992:OPV

[ZPS^{+}92] C. Zurzolo, C. Polistina, M. Saini, R. Gentile, L. Aloj, G. Migliaccio, S. Bonatti, and L. Nitsch. Opposite polarity of virus budding and of viral envelope glycoprotein distribution in epithelial cells derived from different tissues. Journal of Cell Biology, 117(3):551–??, May 1992. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/117/3/551.

Zhang:1994:EDA

[ZRW^{+}94] C. J. Zhang, A. G. Rosenwald, M. C. Willingham, S. Skuntz, J. Clark, and R. A. Kahn. Expression of a dominant allele of human ARF1 inhibits membrane traffic in vivo. Journal of Cell Biology, 124(3):289–??, February 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/124/3/289.

Zhang:1994:FAD

[ZS94] Y. Zhang and W. J. Snell. Flagellar adhesion-dependent regulation of Chlamydomonas adenylyl cyclase in vitro: a possible role for protein kinases in sexual signaling. Journal of Cell Biology, 125(3):617–??, May 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/125/3/617.

Zhu:1991:PES

[ZSC91] Y. Y. Zhu, R. J. Schwartz, and M. T. Crow. Phorbol esters selectively downregulate contractile protein gene expression in terminally differentiated myotubes through transcriptional repression and message destabilization. Journal of Cell Biology, 115(3):745–??, November 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/115/3/745.

Zhang:1993:CFA

[ZSH^{+}93] R. Zhang, W. Skach, H. Hasegawa, A. N. van Hoek, and A. S. Verkman. Cloning, functional analysis and cell localization of a kidney proximal tubule water transporter homologous to CHIP28. Journal of Cell Biology, 120(2): 359–??, January 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/120/2/359.

Zhong:1993:MAR

[ZSM^{+}93] Y. Zhong, T. Saitoh, T. Minase, N. Sawada, K. Enomoto, and M. Mori. Monoclonal antibody 7H6 reacts with a novel tight junction-associated protein distinct from ZO-1, cingulin and ZO-2. Journal of Cell Biology, 120(2):477–??, January 1993. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/120/2/477.

Zaar:1991:PMP

[ZVF91] K. Zaar, A. Völkl, and H. D. Fahimi. Purification of marginal plates from bovine renal peroxisomes: identification with L-alpha-hydroxyacid oxidase B. Journal of Cell Biology, 113(1):113–??, April 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://

jcb.rupress.org/content/113/1/113.

Zakeri:1990:CIH

[ZWW90] Z. F. Zakeri, W. J. Welch, and D. J. Wolgemuth. Characterization and inducibility of hsp 70 proteins in the male mouse germ line. Journal of Cell Biology, 111(5):1785–??, November 1990. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/111/5/1785.

Zheng:1991:EAT

[ZZL^{+}91] L. M. Zheng, A. Zychlinsky, C. C. Liu, D. M. Ojcius, and J. D. Young. Extracellular ATP as a trigger for apoptosis or programmed cell death. Journal of Cell Biology, 112(2): 279–??, January 1991. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/112/2/279.

Zheng:1994:LRS

[ZZP94] J. Q. Zheng, Z. Zheng, and M. Poo. Long-range signaling in growing neurons after local elevation of cyclic AMPdependent activity. Journal of Cell Biology, 127(6):1693– ??, December 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/127/6/1693.

Zimmer:1994:TSE

[ZZR94] A. Zimmer, A. M. Zimmer, and K. Reynolds. Tissue specific expression of the retinoic acid receptor-beta 2: regulation by short open reading frames in the 5’-noncoding region. Journal of Cell Biology, 127(4):1111–??, November 1994. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/127/4/